A Complete Bibliography of the Journal of Statistical Physics: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

12 October 2024 Version 1.08

Title word cross-reference

(1 + 1) [FT93, Tan92]. (2 + 1) [BK96b]. (Δϕ)^{4} [PO93]. 0 [HS95b]. 1 [FLS95, Ken94, KE97, LPT96, NM99, Shi93a]. 1/(−j)^{2} [Mar92b]. 1/2 [GS93c, MKK97, Cho94, PBP97, Pok93a]. 1/D [Gar94, Gar96]. 1/f [KGM^{+}93]. 1/r^{2} [For93, FZ91]. 12 [Bat98, TLW91]. 16 [KS93c]. 2 [AC97c, BMO95, CMP95, Che92, Che93, Cho94, FLS95, HP91, HDS98, Iof94, Ser98, TF99]. 2/3 [dSCT91]. 3 [Bat98, BQ90a, CJK98, KM93, Ked93, MOT90, RB94, TLW91]. 4 [CQ90, SS97c, SS94c]. 5 × 5× [BS90b]. > [BLS91, CKS91, DW91, THK^{+}91, ZKB91a]. ^{4} [HS95b]. _{3} [STAJ95]. _{6} [STAJ95]. _{A}g [SO92a]. _{f−g} [SO92a]. _{p} [PPQ90]. A + B → 0 [Koz94, Koz96, KK93b]. A + B → B [ABP96]. A + B → C [QO95]. A + B → ⊘ [LC95b, LC97]. α [JW97]. because [MvR97]. β > 1 [Mar92b]. β > 8π [MK91]. β_{c} [OL97]. C∗ [GMR99]. c = −5 [ABT99]. C^{2} [AZP97, Hay96]. D [BB98, OP90, Sim96, UO91]. d = 1 [LS92]. d = 2 [SSV93]. d ≥ 3 [AZ95]. ϵ [PPW94]. F [Nag95]. G(r) [RST91]. g = 0 [RW91]. Γ = 2 [For90a, For92]. Γ = 4 [TF99]. Γ = 6 [TF99]. H [BCvB92, RRGT97]. J [BC92]. K [SS94c, Gui99]. k = 0 [BP94a]. L^{1} [AC90b, Pol90c]. l^{2} [SY90]. L_{2} [Ers94]. ≤ 2 [GM88, GM96]. m [CK95]. R [Naw98]. Z^{d} [CK99b]. N [BF97, RM97, Ber94a, CBR95, FE98, Kho91b, Mar97, OSB97, PRW99, PS99b, TF92, TF93a, TF93b, dVOS98, dME90, dMM98]. N = 1 [BMO96]. N → ∞ [CCRM95]. O(2) [Aiz94]. O(N) [BKJZJ93, CMPS97, PRW99, SZ99, Bat98]. Ω [CT93]. p [DJB98, Moo97, TF92, TF93b]. ϕ^{4} [JS95, RW96b, Wie98a]. Φ_{4}^{4} [Wie97]. ±J [Ben91, Ben92a, Gro95, JR94, Maz98, SDJ^{+}96, UO91]. Q [BVZ93, BJS99, Bis98, BSVZ94, BSV94, BJS98, CJK98, DLM^{+}90, Gay92, KLMR90, LMR93]. R^{1} [Sza93]. R^{2} [AZP97]. R^{d} [PZ99]. R^{n} [KP91b]. S [dOdOdSB95]. S = 1/2 [Ino90]. S^{2} [Wei95]. SM(2, 4v) [BMO96]. SO(3) [GLR98]. t [BC92]. T = 0 [Ben91, Leg98]. T_{c} [FOS94, Mon94]. U [Mie91]. U_{q}[sl(2)] [GS93c]. V [BC92, HJ90]. ξ [Whe99]. XXZ [GS93c]. xy [LLH92]. Z [Tót94]. Z(n) [Bha90]. Z^{2} [BG93b]. Z^{d} [Niw97]. Z_{N} [Yam96].

-Actions [CK99b]. -Body [FE98, PS99b, CBR95, Mar97]. -bonds [CK95]. -Completeness [Moo97]. -Component [PRW99, dVOS98, dME90]. -dimensional [BK96b, FT93, OP90, OP93, Sim96, Tan92, UO91]. -Dynamics [GMR99]. -Expansion [CT93]. -expansions [PPW94]. -Invariant [PRW99, CMPS97]. -Invariants [GLR98]. -Ising [BSVZ94, BSV94, BJS98, BVZ93, BJS99]. -model [Nag95]. -particle [RM97]. -point [MvR97]. -property [SS94c]. -R-B [SO92a]. -representability [HJ90]. -spin [TF92, TF93b]. -stable [JW97]. -State [CJK98, DJB98, SS97c, DLM^{+}90, Gay92, KM93, Ked93, KLMR90, LMR93, BVZ93]. -Step [Gui99]. -symmetric [Yam96]. -theorem [BCvB92]. -torus [Che92, Che93]. -Trajectory [Wie98a, RW96b]. -Vector [Whe99, dMM98, TF92, TF93a, TF93b]. -velocity [CQ90]. -vertex [KS93c].

0 [BLS91, CKS91, DW91, ZKB91a].

1 [KS90]. 1/d [NFID92]. 13-14 [Ano96j]. 1903-1976 [Lyo95]. 1968 [Ano95b]. 1976 [LHF95]. 1991 [Ano92q]. 1993 [Ano94a]. 1995 [Ano96j, Pod98b]. 1996 [Bry98, Wid98, Sac98]. 1997 [Git98, Opp98a, Opp98b, Rap98, Pod98a].

3 [McK95, Sch95a]. 3rd [Ano90o].

420 [Sre98]. 45th [Ano90o, Ano91q, Ano92q, Ano93o, Ano97g]. 4th [Ano91q].

5th [Ano93n].

62nd [Leb90]. 63rd [Ano90m]. 64th [Ano91o]. 65th [Leb91]. 66th [Leb92b].

67th [Leb92a]. 68th [Leb93a]. 69th [Leb93b]. 6th [Ano93o].

70th [Leb94b]. 71st [Leb94c]. 72nd [Leb95b]. 73rd [Leb95c]. 74th [Leb96]. 75th [Leb97b]. 76th [Leb97c]. 77th [Leb97d]. 78th [Ano98p]. 79th [Leb98].

80th [Ano99-28]. 81st [Ano99-29]. 86 [Ano01].

= [BB98].

abc [HW97b]. abc-gaskets [HW97b]. abelian [Spe93, SM99]. ABF [BPO96]. abilities [EIK92]. abnormal [KC91b]. above [Koz97, vEFK95]. Absence [BK94b, Con98a, HK96, Joy94, Ker93, KY93a, NK99, PS91a, SS97a, Aiz94, BT94, FjLL95, FLL95, GMTB95, MW94]. absolute [DGZ92b]. absolutely [AZP97, DZ94, Joy94]. absorbed [FMS97]. absorber [FGMA93]. Absorbing [MGD98, HT90b, Kło95, dMBD91]. Absorption [KPS98, BNRW93]. abstract [BM92a]. abuses [OO91]. ABV [YB91]. ac [Ger90]. ac-driven [Ger90]. acceleration [BAKK^{+}90, Mil91a]. accelerator [CJ92, Kar94]. accommodating [BDG97]. Accuracy [OTH92]. Accurate [BZW92, Hel98]. Acid [MSS98a]. Acoustic [AS95, FK94a, TKG93]. Acting [BHK98]. action [OdAdA96, Wei93a]. Actions [CK99b, PS97b]. activated [Gra90, Mel93]. Activation [EJ93a]. Active [Con98b]. activity [Pol94]. acyclic [AE90]. adaptive [Nad95]. Addition [Ano94a, TST91]. Additive [BAZ98, BV98, CLLL98, DIK98, GW93a, Num94, TDSR95]. ADE [FT94b]. ADE-O [FT94b]. Adiabatic [MP98b, AAR92, BM96a, CLY92, Sze96]. Adjointness [LPY98]. adjustable [GAA97]. Adsorbed [BCL^{+}99]. adsorbing [zzMZ96]. Adsorption [BKV93, CISS99, Guy91, Ram95, Rei98, dMdO98, BE94, EN92b, FP92, MD97, Per93, Ram93, SP94]. Adult [Jan94]. Advanced [Ano92a]. Advection [BGK98]. affine [HKV91, PW94]. after [CT98b]. Aggregation [BN99, BMPZ98, MP96, SLSA91, vD97, MP94, PN94, TTPH91, WMTR90]. aging [HH95a, Jan94, Pen95a, Ray94, Raa98]. Agreement [GLM95]. air [OAB^{+}96]. Airy [BW99]. Aker [McC95]. algebra [GS93c, Spe97, UI96, VZ92]. Algebraic [AC99, CCO99, EB95a, MVZ97, BT91, AC97d]. algebras [Fod97, FT94b]. Algorithm [BAZ98, CJK98, Hel98, HWD97, BTY91, CPS90, CPS91, CPS92, DDJ^{+}95, GAA97, Mer90, OvRW96, PBSR97, Rie93, SS96a, SS97b]. Algorithm-Independent [HWD97]. Algorithmic [LZ91, Shi93b]. Algorithms [BG97b, Bal92a, BG94, BGR94, BG96, FR90, Gon94, Gro95, HH93, Kaw96, Mar92d, SBH92, Vol94]. Alkanes [IRB^{+}99]. Allen [FL94]. Alloy [CHM99, KPS98, NFL99, Kra94]. Alloy-Type [KPS98]. Alloys [FPL99, JV99, Kik99, CNC94]. Almeida [CCST90, vE90]. Almost [BT91, DN94b, Ger99, Häg96, PV95, vES98]. Almost-sure [DN94b]. alternative [FF95]. Always [GJ99, Mie93]. Amino [MSS98a]. among [MC93, SP95a, Tor91]. amorphous [SL93b]. amphiphilic [TKD97].

Amplitude [Col98, Cra94, MS90a, LMS95]. amplitudes [Gon94]. analogy [BKL97a]. Analyses [Kaw98, Shi90b]. Analysis [AQ98, CBK99, DIK98, Häg98, Jar99, KT99, KK91b, MLM93, PPO99, SBZ98, SW99b, VB98, AGL91, BMA93, BPH^{+}94a, BBM92, BPH94b, BC96, Bov90, BP91a, BMP92, CCC^{+}90, CR94, Deu92, DH92, FS93, GP91, HZLD97, Hol94, Hua97, Kło95, LP96a, LP96b, LPT96, LPT97, tLZ96, Mar92d, MNO97, Mol95, Nay93, NR90, Paj97, PW97, TS94, TLW91, Wan96b, ZM93]. Analytic [HZLD97, KP91a, Leb99, Luo97, SER99, DF93, Pen95c, DN97]. Analytical [CLHS91, EJ93b, GY98, Jus98, ST99, ZHD95, BVHP92, ED96, MK93, BSTV94a, ESB98, FT90c, KPWH95]. Analyticity [AK92, MO96a, Uel99, CO97, Con96, FdlL92a, dlL92]. anchored [zzMZ96]. anchoring [ZM93]. Anderson [AK92, Bov90, EK96, Gra94, MPR98, MD94, NS95, Spe91]. Andrews [War96a, War96b]. angle [CDD94]. angles [CDM93]. angular [KK94b, Mar94b]. Anharmonic [FM99, HS95a, KLR94, VZ92, VZ95]. animal [ON96]. animals [HS90a, Mad95, Mer90, ML92, SF91]. Anisotropic [Fuj92, HW97b, Lüt95, MA98, NOZ99, ST95, SSZ99, Wan96b, BWK91, Fuj90b, GBN92, Hen94, KS97b, Kaw96, PS94a, PS97c, PCG95, RSL90, TC94]. Anisotropy [CHM99, Fuj90a, Fuj91, FSB91, Kho91b, MS94, Pri93]. Annealed [TF92, MKP91]. Annealing [ST97, Fri90]. Annihilating [CT98a, ETW98, Bel93a, Bel95, BL91a]. annihilation [BF95, FA91, LSKB91, Pri93, PCG95, RL91, Ric97, Asl99]. ANNNI [KS93a, STAJ95]. Announcement [Ano90a, Ano92r]. Annual [Ano90o, Ano91q, Ano93o]. Annular [Kin99]. Anomalies [Mol98, Spe91]. Anomalous [GS92, Sid99, WRJ95, bABD90, dSL90, AEA93, Dah96, FTGW96, GS91b, KK91c, Zha92]. anomaly [HS91, HS92b, PVZ94, Sta97b]. ansatz [BO90, LP90b, MF91, Sch93c, Yam96]. antibody [SP91]. antidissipative [Wol92]. Antiferroelectric [Alb98]. Antiferromagnet [SS98]. Antiferromagnetic [FS99, GH94, Ker93, KY93a, PS94a, PBP97, SS97a]. antiferromagnets [Mom94, PSP94]. Aperiodic [SBB98, BGP95a, Hof93, Hof95, Luc93]. Apollonian [MV91]. apparatus [GS90a, SHG91]. appearing [AFNB97]. Application [BPH^{+}94a, Ger99, HWD97, JS95, JL98, LFvW98, OR95a, PPR93, RVW96, YS97, BG94, BJO97, Bha90, ELMD^{+}90, KG95, LS92, MS91, Sco93, Tót90, ZP93]. Applications [Ano94a, Ano97f, BW99, HCW96, Jaf91, LK98, Opp98c, Pod98a, RM97, AGL91, Bob97, Dou95, Gal97, Git92b, Opp97, SS96b, Ste95a, WT93]. Applied [ZFB98, Bak94, KK92]. Approach [Col98, Jus98, PS99b, PS99a, SEW98, Tay99, BT91, BKL97a, BG93b, BCK97, CLD94, DN94a, DK96, DP92, Ebe96, FTGW96, Fog92, FF95, Guy91, HWvB97, Hon96, Jus92, KK94a, KG90, Kot95, LPT97, LB96, MOS90b, MD97, NBM90, OS91, Pal90, Pat93, PAB^{+}93, Pla90, Pol91b, RSGRP97, RR97a, SR95, SK90, Sch92, Shi90a, Spo96, TQGO95, TST91, TDSR95, TKD97, UO91, WKWS95, Wol92, YS97, dSCT91, dMPS95].

Approaches [SACB98]. approaching [Ray94]. Approximate [CJB99, Dah96]. Approximating [KRS96]. Approximation [KRT99, PRW99, SB97b, Wag95, AT90, BDG97, BCS93, BR91, BB95b, BHP96, BP93b, DDM90, HS92b, HS97, HM92b, IW93, KRT97, KvL92, LIF92, MO94, NS95, SB97a, VBF97, vVBE93, vWL95]. Approximations [CGT99, VB98, EJ93b, LS92, NP94]. arbitrarily [Nas91, vE96]. Arbitrary [Leg98, CGK94, Koz96, RW90, VBF97]. Archimedes [Leb97e]. Area [JW98, BCCF92, CC91b, CPP97, EG90]. area-preserving [BCCF92, CC91b]. Argument [AKR98, LM98, PSP94]. arising [Ser96]. arithmetic [FMR94]. arm [OB91a]. Arnold [HSK91, CV93a]. arrangement [Fra94]. array [DDG97, DDG96, LS97, MS90a, WCT91]. Arrays [Jos98]. arrow [Leb97e, Spo92]. Arrows [Opp98b]. Art [Pod98b]. artificial [KK93a, SBH92]. Artuso [Pol91b]. Ashkin [BC98a, BLL94, PV97, SS96a]. Aspects [Dou95, CFJ91, KWG96, LVY92, Par91, RRGT97, ZKB91b, ZKB91a]. assessment [NGB95]. associated [CCRM95, CMVG95, Ole90, Pey91, SY90, WPK95, ZXZY94, dlL97]. associating [KR92, VDH97]. association [HWvB97, OS91]. associative [Fod97, Shu93]. Asymmetric [Alb98, AKK99, BFSV91, DGLS98, EFGM95, RSSS98, SER99, SK99, Spe93, ACJL92, APC^{+}92, BKL97b, DDM92, DJLS93, DEM95, DLS97, FFK94, For90b, FG94, HL97a, JL94, JW93, KO93, Nol92, Oze93, RS91b, Sch94b, Sch93c, Sch97a, Sch97b, Swe97]. asymmetry [ALLZ96, WMS90]. Asymptotic [AE99, Bel93a, Bel95, BL91a, BOV98b, CdC94, CDM93, Ers92, Gar98, Gol99, Joy90, Kot95, LR96, Mer99, Ole90, Shi93a, SCM96, Sut92, Tab96, TV90, BBOC91, Don95, DN94b, Fuj90b, GK95, MLL90, MO96b, Pen95c, PM91, Sin91, BO99b]. asymptotically [HW97b]. Asymptotics [CM98, KPS98, MZ96, Gou97, IM96]. Atomic [FG99, Git98, LMM92, SGP90]. Atoms [WC98, LP90a]. attached [KP94a]. attack [ML92]. attraction [DKKP96, SLSA91]. Attractive [BI99, BEO98, Vuj99, APC^{+}92, BFSV91, Mar97]. attractor [HC92]. attractor-basin [HC92]. attractors [CV96, Cut91a, Cut91b, Por90]. auditory [TKG93]. auditory-nerve [TKG93]. Aurell [Pol91b]. Australian [Ano91a, Ano92b, Ano93a]. Author [Ano98a, Ano99a]. Autocorrelation [ZP99, BvV95, MM97b]. Autocorrelations [JS95]. Autocovariances [CLLL98]. autoimmune [CS90b]. Automata [AAH98, Ano92a, Ano99k, MM98, Moo97, ABF^{+}95, Ahm96, BKM93, BSG95, BBM92, BG93b, BEK91, BKW90, BT92, BT94, BED95, CDM91, CMMC92, CD91a, CFJ91, CW95, DB90a, Elo94, ED90b, ED92a, EB95a, Fri94, KH96, Koh91a, Koh92b, Koh91c, KC91a, KS90, LMS90, Li92, Mey96, Niw97, PS90, Pom93a, RW96a, Rus94, Sch90a, Ste95a, SRC93, TB92, Voo92, WC95, vBE93, ED90a, Rap98]. automata-based [KH96]. automatic [DP91]. Automaton [ERS99, AEGL92, DEP92, EN92a, HC92, HP91, Kar95, KA94, Nas91, Num94, RK93, ZP94, dSM92, vRE93]. automorphisms [Che92, Che93]. Auxiliary [Fuj98]. Avalanche [BP96, HW97a, Man90]. avalanches [GPSS93]. average [FJ92]. averagers [Bre91]. averaging [AM94b, PS91a]. Avoiding [Ano99h, BBG98, Noo98, RRT98, BF96, Bar90, BKV93, CPS90, CPS91, CPS92, GB90, HSS93, HSS95, HA97, Ken94, LM91, LS90a, LMS95, MOS90a, NDF92, O’B90, Pen94a, RR93, RB90, SA94, SA95, SSV93, TvROW96, Tót94, Wu95, vR97a]. Axial [Sta98]. Axiom [Hay93, Liu98, Wei99, Por90, Sim94a, Sim94b]. Axisymmetric [BBW98].

B [CKS91, DW91, SO92a, ZKB91a, BLS91, THK^{+}91, AAH98, SP91]. B- [CKS91, DW91, ZKB91a, BLS91, THK^{+}91]. B-cell [SP91]. B. [Dor94]. back [BK91]. back-jump [BK91]. Backbends [RSW98]. backbone [AE90]. backgammon [FR96]. Background [Mül99, FJ96, PPQ90]. backscattering [vBE93]. bacteria [vK95b]. Baker [CK91]. balance [Che95, EB95a, Fri90, GS97, Jes96, LC95a]. Ballistic [BF95, Ken94, NR98, MP94, Ric97, TH96]. Band [FK94c, KPS98, Klo98, GPS90, Niu91, Ole90]. Band-gap [FK94c]. bands [Tas96]. barrier [ET90]. barriers [Dro96]. Barton [Jan94]. based [Ald93, KH96, KF97, Nad95, WL92, WLC94]. Basic [Liu98, Sta91, BM92a, GW94a]. Basin [DKKP96, HC92]. Basins [AC91, CGM^{+}98]. Bath [FP98, SW99b, MHdA90]. Baxter [War96a, War96b, CCRM95, HH95b, HH96, SMS96]. Bazhanov [CCRM95, HH95b, HH96, SMS96]. BBGKY [KOT98]. bcc [KL91]. BCS [PR94b]. be [BMHH97]. beam [PP95]. bearings [MV91]. Becker [Pen97, Vel98]. Becker-Döring [Pen97, Vel98]. Behavior [AG99, AKR98, AKMR98, BJ99, BO99b, BHW99, CG99a, CJK98, Gar98, KE97, Klo98, KOJ98, KS97c, LK98, NP99b, SSLI97, SI99, dVOS98, ACM95, Alb92, AC90a, Bel95, BKL97b, BOP94, BL91a, CPP94, CdC94, CM95, CP97, Cut91a, Cut91b, Deu92, DR93, Don95, GW93a, GGL90, HMV90, Joy90, Jur95, Ken94, KS93c, Koz96, LM96, Luc93, MZ90, Mas92, MO96b, MNO97, MC92b, MS94, NHT92, OYSK91, Ole90, OTT92, OPB93, PS92a, Pei95, Pom93a, RR93, Rus93, SS96a, SS97b, Shi93a, Sin91, Spo92, SSV93, SRC93, Sut92, WHF92, Wu95, YK91, dOP93, vDL95, vEFK95]. belong [dSCT91]. below [FOS94]. Bénard [BTT98]. Berkooz [Sre98]. Bernoulli [BRT98, LP94, LP96a, LP96b, Por98]. Besov [Eyi95a]. Bessel [BT93a]. Bethe [GM96, AK92, ABPSJ91a, BRZ95, BRZ98, BO90, CCC^{+}90, CCST90, DKMM94, GM88, HS90b, KM93, LP90b, MF91, MKP91, RV91, Sch93c, Yam96, ZP93]. Between [BMSW99, FR99, FMPP99, HLW99, Wid99, BD97, BR91, BLS91, CPPG91, DFZ94, DI93, FJ92, HT90a, Hof96, HY98, HTPH93, Ino90, JW93, KB97b, Mat90, Ném91, PO93, Per95b, Sni95]. beyond [AMF98, Gal95, vD97, vVBE93]. BFACF [CPS91]. BGK [BP93b, Bou99, COA95, HZLD97, QO95, TQGO95, ZHD95]. Bias [RLK98, CCC^{+}90, CCST90]. Biased [BCL^{+}99, BE92, JS99]. Bibliography [Ano92c]. biennial [WRJ95]. Bifractality [AFNB97]. bifurcation [OPdlR95b, VB93, CCC^{+}90]. Bifurcations [DIK98, Göt96, JL98, Jus95, Kus99, Tat98, ACLS94, CCT92, Hal97, JR91b, MS96a, dSVG90]. bilayers [NE95]. Bilinear [ZK93]. Billiard [Gar97, GM97a, KMKdC96, OdAdA96, Sze96, Tab96]. Billiards [Ano96k, Che99b, GG94a, Gut96, HS98, Ami96, ACG96, BDG97, BK96a, Che97b, Dag96, Hay96, Ish96, SS94c, SHW92, TY96]. Bimolecular [LSKB91, KK93b]. Binary [CHM99, CFP91, Häg98, JV99, KK94a, BJO97, CNC94, Com91b, FT94a, GRZ90, KGM92, KK91c, Kra94, MGS94, Num94]. Binding [DFF99, BB95b, BHP96, MO94]. Binomial [BGL99]. Biological [AEA97, Raa98, Tat98, CA93, Jan94, Pen95a]. biologically [Gro95]. biomembranes [Nag95]. Bipartite [GR98]. Bipolaronic [GMR98, GMR99, AAR92, BM96a]. Bird [Wag92]. birth [GKRT94]. Bistability [ZK98]. Bistable [IdRB98, dRIB99, Fer96, FR97, GDJH93]. Bit [Raa98, Pen95a]. Bit-String [Raa98, Pen95a]. blend [FT94a]. Block [Bak98, FM91, GMH98, BMO95, NHT92, O’C93a, OL97, RK93, vE96]. Block-Spin [Bak98, OL97, vE96]. blockage [JL94, Sch93c]. Blow [Naw98]. Blow-up [Naw98]. Blume [BFB94, CO96a, dOdOdSB95]. Blumel [Git98]. Body [FE98, Leg98, PS99b, RS97a, BMS91, Buo90, CBR95, For93, For94, GSH90, Gri94, GKT93, KT94, LP91, Mar97, MSZZ90, VB90b]. Bogoliubov [Sac98]. Bohmian [DGZ92a]. Boltzmann [ADG96, AL95b, APT94, AC90b, ACI91, AM94d, BB91, BP94a, BF91, BBM96, Bob95, Bob97, BC99c, BED95, CC92, CC94a, Cer90, Cer96, CWSD92, CQ90, CP93, Cor95, CM92, Dal97, DE96, DEJ92, ELM95, EML98, GTW95, Gal95, GAA97, GS91a, Gd96, Gou97, HZLD97, HL97b, HR92a, HT90b, IW93, IS96, Lu99, Luo97, MGA95, MGA97, Mor92a, NS92, NGB95, PS98, Pal90, Pet90, Pet93, RRGT97, RBF93, SS92b, SD95, Sni90, Sni95, Sta92, SAB95, TV99, Wag92, Wag95, WH95, Wen97, Wen99, WG95, Zie93, ZHD95, ZHCD95, vVBE93, vdSE99]. bond [BS90c, CR94, Dua90, DB90b, KIKK97, MKP91, Sch92, Tót94, YHHK96, YS97]. bond-diluted [DB90b]. bond-moving [YS97]. bonded [NLT93]. bonding [VDH97]. bonds [CK95, GS95]. bone [Fra94]. Book [Adl93a, Adl93b, AW90a, AEWF91, Ano95a, Ano99c, Ano99d, Ano99m, Ano99g, Ano99i, Ano99h, Ano99f, Ano99b, Ano99j, Ano99e, Ano99k, Ano99l, BSB97, Ben92b, Ben93, Ber90a, Ber90b, Bry98, Dom92, Dom97, Dor93, Eng91, Fam92, GM92, Gan91, Git90b, Git91b, Git91c, Git92b, Git92c, Git92a, Git92d, GS93b, Git93a, Git93b, Git98, Leb92a, Mas92, Muk91, Nos90, Nos93, OG91, Opp91, Opp98a, Opp98b, Opp98c, Pod98b, Pod98a, Rap92b, Rap98, Ros93, Sac98, Shl90, Shl91, Spo92, Sre98, Wei91a, Wei91b, Wei93a, Wei93b, Wes91, Wid98, dB92, vEdH91, vdBV93]. Books [Ano99m]. Bootstrap [GM97b, Bra93, KS93b, Sch90a, vEAD90, vEAD91]. border [BJ90]. borderline [KB97b]. Borisovna [Hei94]. Bose [AB92, GO95, Koj97, MV99, Pen91a, Sch90b, Sto97a, Tót90, Tót91]. boson [HS94b, PY94, Sch94c, vdBDP92, MV99, OPS93]. bosons [Mar97, SS99]. bottles [Cla91a]. Bound [Gol99, MKK97, Bel93a, dAB91b, Con90, Con96, FMP92, GY93, Iof94, Mad95, Red94, Tót91, Weh97a]. boundaries [BCCF92, DDM92, DEM95, JT96, zzMZ96]. Boundary [CL98, DHW99, DHS98a, DP97, GS98, Hei98, IS96, IdRB98, Koj97, NOV99, Pat98, SSV93, Zie93, ABHP90, AM94d, Bax93a, BPO96, BKJZJ93, Cer96, CL97, FLS95, For90a, For91, Fuj96, Gd96, GH94, Hay96, HZLD97, HT90b, Kło95, LP90b, MF91, NGB95, Pat94, Pet90, RS92c, Smi94, SRC93, UO91, WT92, WT93]. Boundary-Driven [DP97, CL97]. Bounded [GSM98, Too94b, LhBBS97, LO97]. Bounds [CMP97, MS96b, Noo98, Reb98, BG92, BFB94, BG90b, CC92, ED92b, HSS93, HSS95, Joy96, KK91a, Kie92, Mar97, Mon94, PPS95]. Boussinesq [EML98]. bowed [KF90]. Bowen [Sim94a]. box [KZ93, Ras93]. braided [DN97]. Brain [MS93c, CV96]. branch [DDJ^{+}95]. branch-and-cut [DDJ^{+}95]. Branched [GC90, Vuj99, HHD96, HS92a]. Branching [CT98a, Mac97, SK98, dHMP99, GPSS93, KM93, MM93, Par93, Sch93a, SO92a, SO92b, Wu95]. break [dME90]. break-collapse [dME90]. Breakdown [Ano99l, MS97]. Breaking [AHR99, BvEN99, EFGM95, KT94, LP91, MW94, MG92]. Bridge [FMPP99, MD97]. Bringing [Joh90]. Brittle [BRT98]. broken [Koh90, SL95, Yam96, Zim93, Zim94]. Brownian [AFNB97, Ber94a, BPH94b, BHP94, BHP97, BP97, Bre91, BL99, Che91, CF97, CDM93, CO96b, DJ93, Dow91b, FMS97, Ger90, Got90, IMS92, KK92, Lan95, McC95, Mol95, Mol98, Pet99, Rei93, SS94a, SS94b, Sch95a, SL93a, Str97, Wil91, dMBD91]. Bruria [Kau95]. bubbles [Rot93]. Bulk [SY95, OB91a, Zha96b]. Burgers [AFNB97, Ber98, BB92b, Der97, JW97, LOP96b, LOP96a, Mol97, Sin91, WX97]. Búttiker [BL90, SH89].

C [Dor94, Pod98b, THK^{+}91]. c=c [PPQ90]. Cahn [BLO97, BMSW99, FL94, IRB^{+}99, Whe99]. Cahn-Type [IRB^{+}99]. cakes [Kru97]. Calculating [BS97, REK91, Smi94, BGR94]. Calculation [Fuj90b, Kei98, Koh91a, FdlL92a, KP91a, Koo95, LLH92, MS90b]. calculations [Bak94, BT93a, WB92]. Calculus [JL98, EKLR94]. Calogero [Cho97]. calottes [GG94b]. CAM [KS93c, LS92]. Cambridge [Bry98, Git98, Opp98b, Pod98b, Rap98, Sac98, Sre98]. Can [Ber97a, HJZM93]. Canonical [CE94, GMR98, ABJM97, KB91, Ném91, YP95]. Cantor [ABJM97, DM90, Orz96, TV90]. Capacitively [Jos98]. capacity [BG92, FMP92, LW93]. Capel [CO96a, dOdOdSB95]. capillaries [KK91c]. Capillary [BMSW99, RW91, RV97, Wee91]. capping [CWP97]. Capture [LVY92]. cardioids [Ste90]. Cardy [Bry98]. caricature [DPS90, FdH94, ST90, ST91]. Carleman [CDPP90]. Carlo [Ano99e, Alb92, ALLZ96, Bak93, BZW92, BM96b, BHJ92, CPS90, CC94b, Deu92, DJB98, DT93, EMHM95, FS99, FLB91, FS93, GAA^{+}93, GMM90, GP93, GB90, HMP96, Heu93, HCW96, JS95, Jes96, Kal91, KS97a, Kom93, KKBS92, KKBS93, KK94b, LS90a, MOS90a, MvR97, MC94, Mar92d, MHB90, MW95, MKZ92, NP99a, OB91a, OvRW96, PdO90, PBSR97, Rie93, RDWW93, ST97, SW99a, SBH92, Sel97, TBK90a, TBK90b, TvROW96, TFD90, Wag92, WB92, YB91, ZK93, dSCT91]. Carpet [Erc97, Sim94c].

Cars [IS99]. Cascades [MKK90, BBF94, KP91b, SS95]. Case [PZ99, Che94, HS94a, Kło95, OPdlR95b, OS95, OS96a, SMD92]. cases [YT90]. Casimir [Pod98a, Sha95]. catalysis [BAK91]. catalytic [GS90c, MS92]. cavities [LhBBS97]. Cavity [PM99, HS95c]. Cayley [BM97, FNW92, dABMR90, dABR91, vdBDP92]. CECAM [Ano94a]. Cell [JVH98, TNN99, CWP97, HA97, KS93b, KK94b, Muc96, SP91]. Cellular [AAH98, Ano99k, BG93b, CD91a, ERS99, Kar95, MM98, Moo97, Rap98, ABF^{+}95, Ahm96, AEGL92, BKM93, BSG95, BBM92, BEK91, BKW90, BT92, BT94, CFJ91, CW95, DEP92, DB90a, EN92a, Elo94, ED90a, ED90b, ED92a, Fri94, HC92, HP91, KH96, Koh91a, Koh92b, Koh91c, KC91a, KS90, KA94, LMS90, Li92, Mey96, Nas91, Niw97, Num94, PS90, Pom93a, RW96a, RK93, Rus94, Sch90a, Ste95a, SRC93, TB92, Voo92, WC95, ZP94]. cellular-automaton [Nas91]. center [APC^{+}92]. centers [Ste90]. central [BFG93, vdH98]. Certain [CK98, MZMQ90]. Cesàro [MM98]. Chain [MKK97, Mül99, WBG98, AC90a, APC^{+}92, BKV93, CdOW95, CB90, CR94, EJ93b, FZ91, GS93c, Guy91, HH90, Ino90, KM93, Ked93, Kna93, LLM95, Luc93, OMM93, RS91b, dSL90]. Chains [BG97a, BEO98, KS97c, LW98a, ACM95, BP94b, BS91a, DN97, GB94, HT90a, HS95a, KLR94, KM96, MF91, OS95, OS96a, RCB90, Sch93a, Sco93, zzMZ96]. Challenges [SAB95]. chance [Sas95b]. changing [PPQ90]. Channel [LY97, Lie93, PAB^{+}93, WK97]. channels [DI93]. Chaos [Adl93a, Cop98, Erc97, Git92a, NS99, RBGW92, AS91b, Ano92p, BRR96, BM92b, BSVZ94, BSV94, DN94a, DEJ95, DGZ92a, Ers92, Git91c, Git92c, Git92d, Git93b, GKW91, Hal97, KB97b, LR92, OTH92, Paj95, Sas95a, Sze96, Git98]. Chaotic [AAR92, BBF94, DI93, DO97, Gal96, Gal97, Git90b, ILD93, Jez96, KOJ98, Leb99, LAT95, Mas92, NHT92, NV98, ZR91, AZP97, ANS93, BNN95, BNN97, BBL96, Bry94, CV96, CLS90, EIK92, Gas92, GBP91, Man93, NNM93, Shi90a, Shi92, SHW92, Zha96a, Git91b]. Chapman [CH94, Sle98]. characteristic [Gra94, MW91, Oku90]. Characteristics [CS91c, TKG93]. Characterization [JG98, Tat98, KY93b, LPT96, PY94]. characters [DKMM94]. charge [FRHP95, JLM93, KK93b, Lüt95, Rus93, WB92]. Charged [BE99, Dal97]. check [EKLR94]. Chemical [RS92b, ZK98, BB97, OS91, OO91, PTN93, Wei91b]. chemistry [Joh90, Opp97, Wei93b, Ano95b]. Chemists [Opp98a]. Chernov [DP97]. chessboard [Ord92]. Chiral [Bax93a, Bax98, BS90a, AYP95, Bax91, Bax93b, Bax96, Dav91, Hon96, MO96a, Mül93, OB96, WG96, WPK95]. Chromosome [DJM99]. Chronotopic [LPT96, LPT97]. Circle [CM98, FT90b, FT90c]. Clarendon [Pod98a]. Clarkson [Ano90o]. Class [AG99, Häg98, MM98, BKK^{+}92, BAP93, Dai90, DS92, FIS96, Gra95, HW90, Hay96, Mon91b, PS97a, PS91b, ST93, Xu95, dSCT91, vD97]. classes [ORG91]. Classical [Ano96k, AC99, BB91, BF91, BHS99, BO99a, BO99b, DC94, HLW99, JM92, Jan95, Kar99, LW93, LVY92, Mie98, Mie99, Nob95, Per95a, dVOS98, AC97c, APT94, BK96a, BCF97, BJL^{+}91, BQ90b, CW96, CS91b, CCG90, DFF96, DP92, DGZ92a, FK97, FJT96, GI92, Gar94, Gar96, Geo95, Git93b, GGP92, Ish96, JLM93, Kar94, KR92, Kie90, KB97a, KY93b, KMKdC96, LM94a, LSS97, Mac95b, Mat90, Nos93, NP94, PY95, Per97, RS92a, SO91, Sch93b, Sni95, Sut92, Ano99b]. Classification [Sai95, FT94b, FST97]. classroom [Adl93b]. Clausius [Jar99]. Clerk [Dom92, Dom97]. cliff [RK96]. climatic [Nic93]. clock [BDDM90, Uen95]. close [BG93a, CG93, SS94a, SY95]. Closed [MSD92, Git90a, Lie93, Muk91]. Closure [EA98, GY98, BVHP92, Lev96, SW91]. clouds [PN94]. clumping [Hor93]. Cluster [Bis98, BNO98, CGTM99, Häg98, Kaw96, Koh91c, Mor90, OP90, Rap92a, RS97a, RNCT98, SC99, TW98, Vol94, ZFB98, BC96, CK95, CT95, Dim90, Gri94, Gri95, Häg96, HS92b, HP91, HWvB97, HH93, Iof94, Joy90, Man90, Mar92d, MHB90, PV95, PV97, Phi91, PB95, Pri92, RM93, RW92, RDWW93, TBK90a, TBK90b, WK97, YZ92]. cluster-memory [RM93]. clustering [BE94, PY95, YB91]. Clusters [JSA98, BK92a, BTY91, Cam91, Eve93, GS95, HKV91, HA97, PO95, Sch93b]. CN [STAJ95]. coagulation [BC90a, CdC94, HRS97, Hor93, KP94b, PCG95, Sle96]. coagulation-diffusion [HRS97]. coagulation-fragmentation [BC90a, CdC94, Sle96]. coalescence [TH96]. Coalescing [ETW98]. Coarsening [Vel98, Pen97]. code [PdO90]. codimension [DMB97]. codimension-one [DMB97]. Coding [BZ90, PS99a]. Coefficient [KPS98, BPH94b, BHP94, RC97]. coefficients [GGL90, HL97a, Pol90b, Pol91a]. coexistence [CGMS96, GLM95, HS90b, LMR93, LLM95, MM96, Mar94c]. Coexisting [BMSW99]. Cohen [LS99]. Cohen-Type [LS99]. coherence [Niu91]. Coherent [FPL99, FG99, GDJH93, Sre98, Sta97b, BLL90, HS91, HS92b, NS95]. coherent-anomaly [HS91, HS92b]. coherent-potential [NS95]. cohesion [WLC94]. collagen [Fra94]. Collapse [BW98b, ON96, OP95, dME90]. collapsed [Fos93]. Collapsing [BNO98, MvR97]. Collective [FP90, vRE93, BBM92, BNS92, SP93, YP95]. colliding [GIT91a]. collimated [PP95]. Collision [BI99, CC94a, Cer90, GMO91]. Collisional [BHS99, CK99a]. collisions [Gou97, MC93]. colloidal [CdS91a, CdS91b, SS94a, SS94b, Sch95a, SLSA91, SHG91, VdSFC97, Wil91]. Colloids [BN99]. color [GS91a, RHA97]. Colored [BAZ98, BZW92, CCT93, HJZM93, Kli91, MT94, NBM90, Rei96a, WB94b, Zie91]. Columnar [DC98]. comb [AC95]. combs [DJ93]. Combustion [KPANG98, MS96b]. Cometary [DLPS99]. commensurate [Gry92, PS94a]. commensurate-incommensurate [Gry92]. Comment [CCT92, GMCP96, Mon97, Raa98, Wee91, GW94b]. Comments [Mar90a, RRGT97]. Compact [JK99, PO95]. Comparative [JSO99, MHdA90]. compared [WC95]. Comparison [BS91a, FR99, Gri95, Kar99, MS90b, RSSS98, RS91a, Gro95, KGM92, SZ91, BvV95]. Competing [GS98, CdOW95, Gry92, SC96]. Competition [KNS98, Spo95]. complementary [Kie90]. Complete [HMY96, Jeo99, CO97].

Completeness [Moo97, CS92]. complex [AGL91, BRR96, BCF97, CDM91, CD90, Edw91, GL93, dOP93]. complexified [Lüt95]. Complexity [Ano99m, Cop98, LMN98, MN99, Fog92, Mac93, MG94, MG96, MM97a, Shi93b, vROS^{+}96]. complicated [HG92]. Component [BBD99, CCF99, Has98, PRW99, dVOS98, AF95a, Alb95, For90a, FJ92, For92, FJ96, HS90b, JMM87, JLM95, MC94, RW90, SP95a, Sut92, dME90]. components [RR97b, THK^{+}91]. Composites [LK98, LB96]. compounds [STAJ95]. compressible [Xu95]. compression [MSG95]. Computation [KA94, PPW94, BZ90, For93]. Computational [LZ98, MN99, SW99b, CWSD92, Mac93, MG96, MM97a]. computationally [Jer90]. Computations [Ano90n, CMMC92]. compute [Bha90]. Computer [MSS98b, NFL99, SP94, BP94b, PB94]. Computers [Cop98]. Computing [BLPP98, BK92b, Din96, Ste99a, DZ94, GK95, Mar90b, SAB95]. concavity [LS91b]. concentrated [Dow91a, FSB91, GGD91, KFK91, Wil91, vdBJ91]. concentration [AMA^{+}90, HS95b, PPQ90]. Concepts [Cop98, Fam96]. Conceptual [HKV91]. concerning [Sin91, dO95]. condensate [Sto97a, Tót91]. Condensation [AHR99, AB92, Pen91a, WT93]. Condensed [Ano99b, KK93b, OO91, Pod96, Sel97, Shl90]. condensing [WT92]. Condition [BW98a, Nag98, Sch98, Che95, For91, Kło95, Koz94, Yos97]. Conditional [CK98]. Conditions [BCO99, CL98, Koj97, NOV99, Pat98, Bax93a, BPO96, BKJZJ93, Cer96, For90a, Fuj96, GH94, HZLD97, LP90b, MF91, NGB95, Pat94, Pet90, RS92c, Smi94, SRC93, Zie93]. Conductance [AM94b, NR98, Weh97a]. conducting [Fig92b]. conduction [Sai96]. Conductivity [Fel98, BEC93, BK95a, BG90b, FTGW96, GS91a, Lan95, LB96, MS94, RC96, SS92a]. conductor [For90b, JT96]. conductor-insulator [For90b]. Conference [Ano90p, Ano95c]. configuration [OL97]. Configurational [DMB97]. configurations [GMMU97]. confined [HSW97, HQSS94, JM92, LhBBS97, LBT97]. conformai [LS90a]. conformai-invariance [LS90a]. Conformal [AC90a, BT90a, PW90, Wei99, BM96b, GB94, PW97, WB92]. conformational [Kho91a]. conformations [NMC^{+}91]. conical [BS95]. Conjecture [Jeo99, BB91, BF91, Böt95, CC91a, MN96]. conjectures [dH94]. Conjugate [BCP98]. Connected [BJS98, PS94b, ZP93]. Connection [Hén92, dMM98, AKV94, Ino90, Pen97, BS90b]. Connective [Noo98, HSS93, HSS95]. connectivities [FAKA97]. connectivity [Uen95]. consensus [Che97a]. Conservation [Bou99, CCO99, KT99, KD98, Lu99, MBD97, NOV99, BC90a, KD90, Yau94]. Conservative [CDG95, Git93b, Mae90]. Conserved [MWA95]. conserving [Koh91c]. considerations [OD96, Sta91]. Considering [Wag98]. consistency [GGM91]. Consistent [FE98, KRT99, VB98, DK96, EKLR94, GAA97, KRT97]. Constant [JT98, Mou99, DEM95, Ge91, HSS93, HSS95, KP94b, Wan96a]. Constants [Noo98, vdH98, Koz96, MS90b]. Constitutive [KW93]. Constrained [BCP96, SS96b, CO97, EJ93b, SSP95, dLPP90]. constraints [Edi93, MS91]. Construction [Bou99, KM93, KOT98, Wie98a, DKS93, RGdG97]. constructions [PW97]. constructive [JLS96]. Contact [BW98a, Blu94, CDD94, GW99, Sal99, Sch98, JK93, KK91a, NDF92, Sch94b, Sch96a, Sim96, Swe97, Wu97b]. Contacts [BBG98, BF96]. containing [HS90b]. Contemporary [Opp95]. continuation [MK93, Pen95c]. continued [Bal92a, Bal92b]. continued-fraction [Bal92a]. Continuity [Wu97a, MRS94]. Continuous [BO99a, BO99b, Guy91, Kik99, Pok93a, RS97a, RRT98, AZP97, CCRM95, Con90, DZ94, GK96, Joy94, Kot95, MW94, Mom96, OMM93, Pra94, Pri93, Wan96a, WH91]. continuous-space [OMM93]. Continuous-Time [RRT98, Kot95, Pra94, Pri93, WH91]. Continuously [Gry92]. Continuum [BK95a, CMPS97, PS97b, VDH97, dlSG99, Bar90, CT95, GS90b, Gra94, KH94, LVY92, MRS94, O’C93a, RV91]. contracting [Ers94]. Contrasting [LSP91]. contribute [Ber97a]. Contributions [Ano91b, Ano91n, Ano98d, Ano98e, Ano98h, Ano98k, Ano98n, Ano99c, Ano99q, Ano99t, Ano99u, Ano99x, Ano99-27, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano90i, Ano90j, Ano90k, Ano90l, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano91i, Ano91j, Ano91k, Ano91l, Ano91m, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano92j, Ano92k, Ano92l, Ano92m, Ano92n, Ano92o, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano93i, Ano93j, Ano93k, Ano93l, Ano93m, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e]. contributions [Ano98b, Ano98c, Ano98f, Ano98g, Ano98i, Ano98j, Ano98l, Ano98m, Ano99o, Ano99p, Ano99r, Ano99s, Ano99v, Ano99w, Ano99y, Ano99z, WKWS95]. controlled [BO91b, CSB97, Wei91b]. controversial [RW91]. Convection [BSdB^{+}98, CD99, FP97, Str94, JR91b, LFtH91, OYSK91]. Convection-enhanced [FP97]. convective [BWK91, CPPG91, OR95b]. Convergence [ACC^{+}98, Aur90, AE99, Bal92a, BCO99, Kei98, LS92, MP98a, PS98, Phi91, Pok93b, RS97a, AC90b, BKM93, CC92, CS91b, LOP96b, LOP96a, MOS91a, Pet93, Wag92]. convergent [MOR97]. Conversion [Sni95, Dua90]. convex [BOP94, Din96, Hay96, HS94a, Sto97b]. Convexity [LS91b, MMSR92]. Convolutions [Por98, LP94, LP96a, LP96b]. Cooperative [PS93c, BE94, EN92b, Pol90b, Pol91a, Rap90]. Coordinates [MV99]. copolymer [TKD97]. Core [BHW99, AB92, ABP96, Sch95b, Tót91, YS93]. cores [BB94]. Coriolis [NAC91, WHS^{+}91]. Corner [AL95a, Bax91, Bax93b, KLT97]. correct [MGA97]. corrected [Sni90]. Correction [McK99, DPS90, SA91]. Corrections [CMP95, DR91, SS97c, BD93, For91, GD93, LF90]. Correlated [JS99, TC93, AK91, BC92, BE92, CM95, DD91, Hav90, KS97a, Koz94, MW90, PS92a, SSV93, TC91, WK97, YHHK96]. Correlation [CLV97, CGTM99, HS95c, Koj97, MS98b, Maz98, PPO99, PIM94, PG99, PG00, Sac98, SO91, Sim98, TW98, BKL97a, BT90b, BR91, CBR95, DH92, ED90a, ED90b, FN95, Fuj90a, Fuj91, GG94a, GH97, HS94a, Hon96, Jez96, Kie92, KNV93, Li90, LBK96, MM90b, MC92a, MW95, MC92b, O’C93b, OTT92, PO93, Pes93, PS91c, RST91, Sen92, Ser96, Sut92, YHHK96, ZPK97, vVBE93, CBR95, CLV98]. Correlations [AC92, AF95a, AC97c, AP99, BD97, Che99b, FP98, FN97, GSM98, GLR98, NM99, SZ98, dVOS98, Aiz94, BK91, BE94, BO91b, Com91b, DBH91, EB95a, For93, JMM87, Jan95, JLM95, JM96, KB97a, KM96, Krá96, LO97, Liv95, MR94a, Mae90, Mil91b, MZ96, PHS^{+}97, RW90, SP95a, SBH92, TF93a, vB90a]. Couette [CC94b, GC92, NAC91, WHS^{+}91]. Coulomb [Mar90a, AC92, AC97c, APT94, AC99, BR92, BMC94, dAB91b, BK94b, BM99, Cho97, DHS98a, FjLL95, FLL95, FM91, For91, FJM92, FJT96, JM92, JLM93, JMP94, Jan95, JT96, JT98, Kie90, Kie92, MKP90, MK91, Mun92, PPNM97, Vie94, Vie95]. Coulomb-type [Kie92]. Coulombic [Fis94, WKWS95]. Coulombland [AC97b]. Counter [Mon04]. Counter-Example [Mon04]. Counterexample [Lef99b, Bra91, LPS94]. Counting [ML92]. Coupled [Jos98, Jus98, LC99, ZFB98, BBC^{+}95, BNS92, BVR93, BLL90, CP93, Cra94, Dol97, Fer96, FR97, HLIM93, Jus95, Kar95, MC92a, NV93, RC97, SM91, WT93, vHW93]. Coupled-Cluster [ZFB98]. Coupling [BG97b, Wie97, MPdlR93, RW96b, Ste97]. Couplings [BOV98b, AE91, BG93d, HT90a, WA90]. Courbage [SAT94]. covariance [BC96]. Coverage [CT98b, dMdO98, BP94b]. covered [MRS94]. covering [Hen97, Kou90]. coverings [KRS96, KP91b]. Crack [Mar98]. created [FK97]. Creating [Bru94]. crisis [Rei96a, Rei96b]. Criteria [HR95, Yos97]. criterion [FdlL92b]. Critical [AC97a, ABHP90, BW98a, BJ99, BAKK^{+}90, CMP97, FKV91, GH94, GB90, HH93, Heu93, HS92c, HY98, JSA98, KS97c, LF93, LMS95, Luc93, MOR97, PS92a, Pin94, PPQ90, PB95, RR93, Rus93, SZ98, Sch90a, SSLI97, SC99, SK98, SI99, WHF92, Wid98, dRIB99, zzMZ96, AG91, AC97c, Alb92, AC90a, AEGL92, BMO95, Bin92, dAB91b, BFB94, BV96, CBR95, CPS91, CMP95, CGMS96, CT96a, Con90, DKMM94, Deu92, FT90b, FT90c, GD93, Git90a, GS97, GGP92, HK96, HS90a, Hen94, HS90b, HKSY96, ILF90, IFR93, JMP94, JT96, KN93, Ken93, KGM92, KZ93, KNV93, KP91a, Koz97, KA94, tLZ96, Mad95, ML91a, MHJ94, MRS94, MLM93, MF92, MKZ92, NHT92, NY95, OBB95, Oze93, RHA97, Row97, RK93, SS96a, SS97b, SSV93, Swe97, Too94a]. critical [Too94b, VZ94, YIK95, MOS91b]. Criticality [GS98, Ste95b, CO97, DR91, Fis94, GS94, GK91, KM96, MKK90, Rap92a, SY95, WKWS95, Zha96b]. crossing [LPPSA92]. Crossings [JL98, KZ93]. Crossover [BI92, FT93, YK91, FOS94, Mik95, OSB97, Pri93]. Crumpling [vR97a]. Crystal [FM99, DFZ94, Fuj92, Gar91, GW90, GW94a, GW95, Jaf91, JR91b, KL91, MS95a, NMC^{+}91, SL95, VZ92, VZ95, vEM92]. Crystalline [SL93b, Tay99, For94, HS95b]. Crystalline-amorphous [SL93b]. crystallization [RW90, SP95b, Süt96, Yep95]. Crystals [DC98, MSS98b, Fra94, KP94a, MMSR92, Nob95, Sza97]. CSM [Cho97]. CTM [War96a]. CTRW [HW90]. Cubic [PRW99, Dia94, FMOU90, RS91a, WHF92, dME90]. Cubic-Invariant [PRW99]. Culture [PdO92]. cumulants [LS91a]. Curie [BG93d, GZ98, MP98a, VZ94, dMP91, dMPZ92]. Curious [BA93]. Current [DB90b, DBB^{+}92, CGK94]. currents [Mar93a]. Curvature [JT98, DOPT93, FL94]. Curved [SD99]. curves [CC91b, HKSY96]. cut [CPS92, DDJ^{+}95]. Cvitanovic [Pol91b]. CVM [TFD90]. cycle [MS90a]. cycles [DKKP96, Ste90, TB92, Voo92]. cyclic [BT94, PB90, RR97b, SRC93, SO92a, SO92b]. cylindric [SS94c]. cylindrical [WH95]. Cyril [Wid98].

D [Pod98b, Rap98, TF99, NFID92, AC97c, BQ90a, BMO95, CJK98, CMP95, FLS95, HP91, HDS98, Iof94, Ken94, LPT96, MOT90, NM99, RB94, Ser98, Shi93a, SS94c]. Daily [JVH98]. Damage [Alb95, HWD97, HDS98, Bag96, GPJ92, Gra95, KB90, MHdA90, Sta94b, JR94]. damage-spreading [Bag96]. dangerous [BT90d]. Data [Ber98, LW98b, BPH^{+}94a, BC94, JW97, LOP96b, LOP96a, tLZ96, Mol97, SF91]. Day [Ano99-30]. Days [Ano96j, EB95b, Ano97h, Ano98q]. Dealing [Nos90]. death [MS90a, RK96]. Debye [BK94b, GN93]. Decay [CCO99, Che99b, GZ97, Krá96, Liv95, RS91b, SZ99, Aiz94, BCFM95, ELMD^{+}90, Gra90, IOT92, JM96, MR94a, Mel93, MZ96, Pen95c, RW90, Zeg94]. Decaying [vEMZ98, Ken94, dHS92]. decimation [MO95]. Decoherence [MP98b]. Decomposition [SW99a, LOP91, OD96, PB94]. decrease [Shi93b]. Decreasing [BQ90b]. Dedication [Ano99n, WF90]. Defect [SI99, Elo94, HT90a, IOT92, JR91b]. defects [BT90a, BWK91, CPPG91, FK97, HS95b]. define [vEM92]. defined [vE96]. Definition [HWD97, Jun98]. definitions [GH97, Pes93]. Deformable [AQ98]. degeneracy [Fuj90b]. degenerate [Mar93b]. Degree [Wag98, ZXZY94]. degrees [BCS93]. Delaunay [BBD99, KK94b, Muc96]. Delay [Kus99, LM92]. delayed [ZR91]. delta [DMP97, Rei93]. delta-kicked [Rei93]. demonstrating [Rap90]. Demonstration [BLPP98]. Dense [JK99, Kaw98, CdS91a, CdS91b, Jag91, NDF92, SGP90, Sni91]. Densities [Kus99, BL91a, Fal92, HW90, Tha95]. Density [BGM98, BM99, Kaw98, Klo98, MY94, NM99, NP94, AK92, AC97a, AC97c, BC90a, Bel93a, Bel95, Bov90, FRHP95, GGM91, HJ90, LMM92, LP90a, MF92, RST91, Sni90, Süt96, Vie94, Vie95, YZ92, dMPS95]. density-corrected [Sni90]. Department [Ano91a]. Dependence [Alb95, TNN99, Cam91, Dan93, JFK91a, JFK91b, RB91]. dependences [BP93a, Pom93b]. Dependent [Alb98, BJ98, Gar98, BEC93, BJL92, Bob93, BC90b, FN97, GPJ92, JMM87, JD93, Joy96, KMKdC96, Pal90, PBSR97, JLM95]. depletion [KK91c]. depolarization [Dou92]. Deposited [CT98b]. Deposition [CGS95, Bel93b]. Deposition-evaporation [CGS95]. derangement [Gil90]. Derivation [APT94, BLO97, LY97, BP97, Cag90, EM94, KD90]. derivations [BGL91]. derivative [RST91]. derived [ADG96, GRZ90, Jur95, Str95]. descendant [BS90a]. described [Sle96, WT93]. describing [BP93a]. Description [FE98, Uen95, vES98, FJ93, FRHP95, RW91, SGP90, SH95]. Descriptions [FR99]. design [ASKK95, KK93a]. Desorption [ZK98, dMdO98, MSZZ90]. detailed [Che95, EB95a, Fri90, Jes96, LC95a, LPT96]. Detecting [Pra94]. Determinant [Shi98]. Determinants [BW99]. Determination [CV96, Voo92, CT95, Ras93]. determine [MHL92]. determining [GMO91]. Deterministic [BO99a, BF95, GKRT94, Mas92, QdL92, SA94, ZKB91a]. developed [CLT90a, CLT90b, Gal97, Shi92]. development [Mil95]. Developments [CCF99]. Deviation [BP98, CLLL98, DA99, LS99, GS97, LP96a, PR94b]. Deviations [Ber98, Che98, LFvdH97, Iof94, SSP95]. devil [AFNB97]. diagonal [PWG97, Sew90]. Diagonalization [KM98]. Diagram [BRZ98, EP97, FT94a, GdH91b, HLIM93, LPW92, Oze93, TFD90, ZP94, dABMR90, dOdOdSB95]. Diagrams [DFF99, GMR98, DFF96, Mon91a, Mon91b, Mon92]. Diamond [Elo99, RHdS^{+}91]. diatomic [MC92b]. dichotomous [KF93, OPdlR95a, OPdlR95b]. dielectric [Fig93, Fig94, JW93, Jur95, MS90b]. Diffeomorphisms [Wei99, Sim94a, Sim94b, Wei95]. Difference [Ném91, SY90]. Differences [Cro98, ZK93]. Different [HA97, SA94, CPPG91, MGS94, CO96a]. differentiability [YZ92]. Differential [AG99, GMH98, JSO99, BM92a, DI93, JLS96, KP92, LM92, MS93a, MT94, WK90]. differently [zzMZ96]. diffraction [RVW96]. Diffuse [Whe99, AM94d, STAJ95, TC94]. Diffusing [Asl99, CT98b, KWG96, MBD97]. Diffusion [AC95, BDG97, BCL^{+}99, BBW98, BSdB^{+}98, BCF90, CP99, DR98, Dou92, FA91, FM93a, FG99, Gas92, IdRB98, JS99, KC91a, KB97c, Mil91a, Mül99, NMHS99, NB90, OAB^{+}96, OS96b, Pae90, QdL92, RSL90, Rom90, SEW98, STV94, SS99, Ste99b, Tor91, WC95, WC96, WF91, ZSP90, dRIB99, vBE93, vdSE99, ABL97, AEA93, BEC93, BSTV94a, BSTV94b, BMS91, BL91b, BT94, BO91b, BB97, CSRPS93, Che90, CV93a, CD91a, CKS91, CW95, CSB97, CV93b, Dah96, DC94, DEM95, DBH91, Dow91b, DR91, Ebe96, FP97, GIT91a, GS92, GHW91, GS91b, GMM90, HW97b, HRS97, Hor93, IOT92, JFK91a, JFK91b, KC91b, KK91c, Koz94, Koz96, KPWH95, KPSW95, LS91a, LC95b, LC97, LSK91, LR96, Maj93, Maj94, Man93, MM93, MS95b, PX91, PS93c, PP95, PPQ90]. diffusion [Pri93, PCG95, QO95, RM93, RR97b, SS94b, Sch95b, SZ96, SCM96, Spo90, SZ91, THK^{+}91, Wei91b, WG95, Xin93, YL96, ZKB91b, bABD90, dHNR92b, dHNR94, vB90a, vD90, vRE93]. Diffusion-annihilation [FA91]. diffusion-controlled [BO91b, CSB97]. diffusion-driven [QO95]. diffusion-influenced [SZ91]. diffusion-limited [BMS91, BL91b, BB97, CKS91, Hor93, KK91c, Koz94, LC95b, LC97, LSK91, Pri93, PCG95, RM93, bABD90]. diffusion-reaction [DBH91]. Diffusional [dMdO98]. Diffusions [dHS98b]. Diffusive [BJ99, BKL97b, BJL^{+}91, FJ93, Ger93a, RK96, Cer96, Com91a, ELS96, PSZ93, Wan96b, YRHMJ92, vB91]. diffusivities [BEK91]. Diffusivity [KG99, SY95]. Dilute [AF99, Fel98, PSP94, SSLI97, SI99, VHR98, BG92, BG93c, CdHM91, GD93, HH93, KS93b, MGS94, Mat94, NFID92, WPSN94]. Diluted [BJS99, HY98, BM97, BSVZ94, DB90b, FN95, KRT97, WHF92]. diluted-field [FN95]. Dimension [CLV98, DA99, Klo98, MS98b, RRT98, Sim94a, Sim94b, Sim98, Tan98, AMA^{+}90, Alb95, ABP96, BJ90, CW96, CFP91, CLV97, CLS90, DN90, DLS97, FA91, FJT96, GD93, GH97, HS90a, JW97, KNV93, Koh92a, Kru92, LP94, Mil91a, Mur94, Pes93, PIM94, Por90, PN96, PN97, Pri92, REK91, RW90, Ser96, SMD92, VB90b, Wei92, ZPK97, Ano99g]. Dimensional [Ano99j, AHR98, Asl99, BC98a, BAZ98, BF99, BBG98, BCCP98, BHS99, BCL^{+}99, BMPZ98, BP98, CISS99, CP99, CL98, DHS98a, DC98, FOS94, GW99, JK99, JT98, Kin99, LM98, LMN98, MPR98, MSS98b, Nag98, OSB97, PRW99, SS97c, ST99, SSLI97, SZ99, SI99, Vuj99, WB99, vdH98, ABF^{+}95, ANV94, AC92, AC97d, ACJL92, AM94a, AORZ95, ABPSJ90a, ABPSJ90b, ABPSJ91b, AE91, AEA93, Bak93, BG93a, BEC93, Bar96, BF96, BB93, Ben91, Ben92a, BG90a, BR92, BT90a, BL93, BAP93, BBM92, Ble92, BRZ96, BCF90, dAB91a, BB94, BK96b, BHJ92, dAB91b, BFB94, BD93, BS91b, BW88, Bur91, CO91, CGK94, CMVG95, CGMS96, CSS95, CP97, CT95, CSB97, Con96, CMPS97, DFZ94, DHV92, DDM92, DEM95, DHP96, Dev91a, DPS90, DN94b, DB90b, DR91]. dimensional [DBB^{+}92, Elo94, EM95, ED92b, ES93, FM93a, FP90, FjLL95, FLL95, FT93, For90a, FM91, FJ92, FJM92, FJ96, Fos93, FG94, FR95, GIT91a, Gar94, Gar96, Gar91, GRZ90, GAA^{+}93, Gon94, GBP91, Gre90, GUJ94, GMMU97, Gry92, HMY96, HMP96, HW97a, HW97b, Heu93, HRS97, HP97, HH90, Hor93, HH96, HC91, HR92b, HPS94, Ino90, JS95, Jer90, JM96, Joy90, KLT97, Ker93, KF90, KPWH95, KPSW95, KKBS92, KKBS93, KK94b, KA94, LPPSA92, LS90a, LMS95, LF90, LBK96, MR94a, MC94, MSD92, MM96, Mar90a, MKP90, MK91, Mar92b, MP94, MP96, MOS91b, Mar94c, MHJ94, MBD97, Mie91, MF92, Mik95, MS95b, MC92b, Mon94, MSG97, Muc96, Mül93, MP93, Nob95, OAB^{+}96, OB91a, OP90, OPS93, OR95a, ON96, OL97, OP93, Pae90, PPS95, PS92b]. dimensional [PS93a, Per90, Per97, PPR93, Pol90a, PS93c, PPQ90, PS97c, Pri94, PPNM97, QdL92, Ray91, Rei96a, Rei96b, RW91, RC96, Rob91, RR97a, Rom90, RBF93, RS92c, Rub91, RDWW93, SA91, SS97b, SP95a, Sch90a, ST90, Sch93c, SD93, SDJ^{+}96, Sim96, Spe97, Spe91, Tan92, Too95, TZ97, UO91, Wan96b, WR97, WLC94, YT90, YT91, YHHK96, YL96, Zeg94, dB91, vB91, ST91]. dimensionality [SS94b]. Dimensions [BC99a, HY98, LFvW98, LD98, SL91, dVOS98, ANHKV93, AZ95, BB92a, BP94b, Bha90, CWSD92, CGS95, Cho94, CF97, Cut91a, Cut91b, Dro96, GHPS96, GW94a, GW95, GC90, GH94, HHD96, HS92a, HS92c, JR94, JMP94, JT97, KP91a, Krá96, Nas91, NY95, OP95, Pes93, PW94, PTN93, Ras93, SO91, Sta94a, TvROW96, VBF97, VWG93, Wee91, YZ92, LP96b, SL97a]. dimer [Alb92, BAK91, Hen97, Jer90, KRS96, Nag95, RHA97]. dimer-dimer [Alb92]. dimer-trimer [BAK91]. dimers [MBD97]. Dipolar [BGW98, Jag91, Jur95, WWW95]. dipole [AF95b, DH92, HTPH93, PO93, PPNM97, VR97b]. Dirac [BK91, Kho91a]. Direct [NGB95, BR91, Wag92]. Directed [BK92a, BRT98, BEO98, BH91, HWD97, JW98, KN93, Piz97, RSW98, AE90, ABPSJ90a, ABPSJ90b, ABPSJ91a, ABPSJ91b, BO90, Cao93, CPP97, CO96b, CD90, DMB97, Dua90, ED92b, Fos93, Gra95, HL97a, HKV91, KIKK97, MZ90, MK93, OS96b, OvR95, OPB93, PB95, PB94, SZ96]. directed-bond [KIKK97]. Direction [Alb98, Mil91b, Bur91, Maj94]. Direction-Dependent [Alb98]. Direction-direction [Mil91b]. directional [Cla91b, TZ97]. director [Mac95a, OR95b]. Dirichlet [Koj97, LPY98, RS92c]. disciples [AC97b]. Discontinuous [BC98a, GGL90, Vai92]. Discrete [BDM90, BS98, BC98b, BC99b, BMR95, CLV98, PS98, Pri93, dMM98, BC90a, BMP90, CW96, CLV97, CQ90, CP93, Cor95, FZ91, Gob92, Hof93, IW93, Ken90, KB91, Mar92b, MO94, Nad95, Ném91, Nev95, OP93, Sle96, Wag95, ZM93, BNN97]. discrete-state [BMP90, CW96]. discrete-velocity [Sle96]. discretizations [DKKP96]. discretized [WRJ95]. Discriminate [HLW99]. disease [CS90b]. Diseases [Sch99]. disjoint [Gri95]. disjoint-occurrence [Gri95]. disk [IM96, KW93, LS90b, PIM94, RVW96, SL95]. Disks [Hel98, LSP91, RC96, SL93b]. Disorder [BJ99, MPR98, MZ90, SK99, Spo95, AM94b, ABPSJ90b, Bov90, DHV92, DP92, Ebe96, GM94, GM95, GKRT94, HC91, Mar93b, MD94, Mun92, Niu91, OSE93, RV91, STAJ95, SSP95, Spe91]. Disordered [Ano99l, Bro98, CW96, Kül98, NK99, ABPSJ91a, Ben93, BMS97, Cao93, Dev91b, Eng91, EK96, GMCP96, GMN94, KB97a, Kül97, NS96, PHS^{+}97, SS92a, Tan94, Weh97b, YK91, Zha96b]. Disparities [Gol99]. Dispersing [Che99b]. Dispersion [CdB93, CHG94, GPS90, SCM96]. Dispersive [GKW91]. Displacement [Kik99]. disrrdered [Mar94a]. Dissipation [Eyi90, GK99, Fin92, FR95, Gal96, GMTB96, Opp94a, Opp94b, Wen97, vK95a]. Dissipative [Ano99d, BR99, BDS97, GD99, GW94d, KK92, Mas92, Per91, Sza97, TP92, Wol92, ZR91, vWL95]. dissolution [Kar95]. Distance [BQ90b, dVOS98, FJ92, OTT92]. distant [DT93]. distributed [Mil92]. Distribution [AC95, BNK98, BCCP98, DLB^{+}98, Hay99, JW98, KB91, MM98, MBF^{+}97, MSG97, SO92a, SO92b, Ald93, AKV94, Buo90, CLT90b, DB90b, Eng92, GM88, GM96, GR92, HTPH93, Man90, Man93, MZ90, Men92, OR91, Rap92a, SK90, TST91, YP95, YS93]. distribution-function [Buo90]. Distributional [Muc96]. Distributions [Bak98, TW98, AM94b, BNN95, CC91c, Cut91a, Cut91b, DF93, DBB^{+}92, GTW95, GK95, JW93, Kot95, PER95c, RW92, TF93a]. disturbed [Zyl90]. Divergence [Zha96b]. diversity [SP91]. divisors [FdlL92a]. DNA [SACB98]. Do [dSCT91, LP91]. Dobrushin [BCF97, SS97a, Yos97]. Does [GJ99, MSG95, RW91]. Domain [BNK98, Jun98, Ber94a, JSC91, KD90, OS95, OS96a, SBH92]. Domains [Hei98, NFL99, BK95a, FdlL92a, dlL92]. Domany [ZP94]. Domb [Wid98, Bar90]. dominant [MS91]. dominated [BLS91]. Döring [Pen97, Vel98]. d’Ottawa [Ano92q]. double [Hal97, HS90b, Pae90, HS92b]. Double-cluster [HS92b]. doubling [dSVG90]. doubly [For90a]. Down [dRIB99, GSCK90]. drag [DB90a]. drift [CV93b, FMS97]. Driven [ADE98, BJ99, DP97, EPRB99, ELS96, Gar98, tL90, SZ98, SSZ99, dlSG99, AM95, ANV94, ALLZ96, BZ95, BJL^{+}91, CL97, DN94b, Ger90, GW93b, GW94c, GDJH93, Kli91, MG92, MHM94, NV93, NBM90, OPdlR95a, OPdlR95b, PSZ93, QO95, RM93, VZL97, Wan96b, WG93, YRHMJ92, YMHMJ93, vB91]. drop [MHL92]. Droplet [KO93, CSS95, HT91, PN94, WT92, WT93]. droplets [ABK94, KO94, MOS91b, MKZ92, Nev95, Sta94a, WF91]. Drug [Sch99]. Drug-Resistant [Sch99]. Drunken [BGL99]. dual [Tab96]. Duality [AT90, Jos98, LK98, Sch97a, BMO96, GS93c, Lüt95]. due [BO91b, Pri93]. Duffing [WB94a, WB94b]. Duhem [Jar99]. during [ALLZ96]. Dynamic [CPS91, JL98, KE97, SS96a, SS97b, dHNR94, BGL91, Hio90, MS93b, MHJ94, Pen91b, Red94, bABD90]. Dynamical [ABPSJ90a, ABPSJ91b, BV98, CJK98, CFL98, CFJ91, DR93, DEJ95, Fel98, GC95, Ger99, Jia99, Kar99, Kaw98, Koh92a, Mar92d, Shl90, Sre98, ABPSJ90b, Aur90, Bal91, BCK97, Cla91a, DKKP96, EKLR94, EIK92, FTGW96, FdH94, GMTB96, Git91c, HMV90, IM96, Jur95, Jus92, KP94b, Mar95, Mar94c, MRC95, PBSR97, RB94, RS92b, Shi90a, Wei92, Wes91, dLPP90]. Dynamically [VZL97]. Dynamics [Ano97f, Ano99b, AMF98, BM92a, BES98, BE98, BBW98, BO99a, BJS98, BJS99, CCM99, DHW99, ESB98, Edw91, Fer98, FL99, GMR99, GGD91, HDS98, KF93, KD98, LS99, LZ98, ML91b, Moo97, Mur94, Pat96, Pet99, Pod98b, PN96, PN97, Rue99, TH96, VZ95, Yos98, dRIB99, ACDD90, BNN95, BNN97, Bav94, BHP94, BHP97, BWK91, BBL96, BBC94b, BVZ93, BDS97, BT93b, But93, CWSD92, CA93, CDG95, CDFG97, Com91a, DHP96, Elo94, FR96, FRHP95, GHPS96, Gas92, GMTB95, Gia91, GL97, Git90b, Git91b, Git92c, Git92a, Git94, Gob92, GVV91, HG92, HH93, Heu93, HP97, Ish96, ILF90, Jag91, Kło95, KO93, LAT95, LV93, Lie93, LLH92, Mae90, MHdA90, Mar92c, MP96, MOS91a, MOS91b, MY94, MND92, MR96, MWA95, MGJ92]. dynamics [MHM94, NNM93, NLT93, OTH92, ORG91, Paj95, Par91, PZ91, Pei95, PTN93, PN94, PPD94, RT90, RM93, RR90, RS92a, Rob91, RRGT97, RZ93, Ros93, SP95b, SS94a, SS94b, Sch95a, SBP^{+}93, Sel97, Shi90b, Shu93, SR93, Spo93, Spo96, Str97, SZ95, Tab96, TBK90a, TBK90b, Wil91, YRHMJ92, YMHMJ93, bABD90, dSL90]. dynamo [IM96]. Dyson [MNO97]. Dzyaloshinsky [AW90b].

early [GW94b, Joh90]. early-time [GW94b]. earthquakes [RK93]. echo [LZ91, Shi93b]. Econophysics [MS97]. Edge [KPS98, Klo98, RK96]. Edwards [vdH98]. Effect [BM96a, DHV92, HC91, KB90, KK93b, OL97, Pod98a, VR97b, BDM90, FT94b, LZ91, MKZ92, MSZZ90, OR95b, PWG97, Sew90, Shi93b, WHS^{+}91].

Effective [CCT96, CSRPS93, DFF99, HS91, HS92b, Lan95, SB97a, SB97b, BC92, Bre91, BG90b, GGL90, vLH97]. Effective-field [HS91, HS92b]. effective-medium [Bre91]. Effectively [JSA98]. Effects [DHS98a, GHW91, IRB^{+}99, IdRB98, ZP99, AS91a, AM94c, BMS91, BHP97, dAB91a, BK95b, BO91b, CCC^{+}90, CCST90, CLT90b, CLY92, CD91b, CV93b, FT93, FG94, GMO91, Guy91, JMP94, KT94, KT91, NBM90, TT94, vEAD90, vEAD91, SS94b]. Efficient [ZFB98, Koh92b]. effusion [Gas92]. Eigenfunctions [NV98]. eigenmodes [GC92]. Eigenstates [Leb99, DP91]. eigenvalue [Kom93, LS91b, PS97a]. Eigenvalues [Bar96, Fuj90b]. Eight [Fuj92, Fuj98]. Eight-Vertex [Fuj98, Fuj92]. Einstein [BCF90, But93, Che91, Pen91a, SH95]. Elastic [FPL99]. elasticity [SC96]. Electric [Gar98, GS95, Jag91, Sut92]. Electrical [Mat94, Opp97]. electrochemistry [FF95]. electroconvection [MdBM91, RHH91]. electrolytes [McC95, VB90b]. electromagnetic [FK94a, LSS97]. Electron [Abd98, NR98, BC92, BMC94, CGK94, KB97a, MM90a, RC97, YHHK96]. electron-proton [MM90a]. electron-screened [RC97]. electronic [YK91]. Electrons [BGM97, Dou92, LSS97]. electrorheological [CS91c, Jag91, Sta91]. Electrostatics [Hel98]. elements [Ber90a, Ber90b, BKW90]. eliminating [ILF92]. Elliptic [FV97, AC91, Ami96]. elliptical [KMKdC96]. elongational [CdHM91, DJ95]. elusive [MO96b]. Embedology [SYC91]. emergence [BLL90]. Emery [BFB94]. enantiomeric [HPS94]. encoding [IP90]. endpoints [CPS90, CPS92, MOS90a]. Energies [GMR98, Mar98, HS95a]. Energy [Alb98, BL94, CCO99, Cro98, Dro96, EK96, JV99, Lu99, ADG96, AZ96, Bax96, BCP96, BCS93, BMC94, Ble90, Ble91, BR91, BJL^{+}91, CO91, CdOW95, Cho97, CS91b, CD91b, DN90, Eyi95b, Fin92, For91, FZ91, Gro95, GR97, Joy96, Ken94, Kou90, Mar94b, Mar97, Mie93, MS95a, MC92b, Ole90, PS94b, Pok93b, RdO94, ST92, ST93, SLSA91, Sni95, ZK93]. Energy-level [BL94]. energy-transport [ADG96]. engineering [Opp97]. Engines [Ano99d]. enhanced [FP97, MS96b, Man93, YMHMJ93, ZKB91b]. Enhancement [Dha97, Mel93, Ger93a]. Ensemble [ABT99, Ano99f, BBC94b, MS93a, Ger93b, KB91]. Ensembles [Wid99, CE94, Elo94, GC95, Gal95, Geo95, LPS94, Ném91, PS97a]. Enskog [AC90b, CC91a, FS93, Hei98, KW97, Los90a, Los90b, Pol90c, Sle98]. Enskog-like [KW97]. Entanglement [vROS^{+}96]. Entropies [Bou99]. Entropy [BS97, BC99c, CC94a, CFL98, Che97b, Dol97, EPRB99, EK99, Gar97, Gas97, GD99, Han96, Jia99, LFvW98, Lu99, Per90, PG99, PG00, Ste99a, Wen97, AKV94, BP93a, BK92b, CC92, Cer97, DMB97, Eng92, Ge91, Ish95, Mür90, NP93, Niw97, PSZ93, Rob91, Rue96, Rue97, Wei92, Wei93a, Wei95]. Enumeration [TB92, Mer90, NFID92]. Environment [Ali99, CvD98, KM98, dHMP99, BP91a, Bra91, GM97b, OS96b, Piz97]. environments [BH91, Dou97, JK93, LM91, Sch92, SA95, SZ96, Wei96]. epsilon [MOR94]. Equation [BC99c, BB98, CM98, CCO99, Col98, DR98, DLPS99, Der97, DA99, EML98, GY98, HL97b, Hei98, Kaw98, LW98b, LY97, Lu99, Luo97, NMHS99, Naw98, PS98, Pal90, Pet99, TV99, WX97, Wen99, Wie98b, WB99, ADG96, APT94, AC90b, ACI91, AM94d, AFNB97, BP94a, Bar96, BRR96, Bec95, BCS^{+}91, BC95, BBM96, BVHP92, Bob95, Bob97, BB92b, BW88, Buo90, BED95, CC91a, CDPP90, CC92, Cer90, CS92, Cer96, CD91a, Dal97, DE96, DEJ92, ELMD^{+}90, EM94, ELM95, FIS96, FT93, FS93, GTW95, GS91a, GW94b, Gre90, HT90b, IW93, IS96, Ken90, KP94b, LM90, LOP96b, LOP96a, LLM95, Los90a, Los90b, LM92, MS93a, McK94, MLM93, MR96, Mol97, NS92, OO91, Opp97, Pen91b, Pet90, Pet93, Pol90c, RSGRP97, RR97b, Sai95, SS92b, Sch97b]. equation [SGP90, SW91, Sin91, Sni90, Sni95, Wag92, Wag95, WRJ95, WB94a, WB94b, Wen97, WG95, vKO97]. Equations [AG99, AF99, BDIV97, GMH98, IVDB98, JSO99, KM91b, KG99, McK99, Vel98, AM94d, BM92a, BC90a, BGL91, BLO97, BJO97, Bob93, BC90b, CNC94, CC94a, CdC94, DKMM94, DI93, DOPT93, For93, GS90b, Gou97, HH90, ILF90, JLS96, KM93, Kli92, KP92, Koh92a, Krá97, LIF92, LM92, McK95, MQ91, MZMQ90, MT94, MJ90, Pen97, Pol94, PS91b, PBSR97, PB95, RZ93, Ste97, WK90, WT93, Xu95, Yam96, dLPP90]. Equidistribution [BZ99]. Equilibria [JV99, BT95, Git90a]. Equilibrium [ACC^{+}98, BHS99, BDK99, FMPP99, HS95b, KP94a, MP98a, NS99, SS94a, Bal91, Ber92, BE92, CH92, CC92, Cer96, DFZ94, DP92, DGZ92b, ELS96, Fuj92, GTW95, GY93, GZ97, HR95, MM90a, MOS90b, MOS91a, MMSR92, MR96, MS95a, Per95a, PW97, Pet93, Por96, Zeg94, CM96]. Equipartition [KLR94, HS95a]. Equivalence [DGLS98, Geo95, LPS94, PSP94, WK97, Zeg90]. Equivalences [GS97]. equivalent [Cho97, GH97]. erased [GB90]. Ergodic [CLLL98, Che90, Che92, Che93, ET90, FL99, GL98, Ser96, Vai92, AZP97, RBF93, SZ95, Wol92]. Ergodicity [Che98, Gal95, RW96a, BK96a, Don99]. Ernst [Kob97]. Erosion [PSR98]. Errata [FLL95, PG00, ST91, Zim94]. Erratum [Ano01, CLV98, ED90a, GM96, HSS95, JLM95, LC97, Pol91a, SL97a, WB99]. Error [NP99a, SP97]. errors [FLB91]. Escape [GW93b, MV97, OPdlR95a, OPdlR95b, SL93a, Ger93a, KO94, MS96a, NBM90, RK96, Zha96a]. Essential [LPY98]. Estimates [Ber98, Wan96a, BHP94, BHP97, CC94a, MP94, SP97]. estimating [MW95]. Estimation [NP99a, SK98, Cut91a, Cut91b, NP93, Ser96]. Eu [Dor94]. Euclidean [Alb95, KNV93, LSK91]. Euler [MW91, Nad95, Oku90, Xu95]. Evaluation [KGM92, Phi91]. evaporation [CGS95, WT93]. Even [Ers94]. Events [BGL99]. Evidence [FGMA93, Ray91, BCF90]. Evolution [AEA97, CNC94, CDMV98, JG98, KLRT97, RLK98, Sta94b, WBG98, DOPT93, Man93, PdO92]. Evolutions [Kar99, FIS96]. evolving [RT90]. Ewald [SL97b]. Exact [AKK99, BMA93, BSG95, BOP94, BBOC91, BKV93, Buo90, CK98, CT95, DDM90, DDJ^{+}95, DJLS93, DEM95, DHP96, DBH91, ERS99, For90a, FJM92, FJ96, GS91a, HH90, HS90b, JL94, Kal91, MC92a, OP93, Pal90, PS94b, Pri92, PCG95, Ric97, SS92b, Sch97b, SDJ^{+}96, TF99, TLW91, YHHK96, CMPS97, DDM92, For94, MF92, NFID92, Sai96, dGN97, BT90a]. Exactly [Shi90a, Bha90, BS90b, Gas92, HM92a, MS90b, OvR95, Pen91a, PS97c, PTN93, SD93, VZ92, vD97]. Examination [SK98]. Example [Mon04, Wen99, MF92]. examples [CA93, HK96, LPS94, Shi90a]. Excess [CCO99]. exchange [KT91, MGJ92]. exchanges [BCS93]. excitability [DI93]. excitations [Aiz94, EJ93a, IFR93]. excited [CP97, Mom94, WB94b]. excitons [HC91]. Excluded [ZP99]. Exclusion [AKK99, AG98, DGLS98, Gui99, Kei98, Nag98, RSSS98, SK99, ACJL92, Bel93a, Bel95, BBM96, DDM92, DJLS93, DEM95, DLS97, EFGM95, FG94, JL94, Mac97, Sch93c, SD93, Sch97a, Sch97b, Str95]. Excursions [DJ95, DJ93]. Existence [AZP97, BC90a, BBD99, BDIV97, DLPS99, Fer96, Fig93, HR92b, HPS94, IVDB98, MS98b, Mou99, PS98, Xin93, AC90b, DEJ92, Pol90c, WL95]. Exit [BOV98b, OS95, OS96a]. Expanding [Wei99, AZP97, Liv95, Lov94, PW97]. Expansion [CT93, Mer99, PG99, RS97a, SZ98, Sle98, Wie97, Wie98b, BD94, BCS^{+}91, Dim90, Gar94, Gar96, KIKK97, Kło95, KS93c, LB96, Mar93b, MOR94, MO96b, MNO97, MD94, NFID92, OP90, Pol90a, Ste97, Süt96, WB94a, WB94b, WR97, Yan94, PG00]. Expansions [KOT98, WB99, BW88, CCT93, CK95, Cra94, GSH90, Iof94, Krá97, Lem95, MOR97, MR94c, Pen95c, Phi91, PPW94]. Expectation [BJ98, Joy96]. Expectations [ZK93]. Experiment [FGMA93, BS91a, vK94, Ram95]. experimental [DDJ^{+}95, KK91c, PAB^{+}93, Ram93, WKWS95]. Experiments [SW99b, ACG96, SHW92]. explained [Phi94]. explanation [dlL92]. Explicit [Kei98, Maj93, MQ91, MZMQ90]. explicitly [Cer94]. Exploiting [SP91]. exploration [Hay96, Sas95a]. explosion [CLY92]. exponent [Bal92a, CPS91, CB90, DHP96, GB90, KNV93, Mad95, MHJ94, MW95, NV93, RCB90, Shi92, vE90]. Exponential [BG93a, BCFM95, BKM93, Pol92, IOT92, JM96, MOS90b, Red94, vD90, MOS91a]. Exponentiality [BOV98b]. Exponentially [BHS99, OS95, OS96a]. Exponents [BW98a, BLPP98, BCP98, Con98b, Gar97, HY98, ANHKV93, ANV94, ABPSJ90a, ABPSJ90b, ABPSJ91b, BK95a, BV96, CT96b, Che97b, CD90, Dah96, DR91, Gon94, HS92c, KLT97, KS92, KK91c, KA94, LF93, LMS95, Lia91, Mar93b, PPS95, PPQ90, PB95, SGH93, zzMZ96]. expression [BPH94b]. Extended [Sch94a, VGC92, Sch95a]. Extension [GY98, MS92, VBF97]. extensions [RV97]. extensive [WA90]. extensively [BGP95b]. External [CMR98, GS98, NOZ99, Pet99, WC98, CCC^{+}90, CCST90, CS92, CP97, DM94, HRS97, Kho91b, KR92, LH94, Pat94, TS94]. Externally [PNT91]. Extinction [Mou99]. Extreme [BNN95]. Extremely [BJS99, BSVZ94].

F [Joh90]. face [BPO96]. Faceted [DHW99]. facets [MS95a]. Facilitated [ST99]. factor [CS90a, MSG95, dHNR94]. Factorization [Shi98, OP90, ZP93]. fails [Häg96]. Failure [BRT98]. Falicov [GJL92, GUJ94, GMMU97, Ken98]. falloff [Mar90a, MKP90, MK91]. Families [MSS98a]. Family [AKK99, Bou99, Hay96]. far [CC91b]. Farey [GUJ94]. fas [BCF90]. Fast [Rie93, vEMZ98, BB91, BCS93, CV93a, Mer90, SL97a, SL97b, SP97, SL91]. Fast-Decaying [vEMZ98]. fate [BHKL95]. faucet [dOP93]. FCC [TFD90]. FCHC [Hén92, vCEBS94]. Features [dMM98]. feedback [ZR91]. Feigenbaum [Pol91b]. Fermi [ACM95, BK91, BG90a, CG99a, CJ92, FST96, HL97c, Kar94, NM99]. Fermion [Cep91, Has98, NM99, HS94b, Kal91, KS97a, KvL92]. Fermionic [War96a, War96b, BK91, KB91, LM94b]. Fermions [PPO99, Sid99, WB99, LM94a, Mar97, Mat94, Per95b, BW88]. ferroelectric [BS95]. Ferrofluid [GZ98]. Ferrofluids [WM98]. Ferromagnet [dMM98, Pat93, Pat94]. ferromagnetic [AG91, CCC^{+}90, CCST90, DKMM94, Koz97, Sal95, UO91]. ferromagnetism [All95, Tas96]. ferromagnets [GD93]. few [Kal91]. few-fermion [Kal91]. Feynman [BC95, MS98a, Ord92]. FHP [Fri94]. Fiber [DS99, GW99, PHS^{+}97]. Fibers [Too99, TKG93]. Fibonacci [IT91]. Fick [LFvdH97, TG95]. Field [AN91b, AC98, AG98, BEC93, BJ99, BRZ98, CT98a, CCF99, CCMS99, CT98b, DJB98, FSB91, Gar98, GR98, GS98, KRT99, Kül98, MP98a, NOZ99, PP91, Yos98, dlSG99, ABHP90, APT94, AEA93, BKK^{+}92, BT91, BM92a, BK96a, BDM90, BRZ96, BK92c, CPP94, CSRPS93, CR94, Das95, Dev91a, DM94, FKV91, Fil94, FN95, FR96, Gar96, GS92, GS93a, GS94, GHW91, Git93a, GS91b, GS95, HS91, HS92b, Jag91, JMM87, JK93, Kho90, Kho91b, KR92, KNV93, KT91, Kül97, LH94, Lem95, Lia91, LS92, LFtH91, Luc93, MS93b, MY94, Men92, MWA95, NS95, NR90, O’C93b, OSE93, OR95b, OP93, Pen91b, PSZ93, Pok93b, PSP94, PS91c, PPD94, Rie93, Sam95, Sch90b, Str94, Sut92, VR97b, WMS90, Wu95, YK91, Zeg90, dMP91, dMPZ92]. field [dHS92]. Field-dependent [BEC93]. Field-induced [AN91b, FSB91, PP91]. field-swept [MS93b]. field-theoretic [KT91]. Fields [CMR98, HLW99, WC98, vES98, AM94a, Ber90a, Ber90b, BJL92, BHKL95, CCC^{+}90, CCST90, CGK94, Dim90, Dou92, EMHM95, GM95, Hav90, LP90a, LSS97, MS96b, MZ96, vEFK95]. Fifth [Ano99-30]. filling [MN96, MV91]. film [PB94, RBB95, RDWW93]. Films [BMSW99, CT98b, McC95, MdBM91]. filtered [CLS90]. filters [BPH^{+}94a]. finance [MS97]. Fingering [WM98]. Finite [AS91a, BF97, BE98, dAB91a, BKMS91, CGT99, CT96a, CD91b, Dan93, DJB98, For91, FG94, FMOU90, GS93a, JMP94, Koh91b, KM96, KT91, KPWH95, KPSW95, LMP99, LFvW98, LP90a, MG92, Mor92b, Mou99, PS97b, PM99, PR90, Rap92b, SS97c, Spe97, TF99, Wei91a, Yos98, dOdOCS95, vEAD90, vEAD91, Alb92, BMC94, BHP97, BK90, BI92, BK95b, BT90d, BT90c, BD93, BKW90, CH94, CR94, CMPS97, DR93, ED92b, FT93, FOS94, GS95, KT94, MOT90, MOR97, OBB95, Oer95, OP90, Sch94b, SLSA91, Smi90, SY90, Wan96b, Wan96a]. Finite- [BF97]. finite-difference [SY90]. Finite-dimensional [Spe97, ED92b]. Finite-Size [DJB98, SS97c, TF99, AS91a, dAB91a, BKMS91, CT96a, CD91b, Dan93, For91, FG94, FMOU90, GS93a, JMP94, KM96, KT91, KPWH95, KPSW95, MG92, Mor92b, PR90, Wei91a, dOdOCS95, vEAD90, Alb92, BHP97, BK90, BI92, BK95b, BT90d, BT90c, BD93, CR94, FT93, FOS94, KT94, Smi90, Wan96b]. finite-size-scaling [CMPS97]. Finite-state [Koh91b]. Finite-temperature [LP90a, BMC94]. Finite-Volume [Yos98, MOT90, OP90, Wan96a]. finitely [Bal91, DFF96]. finitely-many [DFF96]. fire [GK91]. First [AHR98, AE91, BLPP98, BRT98, BAP93, Bin92, Hay99, KG90, LK95, Men92, OS95, OS96a, BK90, BI92, BK95b, BB94, EM95, Ers94, KY93a, KY93b, Lop90, MG92, Mül93, Pat96, PR90, TDSR95, YT90, YL96, dOdOCS95, vK93]. First-Order [AHR98, AE91, BK90, BI92, BK95b, BB94, EM95, KY93a, KY93b, Lop90, MG92, Mül93, PR90, dOdOCS95]. First-Passage [BRT98, Bin92, vK93]. First-passage-time [KG90, TDSR95]. Fisher [Böt95, WRJ95]. Five [Jun98, Cer90, Sta94a]. Five-Moment [Jun98]. Fixed [LM94b, PRW99, AZP97, BPO96, BR92, CPS90, CG93, FST96, Ken90, MOS90a, PPW94, PS92c, Voo92, WR97]. fixed-point [Ken90]. flame [MS96b, PANG^{+}95]. flashing [LR96]. flat [Tas96]. Flatnes [KE97]. flexible [BKV93]. Flexural [KM91a]. flight [FBJ92]. Flip [IM96, Han96, Str95]. Flip-flop [IM96]. flipping [BT93b]. flop [IM96]. Floquet [BJL92, Joy94]. Flory [FT94a]. Flow [BDS99, BKW90, DP97, GK99, Sle98, Ste97, VWG93, Wag98, BS91b, CC94b, CdHM91, CP97, CL97, Edw91, Fig92a, Got90, GC92, KW93, KB97c, ML91a, MSG97, NAC91, OR95b, PTZG91, Pol92, RK96, RSGRP97, Sle96, TS94, WH95, WHS^{+}91, WL92, WLC94, Zha92, dOdOdSB95]. flowing [Sta91]. Flows [CK98, DLPS99, Luo97, AM92, Ben95b, CJ97, DJ95, Dou92, FP97, Hay93, HZLD97, JW97, KWG96, NGB95, RR97a, TQGO95, TC94, dSVG90, ZHCD95, dME90]. Fluctuate [MTG99]. Fluctuating [GP91, BB94, Fil94, GH94, ILF90, KM91b, AS91a]. Fluctuation [AE99, BLS91, BO91b, CLY92, FR95, GMTB96, HT90a, Mae99, ST92, BF97, ED96, Gal96, Opp94a, Opp94b, VZ92]. Fluctuation-dissipation [FR95, Gal96, Opp94b]. Fluctuation-dominated [BLS91]. Fluctuation-induced [HT90a, ED96]. Fluctuations [BB98, Bry97, DBH91, FBJ92, GS94, JVH98, KK98, Sos99, WA90, Zim93, Zim94, dMP91, vB91, ACJL92, Ber94b, BV96, BO91b, Dai90, ELMD^{+}90, Eyi90, ELS96, FKV91, GVV91, GGP92, JLM93, Mom96, MVZ97, OR95b, Piz97, PN96, RV97, SS92a, Shi93a, Sut92, VZ94, VZ95, VR97b, vD90, vWH97]. Fluid [AQ98, BGL91, BLPP98, SER99, Sle98, AL95b, BPH94b, BHP94, CWSD92, COA95, FP90, HSW97, HTPH93, KR92, LO97, MS95b, Nas91, Per97, PTZG91, Rob91, Sle96, WL92, WLC94, Zha91]. Fluids [AZ98, GSM98, Kaw98, BWK91, CDM91, CL95, CS91c, CV93b, ED92a, FST97, HR92a, HH90, HS94b, HS95c, Jag91, JM92, KM91b, KB97a, Koh92b, LW93, LC91, Los90a, Los90b, MY94, NP94, PPR93, SGP90, Sta91, Ste95b, SHG91, VDH97, WKWS95, YS93]. Flux [DDGZ97, WX97, Zif91, Dro96, Eyi95b, GMMU97, MN96, OMM93, Zie91]. Fluxes [AZ98]. foam [NE95]. Fock [BLS94]. Fokker [Bec95, CS92, Gou97, LY97, Pet99, SBP^{+}93, WB94a, WB94b]. Folding [Fer98, Fer94]. follicles [LP91]. Force [CP99, CS92, CP97, HLIM93, NAC91, Sin91, TS94, WHS^{+}91]. Forced [DIK98, GW93a, TDSR95]. Forces [AZ98, BL99, GZ98, IRB^{+}99, Mar97, Pet90, SGP90]. forcing [PAB^{+}93, Wal91]. forest [GK91]. Foreword [Boo92, LDBA91, LOQ95]. form [BS90c, Buo90]. Formal [BGL91]. Formalism [Mol98, PPO99, AvBED97, ABJM97, Ben95b, BG90a, Jus92, KK92, Lop90, PS92c, RS96]. Formation [SK99, All95, ABK94, CLD94, GUJ94, KH96, Mac93, Rot93, SO92a, SO92b]. Forms [LPY98, FT94b, Jur95, TE95]. Formula [Jia99, NM99, SK98, BC95, Böt95, GdH91b, LP94, Lit92, Mor90, RVW96]. Formulas [CDMV98, AKV94, BF97, BW88, Wei92, WB99]. Formulation [Kik99, EN93, HS94b, SMS96]. formulations [FRHP95]. Forrester [War96a, War96b]. Fortuin [KG95]. foundation [VGC92]. foundations [Sas95b, BDIV97]. Four [RRT98, DN94b, GG94b, VDH97]. four-bonding-site [VDH97]. four-dimensional [DN94b]. Fourth [Ano95c, Ano98q]. Fox [GN93]. Fractal [CGM^{+}98, Fam92, Fam96, HKSY96, Rei98, Smi95, Tan98, BMA93, Ber94b, BKV93, CPP97, CKS91, Cut91a, Cut91b, DF93, EMHM95, HKV91, KF97, LSK91, MS96b, McQ97, SA95, WH91, YS97, zzMZ96]. fractal-based [KF97]. fractal-time [WH91]. fractality [TG95]. Fractals [Adl93b, Ben93, Ben95a, Adl93a, Fal92, MG96, RR93, RCB90, RB90, SA94, ZKB91a]. Fraction [Leg98, Bal92a, Bal92b, MRS94]. fractional [FT94b]. Fracture [Ano99l, Eng91]. fragmentation [BC90a, CdC94, Sle96]. framework [FLS96]. France [Ano94a]. Francis [Wid98]. Free [Alb98, AZ96, Bax96, CCO99, CL98, Cro98, Gar97, GMR98, KvL92, MS99, BCP96, BR91, BCF90, CPS92, Che97b, CS91b, DDG97, DDG96, For91, GO95, GR97, MS95a, NS95, PS94b, RW91, Sai95, ST92, ST93]. Freedom [Wag98, BCS93]. freely [MdBM91]. Freezing [JSC91, VHR98, BB91, BF91]. Freidlin [Tót90]. Frenkel [HC91, LH92, Mac95b]. frequencies [Ble90, Ble91, GI92, VdSFC97]. Frequency [Jur95, ZP99, CV93a]. Frequency-independent [Jur95]. Friction [BHP97, HL97a, Pet99, BPH94b, BHP94, Kło95]. frictional [RB91]. frictionless [Ger93a]. Frobenius [Bec95]. Fröhlich [PVZ94]. Front [WX97, KS97b, Xin93]. Fronts [CCO99, Fer96, FR97, PX91, PANG^{+}95]. Frustration [Mie98, dMM98, Mie93, SC96]. Fully [BJS98, CKK99, PdO90, CLT90a, CLT90b, Shi92]. fully-developed [Shi92]. Function [DA99, Maz98, SZ99, ZP99, AE90, AC95, BGP95a, BP93a, BT90b, Bha90, BL93, BR91, BvV95, Buo90, CT95, GN93, GR92, GKT93, Kaw97, MC92a, MC92b, MSG97, MRC95, RST91, RS92c, SK90, Smi90, Tan92, TST91, dSVG90, YS93, vHW93]. Functional [Bax98, BC92, Kaw98, LS99, BFG93, CT93, CCT93, FLS95, LP90a, NP94, PB95, SP95a, SW91]. Functionals [BV98, BCP96, MF92, PS94b]. Functions [Ano01, CGTM99, Gol99, Koj97, PPO99, Sac98, TW98, Aur90, CBR95, CMPS97, Dah96, DH92, ED90a, ED90b, Fin92, GG94a, HWvB97, HTPH93, Hon96, HS95c, Jez96, KM93, Kie92, LC95a, Li90, LBK96, MM97b, O’C93b, OR91, Pen95c, PO93, Rus94, SO91, Sut92, VB90b, WA90, Weh97c, WWW95, YHHK96, Zyl90, vVBE93]. Fundamental [vdBV93]. Fundamentals [Opp98c]. Further [SK98, KvL92, LS92]. further-neighbor [KvL92]. further-neighbors [LS92]. fusion [BPO96]. Future [Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano90i, Ano90j, Ano90k, Ano90l, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91b, Ano91h, Ano91i, Ano91j, Ano91k, Ano91l, Ano91m, Ano91n, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano92j, Ano92k, Ano92l, Ano92m, Ano92n, Ano92o, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano93i, Ano93j, Ano93k, Ano93l, Ano93m, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l]. Future [Ano98m, Ano98n, Ano99o, Ano99p, Ano99q, Ano99r, Ano99s, Ano99t, Ano99u, Ano99v, Ano99w, Ano99x, Ano99y, Ano99z, Ano99-27]. Fuzzy [Ahm96].

G [Sac98, Sre98]. Gacs [dSM92]. Gallavotti [LS99]. games [BB95a]. Gap [CCM99, FK94c]. Gaps [ADE98, Fig93, FK94b]. Gas [Ano92a, BF99, BHS99, BEM99, BMPZ98, BH98, DHS98a, Fel98, JVH98, Koj97, Mie98, Mie99, NM99, PZ99, PG99, PG00, PS99b, Rap98, SEW98, SZ98, SSZ99, TV99, AM95, AN91b, AC92, AF95a, AC97c, AS95, ALLZ96, AB92, AORZ95, BEC93, BK91, BZ95, BR92, BSG95, BMC94, Ble92, BT95, dAB91b, BB92b, BEK91, BMOT90, BT92, BT93b, BED95, CBR95, CC94b, CDM91, CCD^{+}91, CMMC92, Che95, Che94, CT96b, CW95, Dah96, DH92, DEJ95, DP92, ET90, EB95a, FM93a, FJ93, FjLL95, FLL95, FFJ92, FBJ92, FM91, FJM92, GO95, GMM90, Hay92, Hén92, HL97c, HR92a, HP97, HPS94, JFK91a, JFK91b, KS93a, KGM92, KC91a, KSZ97, KB97c, KB91, LLM95, MM97a, MM90a, Mae90, MKP90, MP96, MM97b, MSG97, MR94c, MRC95]. gas [Mun92, Ném91, OR95a, Pal90, Per90, Per95a, PO93, PIM94, Pol94, PPNM97, QdL92, RC96, RZ93, Rot93, RBF93, SP95b, Sch90b, Sni91, Str97, SB97c, TLW91, Tót91, TB95, VWG93, WC95, WT93, Xu95, Yep95, dHNR92a, vBE93, vRE93, vdBDP92, Mar90a]. gas-kinetic [Xu95]. gas-liquid [KGM92, LLM95]. Gases [Häg98, MA98, MV99, dlSG99, AvBED97, Ber92, Bri96b, BvV95, Bun97, BE92, DFHR90, For90b, HM92a, KC91b, LFvdH97, MK91, Mey96, OSE93, Pae90, PS93c, Sha95, Spo90, SY95, vCEBS94, vVBE93]. Gasket [Jez96, HY96]. gaskets [HW97b]. gating [PAB^{+}93]. Gauge [BMOT90, BAKK^{+}90]. Gauge-invariant [BMOT90]. Gaussian [BvEN99, BG96, BG97b, CS91b, FAKA97, FZ91, HLW99, LW98b, LFtH91, Mar92b, Mol97, MSG95, NY95, Rei96a, RV97, VZ94]. Gel [SK98, RGdG97, vD97]. Gelation [Jeo99, RNCT98, BP90]. Genealogical [DJM99]. General [Too94a, AMA^{+}90, BG90a, BB95b, DN94b, HWvB97, HS92c, IOT92, OS95, OS96a, Pet90, Ric97, Uen95, OPdlR95a, Rei96b]. Generalization [HH95a, KG95]. Generalized [AEA97, BLS94, BC90b, BED95, Con98a, DRbA99, Kaw97, Krá97, Mou92, Sch93c, Sos99, Yan94, BPH^{+}94a, Che95, CLS90, FL94, GAA97, GMM90, KK92, MBF^{+}97, NHT92, Pes93, Pol90c, PS91b, RS96, SGP90, Shi90b, SH89, Tót94, Whi94, BL90]. generated [TS94]. generating [MG96, O’C93b]. Generation [NR90, MOS90a]. Generator [Zhi98]. Generators [RS96]. generic [Ble91, KB97a]. generically [Gra95]. Genericity [Wei95]. genetic [Ber97a]. Genetics [CBK99]. Gennes [FT94a]. Genome [AP99, ZM93]. Gentle [GO95]. geodesic [Ami96, Pol92]. geodesics [Weh97b]. Geometric [AZ98, Ben91, LS90b, MW91, PW97]. geometrical [KS92]. geometries [LSK91]. Geometry [Ban99, SCBIR99, GC90, GG94b, JM92, Lit92, PB94, RBB95, RDWW93, TV90]. George [HNS91]. giant [DLM^{+}93b]. Gibbs [GM96, AKMR98, Bal91, BBD99, BRZ95, BRZ96, BK93, BGP95b, CFL98, EKLR94, GR98, GLM95, GM95, GM88, GS97, Hay93, LPY98, LV94, Mae99, OPS93, PY94, PY95, PZ99, Sim94b, dMPZ92, vEdH91]. Gibbsian [Lef99b, LM97, vEFS93, vES98]. Gibbsianness [BCO99]. Ginzburg [BZW92, BRR96, BLO97, BB94, BB98, MHM94, Yau94]. Given [Bou99]. Glass [AC98, DJB98, BL93, CCC^{+}90, CCST90, CHK93, FdH94, Gro95, JR94, NS96, RdO94]. Glasses [Maz98, VHR98, Ben91, Bry96, DDJ^{+}95, FR95, HS91, HS92b, Par93, SSP95, SDJ^{+}96, ZP93, vE90]. Glassy [AMF98, BVR93, CG99a, FR96, FMPP99]. Glauber [GVV91, MHdA90, Pei95, SP95b, SZ95, Yos98]. Global [AC90b, Ber92, DEJ92, GMR99, Pol90c, MLL90, Mie93]. globally [Cra94, Jus95]. globally-coupled [Cra94]. golden [Ble90]. Goldstone [Alb94, MV99, SO91]. Good [KS97c, OB91a]. Goodisman [Opp98a]. Gordon [McK95, McK99]. Grabert [FP98]. Gradient [Nag98, GKRT94, Krá97, Mür90, TC94]. gradients [WT92]. Grain [DHW99]. Grand [GMR98, YP95, ABJM97, Ném91]. grand-canonical [ABJM97]. Grannan [MS92]. Granular [BCCP98, BSdB^{+}98, BDS99, EP97, KW93, RSGRP97, TQGO95]. graph [SF91]. graph-theoretic [SF91]. Graphical [CMR98]. Graphs [ABT99, BHK98, AE90, BP91b, Dua90, MW94, Mon97, Tel90, Wu96, Wu97a]. Grassmannian [BS90c]. Gravitating [BMPZ98, MP96]. Gravitational [BW98b, CT98b, PPD94, Str94]. Gravity [ABT99, Mil91a]. grazing [Gou97]. Great [MR94b]. Greene [FdlL92b]. Griffiths [ACC^{+}98, BM97, BFB94]. Grooves [DHW99]. Gross [PPO99]. Ground [CdOW95, FNW92, GJL92, GMMU97, Koj97, LS97, Mie97, NS96, Ben92a, BRZ96, CW96, DFF96, DDJ^{+}95, For94, Gro95, Mar97, MS91, McK94, PBP97, Pok93b, Pok93a, RHA97, SDJ^{+}96, Weh97b, YHHK96]. Ground-State [Koj97, CdOW95, LS97, NS96, Ben92a, Gro95, Mar97, Pok93b, YHHK96]. Group [BCO99, CGT99, CJB99, Col98, KOT98, LC99, PRW99, SEW98, Bov90, BP91a, CO97, DH92, Ebe96, Gal90, GMTB95, GMTB96, Git96b, GGP92, HK96, HA97, HCW96, Ken93, LC95b, LC97, LM94b, LIF92, MO93, MO95, MNO97, Nay93, O’C93a, PR94a, Per95b, PPW94, SR95, Sch92, Sco93, VZL97, YS97, dlL92, vEFS93, vEFK95]. Group-Theoretic [SEW98]. Groups [FV97, BMR95]. Growing [ADE98, ABK94, KO94]. Growth [AEGL92, CL95, KPANG98, KNS98, MW90, MSS98b, NFL99, PHS^{+}97, RNCT98, Sen92, ANV94, ABK94, BF95, Ber97b, BJL^{+}91, DS92, EN93, Fam92, Fam96, FT90a, FT91, GMCP96, Gar91, GW90, GW94a, GW95, GdH91b, GdH91a, HHD96, HP91, HT91, IT91, KD90, Koh91c, KL91, MG94, Mar94a, MHM94, OP95, PS97c, Pri92, SBH92, Too94a, Too94b, TH96, UW91, Zha92, vD90, ALLZ96]. Guessing [PER95c].

H [Ano99c, Che95, Rap98]. H-theorem [Che95]. habit [Fra94]. Hadamard [KK93a]. Haldane [BBM96, Ino90]. Half [CS92, Dal97, Kło95, GK95, LM96, MN96]. half-filling [MN96]. Half-range [CS92, Kło95]. Half-space [Dal97, GK95, LM96]. Hall [AZ98, BDM90, FT94b, FST97]. Hamilton [BW88, dLPP90]. Hamiltonian [WB99, AW90b, BG93a, BQ90a, BC92, Ben95b, BG94, BE98, CJB99, CL97, Com90, DC98, DMP97, EPRB99, EM94, FRHP95, GGP92, MSD92, MV97, STV94]. Hamiltonians [BF98, DFF99, DF98, KM98, SW98, Zyl90]. Hänggi [FP98]. Hard [BHW99, Fuj90a, Fuj90b, Fuj91, Sze96, AB92, ABP96, BT90b, BPH94b, BHP94, BDS97, CFP91, DE96, FM93a, GGM91, HSW97, JFK91a, JFK91b, KR92, KW93, KP91a, MC93, Mur94, PIM94, Phi91, RST91, RV91, SS94a, SS94b, Sch95a, Sch95b, Tót91, YS93, Zha91, vDL95]. Hard-Core [BHW99, AB92, ABP96, Sch95b, Tót91, YS93]. hard-disk [KW93, PIM94]. Hard-hexagon [Fuj90a, Fuj90b, Fuj91]. hard-rod [Zha91]. hard-sphere [BT90b, BPH94b, BHP94, CFP91, HSW97, Phi91, RV91]. hard-square [JFK91a, JFK91b]. Harmonic [Gar98, Ble90, Ble91, BMS97, Dag96, DPS90, Phi91]. harnesses [Too97]. Hartree [BLS94]. Hartwig [Böt95]. Hausdorff [JW97, KNV93, REK91, Wei92]. Hawking [RS94]. HCP [CHM99]. heaps [BP96]. hearing [Ami96]. Heat [Ano99d, BTT98, CG99a, COA95, LW98b, SW99b, BC95, HR95, LW93, MHdA90, Opp91, Sai96, WHF92]. heat-bath [MHdA90]. Height [KD98, Hen97, Pri94]. Heiles [BR99]. Heisenberg [AW90b, CS91a, CS91b, GS93c, Ino90, Kho91b, PBP97, ZP99]. helical [Sas92]. helicity [Dan93]. Helium [JVH98, Sch93b]. Hemmen [CHK93, Shi90b]. Henon [BR99]. Hepatitis [AAH98]. Herbert [HNS91]. Hermitian [Has98]. Herzfeld [Joh90]. heteroclinic [YT90, YT91]. heterogeneous [BAK91, Bre91, GS91b, Zha92]. hexagon [Fuj90a, Fuj90b, Fuj91]. Hexagonal [Cor95, PK97]. hexagons [KP91a].

Hiemer [Mon04]. Hierarchical [BK93, DD91, EMHM95, Koz97, PPR93, PRW99, Wie98a, dVOS98, BR92, Bov90, BP91a, BK92c, Cao93, LH94, LM94b, LPR91a, LPR91b, MK93, MOR94, MO96b, MNO97, MCC90, Mun92, PP93, PPW94, PSP94, RW96b, RHdS^{+}91, SO91, SR90]. Hierarchically [dHS98b, EJ93b]. hierarchies [Lev96]. Hierarchy [KOT98, ACI91, BPO96, Com91b]. High [Bov98a, FS99, GGM91, Hel98, LM98, LS90a, MOR94, NR90, SZ98, Ste90, ZFB98, Zhi98, ZP99, BBM92, CO91, CFP91, DN90, GAA^{+}93, HS92a, Ken90, Kło95, Koh90, LAT95, MO96b, MNO97, MZ96, NY95, Ray91, YZ92, vEFK95, vE96]. High-density [GGM91]. High-Dimensional [LM98, BBM92, GAA^{+}93, Ray91]. high-energy [CO91]. high-friction [Kło95]. High-Order [ZFB98, Ste90, LAT95]. High-Precision [FS99, LS90a, Koh90]. High-Temperature [SZ98, MOR94, NR90, Ken90, MO96b, MNO97]. Higher [BC99a, KP92, ANHKV93, CGS95, DPS90, GBP91, LS91a]. higher-dimensional [GBP91]. Higher-order [KP92]. highly [NS96]. Hill [McK94]. Hilliard [BLO97, BMSW99]. Histogram [HCW96, NP99a]. Historical [Wid98, Bru94, Row97]. histories [DK96]. Hit [BGL99]. HIV [KS90, PS90]. HIV-1 [KS90]. Hodgkin [LY97]. Hoffman [Whe99]. Holmes [Sre98]. Holstein [AAR92, BM96a, BGM98, LM94a]. homoclinic [DN94a, Hal97]. Homogeneous [CK98, Gun99, Lu99, Sch98, Wen99, ACI91, Bob97, GS91a, MOS91b, YIK95, Los90a]. Homologies [MSS98a]. honeycomb [HS90b, WC95]. Hopf [LM92, SW91]. Hopf-like [LM92]. Hopfield [Shi90b, BG92, BG93c, BGP95b, BvEN99, FMP92, Gay92, GSCK90, GR97, Koh90, Kül97, PST94, PZ91, ST92, Sen92, ST93, Shu93, TC91]. hopping [AvBED97, Pla90, TZ97]. horizon [Ble92, Dah96]. Horizontal [Sta98]. Hove [HL97c]. HPP [DEP92]. Hubbard [BLS94, GR98, MR94a, MF91, Mie91, Ste97, Tas96, Uel99, WG96]. Huggins [FT94a]. hull [Sto97b, Vol94]. Husimi [Mon91a, Mon92]. Huxley [LY97]. Hybrid [Abd98, FAKA97, ST97]. Hydrodynamic [BFSV91, BL99, CdHM91, CM92, CR97, AORZ95, CLT90a, CLT90b, GMO91, GC92, Koh91a, Koh92b, Rot93, WT93]. hydrodynamically [MGA97]. Hydrodynamics [ED90b, ELS96, MS99, SK99, ABF^{+}95, Bob95, DPS90, DBH91, Fri94, GP91, KM91b, NGB95, Pal90, SB97c, ED90a]. hydrogen [NLT93]. hydrogen-bonded [NLT93]. Hyperbolic [McK99, Che92, Che93, LSV93, McK95, Mon97, Pol92, PW94, TV90, Wei92, Wu96, Wu97a, Ano97f]. Hypercube [FT90a, FT91]. Hypergeometric [KIKK97]. hypernetted [LLM95]. hyperscaling [LMS95]. hypothesis [AF95b, CLT90a, Eyi95a, Eyi95b, Gal96, RBF93]. hysteresis [CDD94, KH96, MS93b].

Ice [Elo99, Nag95]. Ideal [JT96, JT97, Rei98]. Ideas [Rue99]. identical [MC93]. Identification [CFL98]. Identities [MPR98, BMO96, War96b]. Identity [Lu99, BG94]. II [WB99, AC97d, BES98, Bax93b, BSTV94b, BO99b, BHP94, BK93, BW88, CCST90, CdC94, CLT90b, FT90c, Ger99, GJL92, GZ97, HS92b, IVDB98, KS97a, KPSW95, Kül98, LP96b, LPT97, LPR91b, Los90b, MOS91b, Mon92, OPdlR95b, OS96a, Rei96b, Sim94b, Too94b, Vie95, War96b, WLC94, Dom97]. Ill [vE96]. Ill-defined [vE96]. illustrated [Dag96]. immiscible [OR95a]. Immobile [LSKB91, BLS91]. immune [CS90b, KS90, SP91]. immunity [Alb95]. immunology [Ahm96, KS93b]. impact [Mar90b]. Impenetrable [Koj97]. imperfect [NB90, OPdlR95b]. Implementation [Ano92a, Hén92, ZFB98]. Implementing [SL91, SL97a]. implicit [KP92]. improve [HJZM93]. Improved [ED92b, LM98, LY97, MN96, ZHCD95]. improvement [Ber97a]. impurities [CLT90b, DMP97, MKP91, Oer95]. Impurity [MSS98b, FLS96, SL95, TZ97]. impurity-perturbed [SL95]. Imry [PSP94]. incoherence [BNS92, SM91]. Incommensurate [KS93a, PS94a, BGM97, Gry92, JM90, PW90]. Incommensurate-commensurate [PS94a]. Incompressible [HL97b, Ben95b, EM94, KWG96, Lia91, ZHCD95]. increase [CLS90, Niw97]. increases [Gác90]. increasing [Mar94b]. Independent [Ano01, HWD97, Cer90, Jur95, Weh97c, YZ92, ZHCD95]. Index [Ano98a, Ano99a]. indices [BMO96]. indifferent [CG93, PS92c]. Induced [BN99, AN91b, AS95, BK95b, BB94, BV94, ED96, FSB91, HT90a, Hor93, MLL90, PP91, PM91]. inequalities [Bob97, CBR95, Gri95, Lem95]. Inequality [LPY98, Zyl90]. Inertial [NBM90, Maj93]. infection [PS90]. inference [Fer94]. Infinite [CD99, CCG90, JSA98, MM93, MGD98, PS91c, BG93a, Ble92, CT96b, CC91c, Dah96, EN92b, Gry92, Hua97, ML91b, Mie91, MGJ92, Pet90, Pok93a, PR90, Weh97b, dMPZ92]. infinite-dimensional [BG93a]. Infinite-order [CCG90, CC91c]. Infinite-range [PS91c, MGJ92, Pet90, PR90]. infinite-ranged [ML91b]. Infinite-scale [MM93]. infinite-volume [Pok93a, dMPZ92]. Infinitely [Rei98, Sal99, Gay92]. Influence [Gou97, Pat98, Böt95, Pat94, WT92]. influenced [SZ91]. Influences [GS98]. Information [FR99, EIK92, Li90]. information-processing [EIK92]. infrared [WR97, PPO99]. inherent [Rus94]. Inhomogeneous [GMR99, KB97b, NMHS99, Per93, SP95b, Wu97b, JLM95, SP93, Sch96a, vK91, Los90b]. Initial [Ber98, Der97, Hei98, AFNB97, BFSV91, BMP90, JW97, Koz94, LOP96b, LOP96a, Mol97]. initially [Koz96, THK^{+}91]. Initio [KFK91]. injection [TTPH91]. input [HRS97, TDSR95]. inspection [BB95a]. instabilities [Cra94, JR91b, Wes91]. Instability [BGM98, DP97, tL90, MO95, Spe91, Hio90, LSS97, PVZ94, YMHMJ93]. Instantaneous [Jeo99]. instanton [NBM90]. Instantons [PS97b, Rut92]. instrument [SBP^{+}93]. Insulator [SBZ98, For90b]. Integrability [BK96a, Dag96, MO96a, MBD97, UI96]. integrable [CC91b, CMVG95]. Integral [BAZ98, FT94b, Gol99, SMD92, BC92, Buo90, GY93, HS94b, Kho90, RZ93]. Integrals [BI99, BG96, CT93, CCT93, OTT92, Phi91]. Integrated [Klo98, dMPS95]. Integration [BAZ98, BG94]. Integrodifferential [BC99a]. intensity [BFG93]. interact [LP91]. Interacting [Ban99, BO99a, BO99b, BL99, Don99, KT99, KS97c, MV99, dHS98b, BFG93, CM96, DdH96, Dai90, FVY92, Fos93, LB94, LM94a, Mar97, Mat94, MKP91, Niu91, OPB93, OP95, Pen91a, Ros93, Sha95, SP94, TvROW96, Too94a, Too94b, Tót90]. Interaction [AQ98, AG98, BPO96, BWK91, FE98, Lef99b, Mer99, RS97a, AS91b, AdAI92, BD94, BT90c, CPP94, Con90, FZ91, KT91, MPdlR93, Mar92b, Mon91a, Mon91b, Mon92, OPS93, PPNM97, SP93, Sch90b, SS92b, SD95, Too95, TB95, VWG93]. Interaction-round-a-face [BPO96]. Interactions [BCO99, CGTM99, LMP99, LD98, RNCT98, SSZ99, SZ99, VHR98, vEMZ98, AW90b, BSG91, BCF97, CdOW95, CdHM91, GL97, Gri94, Jez96, Ker93, KY93b, KvL92, LH92, LSKB91, MSZZ90, NDF92, NP94, OP93, Per97, SC96, Süt96, TLW91]. interated [Göt96]. Interband [KPS98]. Interchangeability [BG90b]. Interface [All95, BSG91, BB98, Dev91a, KD98, Spo93, Str95, Whe99, ACDD90, BJO97, But93, Cag90, DS92, DMRR92, For92, HMP96, HP91, JS95, LM96, Per95a, RW91, SL93b, YMHMJ93, vLH97]. Interfaces [CHM99, AS91a, Blu94, BP91a, BK92c, BK93, BK96b, CDG95, CDFG97, KF90, Pod95b]. Interfacial [BMSW99, CISS99, JV99, MHM94, OB96, RV97, UO91, Fuj90a, Fuj90b, Fuj91, Fuj92, Mor92b, SPR91]. intermediate [FjLL95, FLL95, Hof96]. intermittence [BC95]. Intermittency [BCPV97, Eng92, Iso99, BKL97a, LBK96, Tha95, CLT90b]. intermittent [Ber97b]. Internal [Wag98]. International [Ram95]. interparticle [SD95, SLSA91]. interplanar [HLIM93]. Interplay [GZ98]. Interpolation [WR97, BG94]. Interpretation [Lie93, Kou90]. intersection [AZ95]. Interval [Ste99a, BK92b, CG93, LSV93, RS92b]. intervals [PS92c]. intractable [Jer90]. Intrinsic [Dai90, DGLS98]. Intrinsical [CK99b]. Introduction [DBD97, Fod97, Git94, Git96a, Wid98, Bry94, Bry96, Por96, Por97, Row97, Git91b]. Invalidated [Bis98]. Invariance [AHR99, GJLL99, Wie97, dHNR92a, AC90a, BT90a, GMTB95, GB94, Hen94, KB97a, LS90a, PW90]. Invariant [BV98, CC91b, FFK94, PRW99, dlL97, AZP97, BMOT90, CMPS97, DKKP96, DZ94, Din96, GBP91, HM92b, HKSY96, PS97a, RB94, Tha95, WPK95]. Invariants [BC99b, CK99a, GLR98, ML91a, Ber92, CCRM95, CMVG95, Cer90]. invasion [WL92, WLC94]. Inverse [Mer99, Sac98, WMTR90, dAB91b, CC91a, MK91, Mar92b]. inversion [Mor90]. investigating [Ram93]. Investigation [DJB98, ESB98, YIK95, Alb92, BT90c]. investigations [Sch95a, TFD90]. Inviscid [Der97]. involving [Cag90]. ion [DI93, HTPH93, Lie93, PAB^{+}93]. ion-channel [Lie93]. ion-dipole [HTPH93]. ionic [BHP96, HS95c, SS90, Ste95b, WKWS95]. Ionization [MM90a, Nd97]. ionized [LMM92]. ions [BB95b, LBT97]. Irreducible [GLR98, Kaw97]. Irregular [Hei98, KK98, LhBBS97, NS92]. irrelevant [BT90d].

Irreversibility [Lin93, Gal95, RBGW92]. irreversible [Alb95, CLHS91, Dor93, Eu94, Mil95, Sch94a, VGC92]. Ising [Alb94, ABHP90, ABCP96, ACC^{+}98, ABT99, BT91, Bak93, Bak94, Bak98, BM97, BNK98, Ben91, Ben92a, BMO95, BT90a, BS90b, BL93, BRZ95, BRZ96, BRZ98, BBC^{+}95, BVZ93, BSVZ94, BSV94, BJS98, BJS99, BZ97, BMOT90, CMP95, CGMS96, CMR98, CCMS99, CB94, CL98, CGTM99, Con96, CR94, Das95, DDJ^{+}95, DG90, Dev91a, EJ93b, FA91, FN95, GPJ92, GAA^{+}93, GS97, GS98, GB94, Gry92, HLIM93, HMY96, HMP96, HH93, Heu93, HY96, HDS98, HS92c, HY98, HS97, Iof94, Ish95, JSC91, Jez96, JK93, Joy90, KS97b, KE97, KRT97, KRT99, KF90, Kob97, KvL92, KO93, KA94, LM98, LH94, LF90, Luc93, MRV99, MM96, MHdA90, MOS90b, MOS91b, MO93, Mar94c, MHJ94, Maz98, MHB90, MOR94, MO96b, MOT90, Mik95, MZ96, Mon91a, Mon92]. Ising [Mon94, Mon97, Moo97, MP93, NOZ99, Nev95, Oer95, OL97, PP91, PPS95, Pen91b, PPR93, Pok93b, Pol90a, PSP94, RK90, RT90, Ray91, Rie93, RW92, RDWW93, Sal95, SP93, SSP95, SZ98, SS96b, ST99, SSLI97, Ser98, Shi93a, SDJ^{+}96, SC96, Sta94a, SC99, SI99, TBK90a, TBK90b, UO91, VHR98, WBG98, WHF92, Wu96, ZP93, dSL90, dMP91, dMPZ92]. Ising-like [HS92c, MP93]. Ising-Model [Bak98]. Ising-type [BBC^{+}95]. Isothermal [Sle98, RST91]. Isotropic [BG97b, dO92, dHS98b, Maj94]. issues [Sas95b]. Istanbul [Ano96j]. iterate [Ers94]. Iterated [Ano01, FMR94, Weh97c]. iterations [SY90]. iterative [Ber97c]. itinerant [LM94a]. Izergin [Sac98].

J [Ano01, Opp98a, Sre98]. Jack [FLS95]. Jacobi [WB99, BW88, Klo98]. James [Dom92, Dom97]. jamming [BP94b]. Jams [BH98]. Jan [Ano01]. Janco [AC97b]. Jancovici [RC97]. Jeans [BB91, BF91]. Joel [Ano98o]. John [Bry98]. Join [CPS92]. Join-and-cut [CPS92]. Josephson [Edi93, Jos98, LS97]. Journal [Ano91n, Ano98a, Ano99a, Ano94i, Ano97e, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano90i, Ano90j, Ano90k, Ano90l, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91b, Ano91h, Ano91i, Ano91j, Ano91k, Ano91l, Ano91m, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano92j, Ano92k, Ano92l, Ano92m, Ano92n, Ano92o, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano93i, Ano93j, Ano93k, Ano93l, Ano93m, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano97a, Ano97b, Ano97c, Ano97d, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j]. Journal [Ano98k, Ano98l, Ano98m, Ano98n, Ano99o, Ano99p, Ano99q, Ano99r, Ano99s, Ano99t, Ano99u, Ano99v, Ano99w, Ano99x, Ano99y, Ano99z, Ano99-27]. Joyce [Bar90]. Julia [Ish95, Tan98]. July [Ano96j]. jump [BK91]. jumps [ABL97]. Junction [Tay99, Edi93]. Junctions [Jos98]. June [Ano94a]. Jupiter [MR94b]. justification [DKS93, FdlL92b].

K. [dH94]. Kac [ABCP96, BC95, BZ97, Bov98a, BP98, CMP95, CMP97, FT94b, HS94a]. Kac-like [HS94a]. Kagomé [DG90, Fuj98]. Kagomé-Lattice [Fuj98]. KAM [CJB99, MG95]. Kampen [CT93, WB94a, WB94b]. Kardar [DA99, FT93]. Karl [Joh90]. Kasteleyn [KG95, Whi94]. Kauffman [KB97b, Sta94b]. Kaufman [Kau95]. Kawasaki [Alb94, CCM99, Gia91, MGJ92]. Keeping [FST96]. Kelvin [Sch93b]. kernel [BT93a, KP94b]. kernels [CC94a]. key [Mil95]. Kicked [BF98, DF98, Com90, Com91a, Rei93]. Kimball [GJL92, GUJ94, GMMU97, Ken98]. kind [GKT93]. Kinematics [GP93]. Kinetic [Abd98, Bou99, BDS99, DEP92, DLPS99, Dor93, FE98, GM94, KW93, KL91, Kra94, Mae90, NOV99, SBB98, ST99, WT93, ZK98, ANHKV93, BGL91, BCvB92, BvV95, Cer94, Cer97, EJ93b, Eu94, GK95, HT90b, HT91, JSC91, KW97, Lev96, MM96, OP95, Pol90a, PS91b, PANG^{+}95, Sni91, Sni95, SMD92, VGC92, WT92, Xu95]. Kinetics [CCO99, Kra92, SD99, SC99, UW91, Wei91b, AEGL92, BMS91, BLS91, BO91b, CL95, FA91, KD90, KK91c, MP93, OB91b, bABD90]. Kink [Num94, EN92a]. Kinked [Mar98]. Kinzel [ZP94]. Kirkpatrick [Bov98a, PS91a, Pat96]. Kirkwood [GS90b]. KMS [GMR99]. Knot [CCRM95]. knots [Dia94]. Kohlrausch [Phi94]. Kohn [PVZ94]. Kolmogorov [BBF94, CLT90a, CH94, CK99b, Gar97, HSK91, Lia91, Sta97a]. Kondo [FLS96, WG96]. Kontorova [LH92, Mac95b]. Korepin [Sac98]. Kosterlitz [Mar90a, AC92, AF95b, AC97a, AC97c, MKP90]. KPZ [PS97c, Wie98b]. Kramers [Sai95, Sam99, SGH93, Sos99, Ano99c]. Kubo [Che91, FR90]. Kuramoto [vHW93]. Kurdyumov [dSM92]. Kutta [MZMQ90]. Kutta-like [MZMQ90].

L [Opp98b, Sre98]. Laar [MLM93]. Labyrinthine [JG98]. ladder [FP92]. Ladders [ADE98, EN92b]. Lagrangian [Fer98]. Landau [BZW92, BRR96, BCS93, BLO97, BB94, BB98, DMP97, MS93b, MHM94, Yau94]. Landauer [SH89]. Landscape [RLK98, KLRT97, RdO94]. Landscapes [PSR98]. Landslides [Kru92]. Langevin [ACDD90, BCS^{+}91, Bob93, BC90b, CDG95, ELMD^{+}90, GL93, GMTB95, Kli92, Kło95, MS93b, MND92, OO91, Opp97, RSGRP97, SGP90, vKO97]. Laplace [HTPH93]. Large [Ano90n, Ano01, BE98, Ber98, BP98, CLLL98, Che98, DA99, EP92, Fil94, Iof94, JLM93, Kho91b, LS99, Lem95, Man90, Ros93, SSP95, Sos99, dVOS98, CD90, Dai90, Eyi90, ES93, GS97, Kli92, Koh91b, Koh92b, Lan94, Mar90b, Mie91, Mon91b, MSG97, OSB97, Pen97, PR94b, VdSFC97, Vol94, Weh97c, LP96a]. Large- [Kho91b, OSB97]. Large-Deviation [BP98, CLLL98, DA99, PR94b, LP96a]. Large-Deviations [Ber98]. Large-Distance [dVOS98]. Large-field [Lem95]. Large-scale [EP92, Fil94, Man90, ES93, Mar90b]. largest [Fuj90b]. Lars [Ano95i, Leb95a, Lyo95, Mac95a, Pen95b, Ano95b, Kau95, LHF95]. laser [FGMA93, Sha95]. Lattice [AL95b, AKR98, AKMR98, Ano92a, BF99, Bat98, BEM99, BRZ98, BDK99, BH98, CGT99, CDM91, CWSD92, Col98, FS99, FFJ92, Fuj98, GBCB99, Häg98, HM92a, HL97b, HR92a, Jia99, KS93b, Luo97, MA98, Mer90, Mie98, Mie99, NFID92, NR90, NFL99, O’C93a, OSE93, Pol94, PG99, PG00, PK97, PS99b, SEW98, SER99, SS98, SZ98, SSZ99, SF91, TB95, Vuj99, WH95, Yep95, ZP99, dlSG99, vdSE99, AM95, AK92, AN91b, AZ95, AM94a, ALLZ96, AB92, AORZ95, ABPSJ90b, ABPSJ91a, BK91, BZ95, Bax93b, BB92a, BAKK^{+}90, BGM97, Ber92, BSG95, BS90b, BL94, BRZ95, BT95, Bri96b, BB92b, BEK91, BvV95, BMOT90, BG90b, BT92, BT93b, Bun97, BE92, BED95, CC97, CCC^{+}90, CCST90, CCD^{+}91, CMMC92, Che95, COA95, CW95, DFF96, DG90, Dia94]. lattice [Dim90, DPS90, DEJ95, DFHR90, EG90, EB95a, FR97, Fig92a, Fig93, FK94b, FJ93, For90b, FMOU90, GM94, Gd96, GM88, GM96, Gon94, GMM90, GZ97, HS90a, Hay92, HZLD97, Hén92, HY96, HL97c, HP97, HR92b, HPS94, JFK91a, JFK91b, JD93, Jez96, JM96, KS93a, KIKK97, Ken90, KRS96, KC91a, KC91b, KvL92, KSZ97, KB97c, KB91, LH94, LPS94, LM97, MM97a, Mad95, Mae90, Mar90a, MKP90, Mar90b, MGA95, MGA97, ML92, Mey96, MZ96, Mom94, Mom96, Mor92a, MH92, MKP91, MD97, NS92, Ném91, NGB95, OP90, OPS93, OR95a, ON96, OvR95, OvRW96, Pae90, PS94a, PY94, PB90, Per90, Per93, PO93, PWG97, PS93c, QdL92, QO95, RS91a, Rot93, RHdS^{+}91, Sal95, SP95b, SR90, SD95, SP94, Spo90, SY95, Spo95, SB97c, SAB95, TLW91, Tót91]. lattice [VWG93, WC95, WG96, WLC94, WG95, Zeg90, ZP93, Zie93, ZHD95, ZHCD95, dO92, dHNR92a, vBE93, vCEBS94, vROS^{+}96, vRE93, vVBE93, FK94a, TQGO95, Rap98]. lattice-Boltzmann [HR92a, SD95]. Lattice-Gas [BH98, Mie98, PG99, PG00, HR92a, Yep95, GMM90, KS93a, QdL92]. lattice-point [BL94]. Lattices [Asl99, Jus98, KK98, SCBIR99, BBC^{+}95, BKV93, BLL90, Cao93, Dol97, Fer96, GMN94, HS90b, Jus95, Ken97, MK93, Mil92, MCC90, NS92, OBB95, Pae90, Pen94a, PS94b, PN97, RSL90, Tan94, WC95, WC96, YS97, ZP93, zzMZ96]. Law [Ano01, CCO99, CSS95, GPS90, IT91, LOP96b, LOP96a, LFvdH97, MKP90, MK91, McK94, Niw97, Pen94b, TG95, Weh97c, Yau94, Mar90a]. Laws [Bou99, KT99, NOV99, CDM93, IOT92, KD90, MBD97, TC94]. Layer [MRV99, HT90b, LK95, WT92, vLH97]. Layered [Jos98, BSV94, Bur91, Dou92, PX91, STAJ95]. Layering [DM94, CM96]. Layers [Ser98, Too99, WT93]. lead [RW91]. Leading [Fer98, ZXZY94]. leaky [dOP93]. learned [BMHH97]. Lebesgue [CK99b]. Lebowitz [DP97]. Lebwitz [Ano98o]. Lee [For90b, Koz97, PS93b]. legacy [Sta97a]. Length [IS99, CPS92, FN95, Fuj90a, Fuj91, KNV93, MOS90a, MM90b, PS91c, Zha92]. lengths [Ami96, AM94b]. Letter [Kau95]. letters [Dom92, Dom97]. Level [CLLL98, WC98, AM94a, Ble90, Ble91, BL94, BJL92, EK96, Lop90, Ole90, Per91]. level-2 [Lop90]. Levermore [Jun98]. Levin [dSM92]. lévy [FBJ92, JW97, JW93]. Lévy-stable [JW93]. Liblice [Ano90p, Ano95c]. Liesegang [CLD94]. Life [RK96, Shl91]. Lifetimes [NR98]. Lifshits [BHKL95, HLW99]. Lifshitz [CSS95, Vel98]. Light [KPS98, Fig92b]. like [HS94a, HS92c, KW97, LM92, MZMQ90, MP93, PW97, Sch90a]. Limit [AMF98, BGW98, BEM99, BL99, CC91c, FKST99, FE98, MM98, NMHS99, PS97b, SBZ98, vdH98, APT94, BKK^{+}92, BQ90a, BFSV91, BFG93, CCRM95, CJ92, CR97, DdH96, DN90, DE96, Don95, ELM95, Gay92, Gia91, HL97c, Kho91b, Kie90, Kie92, Kło95, Koz97, Mae90, McK94, MS90a, NY95, OSB97, Pok93b, Pok93a, PW94, Shi93a, SY90, TB92]. limit-cycle [MS90a]. Limit-Periodic [SBZ98, SY90]. limitations [vEFS93]. limited [BMS91, BL91b, BB97, CKS91, Hor93, KK91c, Koz94, LC95b, LC97, LSK91, Pri93, PCG95, RM93, bABD90]. Limiting [GR98, BRZ95, OR91]. Limits [LW98b, BGL91, BGR94, CM92, CMPS97, DEP92, GL97]. Line [Mar98, SZ98, AC97a, AC97c, CCST90, Kra92, MC94, MLM93, Str95, vE90]. Linear [DLM^{+}93b, Jus98, KS97c, Los90a, Los90b, RS97b, BPH^{+}94a, BCvB92, Che92, Che93, GW94c, GK95, Lin93, Mil95, Pet90, Pet93, PM91, Sta92, WCT91, Opp94a, Opp94b]. Linearization [JSO99, BP94a]. Linearized [Luo97, HT90b]. linearly [Cer90]. linearly-independent [Cer90]. Lines [SI99, Dro96, HT90a, KGM92, Tan94, Zie91]. link [Mat90, WPK95]. Linking [TC94]. Liouvillian [Gas92]. Liquid [DC98, LMP99, BG90a, CBR95, FLS96, Jaf91, JR91b, KM91a, KGM92, LLM95, NMC^{+}91, Per95a, RW91, TT94]. liquid-gas [CBR95, Per95a]. Liquid-Vapor [LMP99, RW91, TT94]. Liquids [Ano90p, Ano95c]. Little [PZ91, Shi90b]. lived [ES93]. Local [Eyi95b, Gd96, IVDB98, JSO99, KD98, Lu99, Mol98, AZ95, AE91, ELS96, HR95, Hol94, LIF92, Mie93, PS97a, Rus94, Sch90b, Shi92, Uen95, VR97b]. locality [Bri96a]. Localization [BBC^{+}94a, FK94a, FK94b, Ger99, AvBED97, Ben91, Bun97, Fig92a, Gra94, LB94, Mer92, Nos93]. Localized [FK97, JR91b, BMS97, GKRT94]. locally [Nad95]. Location [Fos93, OL97]. Lockings [DIK98, TDSR95]. Log [LPY98, AF95a, For90b]. log-gas [AF95a]. log-potential [For90b]. logarithm [RW96b]. Logarithmic [BD93, GD93, SS97c, Wan96a]. logistic [Ste90]. London [Wid98]. Long [AKR98, AP99, BHS99, BL99, CP97, CF97, FP98, Fer98, GSM98, IRB^{+}99, KB97a, Lef99b, MM97b, Nic93, SZ99, VHR98, dHNR92b, vB90a, vEMZ98, ABL97, BT90c, ES93, GL97, HW90, Jez96, Ker93, KT91, Koz96, LC91, LS97, Mae90, MZ90, MP94, MR96, MHM94, NP94, OYSK91, OPS93, PWG97, RK90, Sew90, dHNS92]. Long- [AP99]. long-lived [ES93]. Long-Range [AKR98, BL99, IRB^{+}99, Lef99b, SZ99, VHR98, vEMZ98, KB97a, vB90a, ABL97, BT90c, Jez96, Ker93, KT91, LS97, Mae90, NP94, OPS93, PWG97, RK90, Sew90]. Long-Ranged [GSM98]. long-tailed [HW90]. Long-term [Nic93]. Long-Time [FP98, Fer98, CP97, CF97, MM97b, dHNR92b, Koz96, MP94, MR96, MHM94, dHNS92]. long-time-scale [OYSK91]. Longest [SR90]. longitudinal [PS91c]. Looking [SBP^{+}93]. Loop [Bat98, GB90]. Loops [NM99, KLMR90]. Lorentz [AvBED97, BEC93, BF99, Ble92, BMHH97, BT92, BT93b, Bun97, Che94, CT96b, CW95, Dah96, DEJ95, Fel98, KC91a, KC91b, LFvdH97, MM97a, MM97b, MR94c, MRC95, Pal90, PIM94, WC95, dHNR92a, vBE93]. Low [AMA^{+}90, BM99, CCM99, DFF96, DFHR90, Gar96, LM94a, Mom94, Sal95, Süt96, BDM90, BZ97, Con96, Gar94, JR94, Ken97, LM96, MOS90b, MOS91b, Mar92d, Nev95, Par91, Pei95, PN96, PN97, PTN93, SO91].

Low-concentration [AMA^{+}90]. Low-density [Süt96]. Low-dimensional [Gar96, Gar94]. low-field [BDM90]. Low-lying [Mom94]. Low-temperature [DFF96, LM94a, Sal95, BZ97, JR94, LM96, MOS90b, Mar92d, Par91, Pei95]. Low-viscosity [DFHR90]. Lower [Gol99, FMP92, HSS93, HSS95, Iof94, Weh97a]. LSW [NP99b, Pen97]. Lumley [Sre98]. lure [Sta97c]. Lushnikov [BP90]. Luttinger [FLS96]. Lyapunov [BCP98, CT96b, Che97b, CD90, Dah96, DP97, GK96, Gar97, GKC94, LC95a, LPT96, LPT97, Mar93b, NV93, PPS95, SGH93, Shi92, Wei99, vHW93]. lying [Mom94].

M [Pod98a, Sac98]. Ma [PSP94]. Machine [Hén92, ABF^{+}95, BS90b]. Macro [Ano98o]. Macrodynamics [EP92]. macroions [HTPH93]. macromolecules [FSB91, Pod95a, Pod97]. macrophysics [Git92b]. Macroscopic [Gra90, LC95a, Sto97a, Cag90, SH95, GL97]. Macrotransport [Bre91]. Magnetic [GZ98, HLW99, Yos98, BK96a, BHKL95, CGK94, CR94, DM94, Gar96, HMV90, JMM87, LP90a, Luc93, OR95b, YK91]. magnetically [AEWF91]. magnetism [Kun94]. magnetization [AL95a, DN90, KG90, Kli92, MCC90, RS91b, Pat94]. magnetization-energy [DN90]. Magnetized [CCST90]. Magnetizing [CKK99]. magnetofluid [JT97]. Magnetohydrodynamics [CMMC92, SS90]. magnetotactic [vK95b]. Magnets [ZFB98, Gar94, Gar96]. Magnons [MKK97]. Majority [Ken97, Moo97, Che97a, Gal90, HP91, Ken93, TC93, dSCT91, dO92]. Majority-Vote [Moo97, dO92]. males [Ber97a]. Man [BGL99]. Mandelbrot [DM90, Ste90, WL95]. Manev [BDIV97, IVDB98]. manifestations [Git93b]. manifolds [Pol92, dlL97]. Many [BMS91, Leg98, RS97a, Sal99, VB90b, AZP97, BGP95b, Buo90, CV93a, DFF96, For93, For94, Gay92, GSH90, Gri94, GKT93, KT94, LP91, LMM92, Nas91, OR91, OB91a, YIK95]. many-arm [OB91a]. Many-Body [Leg98, RS97a, BMS91, VB90b, Buo90, For93, For94, GSH90, Gri94, GKT93, KT94, LP91]. many-frequency [CV93a]. many-particle [LMM92, OR91, YIK95]. Map [DO97, Jus98, KK98, BBC^{+}95, BLL90, CG93, Dol97, Fer96, FR97, Jus95, LSV93, LBK96, Pey91, PM91, Ste90, TG95, Vai92, WK90]. Mapping [Has98, FMR94, Göt96, ZM93, ZXZY94]. mappings [BG94, Lov94]. Maps [Col98, Hay99, KOJ98, LC99, PS92c, Ste99a, Tan98, Wei99, AC91, BNN97, BCCF92, BK92b, CC91b, CK91, FT90b, FT90c, GRZ90, HG92, HSK91, IP90, KK94a, LB94, Liv95, OD96, PW97, Rad93, Rei96a, Rei96b, RB94, STV94, Tha95, YT90, YT91, vEFK95]. Margulis [Sim94a]. Markov [Bak93, Bak94, Bak98, Che99a, DN97, Ers94, OS95, OS96a, Sch93a, Sco93, vWL95]. Markov-property [Bak93]. Markovian [BP97, BMP90, CH94, Cro98, IMS92]. Marriage [NFID92]. Maslov [Lit92].

Maslova [Hei94]. Mass [AKMR98, BMPZ98, Ald93, HR95, vB91]. massive [BPH94b, BHP94, BCF90, BG96]. Master [MJ90, PBSR97, Sch97b, vKO97, Ano99c]. Materials [BRT98, CCF99, WWW95]. Mathematical [Ano99i, Ano99-31, BB95a, GS90c, Pea95, Pes93]. Mathematics [Ano91a, Bru94]. Mathieu [Ger99]. Matrices [BW99, Klo98, Shi98, TW98, BF97, Bax91, Bax93b, CD90, Dev91b, Eva92, FM91, FN97, GK96, HJ90, Kom93, LMM92, Mar93b, Pey91]. Matrix [HP97, LW98a, Wid99, BC96, FZ97, Fuj90b, Koo95, MD97, PS97a, PB90, Per91, PS91c, Vie94, Vie95, dMPS95]. Matrix-product [HP97]. matter [Kie90, KK93b, Pod96, Sel97]. Mattis [Oze93]. Maximal [NV93, Ish95]. maximize [MSG95]. Maximum [MC93, Wei93a, Eng92, PSZ93, Rob91]. maximum-entropy [Rob91]. Maxwell [Dom97, APT94, FK94c, TV99, Dom92]. Maxwellian [BCCP98]. Mayer [BW88, Joy90, PPNM97, WB99]. McKean [DdH96]. Me [STAJ95]. Mean [AC98, AG98, CCMS99, DJB98, Gar97, GR98, Kül98, Maz98, APT94, BKK^{+}92, BT91, BM92a, Ble90, BR91, BB95b, BHP96, BD93, Cam91, CPP94, Che97b, Dan93, DDG97, DDG96, FKV91, FR96, GS93a, GS94, HS97, Kül97, LS92, NS95, OSE93, Pen91b, PSZ93, PS91c, Sch90b, VBF97, Wu95]. Mean-Field [CCMS99, DJB98, GR98, Kül98, APT94, BKK^{+}92, BT91, BM92a, CPP94, FKV91, FR96, GS93a, GS94, Kül97, LS92, NS95, Pen91b, PS91c, Sch90b, Wu95]. mean-spherical [HS97]. Means [CK98, LY97, Ras93]. Measure [Bis98, Fer94, Ge91, GH97, Sim94a, ZPK97]. measured [Lie93]. measurement [RHH91, Opp98b]. Measurements [Cro98, GS90a]. Measures [CFL98, CT96b, DGLS98, Lef99a, Lef99b, LPY98, AT90, AZP97, BK93, BMP92, CDFG97, CH94, DKKP96, DZ94, Din96, FFK94, GBP91, GS97, Han95, Hay93, HM92b, Ish95, LV94, LM97, MW91, PW97, Rad93, Sim94b, vEdH91, Jia99]. Measuring [Pom93b]. mechanical [BKL97a, MR94b, vK95a]. mechanically [MGS94]. Mechanics [Ano90m, Ano90p, Ano91a, Ano91o, Ano93a, Ano95c, Ano97f, Ano98p, Ano99g, Ano99f, Ano99-28, Ano99-29, Ano99-31, BJ98, CG99b, Leb90, Leb91, Leb92a, Leb92b, Leb93a, Leb93b, Leb94b, Leb94c, Leb95b, Leb95c, Leb96, Leb97b, Leb97c, Leb98, McK99, Opp98a, Opp98c, PS99a, Rue99, AEWF91, Ano91q, Ano92b, Ano92q, Bax95, BMR95, Bri96b, CHG94, DN97, DK96, DGZ92a, Edi93, GMN94, HH96, Joh90, Joy96, Kar94, KY93b, KH94, LMS90, Leb97d, Mac95b, McK95, Muk91, O’C93b, OS91, Omn91, Opp96a, Pea95, Pod97, RR97a, Rue96, Sas95b, Sza97, dMPS95, dB92, vKO97, vdBV93]. mechanics/mathematical [Pea95]. mechanism [CLS90, GL93, RM93]. mechanisms [CO96a, CS91c]. Media [BGW98, BCCP98, BSdB^{+}98, GBCB99, SB97b, SBB98, Ber97b, BK93, BK96b, Bre91, CB90, CdB93, DdH96, DC94, Eng91, EP92, FK94a, GMCP96, GS91b, GGL90, GdH91b, GdH91a, Koh91a, LBT97, ML91a, Mar94a, Nos93, PP93, PX91, PANG^{+}95, SB97a, Sza93, Xin93, Zha92, Zha96b, vK91]. Medium [Bro98, MSS98b, SB97b, Bre91, BMS97, BS91b, CM95, CD91b, ED92b, FIS96, Fig92b, KPW91, NB90, SB97a, SS92a]. Meeting [Ano90o, Ano91a, Ano91q, Ano93a, Ano93o, Leb92a, Ano92b, Ano92p, Pea95, Ano90m, Ano91o, Ano97f, Ano97g, Ano98p, Ano99-28, Ano99-29, Leb90, Leb91, Leb92b, Leb93a, Leb93b, Leb94b, Leb94c, Leb95b, Leb95c, Leb96, Leb97b, Leb97c, Leb98, Leb97d]. Meissner [PWG97, Sew90]. Melbourne [Ano91a]. Mellin [OTT92]. melts [KN93]. Melzer [War96b]. Memberships [MSS98a]. Membrane [AQ98, BKW90, RS94]. membranes [DMB97, Pod95b, WF91]. memoriam [Ano96h]. memories [DD91]. memory [FMP92, Hei94, Kaw97, RM93, Shu93, Vol94]. meniscus [DDM90]. Meromorphic [OTT92]. mesh [DT93]. Mesoscopic [Ber99, CGK94, Mar93a]. Meta [Ano98o]. Metal [SBZ98, AM94c]. metallic [Dev91b]. metals [EK96, Sai96]. metamagnet [Das95]. Metastability [AH98, AG98, CO96a, CL98, MOS90b, MP98a, PS90, Sim96, TT94, Yau94, dABR91, Sco93]. Metastable [KE97, Kli91, Pei95, Pen95c, Sle96, BK94a, ELMD^{+}90, GW93b, Gra90, KO94, Mel93, RT90]. Metastates [Kül97, Kül98]. Meteorological [JVH98]. Method [AQ98, Bak98, CGTM99, Fuj98, JSO99, NP99a, Sac98, Bak93, Bak94, Bha90, BV94, DZ94, Dua90, FS93, GAA97, GMO91, GK95, HS91, HS92b, JS95, Kom93, MHL92, MW95, Mor90, PSZ93, RDWW93, SL97a, SL97b, SMD92, SP97, Sta97b, Wag92, YP95, dME90, SL91, Ano99f]. Methods [Ano99i, Ano99e, Git92b, BMP90, DN97, DT93, EMHM95, Gd96, GB90, Jes96, JLS96, KPSW95, MS90b, NFID92, SF91, War96a]. Metric [McK99, LPR91b, McK95]. Metrics [GTW95, TV99]. Metropolis [MHdA90]. Micro [Ano98o]. Microcanonical [GL98, MF92, TP92, BMOT90]. microphysics [Git92b]. Microscopic [ABPSJ90b, BP97, CLD94, GJLL99, Jar99, Kaw98, SH95, SO92a, SO92b, WL92, WLC94, BPH94b, Cag90, DKS93, Mie90, OS91, PN94]. Microscopic-based [WL92, WLC94]. Microscopically [Cro98]. Microwave [SHW92]. Midwest [Ano93n]. Migration [BC94]. mineral [Fra94]. minima [BCP96, Mie93]. Minimal [BGL98, PK97, RHdS^{+}91, GS93c, Mür90]. Minimum [EK99, KD98, Mie93]. miscible [HR92a]. Misfit [FPL99, NFL99]. Misra [SAT94]. Mix [GJ99]. mixed [Per90]. Mixing [BCO99, BQ90b, AT90, DW91, Pol92, Yos97, Zha92]. mixing-length [Zha92]. mixture [MGS94, Per97, Pol90b, Pol91a, Sha95]. Mixtures [BC98b, BHP96, CFP91, Deu92, RBB95, SS90, WT93]. Mobile [DRbA99, TZ97]. Mobilization [WLC94]. Möbius [AE90, Mor90]. mode [Alb94, HS95a]. Model [Abd98, AEA97, Alb98, ABT99, AC99, Bak98, Bat98, BC99a, BES98, Bax98, BNK98, BHS99, BGM98, BRZ98, BDK99, Bov98a, BvEN99, BH98, CJK98, CL98, CGTM99, CBK99, DP97, DRbA99, ESB98, Fig92b, Fuj98, GW99, GS90a, GMR98, GR98, GS98, HL97b, HDS98, HY98, KS97a, Ken98, KE97, KRT99, Kin99, KNS98, KPS98, KD98, KW98, LY97, MRV99, MPR98, MP98a, NOZ99, NFL99, PPO99, PG99, PG00, Pri92, Raa98, SEW98, SER99, SS97c, SM99, ST99, SK99, Ser98, SW99b, Wat99, Zhi98, ZP99, dVOS98, dlSG99, vdH98, ADG96, AN91a, AK92, Aiz94, Alb94, ABHP90, All95, ABK94, AdAI92, ABPSJ91b, AAR92, BLS94, BM96a, BT91, BJ90, Bak93, Bak94, BZW92, BM96b, Bar90, BQ90a, Bax91, Bax93a, Bax93b, Bax96, BS90a, BB93, BSTV94a]. model [BSTV94b, BC92, Bel93b, BAK91, BMO95, BLL94, BT90a, BMO96, BFG93, BS90b, Bin92, BL94, BRZ95, BRZ96, BMHH97, BT95, dAB91a, BC96, BB94, BP93b, Bov90, BP91a, BG92, BK92c, BG93c, BG93d, BGP95b, BHJ92, BFB94, BG96, BT90c, BD93, BKJZJ93, BMOT90, BP90, BS95, BL90, CCRM95, Cer94, CGMS96, CT96a, CCD^{+}91, CD91a, Cho97, Cho92, CO96a, CLHS91, CDD94, CS91a, CS91b, Con96, CQ90, CC91c, Dan93, DKMM94, DG90, DW91, DDM92, DEM95, DHP96, Dev91a, DM94, EM95, EFGM95, FA91, FMP92, FP90, Fin92, FT90a, FT91, FT94a, FG94, FR96, Fra94, Fuj90a, Fuj90b, Fuj91, Fuj92, Fuj96, FMOU90, GPSS93, Gar91, Gay92, GSCK90, Gia91, GS91b, GPJ92, Gob92, GMM90, GH94, GK91, GS97, Gre90, GJL92, GUJ94]. model [GMMU97, Gry92, GR97, Häg96, HMY96, HZLD97, HP91, HH93, Hén92, Heu93, HH95a, HRS97, Hon96, HS95c, HS97, HH95b, HH96, HT91, HR92b, HPS94, IW93, Ino90, IT91, Iof94, JSC91, Jan94, Jez96, JK93, Joy90, KH96, Kar95, KS93a, KLT97, KS93b, Kho90, Kho91a, Kho91b, KRT97, KW93, KF90, Koh90, KS93c, KM97, KvL92, KT91, KLMR90, KO93, KO94, KL91, Kou90, Koz97, KA94, LOP91, LMM92, LM94a, LH94, LM94b, LH92, LP90b, LLH92, LPR91a, LPR91b, Maj93, Man90, MM96, Mar92a, Mar93a, Mar92b, MW90, MHdA90, MOS91b, MO93, Mar94c, MHJ94, MO96a, MGA97, ML91b, MBD97, MHB90, MOR94, MO96b, MNO97, MOT90, Mie90, Mie91, MD94, MZ96, MC92b, MS90b, MKP91, Mun92, Mür90, Nag95, NHT92, Nas91, NFID92, NS95]. model [NS96, Nol92, OBB95, OB96, Ord92, OP93, Oze93, PP93, PS94a, PPS95, PS91a, PST94, PS92a, PZ91, Pat93, Pat94, Pat96, PR94a, Pen95a, Pen91b, Per91, PV95, PV97, PTZG91, PR94b, PPR93, PS93b, Pok93b, Pol90a, PSP94, PS97c, PTN93, PN94, QdL92, Red94, RW91, RS92a, Rie93, RV97, RDWW93, RK93, Rut92, Sal95, SS96a, SS97b, Sam95, STAJ95, ST95, Sas92, ST92, SS96b, SO91, Sch94c, SS92b, SP91, Sen92, SMS96, SD95, Shi93a, Shu93, Sle96, Smi90, SC96, SCM96, Spe91, Sta94b, Ste97, SH89, Str95, SO92a, Süt96, SZ91, TC91, TBK90a, TBK90b, Tan92, TF92, TF93a, TF93b, TvROW96, TKD97, TZ97, TC93, Uen95, VZ92, VZ94, VZ95, WG96, War96a, War96b, Wee91, WPK95, YB91, Yam96, Yau94, ZHD95, ZHCD95, dABMR90]. model [dABR91, dME90, dMP91, dMPZ92, dO92, dOdOCS95, dHNR92b, dHNR94, vB90a, vD90, vHW93, vK95a]. modeled [COA95]. Modeling [AAH98, Ano99k, CCF99, DLPS99, EA98, FPL99, vdBJ91, CDPP90, RR97a, Tha95]. Models [AC98, AKMR98, ACC^{+}98, Ano99i, BC98a, BC98b, BC99b, Bou99, BG97b, BDS99, BHW99, BP98, CISS99, CCMS99, CGM^{+}98, DFF99, FV97, FS99, GZ98, KOJ98, Kül98, LW98a, LM98, LZ98, Mie98, Mie99, PS98, Pat98, SSLI97, SR93, SZ99, SI99, Uel99, Wag98, Whe99, WC98, Ahm96, AG91, Alb92, Ald93, AvBED97, ABL97, AYP95, BM92a, BM97, Bax95, BB92a, BPO96, Ben92a, BDM90, BDDM90, BLO97, BCPV97, BCK97, BKMS91, BMR95, BK96b, BZ97, BT93b, But93, CMVG95, CB94, CLY92, CS90b, CO97, CD91b, CP93, Cor95, CMPS97, DLM^{+}90, Eng91, FKV91, FNW92, FL94, FK94a, FJ93, FN95, GMCP96, Gar96, GM94, Gd96, GAA^{+}93, GKRT94, GS90c, GS93c, Gri94, Gri95, HM92a, HW97a, HS94a, HY96, HR92a, Ish95, IOT92, JD93, KS97b]. models [Kaw96, Ker93, KGM92, KF97, KW97, Kra94, Kül97, KB97b, LMR93, LF90, Lon93, Lov94, Mac95b, MG94, MR94a, ML91a, Maj94, Mar94a, MOS90b, Mat90, MGA95, Mon91a, Mon92, Mon97, Mou92, MHM94, Mül93, Nag95, Ném91, NR90, O’C93b, OL97, PS92b, PS93a, PB90, PO93, Pok93a, PSP94, PN97, PB95, PO95, PR90, QO95, RHA97, RK90, RT90, Ray91, RW92, SS97a, ST93, Spe93, Spo93, Spo95, Tas96, UO91, Wag95, WPSN94, WK97, Whi94, Wu96, YHHK96, dSCT91, vD97, vEAD90, vEAD91, Ano99j]. Modem [Opp98c]. Modern [Ano99c, Ano99i, Wid98, Bru94, Row97, Sta97c, Ano99d]. Modes [BGK98, FP90, SP93]. Modified [BK91, KOT98, GGM91, LLH92]. modular [KM93]. Modulated [CA93, Sas92, BSG91, BSTV94a, BSTV94b]. modulus [Dan93]. Molecular [GHPS96, GZ98, KW98, LV93, LLH92, Pet99, Pod98b, BHP97, BBC94b, NMC^{+}91, RRGT97, Sel97, Shl90, BHP94]. Molecule [GUJ94]. molecules [LC91, MC92b, RC97]. Moment [Bob97, Jun98, Lev96, Kru92, Wen97, Yan94]. Moments [BT90b, ZP99, MK93, MWA95]. momentum [Mar94b]. Monolayer [BCL^{+}99, UW91]. monomer [Jer90, KRS96]. monomer-dimer [Jer90, KRS96]. monotone [BFSV91, BK92b]. Monotonic [Ste99a]. Monotonicity [O’B90, AG91]. Monte [Alb92, ALLZ96, Bak93, BZW92, BM96b, BHJ92, CPS90, CC94b, Deu92, DJB98, DT93, EMHM95, FS99, FLB91, FS93, GAA^{+}93, GMM90, GP93, GB90, HMP96, Heu93, HCW96, JS95, Jes96, Kal91, KS97a, Kom93, KKBS92, KKBS93, KK94b, LS90a, MOS90a, MvR97, MC94, Mar92d, MHB90, MW95, MKZ92, NP99a, OB91a, OvRW96, PdO90, PBSR97, Rie93, RDWW93, ST97, SW99a, SBH92, Sel97, TBK90a, TBK90b, TvROW96, TFD90, Wag92, WB92, YB91, ZK93, dSCT91, Ano99e]. Moody [FT94b]. Moran [PW97]. Moran-like [PW97]. Moriya [AW90b]. Morphological [tL90]. Morse [ZPK97]. Moser [Cho97, HSK91]. Mostepanko [Pod98a]. Motion [BES98, DOPT93, ESB98, FL94, MS95b, Mol98, Tay99, TNN99, WG93, BJO97, BPH94b, BHP94, Che91, CF97, CDM93, CO96b, Dev91a, Dow91b, FMS97, Ger90, Got90, HL97a, IMS92, KK92, Lan95, McC95, MQ91, MZMQ90, Nob95, Pen91b, Spo93, Str95, Sza93, TC94, dMBD91]. motions [SGP90, vB91]. motivated [Gro95]. Motors [KW98]. movements [Num94]. moving [YS97]. MSA [HTPH93]. Multi [LW98a, Zha92]. multi-length-scale [Zha92]. Multi-Site [LW98a]. Multiattractor [Sos99]. multibaker [TG95]. Multibondic [CJK98]. Multicanonic [CJK98]. Multicanonical [Ber96, JS95]. Multicomponent [SD95, CL95, YP95]. Multidimensional [FM99, GK99, Bal92a, Bal92b, DZ94, ET90, Ers92, Lov94, Che94]. Multifractal [Ber97b, Ber99, CLT90a, CLT90b, HHD96, Hua97, Mol95, Mol98, MCC90, Por98, ABJM97, BMP92, DF93, Eyi95a, FT90b, FT90c, LP96a, LP96b, PW97, Rad93]. Multifractality [Ber97c, MV91, dB91]. multifractals [RS96]. Multigrid [BG97b, BGR94, BG96, BTY91, GP93, JS95]. Multilayer [BDDM90, DLM^{+}90, PBP97]. multineuron [AdAI92]. Multiparticle [PG99, PG00]. Multiple [CH92, BHP96, Sch93a, BPH94b]. multiple-binding [BHP96]. Multiple-time-scale [BPH94b]. multiplicative [BZW92]. Multiplicity [BK94a]. multipoint [Jez96]. multipolar [LSKB91]. Multipole [SL91, SL97b, SL97a, SP97]. Multiscale [KT99, O’C93b, PPO99]. Multiscaling [BLPP98]. multisite [Mon91a, Mon91b, Mon92]. multispeed [CP93]. multistate [BK94a, BBOC91]. Mutation [CBK99]. Mutual [Asl99, Li90]. mutually [LS90a]. my [Mac95a].

N [Pod98a, Sac98, Sta97a, UI96]. narrow [Niu91, RSGRP97]. narrow-band [Niu91]. Natural [BCCF92, GI92, TE95]. nature [DMP97, SP93]. Navier [BEM99, BP93b, CP97, DE96, EM94, ELM95, HL97b]. near [ACDD90, AC97a, AC97c, AC99, BKL97a, BG94, BDM90, BEO98, CBR95, CGMS96, FM91, For92, GGP92, HK96, Hal97, Ken93, KPS98, LK98, Lia91, LBK96, MS94, Pet99, RK90, Rei96a, Rei96b, RW91, SI99, VB93, dMBD91]. near-incompressible [Lia91]. near-to-the [BG94]. nearest [BS90c, Han95, KK91b, Mür90, Ray91, Rub91, Voo92]. nearest-neighbor [BS90c, Ray91]. nearest-particle [Han95]. Nearly [ACM95, AM92, Tas96]. necessary [Jes96]. Néel [PBP97]. Negative [ES93, JT98, LW93, Ole90]. Negative-temperature [ES93]. neighbor [BS90c, KvL92, Mür90, Per97, Ray91, Rub91, Voo92]. Neighborhood [dRIB99, UO91]. neighbors [LS92]. nematic [BS91a, JR91b, OR95b]. neocortical [MS93c]. nerve [TKG93]. nested [AF95b]. nets [FAKA97, KK93a]. network [ASKK95, BDM90, DB90b, DBB^{+}92, ML91a, MLM93, RR90, SO92a, SO92b]. Networks [BJS98, BJS99, DS99, LZ98, MTG99, AM94b, AdAI92, BKK^{+}92, BVZ93, BSVZ94, BSV94, BK94a, Bry96, DD91, Koh91b, KB90, LC95a, LAT95, PdO92, PHS^{+}97, SP93, Shi90b, TC91, Weh97a, dB92]. Neumann [Koj97]. Neural [BJS98, BJS99, MTG99, ASKK95, AdAI92, BZ90, BVZ93, BSVZ94, BSV94, Bry96, DD91, FAKA97, Koh91b, KB90, KK93a, LC95a, LAT95, PdO92, Shi90b, TC91, dB92]. neuron [Lon93]. neuronal [MS93c]. neurons [BK94a]. Neutral [AHR99, Ken98, LZ98, VdSFC97]. Neutrino [EK99]. Neutron [NMC^{+}91, SHG91]. Neveu [PPO99]. next [Per97]. nice [BK96b]. Nina [Hei94]. NMR [Guy91]. NN [CGTM99]. NNN [CGTM99]. No [CPP97, BK96b, IS99]. nodes [Cep91]. Noise [BAZ98, Ber98, CLLL98, CDMV98, DIK98, Git91c, Hor93, JS99, MLL90, PM91, BZW92, BBC^{+}94a, CCT92, CCT93, CCT96, CV93b, Dha97, DN94b, Eyi96, Ger93a, GW93a, HJZM93, KF93, KGM^{+}93, Kli91, MS96a, MS93c, MW90, MT94, NV93, NBM90, OPdlR95a, OPdlR95b, Rei96a, Rei96b, TDSR95, WB94b, ZR91].

Noise-induced [Hor93, MLL90, PM91]. Noisy [Kus99, Rei96a, Rei96b, CA93, GW93b, GW94c, HG92, KK92, SBP^{+}93]. Non [BCCP98, CH94, Don99, Has98, LW98b, LK98, NP99b, SM99, BP97, Bri96a, Fod97, GN93, IMS92, PS94b, Por96, Rub91, VZ94, WKWS95]. Non-Abelian [SM99]. non-associative [Fod97]. non-Coulombic [WKWS95]. non-Debye [GN93]. non-equilibrium [Por96]. Non-ergodicity [Don99]. non-Gaussian [VZ94]. Non-Hermitian [Has98]. non-locality [Bri96a]. Non-Markovian [CH94, BP97, IMS92]. non-nearest-neighbor [Rub91]. Non-Self-Similar [NP99b]. non-simply-connected [PS94b]. nonadditive [CFP91]. nonadiabatic [Yan94]. Nonanalytic [Tan98, HSK91]. nonattracting [Zha96a]. Nonautonomous [Kar99, SW98]. Noncharacteristic [Kło95]. Noncoincidence [Ami96]. nonconstant [CR97]. Nonconventional [DLM^{+}93a]. noncylindrical [TB92]. Nondegenerate [Con98a]. Nonequilibrium [Ano97f, BNK98, BJO97, CG99b, Cro98, CHG94, FR99, GSM98, KSZ97, MSS98b, NS99, RM93, Rue99, Wil91, BFG93, CL97, CE94, ED96, JSC91, JD93, KK93b, LO97, MWA95, Muk91, Opp94a, Opp94b, Rue96, Spo96, TG95, TS94, TKD97, YMHMJ93, Zim93, Zim94, bABD90, Ano99g, Ano99f]. nonergodicity [Pra94]. nonexistence [Xin93]. Nonglobal [BDIV97]. nonhomogeneous [Fig92b, Sza93]. Nonideal [PZ99]. Noninteger [Bis98]. nonintegrable [Mer92, TY96]. noninteracting [YK91]. Nonlinear [AG99, BNS92, Bro98, EPRB99, FL99, Git92c, IFR93, MGS94, McK99, Naw98, Opp94a, Opp94b, Paj95, PTN93, SB97b, SB97c, Ber97c, BMS97, DI93, EJ93a, ELM95, Ger93a, Ger93b, Git90b, Git91c, Git92a, GW93a, Git94, Git96a, GM88, GM96, KM91b, KPW91, LB96, MS93a, McK95, PB95, RK96, SB97a, SH95]. nonlinearities [MY94, Nos90]. nonlinearity [DLM^{+}93b]. nonlinearly [MC92a, WT93, vHW93]. Nonlocal [CPS90, CCO99, BCP96, DOPT93, Li92, Sni95]. Nonlocality [SAT94]. Nonmonotonic [BHW99]. nonnested [FST96]. nonorthogonal [SS94c]. Nonperiodic [GB94, vEMZ98]. Nonphysical [BC99b]. nonprimitive [VB90b]. nonresonant [dlL97]. nonreversible [Han96]. nonslip [HZLD97]. nonstationary [dO95]. Nonsymmetric [BB98, PR90]. Nontrivial [BR92, WR97]. Nonuniform [KP95, Per97, FP90, Zha91]. nonuniformly [LSV93]. Nonuniqueness [Wen99]. nonuniversal [KS93c]. Nonuniversality [MK93, MRS94]. norm [Ers94]. Normal [MV99, HS95a, KM91b, LOP96b, LOP96a]. Note [CdS91a, CdS91b, CG99b, GC92, Jes96, dHMP99, vK94, LV94, Mie93, Pol91b, Rus94, Tel90, Wu95, YZ92]. notion [DDG97, Omn91]. Novel [BCS^{+}91, MZ90, NAC91, Uen95]. November [LHF95]. Nucleation [Gun99, KPANG98, RK90, Bel93b, CO96a, CWP97, MOS91b, SH95]. null [SRC93]. Number [BNK98, CD99, JSA98, AZP97, BP93b, Cam91, DHP96, Dia94, FRHP95, GBP91, HS92a, KRS96, Kna93, Mad95, MC93, O’B90, Weh97b, vE90]. number-theoretic [Kna93]. Numbers [Ano01, MGD98, NV93, Weh97c, zzMZ96]. Numerical [ACG96, BAZ98, BCF90, CvD98, ED96, FdlL92a, FS93, Hay96, MT94, OBB95, QO95, ST99, BB91, BMO95, Bha90, BG93b, BvV95, CO97, DN97, Eva92, FR90, GK95, HKV91, KP92, MOR94, Rap92b, dlL92, BSTV94b, KPSW95].

O [FT94b]. Object [Cop98]. Object-Oriented [Cop98]. Observables [BJ98, GL98, TE95]. Observation [Lan94, VB93, KK91c]. observations [dlL92]. observed [BV94]. obstacle [WH95]. Obstacles [HS98, HLW99, DDG97, DDG96, VWG93]. obtained [FS93]. occupancy [Pae90]. occupied [KZ93]. occurrence [Gri95]. OCP [TF99]. octagonal [dGN97]. octionion [Fod97]. October [Ano92q, LHF95]. Odd [Avr98]. Odins [McC95]. Off [FMPP99, PWG97, Sew90]. Off-diagonal [PWG97, Sew90]. Off-Equilibrium [FMPP99]. Ohmic [LK98]. old [Cla91a, Phi94]. Oliveira [Wei95]. Olympic [RGdG97]. On-Site [LW98a]. One [Ano99g, Ano99j, AHR98, Asl99, Bak93, BAZ98, BF99, BCCP98, BHS99, BCL^{+}99, BRZ96, BMPZ98, BP98, CP99, DHS98a, DA99, DC98, DPS90, Has98, KM97, MP94, Mül93, Nag98, Rub91, ST90, ST99, SZ99, vdH98, ABP96, ABPSJ90a, ABPSJ90b, AE91, BJ90, Bar96, BG90a, BAP93, BCF90, CW96, CGK94, CT95, CMPS97, DDM92, DEM95, DHP96, DLS97, DMB97, Elo94, FA91, FM93a, FP90, GRZ90, GM88, GM96, Gre90, GUJ94, HW97a, HW97b, HRS97, HP97, HH90, Hor93, HS97, HJ90, Ino90, JMM87, JLM95, Joy90, Ker93, Koh92a, Krá96, KPWH95, KPSW95, Kru92, LBK96, MR94a, Mar92b, MP96, MBD97, Mie91, MF92, Mil91a, MC92b, Mon94, MP93, Mur94, OPS93, PPS95, Per90, Per97, Pol90a, PS93c, Pri92, QdL92, Rei96a, Rei96b, RW90]. one [RBF93, RS92c, SP95a, Sch93c, SD93, SMD92, Spe91, Sut92, Too95, TZ97, VB90b, YHHK96, YL96, Zeg94, dABMR90, dB91, vB91, vEM92, ST91]. One- [Asl99, MR94a, Sut92]. One-Component [Has98, JMM87, JLM95, RW90, SP95a]. One-Dimensional [Ano99j, AHR98, BAZ98, BF99, BCCP98, BHS99, BCL^{+}99, BMPZ98, BP98, CP99, DHS98a, DC98, Nag98, ST99, SZ99, vdH98, Bak93, BRZ96, DPS90, MP94, Mül93, Rub91, ST90, ABPSJ90a, ABPSJ90b, AE91, Bar96, BG90a, BAP93, BCF90, CGK94, CT95, CMPS97, DDM92, DEM95, DHP96, Elo94, FM93a, FP90, GRZ90, Gre90, GUJ94, HW97a, HW97b, HRS97, HP97, HH90, Hor93, Ino90, Joy90, Ker93, KPWH95, KPSW95, LBK96, Mar92b, MP96, MBD97, Mie91, MF92, MC92b, Mon94, OPS93, PPS95, Per90, Per97, Pol90a, PS93c, QdL92, Rei96a, Rei96b, RBF93, RS92c, Sch93c, SD93, Spe91, Too95, TZ97, YHHK96, YL96, Zeg94, dB91, vB91, ST91]. one-electron [CGK94]. One-particle [KM97, HJ90]. one-site [GM88, GM96]. one-state [dABMR90]. Onsager [LHF95, Ano95b, Ano95i, Bax95, Böt95, GJLL99, Gal96, Kau95, Kho90, Leb95a, Lyo95, Mac95a, Mil95, Nag95, Pen95b, RC97, Sha95, SH95, UI96, vK91]. onset [MdBM91]. onward [Bax95]. Open [GD99, DDM92, DEM95, Gas97, Ish96, Lie93, Muk91]. Operator [Ger99, Buo90, DP91, Ers94, Jus92, MQ91, MZMQ90, Pol91b, SY90, Sta92]. Operators [BW99, CDMV98, LPY98, Böt95, FK94c, FK94b, Hof93, Hof95, ILF92, Joy94, Sal95, VZ92]. opinion [KH96]. Optimal [BGR94, BG96, BG97b, Mar97]. Optimized [Deu92]. orbit [MR94c, RM97]. orbital [Kun94]. orbits [AC91, MSD92, Mer92, OdAdA96]. Order [AKR98, AHR98, Bax98, DP92, Hel98, KE97, WPSN94, YL96, ZFB98, vEMZ98, AE91, BK90, BI92, BK95b, BB94, Bry97, Cla91a, CCG90, CC91c, EM95, Fal92, Fil94, Gd96, HSW97, KY93a, KY93b, KP92, KK91b, KB97b, LAT95, Lop90, MG92, Mül93, PBP97, PS91a, PWG97, PR90, RV91, STAJ95, Sew90, Ste90, WB90, dOdOCS95]. order-two [Fal92]. Ordered [CHM99, NFL99, AEWF91, Rot93, RS92c, TV90]. Ordering [MP93, SD99, CNC94, Dow91a, PPD94, Sas92, Uen95, VR97b, bABD90]. Organization [AP99, GBCB99, Ray94, RL91]. organized [GS94, GK91, MKK90, NHT92]. Orientation [YT90]. Orientation-preserving [YT90]. orientational [Dow91a, VR97b]. Oriented [BBG98, Cop98, Koo95, BF96, Mil91b, NY95]. Origin [RB91, DGZ92b, Spo92]. origins [Mil95]. Ornstein [BVHP92, DMR97, DR98, GY98, KRT97, KRT99]. Orsay [Ano94a]. Orthogonal [Wid99]. Orthogonality [PO93, Per95b]. Orthorhombic [vdSE99]. Oscillated [BSdB^{+}98]. Oscillating [ABJM97, BB97, GHW91, Men92]. Oscillation [GR92, GP91, WCT91]. oscillations [EJ93a, WRJ95, ZR91]. Oscillator [DIK98, Gar98, Tat98, FR90, SGH93, TDSR95, WB94a, WB94b]. Oscillators [FM99, Ble90, Ble91, BNS92, BVR93, BJL^{+}91, Cra94, Dai90, Ger93a, Ger93b, KLR94, KF93, MS90a, NV93, SM91, vHW93]. oscillatory [MNO97]. osmosis [FFJ92]. Ostwald [AF99, LD98, NP99b]. other [DM90, FZ97, Fod97]. outgoing [GK95]. ovarian [LP91]. overall [Shi93b]. overbarrier [KG90]. Overcoming [SBH92]. Overlapping [Hel98]. overlaps [TF93a]. overshoots [CQ90]. Overview [Dow91a]. Oxford [Pod98a].

P [Git98, Sre98]. packings [LS90b, LSP91, MV91, SS94a, SL93b]. Pair [Lef99b, Maz98, PZ99, TC91, VWG93, BD94, BHP97, HTPH93, OR91, RST91, RW90, Ric97, SP95a, Vie94, Vie95, Mon92]. Pair-correlated [TC91]. Pair-Interaction [Lef99b]. pair-potential [BD94]. pair-reaction [Ric97]. pair-summable [OR91]. Paired [CLV98, CLV97]. Pairing [BCP98, BB95b, PP91]. pairs [KS97a]. Pairwise [BBD99, LD98]. Paper [GW99]. papers [Dom92, Dom97, Wei95]. parabolic [LBK96]. Paradigm [JVH98]. paradoxes [Gal90]. Parallel [Ano93o, Ano97g, BBG98, BJS98, BJS99, Ano90o, Ano91q, Ano92q, BF96, BVZ93, BTY91, DFZ94, HP97, MG94, ML92, PZ91, PdO90, SBH92, WK97]. paramagnetic [GMPS93]. Parameter [Ban99, TNN99, AT90, CR97, Fil94, FN97, KK91b, PS91a, WR97]. parameter-dependent [FN97]. Parameters [Bax98, KE97, MPdlR93, WPSN94]. Parametric [AKK99]. Parisi [DA99, FT93]. Parking [IS99, Kra92]. Part [AC97c, AC97d, KPWH95, KPSW95, Too94a]. Partial [AG99, BNRW93, GMH98, Dow91a, FdlL92b, MS93a]. partially [BOP94, DLM^{+}90, LMR93, OPB93, PB95]. Particle [BHK98, Ban99, Che98, FE98, Pet99, Zhi98, Bel93a, Bel95, BFSV91, BP97, BCF90, BL91a, BL91b, CC97, CDPP90, CR97, EM94, GAA97, Gia91, GL97, GW93b, HT90a, Han95, HRS97, HJ90, Koh91c, KM97, KK91b, LOP91, LMM92, NBM90, Niu91, OR91, PP91, PP95, Pri93, RM97, SL93a, Sza93, YIK95]. particle-conserving [Koh91c]. Particles [Asl99, BE99, BL99, CT98b, DRbA99, Don99, KT99, AL95b, AM94c, AEA93, Ber94a, BFG93, BLS91, BHP97, CdB93, Cor95, Dal97, FM93a, FJ92, FSB91, Geo95, HL97a, LhBBS97, Lan94, MGS94, OAB^{+}96, Pen91a, Pow91, Ros93, Sch95b, SLSA91, SCM96, Str97, TH96, vDL95]. Particlewise [SK99]. particular [CCRM95]. Partition [BGP95a, BL93, Gol99, RS92c, Bha90, HWvB97, MRC95, Smi90, Zyl90]. Partitions [Kin99, AZP97]. Partridge [Jan94]. Passage [BRT98, BAP93, Bin92, KG90, LK95, Men92, TDSR95, YL96, vK93]. Passive [BGK98, CK98, GK99, Ber94b, CLT90b, KWG96, LFtH91, MS95b]. past [PTZG91, VWG93, WH95]. Pasta [ACM95, CG99a]. Path [Gol99, HS94b, AZ96, BG96, Che97b, DDG97, DDG96, Kho90, RHdS^{+}91, SR90]. Pathological [vEFK95]. pathologies [HK96, MO93, MO95, vEFS93]. pathology [RW91]. pathology-free [RW91]. Paths [DLB^{+}98, BK92a, GMN94]. pathways [Fer94]. Patrascioiu [Aiz94]. Pattern [ASKK95, ABK94, Rot93, Cla91b, GP91, Mac93]. Pattern-specific [ASKK95]. Patterns [BSdB^{+}98, IdRB98, JG98, LMN98, MTG99, MdBM91, SL95, WM98, BWK91, BGP95b, CLD94, CPPG91, Gay92, Git92d, TC91, Wal91, vB91]. pausing [HW90]. Pawl [MS98a]. Peierls [AKR98, BGM98, LM98, PVZ94]. Peierls-Fröhlich [PVZ94]. penalty [Ken94]. percolated [Uen95]. percolating [SR90]. Percolation [AM94a, BP94b, BRT98, BGL98, CCMS99, DLB^{+}98, Gan91, HWD97, JSA98, KNS98, LK98, RSW98, Too99, dMM98, AN91b, AG91, AEA93, BK92a, BDM90, BK95a, Ber97b, BG93b, Bra93, BTY91, CPP97, Cho92, CT95, DM90, DB90b, DR91, GS93a, GLM95, GS90b, Gra95, GM97b, HKV91, HA97, HCW96, JK93, KIKK97, KS93b, KZ93, KS92, LPPSA92, LC91, LS92, MRS94, MHB90, MM93, MS94, NY95, Num94, Oku90, Orz96, Pae90, PS92a, PS92b, PS93a, PR94a, Pin94, PS91c, PHS^{+}97, Rom90, RCB90, RB90, SA91, SR95, Sch90a, Sch92, Ste95a, VDH97, WL95, Wu97a, YZ92, vEAD90, vEAD91]. percolation-like [Sch90a]. percolative [GS95, ORG91]. Percus [BT90b, GGM91, HSW97]. Perfect [PS97b, Rob91]. performs [EN92a]. perimeter [Mer90, SF91]. period [GM88, GM96, dSVG90]. Periodic [BFPR99, MR94c, PAB^{+}93, Pom93a, SBZ98, TNN99, TZ97, AC91, BRR96, Ble92, Bur91, Che91, Che94, CT96b, CdB93, DDG97, DDG96, Fig93, FK94c, Fig94, For90a, For91, FM93b, GW94c, IMS92, KF93, MM97b, Mer92, OdAdA96, RM97, Smi94, SY90, TDSR95, WG93, Xin93]. periodically [Com90, GW93a, GW93b, GDJH93, PX91]. periodically-forced [GW93a]. periodicity [BT94, PM91]. permeability [Koh91a, MS94]. Permutative [MM98]. Perron [Bec95]. Persistence [SS99]. Persistent [Ali99, BGL99, Mar93a, DHP96, SCM96]. perspective [Bru94]. Perturbation [BG90a, CDMV98, FST96, GSH90, TS94, Wie98b, BCS^{+}91, Bob93, FdlL92a, JD93, LM92, RW96b]. Perturbations [BE98, FMPP99, Liu98, DFF96, GO95, KMKdC96, MKZ92, SHW94]. perturbative [Hon96, Lia91]. perturbed [BCPV97, IM96, PNT91, SL95]. Phase [AM95, AEGL92, ACC^{+}98, Ano99b, AORZ95, AHR98, BZ95, BG97a, BC99a, BRZ98, CCO99, CCF99, CKK99, DFF99, DIK98, FMS97, FN95, FPL99, FG99, GMR98, GR98, GL97, Git90a, HLIM93, JK99, JV99, KH96, Ken98, LMR93, LMP99, LPW92, MSS98b, Mon91a, Mon92, MGD98, NE95, NV98, Oze93, PS92b, PS93a, PPD94, RBB95, SD99, SD93, SSZ99, Shi92, Tót90, VZ92, Whe99, YRHMJ92, dABMR90, vDL95, AC92, ALLZ96, All95, dAB91a, BK90, BK95b, BZ97, BV94, Cag90, CNC94, Cao93, CCST90, CGMS96, Cho94, Cla91a, CCG90, CC91c, Dai90, DFF96, DDM90, EM95, EP97, Fog92, FT94a, Fos93, FG94, FdH94, FZ91, Gia91, GLM95, Gob92, GdH91b, Gry92, HMV90, HPS94, JR94, Ken90, Ker93, KY93a, KY93b, Koh90, KSZ97, KO94, Kou90, Lit92]. phase [Lop90, MN96, MM96, Mar90a, MKP90, Mar92b, Mar94c, MHB90, Mon91b, MSG95, MG92, Mül93, OR95a, Orz96, PS94a, Pat93, PS92c, Pri92, RS96, RBF93, SS97a, Sch95a, Sch93a, ST90, ST91, SPR91, Ste95b, TDSR95, TFD90, UO91, Uen95, WL95, YB91, ZP94, vEdH91]. Phase-Field [CCF99]. Phase-Separating [MSS98b]. phase-space [MSG95]. Phases [BMSW99, Kna93, BSG91, BK94a, CH92, FjLL95, FLL95, HR92b, LM94a, OO91, PW90, Sas92]. phenomena [ACLS94, BO91a, Fam92, FT90b, FT90c, Git90a, KN93, MOR97, Rap90, Row97, RK93, Too94a, Too94b, Wid98]. Phenomenological [FR99]. Phenomenology [Li92, FT90b]. phenomenon [FK94b, PS93b]. Philosophical [Sas95b]. philosophy [Bru94]. phononic [Mor92a]. phonons [CB94, LM94a, Niu91]. phosphatidylcholine [NE95]. Photonic [Fig94]. Phys [Ano01]. Physical [Ano99k, SACB98, MHB90]. physically [CC94a]. Physicist [Kob97]. physicists [Git94]. Physics [Ano91n, Ano93o, Ano96j, Ano97g, Ano97h, Ano98a, Ano98q, Ano99a, Ano99c, Ano99l, Ano99-30, Ano99-31, Cop98, McQ97, Sas95b, SM99, Ano90o, Ano94i, Ano97e, Bru94, BQ90b, EB95b, Fod97, JR91a, Opp95, Opp97, Pea95, Pod95a, Pod96, Por96, RZ93, Wei93b, Ano90n, Bry98, Git98, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano90i, Ano90j, Ano90k, Ano90l, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91b, Ano91h, Ano91i, Ano91j, Ano91k, Ano91l, Ano91m, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano92j, Ano92k, Ano92l, Ano92m, Ano92n, Ano92o, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano93i, Ano93j, Ano93k, Ano93l, Ano93m, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f].

Physics [Ano94g, Ano94h, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano97a, Ano97b, Ano97c, Ano97d, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano98n, Ano99e, Ano99o, Ano99p, Ano99q, Ano99r, Ano99s, Ano99t, Ano99u, Ano99v, Ano99w, Ano99x, Ano99y, Ano99z, Ano99-27, Git92b]. physiology [Smi95]. picture [RS94, SO91]. pictures [Zeg90]. pieces [BK92b]. Piecewise [BO99a, Jus98, Ste99a, AZP97, Che92, Che93, Din96, Liv95, Lov94, LBK96, PM91]. piecewise-expanding [AZP97]. pile [Man90]. Pinchoff [BBW98]. pining [GKRT94]. Pinning [DMRR92, OYSK91]. pipe [RSGRP97]. Pirogov [KLMR90]. pitchfork [OPdlR95b]. Pixton [Wei95]. places [Cla91a]. Planar [Zhi98, BHJ92, CDM93, Mik95, Str95]. Planck [Bec95, CS92, Gou97, LY97, Pet99, SBP^{+}93, WB94a, WB94b]. plane [BO91a, BF91, CC94b, PW94]. planes [AM95, DFZ94, HLIM93, Wu95]. plans [BB95a]. Plaquette [CGTM99]. Plasma [AC99, Has98, For90a, FJ92, For92, FJ96, GY93, JMM87, MC94, RZ93, RS92c, SP95a, Sut92, JLM95]. Plasmas [VB98, AC97d, FLS95, Nd97, RW90, RC97]. Platoon [SK99]. Pleiotropy [CBK99]. plus [FJ96, KM98]. Point [AZ98, BBD99, NOV99, SZ98, SZ99, SC99, SK98, Wid98, dRIB99, AC97a, AC97c, AS91b, BMO95, BL94, BS95, CO91, CBR95, CGMS96, CB90, CT95, CG93, FM93a, FT90b, FT90c, GPS90, HS95b, Ken90, Ken93, KP91b, Koz97, Leb97e, MvR97, MLM93, OP95, RW91, Row97, RCB90, TFD90, WR97, WB90, YIK95, vD97]. Points [AH98, Naw98, PRW99, AG91, BSG91, BR92, DT93, GGP92, HS90b, JW97, LM94b, ML91a, PPW94, PS92c, Sas92, Sch90a, Voo92]. poisoning [MS92]. Poisson [CM98, CP99, FAKA97, KP91b, KKBS92, KKBS93, KK94b, Muc96]. Poissonian [Der97, HLW99, IM96]. Pol [DIK98]. Polaritons [HQSS94]. polarization [FRHP95]. polarized [AN91b]. polaronic [AAR92, BM96a]. Pollak [FP98]. Pollak-Grabert [FP98]. polluted [GM97b]. polydisperse [SS94a]. Polygonal [HS98]. Polygons [SM99, EG90, Gut96, Mad95]. Polymer [Ano99i, BG97a, BES98, Ebe96, ESB98, KS97c, PS99b, Whi94, BD94, BKV93, Cao93, CB90, CdHM91, CO96b, Deu92, Fer94, FT94a, JR91a, KN93, KC91b, OB91b, RBB95, RCB90, zzMZ96]. polymerization [Don95]. Polymers [BNO98, BP91b, JK99, Mül99, Rei98, Vuj99, BO90, CJ97, CD90, CD91b, DJ95, Dou92, Dow91b, ED92b, FVY92, Fos93, GC90, HS92a, Jaf91, Kho91a, Koo95, MZ90, MK93, Mil91b, NFID92, NMC^{+}91, OB91a, OS96b, Piz97, SZ96]. Polynomial [BMO96, DO97, BD97, SF91, War96b, ZXZY94]. polynomials [BBL96, FLS95, Mer90, MBF^{+}97, SY90]. poor [RW90]. Population [CBK99, GdH91b, GdH91a, BNS92, BVR93, Ray94, SM91]. populations [Cra94, Dai90, PdO92]. porous [BKW90, BS91b, CM95, FIS96, GS91b, Koh91a, ML91a, Zha92]. portrait [HC92, RBF93]. position [HCW96, SR95, vEFS93]. position-space [HCW96, SR95, vEFS93]. positional [Dow91a]. Positive [SGH93, vEMZ98, Wei95]. Positivity [Bis98, Rue96, Rue97]. possessing [Koz97]. Possible [BB94]. Potassium [LY97]. Potential [Don99, NK99, Naw98, SBZ98, AZ96, BD94, BGM97, Bra91, CCT96, Che91, DP91, DF93, DMRR92, For90b, IMS92, MS93b, NS95, OR91, Sni95, Sut92, TB95, WG93, vLH97]. Potentials [BI99, BP98, CMP97, PZ99, ABCP96, CMP95, DE96, Dou92, FM93b, RC97]. Potts [AYP95, BC98a, BM96b, Bax91, Bax93a, Bax93b, Bax96, Bax98, BS90a, BVZ93, dAB91a, BKMS91, BC96, CISS99, CJK98, DKMM94, Dav91, DHP96, DJB98, FMP92, FS99, FT90a, FT91, FMOU90, Gay92, GH94, Gri94, Hon96, KLT97, KM93, Ked93, Kho90, Kho91a, Kin99, KLMR90, Mat90, MO96a, ML91b, OBB95, OB96, SS97a, SS97b, SS97c, SS98, Whi94, WPK95, dABMR90, dABR91, dOdOCS95]. Potts-spin [FT90a, FT91]. Powders [Sta98, MND92]. Power [IT91, MKP90, MK91, BCFM95, Kom93, LF93, Mar90a]. Power-law [MKP90, MK91, Mar90a]. pp [Sre98]. practical [BB92b]. Prandtl [BP93b, CD99]. preasymptotic [CHG94]. precipitation [Kar95]. precipitation/dissolution [Kar95]. Precision [FS99, Koh90, LS90a]. Predicting [GW99]. Prediction [PM99]. predictions [BM96b, Dow91b, LS90a]. Predictive [EA98]. Preface [AKL91, Ano96i, BL97, BHM^{+}93, GL94, LR92, Leb94a, Leb97a, LST98]. Preliminaries [Fer94]. Preliminary [SHG91]. prepared [Cla91b]. presence [BT90d, FdlL92a, GHW91, GS95, OTH92, Rub91, Rue97, WMS90, WH91, YK91]. presented [Bal91]. Preserving [GD99, BCCF92, CC91b, Gas97, YT90]. Press [Bry98, Git98, Opp98b, Pod98b, Pod98a, Rap98, Sac98, Sre98]. Pressure [Mer99, Bav94, GY93, Smi94]. Prigogine [SAT94]. primary [TKG93]. principal [GP91]. Principle [BQ90b, BP98, CLLL98, Che99a, Lef99a, Gal97, GS97, Mür90, PR94b, Rob91, dHNR92a]. Principles [BLPP98, Ber90a, Ber90b, Pod96]. Probabilistic [Bra93, DN94a, Lov94, AEGL92, BKM93, Kot95, LMS90, PS90]. probabilities [LPPSA92, Mil92, Mol97, OS95, OS96a, Pri94]. Probability [DDGZ97, JR91a, Kus99, TF93a, TV99, ABP96, AC95, Bru94, CC91c, GTW95, GM88, GM96, GH97, KK91a, KIKK97, Man93, PER95c, Red94, Ric97, Wu97a]. Problem [BS98, BE99, PM99, Sam99, Sos99, BL94, CO96b, CD90, CD91b, Dal97, FLS96, GS91a, Kom93, LP91, LS91a, LS92, MS96a, Nev95, OS95, OS96a, Phi94, SS96b, WT92]. Problems [BAZ98, Hei98, SS99, Bob97, BMP90, Cag90, CW96, FLS95, GK95, IS96, Opp95, WMTR90, vdBV93]. procedure [ILF92]. Procedures [RSSS98, BP94a]. Process [Ano94a, AG98, BW98a, BBD99, DMR97, DR98, DGLS98, GJ99, RSSS98, Sal99, dMdO98, Alb95, ACJL92, Che97a, CM92, DM90, DJLS93, DLS97, FFK94, GPSS93, GHW91, Hor93, JL94, KK91a, Mac97, MM93, SP95b, Sch94b, Sch96a, Sch93c, SD93, Sch97b, Sim96, SO92a, Swe97, WL95]. Processes [AKK99, Che99a, Gui99, Jar99, Nag98, Sch98, Bel93b, BKL97b, Ber97c, Bre91, BO91b, Che90, CdS91a, CdS91b, DdH96, Don95, ED96, EN93, Ger90, Ger93a, GN93, Gri94, Han96, JK93, KP91b, MJ90, Por96, Sch95b, Sch97a, Shl90, SO92b, Wei91b, Wei93b, WMTR90, Wu97b, YL96, ZSP90, vWL95]. processing [EIK92, SR93]. Product [CDFG97, GK96, HP97]. Production [BC99c, EPRB99, EK99, GD99, LFvW98, CC92, CC94a, Gas97, Han96, Mür90, Rue96, Rue97, Wen97]. products [Hua97, Mar93b, Pey91]. profile [BP93a]. Profiles [BMSW99, BFSV91, DJLS93, DLS97, Pat94]. Program [Ano90m, Ano90n, Ano91o, Ano91p, Ano92p, Ano93n, Ano96j, Ano97f, Ano97g, Ano97h, Ano98q, Ano98p, Ano99-30, Ano99-28, Ano99-29, Leb90, Leb91, Leb92a, Leb92b, Leb93a, Leb93b, Leb94b, Leb94c, Leb95b, Leb95c, Leb96, Leb97b, Leb97c, Leb97d, Leb98]. Programming [Cop98]. Progressive [RGdG97]. Projection [Jus92, DZ94]. projections [Rad93]. projecton [LV94]. Projector [KK92]. Proof [KP94b, MS98b, Reb98, Uen95, BF91, DEJ92, MN96, Wag92, War96b]. Propagating [FR97]. Propagation [Bro98, GBCB99, ACLS94, KS97b, KC91a, PANG^{+}95, Xin93]. propagator [Kho91a]. Proper [Con98a]. Properties [BF98, BDIV97, CBR95, DF98, FL99, For94, GL98, Gui99, KKBS92, KKBS93, RT90, TNN99, VB98, AZ96, AEGL92, Bal92b, BSTV94b, BA93, Ber94b, Ble92, BBC94b, BBOC91, BT92, Che92, Che93, Che94, CK91, Com90, DI93, DF93, ET90, FMR94, Fig92a, FVY92, FT90b, FT90c, GAA97, Ge91, GGM91, GPJ92, GH94, GPS90, Guy91, HRS97, HKSY96, Jez96, KK94b, LP91, LSV93, LS97, LSK91, LPR91a, LPR91b, LS90b, LSP91, LR96, Mar95, MMSR92, Mie90, Muc96, OP90, Oze93, PY95, PS91b, RW92, Rus94, SGP90, SZ95, THK^{+}91, TTPH91, Vai92, WH91, vEFS93]. properties-the [GAA97]. Property [Bak98, KE97, Mae99, Bak93, Bak94, Koz97, SS94c]. Protein [MSS98a, Lie93]. Proton [NLT93, MM90a]. Pruning [BV94]. Pruning-induced [BV94]. pseudogaps [Fig94]. pseudointegrability [Dag96]. pulses [WG93]. pumping [Mel93]. Pure [NS99, TF93a, TKG93, WHF92, vE90, Mon91a]. pure-tone [TKG93]. purity [BRZ95].

quadratic [FT94b, SY90]. qualitative [Mar92c]. quantization [Sze96]. quantized [HS94b]. Quantum [Abd98, AKR98, AKMR98, Ano96k, Ano99b, AC99, ADE98, AZ98, BJ98, BO99a, BO99b, BM92b, BBL96, Bri96a, DDGZ97, DGZ92a, DGZ92b, FM99, FV97, FL99, Gar98, Git93a, Git93b, Git97, GW94d, HLW99, Ish96, Kar99, LFvW98, LPY98, MS99, Mar92c, Mom96, Niu91, Opp98b, Par91, SBB98, WBG98, ZFB98, vWL95, AC97d, AS91b, AC90a, Ami96, APT94, Ano92p, BKL97a, BG90a, BL94, BK94b, BJL^{+}91, Buo90, CdOW95, Che91, CJ92, Com90, Com91a, DFF96, Dev91b, DK96, ELMD^{+}90, Edi93, FKV91, For93, Fri90, FT94b, FST97, GS90a, GSH90, GS93c, GB94, GGP92, Joy96, KN93, Kar94, KG95, Kaw96, KT94, KM96, Krá97, LB94, Lan94, LR92, Lia91, Los90a, Los90b, LVY92, Luc93, Mat90, MQ91, MZMQ90, Mey96, Mom94, MVZ97, O’C93b, Omn91, PY95]. quantum [RW96a, RBGW92, Sam95, SHW94, Sni90, Sto97a, VZ92, VZ95, ZK93, dO95, vK95a, vKO97, Sac98]. Quartic [Naw98]. Quasi [MS99, MKK97, TNN99, Git92d, MP93, WC95]. Quasi-Bound [MKK97]. quasi-lattices [WC95]. quasi-one-dimensional [MP93]. Quasi-Periodic [TNN99]. quasi-regular [Git92d]. Quasiaverage [Pat93]. Quasiclassical [Pra98, ELMD^{+}90]. quasicrystal [LPW92]. Quasicrystalline [Mie97, Mie90]. Quasicrystals [Mie99, Rad99, Bur91]. Quasilocality [Lef99b, Häg96, PV95]. Quasiparticle [VB98, BG90a]. Quasiperiodic [BF98, DF98, WC98, BJL92, GGL90, Hof96, JK93, Koh92a, SHW94]. Quasipotentials [HG92]. Quasistationary [Bob95]. Quenched [AC98, ABT99, BJ99, Mül99, Mun92, Ser98, TF93b, ABPSJ90b, ABPSJ91b, Ebe96, LM91, WA90]. questions [HKV91]. queueing [BKK^{+}92].

R [Git98, SO92a]. Radial [YS93, GR92]. Radiation [EK99, FR99, RS94, Sha95]. Radiative [BFPR99]. radiophysics [Ber90a, Ber90b]. Ramified [Rei98]. Random [ACC^{+}98, Ali99, Ano94a, Ano01, AH98, ABPSJ91a, APC^{+}92, BW99, BGL99, BZ99, Bis98, BRZ98, Bra91, CPP94, CT98a, CP99, CS91a, CvD98, CR94, DT93, Dou97, EN92b, FP92, FMPP99, Gar98, GKRT94, GBCB99, Häg98, JW98, Kho91b, KRT99, Klo98, Kül97, Leb99, Liu98, Maj94, MP98a, MS91, Per91, PV97, Ram95, RRT98, SM99, Shi98, Sin99, TST91, TW98, WX97, Wei96, dMPZ92, dHS92, dHMP99, AN91a, AZ95, AZ96, Ald93, AM94a, AM94b, AK91, ABPSJ90a, ABPSJ90b, ABPSJ91b, AC95, AEA93, BF97, BA93, Ben92a, Ber90a, Ber90b, BRZ96, BD97, BBL96, BC96, BO90, BP91a, BK92c, BK93, BG93d, BK96b, BL91a, BBOC91, BS91b, BH91, BP91b, CB90, Com91b, CD90, CD91b, DdH96, DI93, DC94, DM90, Don95, Dou95, Dou92, DB90b]. random [DBB^{+}92, EMHM95, EN92a, Eva92, ED92b, FP97, FZ97, Fig92a, FK94a, FK94b, FN97, GK96, GBN92, GS92, GMN94, GS90b, GS91b, GPS90, GdH91b, GdH91a, Gri94, Gri95, GZ97, Häg96, HW97b, Hav90, Hof96, HS92c, IW93, KNV93, KWG96, KPW91, Kot95, Kou90, KP91b, Kra92, LM91, LB96, LPW92, LS90b, LSP91, MG96, Mar93b, MW91, MW94, MBF^{+}97, Mil92, MH92, NS95, Nos93, Oer95, OS96b, Orz96, PP93, PPS95, PS97a, Per93, PV95, Piz97, PSP94, PANG^{+}95, Rad93, Ram93, Red94, Rie93, Rom90, Rot93, Rub91, Rue97, SS94a, SGP90, SA95, SP94, SZ96, Tel90, TKD97, VR97b, WC96, WA90, Weh97a, Weh97c, WMS90, WH91, Zeg94, ZM93, dGN97, dMP91, dMPS95, dHNS92, dHNR92b, dHNR94, vWH97]. Random-bond [CR94]. Random-Cluster [Bis98, Häg98, PV97, Gri94, Gri95, Häg96]. random-field [BRZ96]. Random-matrix [Per91]. Random-random [APC^{+}92, ABPSJ90a, ABPSJ90b, ABPSJ91b]. random-resistor [DB90b, DBB^{+}92]. random-site [VR97b]. Randomly [LR96, FVY92, IM96, MM90b, Mil92]. Randomness [CK99b, DGZ92a, Lin93]. Range [AKR98, AP99, BE98, BL99, IRB^{+}99, LMP99, Lef99b, Mer99, SZ99, VHR98, vEMZ98, ABL97, BKL97b, BT90c, CPP94, CS92, Con90, GL97, Jez96, Ker93, KB97a, Kło95, KT91, LS97, Mae90, Maj93, MGJ92, NP94, OPS93, Pet90, PWG97, PR90, PS91c, RK90, Sew90, vB90a]. Ranged [GSM98, ML91b]. Rapaport [Pod98b]. Rapidly [GJ99]. rarefied [CC94b, Sha95]. rat [CV96]. Ratchet [KW98, MS98a]. Rate [BC99c, CCO99, Zha96a, CC92, Ger90, LOP96b, LOP96a, NBM90, RB91, Sch96b]. Rates [DRbA99, MV97, Pen95c, PBSR97, TP92]. ratio [Ble90, Ble91, Dha97, FR95]. Rational [BGM98, Hay99, TY96]. ratios [LMS95]. ray [STAJ95]. Rayleigh [BTT98, ET90]. reactants [LSKB91]. Reacting [DRbA99]. Reaction [BB98, DRbA99, IdRB98, NMHS99, OB91b, PX91, Sch95b, SS99, dRIB99, dB91, Alb92, Alb95, ABP96, BLS91, Che90, CLHS91, CKS91, DW91, DBH91, GMM90, Kho90, KK91c, Koz94, Koz96, KPWH95, KPSW95, LC95b, Par91, QO95, RM93, Ric97, RR97b, Sch96b, THK^{+}91, Tor91, TP92, Xin93, Yan94, ZKB91a, bABD90, vD90, LC97]. Reaction-diffusion [Sch95b, Che90, GMM90, Koz96, KPWH95, KPSW95, RR97b, THK^{+}91, Xin93, vD90]. Reactions [SK98, BL91b, BB97, CSB97, KK93b, LSK91, LSKB91, OO91, Pri93, PCG95, RS92b, SZ91]. reactive [Git90a, PN96]. Real [GMTB95, PR94a, DN94b, HA97, MBF^{+}97, NV93, Nos90]. Real-space [GMTB95, HA97]. real-world [Nos90]. realistic [BD94, CC94a]. realization [STAJ95]. reciprocity [Gal96, Sha95]. recirculating [NGB95]. recombination [Nd97]. reconstructed [HR92b]. rectangle [dGN97]. rectangle-triangle [dGN97]. rectangles [FZ97]. rectangular [OBB95, WG93]. Recurrence [BNN97, BT92, DP91, For93, HSK91, Sim94c, TNN99, dHMP99]. Recurrent [Com91a, MW94, Tel90]. recursions [Ber96]. Red [MR94b]. reduced [LMM92]. Reduction [NM99, WB99, BW88, CK95]. redundant [ILF92]. Reentrant [Mik95]. reference [PPR93]. refined [Eyi95b]. Reflection [Bis98, AM94d, DS92]. reformulation [Ord92]. Regime [EML98, GM95, HT91, PS97c, Sch95a]. Regimes [WCT91, GW94a]. region [Bax96, CGMS96, IFR93, tLZ96, LLM95, MM96, Mar94c, Smi90]. regular [Git92d, JW97]. Regularity [Por98, vEFS93, Hol94]. Regularization [LFvW98]. Reinhard [Git98]. reinterpretation [Sni91]. Related [Ano99-31, Shi98, AM94d, Koh92a, tLZ96, MW91, Nev95, Ste91, TT94]. Relation [BR91, Whe99, Wid99, BB93, BCF90, But93, Che91, HH95b, RW90, SP95a, SH95]. Relations [Bax98, LK98, Wat99, FLS95, KW93, Kru92, Sch97a, Sha95, vK91]. Relationship [DI93, JW93]. relationships [GBN92]. Relative [CFL98, Gol99, LFvW98]. Relativistic [CK99a, DMR97, DR98, VB98, DEJ92, LSS97]. relativity [Bri96a]. Relaxation [AMF98, Hen97, KT99, LB94, Oer95, Pol90b, WWW95, dMdO98, vCEBS94, DR93, GN93, KG90, TDSR95, Too95, WK97, Xu95, YMHMJ93, KE97, Pol91a]. relaxations [CB94]. Relaxed [Yos97]. relevance [Jag91]. remanent [RS91b]. Remark [Aiz94, Sam99, AB92, Ben95b, Hof95, LM96, MR94a, SZ96, Wei95]. Remarks [DDG97, HH95b, Mar92a, Hof93, MO93, dO95]. remembered [Pen95b]. removal [SS92b]. Renewal [Iso99]. Renormalization [BCO99, Bry98, CGT99, CJB99, CO97, Col98, GRZ90, Git96b, GGP92, HW97a, LC95b, LC99, LIF92, OSE93, PRW99, Sch92, Sco93, Sle98, Wie97, dHS98b, AZ95, Bov90, BP91a, DH92, Ebe96, Gal90, GMTB95, GMTB96, HK96, HA97, HCW96, ILF92, Ken93, LM94b, Maj93, MO93, MO95, MNO97, Nay93, O’C93a, PR94a, Per95b, PPW94, SR95, VZL97, YS97, dlL92, vEFS93, vEFK95, LC97]. Renormalization-Group [BCO99, CJB99, PRW99, HA97, MO93, MO95, PPW94, YS97, vEFS93, vEFK95]. Renormalized [BCO99, BT95, Lef99a, Wie97, Sal95]. Renormalizing [FZ97]. Repeller [LPR91a, LPR91b]. repellers [AT90]. replica [Koh90, PST94]. replica-symmetric [PST94]. replication [Rus94]. Reponse [Dor94]. Report [Ram95]. representability [HJ90]. representation [GN93, Hen97, O’C93b, PO93, PV97]. Representations [CMR98, CS91a, Spe97]. reproduction [ZSP90]. Reptation [BES98, ESB98]. repulsion [ABP96, Tót94]. repulsive [Tan94]. requirement [Vol94]. Research [Ano92a, Ano99i, Böt95]. reservoir [MPdlR93]. resistance [IT91, Zha96b]. Resistant [Sch99]. resistivity [Mat94]. resistor [AM94b, DB90b, DBB^{+}92, Weh97a]. resolved [KK91c]. Resonance [Wol92, ANS93, BG93a, DLM^{+}93a, DLM^{+}93b, FGMA93, FM93b, GMPS93, HJZM93, ILD93, KGM^{+}93, Lon93, MS93c, Nic93, NNM93, RS97b, VB93]. Resonances [ACLS94, Ger90, Hal97]. Resonant [Ger93b, Mel93]. resonating [BS90c, YHHK96]. resonating-valence [BS90c]. resonating-valence-bond [YHHK96]. Response [BL90, Eu94, FMPP99, GKT93, ZP93, CS90b, DLM^{+}93b, Ger93b, KS90, Lin93, RS97b, TKG93]. rest [Cor95]. Restricted [Hio90, BK95a, EN93, FJT96, MW90]. Restriction [MRV99]. Restrictions [Mon91b]. Restructuring [ZK98]. result [AZ95, MS92]. Results [BEO98, CK98, Ste99b, BMA93, BM92a, BSTV94a, BSTV94b, BMO95, BLL94, BT90a, BSG95, BG93c, BKV93, CM96, CW95, Cut91a, Cut91b, DDM90, DDJ^{+}95, For90a, FJM92, FJ96, FT90c, GS90c, Gut96, HKV91, HS95a, JL94, Ken93, KPWH95, LP96a, MM96, MF91, OBB95, Piz97, Pri92, Sin91, SCM96, SHG91, Tor91]. retardations [Göt96]. Retrieval [BSVZ94, BSV94, BK94a, GSCK90]. Return [Hay99, CH92, Sim94c, WK90]. reversibility [LV93, SH95]. Reversible [Cro98, RS91a, SZ91, CH94, Don95, Han95, Hor93, KH94, OS95, RB94, RRGT97, TG95]. Review [Ano95a, Ano99c, Ano99d, Ano99g, Ano99i, Ano99h, Ano99f, Ano99b, Ano99j, Ano99e, Ano99k, Ano99l, BSB97, BM99, Bry98, Dom97, Git98, Opp98a, Opp98b, Opp98c, Pod98b, Pod98a, Ram93, Rap98, Sac98, Sre98, Wid98, Adl93a, Adl93b, AEWF91, Ben92b, Ben93, Ber90a, Ber90b, Dom92, Dor93, Eng91, Eu94, Fam92, GM92, Gan91, Git90b, Git91b, Git91c, Git92b, Git92c, Git92a, Git92d, GS93b, Git93a, Git93b, Leb92a, Mas92, Muk91, Nos93, OG91, Opp91, Rap92b, Ros93, Shl90, Shl91, Spo92, Wei91a, Wei91b, Wei93a, Wei93b, Wes91, dB92, vEdH91, vdBV93]. Reviews [Ano99m, AW90a, Nos90]. revised [CC91a]. revisited [AYP95, CHK93, CLT90a, Jan95, SS97b, VGC92, Zha91, dSM92, vWL95]. Reynolds [Hay92]. RG [dOdOdSB95]. Rheology [BS91a, Pow91]. ribbon [OvR95]. ribbons [OvRW96, vROS^{+}96]. Riccati [CP93]. Riccati-coupled [CP93]. rigid [FSB91, SL95]. rigid-disk [SL95]. rigidity [KM91a, KS92]. Rigorous [BG92, BG93c, GS90c, KD90, MP94, NS95, PPS95, SBZ98, BK90, BP91a, dAB91b, FdlL92b, Ken93, Mad95, Mür90, Pes93]. Ring [AHR99, BvV95, CGK94, GM97a, LhBBS97, Sch93c]. ring-bounded [LhBBS97]. Rings [ADE98, Mar93a]. Ripening [LD98, AF99, NP99b]. River [CGM^{+}98]. RNA [Fer98]. Robertson [BP94a]. Rod [SER99, Zha91]. rodlike [BS91a, Pow91]. rods [GM97a, KR92, Mur94]. Role [RLK98, TNN99, CPPG91, Ger90, GGM91, Mil95]. roll [BWK91]. root [VdSFC97]. roots [MBF^{+}97]. Rosenau [Sle98]. rotated [Fuj96, LP90b]. Rotating [Sta98, Got90]. Rotation [Kli92, NV93]. rotational [CF97]. rotations [BB91, BF91]. Rotators [BT94, Zhi98]. Rothman [Rap98]. rotor [Rei93]. Rough [BDK99, Blu94, MM90b]. roughening [ANHKV93, HMP96, PANG^{+}95]. round [BPO96]. Row [PB90, Fuj90b]. row-row [Fuj90b]. Rowlinson [BHW99]. Ruch [BQ90b]. Ruelle [Reb98]. Ruggedness [AMF98]. Ruijsenaars [FV97]. Rule [BCP98, EN92a, Gác90, Gal90, Ken93, Ken97]. rules [Jur95, Nas91]. Run [Mar94a, GMCP96]. Runge [MZMQ90]. Running [Wie97, RW96b].

S [Opp98b, Rap98]. Sailor [BGL99]. Salesman [PM99]. Salsburg [GS90b]. same [dSCT91]. Sample [SS92a, AZ96, SS94b]. Sample-to-sample [SS92a]. samples [PER95c]. sampling [BB95a, FLB91]. sand [BP96, GPSS93, Man90]. Sander [Mar92a]. sandpile [CFJ91, Pri94, Rus93, Spe93]. Sandpiles [MN99, PK97, Kru92]. saturable [FGMA93]. Saturation [Sle98]. sawtooth [Vai92]. Scalar [CK98, GK99, KT99, NOV99, Ber94b, CLT90b, LFtH91, MS95b, NR90, SA91, WR97]. scalar-field [NR90]. Scalars [Con98b]. Scale [Ano90n, BPH94b, EP92, ES93, FIS96, Fil94, ILF90, KB97a, Lan94, Man90, Mar90b, MM93, OYSK91, Ros93, Zha92]. scales [Ber94b, PO93, Per95b, RR97a, SP91]. Scaling [ANHKV93, ANV94, BE94, BLPP98, BV96, CC97, Con98b, DJB98, DLB^{+}98, GBN92, Ge91, GLR98, KPANG98, LW98b, LSK91, MS97, MS94, PSR98, PS97b, PM99, PANG^{+}95, RK93, SS97c, Alb92, Bax96, BBF94, BL94, BK90, BKMS91, BI92, BOP94, BT90d, BT90c, CR94, CMPS97, DN90, DOPT93, DR91, FOS94, FMOU90, GS93a, KK94a, KP91a, KM96, KPWH95, KPSW95, KP94b, LS91a, LF90, MS96a, Mar92a, MZ90, MG92, Mor92b, PR90, Rap92b, RW92, SA91, Smi90, Tan92, TE95, dSVG90, Wan96b, Wei91a, dOdOCS95, Bry98]. Scalings [QO95]. Scars [OdAdA96, TY96]. Scatterers [Fel98, BDG97, GPS90, RSL90, SS94c]. Scattering [BE99, Nos93, Sac98, BV94, Gas92, LhBBS97, NMC^{+}91, RVW96, STAJ95, SHG91]. scenery [dHNS92]. Scheme [CJB99, vdSE99, Fer94]. Schemes [LZ98, KP92, MQ91, MZMQ90, Xu95]. Schlögl [BT95, PN97]. Schrödinger [DP91, Hen94, Hof93, Hof95, Koh92a, Naw98, Shl91]. Schulman [Opp98b]. science [Ben95a, Git96a, Sta97c]. Scientific [Ano99c, Dom92, Dom97, Joh90]. Scope [vEFS93]. screened [RC97]. Screening [AC99, Jan95, RC97, BR92, BK94b, KK93b, Rue90]. search [Eve93]. Second [HSW97, Gd96, GKT93, Pen94b, Swe97, Ano96j, Ram95]. Second-order [HSW97, Gd96]. secular [Pol94]. seen [JMP94, JT96]. Segment [BES98, ESB98]. Segregation [Sta98, ALLZ96, FG94, GHPS96, GL97, KK91c]. Seiler [Aiz94]. seismic [BC94]. Selection [CBK99, Cla91b]. Self [Ano99h, ABL97, BBG98, BBW98, BMPZ98, Con98a, EKLR94, FE98, GIT91a, HKV91, KRT99, LW98a, LM91, LPY98, NP99b, Noo98, Pen94a, Ray94, RRT98, SA95, VB98, BF96, Bar90, BKV93, CPS90, CPS91, CPS92, DF93, FVY92, GS94, GK91, GB90, HSS93, HSS95, HA97, JFK91a, JFK91b, Jur95, Ken94, KRT97, LMS95, MOS90a, MKK90, MP96, Mol97, NHT92, NDF92, O’B90, OPB93, PS91a, PW94, RR93, RB90, SS94b, SA94, SSV93, TF93a, TvROW96, Tót94, Wu95, bABD90, vR97a, AM94b]. Self-Adjointness [LPY98]. Self-affine [HKV91, PW94]. self-averaging [PS91a, AM94b]. Self-Avoiding [Ano99h, BBG98, Noo98, RRT98, LM91, Pen94a, SA95, BF96, Bar90, BKV93, CPS90, CPS91, CPS92, GB90, HA97, Ken94, LMS95, MOS90a, NDF92, O’B90, RR93, RB90, SA94, SSV93, TvROW96, Tót94, Wu95, vR97a]. self-avoiding-walk [HSS93, HSS95]. Self-Consistent [FE98, KRT99, VB98, EKLR94, KRT97]. self-correlations [TF93a]. Self-diffusion [ABL97, GIT91a, JFK91a, JFK91b, SS94b]. Self-Gravitating [BMPZ98, MP96]. self-interacting [FVY92, OPB93]. self-ordering [bABD90]. Self-organization [Ray94]. self-organized [GS94, GK91, MKK90, NHT92]. Self-Similar [BBW98, Con98a, LW98a, DF93, Jur95, Mol97]. Semi [BRT98, Che95]. semi-detailed [Che95]. Semi-Directed [BRT98]. Semiclassical [BR99, Mer92, CJ92, MSD92, OTH92, Sch93b]. semiconductors [ADG96, Cer94, Pla90]. Semidilute [Mül99]. Semiflexible [CJ97, Kho91a]. semigroup [SAT94]. Semiinfinite [CP99]. semipermeable [ET90]. Semistiff [BG97a]. separable [ACM95, LC95a]. separated [Koz96, THK^{+}91]. Separating [MSS98b]. Separation [FPL99, Ken98, AORZ95, CNC94, Gia91, Gob92, HPS94, OR95a, Pat93, Pri92, RBB95, YB91, YRHMJ92]. Sequence [WBG98, ZPK97]. Sequences [Iso99, MSS98a, CS90a, Com91b, Pey91]. Sequential [Ano94a, Ram95, BE94, EN92b, FP92, Kra92, Per93, Ram93, SP94, TST91]. Series [GAA^{+}93, KS93c, AMA^{+}90, BBM92, CLV97, CLS90, Gon94, KIKK97, LB96, MOR97, NR90, Pol90a, Pol94, Pom93b, PPNM97, Sal95, CLV98]. sessile [WF91]. Set [Tan98, AE90, Mol95, Sch94b, Ste90]. Sets [Liu98, AM94a, ABJM97, DM90, Gác90, Ish95, MW91, Orz96, PW94, TV90, Zha96a]. Seven [GW94a]. several [HR92b]. Shape [GS95, DFZ94, Fuj92, Kru97, MHL92, RV97]. Shapes [KO94, KF90, KP94a]. sharp [Cag90, TC94]. Shastry [Ino90]. Shear [CL97, DP97, KW93, MSG97, RK96, Wil91]. sheared [OR95a, PIM94]. Shearing [SHG91, Maj94]. sheet [Orz96]. Shell [AM94c, KOJ98]. Sherrington [Bov98a, PS91a, Pat96]. Shift [DGLS98, DO97, CT96a, Voo92]. Shlosman [Yos97]. Shock [ACJL92, DJLS93, DLS97, BO91a, CP93, FS93, Spe97]. shock-wave [FS93]. Shocks [DGLS98]. Short [AP99, BL99, MS98b, MC92b, MHM94, vK93, CPP94, Con90, DR93, OTT92]. Short- [BL99]. short-distance [OTT92]. Short-Range [AP99, CPP94, Con90]. Short-time [MC92b, DR93]. Shortest [DLB^{+}98]. should [vEM92]. showing [KGM^{+}93, Mae90]. Shuler [dH94]. sided [MS91]. Sierpinski [Erc97, HY96, Sim94c, Jez96]. Sierpiński-Gasket [Jez96]. Sign [vES98]. signal [Dha97, GW94c]. signal-to-noise [Dha97]. signals [AGL91, BMA93]. Significance [LZ98]. Similar [BBW98, Con98a, LW98a, NP99b, DF93, Jur95, Mol97]. similarity [CP93, Eyi95b, Ish95, Sai95]. Simple [DGLS98, Kei98, Luo97, Sin99, Tat98, Too95, ACJL92, ABL97, BA93, CDM91, DJLS93, DLS97, FMOU90, HG92, HZLD97, JL94, KF97, KSZ97, MY94, Nas91, SZ91, WHF92, ZR91, dSCT91, vB90a]. Simplified [KW98, Kar94]. simply [PS94b, ZP93]. Simulated [Fri90, ST97, BKW90]. Simulating [ABF^{+}95, KS90, dOP93, Koh92b, TQGO95]. Simulation [Ano92a, BS91b, Gon94, JSO99, MSS98b, MS90b, Pod98b, AL95b, BP94b, BHJ92, CC94b, FR90, GHPS96, ILD93, Koh91b, MS93b, Man90, OB91a, PdO90, PB94, QdL92, Rap92b, Rie93, SP94, SZ91, Tor91, Wag92, WLC94, YB91]. Simulations [CvD98, SW99a, BM96b, BAKK^{+}90, BDM90, BvV95, DN97, Deu92, FFJ92, GL93, KG95, KFK91, MC94, OP95, QO95, RS91a, SS94b, SBH92, Sch93b, Smi94, VWG93, WH95, Wil91, WL92, Zie93, BES98, Ano99b]. Sinai [Gar97, KLMR90, OMM93, Ste99b]. sine [McK95, McK99]. sine-Gordon [McK95, McK99]. Single [BG97a, TBK90a, TBK90b, Alb95, Eve93, Hor93, Lie93, Sch96a, Tan92]. Single-cluster [TBK90a]. single-component [Alb95]. single-species [Hor93]. single-step [Tan92]. Singular [KG99, LW98b, NMHS99, Kło95, LOP96b, LOP96a, RZ93]. Singularities [Con98a, ABJM97, BM97, Mar95, Ray91]. Singularity [Alb98, BMA93, VdSFC97]. sinusoidal [CP97]. Sinusoidally [DIK98]. Site [LW98a, BG93b, Dua90, GM88, GM96, Han96, HCW96, KRT97, LS92, Sch96a, Sch92, TLW91, VDH97, VR97b, WHF92, FRHP95]. site-bond [Sch92]. site-diluted [KRT97, WHF92]. site-to-bond [Dua90]. sites [MD97]. Six [Alb98, Fuj96, Wat99, BS90a, BS95, JR94, KL91, LP90b, Nol92]. Six-Vertex [Alb98, Wat99, Fuj96, BS90a, BS95, KL91, LP90b, Nol92]. Size [DJB98, JFK91a, JFK91b, PS97b, PM99, SS97c, TF99, TNN99, AS91a, Alb92, BHP97, dAB91a, BK90, BKMS91, BI92, BK95b, BT90d, BT90c, BD93, CT96a, CD91b, CR94, CMPS97, Dan93, FT93, For91, FG94, FOS94, FMOU90, GHPS96, GS93a, HS92a, JMP94, KT94, KM96, KT91, KPWH95, KPSW95, KK94b, tLZ96, MG92, Mor92b, PR90, Rap92b, Rap92a, Smi90, Wan96b, Wei91a, dOdOCS95, vEAD90, vEAD91]. skew [Lan95]. skewed [Bax93a]. slab [Cer96]. slabs [Dev91b]. slider [NHT92, RK93]. slider-block [NHT92, RK93]. sliding [RB91]. slip [AS95]. slit [vK94]. Slope [PSR98]. Slow [BGK98, KPANG98, Aiz94]. Slowing [GSCK90, dRIB99]. Slowly [NK99]. Slyozov [Vel98]. Small [AKMR98, HS98, Yos98, AM94c, Ber94b, BP93b, FdlL92a, HA97, Lem95, OS95, OS96a, PER95c, RR97a, RBF93, Vol94]. small-cell [HA97]. small-field [Lem95]. Small-Mass [AKMR98]. smallest [Dia94]. smectic [MdBM91]. Smoluchowski [Bar96, DRbA99, GW94b, KP94b]. Smooth [Don99, GK99, Mar92b, RLK98, Rue99, KLRT97]. Snider [LM90]. Snurnikov [Mon04]. Sober [BGL99]. Sobolev [Lem95, LPY98, Wan96a]. SOC [BCK97]. Social [Gal90]. Solid [FG99, Pat98, Pet99, AL95b, CDD94, CDG95, CDFG97, MW90, Mor92a, PB90]. Solid-on-Solid [Pat98, CDD94, CDG95, CDFG97, PB90]. solidification [Cla91b]. solitonic [EJ93a]. soluble [Pen91a, PTN93, SD93, Spo95, VZ92, vK95a]. solute [Kar95]. Solution [AKK99, BMP90, ERS99, GY98, TV99, BVHP92, BB95b, BHP96, CLHS91, DDM92, DJLS93, FS93, GS91a, HS90b, MT94, OP93, PST94, Phi94, Ric97, Sch93c, Sch97b, War96a, War96b, dGN97]. Solutions [AG99, BC99a, CM98, DLPS99, EML98, KOT98, LW98b, Luo97, Naw98, PS98, RZ93, SER99, Wen99, AC90b, ACI91, BS91a, CdC94, Cer96, CdHM91, CP93, ELM95, GTW95, GW94b, HZLD97, LOP96b, LOP96a, LLM95, Pal90, Pet90, Pet93, PCG95, Sai95, Sai96, SS92b, Sin91, ZHD95]. Solvable [Ano99j, AC99, Bax95, BB92a, Cer94, Gas92, HM92a, MS90b, OvR95, Shi90a, vD97]. solvation [FRHP95]. solve [MQ91, MZMQ90]. solved [PS97c]. Solvent [KS97c, LBT97]. solvents [OB91a]. solver [Nad95]. Solving [BS90b, FLS95]. Some [BMO95, Cut91a, Cut91b, ETW98, Gal97, Gui99, Hof93, Ken93, Kru92, Lef99a, LPS94, MM96, MO93, Piz97, Shi98, Ste99b, THK^{+}91, WH91, Wei92, dO95, AZ96, Bal92b, CK91, FJT96, GLM95, JM96, Niw97, O’C93b, PW94, RRGT97, Tor91, vEAD90, vEAD91]. SOS [BK96b, CM96, DDM90, DM94, LM96, OP93, vES98]. SOS-type [DDM90]. Space [Ban99, BC99a, BS98, CK99b, KK91c, NV98, AT90, BP94a, Dal97, Fer94, Fog92, GMTB95, GK95, Gra94, HA97, HCW96, Lan94, LM96, Lit92, MV91, MSG95, OMM93, PR94a, Per97, SR95, vEFS93]. Space-and [KK91c]. space-filling [MV91]. spaces [Eyi95a]. Spacing [TW98, Ble90, Ble91]. Spanning [BGL98]. sparse [CD90, Eva92]. Spatial [AHR99, BL91b, DS99, LO97, RL91, vD90, BE94, BRR96, CFJ91, EB95a, PN96, PN97, SBH92, vB90a]. Spatially [Lu99, NMHS99, Bob97]. Spatiotemporal [AGL91, AL95c, ACLS94]. Special [Sch94c]. species [Ber97a, EFGM95, FFK94, FG94, Hor93, RL91, Ric97, SP94, SO92a, SO92b]. Specific [CG99a, ASKK95, WHF92]. spectra [AGL92, LF93, LMM92, MSZZ90, PB90]. Spectral [CCM99, Com90, DIK98, OD96, SBZ98, TKG93, ZPK97, dlL97]. Spectrum [Por98, Wei99, Zhi98, BMA93, BKL97a, BCFM95, BJL92, Cho97, DMP97, Fig93, FK94c, FK94b, GK96, GKC94, Hon96, IMS92, Joy94, Pes93, Pok93a, Por90, Sim94a, Sim94b, Sta92]. Speed [Kei98, WX97, CR97]. speeds [MS96b]. sphere [BT90b, BPH94b, BHP94, CFP91, FJM92, FJ96, GG94b, HSW97, HT90b, PTZG91, Phi91, RV91, dMBD91]. Spheres [BZ99, BDS97, GMO91, GGM91, LSP91, MC93, RST91, SS94a, SS94b, Sch95a, Ste91, SL93b]. Spherical [CT98b, All95, Ber94a, BR91, BB95b, BHP96, BT90c, BD93, CT96a, Dan93, DDG97, DDG96, HS97, Pat93, Pat94, Phi91, Rut92, SMD92, VBF97]. Spherically [BW98b, Fel98]. spike [BZ90]. Spin [AC98, Bak98, BP98, Che98, IS99, Ked93, KE97, LPY98, Maz98, MKK97, RdO94, ST99, SZ99, VHR98, ZP99, BM92a, Ben91, BMO95, BL93, BCF97, Bry96, CCC^{+}90, CCST90, CHK93, CM95, CCG90, CMPS97, DDJ^{+}95, Dou92, FT90a, FT91, FR95, FdH94, GI92, GS93c, GZ97, Han96, HS91, HS92b, HP91, HS97, JR94, JM96, KG95, Kaw96, Ken90, Kli92, Kna93, KM96, Koz97, LZ91, LM97, MZ96, Mon91b, MGJ92, NS96, OL97, Par93, PY95, PBP97, PS92b, PS93a, Pok93a, Pra94, RS92a, RS91b, Sas92, SSP95, Sch94c, Shi93b, SDJ^{+}96, Str95, TF92, TF93b, Wan96a, Zeg90, Zeg94, ZP93, dOdOdSB95, vE90, vE96]. Spin- [KE97, MKK97, GS93c, PBP97, Pok93a, dOdOdSB95]. spin-boson [Sch94c]. spin-echo [LZ91, Shi93b]. spin-flip [Han96, Str95]. Spin-glass [RdO94, CCST90, NS96]. spin-glasses [vE90]. spin-one [HS97]. sping [Gro95]. spinless [GJL92]. Spinodal [AH98, SW99a, LOP91, PB94, RK90, Ray91]. Spins [AKR98, AKMR98, GMH98, Ser98, DHP96]. Spontaneous [AHR99, EFGM95, AL95a]. Spot [MR94b]. Spouge [Jeo99]. Spread [Sch99, Pen94a]. spread-out [Pen94a]. Spreading [HWD97, HDS98, Alb95, Bag96, GPJ92, Gra95, MHdA90, Sta94b]. Spurious [BHP97]. Square [FS99, SS98, VdSFC97, Vuj99, ZP99, ABHP90, BG90b, EG90, HS90b, JFK91a, JFK91b, KIKK97, Ken97, MD97, Sal95, WC95, dO92]. Square-Lattice [SS98, Sal95]. squares [KP91a]. SRB [DKKP96, Jia99, Wei92]. Stability [AC98, ADE98, BBW98, BP91a, BK92c, CCO99, IdRB98, LSS97, Lu99, PS98, PRW99, SM91, Tas96, BG93a, BSG95, BNS92, BMS97, CC92, DN94b, Fer96, GW90, Hio90, MLL90, MG95, Pod97, SHW94, dO95, DFF96]. stability-instability [Hio90]. Stabilization [MGA95]. stabilized [BBC^{+}94a, LVY92]. Stable [Ber98, Mie97, JW97, JW93]. Stacking [PO95, STAJ95, FT90a, FT91]. stadium [OdAdA96]. Staggered [BEK91, Das95]. staircase [AFNB97]. standard [KLMR90]. Star [BB93, HH95b, OB91a]. Star-triangle [BB93, HH95b]. Stat [Ano01]. State [AC98, AE99, CJK98, DJB98, EK99, FG99, Koj97, PZ99, SS97c, SS98, BM96b, Ben92a, BRZ95, BS90c, BMP90, BVZ93, dAB91a, CW96, CdOW95, CSB97, DKMM94, DLM^{+}90, ELMD^{+}90, For94, FMOU90, Gay92, GW93b, GH94, Gro95, HH90, KM93, Ked93, Kli91, Koh91b, KSZ97, KLMR90, LMR93, LS97, Mar97, McK94, NS96, OBB95, PBP97, Pok93b, Pok93a, Ray94, SS97b, Tan92, TP92, Uen95, YHHK96, dABMR90]. States [AKMR98, AHR98, AHR99, DO97, GR98, GMR99, KE97, Klo98, Mie97, MKK97, MGD98, NS99, AK92, AAR92, Bal91, BG93a, Ber92, BRZ96, BCF97, Bov90, BGP95b, BMS97, BE92, CO91, CW96, CL97, DFF96, DI93, DDJ^{+}95, ES93, FNW92, Fri94, GC95, GW95, GM88, GM96, GK95, GKRT94, Gra94, Gra90, GJL92, GMMU97, HRS97, HP97, LMM92, MS91, Mel93, MR96, Mom94, NAC91, OPS93, Par93, PY94, PY95, RHA97, Sch94c, SDJ^{+}96, TG95, TF93a, Weh97b, dMPZ92, dMPS95, vE90]. Static [KMKdC96, LM94a, Niu91, OSE93]. Statics [bABD90]. Stationary [Abd98, Ali99, AHR99, BC99a, BE99, CL97, DP97, Fri94, GW95, IdRB98, SW91, dMdO98, AKV94, DE96, ELM95, GC95, HRS97, Kli91, TG95, TS94]. Statiscal [Ano91b, Ano91n]. Statistical [AEWF91, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano90i, Ano90j, Ano90k, Ano90l, Ano90m, Ano90n, Ano90o, Ano90p, Ano91a, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano91i, Ano91j, Ano91k, Ano91l, Ano91m, Ano91o, Ano91q, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano92j, Ano92k, Ano92l, Ano92m, Ano92n, Ano92o, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano93i, Ano93j, Ano93k, Ano93l, Ano93m, Ano93o, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95c, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96j, Ano97a, Ano97b, Ano97c, Ano97d, Ano97f, Ano97h, Ano98a, Ano98b]. Statistical [Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano98n, Ano98q, Ano98p, Ano99a, Ano99g, Ano99f, Ano99e, Ano99l, Ano99o, Ano99p, Ano99q, Ano99r, Ano99s, Ano99t, Ano99u, Ano99v, Ano99w, Ano99x, Ano99y, Ano99z, Ano99-27, Ano99-30, Ano99-28, Ano99-29, Ano99-31, Ble92, Bry98, Che94, CG99b, DN97, Eng91, EB95b, FMR94, FLB91, Fra94, GMN94, GG94b, KWG96, LSV93, Leb90, LMS90, Leb91, Leb92a, Leb92b, Leb93a, Leb93b, Leb94b, Leb94c, Leb95b, Leb95c, Leb96, Leb97b, Leb97c, Leb98, Mar95, McK95, MR94b, MR96, Opp91, Opp96a, Opp98c, Pea95, Pod95a, Pod95b, Pod97, PG99, PS99a, Rue99, SM99, Shi90a, SCBIR99, Spo96, Sza97, Tai95, TTPH91, TNN99, dB92]. statistical [Ano92b, Ano94i, Ano97e, ACI91, AL95c, Bal92b, BP93a, Bax95, BSTV94b, Ber90a, Ber90b, BMR95, Bri96b, BQ90b, Che92, Che93, CHG94, Edi93, Fer94, Git92b, Git93a, Git97, HH96, Joh90, KY93b, Leb97d, Mac95b, Muk91, O’C93b, OS91, Opp95, Pom93b, Por96, Por97, Rap92b, RR97a, Rue96, Sas95b, SGP90, Str97, dMPS95, vdBV93, Ano92q, Ano97g, McK99, PG00, Opp98a]. Statistical-thermodynamic [Shi90a]. Statistically [CK98]. Statistics [BGL99, Ber99, BMPZ98, CO91, CGK94, CG93, Dev91b, DS99, GK99, MD97, Zie91, APT94, BNN97, BK91, BBM96, BL94, BP96, EK96, GMCP96, Gil90, Kho91a, KC91b, Lie93, MS93a, Mar94a, Mar90b, OPdlR95a, OPdlR95b, PS97a, YL96]. Steady [AHR98, AE99, BCCP98, CSB97, EK99, OMM93, Tan92, Fer96, Fig92a, GW94a, IS96, Ray94]. Steady-State [AE99, Tan92]. stem [MS93c]. Step [Gui99, Koh91b, MS95a, Pat96, Tan92]. sticks [GIT91a]. Sticky [Ste91, YS93]. Stiffness [vE90]. stimuli [TKG93]. stirring [Pei95]. Stochastic [ANS93, BF99, BEM99, Bob93, BvEN99, BB98, CDMV98, DS99, DIK98, DLM^{+}93b, FM93b, GMPS93, JL98, JSO99, KS97b, LS99, Lon93, NNM93, PM99, Shi90b, TDSR95, Tat98, Wei93b, Zhi98, dMdO98, BZW92, Bel93b, BSG95, BC95, BAP93, Bin92, Bre91, BB92b, CDPP90, CSRPS93, CGMS96, CGS95, DI93, Dim90, DLM^{+}93a, FIS96, FGMA93, Gar91, HJZM93, IM96, JW97, JLS96, KGM^{+}93, KP92, KM97, LC95a, MPdlR93, MHL92, Man93, Mar93a, MOS90b, Nic93, Par91, Pla90, Por96, RS92a, RS97b, Sam95, Sch92, Spo93, Spo96, VB93, WK90, ZKB91a, dHNR92a]. stochastically [BCPV97]. stochasticity [ACM95]. Stokes [BEM99, BP93b, CP97, DE96, EM94, ELM95, HL97b, PS93b]. stop [ABK94]. storage [BG92, DD91]. story [Fis94]. straining [CJ97]. stratified [AM92]. Stress [JV99, vVBE93]. Stress-stress [vVBE93]. stresses [Hay92]. Stretched [IOT92]. Stretched-exponential [IOT92]. Strict [AG91, CC92]. strictly [Hay96]. String [Raa98, Pen95a]. strip [Mac97]. stripes [Rot93]. strips [EN92b]. Strong [Ano01, BCO99, BO91b, Zeg94, DC94, KP92, LP90a, OS96b, Pet93, Pok93b, Ste97, Weh97c]. strong-coupling [Ste97]. Strongly [LC99, NOZ99, BC92, Hen94, RC97, SB97a, YHHK96]. Structural [CB94, IdRB98, All95, MG92]. Structure [BFPR99, BNO98, CS90a, CT98b, DGLS98, Fer98, Fin92, Jag91, MGD98, Pri94, BRZ96, BL91b, CLT90a, CLT90b, DM90, FK94c, Fil94, Hof96, KS93a, Lie93, LPR91a, LPR91b, MVZ97, NS96, OTT92, Par93, PS92b, PS93a, dHNR94]. structured [Sch92]. Structures [Sre98, BM96a, BLL90, EP92, Fig93, Fig94, JSC91, JM90, McQ97, REK91, SS94a, Ano99d]. studied [STAJ95]. Studies [ST99, ALLZ96, ED96, KPWH95, KPSW95, NMC^{+}91, SHG91]. Study [FS99, JSO99, JLS96, RDWW93, SACB98, TF99, BG93b, BB92b, CO97, Eva92, GAA^{+}93, GMM90, GGP92, HMP96, Heu93, HA97, Koh90, KS93c, KKBS92, KKBS93, KK94b, LMM92, LC95b, LC97, MvR97, MHdA90, MHB90, MOR94, MKZ92, MG92, OPdlR95b, Pol90a, PS91c, RR90, TvROW96]. Subcritical [Don95, RHH91, VB93]. Subdiffusive [KD98]. subdynamics [CE94]. Subgraph [dME90]. subject [GW93a]. subjected [NAC91]. subordinated [MJ90]. subshifts [CH94]. subspaces [KM97, dlL97]. substitution [ZXZY94]. substitutional [CS90a]. substitutive [Pey91]. substrate [CKS91]. Substrates [BDK99]. Subsystem [BO99b]. Subtleties [tLZ96]. Suggest [MSS98a]. sum [Mie93]. summable [OR91]. sums [Joy90, LS91b, SL97b]. Super [PS97b]. Super-Instantons [PS97b]. superconducting [MR94a, vLH97]. superconductivity [Git97]. Superconductor [GMR98]. Superconductors [GMR99]. superconformal [BMO96]. Superdegenerate [TFD90]. Superdiffusion [AM92]. Superexponential [MG95, Too95]. Superfluid [Leg98, Cho92, McC95]. superlattices [Mik95]. superparamagnets [RS97b]. Superpositions [RS96, Sto97b]. Superstability [Reb98]. Superstable [BBD99, KY93b]. Superstates [Kül98]. superstrong [Dow91b]. superstructure [SBH92]. supersymmetry [Zim93, Zim94]. Supplementation [SF91]. support [vWH97]. supposed [GK91]. sure [DN94b, Häg96, PV95]. Surface [ABCP96, BBW98, BK95b, For92, JT98, KLMR90, Kun94, MOT90, MS95a, PB94, SSLI97, Wei99, YMHMJ93, ZK98, AS91a, AN91a, ANV94, Alb92, AS95, BF95, BG90a, BT90a, CM96, CM95, CLHS91, CWP97, Fam96, FST96, FT90a, FT91, HR92b, KM91a, MMSR92, OB91a, OP93, PS97c, TC94]. Surface-directed [PB94]. Surface-driven [YMHMJ93]. Surface-induced [BK95b]. surface-reconstructed [HR92b]. Surfaces [DDGZ97, Pet99, SD99, Vuj99, FST96, GS90c, HS90b, MM90b, MS91, MS92, Pod95b, vR97a]. surfactant [SPR91]. Surrogate [FRHP95]. Survey [Gut96, Gil90, Git92a, Opp96b]. Survival [ABP96, Mil92, AC95, Jan94, KK91a, NB90, Red94]. Susceptibility [DG90, JW93, Jur95]. suspended [AL95b, BPH94b, BHP94, MdBM91]. suspension [PTZG91]. suspensions [CdS91a, CdS91b, KFK91, Pow91, SHG91, VdSFC97, Wil91, vdBJ91]. Sutherland [Cho97]. Swendsen [GJ99, HDS98, MOS91a, MOS91b, RT90, SS96a, SS97b]. swept [MS93b]. Swindle [MS92]. Symmetric [BJS99, BW98b, Fel98, Kei98, Nag98, DLM^{+}90, Han95, Kie92, KS93c, LMR93, MS96a, PST94, Yam96]. symmetrical [RBB95]. Symmetries [Rad99, AL95c, CPPG91, Hio90, PPNM97]. Symmetry [BvEN99, GJLL99, GKC94, KT94, LS99, Sre98, Wat99, AM94a, BMR95, DS92, EFGM95, Koh90, LP91, MW94, Mom96, MG92, OD96, SL95]. Symmetry-Breaking [BvEN99, LP91, MG92]. Symplectic [Wid99, BG94, Wei95]. Symposium [Ano95i, Ano96k, Ano93n]. Synaptic [MTG99]. synchronization [BNS92, BVR93, SL93a]. synchrony [TKG93]. Synergetics [Wes91]. synthesis [Bry97]. System [AHR99, BR99, Bou99, IdRB98, Jun98, Kar99, KM98, Leg98, MPdlR93, Mou99, dRIB99, Alb95, APT94, BC92, BCK97, BK94b, CDPP90, Cho97, Com91a, Con90, CCG90, DDM90, DBH91, DN94b, For93, GI92, GIT91a, GW93a, GW93b, GW94c, GMM90, GDJH93, GS95, HR95, IM96, Jag91, Koh92a, KK91b, LK95, MC92a, Mür90, Oer95, PP91, Pen91a, PN96, RVW96, RR97b, SP91, Shi90a, THK^{+}91, Tót90, Wan96b, YK91, YP95, YIK95, ZR91]. System-reservoir [MPdlR93]. systematic [Dua90, FLB91]. Systematics [CS90b]. Systems [AKR98, AF99, Ano99k, Ano99l, ADE98, ABL97, BHK98, Ban99, BJ99, BTT98, Ber99, BO99a, BO99b, BV98, BM99, CG99a, CJB99, CMR98, CFL98, Che98, Cro98, EPRB99, ETW98, FMPP99, Gar97, GD99, JT98, Jia99, KT99, LMP99, tL90, LPY98, MS99, MGD98, NK99, PS99b, RS97a, Sos99, Sre98, WT93, AEWF91, AF95a, ABCP96, AS91b, ANS93, AE91, BGP95a, Bal91, BG93a, BFSV91, Ben93, BG90a, BJO97, BAP93, BL94, BJL92, BBC94b, BCF97, BT90c, Bry94, CMP95, CM95, CGS95, Che95, CL97, CV93a, CR97, DFF96, DKKP96, DR93, Dow91a, DR91, Edi93, EM94, EIK92, ELS96, For91, FJT96, FSB91, GMPS93, Gas97, GSH90, Geo95, GLM95, GL97, Git90b, Git90a, Git91c, Git93b, GGD91, GW94d, GKT93, GZ97, GGP92, Han95, HMV90]. systems [Hen94, HS92c, HQSS94, HC91, ILF90, JLM93, JMP94, Jan95, JT96, Jer90, JM96, Jur95, Jus92, Kal91, KK92, Ken90, KGM92, Kie92, KB97a, KGM^{+}93, KY93a, Kli91, Koh91b, KT94, Koz96, KPWH95, KPSW95, LMM92, LPT96, LPS94, LO97, LM97, Man93, MSD92, Mar95, Mar97, Mas92, Mer92, MF92, Mom96, Mon91a, Mon91b, Mon92, Mon94, MV97, Muk91, MP93, NLT93, OR91, OPdlR95a, OPdlR95b, OP90, OPS93, OB91b, PY94, PY95, PBP97, Per91, PSZ93, PWG97, Pol90b, Pol91a, PNT91, Pra94, Rap92b, RB94, Rom90, RS92b, RM97, RBF93, SHW94, Sel97, Shi92, Shl90, SPR91, Ste91, Too94a, Too95, TP92, Vol94, VR97b, Wan96a, Weh97b, Wei92, WHF92, Wol92, YRHMJ92, Zeg90, Zeg94, Zim93, Zim94, dLPP90, dO95, dB91, vB91, vD90]. systems.Part [Too94b].

T [KS93b]. T-cell [KS93b]. table [OAB^{+}96]. tagged [Sza93]. Tail [Mol97, dHNS92]. tailed [HW90]. Tailing [HLW99]. Tails [Too97, AC97d, BHKL95, MZ90, MM97b, dHNR92b]. tangency [YT90, YT91]. tangent [Pol91b]. tangle [CMVG95]. Taylor [Wid98, NAC91, SCM96, WHS^{+}91]. TBA [War96a]. Teacher [Kob97]. Technical [Ano96j]. Techniques [ZFB98, Ram93]. Technology [Ano99i]. Teller [BC98a, BCS93, BLL94, PV97, SS96a]. Temperature [CCM99, CMP97, Cer97, CQ90, FKST99, SS98, SZ98, VHR98, BM96a, BMC94, BZ97, dAB91b, BFB94, BC90b, Cam91, CMP95, CT96a, CSS95, Con90, Con96, DFF96, DHP96, ED92b, ES93, FVY92, HK96, JR94, Ken90, Kra94, LM94a, LM96, LP90a, MK91, Mar92b, MOS90b, MOS91b, Mar92d, MOR94, MO96b, MNO97, MG92, NR90, Par91, Pei95, Sal95, Shi93a, WT92, vEFK95]. temperature-driven [MG92]. Temperatures [Bov98a, Zhi98, ZP99, vEMZ98, BZ95, Ken97, MZ96, Nev95, SO91, vE96]. Temporal [Wal91, SR93]. tension [AS91a, ABCP96, CM95, For92, Fuj90a, Fuj90b, Fuj91, Fuj92, JS95, MMSR92, MOT90, MS95a, Mor92b, OB96]. tensions [KLMR90]. tensor [BHP97]. tent [OD96]. Term [BB98, NMHS99, Nic93, WRJ95, ZXZY94]. terms [MHB90, Uen95]. tesselation [KKBS92, KKBS93]. Test [Mül99, BB91, Cla91a, HL97a, LS90a]. testbench [AF95b]. Testing [TE95]. Tests [FR90, BZW92, PR90]. Their [CG99a, JG98, Böt95, Jag91, Pen97, PPW94, Rom90]. Theorem [Mae99, Mon04, BFG93, BCvB92, Che95, Gal96, GMTB96, RRGT97, SS97a, YT90, vdH98]. Theorems [AE99, CG99b, Blu94, Che90, Opp94a, Opp94b, Ser96]. Theoretic [FR99, SEW98, Kna93, KT91, SF91]. Theoretical [Dow91b, Rue99, Tor91, ZM93, Ano99i]. theories [Bha90, Bre91, FdlL92a, Lev96]. Theory [Ano92a, AGL91, BR99, BJ99, BDM90, BE99, BB97, CT98a, CCMS99, CV93a, CDMV98, FP98, IRB^{+}99, tL90, LC91, NP99b, PS99a, Ram95, RB90, SBB98, SPR91, Shu93, Vel98, Wid98, BLS94, BM96b, BG90a, BS91a, BCvB92, BK90, BvV95, Bry96, Buo90, Cer97, CWP97, DF93, Dor93, Ers92, Eu94, FZ97, FST96, FGMA93, For90b, GGM91, Git92a, Git93a, Git97, Gra90, HS91, HS92b, HSW97, JR91a, JD93, KK92, KP95, KW93, KNV93, KLMR90, LP90a, Lia91, Lin93, Lit92, LM92, MS96a, Maj93, Mar93b, MS91, MY94, MR94b, O’C93b, Oku90, OPdlR95a, Opp91, OSE93, PR94a, Pen97, PPR93, RST91, RS91a, RW96b, RM97, Row97, Sam95, SK90, Sch96b, Sni91, SO92a, SO92b, SZ91, Tót90, TP92, Zha92, dMPS95, vEFS93, vWL95]. There [BMHH97, BK96b, Cer90, Rue90]. Thermal [BEM99, ED92a, RC96, Sch93b, ZK98, LS97, MGA95, MGA97, MSZZ90, Per95a]. thermalization [BP97]. thermalized [Ger93b]. Thermally [EPRB99, BB94]. thermocapillary [BJO97]. thermodynamcis [Sch94a]. Thermodynamic [AvBED97, BGW98, Gay92, PS92c, VB98, BGR94, Don95, Eyi90, GGM91, GPJ92, HJ90, Jus92, Kie90, Kie92, Lop90, Pen95c, Pok93b, RS96, Shi90a, Spo92, Koz97]. Thermodynamics [Ano93n, Ano99d, Bry94, Ked93, MKP91, NDF92, BG93c, BG93d, Dor93, Eu94, Mil95, Opp94a, Opp94b, Opp96a, Opp96b, Pen94b, Pod95b, Por97, Rut92, Shi92, Spo96, VGC92, Ano99j]. Thermoelastic [CCF99]. thermohydrodynamics [CCD^{+}91]. Thermomechanical [FG99]. thermophysics [Tai95]. thermostat [MSG95, Rue97]. thesis [Mac95a]. Theta [CB90, RCB90]. Theta-point [CB90, RCB90]. thickness [Gác90]. Thin [BMSW99, Dev91b, PB94, RBB95]. thin-film [PB94, RBB95]. things [Cla91a]. thinning [Wil91]. Third [Ano90p, Bur91, CLT90a, Ano97h]. thought [Shl91]. Thouless [Mar90a, AC92, AF95b, AC97a, AC97c, CCST90, MKP90, vE90]. Three [CMVG95, CA93, GW99, GM97a, HH96, OR95a, PRW99, SS98, SL91, dVOS98, ABF^{+}95, AC97d, AORZ95, BM96b, BB92a, BB93, BL93, BK92b, dAB91a, BD93, CWSD92, DKMM94, DDM90, Dro96, DB90b, DR91, EM95, FR95, FMOU90, GHPS96, Gar91, GW95, GC90, GH94, Gre90, HMY96, HMP96, Heu93, HC91, HS90b, HPS94, KSZ97, KKBS92, KKBS93, KK94b, LMS95, LF90, MHJ94, Muc96, MSZZ90, OBB95, Pae90, PPR93, Pol92, PS97c, RW91, Rom90, RVW96, RDWW93, RR97b, SA91, SS97b, SL97a, SO91, TvROW96, TLW91, VWG93, Wee91, WR97, dH94]. three-body [MSZZ90]. three-component [HS90b]. Three-Dimensional [GW99, PRW99, CMVG95, HH96, OR95a, ABF^{+}95, AC97d, AORZ95, BB93, BL93, dAB91a, BD93, DB90b, DR91, EM95, FR95, Gar91, Gre90, HMY96, HMP96, Heu93, HC91, HPS94, KKBS92, KKBS93, KK94b, LMS95, LF90, MHJ94, Muc96, Pae90, PPR93, PS97c, RW91, Rom90, RDWW93, SA91, WR97]. three-disk [RVW96]. three-manifolds [Pol92]. three-phase [DDM90]. three-site [TLW91]. Three-State [SS98, BM96b, dAB91a, DKMM94, FMOU90, GH94, KSZ97, OBB95, SS97b]. Threshold [JL98, ACM95, Ben91, Ben92a, BDM90, DB90b, MS94]. thresholds [KLR94, Rom90]. Thue [ZPK97]. Tight [DFF99, MO94]. Tight-Binding [DFF99]. tiling [LPW92, dGN97]. tilings [DMB97, Hof95].

Tilted [PSR98]. Time [BJ98, BRR96, BBM92, CM98, CLV98, FP98, Fer98, Gar98, Gar97, GPJ92, Hay99, JMM87, JD93, KH94, Man93, MTG99, Mol98, RRT98, AZ95, BNN97, BP94a, BAP93, BJL92, BPH94b, BC90b, BMS97, CLV97, CP97, CLS90, CF97, DR93, ED90a, ED90b, GMCP96, GW94b, Gra94, HW90, Hen97, Joy96, KG90, KMKdC96, KK91c, Kot95, Koz96, Lan94, LK95, LV93, Mar94a, MP94, MM97b, MC92a, Men92, MR96, MC92b, MHM94, OYSK91, Pal90, Pom93b, Pra94, PBSR97, Pri93, SP91, TDSR95, WH91, ZHCD95, dHNS92, dHNR92b, JLM95, Leb97e, Spo92, Opp98b]. Time-Dependent [BJ98, Gar98, GPJ92, JMM87, JD93, BJL92, BC90b, Joy96, KMKdC96, Pal90, PBSR97, JLM95]. time-harmonic [BMS97]. time-independent [ZHCD95]. Time-periodic [BRR96]. time-resolved [KK91c]. Time-reversible [KH94]. Time-series [BBM92]. Times [BHS99, BOV98b, TNN99, Bin92, Che97a, Lie93, Oer95, Pen97, Sim94c, Too95, YL96, vK93]. Toda [LH92]. Toeplitz [Böt95, FM91]. tone [TKG93]. tool [Paj97]. Toom [Gác90, BKM93]. top [MD97]. Topics [Ano99-31, FZ97]. topography [KWG96]. Topological [BS97, Ber94b, BT93b, Ste99a, Aiz94, BK92b, CK95, Ge91, NP93, Wei92, Wei95, LPR91a]. Tori [Sin99, HSK91, MG95]. toroidal [MF91]. Torsional [Fer98]. tortuosity [KZ93]. torus [AC91, Che97a, CP97, Che92, Che93, Pin94, STV94]. Totally [SK99, BDG97, DJLS93]. touch [ABK94]. touch-and-stop [ABK94]. TP [GJLL99]. Trace [CDMV98, RB94, Pey91, RVW96, ZXZY94]. Tracer [Spo90, PPQ90, Zha92]. tracers [Bre91, KWG96]. Traffic [BH98, ERS99, Wag98, KW97]. trails [OP95]. trains [BZ90]. trajectories [CC97]. Trajectory [Wie97, Wie98a, MS93a, RW96b, dSVG90]. Transfer [BTT98, LW98a, MD97, PS91c, Bax91, Bax93b, COA95, Dev91b, Fuj90b, Kom93, Koo95, PB90]. Transfer-matrix [MD97]. transform [HTPH93]. Transformation [CK95, BMO95, GM88, KG95, Tab96, GM96]. Transformational [MSS98a]. Transformations [BCO99, CLLL98, AZP97, Din96, GBP91, GS93c, Ken93, MO93, ZR91, vEFS93, vE96]. transforms [Fal92, OTT92, Ras93]. Transience [dHMP99]. Transient [AE99, HMV90, SRC93, ZKB91a]. Transition [CKK99, JK99, SBZ98, AN91a, AC97a, All95, dAB91a, BB94, BV94, Cao93, CM96, CPPG91, CO96a, Cla91a, CCG90, CC91c, Dai90, DM94, EM95, EK96, ED92b, FMS97, For90b, FdH94, FZ91, Git93b, GW94c, Gry92, HMP96, HT91, Kou90, Lop90, MHB90, MG92, OS95, OS96a, ON96, Orz96, PS93b, PBSR97, Rad93, RV91, Rus94, SS97a, Sch95a, ST90, Smi90, TT94, Tót90, TP92, Yan94, bABD90, vEFK95, ST91]. transition-state [TP92]. Transitions [AHR98, BC98a, BG97a, BC99a, FG99, GR98, LMP99, Sal99, SSZ99, Whe99, AM95, AEGL92, AE91, Bag96, BZ95, BBC^{+}95, BK90, BI92, BK95b, Cag90, CCT96, Cho94, Gra95, HMV90, Hio90, KH96, Ker93, KY93a, KY93b, KSZ97, Mül93, Nic93, NE95, PS94a, PS92c, PR90, RS96, Sch93a, SD93, Shi92, SPR91, Ste95b, Uen95, VZ92, dOdOCS95, vEdH91]. Transitivity [McK99, McK95]. translation [CF97]. Translational [AHR99, GMTB95]. Transmission [AAH98, GPS90, KPW91]. Transport [Abd98, BFPR99, BO91a, BF98, BE99, DF98, Hav90, LBT97, NK99, Pet99, SBB98, ADG96, Buo90, CK91, CdS91a, CdS91b, ED96, GAA97, Ger93a, GGL90, GDJH93, HR95, Ish96, Kar95, KS92, Krá97, Lüt95, MGS94, MR96, Nay93, Pla90, Sza97, vCEBS94, vK91]. Transverse [PP95, CR94, HMY96, JK93, Luc93, MM90b, Pok93b]. Trap [DRbA99, GHW91, Red94, Zif91]. Trapping [AK91, AEA93, ZKB91b, APC^{+}92, HC91, Par91, RS91a, Wol92, dB91]. traps [AK91, BNRW93, Mil92, NB90, Rub91, Tor91, WH91, dHS92]. Traveling [PM99, OYSK91, Xin93]. traveling-wave [OYSK91]. Treatment [BMSW99, FG99, LD98, LZ91]. Tree [Ald93, DJM99, Sal99, BM97, Gon94, GUJ94, Häg96, KK94a, MD94, Par93, REK91, dABMR90, dABR91, vdBDP92]. Tree-based [Ald93]. Trees [BGL98, Sch98, FNW92, GRZ90, HS90a, Mad95, MvR97, Mon91a, Mon92, Pen94a, Wu97b]. Trend [Cer96, GTW95]. trends [Sch96b]. Triangle [AH98, BW98a, Sch98, BB93, HH95b, dGN97]. Triangular [Bav94, Bax93b, GM97a, Ken97, KC91a, KvL92, MM97b, Mom94, PS94a, WC95]. trick [PST94]. Tricritical [KP91a, OPB93]. Tridimensional [PPO99]. trimer [BAK91]. Triple [Tay99]. Triple-Junction [Tay99]. triplet [AC90a]. trongly [SB97b]. True [Tót94]. Truncated [Gol99]. Truncated-Path-Integral [Gol99]. Trunov [Pod98a]. truth [Omn91]. Tube [Sta98]. tunneling [GS95, MOT90, Sto97a]. Turbulence [AGL92, BLPP98, Ber98, Eyi96, EA98, GLR98, BCFM95, CLT90a, CLT90b, ES93, Gal97, JT97, Lia91, Rue90, Sre98, Sta97a]. Turbulent [JVH98, SS95, Ber94b, BCPV97, Cho92, DW91, EP92, Fin92, LFtH91, Maj93, Maj94, MS96b, Nay93, RR97a, SW91]. turning [vK95b]. Twist [OvR95, BCCF92, CC91b, Dav91, HSK91]. twisted [BKJZJ93]. Two [AKK99, ABPSJ91b, Asl99, BC98a, BP94a, BF96, BBG98, CISS99, CCF99, CL98, CG99b, Don99, Gon94, HLIM93, HP91, JK99, JT98, Jer90, Jos98, Kin99, KS97c, LD98, LMN98, MPR98, MSS98b, MKK97, Naw98, RC96, SS97c, SZ98, SSLI97, Sin91, SZ99, Ste95a, SI99, Tan94, Vuj99, WC98, Zhi98, AM95, ANV94, AC92, AF95a, ACJL92, AM94a, AORZ95, AEA93, BEC93, BZ95, BP94b, Bel93a, Bel95, Ben91, Ben92a, BR92, BT90a, Bha90, Ble90, Ble91, Ble92, BJL92, BB94, BHJ92, dAB91b, BFB94, BL91a, BL91b, BS91b, Bur91, CO91, CGMS96, CSS95, CGS95, CP97, CF97, CT95, CSB97, Con96, DFZ94, DHV92, Dev91a, DBB^{+}92, EFGM95, ES93, Fal92, FFK94, FjLL95, FLL95, For90a, FM91, FJ92, For92, FJM92, FJ96, Fos93, FG94]. two [GI92, GIT91a, GW94a, GMO91, GPS90, GMMU97, Gry92, HR92b, JR94, JMP94, JS95, JM96, JT97, Jur95, KLT97, KF90, Krá96, KA94, LPPSA92, LK95, LS90a, LMS95, MR94a, MC94, MSD92, MM96, Mar90a, MKP90, MK91, MOS91b, Mar94c, MS91, MS95b, MSG97, NV93, Nob95, OAB^{+}96, OB91a, ON96, OL97, OP95, PS92b, PS93a, Per91, Pok93a, PPQ90, Pri94, Pri93, PPNM97, RHA97, RL91, Ric97, Rob91, RR97a, SA91, SS97b, SP95a, Sch90a, SDJ^{+}96, SP94, TLW91, VWG93, Wan96b, WLC94, YT90, YT91, vK94, HK96]. two- [LMS95, SA91, TLW91]. two-band [GPS90]. two-color [RHA97].

Two-Component [CCF99, AF95a, For90a, FJ92, For92, FJ96, MC94]. Two-Dimensional [Asl99, BC98a, BBG98, CISS99, CL98, JK99, JT98, Kin99, LMN98, MPR98, MSS98b, SS97c, SSLI97, SI99, Vuj99, ABPSJ91b, BF96, Gon94, Jer90, RC96, ANV94, AC92, ACJL92, AM94a, AEA93, BEC93, Ben91, Ben92a, BR92, BT90a, Ble92, BB94, BHJ92, dAB91b, BFB94, BS91b, Bur91, CO91, CGMS96, CSS95, CP97, CSB97, Con96, DFZ94, DHV92, Dev91a, DBB^{+}92, ES93, FjLL95, FLL95, For90a, FM91, FJ92, FJM92, FJ96, Fos93, FG94, GIT91a, GMMU97, Gry92, HR92b, JS95, JM96, KLT97, KF90, KA94, LPPSA92, LS90a, MR94a, MC94, MSD92, MM96, Mar90a, MKP90, MK91, MOS91b, Mar94c, MS95b, MSG97, Nob95, OAB^{+}96, OB91a, ON96, OL97, PS92b, PS93a, PPQ90, Pri94, PPNM97, Rob91, RR97a, SS97b, SP95a, Sch90a, SDJ^{+}96, Wan96b]. two-dimensional [WLC94, YT90, YT91]. two-layer [LK95]. Two-Level [WC98, BJL92, Per91]. Two-Parametric [AKK99]. Two-Particle [Zhi98, Bel93a, Bel95, BL91a, BL91b, Pri93]. two-phase [AORZ95]. Two-Point [SZ98, SZ99, CT95]. two-sided [MS91]. two-slit [vK94]. two-species [FFK94, RL91, Ric97]. two-spin [GI92]. Two-spin-majority [HP91]. Type [CCMS99, Der97, IRB^{+}99, KPS98, LS99, MPR98, SS96a, BCK97, Bob93, BBC^{+}95, CH94, DDM90, Kie92, Mar92a, MM93, Shi90b, Uen95, YT91]. types [SA94].

Uhlenbeck [BS98, DMR97, DR98]. Ulam [ACM95, CG99a]. Ultimate [Mie98]. Ultrametric [AMF98]. Ultrasmall [Jos98]. Unbounded [AKR98, AKMR98, LPY98, BFG93, GM95, PY95]. uncertainty [DGZ92b]. uncorrelated [Rei96b, SSV93]. Uncoupled [KK98]. undergoing [TH96]. Undergraduates [Cop98]. Undirected [BHK98]. uniaxial [GD93, Kho91b]. Unicity [BC98a]. unidimensional [CQ90]. Unification [War96a]. unified [FLS96, LPT97]. Uniform [CS91b, Kie92, JMM87, KK91b]. unimodal [GRZ90, IP90, KK94a]. Uniquely [CLLL98]. Uniqueness [IVDB98, JM96, PY95, PZ99, TV99, BC90a, FN95, GM95, SS97a]. unison [Kli92]. Unitary [Wid99, MQ91, PS97a]. Universal [DA99, FT90b, FT90c, Hal97, IP90, JM90, vWH97, DR93, LMS95]. Universality [BL94, DS92, FJT96, HRS97, PS97a, Gra95, KS92, LPPSA92, ORG91, dSCT91]. universally [RW91]. Université [Ano92q]. University [Ano91a, Ano96j, Bry98, Git98, Opp98b, Pod98b, Rap98, Sac98, Sre98, Ano90o]. unresponsiveness [KS93b]. unstable [BBC94b, PNT91, Süt96]. Unusual [Edi93, LP91]. Update [RSSS98]. Updating [LZ98]. Upper [BFB94, CMP97, Joy96, KK91a, Mon94, Noo98, Tót91, Bel93a, dAB91b, Con90, HS90a, Kie92, Red94]. upsilon [BSG91, Sas92]. Use [Pol90b, Pol91a]. Uses [OO91]. Using [AAH98, ST97, BS90b, CK95, Dah96, EKLR94, FLS95, HWvB97, Koh91a, MHdA90, OD96, RRGT97, Smi94].

V [Pod98a, Sac98, WL95]. vacancies [SSV93]. vacuum [MOT90]. valence [BS90c, YHHK96]. valid [RW91]. Validation [CC94b]. Validity [BCP98, But93, EKLR94, SS90]. Value [Der97, Hei98, BNN95, BMP90, IS96, Swe97]. Values [BJ98]. Vapor [LMP99, RW91, TT94]. Variable [BG97b, BBM92, CPS92, Rus93]. Variable-Coupling [BG97b]. Variables [Ano01, BT90d, EKLR94, NS95, Weh97c, YP95]. Variance [BT93a, Weh97a]. Variation [CGTM99, Mor90]. Variational [Che99a, EA98, Lef99a, Tay99, VBF97, Nev95, SK90, SS96b, TP92, Wei92, GdH91b]. various [Gro95, KGM92]. Varying [NK99]. VBS [FNW92]. Vector [BP98, Whe99, dVOS98, dMM98, BKJZJ93, Koz97, SA91, SO91, TF92, TF93a, TF93b]. vectorizable [MH92]. Vectorized [Eve93, Rie93]. vectors [GM88, GM96]. vehicular [KW97]. velocities [MSG97, Nad95]. Velocity [BS98, BC98b, BC99b, BvV95, GLR98, IMS92, PS98, AFNB97, CSRPS93, CQ90, GS92, IW93, LFtH91, MS96b, MM97b, Sle96, Wag95]. Ventsel’ [Tót90]. Verhulst [WRJ95]. Version [Bov98a, CO96b, Mür90, dMP91]. Versus [BMSW99, dHMP99, Com91a, GMR98, Gon94, LMM92, Lem95, Li90, MOR94]. Vertex [Alb98, Fuj98, Wat99, BS90a, BMR95, BS95, EM95, Fuj92, Fuj96, HH96, KS93c, KL91, LP90b, Nol92, SMS96]. vertical [GIT91a]. Vertices [Fuj98]. very [Koh91b]. Vesicles [BEO98, BOP94, PO95]. VHP [SS92b]. via [Don99, FS93, SS97a, SW91, Sle96, TC94]. vibrated [MND92]. vibrational [MC92a, MC92b]. vibrational-energy-time [MC92b]. Vibrations [KF97]. vicinity [YIK95]. View [NOV99, FT90b, FT90c, RW91]. viewpoint [ACLS94]. VII [vdBV93]. violating [EB95a]. violation [Sto97b]. Virasoro [DKMM94]. Virial [BD94, Pol90b, Pol91a]. virtual [BNRW93]. virus [KS90]. viscosities [Koh92b]. Viscosity [Avr98, DFHR90, VdSFC97]. Viscous [DB90a, WM98]. visited [Ber94a]. visits [CG93]. visual [SR93]. Vlasov [BDIV97, CM98, DdH96, IVDB98]. Vleck [Lit92]. Vol [Dom97, Dou97, Wei91a]. Volume [Ber94a, GD99, Yos98, ZP99, CSS95, Gas97, MRS94, MOT90, MOR97, OP90, Pok93a, Wan96a, Wei96, dMPZ92]. Volume-Preserving [GD99, Gas97]. Voronoi [KNS98, KKBS92, KKBS93]. Vortex [Cho94, BB94, CO91, Cho92]. vortices [CO91, ES93, McC95, Rus93, WCT91]. vorticity [Bun97, Gre90]. Vote [Moo97, Che97a, dO92]. voter [Mou92, ST95]. voting [Gal90]. vs [LSP91].

W [Git98]. Waals [Gia91, KP95, MLM93]. Walk [BGL99, CvD98, RRT98, dHMP99, AZ96, ABPSJ90b, ABPSJ91a, APC^{+}92, Bra91, CS91a, DT93, Dou95, EN92a, GB90, HSS93, HSS95, Ken94, NDF92, TvROW96, WMS90, WH91, Wu95, dHNS92, vWH97, Ano99h]. Walker [AH98, BP94a]. Walks [Ali99, BBG98, BZ99, CT98a, JW98, Noo98, Sin99, AZ95, AK91, ABPSJ90a, ABPSJ91b, BF96, Bar90, BA93, BL91a, BBOC91, BH91, CPS90, CPS91, CPS92, CPP94, Dou97, GBN92, HW97b, Hio90, HA97, KNV93, Kot95, LM91, LS90a, LMS95, MOS90a, MW94, Mil92, O’B90, OPB93, PP93, Pen94a, RR93, RB90, Rub91, SA94, SA95, SSV93, Tel90, Tót94, Wei96, dHS92]. Wall [AC99, BEO98, RNCT98, ACDD90, CM96, FM91, MHL92]. Walls [HLW99, KP94a]. Wand [Mon92]. Wang [BS98, GJ99, HDS98, MOS91a, MOS91b, RT90, SS96a, SS97b]. Wang-Type [SS96a]. Wannier [ADE98]. Ward [MPR98]. Ward-Type [MPR98]. warm [PN94]. was [Mac95a]. Water [IRB^{+}99]. Wave [AS91b, ACLS94, Bro98, FE98, McK99, BO91a, CP93, FS93, McK95, OYSK91, RW91, Wal91, Wee91]. Wavefunctions [Pra98, MSD92, Mer92]. Waveguide [NR98]. wavelength [Cla91b]. Wavelet [Fal92, BMA93, Ras93]. Wavelets [Paj97, Hol94, O’C93a, PO93]. Waves [BMSW99, ZK98, AS95, FK94a, FK97, GPS90, KPW91, Nos93, RV97, Xin93]. Weak [CCT93, CDMV98, Git92d, MPR98, Mar93b, McC95, MD94, Bov90, CCT96, Cer96, DMRR92, Ger90, MS96a, Pet93, Rei96b, Spe91, YK91, vEM92, CdC94, Rei96a]. Weak-disorder [MD94]. weak-noise [MS96a]. Weakly [GMR99, Lef99b, LB96, LM97, Bob93, Ken94]. wedding [Kru97]. wedge [Dag96, GC90, Sze96]. Wehr [Ano01]. weights [CD90]. Weiss [BG93d, GZ98, HNS91, MP98a, VZ94, dMP91, dMPZ92]. Welcome [Ano98o]. well [TV90]. well-ordered [TV90]. Wetting [BN99, BDK99, IRB^{+}99, AN91a, ABHP90, BDDM90, DHV92, DLM^{+}90, HMY96, PS93b, Smi90]. Wetting-Induced [BN99]. Which [MTG99]. white [Rei96a]. Widom [BHW99]. width [RV97]. Wigner [Buo90, RW90]. wih [MKP91]. Wiley [Opp98a]. Will [BGL99]. Wilson [KLMR90]. winding [CDM93]. windy [RK96]. wine [Cla91a]. within [Dan93, DDM90, GGM91, GKRT94, KB91]. Without [BC99b, AK91, Bra91, Cor95, Iof94, Jes96, KD90, Pae90, PST94]. Witten [Mar92a]. work [Joh90]. Workshop [Ano92a, Ano92r, Ano94a, Ano98o, Ram95, Ano91p, Ano90n]. world [DI93, Nos90]. Wulff [DKS93].

X [STAJ95]. X-ray [STAJ95]. XXZ [AW90b, GI92, MKK97]. XY [AC90a, GH94, KM97].

Yang [For90b, Koz97, PS93b]. years [Phi94]. Yevick [BT90b, GGM91, HSW97]. York [Opp98a, Sre98]. Yukawa [BVHP92, GY98, RC97].

Z [BFSV91, SY90]. Zaleski [Rap98]. Zamolodchikov [BQ90a]. zeolites [KB97c]. Zernike [BVHP92, GY98, KRT97, KRT99]. Zero [FKST99, FVY92, SS98, AC97a, AC97c, BKL97b, CSS95, CPP97, DHP96, Kra94, Mol95, Shi93a]. zero-density [AC97a, AC97c]. zero-range [BKL97b]. Zero-temperature [FVY92, DHP96, Shi93a]. Zeros [Smi90, BGP95a, BL93, BD97, PS93b, RS92c]. zeta [Aur90, Dah96]. Zhang [BCK97, DA99, FT93]. zone [KK91c].

References

Ahmed:1998:MHB

[AAH98] E. Ahmed, H. N. Agiza, and S. Z. Hassan. On modeling hepatitis B transmission using cellular automata. Journal of Statistical Physics, 92(3–4):707–712, August 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023001007714.

Aubry:1992:CPB

[AAR92] Serge Aubry, Gilles Abramovici, and Jean-Luc Raimbault. Chaotic polaronic and bipolaronic states in the adiabatic Holstein model. Journal of Statistical Physics, 67(3–4):675–780, May 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049725.

Angelescu:1992:RCH

[AB92] N. Angelescu and M. Bundaru. A remark on the condensation in the hard-core lattice Bose gas. Journal of Statistical Physics, 69(3–4):897–903, November 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01050441.

Alberti:1996:STI

[ABCP96] G. Alberti, G. Bellettini, M. Cassandro, and E. Presutti. Surface tension in Ising systems with Kac potentials. Journal of Statistical Physics, 82(3–4):743–796, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179792.

Abdallah:1998:HKQ

[Abd98] Naoufel Ben Abdallah. A hybrid kinetic–quantum model for stationary electron transport. Journal of Statistical Physics, 90(3– 4):627–662, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023216701688.

Adler:1995:STD

[ABF^{+}95] Christopher Adler, Bruce Boghosian, Eirik G. Flekkøy, Norman Margolus, and Daniel H. Rothman. Simulating three-dimensional hydrodynamics on a cellular automata machine. Journal of Statistical Physics, 81(1–2):105–128, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179971.

Albano:1990:CWS

[ABHP90] E. V. Albano, K. Binder, D. W. Heermann, and W. Paul. Critical wetting in the square Ising model with a boundary field. Journal of Statistical Physics, 61(1–2):161–178, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01013958.

Arneodo:1997:OSC

[ABJM97] A. Arneodo, E. Bacry, S. Jaffard, and J. F. Muzy. Oscillating singularities on Cantor sets: A grand-canonical multifractal formalism. Journal of Statistical Physics, 87(1–2):179–209, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02181485.

Andrienko:1994:PFG

[ABK94] Yu. A. Andrienko, N. V. Brilliantov, and P. L. Krapivsky. Pattern formation by growing droplets: The touch-and-stop model of growth. Journal of Statistical Physics, 75(3–4):507–523, May 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02186870.

Asselah:1997:SDS

[ABL97] A. Asselah, R. Brito, and J. L. Lebowitz. Self-diffusion in simple models: Systems with long-range jumps. Journal of Statistical Physics, 87(5–6):1131–1144, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02181276.

Arora:1996:SPO

[ABP96] D. Arora, D. P. Bhatia, and M. A. Prasad. Survival probability in one dimension for the A + B → B reaction with hard-core repulsion. Journal of Statistical Physics, 84(3–4):697–711, August 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179654.

Aslangul:1990:DEO

[ABPSJ90a] Claude Aslangul, Marc Barthelemy, Noëlle Pottier, and Daniel Saint-James. Dynamical exponents for one-dimensional randomrandom directed walks. Journal of Statistical Physics, 59(1– 2):11–21, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01015561.

Aslangul:1990:MDE

[ABPSJ90b] Claude Aslangul, Marc Barthelemy, Noëlle Pottier, and Daniel Saint-James. Microscopic dynamical exponents for randomrandom directed walk on a one-dimensional lattice with quenched disorder. Journal of Statistical Physics, 61(1–2):403–413, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01013972.

Aslangul:1991:RWD

[ABPSJ91a] C. Aslangul, M. Barthélémy, N. Pottier, and D. Saint-James. Random walk on a disordered directed Bethe lattice. Journal of Statistical Physics, 65(3–4):695–713, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01053749.

Aslangul:1991:TDR

[ABPSJ91b] Claude Aslangul, Marc Barthélémy, Noëlle Pottier, and Daniel Saint-James. Two-dimensional random-random walks: Dynamical exponents in a quenched directed model. Journal of Statistical Physics, 65(3–4):673–693, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053748.

Anagnostopoulos:1999:IMQ

[ABT99] K. N. Anagnostopoulos, P. Bialas, and G. Thorleifsson. The Ising model on a quenched ensemble of c = −5 gravity graphs. Journal of Statistical Physics, 94(3–4):321–345, February 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004583901498.

Alcaraz:1990:CIC

[AC90a] Francisco C. Alcaraz and Clisthenis P. Constantinidis. Conformal invariance and the critical behavior of the triplet XY quantum chain. Journal of Statistical Physics, 60(5–6):639–657, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01025986.

Arkeryd:1990:GEE

[AC90b] Leif Arkeryd and Carlo Cercignani. Global existence in L^{1} for the Enskog equation and convergence of the solutions to solutions of the Boltzmann equation. Journal of Statistical Physics, 59 (3–4):845–867, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01025854.

Amadasi:1991:BPO

[AC91] Luciano Amadasi and Mario Casartelli. Basins of periodic orbits for elliptic maps of the torus. Journal of Statistical Physics, 65(1– 2):363–372, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01329864.

Alastuey:1992:CKT

[AC92] A. Alastuey and F. Cornu. Correlations in the Kosterlitz– Thouless phase of the two-dimensional Coulomb gas. Journal of Statistical Physics, 66(1–2):165–231, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01060065.

Aslangul:1995:DRC

[AC95] Claude Aslangul and Petr Chvosta. Diffusion on a random comb: Distribution function of the survival probability. Journal of Statistical Physics, 78(5–6):1403–1428, March 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02180137.

Alastuey:1997:CLN

[AC97a] A. Alastuey and F. Cornu. Critical line near the zero-density critical point of the Kosterlitz–Thouless transition. Journal of Statistical Physics, 87(3–4):891–895, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02181249.

Alastuey:1997:JDC

[AC97b] A. Alastuey and F. Cornu. Janco’s disciples in Coulombland. Journal of Statistical Physics, 89(1–2):5, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02770750.

Alastuey:1997:PCC

[AC97c] A. Alastuey and F. Cornu. Part I. The 2D classical Coulomb gas near the zero-density Kosterlitz–Thouless critical point: Correlations and critical line. Journal of Statistical Physics, 89 (1–2):6–19, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02770751.

Alastuey:1997:PIA

[AC97d] A. Alastuey and F. Cornu. Part II. Algebraic tails in threedimensional quantum plasmas. Journal of Statistical Physics, 89 (1–2):20–35, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02770752.

Aizenman:1998:SQS

[AC98] M. Aizenman and P. Contucci. On the stability of the quenched state in mean–field spin–glass models. Journal of Statistical Physics, 92(5–6):765–783, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023080223894.

Aqua:1999:CQA

[AC99] J. N. Aqua and F. Cornu. Classical and quantum algebraic screening in a Coulomb plasma near a wall: A solvable model. Journal of Statistical Physics, 97(1–2):173–207, October 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004667018492.

Alexander:1998:CER

[ACC^{+}98] K. S. Alexander, F. Cesi, L. Chayes, C. Maes, and F. Martinelli. Convergence to equilibrium of random Ising models in the Griffiths phase. Journal of Statistical Physics, 92(3–4):337–351, August 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023077101354.

Abraham:1990:LDI

[ACDD90] Douglas Abraham, Pierre Collet, Joël De Coninck, and François Dunlop. Langevin dynamics of an interface near a wall. Journal of Statistical Physics, 61(3–4):509–532, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01027290.

Artuso:1996:NEB

[ACG96] Roberto Artuso, Giulio Casati, and Italo Guarneri. Numerical experiments on billiards. Journal of Statistical Physics, 83 (1–2):145–166, April 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183643.

Arkeryd:1991:HBH

[ACI91] L. Arkeryd, S. Caprino, and N. Ianiro. The homogeneous Boltzmann hierarchy and statistical solutions to the homogeneous Boltzmann equation. Journal of Statistical Physics, 63 (1–2):345–361, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01026609.

Alexander:1992:SFT

[ACJL92] Francis J. Alexander, Zheming Cheng, Steven A. Janowsky, and Joel L. Lebowitz. Shock fluctuations in the two-dimensional asymmetric simple exclusion process. Journal of Statistical Physics, 68(5–6):761–785, September 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048875.

Aubry:1994:WPP

[ACLS94] Nadine Aubry, Fernando Carbone, Ricardo Lima, and Said Slimani. Wave propagation phenomena from a spatiotemporal viewpoint: Resonances and bifurcations. Journal of Statistical Physics, 76(3–4):1005–1043, August 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02188696.

Alabiso:1995:NSB

[ACM95] Carlo Alabiso, Mario Casartelli, and Paolo Marenzoni. Nearly separable behavior of Fermi–Pasta–Ulam chains through the stochasticity threshold. Journal of Statistical Physics, 79(1– 2):451–471, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179398.

Arenzon:1992:MIM

[AdAI92] J. J. Arenzon, R. M. C. de Almeida, and J. R. Iglesias. A multineuron interaction model for neural networks. Journal of Statistical Physics, 69(1–2):385–409, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053798.

Asch:1998:SDS

[ADE98] Joachim Asch, Pierre Duclos, and Pavel Exner. Stability of driven systems with growing gaps, quantum rings, and Wannier ladders. Journal of Statistical Physics, 92(5–6):1053– 1070, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023000828437.

Abdallah:1996:ETM

[ADG96] N. Ben Abdallah, P. Degond, and S. Genieys. An energytransport model for semiconductors derived from the Boltzmann equation. Journal of Statistical Physics, 84(1–2):205–231, July 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02179583.

Adler:1993:BRC

[Adl93a] Joan Adler. Book review: Chaos and fractals. Journal of Statistical Physics, 73(3–4):807, November 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01054354.

Adler:1993:BRF

[Adl93b] Joan Adler. Book review: Fractals for the classroom. Journal of Statistical Physics, 73(3–4):809, November 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01054355.

Arrowsmith:1990:MFS

[AE90] D. K. Arrowsmith and J. W. Essam. Möbius function for the set of acyclic directed backbone graphs. Journal of Statistical Physics, 58(3–4):553–574, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01112762.

Asorey:1991:FOT

[AE91] M. Asorey and J. G. Esteve. First-order transitions in onedimensional systems with local couplings. Journal of Statistical Physics, 65(3–4):483–494, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053741.

Ayton:1999:ACT

[AE99] Gary Ayton and Denis J. Evans. On the asymptotic convergence of the transient and steady-state fluctuation theorems. Journal of Statistical Physics, 97(3–4):811–815, November 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004679628622.

Avellaneda:1993:TPA

[AEA93] Marco Avellaneda, Frank Elliott, Jr., and Christopher Apelian. Trapping, percolation, and anomalous diffusion of particles in a two-dimensional random field. Journal of Statistical Physics, 72(5–6):1227–1304, September 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01048187.

Agiza:1997:GMB

[AEA97] H. N. Agiza, M. F. Elettreby, and E. Ahmed. On a generalized model of biological evolution. Journal of Statistical Physics, 88 (3–4):985–989, August 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000015183.65230.c3.

Alexander:1992:PTP

[AEGL92] F. J. Alexander, I. Edrei, P. L. Garrido, and J. L. Lebowitz. Phase transitions in a probabilistic cellular automaton: Growth kinetics and critical properties. Journal of Statistical Physics, 68 (3–4):497–514, August 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01341759.

Aharony:1991:BRS

[AEWF91] Amnon Aharony, Ora Entin-Wohlman, and Victor Fleurov. Book review: Statistical mechanics of magnetically ordered systems. Journal of Statistical Physics, 62(1–2):491–492, January 1991.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020885.

Alastuey:1995:CTC

[AF95a] A. Alastuey and P. J. Forrester. Correlations in two-component log-gas systems. Journal of Statistical Physics, 81(3–4):579– 627, November 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179249.

Alastuey:1995:TND

[AF95b] A. Alastuey and P. J. Forrester. A testbench for the nested dipole hypothesis of Kosterlitz and Thouless. Journal of Statistical Physics, 79(3–4):503–523, May 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02184869.

Alikakos:1999:EOR

[AF99] Nicholas D. Alikakos and Giorgio Fusco. The equations of Ostwald Ripening for dilute systems. Journal of Statistical Physics, 95(5–6):851–866, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004594131850.

Aurell:1997:BDS

[AFNB97] E. Aurell, U. Frisch, A. Noullez, and M. Blank. Bifractality of the devil’s staircase appearing in the Burgers equation with Brownian initial velocity. Journal of Statistical Physics, 88(5–6): 1151–1164, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02732429.

Aizenman:1991:SMC

[AG91] Michael Aizenman and Geoffrey Grimmett. Strict monotonicity for critical points in percolation and ferromagnetic models. Journal of Statistical Physics, 63(5–6):817–835, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01029985.

Asselah:1998:MEP

[AG98] Amine Asselah and Giambattista Giacomin. Metastability for the exclusion process with mean–field interaction. Journal of Statistical Physics, 93(5–6):1051–1110, December 1998.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033153.16878.b0.

Abdel-Gawad:1999:BSC

[AG99] H. I. Abdel-Gawad. On the behavior of solutions of a class of nonlinear partial differential equations. Journal of Statistical Physics, 97(1–2):395–407, October 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004683522126.

Aubry:1991:SAC

[AGL91] Nadine Aubry, Régis Guyonnet, and Ricardo Lima. Spatiotemporal analysis of complex signals: Theory and applications. Journal of Statistical Physics, 64(3–4):683–739, August 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01048312.

Aubry:1992:TS

[AGL92] Nadine Aubry, Régis Guyonnet, and Ricardo Lima. Turbulence spectra. Journal of Statistical Physics, 67(1–2):203–228, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049031.

Arndt:1998:MSP

[AH98] Peter F. Arndt and Thomas Heinzel. Metastability and spinodal points for a random Walker on a triangle. Journal of Statistical Physics, 92(5–6):837–864, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023036408873.

Ahmed:1996:FCA

[Ahm96] E. Ahmed. Fuzzy cellular automata models in immunology. Journal of Statistical Physics, 85(1–2):291–294, October 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02175567.

Arndt:1998:FOP

[AHR98] Peter F. Arndt, Thomas Heinzel, and Vladimir Rittenberg. First-order phase transitions in one-dimensional steady states.

Journal of Statistical Physics, 90(3–4):783–815, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023229004414.

Arndt:1999:SBT

[AHR99] Peter F. Arndt, Thomas Heinzel, and Vladimir Rittenberg. Spontaneous breaking of translational invariance and spatial condensation in stationary states on a ring. I. The neutral system. Journal of Statistical Physics, 97(1–2):1–65, October 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004670916674.

Aizenman:1994:SDC

[Aiz94] Michael Aizenman. On the slow decay of O(2) correlations in the absence of topological excitations: Remark on the Patrascioiu– Seiler model. Journal of Statistical Physics, 77(1–2):351–359, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186846.

Argyrakis:1991:TTC

[AK91] Panos Argyrakis and Klaus W. Kehr. Trapping without traps by correlated random walks. Journal of Statistical Physics, 63 (1–2):399–404, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01026612.

Acosta:1992:ADS

[AK92] Victor Acosta and Abel Klein. Analyticity of the density of states in the Anderson model on the Bethe lattice. Journal of Statistical Physics, 69(1–2):277–305, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053794.

Alimohammadi:1999:TPF

[AKK99] M. Alimohammadi, V. Karimipour, and M. Khorrami. A twoparametric family of asymmetric exclusion processes and its exact solution. Journal of Statistical Physics, 97(1–2):373–394, October 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004631505288.

Aizenman:1991:P

[AKL91] M. Aizenman, M. Kalos, and J. L. Lebowitz. Preface. Journal of Statistical Physics, 63(5–6):809, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01029983;   http://link.springer.com/article/
10.1007/BF01029983.

Albeverio:1998:SMB

[AKMR98] S. Albeverio, Yu. G. Kondratiev, R. A. Minlos, and A. L. Rebenko. Small-mass behavior of quantum Gibbs states for lattice models with unbounded spins. Journal of Statistical Physics, 92(5–6):1153–1172, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023009130254.

Albeverio:1998:PAL

[AKR98] S. Albeverio, A. Yu. Kondratiev, and A. L. Rebenko. Peierls argument and long-range order behavior of quantum lattice systems with unbounded spins. Journal of Statistical Physics, 92 (5–6):1137–1152, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023056913416.

Arkhipov:1994:CFE

[AKV94] Y. Arkhipov, A. Klar, and V. Vedenyapin. On the connection of the formulas for entropy and stationary distribution. Journal of Statistical Physics, 77(5–6):1027–1037, December 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183149.

Abraham:1995:CSM

[AL95a] D. B. Abraham and F. T. Latrémolière. Corner spontaneous magnetization. Journal of Statistical Physics, 81(3–4):539– 559, November 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179247.

Aidun:1995:LBS

[AL95b] Cyrus K. Aidun and Yannan Lu. Lattice Boltzmann simulation of solid particles suspended in fluid. Journal of Statistical Physics, 81(1–2):49–61, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179967.

Aubry:1995:SSS

[AL95c] Nadine Aubry and Ricardo Lima. Spatiotemporal and statistical symmetries. Journal of Statistical Physics, 81(3–4): 793–828, November 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179258.

Albano:1992:CBD

[Alb92] Ezequiel V. Albano. The critical behavior of dimer-dimer surface reaction models. Monte Carlo and finite-size scaling investigation. Journal of Statistical Physics, 69(3–4):643–666, November 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01050429.

Albanese:1994:GMK

[Alb94] Claudio Albanese. A Goldstone mode in the Kawasaki–Ising model. Journal of Statistical Physics, 77(1–2):77–87, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02186833.

Albano:1995:DSS

[Alb95] Ezequiel V. Albano. Damage spreading in a single-component irreversible reaction process: Dependence of the system’s immunity on the Euclidean dimension. Journal of Statistical Physics, 78(3–4):1147–1155, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02183707.

Albertini:1998:DDF

[Alb98] Giuseppe Albertini. Direction-dependent free energy singularity of the antiferroelectric asymmetric six-vertex model. Journal of Statistical Physics, 90(3–4):853–871, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023285222161.

Aldous:1993:TBM

[Ald93] David Aldous. Tree-based models for random distribution of mass. Journal of Statistical Physics, 73(3–4):625–641, November 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01054343.

Alili:1999:PRW

[Ali99] S. Alili. Persistent random walks in stationary environment. Journal of Statistical Physics, 94(3–4):469–494, February 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004596204224.

Allen:1995:IFS

[All95] Scott Allen. Interface formation and a structural phase transition for the spherical model of ferromagnetism. Journal of Statistical Physics, 79(1–2):165–181, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179385.

Alexander:1996:MCS

[ALLZ96] F. J. Alexander, C. A. Laberge, J. L. Lebowitz, and R. K. P. Zia. Monte Carlo studies of a driven lattice gas. I. Growth and asymmetry during phase segregation. Journal of Statistical Physics, 82(3–4):1133–1158, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179806.

Avellaneda:1992:SNS

[AM92] Marco Avellaneda and Andrew J. Majda. Superdiffusion in nearly stratified flows. Journal of Statistical Physics, 69(3–4): 689–729, November 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01050431.

Alexander:1994:PLS

[AM94a] Kenneth S. Alexander and Stanislav A. Molchanov. Percolation of level sets for two-dimensional random fields with lattice symmetry. Journal of Statistical Physics, 77(3–4):627–643, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179453.

Angulo:1994:CDR

[AM94b] Rafael F. Angulo and Ernesto Medina. Conductance distributions in random resistor networks. Self-averaging and disorder lengths. Journal of Statistical Physics, 75(1–2):135–151, April 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02186283.

Anzaldo-Meneses:1994:SES

[AM94c] A. M. Anzaldo-Meneses. Shell effects in small metal particles. Journal of Statistical Physics, 75(1–2):297–315, April 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02186290.

Arkeryd:1994:DRB

[AM94d] L. Arkeryd and N. Maslova. On diffuse reflection at the boundary for the Boltzmann equation and related equations. Journal of Statistical Physics, 77(5–6):1051–1077, December 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183152.

Achahbar:1995:PTD

[AM95] A. Achahbar and J. Marro. Phase transitions in a driven lattice gas in two planes. Journal of Statistical Physics, 78(5–6): 1493–1520, March 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02180140.

Adler:1990:LCS

[AMA^{+}90] Joan Adler, Yigal Meir, Amnon Aharony, A. B. Harris, and Lior Klein. Low-concentration series in general dimension. Journal of Statistical Physics, 58(3–4):511–538, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01112760.

Appignanesi:1998:GRD

[AMF98] Gustavo A. Appignanesi, Ruben A. Montani, and Ariel Fernández. Glassy relaxation dynamics and ruggedness beyond the ultrametric limit. Journal of Statistical Physics, 91(3–4):669– 677, May 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023033812314.

Amiran:1996:NGL

[Ami96] Edoh Y. Amiran. Noncoincidence of geodesic lengths and hearing elliptic quantum billiards. Journal of Statistical Physics, 85(3– 4):455–470, November 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02174214.

Abraham:1991:WTR

[AN91a] D. B. Abraham and C. M. Newman. The wetting transition in a random surface model. Journal of Statistical Physics, 63(5– 6):1097–1111, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01030001.

Aertsens:1991:FIP

[AN91b] Marc Aertsens and Jan Naudts. Field-induced percolation in a polarized lattice gas. Journal of Statistical Physics, 62(3–4): 609–630, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01017976.

Ala-Nissila:1993:SEK

[ANHKV93] T. Ala-Nissila, T. Hjelt, J. M. Kosterlitz, and O. Venäläinen. Scaling exponents for kinetic roughening in higher dimensions. Journal of Statistical Physics, 72(1–2):207–225, July 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01048047.

Anonymous:1990:A

[Ano90a] Anonymous. Announcement. Journal of Statistical Physics, 59(5–6):1603, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01334767.

Anonymous:1990:FCBa

[Ano90b] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 58(1–2):405–406, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020305.

Anonymous:1990:FCBb

[Ano90c] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 58(3–4):793–794, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01112775.

Anonymous:1990:FCBc

[Ano90d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 58(5–6):1283–1284, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01026579.

Anonymous:1990:FCBd

[Ano90e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 59(1–2):539–540, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01015585.

Anonymous:1990:FCBe

[Ano90f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 59(3–4):1091–1092, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01025866.

Anonymous:1990:FCBf

[Ano90g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 60(1–2):285–286, July 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01013680.

Anonymous:1990:FCBg

[Ano90h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 60(3–4):527–528, August 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01314935.

Anonymous:1990:FCBh

[Ano90i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 60(5–6):903–904, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01026005.

Anonymous:1990:FCBi

[Ano90j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 61(1–2):507–508, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01013980.

Anonymous:1990:FCBj

[Ano90k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 61(3–4):959–960, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01027317.

Anonymous:1990:FCBk

[Ano90l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 61(5–6):1325–1326, December 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014378.

Anonymous:1990:PSM

[Ano90m] Anonymous. Program of the 63rd Statistical Mechanics Meeting. Journal of Statistical Physics, 60(5–6):895–902, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01026004.

Anonymous:1990:PWL

[Ano90n] Anonymous. Program of the Workshop on Large–Scale Computations in Statistical Physics. Journal of Statistical Physics, 60(5–6):891–893, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01026003.

Anonymous:1990:SPP

[Ano90o] Anonymous. Statistical physics at the 45th parallel: 3rd annual meeting, Clarkson University. Journal of Statistical Physics, 59 (1–2):529–532, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01015582.

Anonymous:1990:TLC

[Ano90p] Anonymous. Third liblice conference on the statistical mechanics of liquids. Journal of Statistical Physics, 61(5–6):1305–1307, December 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014377.

Anonymous:1991:ASM

[Ano91a] Anonymous. Australian Statistical Mechanics Meeting: Mathematics Department, University of Melbourne. Journal of Statistical Physics, 63(1–2):425–426, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01026616.

Anonymous:1991:FCBf

[Ano91b] Anonymous. Future contributions to Journal of Statiscal Physics. Journal of Statistical Physics, 63(5–6):1283–1284, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01030011;   http://link.
springer.com/article/10.1007/BF01030011.

Anonymous:1991:FCBa

[Ano91c] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 62(1–2):507–508, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020892.

Anonymous:1991:FCBb

[Ano91d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 62(3–4):885–886, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017991.

Anonymous:1991:FCBc

[Ano91e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 62(5–6):1271–1272, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01128188.

Anonymous:1991:FCBd

[Ano91f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 63(1–2):431–432, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01026618.

Anonymous:1991:FCBe

[Ano91g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 63(3–4):807–808, May 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01029214.

Anonymous:1991:FCBg

[Ano91h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 64(1–2):479–480, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01057891.

Anonymous:1991:FCBh

[Ano91i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 64(3–4):897–898, August 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048323.

Anonymous:1991:FCBi

[Ano91j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 64(5–6):1163–1164, September 1991. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/BF01048820.

Anonymous:1991:FCBj

[Ano91k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 65(1–2):421–422, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01329874.

Anonymous:1991:FCBk

[Ano91l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 65(3–4):835–836, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01053761.

Anonymous:1991:FCBl

[Ano91m] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 65(5–6):1303–1305, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049613.

Anonymous:1991:FCJ

[Ano91n] Anonymous. Future contributions to journal of statiscal physics. Journal of Statistical Physics, 63(5–6):1283–1284, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01030011.

Anonymous:1991:PSM

[Ano91o] Anonymous. Program of the 64th Statistical Mechanics Meeting. Journal of Statistical Physics, 63(3–4):797–805, May 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01029213.

Anonymous:1991:PW

[Ano91p] Anonymous. Program of the workshop. Journal of Statistical Physics, 62(5–6):1267–1270, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01128187.

Anonymous:1991:SMP

[Ano91q] Anonymous. Statistical mechanics at the 45th parallel: 4th annual meeting. Journal of Statistical Physics, 63(1–2):427–430, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01026617.

Anonymous:1992:ARW

[Ano92a] Anonymous. Advanced research workshop on lattice gas automata theory, implementation, and simulation. Journal of Statistical Physics, 68(3–4):351–352, August 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01341752.

Anonymous:1992:ASM

[Ano92b] Anonymous. Australian statistical mechanics meeting. Journal of Statistical Physics, 67(3–4):839–840, May 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01049733.

Anonymous:1992:B

[Ano92c] Anonymous. Bibliography. Journal of Statistical Physics, 68(3– 4):611–669, August 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01341766.

Anonymous:1992:FCBa

[Ano92d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 66(1–2):687–688, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01060091.

Anonymous:1992:FCBb

[Ano92e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 66(3–4):1187–1188, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01055727.

Anonymous:1992:FCBc

[Ano92f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 66(5–6):1691–1692, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01054447.

Anonymous:1992:FCBd

[Ano92g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 67(1–2):431–432, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01049045.

Anonymous:1992:FCBe

[Ano92h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 67(3–4):841–842, May 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01049734.

Anonymous:1992:FCBf

[Ano92i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 67(5–6):1223–1224, June 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01049018.

Anonymous:1992:FCBg

[Ano92j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 68(1–2):345–346, July 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01048849.

Anonymous:1992:FCBh

[Ano92k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 68(3–4):671–672, August 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01341767.

Anonymous:1992:FCBi

[Ano92l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 68(5–6):1147–1148, September 1992. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/BF01048891.

Anonymous:1992:FCBj

[Ano92m] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 69(1–2):451–452, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01053803.

Anonymous:1992:FCBk

[Ano92n] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 69(3–4):915–916, November 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01050444.

Anonymous:1992:FCBl

[Ano92o] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 69(5–6):1175–1176, December 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01058772.

Anonymous:1992:PQC

[Ano92p] Anonymous. Program of the quantum chaos meeting. Journal of Statistical Physics, 68(1–2):5–6, July 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048835.

Anonymous:1992:SMP

[Ano92q] Anonymous. Statistical mechanics at the 45th parallel: Université d’Ottawa, October 4–5, 1991. Journal of Statistical Physics, 66(5–6):1685–1689, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01054446.

Anonymous:1992:WA

[Ano92r] Anonymous. Workshop announcement. Journal of Statistical Physics, 67(1–2):419, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01049043.

Anonymous:1993:ASM

[Ano93a] Anonymous. Australian statistical mechanics meeting. Journal of Statistical Physics, 71(3–4):825–826, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01058450.

Anonymous:1993:FCBa

[Ano93b] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 70(1–2):513–514, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01053985.

Anonymous:1993:FCBb

[Ano93c] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 70(3–4):1083–1084, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01053612.

Anonymous:1993:FCBc

[Ano93d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 70(5–6):1405–1406, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049444.

Anonymous:1993:FCBd

[Ano93e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 71(1–2):359–360, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01048108.

Anonymous:1993:FCBe

[Ano93f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 71(3–4):837–838, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01058452.

Anonymous:1993:FCBf

[Ano93g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 71(5–6):1237–1238, June 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01049973.

Anonymous:1993:FCBg

[Ano93h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 72(1–2):415–416, July 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01048058.

Anonymous:1993:FCBh

[Ano93i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 72(3–4):855–856, August 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048036.

Anonymous:1993:FCBi

[Ano93j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 72(5–6):1409–1410, September 1993. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/BF01048195.

Anonymous:1993:FCBj

[Ano93k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 73(1–2):459–460, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01052777.

Anonymous:1993:FCBk

[Ano93l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 73(3–4):811, November 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01054356.

Anonymous:1993:FCBl

[Ano93m] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 73(5–6):975–976, December 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01052818.

Anonymous:1993:PMT

[Ano93n] Anonymous. Program of the 5th Midwest Thermodynamics Symposium. Journal of Statistical Physics, 71(1–2):351–353, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048106.

Anonymous:1993:SPP

[Ano93o] Anonymous. Statistical physics at the 45th parallel: 6th annual meeting. Journal of Statistical Physics, 71(1–2):355–357, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048107.

Anonymous:1994:CWA

[Ano94a] Anonymous. CECAM Workshop on Applications of the Random Sequential Addition Process, Orsay, France, June 14–25, 1993. Journal of Statistical Physics, 74(1–2):465–466, January 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02186822.

Anonymous:1994:FCBa

[Ano94b] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 74(1–2):467–468, January 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186823.

Anonymous:1994:FCBb

[Ano94c] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 75(1–2):355–356, April 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02186294.

Anonymous:1994:FCBc

[Ano94d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 75(3–4):779–780, May 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02186884.

Anonymous:1994:FCBd

[Ano94e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 75(5–6):1207–1208, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02186767.

Anonymous:1994:FCBe

[Ano94f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 76(1–2):743–744, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02188684.

Anonymous:1994:FCBf

[Ano94g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 76(3–4):1099–1100, August 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02188702.

Anonymous:1994:FCBg

[Ano94h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 76(5–6):1519–1520, September 1994. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/BF02187077.

Anonymous:1994:FCJ

[Ano94i] Anonymous. Future contributions to journal of statistical physics. Journal of Statistical Physics, 74(3–4):953–954, February 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02188589.

Anonymous:1995:BR

[Ano95a] Anonymous. Book review. Journal of Statistical Physics, 79 (3–4):791–792, May 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02184886.

Anonymous:1995:CLO

[Ano95b] Anonymous. Chemistry 1968. Lars Onsager. Journal of Statistical Physics, 78(1–2):641–660, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183372.

Anonymous:1995:FLC

[Ano95c] Anonymous. Fourth liblice conference on the statistical mechanics of liquids. Journal of Statistical Physics, 78(5–6):1637–1640, March 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02180149.

Anonymous:1995:FCBa

[Ano95d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 79(1–2):501–502, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02179403.

Anonymous:1995:FCBb

[Ano95e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 79(3–4):797–798, May 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02184888.

Anonymous:1995:FCBc

[Ano95f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 79(5–6):1033–1034, June 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02181216.

Anonymous:1995:FCBd

[Ano95g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 81(1–2):537–538, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179992.

Anonymous:1995:FCBe

[Ano95h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 81(5–6):1037–1038, December 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179302.

Anonymous:1995:LOS

[Ano95i] Anonymous. The lars Onsager symposium. Journal of Statistical Physics, 78(1–2):4–6, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02183336.

Anonymous:1996:FCBa

[Ano96a] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 82(3–4):1217–1218, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179811.

Anonymous:1996:FCBb

[Ano96b] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 82(5–6):1679–1680, March 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183401.

Anonymous:1996:FCBc

[Ano96c] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 82(5–6):1679–1680, March 1996. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF02183401;   http://link.
springer.com/article/10.1007/BF02183401.

Anonymous:1996:FCBd

[Ano96d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 83(1–2):289–290, April 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02183651.

Anonymous:1996:FCBe

[Ano96e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 83(3–4):793–794, May 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02183752.

Anonymous:1996:FCBf

[Ano96f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 85(1–2):295–296, October 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02175568.

Anonymous:1996:FCBg

[Ano96g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 85(3–4):523, November 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02174221.

Anonymous:1996:M

[Ano96h] Anonymous. In memoriam. Journal of Statistical Physics, 85(5– 6):799–805, December 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02199363.

Anonymous:1996:P

[Ano96i] Anonymous. Preface. Journal of Statistical Physics, 83(1–2):1–2, April 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183635.

Anonymous:1996:PSS

[Ano96j] Anonymous. Program of the Second Statistical Physics Days Istanbul Technical University 13-14 July 1995. Journal of Statistical Physics, 83(5–6):1267–1270, June 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179566.

Anonymous:1996:SCQ

[Ano96k] Anonymous. Symposium on classical and quantum billiards. Journal of Statistical Physics, 83(1–2):3–5, April 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183636.

Anonymous:1997:FCBa

[Ano97a] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 86(3–4):905–906, February 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02199124.

Anonymous:1997:FCBb

[Ano97b] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 86(5–6):1405–1406, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183634.

Anonymous:1997:FCBc

[Ano97c] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 88(1–2):535–536, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02508486.

Anonymous:1997:FCBd

[Ano97d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 88(3–4):997–998, August 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000015506.46239.f2.

Anonymous:1997:FCJ

[Ano97e] Anonymous. Future contributions to journal of statistical physics. Journal of Statistical Physics, 88(5–6):1425–1426, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/BF02732446.

Anonymous:1997:PHD

[Ano97f] Anonymous. Program of the Hyperbolic Dynamics and Applications to Nonequilibrium Statistical Mechanics Meeting. Journal of Statistical Physics, 87(1–2):469–470, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02181500.

Anonymous:1997:PSP

[Ano97g] Anonymous. Program of the Statistical Physics at the 45th Parallel Meeting. Journal of Statistical Physics, 87(3–4):967–969, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02181263.

Anonymous:1997:PTS

[Ano97h] Anonymous. Program of the Third Statistical Physics Days. Journal of Statistical Physics, 86(5–6):1401–1404, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02183633.

Anonymous:1998:AIJ

[Ano98a] Anonymous. Author index for journal of statistical physics (1998). Journal of Statistical Physics, 93(5–6):1169–1182, December 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000033226.16718.41.

Anonymous:1998:FCBa

[Ano98b] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 90(1–2):517–518, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023236725653.

Anonymous:1998:FCBb

[Ano98c] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 90(3–4):1073–1074, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023214128521.

Anonymous:1998:FCBc

[Ano98d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 90(5–6):1505–1506, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023268403812.

Anonymous:1998:FCBd

[Ano98e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 90(5–6):1505–1506, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1023268403812.

Anonymous:1998:FCBe

[Ano98f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 91(1–2):497–498, April 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1023025112373.

Anonymous:1998:FCBf

[Ano98g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 91(3–4):827–828, May 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1023050215948.

Anonymous:1998:FCBg

[Ano98h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 91(5–6):1063–1064, June 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1023044403286.

Anonymous:1998:FCBh

[Ano98i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 92(1–2):335–336, July 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1023064106941.

Anonymous:1998:FCBi

[Ano98j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 92(3–4):727–728, August 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023005108623.

Anonymous:1998:FCBj

[Ano98k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 92(5–6):1225–1226, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1023073417050.

Anonymous:1998:FCBk

[Ano98l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 93(1–2):405–406, October 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1023/B%3AJOSS.0000026859.63373.f8.pdf.

Anonymous:1998:FCBl

[Ano98m] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 93(3–4):1019, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000033229.10606.9d.

Anonymous:1998:FCBm

[Ano98n] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 93(5–6):1183–1184, December 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000033227.38970.97.

Anonymous:1998:MMM

[Ano98o] Anonymous. Micro, macro, meta: A workshop to welcome Joel Lebwitz. Journal of Statistical Physics, 90(5–6):1501–1503, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023216419742.

Anonymous:1998:PSM

[Ano98p] Anonymous. Program of the 78th Statistical Mechanics Meeting. Journal of Statistical Physics, 91(1–2):485–495, April 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023073028303.

Anonymous:1998:PFS

[Ano98q] Anonymous. Program of the Fourth Statistical Physics Days. Journal of Statistical Physics, 90(1–2):511–515, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023284608815.

Anonymous:1999:AIJ

[Ano99a] Anonymous. Author index for journal of statistical physics (1999). Journal of Statistical Physics, 97(5–6):1033–1044, December 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017222332682.

Anonymous:1999:BRBh

[Ano99b] Anonymous. Book review: Classical and Quantum Dynamics in Condensed Phase Simulations. Journal of Statistical Physics, 97 (1–2):419–420, October 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1017258823035.

Anonymous:1999:BRBa

[Ano99c] Anonymous. Book review: Master of Modern Physics. The Scientific Contributions of H. A. Kramers. Journal of Statistical Physics, 94(1–2):269–274, January 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1017248015610.

Anonymous:1999:BRBb

[Ano99d] Anonymous. Book review: Modern Thermodynamics: From Heat Engines to Dissipative Structures. Journal of Statistical Physics, 94(5–6):1055–1056, March 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1017224020327.

Anonymous:1999:BRBj

[Ano99e] Anonymous. Book review: Monte Carlo Methods in Statistical Physics. Journal of Statistical Physics, 97(5–6):1029–1030, December 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017253831773.

Anonymous:1999:BRBg

[Ano99f] Anonymous. Book review: Nonequilibrium Statistical Mechanics. Ensemble Method. Journal of Statistical Physics, 95(3–4):811– 813, May 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017268031424.

Anonymous:1999:BRBd

[Ano99g] Anonymous. Book review: Nonequilibrium Statistical Mechanics in One Dimension. Journal of Statistical Physics, 95(1– 2):511–512, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1017206317795.

Anonymous:1999:BRBf

[Ano99h] Anonymous. Book review: The Self-Avoiding Walk. Journal of Statistical Physics, 95(3–4):805–809, May 1999. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1017266514586.

Anonymous:1999:BRBe

[Ano99i] Anonymous. Book review: Theoretical and Mathematical Models in Polymer Research, Modern Methods in Polymer Research and Technology. Journal of Statistical Physics, 95(1–2):513–516, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017258301865.

Anonymous:1999:BRBi

[Ano99j] Anonymous. Book review: Thermodynamics of One-Dimensional Solvable Models. Journal of Statistical Physics, 97(3–4):827– 828, November 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1017286429531.

Anonymous:1999:BRC

[Ano99k] Anonymous. Book review: Cellular automata modeling of physical systems. Journal of Statistical Physics, 97(5–6):1031– 1032, December 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1017270215844.

Anonymous:1999:BRS

[Ano99l] Anonymous. Book review: Statistical physics of fracture and breakdown in disordered systems. Journal of Statistical Physics, 94(5–6):1057–1059, March 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017251604398.

Anonymous:1999:BRBc

[Ano99m] Anonymous. Book reviews: Books on complexity. Journal of Statistical Physics, 95(1–2):507–510, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1017259500957.

Anonymous:1999:D

[Ano99n] Anonymous. Dedication. Journal of Statistical Physics, 95(5–6): 825–826, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1017238014103;   http:
//link.springer.com/article/10.1023/A%3A1017238014103.

Anonymous:1999:FCBa

[Ano99o] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 94(1–2):275–276, January 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017275632449.

Anonymous:1999:FCBb

[Ano99p] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 94(3–4):723–724, February 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017227825605.

Anonymous:1999:FCBc

[Ano99q] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 94(5–6):1065–1066, March 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017255705306.

Anonymous:1999:FCBd

[Ano99r] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 95(1–2):1065–1066, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1017214519612.

Anonymous:1999:FCBe

[Ano99s] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 95(3–4):823–824, May 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1017272132333.

Anonymous:1999:FCBf

[Ano99t] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 95(5–6):1505–1506, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1017290009076.

Anonymous:1999:FCBg

[Ano99u] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 95(5–6):1505–1506, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017290009076.

Anonymous:1999:FCBh

[Ano99v] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 96(1–2):459–460, July 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1017291505867.

Anonymous:1999:FCBi

[Ano99w] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 96(3–4):913–914, August 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017201027925.

Anonymous:1999:FCBj

[Ano99x] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 96(5–6):1365–1366, September 1999. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1017255321817.

Anonymous:1999:FCBk

[Ano99y] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 97(1–2):427–428, October 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017291723943.

Anonymous:1999:FCBl

[Ano99z] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 97(3–4):829, November 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017235813601.

Anonymous:1999:FCBm

[Ano99-27] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 97(5–6):1045–1046, December 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017274316752.

Anonymous:1999:PSMa

[Ano99-28] Anonymous. Program of the 80th Statistical Mechanics Meeting. Journal of Statistical Physics, 95(3–4):815–821, May 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1017220115495.

Anonymous:1999:PSMb

[Ano99-29] Anonymous. Program of the 81st Statistical Mechanics Meeting. Journal of Statistical Physics, 97(1–2):0, October 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1017239707105.

Anonymous:1999:PFS

[Ano99-30] Anonymous. Program of the Fifth Statistical Physics Day. Journal of Statistical Physics, 94(5–6):1061–1064, March 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1017203721236.

Anonymous:1999:SMR

[Ano99-31] Anonymous. Statistical mechanics and related topics in mathematical physics. Journal of Statistical Physics, 95(1–2):519–521, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017210402774.

Anonymous:2001:ESL

[Ano01] Anonymous. Erratum on ‘A Strong Law of Large Numbers for Iterated Functions of Independent Random Variables,’ Jan Wehr, J. Stat. Phys. 86:1373 (1997). Journal of Statistical Physics, 104(3–4):901, August 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1017339609610. See [Weh97c].

Anishchenko:1993:SRC

[ANS93] V. S. Anishchenko, A. B. Neiman, and M. A. Safanova. Stochastic resonance in chaotic systems. Journal of Statistical Physics, 70(1–2):183–196, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01053962.

Ala-Nissila:1994:SED

[ANV94] T. Ala-Nissila and O. Venäläinen. Scaling exponents for driven two-dimensional surface growth. Journal of Statistical Physics, 76(3–4):1083–1088, August 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02188700.

Appert:1995:PST

[AORZ95] Cécile Appert, John F. Olson, Daniel H. Rothman, and Stéphane Zaleski. Phase separation in a three-dimensional, two-phase, hydrodynamic lattice gas. Journal of Statistical Physics, 81(1–2): 181–197, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179975.

Almirantis:1999:LSR

[AP99] Y. Almirantis and A. Provata. Long- and short-range correlations in genome organization. Journal of Statistical Physics, 97(1– 2):233–262, October 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004671119400.

Aslangul:1992:RRW

[APC^{+}92] C. Aslangul, N. Pottier, P. Chvosta, D. Saint-James, and L. Skála. Random-random walk on an asymmetric chain with a trapping attractive center. Journal of Statistical Physics, 69 (1–2):17–34, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053780.

Angelescu:1994:DCL

[APT94] N. Angelescu, M. Pulvirenti, and A. Teta. Derivation and classical limit of the mean-field equation for a quantum Coulomb system: Maxwell–Boltzmann statistics. Journal of Statistical Physics, 74(1–2):147–165, January 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02186811.

Aidun:1998:NMA

[AQ98] Cyrus K. Aidun and Dewei W. Qi. A new method for analysis of the fluid interaction with a deformable membrane. Journal of Statistical Physics, 90(1–2):145–158, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023299617476.

Abraham:1991:FSE

[AS91a] D. B. Abraham and N. M. Svrakić. Finite-size effects in surface tension. I. Fluctuating interfaces. Journal of Statistical Physics, 63(5–6):1077–1096, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01030000.

Albeverio:1991:WCQ

[AS91b] S. Albeverio and P. Seba. Wave chaos in quantum systems with point interaction. Journal of Statistical Physics, 64(1–2):369– 383, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01057882.

Aleksandrov:1995:AGS

[AS95] O. E. Aleksandrov and V. D. Seleznev. Acoustic gas slip induced by surface waves. Journal of Statistical Physics, 78(1–2):161–167, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183344.

Anderle:1995:PSN

[ASKK95] Martin Anderle, Herwig Schweng, Karl E. Kürten, and Karl W. Kratky. Pattern-specific neural network design. Journal of Statistical Physics, 81(3–4):843–849, November 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179261.

Aslangul:1999:MAT

[Asl99] Claude Aslangul. Mutual Annihilation of two diffusing particles in one- and two-dimensional lattices. Journal of Statistical Physics, 94(1–2):219–240, January 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004567430632.

Abenda:1990:DPS

[AT90] S. Abenda and G. Turchetti. Duality in parameter space and approximation of measures for mixing repellers. Journal of Statistical Physics, 61(1–2):293–310, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01013966.

Aurell:1990:CDZ

[Aur90] Erik Aurell. Convergence of dynamical zeta functions. Journal of Statistical Physics, 58(5–6):967–995, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01026559.

Appert:1997:TFL

[AvBED97] C. Appert, H. van Beijeren, M. H. Ernst, and J. R. Dorfman. Thermodynamic formalism and localization in Lorentz gases and hopping models. Journal of Statistical Physics, 87 (5–6):1253–1271, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181283.

Avron:1998:OV

[Avr98] J. E. Avron. Odd viscosity. Journal of Statistical Physics, 92 (3–4):543–557, August 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023084404080.

Agmon:1990:BR

[AW90a] Noam Agmon and George H. Weiss. Book reviews. Journal of Statistical Physics, 58(1–2):397–401, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020303.

Alcaraz:1990:HXH

[AW90b] F. C. Alcaraz and W. F. Wreszinski. The Heisenberg XXZ Hamiltonian with Dzyaloshinsky–Moriya interactions. Journal of Statistical Physics, 58(1–2):45–56, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020284.

Au-Yang:1995:CPM

[AYP95] Helen Au-Yang and Jacques H. H. Perk. The chiral Potts models revisited. Journal of Statistical Physics, 78(1–2):17–78, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183338.

Albeverio:1995:RRI

[AZ95] Sergio Albeverio and Xian Yin Zhou. A renormalization result for the intersection local time of lattice random walks in d ≥ 3 dimensions. Journal of Statistical Physics, 80(3–4):603–624, August 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02178550.

Albeverio:1996:FES

[AZ96] Sergio Albeverio and Xian Yin Zhou. Free energy and some sample path properties of a random walk with random potential. Journal of Statistical Physics, 83(3–4):573–622, May 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02183741.

Avron:1998:GFP

[AZ98] J. E. Avron and P. G. Zograf. Geometric forces on point fluxes in quantum Hall fluids. Journal of Statistical Physics, 92(5–6): 1193–1201, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023013231163.

Adl-Zarabi:1997:EME

[AZP97] Kourosh Adl-Zarabi and Harald Proppe. Existence of many ergodic absolutely continuous invariant measures for piecewiseexpanding C^{2} chaotic transformations in R^{2} on a fixed number of partitions. Journal of Statistical Physics, 89(3–4):537– 548, November 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02765534.

Ben-Abraham:1993:CPS

[BA93] Shelomo I. Ben-Abraham. Curious properties of simple random walks. Journal of Statistical Physics, 73(1–2):441–445, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01052773.

ben-Avraham:1990:SDD

[bABD90] Daniel ben Avraham, Martin A. Burschka, and Charles R. Doering. Statics and dynamics of a diffusion-limited reaction: Anomalous kinetics, nonequilibrium self-ordering, and a dynamic transition. Journal of Statistical Physics, 60(5–6):695–728, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01025990.

Bagnoli:1996:DST

[Bag96] Franco Bagnoli. On damage-spreading transitions. Journal of Statistical Physics, 85(1–2):151–164, October 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02175559.

Ben-Avraham:1991:DTM

[BAK91] D. Ben-Avraham and J. Köhler. The dimer-trimer model for heterogeneous catalysis. Journal of Statistical Physics, 65(5–6): 839–848, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049584.

Baker:1993:MPM

[Bak93] George A. Baker, Jr. A Markov-property Monte Carlo method: One-dimensional Ising model. Journal of Statistical Physics, 72 (3–4):621–641, August 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048026.

Baker:1994:MPM

[Bak94] George A. Baker, Jr. The Markov property method applied to Ising model calculations. Journal of Statistical Physics, 77(5–6): 955–976, December 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183146.

Baker:1998:IMB

[Bak98] George A. Baker, Jr. Ising-model, block-spin distributions by the Markov property method. Journal of Statistical Physics, 93(3– 4):573–582, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000033242.95755.4e.

Ben-Av:1990:CAL

[BAKK^{+}90] R. Ben-Av, D. Kandel, E. Katznelson, P. G. Lauwers, and S. Solomon. Critical acceleration of lattice gauge simulations. Journal of Statistical Physics, 58(1–2):125–139, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020288.

Baladi:1991:GSE

[Bal91] V. Baladi. Gibbs states and equilibrium states for finitely presented dynamical systems. Journal of Statistical Physics, 62(1– 2):239–256, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020868.

Baldwin:1992:CEM

[Bal92a] P. R. Baldwin. A convergence exponent for multidimensional continued-fraction algorithms. Journal of Statistical Physics, 66(5–6):1507–1526, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01054431.

Baldwin:1992:MCF

[Bal92b] P. R. Baldwin. A multidimensional continued fraction and some of its statistical properties. Journal of Statistical Physics, 66(5– 6):1463–1505, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01054430.

Bandt:1999:GPS

[Ban99] Christoph Bandt. The geometry of a parameter space of interacting particle systems. Journal of Statistical Physics, 96(3– 4):883–906, August 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004614827017.

Bhatia:1993:FPT

[BAP93] D. P. Bhatia, D. Arora, and M. A. Prasad. First passage time for a class of one-dimensional stochastic systems. Journal of Statistical Physics, 71(5–6):1191–1200, June 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01049967.

Barrett:1990:DJM

[Bar90] A. J. Barrett. On the Domb–Joyce model for self-avoiding walks in the continuum. Journal of Statistical Physics, 58(3– 4):617–626, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01112766.

Barcilon:1996:EOD

[Bar96] Victor Barcilon. Eigenvalues of the one-dimensional Smoluchowski equation. Journal of Statistical Physics, 82(1–2): 267–296, January 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02189231.

Batchelor:1998:LML

[Bat98] M. T. Batchelor. The O(n) loop model on the 3–12 lattice. Journal of Statistical Physics, 92(5–6):1203–1208, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023065215233.

Bavaud:1994:TDU

[Bav94] François Bavaud. Triangular dynamics under pressure. Journal of Statistical Physics, 76(1–2):645–660, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02188679.

Baxter:1991:CTM

[Bax91] R. J. Baxter. Corner transfer matrices of the chiral Potts model. Journal of Statistical Physics, 63(3–4):433–453, May 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01029194.

Baxter:1993:CPM

[Bax93a] R. J. Baxter. Chiral Potts model with skewed boundary conditions. Journal of Statistical Physics, 73(3–4):461–495, November 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01054336.

Baxter:1993:CTM

[Bax93b] R. J. Baxter. Corner transfer matrices of the chiral Potts model. II. The triangular lattice. Journal of Statistical Physics, 70(3– 4):535–582, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053584.

Baxter:1995:SMS

[Bax95] R. J. Baxter. Solvable models in statistical mechanics, from Onsager onward. Journal of Statistical Physics, 78(1–2):7–16, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183337.

Baxter:1996:FEC

[Bax96] R. J. Baxter. Free energy of the chiral Potts model in the scaling region. Journal of Statistical Physics, 82(5–6):1219–1234, March 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183382.

Baxter:1998:FRO

[Bax98] R. J. Baxter. Functional relations for the order parameters of the chiral Potts model. Journal of Statistical Physics, 91(3–4):499– 524, May 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023096408679.

Bao:1998:IAN

[BAZ98] Jingdong Bao, Yasuhisa Abe, and Yizhong Zhuo. An integral algorithm for numerical integration of one-dimensional additive colored noise problems. Journal of Statistical Physics, 90(3–4): 1037–1045, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023201725795.

Baldan:1991:CFF

[BB91] Oscar Baldan and Giancarlo Benettin. Classical ‘freezing’ of fast rotations. A numerical test of the Boltzmann–Jeans conjecture. Journal of Statistical Physics, 62(1–2):201–219, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020866.

Bazhanov:1992:NSL

[BB92a] V. V. Bazhanov and R. J. Baxter. New solvable lattice models in three dimensions. Journal of Statistical Physics, 69(3–4): 453–485, November 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01050423.

Brieger:1992:SLG

[BB92b] Leesa Brieger and Ernesto Bonomi. A stochastic lattice gas for Burgers’ equation: A practical study. Journal of Statistical Physics, 69(3–4):837–855, November 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01050436.

Bazhanov:1993:STR

[BB93] V. V. Bazhanov and R. J. Baxter. Star-triangle relation for a three-dimensional model. Journal of Statistical Physics, 71 (5–6):839–864, June 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049952.

Bormann:1994:PFO

[BB94] Dierk Bormann and Hans Beck. Possible first-order transition in the two-dimensional Ginzburg–Landau model induced by thermally fluctuating vortex cores. Journal of Statistical Physics, 76 (1–2):361–395, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02188667.

Barak:1995:MGS

[BB95a] Liana Barak and Carol Braester. Mathematical games and sampling inspection plans. Journal of Statistical Physics, 79 (3–4):775–787, May 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02184884.

Blum:1995:GSB

[BB95b] L. Blum and O. Bernard. The general solution of the binding mean spherical approximation for pairing ions. Journal of Statistical Physics, 79(3–4):569–583, May 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02184871.

Bussemaker:1997:TDL

[BB97] Harmen J. Bussemaker and Ricardo Brito. Theory for diffusionlimited oscillating chemical reactions. Journal of Statistical Physics, 87(5–6):1165–1178, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02181278.

Brassesco:1998:IFS

[BB98] Stella Brassesco and Paolo Buttà. Interface fluctuations for the D = 1 stochastic Ginzburg–Landau equation with nonsymmetric reaction term. Journal of Statistical Physics, 93(5–6): 1111–1142, December 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000033154.54515.e8.

Blanchard:1994:LSN

[BBC^{+}94a] Ph. Blanchard, G. Bolz, M. Cini, G. F. De Angelis, and M. Serva. Localization stabilized by noise. Journal of Statistical Physics, 75(3–4):749–755, May 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186879.

Bokhove:1994:EPM

[BBC94b] O. Bokhove, C. Bruin, and A. Compagner. Ensemble properties and molecular dynamics of unstable systems. Journal of Statistical Physics, 74(1–2):55–73, January 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02186806.

Boldrighini:1995:ITT

[BBC^{+}95] C. Boldrighini, L. A. Bunimovich, G. Cosimi, S. Frigio, and A. Pellegrinotti. Ising-type transitions in coupled map lattices. Journal of Statistical Physics, 80(5–6):1185–1205, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179868.

Bertin:1999:EDP

[BBD99] Etienne Bertin, Jean-Michel Billiot, and Rémy Drouilhet. Existence of Delaunay pairwise Gibbs point process with superstable component. Journal of Statistical Physics, 95(3–4):719– 744, May 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004551527790.

Biferale:1994:CCK

[BBF94] L. Biferale, M. Blank, and U. Frisch. Chaotic cascades with Kolmogorov 1941 scaling. Journal of Statistical Physics, 75(5– 6):781–795, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186743.

Barkema:1998:TDO

[BBG98] G. T. Barkema, U. Bastolla, and P. Grassberger. Twodimensional oriented self-avoiding walks with parallel contacts. Journal of Statistical Physics, 90(5–6):1311–1324, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023287513382.

Bogomolny:1996:QCD

[BBL96] E. Bogomolny, O. Bohigas, and P. Leboeuf. Quantum chaotic dynamics and random polynomials. Journal of Statistical Physics, 85(5–6):639–679, December 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02199359.

Binder:1992:TSA

[BBM92] P.-M. Binder, B. Buck, and V. A. Macaulay. Time-series analysis of a collective variable in high-dimensional cellular automata. Journal of Statistical Physics, 68(5–6):1127–1130, September 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01048888.

Bhaduri:1996:HES

[BBM96] R. K. Bhaduri, R. S. Bhalerao, and M. V. N. Murthy. Haldane exclusion statistics and the Boltzmann equation. Journal of Statistical Physics, 82(5–6):1659–1668, March 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183398.

Briozzo:1991:EAP

[BBOC91] Carlos B. Briozzo, Carlos E. Budde, Omar Osenda, and Manuel O. Cáceres. Exact and asymptotic properties of multistate random walks. Journal of Statistical Physics, 65(1–2): 167–182, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01329854.

Bernoff:1998:ASD

[BBW98] Andrew J. Bernoff, Andrea L. Bertozzi, and Thomas P. Witelski. Axisymmetric surface diffusion: Dynamics and stability of self-similar pinchoff. Journal of Statistical Physics, 93(3–4): 725–776, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000033251.81126.af.

Ball:1990:DCF

[BC90a] J. M. Ball and J. Carr. The discrete coagulation-fragmentation equations: Existence, uniqueness, and density conservation. Journal of Statistical Physics, 61(1–2):203–234, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013961.

Brey:1990:GLE

[BC90b] J. J. Brey and J. Casado. Generalized Langevin equations with time-dependent temperature. Journal of Statistical Physics, 61(3–4):713–722, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01027298.

Belinicher:1992:FIE

[BC92] V. I. Belinicher and M. V. Chertkov. Functional integral and effective Hamiltonian t–J–V model of strongly correlated electron system. Journal of Statistical Physics, 69(1–2):231–245, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01053792.

Bonomi:1994:MSD

[BC94] Ernesto Bonomi and Gabriella Cabitza. Migration of seismic data. Journal of Statistical Physics, 76(1–2):703–723, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02188682.

Bertini:1995:SHE

[BC95] Lorenzo Bertini and Nicoletta Cancrini. The stochastic heat equation: Feynman–Kac formula and intermittence. Journal of Statistical Physics, 78(5–6):1377–1401, March 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02180136.

Borgs:1996:CMP

[BC96] C. Borgs and J. T. Chayes. The covariance matrix of the Potts model: A random cluster analysis. Journal of Statistical Physics, 82(5–6):1235–1297, March 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183383.

Baker:1998:UDT

[BC98a] T. Baker and L. Chayes. On the unicity of discontinuous transitions in the two-dimensional Potts and Ashkin–Teller models. Journal of Statistical Physics, 93(1–2):1–15, October 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1023/B%3AJOSS.0000026725.69382.86.pdf.

Bobylev:1998:DVM

[BC98b] Alexander V. Bobylev and Carlo Cercignani. Discrete velocity models for mixtures. Journal of Statistical Physics, 91(1– 2):327–341, April 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023052423760.

Bates:1999:IMP

[BC99a] Peter W. Bates and Adam Chmaj. An integrodifferential model for phase transitions: Stationary solutions in higher space dimensions. Journal of Statistical Physics, 95(5–6):1119–1139, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004514803625.

Bobylev:1999:DVM

[BC99b] Alexander V. Bobylev and Carlo Cercignani. Discrete velocity models without nonphysical invariants. Journal of Statistical Physics, 97(3–4):677–686, November 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004615309058.

Bobylev:1999:REP

[BC99c] Alexander V. Bobylev and Carlo Cercignani. On the rate of entropy production for the Boltzmann equation. Journal of Statistical Physics, 94(3–4):603–618, February 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004556522879.

Berretti:1992:NBA

[BCCF92] A. Berretti, A. Celletti, L. Chierchia, and C. Falcolini. Natural boundaries for area-preserving twist maps. Journal of Statistical Physics, 66(5–6):1613–1630, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01054437.

Benedetto:1998:NMS

[BCCP98] D. Benedetto, E. Caglioti, J. A. Carrillo, and M. Pulvirenti. A non–Maxwellian steady distribution for one-dimensional granular media. Journal of Statistical Physics, 91(5–6):979–990, June 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023032000560.

Boldrighini:1990:DER

[BCF90] C. Boldrighini, G. C. Cosimi, and S. Frigio. Diffusion and Einstein relation for a massive particle in a one-dimensional free fas: Numerical evidence. Journal of Statistical Physics, 59(5– 6):1241–1250, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01334749.

Borgs:1997:DSC

[BCF97] C. Borgs, J. T. Chayes, and J. Fröhlich. Dobrushin states for classical spin systems with complex interactions. Journal of Statistical Physics, 89(5–6):895–928, December 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02764213.

Bercovici:1995:EDP

[BCFM95] H. Bercovici, P. Constantin, C. Foias, and O. P. Manley. Exponential decay of the power spectrum of turbulence. Journal of Statistical Physics, 80(3–4):579–602, August 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02178549.

Blanchard:1997:DSA

[BCK97] Ph. Blanchard, B. Cessac, and T. Krüger. A dynamical system approach to SOC models of Zhang’s type. Journal of Statistical Physics, 88(1–2):307–318, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02508473.

Benichou:1999:BDO

[BCL^{+}99] O. Bénichou, A. M. Cazabat, A. Lemarchand, M. Moreau, and G. Oshanin. Biased diffusion in a one-dimensional adsorbed monolayer. Journal of Statistical Physics, 97(1–2):351–371, October 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004679421218.

Bertini:1999:RGT

[BCO99] Lorenzo Bertini, Emilio N. M. Cirillo, and Enzo Olivieri. Renormalization-group transformations under strong mixing conditions: Gibbsianness and convergence of renormalized interactions. Journal of Statistical Physics, 97(5–6):831–915, December 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004620929047.

Bellettini:1996:CMN

[BCP96] G. Bellettini, M. Cassandro, and E. Presutti. Constrained minima of nonlocal free energy functionals. Journal of Statistical Physics, 84(5–6):1337–1349, September 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02174133.

Bonetto:1998:VCP

[BCP98] F. Bonetto, E. G. D. Cohen, and C. Pugh. On the validity of the conjugate pairing rule for Lyapunov exponents. Journal of Statistical Physics, 92(3–4):587–627, August 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023040621826.

Biferale:1997:ISP

[BCPV97] L. Biferale, M. Cencini, D. Pierotti, and A. Vulpiani. Intermittency in stochastically perturbed turbulent models. Journal of Statistical Physics, 88(5–6):1117–1138, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02732427.

Bender:1991:NPE

[BCS^{+}91] Carl Bender, Fred Cooper, L. M. Simmons, Jr., Pinaki Roy, and Greg Kilcup. Novel perturbation expansion for the Langevin equation. Journal of Statistical Physics, 64(1–2):395–428, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01057884.

Benettin:1993:LTA

[BCS93] Giancarlo Benettin, Andrea Carati, and Paolo Sempio. On the Landau–Teller approximation for energy exchanges with fast degrees of freedom. Journal of Statistical Physics, 73(1–2):175–192, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01052756.

Blawzdziewicz:1992:TLK

[BCvB92] J. Bławzdziewicz, B. Cichocki, and H. van Beijeren. H-theorem for a linear kinetic theory. Journal of Statistical Physics, 66(1– 2):607–633, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01060084.

Brankov:1993:LFS

[BD93] Jordan G. Brankov and Daniel M. Danchev. Logarithmic finite-size corrections in the three-dimensional mean spherical model. Journal of Statistical Physics, 71(3–4):775–798, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01058447.

Barrett:1994:VEP

[BD94] A. J. Barrett and C. Domb. Virial expansion for a polymer with a realistic pair-potential interaction. Journal of Statistical Physics, 77(1–2):491–500, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02186854.

Bleher:1997:CBZ

[BD97] Pavel Bleher and Xiaojun Di. Correlations between zeros of a random polynomial. Journal of Statistical Physics, 88(1–2):269– 305, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02508472.

Berlier:1990:MWC

[BDDM90] Karl Berlier, Joël De Coninck, François Dunlop, and Frédéric Menu. Multilayer wetting in clock models. Journal of Statistical Physics, 61(1–2):179–186, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01013959.

Bardos:1997:DAB

[BDG97] Claude Bardos, Laurent Dumas, and François Golse. Diffusion approximation for billiards with totally accommodating scatterers. Journal of Statistical Physics, 86(1–2):351–375, January 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02180210.

Bobylev:1997:VME

[BDIV97] A. V. Bobylev, P. Dukes, R. Illner, and H. D. Victory, Jr. On Vlasov–Manev equations. I: Foundations, properties, and nonglobal existence. Journal of Statistical Physics, 88(3–4):885–911, August 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000015177.60491.3c.

Borgs:1999:ELM

[BDK99] C. Borgs, J. DeConinck, and R. Kotecky’. An equilibrium lattice model of wetting on rough substrates. Journal of Statistical Physics, 94(3–4):299–320, February 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004531917428.

Bergman:1990:DNM

[BDM90] David J. Bergman, Edgardo Duering, and Michael Murat. Discrete network models for the low-field Hall effect near a percolation threshold: Theory and simulations. Journal of Statistical Physics, 58(1–2):1–43, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020283.

Brey:1997:DDH

[BDS97] J. Javier Brey, James W. Dufty, and Andrés Santos. Dissipative dynamics for hard spheres. Journal of Statistical Physics, 87(5– 6):1051–1066, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181270.

Brey:1999:KMG

[BDS99] J. Javier Brey, James W. Dufty, and Andrés Santos. Kinetic models for granular flow. Journal of Statistical Physics, 97(1– 2):281–322, October 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004675320309.

Bussemaker:1992:BLG

[BE92] H. J. Bussemaker and M. H. Ernst. Biased lattice gases with correlated equilibrium states. Journal of Statistical Physics, 68 (3–4):431–455, August 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01341757.

Bartelt:1994:SSC

[BE94] M. C. Bartelt and J. W. Evans. Scaling of spatial correlations in cooperative sequential adsorption with clustering. Journal of Statistical Physics, 76(3–4):867–876, August 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02188690.

Benisti:1998:FRL

[BE98] D. Bénisti and D. F. Escande. Finite range of large perturbations in Hamiltonian dynamics. Journal of Statistical Physics, 92(5– 6):909–972, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023092526620.

Busoni:1999:SST

[BE99] G. Busoni and H. Emamirad. Stationary scattering theory for a charged particles transport problem. Journal of Statistical Physics, 96(1–2):377–401, July 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004536803141.

Baranyai:1993:FDC

[BEC93] András Baranyai, Denis J. Evans, and E. G. D. Cohen. Fielddependent conductivity and diffusion in a two-dimensional Lorentz gas. Journal of Statistical Physics, 70(5–6):1085–1098, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049423.

Beck:1995:PFE

[Bec95] Christian Beck. From the Perron–Frobenius equation to the Fokker–Planck equation. Journal of Statistical Physics, 79(5– 6):875–894, June 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181207.

Bussemaker:1995:GBE

[BED95] H. J. Bussemaker, M. H. Ernst, and J. W. Dufty. Generalized Boltzmann equation for lattice gas automata. Journal of Statistical Physics, 78(5–6):1521–1554, March 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02180141.

Brito:1991:SDL

[BEK91] R. Brito, M. H. Ernst, and T. R. Kirkpatrick. Staggered diffusivities in lattice gas cellular automata. Journal of Statistical Physics, 62(1–2):283–295, January 1991. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020871.

Belitsky:1993:AUB

[Bel93a] V. Belitsky. Asymptotic upper bound of density for two-particle annihilating exclusion. Journal of Statistical Physics, 73(3–4): 671–694, November 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01054345.

Belitsky:1993:SMD

[Bel93b] Vladimir Belitsky. A stochastic model of deposition processes with nucleation. Journal of Statistical Physics, 70(5–6): 1233–1254, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049430.

Belitsky:1995:ABD

[Bel95] V. Belitsky. Asymptotic behavior of the density for two-particle annihilating exclusion. Journal of Statistical Physics, 78(3–4): 937–961, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183695.

Benois:1999:NSL

[BEM99] O. Benois, R. Esposito, and R. Marra. Navier–Stokes limit for a thermal stochastic lattice gas. Journal of Statistical Physics, 96 (3–4):653–713, August 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004618924291.

Bendisch:1991:GLT

[Ben91] J. Bendisch. Geometric localization of the threshold in twodimensional Ising ±J spin glasses for T = 0. Journal of Statistical Physics, 62(1–2):435–442, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020879.

Bendisch:1992:GST

[Ben92a] J. Bendisch. On the ground-state threshold in random twodimensional Ising ±J models. Journal of Statistical Physics, 67 (5–6):1209–1217, June 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049016.

Bendler:1992:BR

[Ben92b] John Bendler. Book review. Journal of Statistical Physics, 69(1– 2):439–441, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053801.

Bendler:1993:BRF

[Ben93] John T. Bendler. Book review: Fractals and disordered systems. Journal of Statistical Physics, 71(5–6):1233–1235, June 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01049972.

Bendler:1995:FS

[Ben95a] John Bendler. Fractals in science. Journal of Statistical Physics, 81(3–4):857–860, November 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179263.

Benedetto:1995:RHF

[Ben95b] Dario Benedetto. A remark on the Hamiltonian formalism for incompressible flows. Journal of Statistical Physics, 79(3–4):743– 747, May 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02184880.

Brak:1998:NRD

[BEO98] R. Brak, J. W. Essam, and A. L. Owczarek. New results for directed vesicles and chains near an attractive wall. Journal of Statistical Physics, 93(1–2):155–192, October 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1023/B%3AJOSS.0000026731.35385.93.pdf.

Beran:1990:BRPa

[Ber90a] Mark J. Beran. Book review: Principles of statistical radiophysics, 3, elements of random fields. Journal of Statistical Physics, 59(1–2):533–534, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01015583.

Beran:1990:BRPb

[Ber90b] Mark J. Beran. Book review: Principles of statistical radiophysics, 4, elements of random fields. Journal of Statistical Physics, 61(3–4):957–958, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01027316.

Bernardin:1992:GIE

[Ber92] D. Bernardin. Global invariants and equilibrium states in lattice gases. Journal of Statistical Physics, 68(3–4):457–495, August 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01341758.

Berezhkovskii:1994:VDV

[Ber94a] A. M. Berezhkovskii. Volume of the domain visited by n spherical Brownian particles. Journal of Statistical Physics, 76(3–4): 1089–1097, August 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02188701.

Bershadskii:1994:TFP

[Ber94b] A. Bershadskii. Topological and fractal properties of turbulent passive scalar fluctuations at small scales. Journal of Statistical Physics, 77(3–4):909–914, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179468.

Berg:1996:MR

[Ber96] Bernd A. Berg. Multicanonical recursions. Journal of Statistical Physics, 82(1–2):323–342, January 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02189233.

Bernardes:1997:CMC

[Ber97a] Américo T. Bernardes. Can males contribute to the genetic improvement of a species? Journal of Statistical Physics, 86(1– 2):431–439, January 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02180214.

Bershadskii:1997:MPG

[Ber97b] A. Bershadskii. Multifractal percolation and growth in intermittent media. Journal of Statistical Physics, 87(3–4):607–611, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02181239.

Bershadskii:1997:MNI

[Ber97c] A. Bershadskii. Multifractality of nonlinear iterative processes. Journal of Statistical Physics, 89(3–4):869–875, November 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02765548.

Bertoin:1998:LDE

[Ber98] Jean Bertoin. Large-deviations estimates in Burgers turbulence with stable noise initial data. Journal of Statistical Physics, 91 (3–4):655–667, May 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023081728243.

Bershadskii:1999:MSM

[Ber99] A. Bershadskii. Multifractal statistics of mesoscopic systems. Journal of Statistical Physics, 94(5–6):725–737, March 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004549829897.

Baumgartner:1998:SMR

[BES98] A. Baumgärtner, U. Ebert, and L. Schäfer. Segment motion in the reptation model of polymer dynamics. II. Simulations. Journal of Statistical Physics, 90(5–6):1375–1400, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023291714290.

Benettin:1991:CFP

[BF91] Giancarlo Benettin and Francesco Fassò. Classical ‘freezing’ of plane rotations: A proof of the Boltzmann–Jeans conjecture. Journal of Statistical Physics, 63(3–4):737–760, May 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01029209.

Belitsky:1995:BAD

[BF95] Vladimir Belitsky and Pablo A. Ferrari. Ballistic annihilation and deterministic surface growth. Journal of Statistical Physics, 80(3–4):517–543, August 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02178546.

Barkema:1996:TDO

[BF96] G. T. Barkema and S. Flesia. Two-dimensional oriented selfavoiding walks with parallel contacts. Journal of Statistical Physics, 85(3–4):363–381, November 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02174210.

Baker:1997:FFF

[BF97] T. H. Baker and P. J. Forrester. Finite- N fluctuation formulas for random matrices. Journal of Statistical Physics, 88(5–6): 1371–1386, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02732439.

Bievre:1998:TPK

[BF98] S. De Bièvre and G. Forni. Transport properties of kicked and quasiperiodic Hamiltonians. Journal of Statistical Physics, 90(5– 6):1201–1223, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023227327494.

Barkai:1999:SOD

[BF99] E. Barkai and V. Fleurov. Stochastic one-dimensional Lorentz gas on a lattice. Journal of Statistical Physics, 96(1–2):325–359, July 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004532702233.

Braga:1994:UBC

[BFB94] Gastão A. Braga, S. J. Ferreira, and F. C. Sá Barreto. Upper bounds on the critical temperature for the two-dimensional Blume–Emery–Griffiths model. Journal of Statistical Physics, 76 (3–4):819–834, August 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02188687.

Bezandry:1993:FCL

[BFG93] Paul Hubert Bezandry, Xavier Fernique, and Gaston Giroux. A functional central limit theorem for a nonequilibrium model of interacting particles with unbounded intensity. Journal of Statistical Physics, 72(1–2):329–353, July 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048053.

Bal:1999:RTP

[BFPR99] Guillaume Bal, Albert Fannjiang, George Papanicolaou, and Leonid Ryzhik. Radiative transport in a periodic structure. Journal of Statistical Physics, 95(1–2):479–494, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004598015978.

Benassi:1991:AAP

[BFSV91] A. Benassi, J. P. Fouque, E. Saada, and M. E. Vares. Asymmetric attractive particle systems on z: Hydrodynamic limit for monotone initial profiles. Journal of Statistical Physics, 63 (3–4):719–735, May 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01029208.

Benfatto:1990:PTF

[BG90a] G. Benfatto and G. Gallavotti. Perturbation theory of the Fermi surface in a quantum liquid. A general quasiparticle formalism and one-dimensional systems. Journal of Statistical Physics, 59 (3–4):541–664, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01025844.

Bruno:1990:IBE

[BG90b] O. Bruno and K. Golden. Interchangeability and bounds on the effective conductivity of the square lattice. Journal of Statistical Physics, 61(1–2):365–386, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01013970.

Bovier:1992:RBS

[BG92] Anton Bovier and Véronique Gayrard. Rigorous bounds on the storage capacity of the dilute Hopfield model. Journal of Statistical Physics, 69(3–4):597–627, November 1992. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01050427.

Bambusi:1993:ESS

[BG93a] Dario Bambusi and Antonio Giorgilli. Exponential stability of states close to resonance in infinite-dimensional Hamiltonian systems. Journal of Statistical Physics, 71(3–4):569–606, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01058438.

Blanchard:1993:CAA

[BG93b] Ph. Blanchard and D. Gandolfo. Cellular automata approach to site percolation on Z^{2}. A numerical study. Journal of Statistical Physics, 73(1–2):399–408, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01052768.

Bovier:1993:RRT

[BG93c] Anton Bovier and Véronique Gayrard. Rigorous results on the thermodynamics of the dilute Hopfield model. Journal of Statistical Physics, 72(1–2):79–112, July 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048041.

Bovier:1993:TCW

[BG93d] Anton Bovier and Véronique Gayrard. The thermodynamics of the Curie–Weiss model with random couplings. Journal of Statistical Physics, 72(3–4):643–664, August 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048027.

Benettin:1994:HIN

[BG94] Giancarlo Benettin and Antonio Giorgilli. On the Hamiltonian interpolation of near-to-the identity symplectic mappings with application to symplectic integration algorithms. Journal of Statistical Physics, 74(5–6):1117–1143, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02188219.

Brandt:1996:OMA

[BG96] A. Brandt and M. Galun. Optimal multigrid algorithms for the massive Gaussian model and path integrals. Journal of Statistical Physics, 82(5–6):1503–1518, March 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183393.

Bastolla:1997:PTS

[BG97a] Ugo Bastolla and Peter Grassberger. Phase transitions of single semistiff polymer chains. Journal of Statistical Physics, 89(5–6): 1061–1078, December 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02764222.

Brandt:1997:OMA

[BG97b] A. Brandt and M. Galun. Optimal multigrid algorithms for variable-coupling isotropic Gaussian models. Journal of Statistical Physics, 88(3–4):637–664, August 1997. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000015166.92664.d8.

Bernard:1998:SMP

[BGK98] Denis Bernard, Krzysztof Gawedzki, and Antti Kupiainen. Slow modes in passive advection. Journal of Statistical Physics, 90(3– 4):519–569, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023212600779.

Bardos:1991:FDL

[BGL91] Claude Bardos, François Golse, and David Levermore. Fluid dynamic limits of kinetic equations. I. Formal derivations. Journal of Statistical Physics, 63(1–2):323–344, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01026608.

Bezuidenhout:1998:PMS

[BGL98] Carol Bezuidenhout, Geoffrey Grimmett, and Armin Löffler. Percolation and minimal spanning trees. Journal of Statistical Physics, 92(1–2):1–34, July 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023092317419.

Bauer:1999:SPE

[BGL99] M. Bauer, C. Godrèche, and J. M. Luck. Statistics of persistent events in the binomial random walk: Will the drunken sailor hit the Sober man? Journal of Statistical Physics, 96(5–6): 963–1019, September 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004636216365.

Benfatto:1997:ELI

[BGM97] G. Benfatto, G. Gentile, and V. Mastropietro. Electrons in a lattice with an incommensurate potential. Journal of Statistical Physics, 89(3–4):655–708, November 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02765540.

Benfatto:1998:PIH

[BGM98] G. Benfatto, G. Gentile, and V. Mastropietro. Peierls instability for the Holstein model with rational density. Journal of Statistical Physics, 92(5–6):1071–1113, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023052812507.

Baake:1995:PFZ

[BGP95a] Michael Baake, Uwe Grimm, and Carmelo Pisani. Partition function zeros for aperiodic systems. Journal of Statistical Physics, 78 (1–2):285–297, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183349.

Bovier:1995:GSH

[BGP95b] Anton Bovier, Véronique Gayrard, and Pierre Picco. Gibbs states of the Hopfield model with extensively many patterns. Journal of Statistical Physics, 79(1–2):395–414, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179395.

Brandt:1994:OMA

[BGR94] A. Brandt, M. Galun, and D. Ron. Optimal multigrid algorithms for calculating thermodynamic limits. Journal of Statistical Physics, 74(1–2):313–348, January 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02186816.

Banerjee:1998:TLD

[BGW98] S. Banerjee, R. B. Griffiths, and M. Widom. Thermodynamic limit for dipolar media. Journal of Statistical Physics, 93(1– 2):109–141, October 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1023/B%3AJOSS.0000026729.83187.79.
pdf.

Buffet:1991:DRW

[BH91] E. Buffet and P. Hannigan. Directed random walks in random environments. Journal of Statistical Physics, 65(3–4): 645–672, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053747.

Braun:1998:TJL

[BH98] Oleg Braun and Bambi Hu. Traffic jams in a lattice-gas model. Journal of Statistical Physics, 92(3–4):629–649, August 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023092705897.

Bhanot:1990:NMC

[Bha90] Gyan Bhanot. A numerical method to compute exactly the partition function with application to Z(n) theories in two dimensions. Journal of Statistical Physics, 60(1–2):55–75, July 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01013669.

Bowen:1992:MCS

[BHJ92] Chris Bowen, D. L. Hunter, and Naeem Jan. Monte Carlo simulation of the two-dimensional planar model. Journal of Statistical Physics, 69(5–6):1097–1113, December 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01058763.

Bandt:1998:PSA

[BHK98] C. Bandt, K. P. Hadeler, and F. Kriese. Particle systems acting on undirected graphs. Journal of Statistical Physics, 91 (3–4):571–586, May 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023073426426.

Broderix:1995:FLT

[BHKL95] Kurt Broderix, Dirk Hundertmark, Werner Kirsch, and Hajo Leschke. The fate of Lifshits tails in magnetic fields. Journal of Statistical Physics, 80(1–2):1–22, July 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02178350.

Bulsara:1993:P

[BHM^{+}93] Adi Bulsara, Peter Hänggi, Fabio Marchesoni, Frank Moss, and Michael Shlesinger. Preface. Journal of Statistical Physics, 70 (1–2):1–2, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053949.

Bocquet:1994:BMMb

[BHP94] Lydéric Bocquet, Jean-Pierre Hansen, and Jaroslaw Piasecki. On the Brownian motion of a massive sphere suspended in a hard-sphere fluid. II. Molecular dynamics estimates of the friction coefficient. Journal of Statistical Physics, 76(1–2):527–548, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02188674.

Blum:1996:SMB

[BHP96] L. Blum, M. F. Holovko, and I. A. Protsykevych. A solution of the multiple-binding mean spherical approximation for ionic mixtures. Journal of Statistical Physics, 84(1–2):191–204, July 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179582.

Bocquet:1997:FTP

[BHP97] Lydéric Bocquet, Jean-Pierre Hansen, and Jaroslaw Piasecki. Friction tensor for a pair of Brownian particles: Spurious finitesize effects and molecular dynamics estimates. Journal of Statistical Physics, 89(1–2):321–346, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02770768.

Benettin:1999:ELE

[BHS99] Giancarlo Benettin, Poul Hjorth, and Paolo Sempio. Exponentially long equilibrium times in a one-dimensional collisional model of classical gas. Journal of Statistical Physics, 94(5– 6):871–891, March 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004583016693.

Brightwell:1999:NBH

[BHW99] Graham R. Brightwell, Olle Häggström, and Peter Winkler. Nonmonotonic behavior in hard-core and Widom–Rowlinson models. Journal of Statistical Physics, 94(3–4):415–435, February 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004592103315.

Borgs:1992:CFS

[BI92] Christian Borgs and John Z. Imbrie. Crossover finite-size scaling at first-order transitions. Journal of Statistical Physics, 69(3–4): 487–537, November 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01050424.

Bobylev:1999:CIA

[BI99] Alexander V. Bobylev and Reinhard Illner. Collision integrals for attractive potentials. Journal of Statistical Physics, 95(3–4):633– 649, May 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004543325973.

Binder:1992:FPT

[Bin92] P.-M. Binder. First-passage times in a critical stochastic model. Journal of Statistical Physics, 67(3–4):827–832, May 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01049730.

Biskup:1998:RPR

[Bis98] Marek Biskup. Reflection positivity of the random-cluster measure invalidated for noninteger q. Journal of Statistical Physics, 92(3–4):369–375, August 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023076202262.

Baker:1990:BMO

[BJ90] George A. Baker, Jr. and J. D. Johnson. The border model in one dimension. Journal of Statistical Physics, 58(3–4):467–474, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01112756.

Barbaroux:1998:EVO

[BJ98] J. M. Barbaroux and A. Joye. Expectation values of observables in time-dependent quantum mechanics. Journal of Statistical Physics, 90(5–6):1225–1249, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023279311564.

Becker:1999:FTC

[BJ99] V. Becker and H. K. Janssen. Field theory of critical behavior in driven diffusive systems with quenched disorder. Journal of Statistical Physics, 96(3–4):817–859, August 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004610726108.

Bunimovich:1991:DEG

[BJL^{+}91] L. Bunimovich, H. R. Jauslin, J. L. Lebowitz, A. Pellegrinotti, and P. Nielaba. Diffusive energy growth in classical and quantum driven oscillators. Journal of Statistical Physics, 62(3–4): 793–817, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01017984.

Blekher:1992:FST

[BJL92] P. M. Blekher, H. R. Jauslin, and J. L. Lebowitz. Floquet spectrum for two-level systems in quasiperiodic time-dependent fields. Journal of Statistical Physics, 68(1–2):271–310, July 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01048846.

Bhagavatula:1997:NIE

[BJO97] Ravi Bhagavatula, David Jasnow, and T. Ohta. Nonequilibrium interface equations: An application to thermocapillary motion in binary systems. Journal of Statistical Physics, 88(5–6): 1013–1031, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02732424.

Bolle:1998:PDF

[BJS98] D. Bollé, G. Jongen, and G. M. Shim. Parallel dynamics of fully connected q-Ising neural networks. Journal of Statistical Physics, 91(1–2):125–153, April 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023088004195.

Bolle:1999:PDE

[BJS99] D. Bollé, G. Jongen, and G. M. Shim. Parallel dynamics of extremely diluted symmetric Q-Ising neural networks. Journal of Statistical Physics, 96(3–4):861–882, August 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004662710179.

Borgs:1990:RTF

[BK90] Christian Borgs and Roman Kotecký. A rigorous theory of finitesize scaling at first-order phase transitions. Journal of Statistical Physics, 61(1–2):79–119, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01013955.

Barszczak:1991:MFD

[BK91] T. Barszczak and R. Kutner. Modified Fermi–Dirac statistics of fermionic lattice gas by the back-jump correlations. Journal of Statistical Physics, 62(1–2):389–397, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020874.

Balents:1992:DPP

[BK92a] Leon Balents and Mehran Kardar. Directed paths on percolation clusters. Journal of Statistical Physics, 67(1–2):1–11, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01049024.

Block:1992:CTE

[BK92b] Louis Block and James Keesling. Computing the topological entropy of maps of the interval with three monotone pieces. Journal of Statistical Physics, 66(3–4):755–774, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01055699.

Bovier:1992:SHI

[BK92c] Anton Bovier and Christof Külske. Stability of hierarchical interfaces in a random field model. Journal of Statistical Physics, 69 (1–2):79–110, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053784.

Bovier:1993:HIR

[BK93] Anton Bovier and Christof Külske. Hierarchical interfaces in random media II: The Gibbs measures. Journal of Statistical Physics, 73(1–2):253–266, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01052760.

Bos:1994:MMR

[BK94a] S. Bös and R. Kühn. Multiplicity of metastable retrieval phases in networks of multistate neurons. Journal of Statistical Physics, 76(5–6):1495–1504, September 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02187073.

Brydges:1994:ADS

[BK94b] David C. Brydges and Georg Keller. Absence of Debye screening in the quantum Coulomb system. Journal of Statistical Physics, 76(1–2):285–297, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02188662.

Berkowitz:1995:CPC

[BK95a] Brian Berkowitz and Rosemary Knight. Continuum percolation conductivity exponents in restricted domains. Journal of Statistical Physics, 80(5–6):1415–1423, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179877.

Borgs:1995:SIF

[BK95b] C. Borgs and R. Kotecký. Surface-induced finite-size effects for first-order phase transitions. Journal of Statistical Physics, 79 (1–2):43–115, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179383.

Berglund:1996:IEC

[BK96a] N. Berglund and H. Kunz. Integrability and ergodicity of classical billiards in a magnetic field. Journal of Statistical Physics, 83 (1–2):81–126, April 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183641.

Bovier:1996:TNN

[BK96b] Anton Bovier and Christof Külske. There are no nice interfaces in (2+1)-dimensional SOS models in random media. Journal of Statistical Physics, 83(3–4):751–759, May 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183747.

Brezin:1993:VMT

[BKJZJ93] E. Brézin, E. Korutcheva, Th. Jolicoeur, and J. Zinn-Justin. O(N) vector model with twisted boundary conditions. Journal of Statistical Physics, 70(3–4):583–598, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01053585.

Baccelli:1992:MFL

[BKK^{+}92] F. Baccelli, F. I. Karpelevich, M. Ya. Kelbert, A. A. Puhalskii, A. N. Rybko, and Yu. M. Suhov. A mean-field limit for a class of queueing networks. Journal of Statistical Physics, 66(3–4): 803–825, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01055703.

Bene:1997:NAC

[BKL97a] Julius Bene, Zoltán Kaufmann, and Hans Lustfeld. New approach to the correlation spectrum near intermittency: A quantum mechanical analogy. Journal of Statistical Physics, 89(3–4): 605–632, November 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02765538.

Benois:1997:DBA

[BKL97b] O. Benois, A. Koukkous, and C. Landim. Diffusive behavior of asymmetric zero-range processes. Journal of Statistical Physics, 87(3–4):577–591, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181237.

Berezner:1993:ECT

[BKM93] S. A. Berezner, M. Krutina, and V. A. Malyshev. Exponential convergence of Toom’s probabilistic cellular automata. Journal of Statistical Physics, 73(5–6):927–944, December 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01052816.

Borgs:1991:FSS

[BKMS91] Christian Borgs, Roman Kotecký, and Salvador Miracle-Solé. Finite-size scaling for Potts models. Journal of Statistical Physics, 62(3–4):529–551, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01017971.

Bubanja:1993:AFS

[BKV93] Vladimir Bubanja, Milan Knezević, and Jean Vannimenus. Adsorption of a flexible self-avoiding polymer chain: Exact results on fractal lattices. Journal of Statistical Physics, 71(1–2):1–21, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048085.

Brosa:1990:FTP

[BKW90] U. Brosa, C. Kiittner, and U. Werner. Flow through a porous membrane simulated by cellular automata and by finite elements. Journal of Statistical Physics, 60(5–6):875–887, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01026001.

Buttiker:1990:RBL

[BL90] M. Büttiker and R. Landauer. Response to ‘The Búttiker– Landauer model generalized’. Journal of Statistical Physics, 58 (1–2):371–373, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020298. See [SH89].

Bramson:1991:ABD

[BL91a] Maury Bramson and Joel L. Lebowitz. Asymptotic behavior of densities for two-particle annihilating random walks. Journal of Statistical Physics, 62(1–2):297–372, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020872.

Bramson:1991:SSD

[BL91b] Maury Bramson and Joel L. Lebowitz. Spatial structure in diffusion-limited two-particle reactions. Journal of Statistical Physics, 65(5–6):941–951, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01049591.

Bhanot:1993:PFZ

[BL93] Gyan Bhanot and Jan Lacki. Partition function zeros and the three-dimensional Ising spin glass. Journal of Statistical Physics, 71(1–2):259–267, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048099.

Bleher:1994:ELS

[BL94] Pavel M. Bleher and Joel L. Lebowitz. Energy-level statistics of model quantum systems: Universality and scaling in a latticepoint problem. Journal of Statistical Physics, 74(1–2):167–217, January 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186812.

Blum:1997:P

[BL97] Lesser Blum and Joel L. Lebowitz. Preface. Journal of Statistical Physics, 89(1–2):1–3, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02770749.

Butta:1999:HLB

[BL99] Paolo Buttà and Joel L. Lebowitz. Hydrodynamic limit of Brownian particles interacting with short- and long-range forces. Journal of Statistical Physics, 94(3–4):653–694, February 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004512807858.

Bleher:1990:ELS

[Ble90] P. M. Bleher. The energy level spacing for two harmonic oscillators with golden mean ratio of frequencies. Journal of Statistical Physics, 61(3–4):869–876, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01027305.

Bleher:1991:ELS

[Ble91] P. M. Bleher. The energy level spacing for two harmonic oscillators with generic ratio of frequencies. Journal of Statistical Physics, 63(1–2):261–283, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01026603.

Bleher:1992:SPT

[Ble92] P. M. Bleher. Statistical properties of two-dimensional periodic Lorentz gas with infinite horizon. Journal of Statistical Physics, 66(1–2):315–373, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01060071.

Bunimovich:1990:ECS

[BLL90] L. A. Bunimovich, A. Lambert, and R. Lima. The emergence of coherent structures in coupled map lattices. Journal of Statistical Physics, 61(1–2):253–262, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01013964.

Benyoussef:1994:MRA

[BLL94] Abdelilah Benyoussef, Lahoussine Laanait, and Mohamed Loulidi. More results on the Ashkin–Teller model. Journal of Statistical Physics, 74(5–6):1185–1193, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02188222.

Bertini:1997:DCH

[BLO97] L. Bertini, C. Landim, and S. Olla. Derivation of Cahn–Hilliard equations from Ginzburg–Landau models. Journal of Statistical Physics, 88(1–2):365–381, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02508476.

Belinicher:1998:CSE

[BLPP98] Victor I. Belinicher, Victor S. L’vov, Anna Pomyalov, and Itamar Procaccia. Computing the scaling exponents in fluid turbulence from first principles: Demonstration of multiscaling. Journal of Statistical Physics, 93(3–4):797–832, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033164.35116.35.

Blumen:1991:FDK

[BLS91] A. Blumen, S. Luding, and I. M. Sokolov. Fluctuation-dominated kinetics in the a+b->0 reaction between immobile particles. Journal of Statistical Physics, 65(5–6):849–857, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01049585.

Bach:1994:GHF

[BLS94] Volker Bach, Elliott H. Lieb, and Jan Philip Solovej. Generalized Hartree–Fock theory and the Hubbard model. Journal of Statistical Physics, 76(1–2):3–89, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02188656.

Blum:1994:CTR

[Blu94] L. Blum. Contact theorems for rough interfaces. Journal of Statistical Physics, 75(5–6):971–980, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02186753.

Bagarello:1992:DMF

[BM92a] F. Bagarello and G. Morchio. Dynamics of mean-field spin models from basic results in abstract differential equations. Journal of Statistical Physics, 66(3–4):849–866, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01055705.

Blumel:1992:QC

[BM92b] R. Blümel and J. B. Mehl. Quantum chaos. Journal of Statistical Physics, 68(1–2):311–319, July 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01048847.

Baesens:1996:ETP

[BM96a] C. Baesens and R. S. MacKay. Effect of temperature on polaronic and bipolaronic structures of the adiabatic Holstein model. Journal of Statistical Physics, 85(3–4):471–488, November 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02174215.

Barkema:1996:MCS

[BM96b] G. T. Barkema and John McCabe. Monte Carlo simulations of conformal theory predictions for the three-state Potts model. Journal of Statistical Physics, 84(5–6):1067–1075, September 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02174128.

Barata:1997:GSD

[BM97] J. C. A. Barata and D. H. U. Marchetti. Griffiths’ singularities in diluted Ising models on the Cayley tree. Journal of Statistical Physics, 88(1–2):231–268, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02508471.

Brydges:1999:CSL

[BM99] David C. Brydges and Ph. A. Martin. Coulomb systems at low density: A review. Journal of Statistical Physics, 96(5–6): 1163–1330, September 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004600603161.

Bacry:1993:SSF

[BMA93] E. Bacry, J. F. Muzy, and A. Arnéodo. Singularity spectrum of fractal signals from wavelet analysis: Exact results. Journal of Statistical Physics, 70(3–4):635–674, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01053588.

Betbeder-Matibet:1994:CEF

[BMC94] Odile Betbeder-Matibet and Monique Combescot. On the Coulomb energy of a finite-temperature electron gas. Journal of Statistical Physics, 75(5–6):953–969, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02186752.

Bobylev:1997:TML

[BMHH97] A. V. Bobylev, Frank A. Maaø, Alex Hansen, and E. H. Hauge. There is more to be learned from the Lorentz model. Journal of Statistical Physics, 87(5–6):1205–1228, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02181280.

Benfatto:1995:SNR

[BMO95] G. Benfatto, E. Marinari, and E. Olivieri. Some numerical results on the block spin transformation for the 2D Ising model at the critical point. Journal of Statistical Physics, 78(3–4):731–757, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183686.

Berkovich:1996:PII

[BMO96] Alexander Berkovich, Barry M. McCoy, and William P. Orrick. Polynomial identities, indices, and duality for the N = 1 superconformal model SM(2, 4v). Journal of Statistical Physics, 83 (5–6):795–837, June 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179546.

Brower:1990:GIL

[BMOT90] Richard C. Brower, K. J. M. Moriarty, Peter Orland, and Pablo Tamayo. Gauge-invariant lattice gas for the microcanonical Ising model. Journal of Statistical Physics, 58(1–2):141–157, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020289.

Boffi:1990:SMD

[BMP90] V. C. Boffi, F. Malvagi, and G. C. Pomraning. Solution methods for discrete-state Markovian initial value problems. Journal of Statistical Physics, 60(3–4):445–472, August 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01314930.

Brown:1992:MAM

[BMP92] G. Brown, G. Michon, and J. Peyrière. On the multifractal analysis of measures. Journal of Statistical Physics, 66(3–4):775–790, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01055700.

Bonvin:1998:SMA

[BMPZ98] J. C. Bonvin, Ph. A. Martin, J. Piasecki, and X. Zotos. Statistics of mass aggregation in a self-gravitating one-dimensional gas. Journal of Statistical Physics, 91(1–2):177–197, April 1998.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023092105104.

Boukraa:1995:DSG

[BMR95] S. Boukraa, J.-M. Maillard, and G. Rollet. Discrete symmetry groups of vertex models in statistical mechanics. Journal of Statistical Physics, 78(5–6):1195–1251, March 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02180130.

Berezhkovskii:1991:MBE

[BMS91] A. M. Berezhkovskii, Yu. A. Makhnovskii, and R. A. Suris. Many-body effects in diffusion-limited kinetics. Journal of Statistical Physics, 65(5–6):1025–1041, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01049596.

Bronski:1997:STH

[BMS97] Jared C. Bronski, David W. McLaughlin, and Michael J. Shelley. On the stability of time-harmonic localized states in a disordered nonlinear medium. Journal of Statistical Physics, 88(5–6): 1077–1115, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02732426.

Binder:1999:IPB

[BMSW99] K. Binder, M. Müller, F. Schmid, and A. Werner. Interfacial profiles between coexisting phases in thin films: Cahn–Hilliard treatment versus capillary waves. Journal of Statistical Physics, 95 (5–6):1045–1068, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004510702716.

Beysens:1999:WIA

[BN99] D. Beysens and T. Narayanan. Wetting-induced aggregation of colloids. Journal of Statistical Physics, 95(5–6):997–1008, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004506601807.

Ben-Naim:1998:DND

[BNK98] E. Ben-Naim and P. L. Krapivsky. Domain number distribution in the nonequilibrium Ising model. Journal of Statistical Physics, 93(3–4):583–601, November 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033243.27556.99.

Balakrishnan:1995:EVD

[BNN95] V. Balakrishnan, C. Nicolis, and G. Nicolis. Extreme value distributions in chaotic dynamics. Journal of Statistical Physics, 80 (1–2):307–336, July 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02178361.

Balakrishnan:1997:RTS

[BNN97] V. Balakrishnan, G. Nicolis, and C. Nicolis. Recurrence time statistics in chaotic dynamics. I. Discrete time maps. Journal of Statistical Physics, 86(1–2):191–212, January 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02180204.

Brak:1998:CSC

[BNO98] R. Brak, P. P. Nidras, and A. L. Owczarek. Cluster structure of collapsing polymers. Journal of Statistical Physics, 91(1–2):75– 93, April 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023083903287.

Ben-Naim:1993:PAV

[BNRW93] E. Ben-Naim, S. Redner, and G. H. Weiss. Partial absorption and ‘virtual’ traps. Journal of Statistical Physics, 71(1–2):75–88, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048089.

Bonilla:1992:NSI

[BNS92] Luis L. Bonilla, John C. Neu, and Renato Spigler. Nonlinear stability of incoherence and collective synchronization in a population of coupled oscillators. Journal of Statistical Physics, 67 (1–2):313–330, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049037.

Bouchaud:1990:BAR

[BO90] J. P. Bouchaud and H. Orland. On the Bethe ansatz for random directed polymers. Journal of Statistical Physics, 61(3–4):

877–884, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01027306.

Bashkirov:1991:TPP

[BO91a] A. G. Bashkirov and A. V. Orlov. Transport phenomena in a plane shock wave. Journal of Statistical Physics, 64(1–2):429– 436, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01057885.

Burlatsky:1991:FKD

[BO91b] S. F. Burlatsky and G. S. Oshanin. Fluctuation kinetics of diffusion-controlled processes: Strong effects due to correlations and fluctuations. Journal of Statistical Physics, 65(5–6):1095– 1107, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049600.

Blanchard:1999:IQCa

[BO99a] Ph. Blanchard and R. Olkiewicz. Interacting quantum and classical continuous systems I. The piecewise deterministic dynamics. Journal of Statistical Physics, 94(5–6):913–931, March 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004587117602.

Blanchard:1999:IQCb

[BO99b] Ph. Blanchard and R. Olkiewicz. Interacting quantum and classical continuous systems II. Asymptotic behavior of the quantum subsystem. Journal of Statistical Physics, 94(5–6):933–953, March 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004539201672.

Bobryk:1993:SEL

[Bob93] Roman V. Bobryk. Stochastic equations of the Langevin type under a weakly dependent perturbation. Journal of Statistical Physics, 70(3–4):1045–1056, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053607.

Bobylev:1995:QHB

[Bob95] A. V. Bobylev. Quasistationary hydrodynamics for the Boltzmann equation. Journal of Statistical Physics, 80(5–6):1063– 1083, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179864.

Bobylev:1997:MIB

[Bob97] A. V. Bobylev. Moment inequalities for the Boltzmann equation and applications to spatially homogeneous problems. Journal of Statistical Physics, 88(5–6):1183–1214, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02732431.

Boon:1992:F

[Boo92] Jean Pierre Boon. Foreword. Journal of Statistical Physics, 68 (3–4):347–349, August 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01341751.

Brak:1994:ESB

[BOP94] R. Brak, A. L. Owczarek, and T. Prellberg. Exact scaling behavior of partially convex vesicles. Journal of Statistical Physics, 76(5–6):1101–1128, September 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02187057.

Bottcher:1995:OFF

[Böt95] Albrecht Böttcher. The Onsager formula, the Fisher–Hartwig conjecture, and their influence on research into Toeplitz operators. Journal of Statistical Physics, 78(1–2):575–584, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183366.

Bouchut:1999:CBM

[Bou99] F. Bouchut. Construction of BGK models with a family of kinetic entropies for a given system of conservation laws. Journal of Statistical Physics, 95(1–2):113–170, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004525427365.

Bovier:1990:DSA

[Bov90] Anton Bovier. The density of states in the Anderson model at weak disorder: A renormalization group analysis of the hierarchical model. Journal of Statistical Physics, 59(3–4):745–779, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01025849.

Bovier:1998:KVS

[Bov98a] Anton Bovier. The Kac version of the Sherrington–Kirkpatrick model at high temperatures. Journal of Statistical Physics, 91 (1–2):459–474, April 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023064826485.

Brassesco:1998:CAE

[BOV98b] S. Brassesco, E. Olivieri, and M. E. Vares. Couplings and asymptotic exponentiality of exit times. Journal of Statistical Physics, 93(1–2):393–404, October 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1023/B%3AJOSS.0000026739.
46334.05.pdf.

Buffet:1990:LMG

[BP90] E. Buffet and J. V. Pulé. On Lushnikov’s model of gelation. Journal of Statistical Physics, 58(5–6):1041–1058, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01026562.

Bovier:1991:SIR

[BP91a] Anton Bovier and Pierre Picco. Stability of interfaces in a random environment. A rigorous renormalization group analysis of a hierarchical model. Journal of Statistical Physics, 62(1–2): 177–199, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020865.

Buffet:1991:PRG

[BP91b] E. Buffet and J. V. Pulé. Polymers and random graphs. Journal of Statistical Physics, 64(1–2):87–110, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01057869.

Bandt:1993:EPF

[BP93a] Christoph Bandt and Bernd Pompe. The entropy profile — a function describing statistical dependences. Journal of Statistical Physics, 70(3–4):967–983, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053603.

Bouchut:1993:BMS

[BP93b] François Bouchut and Benoít Perthame. A BGK model for small Prandtl number in the Navier–Stokes approximation. Journal of Statistical Physics, 71(1–2):191–207, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01048094.

Banach:1994:TLP

[BP94a] Zbigniew Banach and Slawomir Piekarski. Two linearization procedures for the Boltzmann equation in a k = 0 Robertson– Walker space–time. Journal of Statistical Physics, 76(5–6):1415– 1437, September 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02187069.

Becklehimer:1994:PCJ

[BP94b] Jeffrey L. Becklehimer and Ras B. Pandey. Percolation of chains and jamming coverage in two dimensions by computer simulation. Journal of Statistical Physics, 75(3–4):765–771, May 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02186881.

Buchholtz:1996:ASS

[BP96] Volkhard Buchholtz and Thorsten Pöschel. Avalanche statistics of sand heaps. Journal of Statistical Physics, 84(5–6):1373– 1378, September 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02174136.

Bocquet:1997:MDN

[BP97] Lydéric Bocquet and Jaroslaw Piasecki. Microscopic derivation of non-Markovian thermalization of a Brownian particle. Journal of Statistical Physics, 87(5–6):1005–1035, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02181268.

Butta:1998:LDP

[BP98] Paolo Buttà and Pierre Picco. Large-deviation principle for one-dimensional vector spin models with Kac potentials. Journal of Statistical Physics, 92(1–2):101–150, July 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023095619236.

Barnes:1994:AGL

[BPH^{+}94a] S. E. Barnes, M. Peter, L. Hoffmann, A. A. Manuel, and A. Shukla. Application of generalized linear filters in data analysis. Journal of Statistical Physics, 76(1–2):679–701, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02188681.

Bocquet:1994:BMMa

[BPH94b] Lydéric Bocquet, Jaroslaw Piasecki, and Jean-Pierre Hansen. On the Brownian motion of a massive sphere suspended in a hardsphere fluid. I. Multiple-time-scale analysis and microscopic expression for the friction coefficient. Journal of Statistical Physics, 76(1–2):505–526, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02188673.

Behrend:1996:IRF

[BPO96] Roger E. Behrend, Paul A. Pearce, and David L. O’Brien. Interaction-round-a-face models with fixed boundary conditions: The ABF fusion hierarchy. Journal of Statistical Physics, 84 (1–2):1–48, July 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179576.

Baxter:1990:HLZ

[BQ90a] R. J. Baxter and G. R. W. Quispel. Hamiltonian limit of the 3D Zamolodchikov model. Journal of Statistical Physics, 58(3– 4):411–430, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01112753.

Busch:1990:RPD

[BQ90b] Paul Busch and Ralf Quadt. On Ruch’s principle of decreasing mixing distance in classical statistical physics. Journal of Statistical Physics, 61(1–2):311–328, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01013967.

Blum:1991:RBF

[BR91] L. Blum and Yaakov Rosenfeld. Relation between the free energy and the direct correlation function in the mean spherical approximation. Journal of Statistical Physics, 63(5–6):1177–1190, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01030005.

Benfatto:1992:NFP

[BR92] G. Benfatto and J. Renn. Nontrivial fixed points and screening in the hierarchical two-dimensional Coulomb gas. Journal of Statistical Physics, 67(5–6):957–980, June 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01049006.

Bag:1999:STD

[BR99] Bidhan Chandra Bag and Deb Shankar Ray. A semiclassical theory of a dissipative Henon–Heiles system. Journal of Statistical Physics, 96(1–2):271–302, July 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004528601324.

Bramson:1991:RWR

[Bra91] Maury Bramson. Random walk in random environment: A counterexample without potential. Journal of Statistical Physics, 62(3–4):863–875, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01017987.

Branco:1993:PBP

[Bra93] N. S. Branco. Probabilistic bootstrap percolation. Journal of Statistical Physics, 70(3–4):1035–1044, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01053606.

Brenner:1991:MPB

[Bre91] Howard Brenner. Macrotransport processes: Brownian tracers as stochastic averagers in effective-medium theories of heterogeneous media. Journal of Statistical Physics, 62(5–6):1095–1119, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01128179.

Bricmont:1996:QNL

[Bri96a] J. Bricmont. Quantum non-locality and relativity. Journal of Statistical Physics, 82(3–4):1213–1216, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179810.

Bricmont:1996:SML

[Bri96b] J. Bricmont. The statistical mechanics of lattice gases. Journal of Statistical Physics, 82(1–2):453–454, January 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02189241.

Bronski:1998:NWP

[Bro98] Jared C. Bronski. Nonlinear wave propagation in a disordered medium. Journal of Statistical Physics, 92(5–6):995– 1015, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023096627528.

Bazhenov:1996:TPS

[BRR96] M. Bazhenov, M. Rabinovich, and L. Rubchinsky. Time-periodic spatial chaos in the complex Ginzburg–Landau equation. Journal of Statistical Physics, 83(5–6):1165–1181, June 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179556.

Berlyand:1998:FPP

[BRT98] L. Berlyand, M. D. Rintoul, and S. Torquato. First-passage percolation, semi-directed Bernoulli percolation, and failure in brittle materials. Journal of Statistical Physics, 91(3–4):603–623, May 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023077627335.

Brush:1994:CMP

[Bru94] Stephen G. Brush. Creating modern probability. Its mathematics, physics and philosophy in historical perspective. Journal of Statistical Physics, 77(5–6):1105–1107, December 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183156.

Bryngelson:1994:TCS

[Bry94] Joseph Bryngelson. Thermodynamics of chaotic systems: An introduction. Journal of Statistical Physics, 75(3–4):775–777, May 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186883.

Bryngelson:1996:ITS

[Bry96] Joseph D. Bryngelson. An introduction to the theory of spin glasses and neural networks. Journal of Statistical Physics, 85(3– 4):519–520, November 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02174219.

Bryngelson:1997:FON

[Bry97] Joseph D. Bryngelson. Fluctuations and order: The new synthesis. Journal of Statistical Physics, 87(3–4):963–964, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02181261.

Bryngelson:1998:BRB

[Bry98] Joseph D. Bryngelson. Book review: Scaling And Renormalization In Statistical Physics. John Cardy, Cambridge University Press, Cambridge, 1996. Journal of Statistical Physics, 90(3–4): 1071–1072, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023261911683.

Bleher:1995:PLG

[BRZ95] P. M. Bleher, J. Ruiz, and V. A. Zagrebnov. On the purity of the limiting Gibbs state for the Ising model on the Bethe lattice. Journal of Statistical Physics, 79(1–2):473–482, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179399.

Bleher:1996:ODR

[BRZ96] P. M. Bleher, J. Ruiz, and V. A. Zagrebnov. One-dimensional random-field Ising model: Gibbs states and structure of ground states. Journal of Statistical Physics, 84(5–6):1077– 1093, September 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02174129.

Bleher:1998:PDR

[BRZ98] P. M. Bleher, J. Ruiz, and V. A. Zagrebnov. On the phase diagram of the random field Ising model on the Bethe lattice. Journal of Statistical Physics, 93(1–2):33–78, October 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1023/B%3AJOSS.0000026727.43077.49.pdf.

Bazhanov:1990:CPM

[BS90a] V. V. Bazhanov and Yu. G. Stroganov. Chiral Potts model as a descendant of the six-vertex model. Journal of Statistical Physics, 59(3–4):799–817, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01025851.

Bhanot:1990:SIM

[BS90b] Gyan Bhanot and Srikanth Sastry. Solving the Ising model exactly on a 5×5×4 lattice using the Connection Machine. Journal of Statistical Physics, 60(3–4):333–346, August 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01314924.

Blum:1990:NNR

[BS90c] Thomas Blum and Yonathan Shapir. The nearest-neighbor resonating-valence bond state in a Grassmannian form. Journal of Statistical Physics, 59(1–2):333–355, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01015573.

Berry:1991:RNS

[BS91a] G. C. Berry and Mohan Srinivasarao. Rheology of nematic solutions of rodlike chains: Comparison of theory and experiment. Journal of Statistical Physics, 62(5–6):1041–1058, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01128176.

Brosa:1991:SFT

[BS91b] U. Brosa and D. Stauffer. Simulation of flow through a twodimensional random porous medium. Journal of Statistical Physics, 63(1–2):405–409, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01026613.

Bukman:1995:CPF

[BS95] Dirk Jan Bukman and Joel D. Shore. The conical point in the ferroelectric six-vertex model. Journal of Statistical Physics, 78(5– 6):1277–1309, March 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02180132.

Baldwin:1997:CTE

[BS97] Stewart L. Baldwin and Edward E. Slaminka. Calculating topological entropy. Journal of Statistical Physics, 89(5–6):1017– 1033, December 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02764219.

Bicout:1998:WUP

[BS98] D. J. Bicout and Attila Szabo. On the Wang–Uhlenbeck problem in discrete velocity space. Journal of Statistical Physics, 91(5– 6):1047–1054, June 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023088118307.

Beenakker:1997:BR

[BSB97] C. W. J. Beenakker, J. M. H. Levelt Sengers, and S. G. Brush. Book review. Journal of Statistical Physics, 89(5–6): 1097–1103, December 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02764225.

Bizon:1998:CDP

[BSdB^{+}98] C. Bizon, M. D. Shattuck, John R. de Bruyn, J. B. Swift, W. D. McCormick, and Harry L. Swinney. Convection and diffusion in patterns in oscillated granular media. Journal of Statistical Physics, 93(3–4):449–465, November 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033158.05927.3f.

Bassler:1991:IIM

[BSG91] Kevin E. Bassler, Kazuo Sasaki, and Robert B. Griffiths. Interface interactions in modulated phases, and upsilon points. Journal of Statistical Physics, 62(1–2):45–88, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020859.

Bernardin:1995:ESR

[BSG95] D. Bernardin and O. E. Sero-Guillaume. Exact stability results in stochastic lattice gas cellular automata. Journal of Statistical Physics, 81(1–2):409–443, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179987.

Bazzani:1994:MMDa

[BSTV94a] A. Bazzani, S. Siboni, G. Turchetti, and S. Vaienti. A model of modulated diffusion. I. Analytical results. Journal of Statistical Physics, 76(3–4):929–967, August 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02188693.

Bazzani:1994:MMDb

[BSTV94b] A. Bazzani, S. Siboni, G. Turchetti, and S. Vaienti. A model of modulated diffusion. II. Numerical results on statistical properties. Journal of Statistical Physics, 76(3–4):969–984, August 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02188694.

Bolle:1994:RCL

[BSV94] D. Bollé, G. M. Shim, and B. Vinck. Retrieval and chaos in layered q-Ising neural networks. Journal of Statistical Physics, 74(3– 4):583–606, February 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02188572.

Bolle:1994:RCE

[BSVZ94] D. Bollé, G. M. Shim, B. Vinck, and V. A. Zagrebnov. Retrieval and chaos in extremely diluted q-Ising neural networks. Journal of Statistical Physics, 74(3–4):565–582, February 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02188571.

Berche:1990:CIS

[BT90a] Bertrand Berche and Loïc Turban. Conformal invariance and surface defects in the two-dimensional Ising model. Exact results. Journal of Statistical Physics, 60(1–2):167–180, July 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013672.

Berger:1990:MPY

[BT90b] N. E. Berger and V. Twersky. Moments of the Percus–Yevick hard-sphere correlation function. Journal of Statistical Physics, 61(5–6):1187–1201, December 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014371.

Brankov:1990:IFS

[BT90c] Jordan G. Brankov and Nicholai S. Tonchev. An investigation of finite-size scaling for systems with long-range interaction: The spherical model. Journal of Statistical Physics, 59 (5–6):1431–1450, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01334758.

Brankov:1990:FSS

[BT90d] Jordan G. Brankov and Nicholai S. Tonchev. On finite-size scaling in the presence of dangerous irrelevant variables. Journal of Statistical Physics, 60(3–4):519–526, August 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01314934.

Bagarello:1991:AMF

[BT91] F. Bagarello and C. Trapani. ‘almost’ mean-field Ising model: An algebraic approach. Journal of Statistical Physics, 65(3–4): 469–482, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053740.

Bunimovich:1992:RPL

[BT92] L. A. Bunimovich and S. E. Troubetzkoy. Recurrence properties of Lorentz lattice gas cellular automata. Journal of Statistical Physics, 67(1–2):289–302, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01049035.

Basor:1993:VCB

[BT93a] Estelle L. Basor and Craig A. Tracy. Variance calculations and the Bessel kernel. Journal of Statistical Physics, 73(1–2):415–421, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01052770.

Bunimovich:1993:TDF

[BT93b] L. A. Bunimovich and S. E. Troubetzkoy. Topological dynamics of flipping Lorentz lattice gas models. Journal of Statistical Physics, 72(1–2):297–307, July 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01048051.

Bunimovich:1994:RPA

[BT94] L. A. Bunimovich and S. E. Troubetzkoy. Rotators, periodicity, and absence of diffusion in cyclic cellular automata. Journal of Statistical Physics, 74(1–2):1–10, January 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02186804.

Boghosian:1995:RES

[BT95] Bruce M. Boghosian and Washington Taylor. Renormalized equilibria of a Schlögl model lattice gas. Journal of Statistical Physics, 81(1–2):295–317, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179980.

Benzi:1998:HTR

[BTT98] R. Benzi, F. Toschi, and R. Tripiccione. On the heat transfer in Rayleigh–Bénard systems. Journal of Statistical Physics, 93(3– 4):901–918, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000033168.36971.59.

Brower:1991:PMA

[BTY91] R. C. Brower, Pablo Tamayo, and Bryant York. A parallel multigrid algorithm for percolation clusters. Journal of Statistical Physics, 63(1–2):73–88, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01026593.

Bunimovich:1997:LVL

[Bun97] L. A. Bunimovich. On localization of vorticity in Lorentz lattice gases. Journal of Statistical Physics, 87(1–2):449–457, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02181496.

Buot:1990:EIO

[Buo90] F. A. Buot. Exact integral operator form of the Wigner distribution-function equation in many-body quantum transport theory. Journal of Statistical Physics, 61(5–6):1223–1256, December 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014373.

Burkov:1991:LTD

[Bur91] S. E. Burkov. Are layered two-dimensional quasicrystals periodic in the third direction? Journal of Statistical Physics, 65(1– 2):395–401, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01329868.

Butta:1993:VER

[But93] Paolo Buttà. On the validity of an Einstein relation in models of interface dynamics. Journal of Statistical Physics, 72(5–6): 1401–1406, September 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048193.

Breymann:1994:PIP

[BV94] Wolfgang Breymann and Jürgen Vollmer. Pruning-induced phase transition observed by a scattering method. Journal of Statistical Physics, 76(5–6):1439–1465, September 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02187070.

Broidioi:1996:SFC

[BV96] M. Broidioi and M. Van Canneyt. Scaling of fluctuations and critical exponents. Journal of Statistical Physics, 82(1–2):115–129, January 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02189227.

Bobylev:1998:AIF

[BV98] A. V. Bobylev and H. D. Victory. Additive invariant functionals for dynamical systems. Journal of Statistical Physics, 92 (1–2):269–299, July 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023003921053.

Bovier:1999:SSB

[BvEN99] Anton Bovier, Aernout C. D. van Enter, and Beat Niederhauser. Stochastic symmetry-breaking in a Gaussian Hopfield model. Journal of Statistical Physics, 95(1–2):181–213, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1004529528273.

Blum:1992:ASO

[BVHP92] L. Blum, F. Vericat, and J. N. Herrera-Pacheco. On the analytical solution of the Ornstein–Zernike equation with Yukawa closure. Journal of Statistical Physics, 66(1–2):249–262, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01060067.

Bonilla:1993:GSP

[BVR93] L. L. Bonilla, C. J. Pérez Vicente, and J. M. Rubí. Glassy synchronization in a population of coupled oscillators. Journal of Statistical Physics, 70(3–4):921–937, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01053600.

Brito:1995:VAF

[BvV95] R. Brito and G. A. van Velzen. Velocity autocorrelation function in lattice gases from the ring kinetic theory. Comparison with numerical simulations. Journal of Statistical Physics, 80 (3–4):565–578, August 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02178548.

Bolle:1993:PDS

[BVZ93] D. Bollé, B. Vinck, and V. A. Zagrebnov. On the parallel dynamics of the Q-state Potts and Q-Ising neural networks. Journal of Statistical Physics, 70(5–6):1099–1119, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01049424.

Brydges:1988:MEH

[BW88] D. C. Brydges and J. D. Wright. Mayer expansions and the Hamilton–Jacobi equation. II. Fermions, dimensional reduction formulas. Journal of Statistical Physics, 51(3–4):435–456, May 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.

1007/BF01028465. See erratum [BW88].

Barsky:1998:CEC

[BW98a] David J. Barsky and C. Chris Wu. Critical exponents for the contact process under the triangle condition. Journal of Statistical Physics, 91(1–2):95–124, April 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023036020125.

Brenner:1998:SSG

[BW98b] Michael P. Brenner and Thomas P. Witelski. On spherically symmetric gravitational collapse. Journal of Statistical Physics, 93(3–4):863–899, November 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033167.19114.b8.

Basor:1999:DAO

[BW99] Estelle L. Basor and Harold Widom. Determinants of Airy operators and applications to random matrices. Journal of Statistical Physics, 96(1–2):1–20, July 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004539513619.

Bodenschatz:1991:IDD

[BWK91] Eberhard Bodenschatz, Andreas Weber, and Lorenz Kramer. Interaction and dynamics of defects in convective roll patterns of anisotropic fluids. Journal of Statistical Physics, 64(5–6): 1007–1015, September 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048810.

Bialek:1990:CCN

[BZ90] William Bialek and A. Zee. Coding and computation with neural spike trains. Journal of Statistical Physics, 59(1–2):103–115, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01015565.

Bassler:1995:PTD

[BZ95] Kevin E. Bassler and R. K. P. Zia. Phase transitions in a driven lattice gas at two temperatures. Journal of Statistical Physics, 80 (3–4):499–515, August 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02178545.

Bovier:1997:LTP

[BZ97] Anton Bovier and Milos Zahradník. The low-temperature phase of Kac–Ising models. Journal of Statistical Physics, 87(1– 2):311–332, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181490.

Bertacchi:1999:ERW

[BZ99] D. Bertacchi and F. Zucca. Equidistribution of random walks on spheres. Journal of Statistical Physics, 94(1–2):91–111, January 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004559228814.

Bao:1992:AMC

[BZW92] Jingdong Bao, Yizhong Zhuo, and Xizhen Wu. Accurate Monte Carlo tests of the stochastic Ginzburg–Landau model with multiplicative colored noise. Journal of Statistical Physics, 66(5–6): 1653–1658, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01054441.

Chialvo:1993:MNB

[CA93] Dante R. Chialvo and A. Vania Apkarian. Modulated noisy biological dynamics: Three examples. Journal of Statistical Physics, 70(1–2):375–391, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053974.

Caginalp:1990:MDM

[Cag90] G. Caginalp. A microscopic derivation of macroscopic sharp interface problems involving phase transitions. Journal of Statistical Physics, 59(3–4):869–884, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01025855.

Cammarota:1991:TDM

[Cam91] C. Cammarota. On the temperature dependence of the mean number of clusters. Journal of Statistical Physics, 63(3–4):783– 790, May 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01029211.

Cao:1993:PTD

[Cao93] M. S. Cao. The phase transition of the directed polymer on disordered hierarchical lattices. Journal of Statistical Physics, 71(1–2):51–59, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048087.

Chakrabarti:1990:TPE

[CB90] B. K. Chakrabarti and Somendra M. Bhattacharjee. Theta-point exponent for polymer chain in random media. Journal of Statistical Physics, 58(1–2):383–388, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020300.

Chiolero:1994:SRP

[CB94] A. Chiolero and D. Baeriswyl. Structural relaxations, phonons, and Ising models. Journal of Statistical Physics, 76(1–2):347– 360, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02188666.

Coppersmith:1999:APG

[CBK99] S. N. Coppersmith, Robert D. Blank, and Leo P. Kadanoff. Analysis of a population genetics model with mutation, selection, and pleiotropy. Journal of Statistical Physics, 97(3–4): 429–457, November 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004678222262.

Caprio:1995:PBC

[CBR95] D. Di Caprio, J. P. Badiali, and V. Russier. Properties of the n-body correlation functions near the liquid-gas critical point. Correlation inequalities. Journal of Statistical Physics, 80(5–6): 1241–1278, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179870.

Cannone:1991:ICR

[CC91a] Marco Cannone and Carlo Cercignani. The inverse conjecture for the revised Enskog equation. Journal of Statistical Physics, 63(1–2):363–387, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01026610.

Celletti:1991:ICA

[CC91b] Alessandra Celletti and Luigi Chierchia. Invariant curves for area-preserving twist maps far from integrable. Journal of Statistical Physics, 65(3–4):617–643, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01053746.

Costin:1991:LPD

[CC91c] O. Costin and R. D. Costin. Limit probability distributions for an infinite-order phase transition model. Journal of Statistical Physics, 64(1–2):193–205, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01057873.

Carlen:1992:SEP

[CC92] E. A. Carlen and M. C. Carvalho. Strict entropy production bounds and stability of the rate of convergence to equilibrium for the Boltzmann equation. Journal of Statistical Physics, 67 (3–4):575–608, May 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049721.

Carlen:1994:EPE

[CC94a] E. A. Carlen and M. C. Carvalho. Entropy production estimates for Boltzmann equations with physically realistic collision kernels. Journal of Statistical Physics, 74(3–4):743–782, February 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02188578.

Cercignani:1994:VMC

[CC94b] Carlo Cercignani and Stefano Cortese. Validation of a Monte Carlo simulation of the plane Couette flow of a rarefied gas. Journal of Statistical Physics, 75(5–6):817–838, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02186745.

Cao:1997:SPT

[CC97] Meng-She Cao and E. G. D. Cohen. Scaling of particle trajectories on a lattice. Journal of Statistical Physics, 87(1–2):147–178, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02181484.

Carlson:1990:BLSa

[CCC^{+}90] J. M. Carlson, J. T. Chayes, L. Chayes, J. P. Sethna, and D. J. Thouless. Bethe lattice spin glass: The effects of a ferromagnetic bias and external fields. I. Bifurcation analysis. Journal of Statistical Physics, 61(5–6):987–1067, December 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01014364.

Chen:1991:LGM

[CCD^{+}91] Shiyi Chen, Hudong Chen, Gary D. Doolen, Semion Gutman, and Minxu Lee. A lattice gas model for thermohydrodynamics. Journal of Statistical Physics, 62(5–6):1121–1151, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01128180.

Charach:1999:DPF

[CCF99] Ch. Charach, C. K. Chen, and P. C. Fife. Developments in phasefield modeling of thermoelastic and two-component materials. Journal of Statistical Physics, 95(5–6):1141–1164, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004566820463.

Costin:1990:IOP

[CCG90] O. Costin, R. D. Costin, and C. P. Grünfeld. Infinite-order phase transition in a classical spin system. Journal of Statistical Physics, 59(5–6):1531–1546, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01334762.

Cancrini:1999:SGK

[CCM99] N. Cancrini, F. Cesi, and F. Martinelli. The spectral gap for the Kawasaki dynamics at low temperature. Journal of Statistical Physics, 95(1–2):215–271, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004581512343.

Chayes:1999:MFT

[CCMS99] L. Chayes, A. Coniglio, J. Machta, and K. Shtengel. Mean-field theory for percolation models of the Ising type. Journal of Statistical Physics, 94(1–2):53–66, January 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004555127906.

Carlen:1999:ARD

[CCO99] E. A. Carlen, M. C. Carvalho, and E. Orlandi. Algebraic rate of decay for the excess free energy and stability of fronts for a nonlocal phase kinetics equation with a conservation law. I. Journal of Statistical Physics, 95(5–6):1069–1117, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004562719554.

Cerchiai:1995:KIA

[CCRM95] B. L. Cerchiai, P. Cotta-Ramusino, and M. Martellini. Knot invariants associated with a particular N → ∞ continuous limit of the Baxter–Bazhanov model. Journal of Statistical Physics, 81(3–4):629–645, November 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179250.

Carlson:1990:BLSb

[CCST90] J. M. Carlson, J. T. Chayes, J. P. Sethna, and D. J. Thouless. Bethe lattice spin glass: The effects of a ferromagnetic bias and external fields. II. Magnetized spin-glass phase and the de Almeida–Thouless line. Journal of Statistical Physics, 61(5–6): 1069–1084, December 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014365.

Calisto:1992:CNB

[CCT92] H. Calisto, E. Cerda, and E. Tirapegui. Comment on noise and bifurcations. Journal of Statistical Physics, 69(5–6):1115– 1122, December 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01058764.

Calisto:1993:WNE

[CCT93] H. Calisto, E. Cerda, and E. Tirapegui. Weak noise expansions through functional integrals for colored noise. Journal of Statistical Physics, 71(3–4):513–528, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01058435.

Calisto:1996:EPW

[CCT96] H. Calisto, E. Cerda, and E. Tirapegui. Effective potential and weak noise transitions. Journal of Statistical Physics, 82(3–4): 1015–1045, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179800.

Cook:1990:LEL

[CD90] J. Cook and B. Derrida. Lyapunov exponents of large, sparse random matrices and the problem of directed polymers with complex random weights. Journal of Statistical Physics, 61(5–6): 961–986, December 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014363.

Chopard:1991:CAM

[CD91a] B. Chopard and M. Droz. Cellular automata model for the diffusion equation. Journal of Statistical Physics, 64(3–4):859–892, August 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048321.

Cook:1991:FSE

[CD91b] J. Cook and B. Derrida. Finite-size effects in random energy models and in the problem of polymers in a random medium. Journal of Statistical Physics, 63(3–4):505–539, May 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01029198.

Constantin:1999:IPN

[CD99] Peter Constantin and Charles R. Doering. Infinite Prandtl number convection. Journal of Statistical Physics, 94(1–2):159–172, January 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004511312885.

Claes:1993:DPP

[CdB93] I. Claes and C. Van den Broeck. Dispersion of particles in periodic media. Journal of Statistical Physics, 70(5–6):1215–1231, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049429.

Carr:1994:ABS

[CdC94] J. Carr and F. P. da Costa. Asymptotic behavior of solutions to the coagulation-fragmentation equations. II. Weak fragmentation. Journal of Statistical Physics, 77(1–2):89–123, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186834.

Collet:1994:CAH

[CDD94] P. Collet, J. De Coninck, and F. Dunlop. Contact angle hysteresis in a solid-on-solid model. Journal of Statistical Physics, 75(1–2):37–49, April 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186279.

Collet:1997:PMD

[CDFG97] Pierre Collet, François Dunlop, Damien P. Foster, and Thierry Gobron. Product measures and dynamics for solid-on-solid interfaces. Journal of Statistical Physics, 89(3–4):509–536, November 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02765533.

Collet:1995:CLD

[CDG95] Pierre Collet, François Dunlop, and Thierry Gobron. Conservative Langevin dynamics of solid-on-solid interfaces. Journal of Statistical Physics, 79(1–2):215–229, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179387.

Chavez:1991:HID

[CdHM91] A. E. Chávez, M. López de Haro, and O. Manero. Hydrodynamic interactions of dilute polymer solutions in elongational flow. Journal of Statistical Physics, 62(5–6):1255–1266, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01128186.

Chen:1991:LGA

[CDM91] Shiyi Chen, G. D. Doolen, and W. H. Matthaeus. Lattice gas automata for simple and complex fluids. Journal of Statistical Physics, 64(5–6):1133–1162, September 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048819.

Comtet:1993:ALW

[CDM93] A. Comtet, J. Desbois, and C. Monthus. Asymptotic laws for the winding angles of planar Brownian motion. Journal of Statistical Physics, 73(1–2):433–440, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01052772.

Cvitanovic:1998:TFS

[CDMV98] Predrag Cvitanović, C. P. Dettmann, Ronnie Mainieri, and Gábor Vattay. Trace formulas for stochastic evolution operators: Weak noise perturbation theory. Journal of Statistical Physics, 93(3–4):981–999, November 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033173.38345.f9.

Carneiro:1995:GSE

[CdOW95] C. E. I. Carneiro, M. J. de Oliveira, and W. F. Wreszinski. Ground-state energy of a quantum chain with competing interactions. Journal of Statistical Physics, 79(1–2):347–376, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179393.

Caprino:1990:SPS

[CDPP90] S. Caprino, A. DeMasi, E. Presutti, and M. Pulvirenti. A stochastic particle system modeling the Carleman equation. Journal of Statistical Physics, 59(1–2):535–537, April 1990.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01015584.

Cohen:1991:NTPa

[CdS91a] E. G. D. Cohen and I. M. de Schepper. Note on transport processes in dense colloidal suspensions. Journal of Statistical Physics, 63(1–2):241–248, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01026601.

Cohen:1991:NTPb

[CdS91b] E. G. D. Cohen and I. M. de Schepper. Note on transport processes in dense colloidal suspensions. Journal of Statistical Physics, 65(1–2):419, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01329873.

Coveney:1994:CNE

[CE94] Peter V. Coveney and Allan K. Evans. Canonical nonequilibrium ensembles and subdynamics. Journal of Statistical Physics, 77(1– 2):229–258, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186842.

Ceperley:1991:FN

[Cep91] D. M. Ceperley. Fermion nodes. Journal of Statistical Physics, 63 (5–6):1237–1267, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01030009.

Cercignani:1990:TMT

[Cer90] Carlo Cercignani. Are there more than five linearly-independent collision invariants for the Boltzmann equation? Journal of Statistical Physics, 58(5–6):817–823, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01026552.

Cercignani:1994:ESK

[Cer94] Carlo Cercignani. An explicitly solvable kinetic model for semiconductors. Journal of Statistical Physics, 77(5–6):1039– 1048, December 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183150.

Cercignani:1996:TEW

[Cer96] C. Cercignani. Trend to equilibrium of weak solutions of the Boltzmann equation in a slab with diffusive boundary conditions. Journal of Statistical Physics, 84(3–4):875–888, August 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179661.

Cercignani:1997:TEK

[Cer97] Carlo Cercignani. Temperature, entropy, and kinetic theory. Journal of Statistical Physics, 87(5–6):1097–1109, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02181273.

Cichocki:1997:LTT

[CF97] B. Cichocki and B. U. Felderhof. Long-time translation and rotational Brownian motion in two dimensions. Journal of Statistical Physics, 87(5–6):989–1003, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02181267.

Christensen:1991:DSA

[CFJ91] Kim Christensen, Hans C. Fogedby, and Henrik Jeldtoft Jensen. Dynamical and spatial aspects of sandpile cellular automata. Journal of Statistical Physics, 63(3–4):653–684, May 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01029204.

Chazottes:1998:REI

[CFL98] J.-R. Chazottes, E. Floriani, and R. Lima. Relative entropy and identification of Gibbs measures in dynamical systems. Journal of Statistical Physics, 90(3–4):697–725, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023220802597.

Carmesin:1991:BNH

[CFP91] H.-O. Carmesin, H. L. Frisch, and J. K. Percus. Binary nonadditive hard-sphere mixtures at high dimension. Journal of Statistical Physics, 63(3–4):791–795, May 1991. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01029212.

Collet:1993:SCV

[CG93] P. Collet and A. Galves. Statistics of close visits to the indifferent fixed point of an interval map. Journal of Statistical Physics, 72 (3–4):459–478, August 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048020.

Carati:1999:SHF

[CG99a] Andrea Carati and Luigi Galgani. On the specific heat of Fermi– Pasta–Ulam systems and their glassy behavior. Journal of Statistical Physics, 94(5–6):859–869, March 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004531032623.

Cohen:1999:NTT

[CG99b] E. G. D. Cohen and G. Gallavotti. Note on two theorems in nonequilibrium statistical mechanics. Journal of Statistical Physics, 96(5–6):1343–1349, September 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004604804070.

Cattaneo:1994:SOE

[CGK94] Alberto S. Cattaneo, Andrea Gamba, and Igor V. Kolokolov. Statistics of the one-electron current in a one-dimensional mesoscopic ring at arbitrary magnetic fields. Journal of Statistical Physics, 76(3–4):1065–1074, August 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02188698.

Cieplak:1998:MFR

[CGM^{+}98] Marek Cieplak, Achille Giacometti, Amos Maritan, Andrea Rinaldo, Ignacio Rodriguez-Iturbe, and Jayanth R. Banavar. Models of fractal river basins. Journal of Statistical Physics, 91 (1–2):1–15, April 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023069201470.

Cesi:1996:TDS

[CGMS96] F. Cesi, G. Guadagni, F. Martinelli, and R. H. Schonmann. On the two-dimensional stochastic Ising model in the phase coexistence region near the critical point. Journal of Statistical Physics, 85(1–2):55–102, October 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02175556.

Chen:1995:DES

[CGS95] Niu-Niu Chen, M. D. Grynberg, and R. B. Stinchcombe. Deposition-evaporation stochastic systems in two and higher dimensions. Journal of Statistical Physics, 78(3–4):971–1009, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183697.

Cacciuto:1999:RGF

[CGT99] Angelo Cacciuto, Eric Gregory, and Alex Travesset. The renormalization group and its finite lattice approximations. Journal of Statistical Physics, 97(3–4):541–574, November 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004659124079.

Cirillo:1999:CFC

[CGTM99] E. N. M. Cirillo, G. Gonnella, M. Troccoli, and A. Maritan. Correlation functions by cluster variation method for Ising model with NN, NNN, and plaquette interactions. Journal of Statistical Physics, 94(1–2):67–89, January 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004507211976.

Carey:1992:MPR

[CH92] A. L. Carey and L. R. Hume. Multiple phases and return to equilibrium. Journal of Statistical Physics, 68(5–6):787–828, September 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048876.

Courbage:1994:NMR

[CH94] M. Courbage and D. Hamdan. Non-Markovian reversible Chapman–Kolmogorov measures on subshifts of finite type. Journal of Statistical Physics, 74(5–6):1281–1292, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02188231.

Chen:1990:ETR

[Che90] Mu-Fa Chen. Ergodic theorems for reaction-diffusion processes. Journal of Statistical Physics, 58(5–6):939–966, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01026558.

Chen:1991:KER

[Che91] Yong-Cong Chen. Kubo–Einstein relation for quantum Brownian motion in a periodic potential. Journal of Statistical Physics, 65(3–4):761–771, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01053753.

Chernov:1992:ESP

[Che92] N. I. Chernov. Ergodic and statistical properties of piecewise linear hyperbolic automorphisms of the 2-torus. Journal of Statistical Physics, 69(1–2):111–134, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053785.

Chernov:1993:ESP

[Che93] N. I. Chernov. Ergodic and statistical properties of piecewise linear hyperbolic automorphisms of the 2-torus. Journal of Statistical Physics, 71(1–2):341–347, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048104.

Chernov:1994:SPP

[Che94] N. I. Chernov. Statistical properties of the periodic Lorentz gas. Multidimensional case. Journal of Statistical Physics, 74 (1–2):11–53, January 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186805.

Chen:1995:HTG

[Che95] Hudong Chen. H-theorem and generalized semi-detailed balance condition for lattice gas systems. Journal of Statistical Physics, 81(1–2):347–359, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179983.

Chen:1997:CTM

[Che97a] Dayue Chen. The consensus times of the majority vote process on a torus. Journal of Statistical Physics, 86(3–4):779–802, February 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02199120.

Chernov:1997:ELE

[Che97b] N. Chernov. Entropy, Lyapunov exponents, and mean free path for billiards. Journal of Statistical Physics, 88(1–2):1–29, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02508462.

Chen:1998:LDE

[Che98] Jinwen Chen. Large deviations and ergodicity for spin particle systems. Journal of Statistical Physics, 91(1–2):369–393, April 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023056524668.

Chen:1999:VPM

[Che99a] Jinwen Chen. A variational principle for Markov processes. Journal of Statistical Physics, 96(5–6):1359–1364, September 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004608904978.

Chernov:1999:DCD

[Che99b] N. Chernov. Decay of correlations and dispersing billiards. Journal of Statistical Physics, 94(3–4):513–556, February 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004500405132.

Cushman:1994:NSM

[CHG94] John H. Cushman, Xiaolong Hu, and Timothy R. Ginn. Nonequilibrium statistical mechanics of preasymptotic dispersion. Journal of Statistical Physics, 75(5–6):859–878, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02186747.

Celik:1993:VHS

[CHK93] T. Celik, U. H. E. Hansmann, and M. Katoot. The van Hemmen spin glass revisited. Journal of Statistical Physics, 73(3–4): 775–788, November 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01054350.

Cahn:1999:AIO

[CHM99] J. W. Cahn, S. C. Han, and G. B. McFadden. Anisotropy of interfaces in an ordered HCP binary alloy. Journal of Statistical Physics, 95(5–6):1337–1360, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004583324097.

Chorin:1992:VMS

[Cho92] Alexandre Joel Chorin. A vortex model with superfluid and turbulent percolation. Journal of Statistical Physics, 69(1–2):67–78, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01053783.

Chorin:1994:VPT

[Cho94] Alexandre J. Chorin. Vortex phase transitions in 2 1/2 dimensions. Journal of Statistical Physics, 76(3–4):835–856, August 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02188688.

Choquard:1997:CSE

[Cho97] Ph. Choquard. Coulomb system equivalent to the energy spectrum of the Calogero–Sutherland–Moser (CSM) model. Journal of Statistical Physics, 89(1–2):61–68, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02770754.

Carlon:1999:IAT

[CISS99] E. Carlon, F. Iglói, W. Selke, and F. Szalma. Interfacial adsorption in two-dimensional Potts models. Journal of Statistical Physics, 96(3–4):531–543, August 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004542105635.

Chu:1992:SLQ

[CJ92] G. Chu and Jorge V. José. The semiclassical limit of a quantum Fermi accelerator. Journal of Statistical Physics, 68(1–2):153– 174, July 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048840.

Chan:1997:SPS

[CJ97] Terence Chan and Kalvis M. Jansons. Semiflexible polymers in straining flows. Journal of Statistical Physics, 88(1–2):145–176, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02508468.

Chandre:1999:AKR

[CJB99] C. Chandre, H. R. Jauslin, and G. Benfatto. An approximate KAM–renormalization-group scheme for Hamiltonian systems. Journal of Statistical Physics, 94(1–2):241–251, January 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004519514702.

Carroll:1998:DBM

[CJK98] Malcolm S. Carroll, Wolfhard Janke, and Stefan Kappler. Dynamical behavior of the multibondic and multicanonic algorithm in the 3D q-state Potts model. Journal of Statistical Physics, 90(5–6):1277–1293, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023283412473.

Childress:1991:STP

[CK91] Stephen Childress and Isaac Klapper. On some transport properties of Baker’s maps. Journal of Statistical Physics, 63(5–6):897– 914, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01029990.

Chiu:1995:TTR

[CK95] Hung-Chang Chiu and David A. Kofke. Transformation and topological reduction of cluster expansions using m-bonds. Journal of Statistical Physics, 78(3–4):877–892, February 1995.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02183692.

Ching:1998:ERC

[CK98] Emily S. C. Ching and Robert H. Kraichnan. Exact results for conditional means of a passive scalar in certain statistically homogeneous flows. Journal of Statistical Physics, 93(3–4): 787–795, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000033163.33811.5a.

Cercignani:1999:RCI

[CK99a] C. Cercignani and G. M. Kremer. On relativistic collisional invariants. Journal of Statistical Physics, 96(1–2):439–445, July 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1004545104959.

Courbage:1999:IRK

[CK99b] M. Courbage and B. Kamiński. Intrinsical randomness of Kolmogorov Z^{d}-actions on a Lebesgue space. Journal of Statistical Physics, 97(3–4):781–792, November 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004623510875.

Contucci:1999:FMP

[CKK99] Pierluigi Contucci, Peter Kleban, and Andreas Knauf. A fully magnetizing phase transition. Journal of Statistical Physics, 97(3–4):523–539, November 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004607107241.

Clement:1991:DLR

[CKS91] E. Clément, R. Kopelman, and L. M. Sander. The diffusionlimited reaction a+b->0 on a fractal substrate. Journal of Statistical Physics, 65(5–6):919–924, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01049589.

Chen:1995:GKM

[CL95] Shiyi Chen and Turab Lookman. Growth kinetics in multicomponent fluids. Journal of Statistical Physics, 81(1–2):223–235, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179977.

Chernov:1997:SNS

[CL97] N. I. Chernov and J. L. Lebowitz. Stationary nonequilibrium states in boundary-driven Hamiltonian systems: Shear flow. Journal of Statistical Physics, 86(5–6):953–990, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02183610.

Cirillo:1998:MTD

[CL98] Emilio N. M. Cirillo and Joel L. Lebowitz. Metastability in the two-dimensional Ising model with free boundary conditions. Journal of Statistical Physics, 90(1–2):211–226, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023255802455.

Cladis:1991:DTP

[Cla91a] P. E. Cladis. A dynamical test of phase transition order: New things in old places or old wine in new bottles. Journal of Statistical Physics, 62(5–6):899–925, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01128168.

Cladis:1991:PPW

[Cla91b] P. E. Cladis. A prepared pattern with wavelength selection in directional solidification. Journal of Statistical Physics, 64(5–6): 1103–1119, September 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048817.

Chopard:1994:MAF

[CLD94] Bastien Chopard, Pascal Luthi, and Michel Droz. Microscopic approach to the formation of Liesegang patterns. Journal of Statistical Physics, 76(1–2):661–677, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02188680.

Clement:1991:ASI

[CLHS91] E. Clément, P. Leroux-Hugon, and L. M. Sander. Analytical solution of an irreversible surface reaction model. Journal of Statistical Physics, 65(5–6):925–939, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01049590.

Carmona:1998:LLD

[CLLL98] S. C. Carmona, C. Landim, A. Lopes, and S. Lopes. A level 1 large-deviation principle for the autocovariances of uniquely ergodic transformations with additive noise. Journal of Statistical Physics, 91(1–2):395–421, April 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023008608738.

Chennaoui:1990:MIG

[CLS90] A. Chennaoui, J. Liebler, and H. G. Schuster. The mechanism of the increase of the generalized dimension of a filtered chaotic time series. Journal of Statistical Physics, 59(5–6):1311–1328, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01334753.

Chechetkin:1990:MSFa

[CLT90a] V. R. Chechetkin, V. S. Lutovinov, and A. Yu. Turygin. Multifractal structure of fully developed hydrodynamic turbulence. I. Kolmogorov’s third hypothesis revisited. Journal of Statistical Physics, 61(3–4):573–588, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01027292.

Chechetkin:1990:MSFb

[CLT90b] V. R. Chechetkin, V. S. Lutovinov, and A. Yu. Turygin. Multifractal structure of fully developed hydrodynamic turbulence. II. Intermittency effects in the distribution of passive scalar impurities. Journal of Statistical Physics, 61(3–4):589–605, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01027293.

Celletti:1997:CDP

[CLV97] Alessandra Celletti, Victoria M. Bajo Lorenzana, and Alessandro E. P. Villa. Correlation dimension for paired discrete time series. Journal of Statistical Physics, 89(3–4):877–884, November 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/

10.1007/BF02765549. See erratum [CLV98].

Celletti:1998:ECD

[CLV98] Alessandra Celletti, Victoria M. Bajo Lorenzana, and Alessandro E. P. Villa. Erratum: Correlation Dimension for Paired Discrete Time Series. Journal of Statistical Physics, 92(1–2):331–332, July 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023060006032. See [CLV97].

Chou:1992:FEM

[CLY92] Dong-Pao Chou, Thomas Lackner, and Sidney Yip. Fluctuation effects in models of adiabatic explosion. Journal of Statistical Physics, 69(1–2):193–215, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053790.

Costantini:1992:HLB

[CM92] C. Costantini and R. Marra. Hydrodynamic limits for the Boltzmann process. Journal of Statistical Physics, 67(1–2):229–249, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049032.

Chayes:1995:BST

[CM95] L. Chayes and J. Machta. On the behavior of the surface tension for spin systems in a correlated porous medium. Journal of Statistical Physics, 79(1–2):117–164, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179384.

Cesi:1996:LTS

[CM96] Filippo Cesi and Fabio Martinelli. On the layering transition of an SOS surface interacting with a wall. I. Equilibrium results. Journal of Statistical Physics, 82(3–4):823–913, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179794.

Caglioti:1998:TAS

[CM98] E. Caglioti and C. Maffei. Time asymptotics for solutions of Vlasov–Poisson equation in a circle. Journal of Statistical Physics, 92(1–2):301–323, July 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023055905124.

Chen:1992:MCL

[CMMC92] Shiyi Chen, Daniel O. Martínez, W. H. Matthaeus, and Hudong Chen. Magnetohydrodynamics computations with lattice gas automata. Journal of Statistical Physics, 68(3–4):533–556, August 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01341761.

Cassandro:1995:CCT

[CMP95] M. Cassandro, R. Marra, and E. Presutti. Corrections to the critical temperature in 2D Ising systems with Kac potentials. Journal of Statistical Physics, 78(3–4):1131–1138, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02183705.

Cassandro:1997:UBC

[CMP97] M. Cassandro, R. Marra, and E. Presutti. Upper bounds on the critical temperature for Kac potentials. Journal of Statistical Physics, 88(3–4):537–566, August 1997. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000015163.27899.8f.

Cucchieri:1997:CLE

[CMPS97] Attilio Cucchieri, Tereza Mendes, Andrea Pelissetto, and Alan D. Sokal. Continuum limits and exact finite-size-scaling functions for one-dimensional O(N)-invariant spin models. Journal of Statistical Physics, 86(3–4):581–673, February 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02199114.

Chayes:1998:GRI

[CMR98] L. Chayes, J. Machta, and O. Redner. Graphical representations for Ising systems in external fields. Journal of Statistical Physics, 93(1–2):17–32, October 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1023/B%3AJOSS.0000026726.43558.80.pdf.

Cerchiai:1995:TDI

[CMVG95] B. L. Cerchiai, M. Martellini, and F. Valz-Gris. Threedimensional integrable models and associated tangle invariants. Journal of Statistical Physics, 78(3–4):1083–1109, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183703.

Cahn:1994:EEP

[CNC94] J. W. Cahn and A. Novick-Cohen. Evolution equations for phase separation and ordering in binary alloys. Journal of Statistical Physics, 76(3–4):877–909, August 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02188691.

Campbell:1991:STD

[CO91] L. J. Campbell and Kevin O’Neil. Statistics of two-dimensional point vortices and high-energy vortex states. Journal of Statistical Physics, 65(3–4):495–529, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053742.

Cirillo:1996:MNB

[CO96a] Emilio N. M. Cirillo and Enzo Olivieri. Metastability and nucleation for the Blume–Capel model. Different mechanisms of transition. Journal of Statistical Physics, 83(3–4):473–554, May 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02183739.

Conlon:1996:BMV

[CO96b] Joseph G. Conlon and Peder A. Olsen. A Brownian motion version of the directed polymer problem. Journal of Statistical Physics, 84(3–4):415–454, August 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179650.

Cirillo:1997:RGC

[CO97] Emilio N. M. Cirillo and Enzo Olivieri. Renormalization group at criticality and complete analyticity of constrained models: A numerical study. Journal of Statistical Physics, 86(5–6):1117–1151, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183617.

Chen:1995:HTL

[COA95] Y. Chen, H. Ohashi, and M. Akiyama. Heat transfer in lattice BGK modeled fluid. Journal of Statistical Physics, 81(1– 2):71–85, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179969.

Collet:1998:AEL

[Col98] P. Collet. Amplitude equation for lattice maps, a renormalization group approach. Journal of Statistical Physics, 90(5–6): 1075–1105, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023212925677.

Combescure:1990:SPP

[Com90] M. Combescure. Spectral properties of a periodically kicked quantum Hamiltonian. Journal of Statistical Physics, 59(3– 4):679–690, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01025846.

Combescure:1991:RVD

[Com91a] M. Combescure. Recurrent versus diffusive dynamics for a kicked quantum system. Journal of Statistical Physics, 62(3–4):779–791, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017983.

Compagner:1991:HCR

[Com91b] Aaldert Compagner. The hierarchy of correlations in random binary sequences. Journal of Statistical Physics, 63(5–6):883– 896, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01029989.

Conlon:1990:UBC

[Con90] Joseph G. Conlon. An upper bound on the critical temperature for a continuous system with short-range interaction. Journal of Statistical Physics, 58(1–2):265–293, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020294.

Contucci:1996:ABT

[Con96] Pierluigi Contucci. An analyticity bound for two-dimensional Ising model at low temperature. Journal of Statistical Physics, 82(5–6):1647–1657, March 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02183397.

Constantin:1998:APN

[Con98a] Peter Constantin. Absence of proper nondegenerate generalized self-similar singularities. Journal of Statistical Physics, 93(3–4): 777–786, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000033162.96173.50.

Constantin:1998:SEA

[Con98b] Peter Constantin. Scaling exponents for active scalars. Journal of Statistical Physics, 90(3–4):571–595, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023264617618.

Coppersmith:1998:CCC

[Cop98] S. N. Coppersmith. Chaos, complexity, and computers: Objectoriented programming and physics concepts for undergraduates. Journal of Statistical Physics, 93(3–4):1009–1018, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000033175.39924.e7.

Cornille:1995:HDB

[Cor95] H. Cornille. Hexagonal discrete Boltzmann models with and without rest particles. Journal of Statistical Physics, 81(1–2): 335–346, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179982.

Cornille:1993:RCS

[CP93] H. Cornille and T. Platkowski. Riccati-coupled similarity shock wave solutions for multispeed discrete Boltzmann models. Journal of Statistical Physics, 71(3–4):719–740, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01058444.

Chen:1997:LTB

[CP97] Zhi-Min Chen and W. G. Price. Long-time behavior of Navier– Stokes flow on a two-dimensional torus excited by an external sinusoidal force. Journal of Statistical Physics, 86(1–2):301–335, January 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02180208.

Chvosta:1999:ODD

[CP99] Petr Chvosta and Noëlle Pottier. One-dimensional diffusion in a semiinfinite Poisson random force. Journal of Statistical Physics, 97(1–2):323–349, October 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004627404379.

Caracciolo:1994:RWS

[CPP94] Sergio Caracciolo, Giorgo Parisi, and Andrea Pelissetto. Random walks with short-range interaction and mean-field behavior. Journal of Statistical Physics, 77(3–4):519–543, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179448.

Chayes:1997:NDF

[CPP97] L. Chayes, Robin Pemantle, and Yuval Peres. No directed fractal percolation in zero area. Journal of Statistical Physics, 88(5–6): 1353–1362, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02732437.

Ciliberto:1991:RDT

[CPPG91] S. Ciliberto, E. Pampaloni, and C. Perez-Garcia. The role of defects in the transition between different symmetries in convective patterns. Journal of Statistical Physics, 64(5–6):1045– 1058, September 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048813.

Caracciolo:1990:NMC

[CPS90] Sergio Caracciolo, Andrea Pelissetto, and Alan D. Sokal. Nonlocal Monte Carlo algorithm for self-avoiding walks with fixed endpoints. Journal of Statistical Physics, 60(1–2):1–53, July 1990.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013668.

Caracciolo:1991:DCE

[CPS91] Sergio Caracciolo, Andrea Pelissetto, and Alan D. Sokal. Dynamic critical exponent of the BFACF algorithm for self-avoiding walks. Journal of Statistical Physics, 63(5–6):857–865, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01029987.

Caracciolo:1992:JCA

[CPS92] Sergio Caracciolo, Andrea Pelissetto, and Alan D. Sokal. Joinand-cut algorithm for self-avoiding walks with variable length and free endpoints. Journal of Statistical Physics, 67(1–2): 65–111, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049027.

Cornille:1990:TOV

[CQ90] Henri Cornille and Yue-Hong Qian. Temperature overshoots for a 4-velocity unidimensional discrete Boltzmann model. Journal of Statistical Physics, 61(3–4):683–712, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01027297.

Crisanti:1994:RBI

[CR94] A. Crisanti and H. Rieger. Random-bond Ising chain in a transverse magnetic field: A finite-size scaling analysis. Journal of Statistical Physics, 77(5–6):1087–1098, December 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183154.

Covert:1997:HLP

[CR97] Paul Covert and Fraydoun Rezakhanlou. Hydrodynamic limit for particle systems with nonconstant speed parameter. Journal of Statistical Physics, 88(1–2):383–426, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02508477.

Crawford:1994:AEI

[Cra94] John David Crawford. Amplitude expansions for instabilities in populations of globally-coupled oscillators. Journal of Statistical Physics, 74(5–6):1047–1084, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02188217.

Crooks:1998:NMF

[Cro98] Gavin E. Crooks. Nonequilibrium measurements of free energy differences for microscopically reversible Markovian systems. Journal of Statistical Physics, 90(5–6):1481–1487, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1023208217925.

Cheng:1990:SFS

[CS90a] Zheming Cheng and Robert Savit. Structure factor of substitutional sequences. Journal of Statistical Physics, 60(3–4):383–393, August 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01314927.

Chowdhury:1990:SMI

[CS90b] Debashish Chowdhury and Dietrich Stauffer. Systematics of the models of immune response and autoimmune disease. Journal of Statistical Physics, 59(3–4):1019–1042, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01025860.

Conlon:1991:RWR

[CS91a] Joseph G. Conlon and Jan Philip Solovej. Random walk representations of the Heisenberg model. Journal of Statistical Physics, 64(1–2):251–270, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01057876.

Conlon:1991:UCF

[CS91b] Joseph G. Conlon and Jan Philip Solovej. Uniform convergence of the free energy of the classical Heisenberg model to that of the Gaussian model. Journal of Statistical Physics, 65(1–2):235–245, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01329858.

Conrad:1991:CME

[CS91c] Hans Conrad and Arnold F. Sprecher. Characteristics and mechanisms of electrorheological fluids. Journal of Statistical Physics, 64(5–6):1073–1091, September 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01048815.

Cercignani:1992:HRC

[CS92] C. Cercignani and C. Sgarra. Half-range completeness for the Fokker–Planck equation with an external force. Journal of Statistical Physics, 66(5–6):1575–1582, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01054434.

Condat:1997:SST

[CSB97] C. A. Condat, G. J. Sibona, and C. E. Budde. Steady state in two-dimensional diffusion-controlled reactions. Journal of Statistical Physics, 89(1–2):369–377, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02770770.

Careta:1993:EDS

[CSRPS93] A. Careta, F. Sagués, L. Ramirez-Piscina, and J. M. Sancho. Effective diffusion in a stochastic velocity field. Journal of Statistical Physics, 71(1–2):235–242, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048097.

Chayes:1995:LLV

[CSS95] L. Chayes, R. H. Schonmann, and G. Swindle. Lifshitz’ law for the volume of a two-dimensional droplet at zero temperature. Journal of Statistical Physics, 79(5–6):821–831, June 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02181205.

Calisto:1993:EVK

[CT93] H. Calisto and E. Tirapegui. Ω-expansion of van Kampen through functional integrals. Journal of Statistical Physics, 71 (3–4):683–703, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01058442.

Cinlar:1995:EDT

[CT95] E. Cinlar and S. Torquato. Exact determination of the twopoint cluster function for one-dimensional continuum percolation. Journal of Statistical Physics, 78(3–4):827–839, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183690.

Chamati:1996:FSS

[CT96a] H. Chamati and N. S. Tonchev. Finite-size shift of the critical temperature in the spherical model. Journal of Statistical Physics, 83(5–6):1211–1218, June 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179559.

Chernov:1996:MIL

[CT96b] N. I. Chernov and S. Troubetzkoy. Measures with infinite Lyapunov exponents for the periodic Lorentz gas. Journal of Statistical Physics, 83(1–2):193–202, April 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183645.

Cardy:1998:FTB

[CT98a] John L. Cardy and Uwe C. Täuber. Field theory of branching and annihilating random walks. Journal of Statistical Physics, 90(1–2):1–56, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023233431588.

Choi:1998:CSF

[CT98b] Ho Suk Choi and Julian Talbot. Coverage and structure of films of spherical particles deposited after diffusing in a gravitational field. Journal of Statistical Physics, 92(5–6):891–908, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023040509781.

Cutler:1991:SRBa

[Cut91a] C. D. Cutler. Some results on the behavior and estimation of the fractal dimensions of distributions on attractors. Journal of Statistical Physics, 62(3–4):651–708, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01017978.

Cutler:1991:SRBb

[Cut91b] C. D. Cutler. Some results on the behavior and estimation of the fractal dimensions of distributions on attractors. Journal of Statistical Physics, 65(1–2):417–418, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01329872.

Chirikov:1993:TFA

[CV93a] B. V. Chirikov and V. V. Vecheslavov. Theory of fast Arnold diffusion in many-frequency systems. Journal of Statistical Physics, 71(1–2):243–258, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048098.

Crisanti:1993:END

[CV93b] A. Crisanti and A. Vulpiani. On the effects of noise and drift on diffusion in fluids. Journal of Statistical Physics, 70(1–2): 197–211, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053963.

Celletti:1996:DCA

[CV96] Alessandra Celletti and Alessandro E. P. Villa. Determination of chaotic attractors in the rat brain. Journal of Statistical Physics, 84(5–6):1379–1385, September 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02174137.

Conlon:1998:NSR

[CvD98] Joseph G. Conlon and Brian von Dohlen. Numerical simulations of random walk in random environment. Journal of Statistical Physics, 92(3–4):571–586, August 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023088504988.

Cohen:1995:NRD

[CW95] E. G. D. Cohen and F. Wang. New results for diffusion in Lorentz lattice gas cellular automata. Journal of Statistical Physics, 81 (1–2):445–466, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179988.

Canright:1996:DGS

[CW96] Geoff Canright and Greg Watson. Disordered ground states for classical discrete-state problems in one dimension. Journal of Statistical Physics, 84(5–6):1095–1131, September 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02174130.

Coutsias:1997:NTC

[CWP97] E. A. Coutsias, M. J. Wester, and A. S. Perelson. A nucleation theory of cell surface capping. Journal of Statistical Physics, 87 (5–6):1179–1203, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181279.

Chen:1992:LBC

[CWSD92] Shiyi Chen, Zheng Wang, Xiaowen Shan, and Gary D. Doolen. Lattice Boltzmann computational fluid dynamics in three dimensions. Journal of Statistical Physics, 68(3–4):379–400, August 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01341754.

Derrida:1999:ULD

[DA99] B. Derrida and C. Appert. Universal large-deviation function of the Kardar–Parisi–Zhang equation in one dimension. Journal of Statistical Physics, 94(1–2):1–30, January 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004599526997.

Bonfim:1991:FSE

[dAB91a] O. F. de Alcantara Bonfim. Finite-size effects and phase transition in the three-dimensional three-state Potts model. Journal of Statistical Physics, 62(1–2):105–115, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020861.

Braga:1991:NRU

[dAB91b] Gastão de Almeida Braga. A new rigorous upper bound for the inverse critical temperature of the two-dimensional Coulomb gas. Journal of Statistical Physics, 63(1–2):285–289, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01026604.

deAguiar:1990:PDO

[dABMR90] F. S. de Aguiar, F. A. Bosco, A. S. Martinez, and S. Goulart Rosa, Jr. Phase diagram of the one-state Potts model on the Cayley tree. Journal of Statistical Physics, 58(5–6):1231–1238, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01026573.

deAguiar:1991:MPM

[dABR91] F. S. de Aguiar, L. B. Bernardes, and S. Goulart Rosa, Jr. Metastability in the Potts model on the Cayley tree. Journal of Statistical Physics, 64(3–4):673–682, August 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01048311.

Dagaeff:1996:IPB

[Dag96] T. Dagaeff. Integrability and pseudointegrability in billiards illustrated by the harmonic wedge. Journal of Statistical Physics, 83(1–2):39–70, April 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183639.

Dahlqvist:1996:LEA

[Dah96] Per Dahlqvist. Lyapunov exponents and anomalous diffusion of a Lorentz gas with infinite horizon using approximate zeta functions. Journal of Statistical Physics, 84(3–4):773–795, August 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179657.

Daido:1990:IFP

[Dai90] Hiroaki Daido. Intrinsic fluctuations and a phase transition in a class of large populations of interacting oscillators. Journal of Statistical Physics, 60(5–6):753–800, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01025993.

Dalitz:1997:HSP

[Dal97] Ch. Dalitz. Half-space problem of the Boltzmann equation for charged particles. Journal of Statistical Physics, 88(1–2):129– 144, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02508467.

Danchev:1993:FSD

[Dan93] Daniel Danchev. Finite-size dependence of the helicity modulus within the mean spherical model. Journal of Statistical Physics, 73(1–2):267–292, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01052761.

Dasgupta:1995:IMU

[Das95] Subinary Dasgupta. Ising metamagnet under staggered field. Journal of Statistical Physics, 81(3–4):837–842, November 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179260.

Davies:1991:TCP

[Dav91] Brian Davies. A twist on chiral Potts. Journal of Statistical Physics, 62(1–2):89–104, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020860.

Duarte:1990:VDC

[DB90a] J. A. M. S. Duarte and U. Brosa. Viscous drag by cellular automata. Journal of Statistical Physics, 59(1–2):501–508, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01015579.

Duering:1990:CDT

[DB90b] Edgardo Duering and David J. Bergman. Current distribution on a three-dimensional, bond-diluted, random-resistor network at the percolation threshold. Journal of Statistical Physics, 60 (3–4):363–381, August 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01314926.

denBroeck:1991:RTM

[dB91] C. Van den Broeck. Reaction, trapping, and multifractality in one-dimensional systems. Journal of Statistical Physics, 65(5– 6):971–990, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049593.

denBroeck:1992:BRS

[dB92] C. Van den Broeck. Book review: Statistical mechanics of neural networks. Journal of Statistical Physics, 66(5–6):1683–1684, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01054445.

Duering:1992:CDT

[DBB^{+}92] E. Duering, R. Blumenfeld, D. J. Bergman, A. Aharony, and M. Murat. Current distributions in a two-dimensional randomresistor network. Journal of Statistical Physics, 67(1–2):113–121, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049028.

Dufty:1997:I

[DBD97] J. Dufty, R. Brito, and J. R. Dorfman. Introduction. Journal of Statistical Physics, 87(5–6):975–979, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02181265.

Doering:1991:FCD

[DBH91] Charles R. Doering, Martin A. Burschka, and Werner Horsthemke. Fluctuations and correlations in a diffusion-reaction system: Exact hydrodynamics. Journal of Statistical Physics, 65(5– 6):953–970, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049592.

Deem:1994:CDS

[DC94] Michael W. Deem and David Chandler. Classical diffusion in strong random media. Journal of Statistical Physics, 76(3–4): 911–927, August 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02188692.

Dhulst:1998:ODH

[DC98] R. D’hulst and A. Caillé. One-dimensional Hamiltonian for columnar liquid crystals. Journal of Statistical Physics, 92(5– 6):865–889, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023088425711.

Deshpande:1991:HNN

[DD91] Varsha Deshpande and Chandan Dasgupta. Hierarchical neural networks for the storage of correlated memories. Journal of Statistical Physics, 64(3–4):755–779, August 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01048314.

Dumas:1996:MFP

[DDG96] H. S. Dumas, L. Dumas, and F. Golse. On the mean free path for a periodic array of spherical obstacles. Journal of Statistical Physics, 82(5–6):1385–1407, March 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183388.

Dumans:1997:RNM

[DDG97] H. S. Dumans, L. Dumas, and F. Golse. Remarks on the notion of mean free path for a periodic array of spherical obstacles. Journal of Statistical Physics, 87(3–4):943–950, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02181255.

Daumer:1997:QPF

[DDGZ97] M. Daumer, D. Dürr, S. Goldstein, and N. Zanghi. On the quantum probability flux through surfaces. Journal of Statistical Physics, 88(3–4):967–977, August 1997. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000015181.86864.fb.

DaiPra:1996:MVL

[DdH96] Paolo Dai Pra and Frank den Hollander. McKean–Vlasov limit for interacting random processes in random media. Journal of Statistical Physics, 84(3–4):735–772, August 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179656.

DeSimone:1995:EGS

[DDJ^{+}95] C. De Simone, M. Diehl, M. Jünger, P. Mutzel, G. Reinelt, and G. Rinaldi. Exact ground states of Ising spin glasses: New experimental results with a branch-and-cut algorithm. Journal of Statistical Physics, 80(1–2):487–496, July 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02178370.

DeConinck:1990:ERM

[DDM90] Joël De Coninck, François Dunlop, and Frédéric Menu. Exact results for a meniscus in a three-phase system within an SOS-type approximation. Journal of Statistical Physics, 61(5–6): 1121–1139, December 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014368.

Derrida:1992:ESO

[DDM92] B. Derrida, E. Domany, and D. Mukamel. An exact solution of a one-dimensional asymmetric exclusion model with open boundaries. Journal of Statistical Physics, 69(3–4):667–687, November 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01050430.

DiMeo:1996:NSL

[DE96] M. Di Meo and R. Esposito. The Navier–Stokes limit of the stationary Boltzmann equation for hard potentials. Journal of Statistical Physics, 84(3–4):859–873, August 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179660.

Dudynski:1992:GEP

[DEJ92] Marek Dudyński and Maria L. Ekiel-Jezewska. Global existence proof for relativistic Boltzmann equation. Journal of Statistical Physics, 66(3–4):991–1001, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01055712.

Dorfman:1995:DCL

[DEJ95] J. R. Dorfman, M. H. Ernst, and D. Jacobs. Dynamical chaos in the Lorentz lattice gas. Journal of Statistical Physics, 81(1– 2):497–513, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179990.

Derrida:1995:EDC

[DEM95] B. Derrida, M. R. Evans, and K. Mallick. Exact diffusion constant of a one-dimensional asymmetric exclusion model with open boundaries. Journal of Statistical Physics, 79(5–6):833–874, June 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02181206.

DeMasi:1992:KLH

[DEP92] Anna De Masi, Raffaele Esposito, and Errico Presutti. Kinetic limits of the HPP cellular automaton. Journal of Statistical Physics, 66(1–2):403–464, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01060074.

Dermoune:1997:IBE

[Der97] A. Dermoune. The inviscid Burgers equation with initial value of Poissonian type. Journal of Statistical Physics, 88(3–4):873–883, August 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000015176.78898.cd.

Deutsch:1992:OAC

[Deu92] Hans-Peter Deutsch. Optimized analysis of the critical behavior in polymer mixtures from Monte Carlo simulations. Journal of Statistical Physics, 67(5–6):1039–1082, June 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01049009.

Devillard:1991:IMT

[Dev91a] Pierre Devillard. Interface motion in a two-dimensional Ising model with a field. Journal of Statistical Physics, 62(1–2): 443–451, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020880.

Devillard:1991:STM

[Dev91b] Pierre Devillard. Statistics of transfer matrices for disordered quantum thin metallic slabs. Journal of Statistical Physics, 62 (1–2):373–387, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020873.

Dettmann:1993:PTA

[DF93] C. P. Dettmann and N. E. Frankel. potential theory and analytic properties of self-similar fractal and multifractal distributions. Journal of Statistical Physics, 72(1–2):241–275, July 1993.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01048049.

DeBievre:1998:TPK

[DF98] S. De Bièvre and G. Forni. Transport properties of kicked and quasiperiodic Hamiltonians. Journal of Statistical Physics, 90(5– 6):1201–1223, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023227327494.

Datta:1996:LTP

[DFF96] Nilanjana Datta, Roberto Fernández, and Jürg Fröhlich. Lowtemperature phase diagrams of quantum lattice systems. I. Stability for quantum perturbations of classical systems with finitely-many ground states. Journal of Statistical Physics, 84 (3–4):455–534, August 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179651.

Datta:1999:EHP

[DFF99] Nilanjana Datta, Roberto Fernández, and Jürg Fröhlich. Effective Hamiltonians and phase diagrams for tight-binding models. Journal of Statistical Physics, 96(3–4):545–611, August 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004594122474.

Dubrulle:1990:LVL

[DFHR90] B. Dubrulle, U. Frisch, M. Hénon, and J.-P. Rivet. Low-viscosity lattice gases. Journal of Statistical Physics, 59(5–6):1187–1226, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01334747.

DeConinck:1994:EST

[DFZ94] J. De Coninck, J. Fruttero, and A. Ziermann. The equilibrium shape of a two-dimensional crystal between parallel planes. Journal of Statistical Physics, 74(5–6):1255–1264, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02188228.

Debauche:1990:SKL

[DG90] M. Débauche and H. Giacomini. Susceptibility of the Kagomé lattice Ising model. Journal of Statistical Physics, 58(5–6): 1127–1135, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01026567.

Derrida:1998:SEM

[DGLS98] B. Derrida, S. Goldstein, J. L. Lebowitz, and E. R. Speer. Shift equivalence of measures and the intrinsic structure of shocks in the asymmetric simple exclusion process. Journal of Statistical Physics, 93(3–4):547–571, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033241.54615.43.

deGier:1997:ESO

[dGN97] J. de Gier and B. Nienhuis. The exact solution of an octagonal rectangle-triangle random tiling. Journal of Statistical Physics, 87(1–2):415–437, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181494.

Durr:1992:QCC

[DGZ92a] Detlef Dürr, Sheldon Goldstein, and Nino Zanghi. Quantum chaos, classical randomness, and Bohmian mechanics. Journal of Statistical Physics, 68(1–2):259–270, July 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01048845.

Durr:1992:QEO

[DGZ92b] Detlef Dürr, Sheldon Goldstein, and Nino Zanghí. Quantum equilibrium and the origin of absolute uncertainty. Journal of Statistical Physics, 67(5–6):843–907, June 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01049004.

Dimock:1992:RGA

[DH92] J. Dimock and T. R. Hurd. A renormalization group analysis of correlation functions for the dipole gas. Journal of Statistical Physics, 66(5–6):1277–1318, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01054423.

denHollander:1994:TCK

[dH94] F. den Hollander. On three conjectures by K. E. Shuler. Journal of Statistical Physics, 75(5–6):891–918, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02186749.

Dhara:1997:ESN

[Dha97] Asish K. Dhara. Enhancement of signal-to-noise ratio. Journal of Statistical Physics, 87(1–2):251–271, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02181487.

denHollander:1999:NTV

[dHMP99] F. den Hollander, M. V. Menshikov, and S. Yu. Popov. A note on transience versus recurrence for a branching random walk in random environment. Journal of Statistical Physics, 95(3–4):587– 614, May 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004539225064.

denHollander:1992:IPS

[dHNR92a] F. den Hollander, J. Naudts, and F. Redig. Invariance principle for the stochastic Lorentz lattice gas. Journal of Statistical Physics, 66(5–6):1583–1598, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01054435.

denHollander:1992:LTTb

[dHNR92b] F. den Hollander, J. Naudts, and F. Redig. Long-time tails in a random diffusion model. Journal of Statistical Physics, 69(3–4): 731–762, November 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01050432.

denHollander:1994:DSF

[dHNR94] Frank den Hollander, Jan Naudts, and Frank Redig. Dynamic structure factor in a random diffusion model. Journal of Statistical Physics, 76(5–6):1267–1285, September 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02187062.

denHollander:1992:LTTa

[dHNS92] F. den Hollander, J. Naudts, and P. Scheunders. A long-time tail for random walk in random scenery. Journal of Statistical Physics, 66(5–6):1527–1555, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01054432.

Derrida:1996:EEN

[DHP96] Bernard Derrida, Vincent Hakim, and Vincent Pasquier. Exact exponent for the number of persistent spins in the zerotemperature dynamics of the one-dimensional Potts model. Journal of Statistical Physics, 85(5–6):763–797, December 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02199362.

denHollander:1992:RWR

[dHS92] Frank den Hollander and Kurt E. Shuler. Random walks in a random field of decaying traps. Journal of Statistical Physics, 67(1–2):13–31, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049025.

Dean:1998:BEO

[DHS98a] D. S. Dean, R. R. Horgan, and D. Sentenac. Boundary effects in the one-dimensional Coulomb gas. Journal of Statistical Physics, 90(3–4):899–926, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023241407140.

denHollander:1998:RHI

[dHS98b] F. den Hollander and J. M. Swart. Renormalization of hierarchically interacting isotropic diffusions. Journal of Statistical Physics, 93(1–2):243–291, October 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%
3AJOSS.0000026734.93723.b9.pdf.

Derrida:1992:EDT

[DHV92] B. Derrida, V. Hakim, and J. Vannimenus. Effect of disorder on two-dimensional wetting. Journal of Statistical Physics, 66(5– 6):1189–1213, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01054419.

Dash:1999:DFG

[DHW99] J. G. Dash, V. A. Hodgkin, and J. S. Wettlaufer. Dynamics of faceted grain boundary grooves. Journal of Statistical Physics, 95 (5–6):1311–1322, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004579223189.

DeFelice:1993:CSR

[DI93] Louis J. DeFelice and Aurora Isaac. Chaotic states in a random world: Relationship between the nonlinear differential equations of excitability and the stochastic properties of ion channels. Journal of Statistical Physics, 70(1–2):339–354, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01053972.

Diao:1994:NSK

[Dia94] Yuanan Diao. The number of smallest knots on the cubic lattice. Journal of Statistical Physics, 74(5–6):1247–1254, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02188227.

Doi:1998:SAS

[DIK98] Shinji Doi, Junko Inoue, and Sadatoshi Kumagai. Spectral analysis of stochastic phase lockings and stochastic bifurcations in the sinusoidally forced van der Pol oscillator with additive noise. Journal of Statistical Physics, 90(5–6):1107–1127, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023271109747.

Dimock:1990:CES

[Dim90] J. Dimock. A cluster expansion for stochastic lattice fields. Journal of Statistical Physics, 58(5–6):1181–1207, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01026571.

Ding:1996:CIM

[Din96] J. Ding. Computing invariant measures of piecewise convex transformations. Journal of Statistical Physics, 83(3–4):623–635, May 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183742.

Dean:1993:BEC

[DJ93] David S. Dean and Kalvis M. Jansons. Brownian excursions on combs. Journal of Statistical Physics, 70(5–6):1313–1332, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049434.

Dean:1995:EPE

[DJ95] David S. Dean and Kalvis M. Jansons. Excursions for polymers in elongational flows. Journal of Statistical Physics, 79 (1–2):265–297, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179390.

Dillmann:1998:FSS

[DJB98] O. Dillmann, W. Janke, and K. Binder. Finite-size scaling in the p-state mean-field Potts glass: A Monte Carlo investigation. Journal of Statistical Physics, 92(1–2):57–100, July 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023043602398.

Derrida:1993:EST

[DJLS93] B. Derrida, S. A. Janowsky, J. L. Lebowitz, and E. R. Speer. Exact solution of the totally asymmetric simple exclusion process: Shock profiles. Journal of Statistical Physics, 73(5–6): 813–842, December 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01052811.

Derrida:1999:GTC

[DJM99] B. Derrida and B. Jung-Muller. The genealogical tree of a chromosome. Journal of Statistical Physics, 94(3–4):277–298, February 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004579800589.

Dowker:1996:CHA

[DK96] Fay Dowker and Adrian Kent. On the consistent histories approach to quantum mechanics. Journal of Statistical Physics, 82(5–6):1575–1646, March 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02183396.

Diamond:1996:BAC

[DKKP96] Phil Diamond, Anthony Klemm, Peter Kloeden, and Aleksej Pokrovskii. Basin of attraction of cycles of discretizations of dynamical systems with SRB invariant measures. Journal of Statistical Physics, 84(3–4):713–733, August 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179655.

Dasmahapatra:1994:VCB

[DKMM94] Srinandan Dasmahapatra, Rinat Kedem, Barry M. McCoy, and Ezer Melzer. Virasoro characters from Bethe equations for the critical ferromagnetic three-state Potts model. Journal of Statistical Physics, 74(1–2):239–274, January 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02186814.

Dobrushin:1993:MJW

[DKS93] R. L. Dobrushin, R. Kotecký, and S. B. Shlosman. A microscopic justification of the Wulff construction. Journal of Statistical Physics, 72(1–2):1–14, July 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01048037.

Dokholyan:1998:SDS

[DLB^{+}98] Nikolay V. Dokholyan, Youngki Lee, Sergey V. Buldyrev, Shlomo Havlin, Peter R. King, and H. Eugene Stanley. Scaling of the distribution of shortest paths in percolation. Journal of Statistical Physics, 93(3–4):603–613, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033244.13545.da.

delaLlave:1992:RGE

[dlL92] Rafael de la Llave. A renormalization group explanation of numerical observations of analyticity domains. Journal of Statistical Physics, 66(5–6):1631–1634, March 1992. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01054438.

delaLlave:1997:IMA

[dlL97] Rafael de la Llave. Invariant manifolds associated to nonresonant spectral subspaces. Journal of Statistical Physics, 87(1– 2):211–249, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181486.

Dunlop:1990:MWP

[DLM^{+}90] François Dunlop, Lahoussine Laanait, Alain Messager, Salvador Miracle-Sole, and Jean Ruiz. Multilayer wetting in partially symmetric q-state models. Journal of Statistical Physics, 59(5– 6):1383–1396, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01334756.

Dykman:1993:NSR

[DLM^{+}93a] M. I. Dykman, D. G. Luchinsky, R. Mannella, P. V. E. McClintock, N. D. Stein, and N. G. Stocks. Nonconventional stochastic resonance. Journal of Statistical Physics, 70(1–2):479–499, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01053983.

Dykman:1993:SRL

[DLM^{+}93b] M. I. Dykman, D. G. Luchinsky, R. Mannella, P. V. E. McClintock, N. D. Stein, and N. G. Stocks. Stochastic resonance: Linear response and giant nonlinearity. Journal of Statistical Physics, 70 (1–2):463–478, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053982.

deLeeuw:1990:HEC

[dLPP90] Simon W. de Leeuw, John W. Perram, and Henrik G. Petersen. Hamilton’s equations for constrained dynamical systems. Journal of Statistical Physics, 61(5–6):1203–1222, December 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014372.

Degond:1999:ESK

[DLPS99] Pierre Degond, José L. López, Frédéric Poupaud, and Christian Schmeiser. Existence of solutions of a kinetic equation modeling cometary flows. Journal of Statistical Physics, 96(1–2):361–376, July 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004584719071.

Derrida:1997:SPA

[DLS97] B. Derrida, J. L. Lebowitz, and E. R. Speer. Shock profiles for the asymmetric simple exclusion process in one dimension. Journal of Statistical Physics, 89(1–2):135–167, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02770758.

delosSantos:1999:CFM

[dlSG99] F. de los Santos and P. L. Garrido. Continuum field model of driven lattice gases. Journal of Statistical Physics, 96(1–2):303– 324, July 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004580618162.

Dekking:1990:SMP

[DM90] F. M. Dekking and R. W. J. Meester. On the structure of Mandelbrot’s percolation process and other random Cantor sets. Journal of Statistical Physics, 58(5–6):1109–1126, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01026566.

Dinaburg:1994:LTS

[DM94] Efim I. Dinaburg and Alexander E. Mazel. Layering transition in SOS model with external magnetic field. Journal of Statistical Physics, 74(3–4):533–563, February 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02188570.

Destainville:1997:CEC

[DMB97] N. Destainville, R. Mosseri, and F. Bailly. Configurational entropy of codimension-one tilings and directed membranes. Journal of Statistical Physics, 87(3–4):697–754, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02181243.

deMonvel-Berthier:1991:BMN

[dMBD91] Anne Boutet de Monvel-Berthier and Petre Dita. Brownian motion near an absorbing sphere. Journal of Statistical Physics, 62(3–4):729–736, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01017980.

deMendonca:1998:SCS

[dMdO98] J. Ricardo G. de Mendonça and Mário J. de Oliveira. Stationary coverage of a stochastic adsorption–desorption process with diffusional relaxation. Journal of Statistical Physics, 92(3–4):651–658, August 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023044822735.

deMagalhaes:1990:CCM

[dME90] A. C. N. de Magalhães and J. W. Essam. The n-component cubic model and flows: Subgraph break-collapse method. Journal of Statistical Physics, 58(5–6):1059–1082, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01026563.

deMagalhaes:1998:DVF

[dMM98] A. C. N. de Magalhães and A. M. Mariz. The discrete n-vector ferromagnet: Connection to a percolation with frustration features. Journal of Statistical Physics, 90(3–4):827–851, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023233205322.

deMatos:1991:FCW

[dMP91] J. M. G. Amaro de Matos and J. Fernando Perez. Fluctuations in the Curie–Weiss version of the random field Ising model. Journal of Statistical Physics, 62(3–4):587–608, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01017975.

Dorlas:1997:NSL

[DMP97] T. C. Dorlas, N. Macris, and J. V. Pulé. The nature of the spectrum for a Landau Hamiltonian with delta impurities. Journal of Statistical Physics, 87(3–4):847–875, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02181247.

deMonvel:1995:SMA

[dMPS95] A. Boutet de Monvel, L. Pastur, and M. Shcherbina. On the statistical mechanics approach in the random matrix theory: Integrated density of states. Journal of Statistical Physics, 79 (3–4):585–611, May 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02184872.

deMatos:1992:RIV

[dMPZ92] J. M. G. Amaro de Matos, A. E. Patrick, and V. A. Zagrebnov. Random infinite-volume Gibbs states for the Curie–Weiss random field Ising model. Journal of Statistical Physics, 66(1– 2):139–164, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01060064.

Debbasch:1997:ROU

[DMR97] F. Debbasch, K. Mallick, and J. P. Rivet. Relativistic Ornstein– Uhlenbeck process. Journal of Statistical Physics, 88(3–4): 945–966, August 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000015180.16261.53.

Dunlop:1992:PIW

[DMRR92] François Dunlop, Jacques Magnen, Vincent Rivasseau, and Philippe Roche. Pinning of an interface by a weak potential. Journal of Statistical Physics, 66(1–2):71–98, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01060060.

DeConinck:1990:MES

[DN90] Joël De Coninck and Charles M. Newman. The magnetizationenergy scaling limit in high dimension. Journal of Statistical Physics, 59(5–6):1451–1467, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01334759.

Daems:1994:PAH

[DN94a] D. Daems and G. Nicolis. Probabilistic approach to homoclinic chaos. Journal of Statistical Physics, 76(5–6):1287– 1305, September 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02187063.

Doyle:1994:ASA

[DN94b] M. M. Doyle and N. Sri Namachchivaya. Almost-sure asymptotic stability of a general four-dimensional system driven by real noise. Journal of Statistical Physics, 75(3–4):525–552, May 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02186871.

Desbois:1997:SMB

[DN97] Jean Desbois and Sergei Nechaev. Statistical mechanics of braided Markov chains: I. Analytic methods and numerical simulations. Journal of Statistical Physics, 88(1–2):201–229, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02508470.

deOliveira:1992:IMV

[dO92] M. J. de Oliveira. Isotropic majority-vote model on a square lattice. Journal of Statistical Physics, 66(1–2):273–281, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01060069.

deOliveira:1995:SRC

[dO95] César R. de Oliveira. Some remarks concerning stability for nonstationary quantum systems. Journal of Statistical Physics, 78(3–4):1055–1066, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02183701.

Driebe:1997:PSS

[DO97] Dean J. Driebe and Gonzalo E. Ordóñez. Polynomial shift states of a chaotic map. Journal of Statistical Physics, 89(5–6): 1087–1096, December 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02764224.

deOliveira:1995:FSS

[dOdOCS95] P. M. C. de Oliveira, S. M. Moss de Oliveira, C. E. Cordeiro, and D. Stauffer. Finite-size scaling for first-order transitions:

Potts model. Journal of Statistical Physics, 80(5–6):1433– 1442, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179879.

deOliveira:1995:SBC

[dOdOdSB95] S. Moss de Oliveira, P. M. C. de Oliveira, and F. C. de Sá Barreto. The spin- S Blume–Capel RG flow diagram. Journal of Statistical Physics, 78(5–6):1619–1627, March 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02180146.

Dolgopyat:1997:ECM

[Dol97] D. Dolgopyat. Entropy of coupled map lattices. Journal of Statistical Physics, 86(1–2):377–389, January 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02180211.

Domb:1992:BRS

[Dom92] Cyril Domb. Book review: The scientific letters and papers of James Clerk Maxwell. Journal of Statistical Physics, 67(3– 4):837–838, May 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049732.

Domb:1997:BRS

[Dom97] C. Domb. Book review: The scientific letters and papers of James Clerk Maxwell, Vol. II. Journal of Statistical Physics, 88(5–6): 1419–1422, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02732444.

Dong:1995:SAB

[Don95] Han Dong. Subcritical asymptotic behavior in the thermodynamic limit of reversible random polymerization processes. Journal of Statistical Physics, 80(1–2):389–404, July 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02178365.

Donnay:1999:NET

[Don99] Victor J. Donnay. Non-ergodicity of two particles interacting via a smooth potential. Journal of Statistical Physics, 96(5–6): 1021–1048, September 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004688200435.

deOliveira:1993:SCB

[dOP93] P. M. C. de Oliveira and T. J. P. Penna. Simulating the complex behavior of a leaky faucet. Journal of Statistical Physics, 73(3– 4):789–798, November 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01054351.

DeMasi:1993:MCS

[DOPT93] A. De Masi, E. Orlandi, E. Presutti, and L. Triolo. Motion by curvature by scaling nonlocal evolution equations. Journal of Statistical Physics, 73(3–4):543–570, November 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01054339.

Dorfman:1993:BRK

[Dor93] J. R. Dorfman. Book review: Kinetic theory and irreversible thermodynamics. Journal of Statistical Physics, 73(3–4):799– 801, November 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01054352.

Dorfman:1994:RBC

[Dor94] J. Robert Dorfman. Reponse to B. C. Eu. Journal of Statistical Physics, 76(5–6):1517–1518, September 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02187076.

Doussal:1992:DLR

[Dou92] Pierre Le Doussal. Diffusion in layered random flows, polymers, electrons in random potentials, and spin depolarization in random fields. Journal of Statistical Physics, 69(5–6):917–954, December 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01058756.

Douglas:1995:AAR

[Dou95] Jack F. Douglas. Aspects and applications of the random walk. Journal of Statistical Physics, 79(1–2):497–500, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179402.

Douglas:1997:RWR

[Dou97] Jack F. Douglas. Random walks and random environments, vol. 2, random environments. Journal of Statistical Physics, 87 (3–4):961–962, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181260.

Dowell:1991:OPO

[Dow91a] F. Dowell. Overview of partial orientational and positional ordering in concentrated systems. Journal of Statistical Physics, 62 (5–6):927–944, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01128169.

Dowell:1991:TPD

[Dow91b] F. Dowell. Theoretical predictions of diffusion from Brownian motion in superstrong polymers. Journal of Statistical Physics, 62(5–6):1059–1071, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01128177.

Delyon:1991:RES

[DP91] F. Delyon and J. Peyrière. Recurrence of the eigenstates of a Schrödinger operator with automatic potential. Journal of Statistical Physics, 64(1–2):363–368, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01057881.

Driebe:1992:ODA

[DP92] D. J. Driebe and T. Petrosky. Order and disorder in the approach to equilibrium of a classical gas. Journal of Statistical Physics, 67(1–2):369–394, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049040.

Dellago:1997:LIB

[DP97] Ch. Dellago and H. A. Posch. Lyapunov instability of the boundary-driven Chernov–Lebowitz model for stationary shear flow. Journal of Statistical Physics, 88(3–4):825–842, August 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOSS.0000015174.26700.7e.

Dobrushin:1990:ODH

[DPS90] R. L. Dobrushin, A. Pellegrinotti, and Yu. M. Suhov. Onedimensional harmonic lattice caricature of hydrodynamics: A higher correction. Journal of Statistical Physics, 61(1–2): 387–402, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01013971.

Duering:1991:CSD

[DR91] Edgardo Duering and H. Eduardo Roman. Corrections to scaling for diffusion exponents on three-dimensional percolation systems at criticality. Journal of Statistical Physics, 64(3–4):851–858, August 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048320.

Diehl:1993:DRU

[DR93] H. W. Diehl and U. Ritschel. Dynamical relaxation and universal short-time behavior of finite systems. Journal of Statistical Physics, 73(1–2):1–20, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01052748.

Debbasch:1998:DER

[DR98] F. Debbasch and J. P. Rivet. A diffusion equation from the relativistic Ornstein–Uhlenbeck process. Journal of Statistical Physics, 90(5–6):1179–1199, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023275210656.

Donev:1999:GSM

[DRbA99] Aleksandar Donev, Jeff Rockwell, and Daniel ben Avraham. Generalized von Smoluchowski model of reaction rates, with reacting particles and a mobile trap. Journal of Statistical Physics, 95(1–2):97–112, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004573310526.

deRueda:1999:CSD

[dRIB99] J. M. Reyes de Rueda, G. G. Izús, and C. H. Borzi. Critical slowing down on the dynamics of a bistable reaction–diffusion system in the neighborhood of its critical point. Journal of Statistical Physics, 97(3–4):803–809, November 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004627611784.

Drossel:1996:EBF

[Dro96] Barbara Drossel. Energy barriers for flux lines in three dimensions. Journal of Statistical Physics, 82(1–2):431–441, January 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02189238.

Devillard:1992:UCI

[DS92] P. Devillard and H. Spohn. Universality class of interface growth with reflection symmetry. Journal of Statistical Physics, 66(3–4): 1089–1099, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01055718.

Dodson:1999:SSS

[DS99] C. T. J. Dodson and W. W. Sampson. Spatial statistics of stochastic fiber networks. Journal of Statistical Physics, 96 (1–2):447–458, July 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004597121797.

daSilva:1991:DSM

[dSCT91] Luciano R. da Silva, Uriel M. S. Costa, and Constantino Tsallis. Do the simple and 2/3 majority models belong to the same universality class?: A Monte Carlo approach. Journal of Statistical Physics, 65(1–2):387–393, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01329867.

daSilva:1990:ADI

[dSL90] J. M. Nunes da Silva and E. J. S. Lage. Anomalous dynamics in the Ising chain. Journal of Statistical Physics, 58(1–2):115–124, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020287.

deSa:1992:GKL

[dSM92] Paula Gonzaga de Sá and Christian Maes. The Gacs– Kurdyumov–Levin automaton revisited. Journal of Statistical Physics, 67(3–4):507–522, May 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01049718.

Vieira:1990:TSF

[dSVG90] Maria C. de Sousa Vieira and Gemunu H. Gunaratne. The trajectory scaling function for period doubling bifurcations in flows. Journal of Statistical Physics, 58(5–6):1245–1256, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01026575.

Dimov:1993:RWD

[DT93] I. Dimov and O. Tonev. Random walk on distant mesh points Monte Carlo methods. Journal of Statistical Physics, 70(5–6): 1333–1342, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049435.

Duarte:1990:MSS

[Dua90] J. A. M. S. Duarte. A method for systematic site-to-bond conversion of directed graphs. Journal of Statistical Physics, 58(1– 2):389–393, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020301.

daVeiga:1998:LDB

[dVOS98] P. A. Faria da Veiga, M. O’Carroll, and R. Schor. On the largedistance behavior of correlations for a hierarchical n-component classical vector model in three dimensions. Journal of Statistical Physics, 92(1–2):47–55, July 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023091418328.

Deering:1991:MTM

[DW91] William D. Deering and Bruce J. West. A model of turbulent mixing in the a+b->0 reaction. Journal of Statistical Physics, 65(5–6):1247–1260, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01049610.

Ding:1994:PMC

[DZ94] Jiu Ding and Aihui Zhou. The projection method for computing multidimensional absolutely continuous invariant measures.

Journal of Statistical Physics, 77(3–4):899–908, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179467.

Eyink:1998:PTM

[EA98] Gregory L. Eyink and Francis J. Alexander. Predictive turbulence modeling by variational closure. Journal of Statistical Physics, 91(1–2):221–283, April 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023096206013.

Ernst:1995:ASC

[EB95a] M. H. Ernst and H. J. Bussemaker. Algebraic spatial correlations in lattice gas automata violating detailed balance. Journal of Statistical Physics, 81(1–2):515–536, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179991.

Erzan:1995:SPD

[EB95b] Ayse Erzan and Nihat Berker. Statistical physics days. Journal of Statistical Physics, 78(3–4):1189–1191, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183718.

Ebert:1996:PDQ

[Ebe96] Ute Ebert. Polymer diffusion in quenched disorder: A renormalization group approach. Journal of Statistical Physics, 82(1– 2):183–265, January 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02189230.

Ernst:1990:EHT

[ED90a] M. H. Ernst and J. W. Dufty. Erratum: Hydrodynamics and time correlation functions for cellular automata. Journal of Statistical Physics, 61(1–2):505, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01013979. See [ED90b].

Ernst:1990:HTC

[ED90b] M. H. Ernst and J. W. Dufty. Hydrodynamics and time correlation functions for cellular automata. Journal of Statistical Physics, 58(1–2):57–86, January 1990. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020285. See erra-

tum [ED90a].

Ernst:1992:TCA

[ED92a] M. H. Ernst and Shankar P. Das. Thermal cellular automata fluids. Journal of Statistical Physics, 66(1–2):465–483, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01060075.

Evans:1992:IBT

[ED92b] M. R. Evans and B. Derrida. Improved bounds for the transition temperature of directed polymers in a finite-dimensional random medium. Journal of Statistical Physics, 69(1–2):427–437, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01053800.

Elston:1996:NAS

[ED96] T. C. Elston and Charles R. Doering. Numerical and analytical studies of nonequilibrium fluctuation-induced transport processes. Journal of Statistical Physics, 83(3–4):359–383, May 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02183737.

Edis:1993:UCQ

[Edi93] Taner Edis. Unusual constraints in the quantum statistical mechanics of Josephson junction systems. Journal of Statistical Physics, 71(1–2):313–325, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01048102.

Edwards:1991:DCF

[Edw91] S. F. Edwards. Dynamics of complex flow. Journal of Statistical Physics, 62(5–6):889–897, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01128167.

Evans:1995:AEM

[EFGM95] M. R. Evans, D. P. Foster, C. Godrèche, and D. Mukamel. Asymmetric exclusion model with two species: Spontaneous symmetry breaking. Journal of Statistical Physics, 80(1–2):69–102, July 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02178354.

Enting:1990:ASL

[EG90] I. G. Enting and A. J. Guttmann. On the area of square lattice polygons. Journal of Statistical Physics, 58(3–4):475–484, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01112757.

Evans:1992:IPA

[EIK92] Nolan W. Evans, Reinhard Illner, and Hon C. Kwan. On information-processing abilities of chaotic dynamical systems. Journal of Statistical Physics, 66(1–2):549–561, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01060080.

Ebeling:1993:ANO

[EJ93a] Werner Ebeling and Martin Jenssen. Activation by nonlinear oscillations and solitonic excitations. Journal of Statistical Physics, 70(1–2):49–60, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053953.

Eisinger:1993:AAH

[EJ93b] S. Eisinger and J. Jäckle. Analytical approximations for the hierarchically constrained kinetic Ising chain. Journal of Statistical Physics, 73(3–4):643–670, November 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01054344.

Evangelou:1996:ELS

[EK96] S. N. Evangelou and D. E. Katsanos. Energy level statistics in disordered metals with an Anderson transition. Journal of Statistical Physics, 85(5–6):525–550, December 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02199355.

Essex:1999:MEP

[EK99] Christopher Essex and Dallas C. Kennedy. Minimum entropy production of neutrino radiation in the steady state.

Journal of Statistical Physics, 94(1–2):253–267, January 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004571531540.

Escande:1994:SCC

[EKLR94] Dominique Escande, Holger Kantz, Roberto Livi, and Stefano Ruffo. Self-consistent check of the validity of Gibbs calculus using dynamical variables. Journal of Statistical Physics, 76(1–2):605– 626, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02188677.

Esposito:1995:NSL

[ELM95] R. Esposito, J. L. Lebowitz, and R. Marra. The Navier–Stokes limit of stationary solutions of the nonlinear Boltzmann equation. Journal of Statistical Physics, 78(1–2):389–412, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02183355.

Eckern:1990:QLE

[ELMD^{+}90] U. Eckern, W. Lehr, A. Menzel-Dorwarth, F. Pelzer, and A. Schmid. The quasiclassical Langevin equation and its application to the decay of a metastable state and to quantum fluctuations. Journal of Statistical Physics, 59(3–4):885–934, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01025856.

Eloranta:1994:DDE

[Elo94] Kari Eloranta. The dynamics of defect ensembles in onedimensional cellular automata. Journal of Statistical Physics, 76(5–6):1377–1398, September 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02187067.

Eloranta:1999:DI

[Elo99] Kari Eloranta. Diamond ice. Journal of Statistical Physics, 96(5–6):1091–1109, September 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004644418182.

Eyink:1996:HFO

[ELS96] Gregory L. Eyink, Joel L. Lebowitz, and Herbert Spohn. Hydrodynamics and fluctuations outside of local equilibrium: Driven diffusive systems. Journal of Statistical Physics, 83(3–4):385– 472, May 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183738.

Esposito:1994:DIN

[EM94] R. Esposito and R. Marra. On the derivation of the incompressible Navier–Stokes equation for Hamiltonian particle systems. Journal of Statistical Physics, 74(5–6):981–1004, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02188213.

Elout:1995:FOP

[EM95] M. O. Elout and W. J. A. Maaskant. A first-order phase transition in a three-dimensional vertex model. Journal of Statistical Physics, 80(3–4):919–927, August 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02178560.

Elliott:1995:HMC

[EMHM95] Frank W. Elliott, Jr., Andrew J. Majda, David J. Horntrop, and Richard M. McLaughlin. Hierarchical Monte Carlo methods for fractal random fields. Journal of Statistical Physics, 81(3–4): 717–736, November 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179254.

Esposito:1998:SBE

[EML98] R. Esposito, R. Marra, and J. L. Lebowitz. Solutions to the Boltzmann equation in the Boussinesq regime. Journal of Statistical Physics, 90(5–6):1129–1178, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023223226585.

Eloranta:1992:KCA

[EN92a] Kari Eloranta and Esa Nummelin. The kink of cellular automaton rule 18 performs a random walk. Journal of Statistical Physics, 69(5–6):1131–1136, December 1992. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01058766.

Evans:1992:RCS

[EN92b] J. W. Evans and R. S. Nord. Random and cooperative sequential adsorption on infinite ladders and strips. Journal of Statistical Physics, 69(1–2):151–162, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053787.

Esipov:1993:NFR

[EN93] Sergei E. Esipov and T. J. Newman. New formulation of restricted growth processes. Journal of Statistical Physics, 70(3– 4):691–702, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053590.

Englman:1991:BRS

[Eng91] Robert Englman. Book review: Statistical models for the fracture of disordered media. Journal of Statistical Physics, 65(3–4): 825–826, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053759.

Englman:1992:IME

[Eng92] Robert Englman. Intermittency from maximum entropy distribution. Journal of Statistical Physics, 66(5–6):1383–1395, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01054427.

Ershov:1992:MLS

[EP92] Sergey V. Ershov and Alexey B. Potapov. Macrodynamics: Large-scale structures in turbulent media. Journal of Statistical Physics, 69(3–4):763–779, November 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01050433.

Esipov:1997:GPD

[EP97] Sergei E. Esipov and Thorsten Pöschel. The granular phase diagram. Journal of Statistical Physics, 86(5–6):1385–1395, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02183630.

Eckmann:1999:EPN

[EPRB99] Jean-Pierre Eckmann, Claude-Alain Pillet, and Luc Rey-Bellet. Entropy production in nonlinear, thermally driven Hamiltonian systems. Journal of Statistical Physics, 95(1–2):305–331, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1004537730090.

Ercai:1997:CSC

[Erc97] Chen Ercai. Chaos for the Sierpinski carpet. Journal of Statistical Physics, 88(3–4):979–984, August 1997. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000015182.90436.5b.

Ershov:1992:ATM

[Ers92] Sergey V. Ershov. Asymptotic theory of multidimensional chaos. Journal of Statistical Physics, 69(3–4):781–812, November 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01050434.

Ershov:1994:EFI

[Ers94] Sergey V. Ershov. Even the first iterate of a Markov operator is contracting in an L_{2} norm. Journal of Statistical Physics, 74(3– 4):783–813, February 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02188579.

Evans:1999:ESC

[ERS99] M. R. Evans, N. Rajewsky, and E. R. Speer. Exact solution of a cellular automaton for traffic. Journal of Statistical Physics, 95(1–2):45–96, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004521326456.

Eyink:1993:NTS

[ES93] G. L. Eyink and H. Spohn. Negative-temperature states and large-scale, long-lived vortices in two-dimensional turbulence. Journal of Statistical Physics, 70(3–4):833–886, February 1993.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01053597.

Ebert:1998:SMR

[ESB98] U. Ebert, L. Schäfer, and A. Baumgärtner. Segment motion in the reptation model of polymer dynamics. I. Analytical investigation. Journal of Statistical Physics, 90(5–6):1325–1373, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023239630220.

Erdos:1990:EPM

[ET90] L. Erdős and D. Q. Tuyen. Ergodic properties of the multidimensional Rayleigh gas with a semipermeable barrier. Journal of Statistical Physics, 59(5–6):1589–1602, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01334766.

Ermakov:1998:SAC

[ETW98] Alexei Ermakov, Bálint Tóth, and Wendelin Werner. On some annihilating and coalescing systems. Journal of Statistical Physics, 91(5–6):845–870, June 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023071714672.

Eu:1994:RRK

[Eu94] Byung Chan Eu. Response to review of kinetic theory and irreversible thermodynamics. Journal of Statistical Physics, 76(5–6): 1511–1515, September 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02187075.

Evangelou:1992:NSS

[Eva92] S. N. Evangelou. A numerical study of sparse random matrices. Journal of Statistical Physics, 69(1–2):361–383, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01053797.

Evertz:1993:VSS

[Eve93] Hans Gerd Evertz. Vectorized search for single clusters. Journal of Statistical Physics, 70(3–4):1075–1079, February 1993.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01053610.

Eyink:1990:DLT

[Eyi90] Gregory L. Eyink. Dissipation and large thermodynamic fluctuations. Journal of Statistical Physics, 61(3–4):533–572, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01027291.

Eyink:1995:BSM

[Eyi95a] Gregory L. Eyink. Besov spaces and the multifractal hypothesis. Journal of Statistical Physics, 78(1–2):353–375, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02183353.

Eyink:1995:LEF

[Eyi95b] Gregory L. Eyink. Local energy flux and the refined similarity hypothesis. Journal of Statistical Physics, 78(1–2):335–351, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183352.

Eyink:1996:TN

[Eyi96] Gregory L. Eyink. Turbulence noise. Journal of Statistical Physics, 83(5–6):955–1019, June 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179551.

Family:1991:DAK

[FA91] Fereydoon Family and Jacques G. Amar. Diffusion-annihilation and the kinetics of the Ising model in one dimension. Journal of Statistical Physics, 65(5–6):1235–1246, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01049609.

Fournou:1997:HNN

[FAKA97] E. Fournou, P. Argyrakis, B. Kargas, and P. A. Anninos. Hybrid neural nets with Poisson and Gaussian connectivities. Journal of Statistical Physics, 89(3–4):847–867, November 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02765547.

Falconer:1992:WTO

[Fal92] K. J. Falconer. Wavelet transforms and order-two densities of fractals. Journal of Statistical Physics, 67(3–4):781–793, May 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01049726.

Family:1992:BRF

[Fam92] Fereydoon Family. Book review: Fractal growth phenomena. Journal of Statistical Physics, 66(1–2):683–686, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01060090.

Family:1996:FCS

[Fam96] Fereydoon Family. Fractal concepts in surface growth. Journal of Statistical Physics, 83(5–6):1255–1259, June 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179563.

Fogedby:1992:FLF

[FBJ92] H. C. Fogedby, T. Bohr, and H. J. Jensen. Fluctuations in a lévy flight gas. Journal of Statistical Physics, 66(1–2):583–593, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01060082.

Frigessi:1994:DPT

[FdH94] Arnoldo Frigessi and Frank den Hollander. A dynamical phase transition in a caricature of a spin glass. Journal of Statistical Physics, 75(3–4):585–625, May 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02186873.

Falcolini:1992:NCD

[FdlL92a] Corrado Falcolini and Rafael de la Llave. Numerical calculation of domains of analyticity for perturbation theories in the presence of small divisors. Journal of Statistical Physics, 67(3–4):645–666, May 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049723.

Falcolini:1992:RPJ

[FdlL92b] Corrado Falcolini and Rafael de la Llave. A rigorous partial justification of Greene’s criterion. Journal of Statistical Physics, 67(3–4):609–643, May 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049722.

Firpo:1998:KLB

[FE98] Marie-Christine Firpo and Yves Elskens. Kinetic limit of nbody description of wave–particle self-consistent interaction. Journal of Statistical Physics, 93(1–2):193–209, October 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1023/B%3AJOSS.0000026732.51044.87.pdf.

Felderhof:1998:DCD

[Fel98] B. U. Felderhof. Dynamical conductivity of the dilute Lorentz gas with spherically symmetric scatterers. Journal of Statistical Physics, 93(1–2):307–329, October 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%
3AJOSS.0000026736.41183.8b.pdf.

Fernandez:1994:MSP

[Fer94] Ariel Fernández. A measure on the space of polymer folding pathways: Preliminaries for a new scheme of statistical inference. Journal of Statistical Physics, 77(5–6):1079–1085, December 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183153.

Fernandez:1996:ESS

[Fer96] Bastien Fernandez. Existence and stability of steady fronts in bistable coupled map lattices. Journal of Statistical Physics, 82(3–4):931–950, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179796.

Fernandez:1998:LSL

[Fer98] Ariel Fernández. The Lagrangian structure of long-time torsional dynamics leading to RNA folding. Journal of Statistical Physics, 92(1–2):237–267, July 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023051804215.

Forland:1995:AAE

[FF95] Katrine Seip Førland and Tormod Førland. An alternative approach to electrochemistry. Journal of Statistical Physics, 78(1– 2):513–529, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183362.

Flekkoy:1992:LGS

[FFJ92] E. G. Flekkøy, J. Feder, and T. Jøssang. Lattice gas simulations of osmosis. Journal of Statistical Physics, 68(3–4):515–532, August 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01341760.

Ferrari:1994:IMT

[FFK94] P. A. Ferrari, L. R. G. Fontes, and Y. Kohayakawa. Invariant measures for a two-species asymmetric process. Journal of Statistical Physics, 76(5–6):1153–1177, September 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02187059.

Foster:1994:FSE

[FG94] D. P. Foster and C. Godrèche. Finite-size effects for phase segregation in a two-dimensional asymmetric exclusion model with two species. Journal of Statistical Physics, 76(5–6):1129– 1151, September 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02187058.

Fried:1999:CSS

[FG99] Eliot Fried and Morton E. Gurtin. Coherent solid–state phase transitions with atomic diffusion: A thermomechanical treatment. Journal of Statistical Physics, 95(5–6):1361–1427, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1004535408168.

Fioretti:1993:ESR

[FGMA93] A. Fioretti, L. Guidoni, R. Mannella, and E. Arimondo. Evidence of stochastic resonance in a laser with saturable absorber:

Experiment and theory. Journal of Statistical Physics, 70(1– 2):403–412, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053976.

Figotin:1992:LPR

[Fig92a] Alexander Figotin. The localization properties of a random steady flow on a lattice. Journal of Statistical Physics, 66(5– 6):1599–1612, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01054436.

Figotin:1992:MNM

[Fig92b] Alexander Figotin. Model of a nonhomogeneous medium conducting light. Journal of Statistical Physics, 69(5–6):969– 993, December 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01058758.

Figotin:1993:EGS

[Fig93] A. Figotin. Existence of gaps in the spectrum of periodic dielectric structures on a lattice. Journal of Statistical Physics, 73(3– 4):571–585, November 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01054340.

Figotin:1994:PPP

[Fig94] Alexander Figotin. Photonic pseudogaps for periodic dielectric structures. Journal of Statistical Physics, 74(1–2):433–446, January 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186819.

Filippov:1994:LSS

[Fil94] A. E. Filippov. Large-scale structure of fluctuating order parameter field. Journal of Statistical Physics, 75(1–2):241–252, April 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186288.

Finjord:1992:SFM

[Fin92] Jan Finjord. Structure functions in a model of turbulent energy dissipation. Journal of Statistical Physics, 68(5–6):749– 760, September 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048874.

Fisher:1994:SCC

[Fis94] Michael E. Fisher. The story of Coulombic criticality. Journal of Statistical Physics, 75(1–2):1–36, April 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02186278.

Feng:1996:CSE

[FIS96] Shui Feng, Ian Iscoe, and Timo Seppäläinen. A class of stochastic evolutions that scale to the porous medium equation. Journal of Statistical Physics, 85(3–4):513–517, November 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02174218.

Forrester:1992:ADB

[FJ92] P. J. Forrester and B. Jancovici. On the average distance between particles in the two-dimensional two-component plasma. Journal of Statistical Physics, 69(1–2):163–178, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01053788.

Fiig:1993:DDL

[FJ93] Thomas Fiig and Henrik Jeldtoft Jensen. Diffusive description of lattice gas models. Journal of Statistical Physics, 71(3–4):653– 682, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01058441.

Forrester:1996:TDT

[FJ96] P. J. Forrester and B. Jancovici. The two-dimensional twocomponent plasma plus background on a sphere: Exact results. Journal of Statistical Physics, 84(3–4):337–357, August 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179646.

Fisher:1995:AIP

[FjLL95] Michael E. Fisher, Xiao jun Li, and Yan Levin. On the absence of intermediate phases in the two-dimensional Coulomb gas. Journal of Statistical Physics, 79(1–2):1–11, April 1995.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF02179380. See erratum [FLL95].

Forrester:1992:TDC

[FJM92] P. J. Forrester, B. Jancovici, and J. Madore. The twodimensional Coulomb gas on a sphere: Exact results. Journal of Statistical Physics, 69(1–2):179–192, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01053789.

Forrester:1996:USC

[FJT96] P. J. Forrester, B. Jancovici, and G. Téllez. Universality in some classical Coulomb systems of restricted dimension. Journal of Statistical Physics, 84(3–4):359–378, August 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179647.

Figotin:1994:LEA

[FK94a] Alexander Figotin and Abel Klein. Localization of electromagnetic and acoustic waves in random media. Lattice models. Journal of Statistical Physics, 76(3–4):985–1003, August 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02188695.

Figotin:1994:LPG

[FK94b] Alexander Figotin and Abel Klein. Localization phenomenon in gaps of the spectrum of random lattice operators. Journal of Statistical Physics, 75(5–6):997–1021, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02186755.

Figotin:1994:BGS

[FK94c] Alexander Figotin and Peter Kuchment. Band-gap structure of the spectrum of periodic Maxwell operators. Journal of Statistical Physics, 74(1–2):447–455, January 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02186820.

Figotin:1997:LCW

[FK97] Alexander Figotin and Abel Klein. Localized classical waves created by defects. Journal of Statistical Physics, 86(1–2):165–177, January 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02180202.

Feldman:1999:TZL

[FKST99] Joel Feldman, Horst Knörrer, Manfred Salmhofer, and Eugene Trubowitz. The temperature zero limit. Journal of Statistical Physics, 94(1–2):113–157, January 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004523616519.

Fannes:1991:CFQ

[FKV91] M. Fannes, A. Kossakowski, and A. Verbeure. Critical fluctuations for quantum mean-field models. Journal of Statistical Physics, 65(3–4):801–811, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053756.

Fife:1994:MCG

[FL94] Paul C. Fife and Andrew A. Lacey. Motion by curvature in generalized Cahn–Allen models. Journal of Statistical Physics, 77 (1–2):173–181, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186837.

Fidaleo:1999:EPQ

[FL99] Francesco Fidaleo and Carlangelo Liverani. Ergodic properties for a quantum nonlinear dynamics. Journal of Statistical Physics, 97(5–6):957–1009, December 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004662014026.

Ferrenberg:1991:SSE

[FLB91] Alan M. Ferrenberg, D. P. Landau, and K. Binder. Statistical and systematic errors in Monte Carlo sampling. Journal of Statistical Physics, 63(5–6):867–882, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01029988.

Fisher:1995:EAI

[FLL95] Michael E. Fisher, Xiaojun Li, and Yan Levin. Errata: On the absence of intermediate phases in the two-dimensional Coulomb gas. Journal of Statistical Physics, 81(3–4):865, November 1995.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF02179265. See [FjLL95].

Fendley:1995:SPB

[FLS95] P. Fendley, F. Lesage, and H. Saleur. Solving 1D plasmas and 2D boundary problems using Jack polynomials and functional relations. Journal of Statistical Physics, 79(5–6):799–819, June 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02181204.

Fendley:1996:UFK

[FLS96] P. Fendley, F. Lesage, and H. Saleur. A unified framework for the Kondo problem and for an impurity in a Luttinger liquid. Journal of Statistical Physics, 85(1–2):211–249, October 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02175563.

Forrester:1991:BTM

[FM91] P. J. Forrester and T. M. Morrow. Block Toeplitz matrices and the two-dimensional Coulomb gas near a wall. Journal of Statistical Physics, 63(1–2):1–23, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01026589.

Fernandez:1993:DOD

[FM93a] Julio F. Fernández and J. Marro. Diffusion in a one-dimensional gas of hard point particles. Journal of Statistical Physics, 71 (1–2):225–233, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048096.

Fronzoni:1993:SRP

[FM93b] L. Fronzoni and R. Mannella. Stochastic resonance in periodic potentials. Journal of Statistical Physics, 70(1–2):501–512, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01053984.

Faris:1999:QCM

[FM99] William G. Faris and Robert A. Minlos. A quantum crystal with multidimensional anharmonic oscillators. Journal of Statistical Physics, 94(3–4):365–387, February 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004588002407.

Fukugita:1990:FSS

[FMOU90] M. Fukugita, H. Mino, M. Okawa, and A. Ukawa. Finite-size scaling of the three-state Potts model on a simple cubic lattice. Journal of Statistical Physics, 59(5–6):1397–1429, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01334757.

Ferrari:1992:LBM

[FMP92] Pablo A. Ferrari, Servet Martínez, and Pierre Picco. A lower bound for the memory capacity in the Potts–Hopfield model. Journal of Statistical Physics, 66(5–6):1643–1652, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01054440.

Franz:1999:RGS

[FMPP99] Silvio Franz, Marc Mézard, Giorgio Parisi, and Luca Peliti. The response of glassy systems to random perturbations: A bridge between equilibrium and off-equilibrium. Journal of Statistical Physics, 97(3–4):459–488, November 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004602906332.

Feix:1994:SPI

[FMR94] M. R. Feix, A. Muriel, and J. L. Rouet. Statistical properties of an iterated arithmetic mapping. Journal of Statistical Physics, 76(1–2):725–741, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02188683.

Ferrari:1997:PTA

[FMS97] P. A. Ferrari, S. Martinez, and J. San Martín. Phase transition for absorbed Brownian motion with drift. Journal of Statistical Physics, 86(1–2):213–231, January 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02180205.

Fontes:1995:PUC

[FN95] L. R. G. Fontes and E. Jordão Neves. Phase uniqueness and correlation length in diluted-field Ising models. Journal of Statistical Physics, 80(5–6):1327–1339, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179873.

Forrester:1997:CPD

[FN97] P. J. Forrester and T. Nagao. Correlations for parameterdependent random matrices. Journal of Statistical Physics, 89 (1–2):69–110, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02770755.

Fannes:1992:GSV

[FNW92] M. Fannes, B. Nachtergaele, and R. F. Werner. Ground states of VBS models on Cayley trees. Journal of Statistical Physics, 66(3– 4):939–973, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01055710.

Foda:1997:IOO

[Fod97] Omar Foda. Introduction to octionion and other non-associative algebras in physics. Journal of Statistical Physics, 87(3–4):965– 966, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02181262.

Fogedby:1992:PSA

[Fog92] Hans C. Fogedby. On the phase space approach to complexity. Journal of Statistical Physics, 69(1–2):411–425, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01053799.

Forrester:1990:ERT

[For90a] P. J. Forrester. Exact results for the two-dimensional, twocomponent plasma at Γ = 2 in doubly periodic boundary conditions. Journal of Statistical Physics, 61(5–6):1141–1160, December 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014369.

Forrester:1990:YLT

[For90b] P. J. Forrester. Yang–Lee theory and the conductor-insulator transition in asymmetric log-potential lattice gases. Journal of Statistical Physics, 60(1–2):203–220, July 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01013674.

Forrester:1991:FSC

[For91] P. J. Forrester. Finite-size corrections to the free energy of Coulomb systems with a periodic boundary condition. Journal of Statistical Physics, 63(3–4):491–504, May 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01029197.

Forrester:1992:STT

[For92] P. J. Forrester. Surface tension for the two-component plasma at Γ = 2 near an interface. Journal of Statistical Physics, 67 (3–4):433–448, May 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049715.

Forrester:1993:REC

[For93] P. J. Forrester. Recurrence equations for the computation of correlations in the 1/r^{2} quantum many-body system. Journal of Statistical Physics, 72(1–2):39–50, July 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048039.

Forrester:1994:PEC

[For94] P. J. Forrester. Properties of an exact crystalline many-body ground state. Journal of Statistical Physics, 76(1–2):331–346, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02188665.

Foster:1993:LCP

[Fos93] Damien P. Foster. Location of the collapsed phase for twodimensional, directed, interacting polymers. Journal of Statistical Physics, 70(3–4):1029–1034, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01053605.

Freire:1994:DCF

[FOS94] F. Freire, Denjoe O’Connor, and C. R. Stephens. Dimensional crossover and finite-size scaling below T_{c}. Journal of Statistical Physics, 74(1–2):219–238, January 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02186813.

Fesjian:1990:CMO

[FP90] S. Fesjian and J. K. Percus. Collective modes in a onedimensional nonuniform fluid model. Journal of Statistical Physics, 60(5–6):659–668, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01025987.

Fan:1992:RSA

[FP92] Y. Fan and J. K. Percus. Random sequential adsorption on a ladder. Journal of Statistical Physics, 66(1–2):263–271, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01060068.

Fannjiang:1997:CED

[FP97] Albert Fannjiang and George Papanicolaou. Convectionenhanced diffusion for random flows. Journal of Statistical Physics, 88(5–6):1033–1076, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02732425.

Farago:1998:LTB

[FP98] Jean Farago and Michel Peyrard. Long-time bath correlations in the Pollak-Grabert–Hänggi theory. Journal of Statistical Physics, 91(3–4):733–757, May 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023037913222.

Fratzl:1999:MPS

[FPL99] Peter Fratzl, Oliver Penrose, and Joel L. Lebowitz. Modeling of phase separation in alloys with coherent elastic misfit. Journal of Statistical Physics, 95(5–6):1429–1503, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004587425006.

Fox:1990:TNS

[FR90] Ronald F. Fox and Rajarshi Roy. Tests of numerical simulation algorithms for the Kubo oscillator. Journal of Statistical Physics, 58(1–2):395–396, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020302.

Franz:1995:FDR

[FR95] Silvio Franz and Heiko Rieger. Fluctuation-dissipation ratio in three-dimensional spin glasses. Journal of Statistical Physics, 79 (3–4):749–758, May 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02184881.

Franz:1996:GMF

[FR96] Silvio Franz and Felix Ritort. Glassy mean-field dynamics of the backgammon model. Journal of Statistical Physics, 85(1– 2):131–150, October 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02175558.

Fernandez:1997:PFB

[FR97] Bastien Fernandez and Laurent Raymond. Propagating fronts in a bistable coupled map lattice. Journal of Statistical Physics, 86 (1–2):337–350, January 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02180209.

Fort:1999:CBI

[FR99] J. Fort and P. Roura. A comparison between information– theoretic and phenomenological descriptions of nonequilibrium radiation. Journal of Statistical Physics, 97(5–6):941–955, December 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004610029956.

Fratzl:1994:SMH

[Fra94] Peter Fratzl. Statistical model of the habit and arrangement of mineral crystals in the collagen of bone. Journal of Statistical Physics, 77(1–2):125–143, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02186835.

Friedman:1995:SHD

[FRHP95] Harold L. Friedman, Fernando O. Raineri, Fumio Hirata, and Baw-Ching Perng. Surrogate Hamiltonian description of solvation dynamics. Site number density and polarization charge density formulations. Journal of Statistical Physics, 78(1–2): 239–266, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183347.

Frigerio:1990:SAQ

[Fri90] Alberto Frigerio. Simulated annealing and quantum detailed balance. Journal of Statistical Physics, 58(1–2):325–354, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020296.

Fritz:1994:SSH

[Fri94] J. Fritz. Stationary states and hydrodynamics of FHP cellular automata. Journal of Statistical Physics, 77(1–2):53–76, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186832.

Frezzotti:1993:NAS

[FS93] Aldo Frezzotti and Carlo Sgarra. Numerical analysis of a shockwave solution of the Enskog equation obtained via a Monte Carlo method. Journal of Statistical Physics, 73(1–2):193–207, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01052757.

Ferreira:1999:APM

[FS99] Sabino José Ferreira and Alan D. Sokal. Antiferromagnetic Potts models on the square lattice: A high-precision Monte Carlo study. Journal of Statistical Physics, 96(3–4):461–530, August 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1004599121565.

Fuller:1991:FIA

[FSB91] Gerald G. Fuller, Kaye Smith, and Wesley R. Burghardt. Fieldinduced anisotropy in concentrated systems of rigid particles and macromolecules. Journal of Statistical Physics, 62(5–6):

1025–1039, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01128175.

Feldman:1996:PTA

[FST96] Joel Feldman, Manfred Salmhofer, and Eugene Trubowitz. Perturbation theory around nonnested Fermi surfaces. I. Keeping the Fermi surface fixed. Journal of Statistical Physics, 84(5–6): 1209–1336, September 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02174132.

Frohlich:1997:CQH

[FST97] Jürg Fröhlich, Urban M. Studer, and Emmanuel Thiran. A classification of quantum Hall fluids. Journal of Statistical Physics, 86(3–4):821–897, February 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02199122.

Forrest:1990:HSP

[FT90a] Bruce M. Forrest and Lei-Han Tang. Hypercube stacking: A Potts-spin model for surface growth. Journal of Statistical Physics, 60(1–2):181–202, July 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01013673.

Fourcade:1990:UMPa

[FT90b] B. Fourcade and A.-M. S. Tremblay. Universal multifractal properties of circle maps from the point of view of critical phenomena I. Phenomenology. Journal of Statistical Physics, 61(3–4): 607–637, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01027294.

Fourcade:1990:UMPb

[FT90c] B. Fourcade and A.-M. S. Tremblay. Universal multifractal properties of circle maps from the point of view of critical phenomena II. Analytical results. Journal of Statistical Physics, 61(3–4): 639–665, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01027295.

Forrest:1991:HSP

[FT91] Bruce M. Forrest and Lei-Han Tang. Hypercube stacking: A Potts-spin model for surface growth. Journal of Statistical Physics, 65(1–2):413, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01329870.

Forrest:1993:CFS

[FT93] Bruce M. Forrest and Raul Toral. Crossover and finite-size effects in the (1 + 1)-dimensional Kardar–Parisi–Zhang equation. Journal of Statistical Physics, 70(3–4):703–720, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01053591.

Forrest:1994:PDF

[FT94a] Bruce M. Forrest and Raúl Toral. The phase diagram of the Flory–Huggins–de Gennes model of a binary polymer blend. Journal of Statistical Physics, 77(1–2):473–489, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02186853.

Frohlich:1994:IQF

[FT94b] Jürg Fröhlich and Emmanuel Thiran. Integral quadratic forms, Kac–Moody algebras, and fractional quantum Hall effect. An ADE-O classification. Journal of Statistical Physics, 76(1– 2):209–283, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02188661.

Floriani:1996:DAA

[FTGW96] Elena Floriani, György Trefán, Paolo Grigolini, and Bruce J. West. A dynamical approach to anomalous conductivity. Journal of Statistical Physics, 84(5–6):1043–1066, September 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02174127.

Fujimoto:1990:HHMa

[Fuj90a] Masafumi Fujimoto. Hard-hexagon model: Anisotropy of correlation length and interfacial tension. Journal of Statistical Physics, 59(5–6):1355–1381, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01334755.

Fujimoto:1990:HHMb

[Fuj90b] Masafumi Fujimoto. Hard-hexagon model: Calculation of anisotropic interfacial tension from asymptotic degeneracy of largest eigenvalues of row-row transfer matrix. Journal of Statistical Physics, 61(5–6):1295–1304, December 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01014376.

Fujimoto:1991:HHM

[Fuj91] Masafumi Fujimoto. Hard-hexagon model: Anisotropy of correlation length and interfacial tension. Journal of Statistical Physics, 62(1–2):503–504, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020890.

Fujimoto:1992:EVM

[Fuj92] Masafumi Fujimoto. Eight-vertex model: Anisotropic interfacial tension and equilibrium crystal shape. Journal of Statistical Physics, 67(1–2):123–154, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01049029.

Fujimoto:1996:SVM

[Fuj96] Masafumi Fujimoto. Six-vertex model with rotated boundary conditions. Journal of Statistical Physics, 82(5–6):1519–1539, March 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183394.

Fujimoto:1998:AVM

[Fuj98] Masafumi Fujimoto. Auxiliary vertices method for Kagomélattice eight-vertex model. Journal of Statistical Physics, 90(1– 2):363–388, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023272206089.

Felder:1997:EQG

[FV97] Giovanni Felder and Alexander Varchenko. Elliptic quantum groups and Ruijsenaars models. Journal of Statistical Physics, 89(5–6):963–980, December 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02764216.

Foster:1992:ZTP

[FVY92] Damien P. Foster, Carlo Vanderzande, and Julia Yeomans. Zerotemperature properties of randomly self-interacting polymers. Journal of Statistical Physics, 69(3–4):857–868, November 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01050437.

Frohlich:1991:PTD

[FZ91] J. Fröhlich and B. Zegarliński. The phase transition in the discrete Gaussian chain with 1/r^{2} interaction energy. Journal of Statistical Physics, 63(3–4):455–485, May 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01029195.

Feinberg:1997:RRO

[FZ97] Joshua Feinberg and A. Zee. Renormalizing rectangles and other topics in random matrix theory. Journal of Statistical Physics, 87(3–4):473–504, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181233.

Gofman:1993:SMC

[GAA^{+}93] Misha Gofman, Joan Adler, Amnon Aharony, A. B. Harris, and Dietrich Stauffer. Series and Monte Carlo study of highdimensional Ising models. Journal of Statistical Physics, 71(5– 6):1221–1230, June 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049970.

Garcia:1997:PMA

[GAA97] Alejandro L. Garcia, Francis J. Alexander, and Berni J. Alder. A particle method with adjustable transport properties-the generalized consistent Boltzmann algorithm. Journal of Statistical Physics, 89(1–2):403–409, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02770772.

Gacs:1990:TRI

[Gác90] Peter Gács. A toom rule that increases the thickness of sets. Journal of Statistical Physics, 59(1–2):171–193, April 1990.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01015567.

Galam:1990:SPM

[Gal90] Serge Galam. Social paradoxes of majority rule voting and renormalization group. Journal of Statistical Physics, 61(3–4): 943–951, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01027314.

Gallavotti:1995:EEI

[Gal95] Giovanni Gallavotti. Ergodicity, ensembles, irreversibility in Boltzmann and beyond. Journal of Statistical Physics, 78(5– 6):1571–1589, March 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02180143.

Gallavotti:1996:CHO

[Gal96] Giovanni Gallavotti. Chaotic hypothesis: Onsager reciprocity and fluctuation-dissipation theorem. Journal of Statistical Physics, 84(5–6):899–925, September 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02174123.

Gallavotti:1997:CPS

[Gal97] Giovanni Gallavotti. Chaotic principle: Some applications to developed turbulence. Journal of Statistical Physics, 86(5–6): 907–934, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183608.

Gandolfi:1991:BRP

[Gan91] Alberto Gandolfi. Book review: Percolation. Journal of Statistical Physics, 64(3–4):893–895, August 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048322.

Garrod:1991:SMT

[Gar91] Claude Garrod. A stochastic model of three-dimensional crystal growth. Journal of Statistical Physics, 63(5–6):987–1003, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01029994.

Garanin:1994:ELD

[Gar94] D. A. Garanin. The 1/D expansion for low-dimensional classical magnets. Journal of Statistical Physics, 74(1–2):275–311, January 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186815.

Garanin:1996:ECM

[Gar96] D. A. Garanin. The 1/D expansion for classical magnets: Lowdimensional models with magnetic field. Journal of Statistical Physics, 83(5–6):907–931, June 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179549.

Garrido:1997:KSE

[Gar97] P. L. Garrido. Kolmogorov–Sinai entropy, Lyapunov exponents, and mean free time in billiard systems. Journal of Statistical Physics, 88(3–4):807–824, August 1997. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000015173.74708.2a.

Garnier:1998:ABQ

[Gar98] Josselin Garnier. Asymptotic behavior of the quantum harmonic oscillator driven by a random time-dependent electric field. Journal of Statistical Physics, 93(1–2):211–241, October 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1023/B%3AJOSS.0000026733.10660.76.pdf.

Gaspard:1992:DEC

[Gas92] Pierre Gaspard. Diffusion, effusion, and chaotic scattering: An exactly solvable Liouvillian dynamics. Journal of Statistical Physics, 68(5–6):673–747, September 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048873.

Gaspard:1997:EPO

[Gas97] Pierre Gaspard. Entropy production in open volume-preserving systems. Journal of Statistical Physics, 88(5–6):1215–1240, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/BF02732432.

Gayrard:1992:TLS

[Gay92] V. Gayrard. Thermodynamic limit of the q-state Potts–Hopfield model with infinitely many patterns. Journal of Statistical Physics, 68(5–6):977–1011, September 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048882.

Guttmann:1990:CEL

[GB90] A. J. Guttmann and R. J. Bursill. Critical exponent for the loop erased self-avoiding walk by Monte Carlo methods. Journal of Statistical Physics, 59(1–2):1–9, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01015560.

Grimm:1994:NIQ

[GB94] Uwe Grimm and Michael Baake. Nonperiodic Ising quantum chains and conformal invariance. Journal of Statistical Physics, 74(5–6):1233–1245, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02188226.

Grosfils:1999:POL

[GBCB99] Patrick Grosfils, Jean Pierre Boon, E. G. D. Cohen, and L. A. Bunimovich. Propagation and organization in lattice random media. Journal of Statistical Physics, 97(3–4):575–608, November 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004611208149.

Gandjbakhche:1992:SRA

[GBN92] A. H. Gandjbakhche, R. F. Bonner, and R. Nossal. Scaling relationships for anisotropic random walks. Journal of Statistical Physics, 69(1–2):35–53, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01053781.

Gora:1991:NIM

[GBP91] P. Góra, A. Boyarsky, and H. Proppe. On the number of invariant measures for higher-dimensional chaotic transformations. Journal of Statistical Physics, 62(3–4):709–728, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017979.

Gaunt:1990:BPW

[GC90] D. S. Gaunt and S. A. Colby. Branched polymers in a wedge geometry in three dimensions. Journal of Statistical Physics, 58(3– 4):539–551, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01112761.

Gwa:1992:NHE

[GC92] Leh-Hun Gwa and E. G. D. Cohen. Note on the hydrodynamic eigenmodes of Couette flow. Journal of Statistical Physics, 69(5–6):1123–1130, December 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01058765.

Gallavotti:1995:DES

[GC95] G. Gallavotti and E. G. D. Cohen. Dynamical ensembles in stationary states. Journal of Statistical Physics, 80(5–6):931– 970, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179860.

Geldart:1993:LCD

[GD93] D. J. W. Geldart and K. De’Bell. Logarithmic corrections for dilute uniaxial ferromagnets at the critical dimension. Journal of Statistical Physics, 73(1–2):409–414, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01052769.

Ginzbourg:1996:LSO

[Gd96] I. Ginzbourg and D. d’Humières. Local second-order boundary methods for lattice Boltzmann models. Journal of Statistical Physics, 84(5–6):927–971, September 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02174124.

Gilbert:1999:EPO

[GD99] T. Gilbert and J. R. Dorfman. Entropy production: From open volume-preserving to dissipative systems. Journal of Statistical Physics, 96(1–2):225–269, July 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004576517254.

Greven:1991:PGRb

[GdH91a] A. Greven and F. den Hollander. Population growth in random media. Journal of Statistical Physics, 65(5–6):1147–1154, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049603.

Greven:1991:PGRa

[GdH91b] A. Greven and F. den Hollander. Population growth in random media. I. Variational formula and phase diagram. Journal of Statistical Physics, 65(5–6):1123–1146, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01049602.

Grossmann:1993:CTP

[GDJH93] Frank Grossmann, Thomas Dittrich, Peter Jung, and Peter Hänggi. Coherent transport in a periodically driven bistable system. Journal of Statistical Physics, 70(1–2):229–245, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01053965.

Ge:1991:SPM

[Ge91] Yuzhen Ge. Scaling properties of the measure of constant topological entropy. Journal of Statistical Physics, 63(1–2):131–140, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01026596.

Georgii:1995:EEC

[Geo95] Hans-Otto Georgii. The equivalence of ensembles for classical systems of particles. Journal of Statistical Physics, 80(5–6): 1341–1378, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179874.

Gerasimov:1990:RWR

[Ger90] A. L. Gerasimov. The role of weak resonances in ac-driven Brownian motion and rate processes. Journal of Statistical Physics, 60 (3–4):485–500, August 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01314932.

Gerasimov:1993:DTE

[Ger93a] A. Gerasimov. Diffusive transport enhancement and escape processes in frictionless nonlinear oscillators with noise. Journal of Statistical Physics, 72(3–4):555–570, August 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01048024.

Gerasimov:1993:RRT

[Ger93b] A. Gerasimov. Resonant response of a thermalized ensemble of nonlinear oscillators. Journal of Statistical Physics, 70(3–4): 939–948, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053601.

Germinet:1999:DLI

[Ger99] François Germinet. Dynamical localization II with an application to the almost Mathieu operator. Journal of Statistical Physics, 95(1–2):273–286, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004533629182.

Garrido:1994:BCF

[GG94a] Pedro L. Garrido and Giovanni Gallavotti. Billiards correlation functions. Journal of Statistical Physics, 76(1–2):549–585, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02188675.

Giarritta:1994:SGF

[GG94b] S. Prestipino Giarritta and P. V. Giaquinta. Statistical geometry of four calottes on a sphere. Journal of Statistical Physics, 75 (5–6):1093–1118, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186758.

Graham:1991:DCS

[GGD91] A. L. Graham, M. Gottlieb, and F. Dowell. Dynamics of concentrated systems. Journal of Statistical Physics, 62(5–6):887–888, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01128166.

Golden:1990:DBE

[GGL90] K. Golden, S. Goldstein, and J. L. Lebowitz. Discontinuous behavior of effective transport coefficients in quasiperiodic media. Journal of Statistical Physics, 58(3–4):669–684, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01112770.

Giaquinta:1991:HDP

[GGM91] P. V. Giaquinta, G. Giunta, and G. Malescio. High-density properties of hard spheres within a modified Percus–Yevick theory: The role of thermodynamic consistency. Journal of Statistical Physics, 63(1–2):141–165, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01026597.

Gyorgyi:1992:RGS

[GGP92] Géza Györgyi, Robert Graham, and R. E. Prange. Renormalization group study of quantum fluctuations near classical critical points of Hamiltonian systems. Journal of Statistical Physics, 68 (1–2):175–188, July 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048841.

Gottlob:1994:XMT

[GH94] Aloysius P. Gottlob and Martin Hasenbusch. The XY model and the three-state antiferromagnetic Potts model in three dimensions: Critical properties from fluctuating boundary conditions. Journal of Statistical Physics, 77(3–4):919–930, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179470.

Guerin:1997:EDC

[GH97] Charles-Antoine Guerin and Matthias Holschneider. On equivalent definitions of the correlation dimension for a probability measure. Journal of Statistical Physics, 86(3–4):707–720, February 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02199116.

Gallas:1996:MDS

[GHPS96] Jason A. C. Gallas, Hans J. Herrmann, Thorsten Pöschel, and Stefan Sokolowski. Molecular dynamics simulation of size segregation in three dimensions. Journal of Statistical Physics, 82(1– 2):443–450, January 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02189239.

Gitterman:1991:EOF

[GHW91] Moshe Gitterman, Shlomo Havlin, and George H. Weiss. Effects of an oscillating field on a diffusion process in the presence of a trap. Journal of Statistical Physics, 63(1–2):315–322, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01026607.

Gade:1992:NFC

[GI92] S. Gade and E. Gade III. Natural frequencies of the classical two-spin XXZ system. Journal of Statistical Physics, 68(5–6): 1143–1145, September 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048890.

Giacomin:1991:VWL

[Gia91] Giambattista Giacomin. Van der Waals limit and phase separation in a particle model with Kawasaki dynamics. Journal of Statistical Physics, 65(1–2):217–234, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01329857.

Gillis:1990:SDS

[Gil90] J. Gillis. The statistics of derangement — a survey. Journal of Statistical Physics, 58(3–4):575–578, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01112763.

Gitterman:1990:PEC

[Git90a] M. Gitterman. Phase equilibria and critical phenomena in closed reactive systems. Journal of Statistical Physics, 58(3–4):707–748, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01112772.

Gitterman:1990:BRC

[Git90b] Moshe Gitterman. Book review: Chaotic dynamics of nonlinear systems. Journal of Statistical Physics, 61(3–4):953–955, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01027315.

Galves:1991:SDT

[GIT91a] A. Galves, N. Ianiro, and L. Triolo. Self-diffusion in a twodimensional system of colliding vertical sticks. Journal of Statistical Physics, 65(1–2):205–215, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01329856.

Gitterman:1991:BRB

[Git91b] M. Gitterman. Book review: Chaotic dynamics — An introduction. Journal of Statistical Physics, 62(3–4):883–884, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017990.

Gitterman:1991:BRN

[Git91c] M. Gitterman. Book review: Noise and chaos in nonlinear dynamical systems. Journal of Statistical Physics, 62(3–4):879–881, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017989.

Gitterman:1992:BRS

[Git92a] Moshe Gitterman. Book review: A survey of nonlinear dynamics (‘chaos theory’). Journal of Statistical Physics, 69(5–6):1155, December 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01058770.

Gitterman:1992:BRB

[Git92b] Moshe Gitterman. Book review: From microphysics to macrophysics. Methods and applications of statistical physics. Journal of Statistical Physics, 69(5–6):1157–1158, December 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01058771.

Gitterman:1992:BRN

[Git92c] Moshe Gitterman. Book review: Nonlinear dynamics and chaos. Journal of Statistical Physics, 69(5–6):1153, December 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01058769.

Gitterman:1992:BRW

[Git92d] Moshe Gitterman. Book review: Weak chaos and quasi-regular patterns. Journal of Statistical Physics, 67(3–4):833–836, May 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01049731.

Gitterman:1993:BRQ

[Git93a] Moshe Gitterman. Book review: Quantum and statistical field theory. Journal of Statistical Physics, 70(5–6):1401–1402, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01049442.

Gitterman:1993:BRT

[Git93b] Moshe Gitterman. Book review: The transition to chaos in conservative classical systems: Quantum manifestations. Journal of Statistical Physics, 72(1–2):413–414, July 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01048057.

Gitterman:1994:IND

[Git94] Moshe Gitterman. Introduction to nonlinear dynamics for physicists. Journal of Statistical Physics, 74(5–6):1327–1328, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02188237.

Gitterman:1996:INS

[Git96a] Moshe Gitterman. Introduction to nonlinear science. Journal of Statistical Physics, 83(5–6):1261–1262, June 1996. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF02179564;   http://link.springer.com/article/
10.1007/BF02179564.

Gitterman:1996:RG

[Git96b] Moshe Gitterman. Renormalization group. Journal of Statistical Physics, 84(3–4):895–896, August 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179663.

Gitterman:1997:QST

[Git97] Moshe Gitterman. Quantum statistical theory of superconductivity. Journal of Statistical Physics, 88(5–6):1423–1424, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02732445.

Gitterman:1998:BRB

[Git98] Moshe Gitterman. Book review: Chaos in Atomic Physics. R. Blumel and W. P. Reinhard, Cambridge University Press, 1997. Journal of Statistical Physics, 91(1–2):477–479, April 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023068927394.

Gore:1999:SWP

[GJ99] Vivek K. Gore and Mark R. Jerrum. The Swendsen–Wang process does not always mix rapidly. Journal of Statistical Physics, 97(1–2):67–86, October 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004610900745.

Gruber:1992:GSS

[GJL92] C. Gruber, J. Jedrzejewski, and P. Lemberger. Ground states of the spinless Falicov–Kimball model. II. Journal of Statistical Physics, 66(3–4):913–938, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01055709.

Gabrielli:1999:OSM

[GJLL99] D. Gabrielli, G. Jona-Lasinio, and C. Landim. Onsager symmetry from microscopic TP invariance. Journal of Statistical Physics, 96(3–4):639–652, August 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004550307453.

Grassberger:1991:FFM

[GK91] Peter Grassberger and Holger Kantz. On a forest fire model with supposed self-organized criticality. Journal of Statistical Physics, 63(3–4):685–700, May 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01029205.

Golse:1995:NMC

[GK95] François Golse and Axel Klar. A numerical method for computing asymptotic states and outgoing distributions for kinetic linear half-space problems. Journal of Statistical Physics, 80(5–6): 1033–1061, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179863.

Gamba:1996:LSC

[GK96] Andrea Gamba and Igor V. Kolokolov. The Lyapunov spectrum of a continuous product of random matrices. Journal of Statistical Physics, 85(3–4):489–499, November 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02174216.

Gamba:1999:DSP

[GK99] Andrea Gamba and Igor V. Kolokolov. Dissipation statistics of a passive scalar in a multidimensional smooth flow. Journal of Statistical Physics, 94(5–6):759–777, March 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004522830805.

Gupalo:1994:SLS

[GKC94] D. Gupalo, A. S. Kaganovich, and E. G. D. Cohen. Symmetry of Lyapunov spectrum. Journal of Statistical Physics, 74(5–6): 1145–1159, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02188220.

Gorshkov:1994:RPL

[GKRT94] K. A. Gorshkov, L. N. Korzinov, M. I. Rabinovich, and L. S. Tsimring. Random pining of localized states and the birth of deterministic disorder within gradient models. Journal of Statistical Physics, 74(5–6):1033–1045, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02188216.

Gu:1993:RFS

[GKT93] Xiao-Yue Gu, G. Kalman, and Z. C. Tao. Response function of the second kind in many-body systems. Journal of Statistical Physics, 70(3–4):887–898, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053598.

Glazier:1991:DC

[GKW91] James A. Glazier, Paul Kolodner, and Hugh Williams. Dispersive chaos. Journal of Statistical Physics, 64(5–6):945–960, September 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048806.

Gausterer:1993:MCL

[GL93] H. Gausterer and Sean Lee. The mechanism of complex Langevin simulations. Journal of Statistical Physics, 73(1–2):147–157, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01052754.

Gruber:1994:P

[GL94] Christian Gruber and Joel L. Lebowitz. Preface. Journal of Statistical Physics, 76(1–2):1, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02188655.

Giacomin:1997:PSD

[GL97] Giambattista Giacomin and Joel L. Lebowitz. Phase segregation dynamics in particle systems with long range interactions. I. Macroscopic limits. Journal of Statistical Physics, 87(1–2):37– 61, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02181479.

Giardina:1998:EPM

[GL98] Cristian Giardinà and Roberto Livi. Ergodic properties of microcanonical observables. Journal of Statistical Physics, 91(5– 6):1027–1045, June 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023036101468.

Giacomin:1995:APP

[GLM95] G. Giacomin, J. L. Lebowitz, and C. Maes. Agreement percolation and phase coexistence in some Gibbs systems. Journal of Statistical Physics, 80(5–6):1379–1403, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179875.

Grossmann:1998:SII

[GLR98] Siegfried Grossmann, Detlef Lohse, and Achim Reeh. Scaling of the irreducible SO(3)-invariants of velocity correlations in turbulence. Journal of Statistical Physics, 93(3–4):715–724, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000033250.57019.46.

Goles:1988:OSD

[GM88] Eric Goles and Servet Martinez. The one-site distribution of Gibbs states on Bethe lattice are probability vectors of period ≤ 2 for a nonlinear transformation. Journal of Statistical Physics, 52(1–2):267–285, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01016414. See erratum [GM96].

Gallavotti:1992:BR

[GM92] Giovanni Gallavotti and Neal Madras. Book review. Journal of Statistical Physics, 69(1–2):443–449, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01053802.

Garrido:1994:KLM

[GM94] P. L. Garrido and J. Marro. Kinetic lattice models of disorder. Journal of Statistical Physics, 74(3–4):663–686, February 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02188575.

Gielis:1995:URG

[GM95] G. Gielis and C. Maes. The uniqueness regime of Gibbs fields with unbounded disorder. Journal of Statistical Physics, 81(3– 4):829–835, November 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179259.

Goles:1996:EOS

[GM96] Eric Goles and Servet Martínez. Erratum: The one-site distribution of Gibbs states on Bethe lattice are probability vectors of period ≤ 2 for a nonlinear transformation. Journal of Statistical Physics, 82(1–2):451, January 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02189240.

Glashow:1997:TRR

[GM97a] Sheldon Lee Glashow and Laurence Mittag. Three rods on a ring and the triangular billiard. Journal of Statistical Physics, 87(3–4):937–941, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181254.

Gravner:1997:BPP

[GM97b] Janko Gravner and Elaine McDonald. Bootstrap percolation in a polluted environment. Journal of Statistical Physics, 87 (3–4):915–927, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181252.

Gabrielli:1996:CRT

[GMCP96] A. Gabrielli, M. Marsili, R. Cafiero, and L. Pietronero. Comment on the run time statistics in models of growth in disordered media. Journal of Statistical Physics, 84(3–4):889–893, August 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179662.

Goldenfeld:1998:BSP

[GMH98] Nigel Goldenfeld, Alan McKane, and Qing Hou. Block spins for partial differential equations. Journal of Statistical Physics, 93(3–4):699–714, November 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033249.19382.d9.

Gonzalez-Miranda:1990:MCS

[GMM90] J. M. González-Miranda and J. Marro. Monte Carlo study of the generalized reaction-diffusion lattice-gas model system. Journal of Statistical Physics, 61(5–6):1283–1293, December 1990.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014375.

Gruber:1997:GSF

[GMMU97] Christian Gruber, Nicolas Macris, Alain Messager, and Daniel Ueltschi. Ground states and flux configurations of the twodimensional Falicov–Kimball model. Journal of Statistical Physics, 86(1–2):57–108, January 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02180199.

Giacometti:1994:SMR

[GMN94] Achille Giacometti, Amos Maritan, and Hisao Nakanishi. Statistical mechanics of random paths on disordered lattices. Journal of Statistical Physics, 75(3–4):669–706, May 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02186876.

Gelbard:1991:NMD

[GMO91] Fred Gelbard, Lisa A. Mondy, and Steven E. Ohrt. A new method for determining hydrodynamic effects on the collision of two spheres. Journal of Statistical Physics, 62(5–6):945–960, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01128170.

Gammaitoni:1993:SRP

[GMPS93] L. Gammaitoni, M. Martinelli, L. Pardi, and S. Santucci. Stochastic resonance in paramagnetic resonance systems. Journal of Statistical Physics, 70(1–2):425–435, January 1993. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01053979.

Gerisch:1998:CVG

[GMR98] T. Gerisch, R. Münzner, and A. Rieckers. Canonical versus grand–canonical free energies and phase diagrams of a bipolaronic superconductor model. Journal of Statistical Physics, 93(5–6):1021–1049, December 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033152.73539.a2.

Gerisch:1999:GDK

[GMR99] Thomas Gerisch, Roland Münzner, and Alfred Rieckers. Global C∗-dynamics and its KMS states of weakly inhomogeneous bipolaronic superconductors. Journal of Statistical Physics, 97(3–4): 751–779, November 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004671426805.

Giacometti:1995:RSR

[GMTB95] Achille Giacometti, Amos Maritan, Flavio Toigo, and Jayanth R. Banavar. Real-space renormalization group for Langevin dynamics in absence of translational invariance. Journal of Statistical Physics, 79(3–4):649–668, May 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02184874.

Giacometti:1996:FDT

[GMTB96] Achille Giacometti, Amos Maritan, Flavio Toigo, and Jayanth R. Banavar. Fluctuation–dissipation theorem and the dynamical renormalization group. Journal of Statistical Physics, 82(5–6): 1669–1674, March 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183399.

Glockle:1993:FFR

[GN93] Walter G. Glöckle and Theo F. Nonnenmacher. Fox function representation of non-Debye relaxation processes. Journal of Statistical Physics, 71(3–4):741–757, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01058445.

Gielerak:1995:GPF

[GO95] Roman Gielerak and Robert Olkiewicz. Gentle perturbations of the free Bose gas. I. Journal of Statistical Physics, 80(3– 4):875–918, August 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02178559.

Gobron:1992:DMP

[Gob92] T. Gobron. On a discrete model of phase separation dynamics. Journal of Statistical Physics, 69(5–6):995–1024, December 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01058759.

Golden:1999:ALB

[Gol99] Sidney Golden. Asymptotic lower bound for the relative disparities of truncated-path-integral partition functions. Journal of Statistical Physics, 95(1–2):495–502, April 1999. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004550100048.

Goncalves:1994:TDL

[Gon94] N. J. A. P. Gonçalves. Two-dimensional lattice tree exponents and amplitudes: Simulation algorithms versus series. Journal of Statistical Physics, 74(3–4):909–917, February 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02188584.

Gotoh:1990:BMR

[Got90] Toshiyuki Gotoh. Brownian motion in a rotating flow. Journal of Statistical Physics, 59(1–2):371–402, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01015575.

Gotze:1996:BIM

[Göt96] W. Götze. Bifurcations of an interated mapping with retardations. Journal of Statistical Physics, 83(5–6):1183–1197, June 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02179557.

Goudon:1997:BEF

[Gou97] T. Goudon. On Boltzmann equations and Fokker–Planck asymptotics: Influence of grazing collisions. Journal of Statistical Physics, 89(3–4):751–776, November 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02765543.

Garcia:1991:FHP

[GP91] Alejandro García and Cécile Penland. Fluctuating hydrodynamics and principal oscillation pattern analysis. Journal of Statistical Physics, 64(5–6):1121–1132, September 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01048818.

Grabenstein:1993:KMM

[GP93] Martin Grabenstein and Klaus Pinn. Kinematics of multigrid Monte Carlo. Journal of Statistical Physics, 71(3–4):607–640, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01058439.

Glotzer:1992:TDT

[GPJ92] Sharon C. Glotzer, Peter H. Poole, and Naeem Jan. Timedependent thermodynamic properties of the Ising model from damage spreading. Journal of Statistical Physics, 68(5–6):895– 910, September 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048879.

Gredeskul:1990:TPR

[GPS90] S. A. Gredeskul, L. A. Pastur, and P. Seba. Transmission properties of random point scatterers for waves with two-band dispersion law. Journal of Statistical Physics, 58(5–6):795–816, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01026551.

Garcia-Pelayo:1993:BPM

[GPSS93] Ricardo García-Pelayo, Iván Salazar, and William C. Schieve. A branching process model for sand avalanches. Journal of Statistical Physics, 72(1–2):167–187, July 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048045.

Gonchar:1992:ORD

[GR92] N. S. Gonchar and A. B. Rudyk. Oscillation of the radial distribution function. Journal of Statistical Physics, 68(5–6):1065– 1087, September 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048885.

Guerberoff:1997:FEH

[GR97] G. R. Guerberoff and G. A. Raggio. On the free energy of the Hopfield model. Journal of Statistical Physics, 87(1–2):333–361, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02181491.

Gerisch:1998:LGS

[GR98] T. Gerisch and A. Rieckers. Limiting Gibbs states and phase transitions of a bipartite mean-field Hubbard model. Journal of Statistical Physics, 91(3–4):759–785, May 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023089930061.

Graham:1990:MTA

[Gra90] Robert Graham. Macroscopic theory of activated decay of metastable states. Journal of Statistical Physics, 60(5–6):675– 694, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01025989.

Graf:1994:ALS

[Gra94] Gian Michele Graf. Anderson localization and the space–time characteristic of continuum states. Journal of Statistical Physics, 75(1–2):337–346, April 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186292.

Grassberger:1995:DST

[Gra95] Peter Grassberger. Are damage spreading transitions generically in the universality class of directed percolation? Journal of Statistical Physics, 79(1–2):13–23, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179381.

Gregorio:1990:ODM

[Gre90] Salvatore De Gregorio. On a one-dimensional model for the three-dimensional vorticity equation. Journal of Statistical Physics, 59(5–6):1251–1263, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01334750.

Grimmett:1994:PMR

[Gri94] Geoffrey Grimmett. Potts models and random-cluster processes with many-body interactions. Journal of Statistical Physics, 75 (1–2):67–121, April 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186281.

Grimmett:1995:CDO

[Gri95] Geoffrey Grimmett. Comparison and disjoint-occurrence inequalities for random-cluster models. Journal of Statistical Physics, 78(5–6):1311–1324, March 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02180133.

Gropengiesser:1995:GSE

[Gro95] Uwe Gropengiesser. The ground-state energy of the ±J sping glass. A comparison of various biologically motivated algorithms. Journal of Statistical Physics, 79(5–6):1005–1012, June 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02181213.

Grynberg:1992:CIC

[Gry92] Marcelo D. Grynberg. Continuously infinite commensurateincommensurate phase transition of a two-dimensional competing Ising model. Journal of Statistical Physics, 69(3–4): 869–878, November 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01050438.

Ge:1990:RBT

[GRZ90] Yuzhen Ge, Edmond Rusjan, and Paul Zweifel. Renormalization of binary trees derived from one-dimensional unimodal maps. Journal of Statistical Physics, 59(5–6):1265–1295, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01334751.

Gaveau:1990:MAQ

[GS90a] Bernard Gaveau and L. S. Schulman. Model apparatus for quantum measurements. Journal of Statistical Physics, 58(5– 6):1209–1230, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01026572.

Given:1990:KSE

[GS90b] James A. Given and George Stell. The Kirkwood–Salsburg equations for random continuum percolation. Journal of Statistical Physics, 59(3–4):981–1018, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01025859.

Grannan:1990:RRM

[GS90c] E. R. Grannan and G. Swindle. Rigorous results on mathematical models of catalytic surfaces. Journal of Statistical Physics, 61(5–6):1085–1103, December 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014366. See extension

[MS92].

Garzo:1991:ESB

[GS91a] V. Garzó and A. Santos. Exact solution of the Boltzmann equation in the homogeneous color conductivity problem. Journal of Statistical Physics, 65(3–4):747–760, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01053752.

Glimm:1991:RFM

[GS91b] James Glimm and David H. Sharp. A random field model for anomalous diffusion in heterogeneous porous media. Journal of Statistical Physics, 62(1–2):415–424, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020877.

Gaveau:1992:ADR

[GS92] B. Gaveau and L. S. Schulman. Anomalous diffusion in a random velocity field. Journal of Statistical Physics, 66(1–2):375–383, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01060072.

Gaveau:1993:FSS

[GS93a] B. Gaveau and L. S. Schulman. Finite-size scaling for meanfield percolation. Journal of Statistical Physics, 70(3–4):613–634, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01053587.

Gitterman:1993:BR

[GS93b] M. Gitterman and John L. Spouge. Book review. Journal of Statistical Physics, 73(1–2):449–452, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01052775.

Grimm:1993:SHC

[GS93c] Uwe Grimm and Gunter Schütz. The spin-1/2 XXZ Heisenberg chain, the quantum algebra U_{q}[sl(2)], and duality transformations for minimal models. Journal of Statistical Physics, 71 (5–6):923–966, June 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049955.

Gaveau:1994:FMF

[GS94] B. Gaveau and L. S. Schulman. Fluctuations in mean-field selforganized criticality. Journal of Statistical Physics, 74(3–4): 607–630, February 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02188573.

Gupta:1995:SCP

[GS95] Abhijit Kar Gupta and Asok K. Sen. Shape of the clusters of a percolative system in the presence of tunneling bonds and a finite electric field. Journal of Statistical Physics, 80(5–6): 1425–1431, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179878.

Greenwood:1997:ELD

[GS97] Priscilla E. Greenwood and Jiaming Sun. Equivalences of the large deviation principle for Gibbs measures and critical balance in the Ising model. Journal of Statistical Physics, 86(1– 2):149–164, January 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02180201.

Greenwood:1998:CCI

[GS98] Priscilla E. Greenwood and Jiaming Sun. On criticality for competing influences of boundary and external field in the Ising model. Journal of Statistical Physics, 92(1–2):35–45, July 1998.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023039401489.

Ghosh:1990:SRH

[GSCK90] M. Ghosh, A. K. Sen, B. K. Chakrabarti, and G. A. Kohring. Slowing down of retrieval in the Hopfield model. Journal of Statistical Physics, 61(1–2):501–504, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01013978.

Gelfand:1990:PEQ

[GSH90] Martin P. Gelfand, Rajiv R. P. Singh, and David A. Huse. Perturbation expansions for quantum many-body systems. Journal of Statistical Physics, 59(5–6):1093–1142, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01334744.

Garcia:1998:LRC

[GSM98] Alejandro L. Garcia, G. Sonnino, and M. Malek Mansour. Longranged correlations in bounded nonequilibrium fluids. Journal of Statistical Physics, 90(5–6):1489–1492, March 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023260201995.

Gabetta:1995:MPD

[GTW95] G. Gabetta, G. Toscani, and B. Wennberg. Metrics for probability distributions and the trend to equilibrium for solutions of the Boltzmann equation. Journal of Statistical Physics, 81(5–6): 901–934, December 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179298.

Guiol:1999:SPS

[Gui99] H. Guiol. Some properties of k-step exclusion processes. Journal of Statistical Physics, 94(3–4):495–511, February 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004548321062.

Gruber:1994:MFF

[GUJ94] C. Gruber, D. Ueltschi, and J. Jedrzejewski. Molecule formation and the Farey tree in the one-dimensional Falicov–Kimball model. Journal of Statistical Physics, 76(1–2):125–157, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02188658.

Gunton:1999:HN

[Gun99] J. D. Gunton. Homogeneous nucleation. Journal of Statistical Physics, 95(5–6):903–923, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004598332758.

Gutkin:1996:BPS

[Gut96] Eugene Gutkin. Billiards in polygons: Survey of recent results. Journal of Statistical Physics, 83(1–2):7–26, April 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183637.

Guyonnet:1991:AEN

[Guy91] Regis Guyonnet. Adsorption effects and NMR properties: Continuous chain approach. Journal of Statistical Physics, 65(1– 2):269–289, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01329861.

Goderis:1991:GDF

[GVV91] D. Goderis, A. Verbeure, and P. Vets. Glauber dynamics of fluctuations. Journal of Statistical Physics, 62(3–4):759–777, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017982.

Gates:1990:SCG

[GW90] D. J. Gates and M. Westcott. On the stability of crystal growth. Journal of Statistical Physics, 59(1–2):73–101, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01015564.

Gitterman:1993:BPF

[GW93a] Moshe Gitterman and George H. Weiss. The behavior of a periodically-forced nonlinear system subject to additive noise. Journal of Statistical Physics, 71(5–6):1213–1220, June 1993.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01049969.

Gitterman:1993:EPD

[GW93b] Moshe Gitterman and George H. Weiss. ‘escape’ of a periodically driven particle from a metastable state in a noisy system. Journal of Statistical Physics, 70(1–2):107–123, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01053957.

Gates:1994:SBR

[GW94a] D. J. Gates and M. Westcott. Seven basic regimes of steady crystal growth in two dimensions. Journal of Statistical Physics, 77 (1–2):199–215, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186839.

Gitterman:1994:CET

[GW94b] Moshe Gitterman and George H. Weiss. A comment on earlytime solutions of the Smoluchowski equation. Journal of Statistical Physics, 74(3–4):947–951, February 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02188588.

Gitterman:1994:TNL

[GW94c] Moshe Gitterman and George H. Weiss. A transition in a noisy linear system driven by a periodic signal. Journal of Statistical Physics, 74(3–4):941–946, February 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02188587.

Grigolini:1994:QDS

[GW94d] Paolo Grigolini and Bruce J. West. Quantum dissipative systems. Journal of Statistical Physics, 77(3–4):951–952, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179474.

Gates:1995:SSC

[GW95] D. J. Gates and M. Westcott. Stationary states of crystal growth in three dimensions. Journal of Statistical Physics, 81(3–4): 681–715, November 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179253.

Gates:1999:PFC

[GW99] D. J. Gates and M. Westcott. Predicting fiber contact in a threedimensional model of paper. Journal of Statistical Physics, 94 (1–2):31–52, January 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004503111067.

Giga:1993:BPI

[GY93] Yoshikazu Giga and Zensho Yoshida. A bound for the pressure integral in a plasma equilibrium. Journal of Statistical Physics, 72(5–6):1375–1389, September 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01048191.

Ginoza:1998:EAS

[GY98] M. Ginoza and M. Yasutomi. An extension of the analytical solution of the Ornstein–Zernike equation with the Yukawa closure. Journal of Statistical Physics, 90(5–6):1475–1480, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023256101086.

Guionnet:1997:DER

[GZ97] Alice Guionnet and Bogusław Zegarliński. Decay to equilibrium in random spin systems on a lattice. II. Journal of Statistical Physics, 86(3–4):899–904, February 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02199123.

Georgii:1998:IMM

[GZ98] Hans-Otto Georgii and Valentin Zagrebnov. On the interplay of magnetic and molecular forces in Curie–Weiss ferrofluid models. Journal of Statistical Physics, 93(1–2):79–107, October 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1023/B%3AJOSS.0000026728.01594.18.pdf.

Hovi:1997:DSA

[HA97] J.-P. Hovi and Amnon Aharony. Different self-avoiding walks on percolation clusters: A small-cell real-space renormalizationgroup study. Journal of Statistical Physics, 86(5–6):1163–1178, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183619.

Haggstrom:1996:ASQ

[Häg96] Olle Häggström. Almost sure quasilocality fails for the randomcluster model on a tree. Journal of Statistical Physics, 84(5–6): 1351–1361, September 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02174134.

Haggstrom:1998:RCA

[Häg98] Olle Häggström. Random-cluster analysis of a class of binary lattice gases. Journal of Statistical Physics, 91(1–2):47–74, April 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1023031819217.

Haller:1997:UHB

[Hal97] G. Haller. Universal homoclinic bifurcations and chaos near double resonances. Journal of Statistical Physics, 86(5–6):1011–1051, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183612.

Handjani:1995:RMS

[Han95] Shirin J. Handjani. The reversible measures for symmetric nearest-particle systems. Journal of Statistical Physics, 80(5–6): 1119–1164, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179866.

Handa:1996:EPP

[Han96] Kenji Handa. Entropy production per site in (nonreversible) spin-flip processes. Journal of Statistical Physics, 83(3–4):555– 571, May 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183740.

Hastings:1998:NHF

[Has98] M. B. Hastings. Non-Hermitian fermion mapping for onecomponent plasma. Journal of Statistical Physics, 90(1–2): 311–326, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023268105180.

Havlin:1990:TRC

[Hav90] Shlomo Havlin. Transport in random correlated fields. Journal of Statistical Physics, 58(3–4):653–668, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01112769.

Hayot:1992:RSL

[Hay92] F. Hayot. Reynolds stresses in a lattice gas. Journal of Statistical Physics, 68(3–4):557–562, August 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01341762.

Haydn:1993:GMA

[Hay93] Nicolai T. A. Haydn. Gibbs measures for axiom a flows. Journal of Statistical Physics, 72(1–2):309–327, July 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01048052.

Hayli:1996:NEF

[Hay96] Avram Hayli. Numerical exploration of a family of strictly convex billiards with boundary of class C^{2}. Journal of Statistical Physics, 83(1–2):71–79, April 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02183640.

Haydn:1999:DFR

[Hay99] Nicolai Haydn. The distribution of the first return time for rational maps. Journal of Statistical Physics, 94(5–6):1027–1036, March 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004543302580.

Huber:1991:EDT

[HC91] D. L. Huber and W. Y. Ching. Effect of disorder on the trapping of Frenkel excitons in three-dimensional systems. Journal of Statistical Physics, 65(5–6):1155–1159, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01049604.

Hanson:1992:ABP

[HC92] James E. Hanson and James P. Crutchfield. The attractorbasin portrait of a cellular automaton. Journal of Statistical Physics, 66(5–6):1415–1462, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01054429.

Hu:1996:HMC

[HCW96] Chin-Kun Hu, Chi-Ning Chen, and F. Y. Wu. Histogram Monte Carlo position-space renormalization group: Applications to the site percolation. Journal of Statistical Physics, 82(3–4): 1199–1206, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179808.

Hinrichsen:1998:DSI

[HDS98] Haye Hinrichsen, Eytan Domany, and Dietrich Stauffer. Damage spreading in a 2D Ising model with Swendsen–Wang dynamics. Journal of Statistical Physics, 91(3–4):807–814, May 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023046115040.

Heintz:1994:MNB

[Hei94] Alexei Heintz. In memory of Nina Borisovna Maslova. Journal of Statistical Physics, 77(5–6):1049–1050, December 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183151.

Heintz:1998:IBV

[Hei98] A. Heintz. On the initial boundary value problems for the Enskog equation in irregular domains. Journal of Statistical Physics, 90(3–4):663–695, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023268718526.

Helsing:1998:HOA

[Hel98] Johan Helsing. A high–order accurate algorithm for electrostatics of overlapping disks. Journal of Statistical Physics, 90(5–6): 1461–1473, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023204117016.

Henon:1992:IFL

[Hén92] M. Hénon. Implementation of the FCHC lattice gas model on the connection machine. Journal of Statistical Physics, 68(3– 4):353–377, August 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01341753.

Henkel:1994:SIS

[Hen94] Malte Henkel. Schrödinger invariance and strongly anisotropic critical systems. Journal of Statistical Physics, 75(5–6):1023– 1061, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/BF02186756.

Henley:1997:RTD

[Hen97] Christopher L. Henley. Relaxation time for a dimer covering with height representation. Journal of Statistical Physics, 89(3– 4):483–507, November 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02765532.

Heuer:1993:CDT

[Heu93] Hans-Otto Heuer. Critical dynamics of the three-dimensional Ising model: A Monte Carlo study. Journal of Statistical Physics, 72(3–4):789–827, August 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01048033.

Hamm:1992:QSN

[HG92] Andreas Hamm and Robert Graham. Quasipotentials for simple noisy maps with complicated dynamics. Journal of Statistical Physics, 66(3–4):689–725, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01055697.

Honnell:1990:EES

[HH90] Kevin G. Honnell and Carol K. Hall. Exact equations of state for one-dimensional chain fluids. Journal of Statistical Physics, 61(3–4):803–842, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01027302.

Hennecke:1993:CDC

[HH93] M. Hennecke and U. Heyken. Critical dynamics of cluster algorithms in the dilute Ising model. Journal of Statistical Physics, 72 (3–4):829–844, August 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048034.

Heumann:1995:GAM

[HH95a] Michael Heumann and Michael Hötzel. Generalization of an aging model. Journal of Statistical Physics, 79(1–2):483–490, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179400.

Hu:1995:RST

[HH95b] Zhan-Ning Hu and Bo-Yu Hou. Remarks on the star-triangle relation in the Baxter–Bazhanov model. Journal of Statistical Physics, 79(3–4):759–764, May 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02184882.

Hu:1996:TDV

[HH96] Zhan-Ning Hu and Bo-Yu Hou. Three-dimensional vertex model in statistical mechanics from Baxter–Bazhanov model. Journal of Statistical Physics, 82(3–4):633–655, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179789.

Halsey:1996:MDB

[HHD96] Thomas C. Halsey, Katsuya Honda, and Bertrand Duplantier. Multifractal dimensions for branched growth. Journal of Statistical Physics, 85(5–6):681–743, December 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02199360.

Hioe:1990:RWS

[Hio90] F. T. Hioe. Restricted walks, stability-instability transitions, and dynamic symmetries. Journal of Statistical Physics, 58(3– 4):627–641, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01112767.

Huber:1990:TRO

[HJ90] Albrecht Huber and Hans-Ulrich Jüttner. On the thermodynamic V -representability of one-particle density matrices. Journal of Statistical Physics, 61(1–2):423–441, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013974.

Hanggi:1993:CCN

[HJZM93] Peter Hänggi, Peter Jung, Christine Zerbe, and Frank Moss. Can colored noise improve stochastic resonance? Journal of Statistical Physics, 70(1–2):25–47, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053952.

Haller:1996:ARG

[HK96] Karl Haller and Tom Kennedy. Absence of renormalization group pathologies near the critical temperature. Two examples. Journal of Statistical Physics, 85(5–6):607–637, December 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02199358.

Hunt:1996:FPC

[HKSY96] Brian R. Hunt, Konstantin M. Khanin, Yakov G. Sinai, and James A. Yorke. Fractal properties of critical invariant curves. Journal of Statistical Physics, 85(1–2):261–276, October 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02175565.

Hede:1991:SAF

[HKV91] B. Hede, J. Kertész, and T. Vicsek. Self-affine fractal clusters: Conceptual questions and numerical results for directed percolation. Journal of Statistical Physics, 64(3–4):829–841, August 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048318.

Handrich:1997:FCD

[HL97a] K. Handrich and F.-P. Ludwig. Friction coefficients and directed motion of asymmetric test particles. Journal of Statistical Physics, 86(5–6):1067–1087, March 1997. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183614.

He:1997:LBM

[HL97b]  Xiaoyi He and Li-Shi Luo.  Lattice Boltzmann model

for the incompressible Navier–Stokes equation. Journal of Statistical Physics, 88(3–4):927–944, August 1997. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000015179.12689.e4.

Ho:1997:FGL

[HL97c] T. G. Ho and L. J. Landau. Fermi gas on a lattice in the van Hove limit. Journal of Statistical Physics, 87(3–4):821–845, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02181246.

Hansen:1993:TCI

[HLIM93] Per Lyngs Hansen, Jesper Lemmich, John Hjort Ipsen, and Ole G. Mouritsen. Two coupled Ising planes: Phase diagram and interplanar force. Journal of Statistical Physics, 73(3–4): 723–749, November 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01054347.

Hupfer:1999:POG

[HLW99] Thomas Hupfer, Hajo Leschke, and Simone Warzel. Poissonian obstacles with Gaussian walls discriminate between classical and quantum Lifshits tailing in magnetic fields. Journal of Statistical Physics, 97(3–4):725–750, November 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004619409967.

Hasslacher:1992:LGE

[HM92a] Brosl Hasslacher and David A. Meyer. Lattice gases and exactly solvable models. Journal of Statistical Physics, 68(3–4):575–590, August 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01341764.

Hunt:1992:AIM

[HM92b] Fern Y. Hunt and Walter M. Miller. On the approximation of invariant measures. Journal of Statistical Physics, 66(1–2):

535–548, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01060079.

Hasenbusch:1996:RTT

[HMP96] M. Hasenbusch, S. Meyer, and M. Pütz. The roughening transition of the three-dimensional Ising interface: A Monte Carlo study. Journal of Statistical Physics, 85(3–4):383–401, November 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02174211.

Hansel:1990:TDB

[HMV90] D. Hansel, C. Meunier, and A. Verga. Transient dynamical behavior and phase transitions in magnetic systems. Journal of Statistical Physics, 61(1–2):329–343, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01013968.

Harris:1996:CWT

[HMY96] A. B. Harris, C. Micheletti, and J. M. Yeomans. Complete wetting in the three-dimensional transverse Ising model. Journal of Statistical Physics, 84(3–4):323–335, August 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179645.

Havlin:1991:GHW

[HNS91] Shlomo Havlin, Ralph Nossal, and Michael Shlesinger. George Herbert Weiss. Journal of Statistical Physics, 65(5–6):837– 838, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049583.

Hof:1993:SRD

[Hof93] A. Hof. Some remarks on discrete aperiodic Schrödinger operators. Journal of Statistical Physics, 72(5–6):1353–1374, September 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048190.

Hof:1995:RSO

[Hof95] A. Hof. A remark on Schrödinger operators on aperiodic tilings. Journal of Statistical Physics, 81(3–4):851–855, November 1995.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179262.

Hof:1996:SIB

[Hof96] A. Hof. On a ‘structure intermediate between quasiperiodic and random’. Journal of Statistical Physics, 84(1–2):309–320, July 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179589.

Hollschneider:1994:MAL

[Hol94] Matthias Hollschneider. More on the analysis of local regularity through wavelets. Journal of Statistical Physics, 77(3–4): 807–840, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179462.

Honecker:1996:PAS

[Hon96] A. Honecker. A perturbative approach to spectrum and correlation functions of the chiral Potts model. Journal of Statistical Physics, 82(3–4):687–741, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179791.

Horsthemke:1993:NIC

[Hor93] Werner Horsthemke. Noise-induced clumping in the onedimensional reversible diffusion-limited single-species coagulation process. Journal of Statistical Physics, 70(1–2):149–162, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01053960.

Hede:1991:TSM

[HP91] B. Hede and V. Privman. Two-spin-majority cellular automaton as a model of 2D cluster and interface growth. Journal of Statistical Physics, 65(1–2):379–385, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01329866.

Honecker:1997:MPS

[HP97] A. Honecker and I. Peschel. Matrix-product states for a onedimensional lattice gas with parallel dynamics. Journal of Statistical Physics, 88(1–2):319–345, July 1997. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02508474.

Huckaby:1994:EEP

[HPS94] Dale A. Huckaby, Radu Pitis, and Masato Shinmi. Existence of enantiomeric phase separation in a three-dimensional lattice gas model. Journal of Statistical Physics, 75(5–6):981–995, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02186754.

Hradil:1994:PCS

[HQSS94] Z. Hradil, A. Quattropani, V. Savona, and P. Schwendimann. Polaritons in confined systems. Journal of Statistical Physics, 76 (1–2):299–305, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02188663.

Holme:1992:LGL

[HR92a] Richard Holme and Daniel H. Rothman. Lattice-gas and lattice-Boltzmann models of miscible fluids. Journal of Statistical Physics, 68(3–4):409–429, August 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01341756.

Huckaby:1992:ESS

[HR92b] Dale A. Huckaby and Franz S. Rys. Existence of several surfacereconstructed phases in a two-dimensional lattice model. Journal of Statistical Physics, 66(5–6):1215–1224, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01054420.

Hafskjold:1995:CLE

[HR95] Bjørn Hafskjold and Signe Kjelstrup Ratkje. Criteria for local equilibrium in a system with transport of heat and mass. Journal of Statistical Physics, 78(1–2):463–494, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183360.

Hinrichsen:1997:UPS

[HRS97] Haye Hinrichsen, Vladimir Rittenberg, and Horatin Simon. Universality properties of the stationary states in the onedimensional coagulation-diffusion model with external particle input. Journal of Statistical Physics, 86(5–6):1203–1235, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02183621.

Hara:1990:UCD

[HS90a] Takashi Hara and Gordon Slade. On the upper critical dimension of lattice trees and lattice animals. Journal of Statistical Physics, 59(5–6):1469–1510, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01334760.

Huckaby:1990:ECS

[HS90b] Dale A. Huckaby and Masato Shinmi. Exact coexistence surfaces containing double critical points for a three-component solution on the Bethe, honeycomb, and square lattices. Journal of Statistical Physics, 60(3–4):347–361, August 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01314925.

Hatano:1991:EFT

[HS91] Naomichi Hatano and Masuo Suzuki. Effective-field theory of spin glasses and the coherent-anomaly method. I. Journal of Statistical Physics, 63(1–2):25–46, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01026590.

Hara:1992:NSB

[HS92a] Takashi Hara and Gordon Slade. The number and size of branched polymers in high dimensions. Journal of Statistical Physics, 67(5–6):1009–1038, June 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01049008.

Hatano:1992:EFT

[HS92b] Naomichi Hatano and Masuo Suzuki. Effective-field theory of spin glasses and the coherent-anomaly method. II. Double-cluster approximation. Journal of Statistical Physics, 66(3–4):897–911, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01055708.

Holovatch:1992:CER

[HS92c] Yu. Holovatch and M. Shpot. Critical exponents of random Ising-like systems in general dimensions. Journal of Statistical Physics, 66(3–4):867–883, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01055706.

Helffer:1994:CKL

[HS94a] Bernard Helffer and Johannes Sjöstrand. On the correlation for Kac-like models in the convex case. Journal of Statistical Physics, 74(1–2):349–409, January 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02186817.

Hoye:1994:PIF

[HS94b] J. S. Høye and G. Stell. Path integral formulation for quantized fermion and boson fluids. Journal of Statistical Physics, 77(1– 2):361–381, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186847.

Henry:1995:NER

[HS95a] B. I. Henry and T. Szeredi. New equipartition results for normal mode energies of anharmonic chains. Journal of Statistical Physics, 78(3–4):1039–1053, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183700.

Hodgdon:1995:ECP

[HS95b] Jennifer A. Hodgdon and Frank H. Stillinger. Equilibrium concentration of point defects in crystalline ^{4}He at 0K. Journal of Statistical Physics, 78(1–2):117–134, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183341.

Hoye:1995:CFC

[HS95c] J. S. Høye and G. Stell. Correlation functions in the cavity model of ionic fluids. Journal of Statistical Physics, 78(3–4):1171–1172, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183710.

Hoye:1997:SOI

[HS97] J. S. Høye and G. Stell. The spin-one Ising model in the meanspherical approximation. Journal of Statistical Physics, 89(1– 2):177–201, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02770760.

Hiemer:1998:PBS

[HS98] Philipp Hiemer and Vadim Snurnikov. Polygonal billiards with small obstacles. Journal of Statistical Physics, 90(1–2):453–466, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/

10.1023/A%3A1023228423836. See counter-example [Mon04].

Hu:1991:RKA

[HSK91] Bambi Hu, Jicong Shi, and Sang-Yoon Kim. Recurrence of Kolmogorov–Arnold–Moser tori in nonanalytic twist maps. Journal of Statistical Physics, 62(3–4):631–649, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017977.

Hara:1993:NLB

[HSS93] Takashi Hara, Gordon Slade, and Alan D. Sokal. New lower bounds on the self-avoiding-walk connective constant. Journal of Statistical Physics, 72(3–4):479–517, August 1993. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF01048021. See erratum [HSS95].

Hara:1995:ENL

[HSS95] Takashi Hara, Gordon Slade, and Alan D. Sokal. Erratum: New lower bounds on the self-avoiding-walk connective constant. Journal of Statistical Physics, 78(3–4):1187–1188, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183717. See [HSS93].

Henderson:1997:SOP

[HSW97] Douglas Henderson, Stefan Sokolowski, and Darsh Wasan. Second-order Percus–Yevick theory for a confined hard-sphere fluid. Journal of Statistical Physics, 89(1–2):233–247, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02770763.

Halsey:1990:FIC

[HT90a] Thomas C. Halsey and Will Toor. Fluctuation-induced couplings between defect lines or particle chains. Journal of Statistical Physics, 61(5–6):1257–1281, December 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01014374.

Hubmer:1990:KBL

[HT90b] G. F. Hubmer and U. M. Titulaer. The kinetic boundary layer for the linearized Boltzmann equation around an absorbing sphere. Journal of Statistical Physics, 59(1–2):441–459, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01015577.

Hubmer:1991:KMD

[HT91] G. F. Hubmer and U. M. Titulaer. A kinetic model for droplet growth in the transition regime. Journal of Statistical Physics, 63(1–2):203–219, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01026599.

Holovko:1993:LTM

[HTPH93] M. F. Holovko, A. D. Trokhymchuk, I. A. Protsykevich, and Douglas J. Henderson. The Laplace transform of the MSA pair distribution functions between macroions in an ion-dipole fluid. Journal of Statistical Physics, 72(5–6):1391–1399, September 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01048192.

Hua:1997:MAI

[Hua97] Fan Ai Hua. Multifractal analysis of infinite products. Journal of Statistical Physics, 86(5–6):1313–1336, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183625.

Havlin:1990:NCL

[HW90] Shlomo Havlin and George H. Weiss. A new class of long-tailed pausing time densities for the CTRW. Journal of Statistical Physics, 58(5–6):1267–1273, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01026577.

Hasty:1997:ROD

[HW97a] Jeff Hasty and Kurt Wiesenfeld. Renormalization of onedimensional avalanche models. Journal of Statistical Physics, 86(5–6):1179–1201, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02183620.

Hattori:1997:ARW

[HW97b] Tetsuya Hattori and Hiroshi Watanabe. Anisotropic random walks and asymptotically one-dimensional diffusion on abcgaskets. Journal of Statistical Physics, 88(1–2):105–128, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02508466.

Hinrichsen:1997:AID

[HWD97] Haye Hinrichsen, Joshua S. Weitz, and Eytan Domany. An algorithm-independent definition of damage spreading — application to directed percolation. Journal of Statistical Physics, 88 (3–4):617–636, August 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000015165.83255.b7.

Hendriks:1997:GAA

[HWvB97] E. M. Hendriks, J. Walsh, and A. R. D. van Bergen. A general approach to association using cluster partition functions. Journal of Statistical Physics, 87(5–6):1287–1306, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02181285.

Higuchi:1996:IML

[HY96] Yasunari Higuchi and Nobuo Yoshida. Ising models on the lattice Sierpinski gasket. Journal of Statistical Physics, 84(1–2):295– 307, July 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179588.

Holovatch:1998:CED

[HY98] Yu. Holovatch and T. Yavors’kii. Critical exponents of the diluted Ising model between dimensions 2 and 4. Journal of Statistical Physics, 92(5–6):785–808, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023032307964.

He:1997:ASS

[HZLD97] Xiaoyi He, Qisu Zou, Li-Shi Luo, and Micah Dembo. Analytic solutions of simple flows and analysis of nonslip boundary conditions for the lattice Boltzmann BGK model. Journal of Statistical Physics, 87(1–2):115–136, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02181482.

Izus:1998:BES

[IdRB98] G. G. Izús, J. Reyes de Rueda, and C. H. Borzi. Boundary effects on the structural stability of stationary patterns in a bistable reaction–diffusion system. Journal of Statistical Physics, 90(1– 2):103–117, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023295416567.

Ivanchenko:1993:NEC

[IFR93] Yu. M. Ivanchenko, A. E. Filippov, and A. V. Radievsky. Nonlinear excitations in the critical region. Journal of Statistical Physics, 71(5–6):1003–1014, June 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01049958.

Ippen:1993:CRS

[ILD93] Erich Ippen, John Lindner, and William L. Ditto. Chaotic resonance: A simulation. Journal of Statistical Physics, 70(1–2): 437–450, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053980.

Ivanchenko:1990:SEC

[ILF90] Yu. M. Ivanchenko, A. A. Lisyanskii, and A. E. Filippov. The scale equations in the critical dynamics of fluctuating systems. Journal of Statistical Physics, 58(1–2):295–323, January 1990.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020295.

Ivanchenko:1992:NRP

[ILF92] Yu. M. Ivanchenko, A. A. Lisyansky, and A. A. Filippov. New renormalization procedure for eliminating redundant operators. Journal of Statistical Physics, 66(3–4):1139–1145, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01055721.

Ito:1996:PAR

[IM96] H. M. Ito and T. Mikami. Poissonian asymptotics of a randomly perturbed dynamical system: Flip-flop of the stochastic disk dynamo. Journal of Statistical Physics, 85(1–2):41–53, October 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02175555.

Igarashi:1992:VSN

[IMS92] A. Igarashi, P. V. E. McClintock, and N. G. Stocks. Velocity spectrum for non-Markovian Brownian motion in a periodic potential. Journal of Statistical Physics, 66(3–4):1059–1070, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01055716.

Inozemtsev:1990:CBO

[Ino90] V. I. Inozemtsev. On the connection between the one-dimensional S = 1/2 Heisenberg chain and Haldane–Shastry model. Journal of Statistical Physics, 59(5–6):1143–1155, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01334745.

Ioffe:1994:LDI

[Iof94] Dmitry Ioffe. Large deviations for the 2D Ising model: A lower bound without cluster expansions. Journal of Statistical Physics, 74(1–2):411–432, January 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02186818.

Ito:1992:SED

[IOT92] H. M. Ito, Y. Ogura, and M. Tomisaki. Stretched-exponential decay laws of general defect diffusion models. Journal of Statistical Physics, 66(1–2):563–582, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01060081.

Isola:1990:UEU

[IP90] S. Isola and A. Politi. Universal encoding for unimodal maps. Journal of Statistical Physics, 61(1–2):263–291, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013965.

Indekeu:1999:WAW

[IRB^{+}99] Joseph O. Indekeu, Karine Ragil, Daniel Bonn, Daniel Broseta, and Jacques Meunier. Wetting of alkanes on water from a Cahntype theory: Effects of long-range forces. Journal of Statistical Physics, 95(5–6):1009–1043, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004558618646.

Illner:1996:BVP

[IS96] R. Illner and J. Struckmeier. Boundary value problems for the steady Boltzmann equation. Journal of Statistical Physics, 85(3– 4):427–454, November 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02174213.

Itoh:1999:PCS

[IS99] Yoshiaki Itoh and Larry Shepp. Parking cars with spin but no length. Journal of Statistical Physics, 97(1–2):209–231, October 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004619102562.

Ishii:1995:IMJ

[Ish95] Yutaka Ishii. Ising models, Julia sets, and similarity of the maximal entropy measures. Journal of Statistical Physics, 78(3–4): 815–825, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183689.

Ishio:1996:QTC

[Ish96] H. Ishio. Quantum transport and classical dynamics in open billiards. Journal of Statistical Physics, 83(1–2):203–214, April 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02183646.

Isola:1999:RSI

[Iso99] Stefano Isola. Renewal sequences and intermittency. Journal of Statistical Physics, 97(1–2):263–280, October 1999. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004623303471.

Iochum:1991:PLG

[IT91] B. Iochum and D. Testard. Power law growth for the resistance in the Fibonacci model. Journal of Statistical Physics, 65(3–4): 715–723, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053750.

Illner:1998:VME

[IVDB98] R. Illner, H. D. Victory, P. Dukes, and A. V. Bobylev. On Vlasov–Manev equations, II: Local existence and uniqueness. Journal of Statistical Physics, 91(3–4):625–654, May 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023029711405.

Illner:1993:RDV

[IW93] Reinhard Illner and Wolfgang Wagner. A random discrete velocity model and approximation of the Boltzmann equation. Journal of Statistical Physics, 70(3–4):773–792, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01053594.

Jaffe:1991:ALC

[Jaf91] Michael Jaffe. Applications of liquid crystal polymers. Journal of Statistical Physics, 62(5–6):985–995, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01128172.

Jaggi:1991:SDD

[Jag91] N. K. Jaggi. Structure and dynamics of a dense dipolar system in an electric field and their relevance to electrorheological fluids. Journal of Statistical Physics, 64(5–6):1093–1102, September 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048816.

Jan:1994:ASP

[Jan94] Naeem Jan. Adult survival in Partridge–Barton model of biological aging. Journal of Statistical Physics, 77(3–4):915– 917, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179469.

Jancovici:1995:CCS

[Jan95] B. Jancovici. Classical Coulomb systems: Screening and correlations revisited. Journal of Statistical Physics, 80(1–2):445–459, July 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02178367.

Jarzynski:1999:MAC

[Jar99] C. Jarzynski. Microscopic analysis of Clausius–Duhem processes. Journal of Statistical Physics, 96(1–2):415–427, July 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1004541004050.

Jensen:1993:TDP

[JD93] Iwan Jensen and Ronald Dickman. Time-dependent perturbation theory for nonequilibrium lattice models. Journal of Statistical Physics, 71(1–2):89–127, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01048090.

Jeon:1999:SCC

[Jeo99] Intae Jeon. Spouge’s conjecture on complete and instantaneous gelation. Journal of Statistical Physics, 96(5–6):1049– 1070, September 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004640317274.

Jerrum:1990:TDM

[Jer90] Mark Jerrum. Two-dimensional monomer-dimer systems are computationally intractable. Journal of Statistical Physics, 59 (3–4):1087–1088, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01025864.

Jesudason:1996:NMC

[Jes96] Christopher G. Jesudason. Note on Monte Carlo methods without necessary detailed balance. Journal of Statistical Physics, 82(3–4):1207–1211, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179809.

Jezewski:1996:CPM

[Jez96] W. Jezewski. Chaotic properties of multipoint correlation functions of an Ising model with long-range interactions on the Sierpiński-gasket lattice. Journal of Statistical Physics, 82(3–4): 1099–1112, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179804.

Jackie:1991:SDS

[JFK91a] J. Jäckie, K. Froböse, and D. Knödler. Size dependence of selfdiffusion in the hard-square lattice gas. Journal of Statistical Physics, 63(1–2):249–260, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01026602.

Jackle:1991:SDS

[JFK91b] J. Jäckle, K. Froböse, and D. Knödler. Size dependence of self-diffusion in the hard-square lattice gas. Journal of Statistical Physics, 65(1–2):415–416, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01329871.

Jones:1998:CLP

[JG98] Ronald E. Jones and Gemunu H. Gunaratne. Characterization of labyrinthine patterns and their evolution. Journal of Statistical Physics, 93(3–4):427–447, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033157.89511.a2.

Jiang:1999:EFS

[Jia99] Miaohua Jiang. The entropy formula for SRB–Measures of lattice dynamical systems. Journal of Statistical Physics, 95(3–4):791– 803, May 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004563930516.

Jitomirskaya:1993:IMQ

[JK93] Svetlana Jitomirskaya and Abel Klein. Ising model in a quasiperiodic transverse field, percolation, and contact processes in quasiperiodic environments. Journal of Statistical Physics, 73(1– 2):319–344, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01052763.

Jacobsen:1999:TCD

[JK99] Jesper Lykke Jacobsen and Jané Kondev. Transition from the compact to the dense phase of two-dimensional polymers. Journal of Statistical Physics, 96(1–2):21–48, July 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004512230458.

Janowsky:1994:ERA

[JL94] S. A. Janowsky and J. L. Lebowitz. Exact results for the asymmetric simple exclusion process with a blockage. Journal of Statistical Physics, 77(1–2):35–51, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02186831.

Jansons:1998:SCA

[JL98] Kalvis M. Jansons and G. D. Lythe. Stochastic calculus: Application to dynamic bifurcations and threshold crossings. Journal of Statistical Physics, 90(1–2):227–251, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023207919293.

Jancovici:1993:LCF

[JLM93] B. Jancovici, J. L. Lebowitz, and G. Manificat. Large charge fluctuations in classical Coulomb systems. Journal of Statistical Physics, 72(3–4):773–787, August 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01048032.

Jancovici:1995:ETD

[JLM95] B. Jancovici, J. L. Lebowitz, and Ph. A. Martin. Erratum: Timedependent correlations in an inhomogeneous one-component plasma. Journal of Statistical Physics, 79(3–4):789, May 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02184885.

Jona-Lasinio:1996:SSD

[JLS96] G. Jona-Lasinio and R. Sénéor. Study of stochastic differential equations by constructive methods. I. Journal of Statistical Physics, 83(5–6):1109–1148, June 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179554.

Jacobs:1990:UIS

[JM90] A. E. Jacobs and David Mukamel. Universal incommensurate structures. Journal of Statistical Physics, 58(3–4):503–510, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01112759.

Jancovici:1992:CCF

[JM92] B. Jancovici and G. Manificat. Classical Coulomb fluids in a confined geometry. Journal of Statistical Physics, 68(5–6):1089– 1103, September 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048886.

Jiang:1996:UED

[JM96] Miaohua Jiang and Alexander E. Mazel. Uniqueness and exponential decay of correlations for some two-dimensional spin lattice systems. Journal of Statistical Physics, 82(3–4):797–821, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179793.

Jancovici:1987:TDC

[JMM87] B. Jancovici, N. Macris, and Ph. A. Martin. Time-dependent correlations for a one-component plasma in a uniform magnetic field. Journal of Statistical Physics, 47(1–2):229–256, April 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF01009044. See erratum [JLM95].

Jancovici:1994:CSS

[JMP94] B. Jancovici, G. Manificat, and C. Pisani. Coulomb systems seen as critical systems: Finite-size effects in two dimensions. Journal of Statistical Physics, 76(1–2):307–329, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02188664.

Johnson:1990:BSM

[Joh90] Karen E. Johnson. Bringing statistical mechanics into chemistry: The early scientific work of Karl F. Herzfeld. Journal of Statistical Physics, 59(5–6):1547–1572, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01334763.

Jose:1998:DTC

[Jos98] Jorge V. José. Duality in two capacitively coupled layered arrays of ultrasmall Josephson junctions. Journal of Statistical Physics, 93(3–4):943–964, November 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033171.04828.d6.

Joyce:1990:ABM

[Joy90] G. S. Joyce. Asymptotic behavior of Mayer cluster sums for the one-dimensional Ising model. Journal of Statistical Physics, 58(3–4):443–465, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01112755.

Joye:1994:AAC

[Joy94] Alain Joye. Absence of absolutely continuous spectrum of Floquet operators. Journal of Statistical Physics, 75(5–6):929–952, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186751.

Joye:1996:UBE

[Joy96] Alain Joye. Upper bounds for the energy expectation in timedependent quantum mechanics. Journal of Statistical Physics, 85(5–6):575–606, December 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02199357.

Jansons:1991:PTP

[JR91a] Kalvis M. Jansons and L. C. G. Rogers. Probability theory and polymer physics. Journal of Statistical Physics, 65(1–2):139–165, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01329853.

Joets:1991:LBD

[JR91b] A. Joets and R. Ribotta. Localized bifurcations and defect instabilities in the convection of a nematic liquid crystal. Journal of Statistical Physics, 64(5–6):981–1005, September 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01048809.

Jan:1994:DLT

[JR94] Naeem Jan and Tane S. Ray. ‘Damage’ in the low-temperature phase of the ±J spin glass in two to six dimensions. Journal of Statistical Physics, 75(5–6):1197–1204, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02186765.

Janke:1995:AMM

[JS95] Wolfhard Janke and Tilman Sauer. Application of the multicanonical multigrid Monte Carlo method to the two-dimensional ϕ^{4}: Autocorrelations and interface tension: Autocorrelations and interface tension. Journal of Statistical Physics, 78(3–4):759–798, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183687.

Janssen:1999:BDC

[JS99] H. K. Janssen and B. Schmittmann. Biased diffusion with correlated noise. Journal of Statistical Physics, 95(3–4):569–586, May 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004562408226.

Jan:1998:INE

[JSA98] Naeem Jan, Dietrich Stauffer, and Amnon Aharony. An infinite number of effectively infinite clusters in critical percolation. Journal of Statistical Physics, 92(1–2):325–330, July 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023008021962.

Jackle:1991:FND

[JSC91] J. Jäckle, R. B. Stinchcombe, and S. Cornell. Freezing of nonequilibrium domain structures in a kinetic Ising model. Journal of Statistical Physics, 62(1–2):425–433, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020878.

Jimenez:1999:SSD

[JSO99] J. C. Jimenez, I. Shoji, and T. Ozaki. Simulation of stochastic differential equations through the local linearization method. A comparative study. Journal of Statistical Physics, 94(3–4): 587–602, February 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004504506041.

Jancovici:1996:CSS

[JT96] B. Jancovici and G. Téllez. Coulomb systems seen as critical systems: Ideal conductor boundaries. Journal of Statistical Physics, 82(3–4):609–632, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179788.

Jordan:1997:IMT

[JT97] Richard Jordan and Bruce Turkington. Ideal magnetofluid turbulence in two dimensions. Journal of Statistical Physics, 87 (3–4):661–695, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181242.

Jancovici:1998:TDC

[JT98] B. Jancovici and G. Téllez. Two-dimensional Coulomb systems on a surface of constant negative curvature. Journal of Statistical Physics, 91(5–6):953–977, June 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023079916489.

Junk:1998:DDL

[Jun98] Michael Junk. Domain of definition of Levermore’s five-moment system. Journal of Statistical Physics, 93(5–6):1143–1167, December 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000033155.07331.d9.

Jurlewicz:1995:FIR

[Jur95] A. Jurlewicz. Frequency-independent rules for the dielectric susceptibility derived from two forms of self-similar dynamical behavior of dipolar systems. Journal of Statistical Physics, 79(5– 6):993–1003, June 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181212.

Just:1992:POA

[Jus92] Wolfram Just. Projection operator approach to the thermodynamic formalism of dynamical systems. Journal of Statistical Physics, 67(1–2):271–287, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01049034.

Just:1995:BGC

[Jus95] Wolfram Just. Bifurcations in globally coupled map lattices. Journal of Statistical Physics, 79(1–2):429–449, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179397.

Just:1998:AAP

[Jus98] Wolfram Just. Analytical approach for piecewise linear coupled map lattices. Journal of Statistical Physics, 90(3–4):727–748, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023272819435.

Johnson:1999:ISI

[JV99] William C. Johnson and P. W. Voorhees. Interfacial stress, interfacial energy, and phase equilibria in binary alloys. Journal of Statistical Physics, 95(5–6):1281–1309, June 1999. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004527106351.

Janosi:1998:THG

[JVH98] I. M. Jánosi, G. Vattay, and A. Harnos. Turbulent helium gas cell as a new paradigm of daily meteorological fluctuations? Journal of Statistical Physics, 93(3–4):919–926, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033169.16298.0a.

Jurlewicz:1993:RBA

[JW93] A. Jurlewicz and K. Weron. A relationship between asymmetric Lévy-stable distributions and the dielectric susceptibility. Journal of Statistical Physics, 73(1–2):69–81, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01052751.

Janicki:1997:HDR

[JW97] A. W. Janicki and W. A. Woyczynski. Hausdorff dimension of regular points in stochastic Burgers flows with Lévy α-stable initial data. Journal of Statistical Physics, 86(1–2):277–299, January 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02180207.

Jonsson:1998:ADD

[JW98] Thordur Jonsson and John F. Wheater. Area distribution for directed random walks. Journal of Statistical Physics, 92(3– 4):713–725, August 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023053024552.

Kutlu:1994:CCE

[KA94] B. Kutlu and N. Aktekin. Computation of critical exponents for two-dimensional Ising model on a cellular automaton. Journal of Statistical Physics, 75(3–4):757–763, May 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02186880.

Kalos:1991:EMC

[Kal91] M. H. Kalos. Exact Monte Carlo for few-fermion systems. Journal of Statistical Physics, 63(5–6):1269–1281, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01030010.

Karner:1994:SFA

[Kar94] Gunther Karner. The simplified Fermi accelerator in classical and quantum mechanics. Journal of Statistical Physics, 77(3–4): 867–879, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179464.

Karapiperis:1995:CAM

[Kar95] T. Karapiperis. Cellular automaton model of precipitation/ dissolution coupled with solute transport. Journal of Statistical Physics, 81(1–2):165–180, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179974.

Karner:1999:QCE

[Kar99] Gunther Karner. Quantum and classical evolutions of a nonautonomous dynamical system — a comparison. Journal of Statistical Physics, 96(3–4):627–637, August 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004598223382.

Kaufman:1995:LLO

[Kau95] Miss Bruria Kaufman. Letter from Lars Onsager to Bruria Kaufman. Journal of Statistical Physics, 78(1–2):585–588, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183367.

Kawashima:1996:CAA

[Kaw96] Naoki Kawashima. Cluster algorithms for anisotropic quantum spin models. Journal of Statistical Physics, 82(1–2):131–153, January 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02189228.

Kawasaki:1997:GIM

[Kaw97] Kyozi Kawasaki. Generalized irreducible memory function. Journal of Statistical Physics, 87(5–6):981–988, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02181266.

Kawasaki:1998:MAD

[Kaw98] Kyozi Kawasaki. Microscopic analyses of the dynamical density functional equation of dense fluids. Journal of Statistical Physics, 93(3–4):527–546, November 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033240.66359.6c.

Koscielny-Bunde:1990:EDN

[KB90] Eva Koscielny-Bunde. Effect of damage in neural networks. Journal of Statistical Physics, 58(5–6):1257–1266, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01026576.

Kutner:1991:DFD

[KB91] R. Kutner and T. Barszczak. Distribution for fermionic discrete lattice gas within the canonical ensemble. Journal of Statistical Physics, 65(3–4):813–821, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053757.

Kirkpatrick:1997:LRC

[KB97a] T. R. Kirkpatrick and D. Belitz. Long-range correlations and generic scale invariance in classical fluids and disordered electron systems. Journal of Statistical Physics, 87(5–6):1307–1323, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02181286.

Kurten:1997:IKM

[KB97b] Karl E. Kürten and Helga Beer. Inhomogeneous Kauffman models at the borderline between order and chaos. Journal of Statistical Physics, 87(3–4):929–935, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02181253.

Kuscer:1997:DZF

[KB97c] Ivan Kuscer and Jan J. M. Beenakker. Diffusion in zeolites as flow of lattice gas. Journal of Statistical Physics, 87(5–6):1083–1095, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02181272.

Kong:1991:DPT

[KC91a] X. P. Kong and E. G. D. Cohen. Diffusion and propagation in triangular Lorentz lattice gas cellular automata. Journal of Statistical Physics, 62(3–4):737–757, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01017981.

Kong:1991:LLG

[KC91b] X. P. Kong and E. G. D. Cohen. Lorentz lattice gases, abnormal diffusion, and polymer statistics. Journal of Statistical Physics, 62(5–6):1153–1171, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01128181.

Kandel:1990:RDD

[KD90] Daniel Kandel and Eytan Domany. Rigorous derivation of domain growth kinetics without conservation laws. Journal of Statistical Physics, 58(3–4):685–706, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01112771.

Koduvely:1998:MSI

[KD98] Hari M. Koduvely and Deepak Dhar. A model of subdiffusive interface dynamics with a local conservation of minimum height. Journal of Statistical Physics, 90(1–2):57–77, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1023291315658.

Keskin:1997:DBS

[KE97] Mustafa Keskin and Riza Erdem. Dynamic behavior of a spin-1 Ising model. I. Relaxation of order parameters and the ‘flatness’ property of metastable states. Journal of Statistical Physics, 89(5–6):1035–1046, December 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02764220.

Kedem:1993:TSP

[Ked93] Rinat Kedem. Thermodynamics of the 3-state Potts spin chain. Journal of Statistical Physics, 71(5–6):903–921, June 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01049954.

Keisling:1998:CSS

[Kei98] John D. Keisling. Convergence speed for simple symmetric exclusion: An explicit calculation. Journal of Statistical Physics, 90(3–4):1003–1013, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023297524887.

Kennedy:1990:FPE

[Ken90] Tom Kennedy. A fixed-point equation for the high-temperature phase of discrete lattice spin systems. Journal of Statistical Physics, 59(1–2):195–220, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01015568.

Kennedy:1993:SRR

[Ken93] Tom Kennedy. Some rigorous results on majority rule renormalization group transformations near the critical point. Journal of Statistical Physics, 72(1–2):15–37, July 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048038.

Kennedy:1994:BBW

[Ken94] Tom Kennedy. Ballistic behavior in a 1D weakly self-avoiding walk with decaying energy penalty. Journal of Statistical Physics, 77(3–4):565–579, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179450.

Kennedy:1997:MRL

[Ken97] Tom Kennedy. Majority rule at low temperatures on the square and triangular lattices. Journal of Statistical Physics, 86(5–6): 1089–1107, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183615.

Kennedy:1998:PSN

[Ken98] Tom Kennedy. Phase separation in the neutral Falicov–Kimball model. Journal of Statistical Physics, 91(5–6):829–843, June 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1023023230602.

Kerimov:1993:APT

[Ker93] Azer Kerimov. Absence of phase transitions in one-dimensional antiferromagnetic models with long-range interactions. Journal of Statistical Physics, 72(3–4):571–620, August 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01048025.

Ko:1990:SBI

[KF90] Lee-Fen Ko and Michael E. Fisher. The shapes of bowed interfaces in the two-dimensional Ising model. Journal of Statistical Physics, 58(1–2):249–264, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020293.

Kapral:1993:DOP

[KF93] Raymond Kapral and Simon J. Fraser. Dynamics of oscillators with periodic dichotomous noise. Journal of Statistical Physics, 70(1–2):61–76, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053954.

Kimball:1997:VSF

[KF97] John C. Kimball and Harry L. Frisch. Vibrations of simple fractal-based models. Journal of Statistical Physics, 89(1–2): 453–468, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02770776.

Kim:1991:TIS

[KFK91] Sangtae Kim, Yuris O. Fuentes, and Seppo J. Karrila. Towards ab initio simulations of concentrated suspensions. Journal of Statistical Physics, 62(5–6):1197–1223, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01128183.

Klik:1990:FPT

[KG90] Ivo Klik and Leon Gunther. First-passage-time approach to overbarrier relaxation of magnetization. Journal of Statistical Physics, 60(3–4):473–484, August 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01314931.

Kawashima:1995:GFK

[KG95] N. Kawashima and J. E. Gubernatis. Generalization of the Fortuin–Kasteleyn transformation and its application to quantum spin simulations. Journal of Statistical Physics, 80(1– 2):169–221, July 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02178358.

Kobayashi:1999:ESD

[KG99] R. Kobayashi and Y. Giga. Equations with singular diffusivity. Journal of Statistical Physics, 95(5–6):1187–1220, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004570921372.

Keskin:1992:ECC

[KGM92] Mustafa Keskin, Mustafa Gençaslan, and Paul H. E. Meijer. Evaluation and comparison of critical lines for various models of gas-liquid binary systems. Journal of Statistical Physics, 66(3– 4):885–896, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01055707.

Kiss:1993:NSS

[KGM^{+}93] László B. Kiss, Zoltán Gingl, Zsuzsanna Márton, János Kertész, Frank Moss, Gabor Schmera, and Adi Bulsara. 1/f noise in systems showing stochastic resonance. Journal of Statistical Physics, 70(1–2):451–462, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053981.

Kum:1994:TRC

[KH94] Oyeon Kum and William G. Hoover. Time-reversible continuum mechanics. Journal of Statistical Physics, 76(3–4):1075–1081, August 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02188699.

Kacperski:1996:PTH

[KH96] K. Kacperski and J. A. Holyst. Phase transitions and hysteresis in a cellular automata-based model of opinion formation. Journal of Statistical Physics, 84(1–2):169–189, July 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179581.

Kholodenko:1990:ORF

[Kho90] A. L. Kholodenko. Onsager’s reaction field for the Potts model from the path integral. Journal of Statistical Physics, 58(1–2): 355–370, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020297.

Kholodenko:1991:PMD

[Kho91a] Arkady L. Kholodenko. Potts model, Dirac propagator, and conformational statistics of semiflexible polymers. Journal of Statistical Physics, 65(1–2):291–316, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01329862.

Khoruzhenko:1991:LLH

[Kho91b] B. A. Khoruzhenko. Large-n limit of the Heisenberg model: Random external field and random uniaxial anisotropy. Journal of Statistical Physics, 62(1–2):21–33, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020857.

Kiessling:1990:CTL

[Kie90] Michael K.-H. Kiessling. A complementary thermodynamic limit for classical Coulomb matter. Journal of Statistical Physics, 59 (5–6):1157–1186, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01334746.

Kiessling:1992:UUB

[Kie92] Michael K.-H. Kiessling. Uniform upper bounds (and thermodynamic limit) for the correlation functions of symmetric Coulombtype systems. Journal of Statistical Physics, 66(5–6):1359–1382, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01054426.

Kikuchi:1999:CDF

[Kik99] Ryoichi Kikuchi. Continuous displacement formulation of alloys. Journal of Statistical Physics, 95(5–6):1323–1336, June 1999.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004531307259.

Katori:1997:HSS

[KIKK97] M. Katori, N. Inui, G. Komatsu, and K. Kameoka. Hypergeometric series in a series expansion of the directed-bond percolation probability on the square lattice. Journal of Statistical Physics, 86(1–2):37–55, January 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02180198.

King:1999:TDP

[Kin99] C. King. Two-dimensional Potts model and annular partitions. Journal of Statistical Physics, 96(5–6):1071–1089, September 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1004692301344.

Katori:1991:UBS

[KK91a] Makoto Katori and Norio Konno. Upper bounds for survival probability of the contact process. Journal of Statistical Physics, 63(1–2):115–130, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01026595.

Konno:1991:AOP

[KK91b] Norio Konno and Makoto Katori. Analysis of the order parameter for uniform nearest particle system. Journal of Statistical Physics, 65(1–2):247–254, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01329859.

Koo:1991:STR

[KK91c] Yong-Eun Lee Koo and Raoul Kopelman. Space-and timeresolved diffusion-limited binary reaction kinetics in capillaries: experimental observation of segregation, anomalous exponents, and depletion zone. Journal of Statistical Physics, 65(5– 6):893–918, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049588.

Kawasaki:1992:PFG

[KK92] Kyozi Kawasaki and Toshihiro Kawakatsu. Projector formalism of generalized Brownian motion theory applied to dissipative and noisy systems. Journal of Statistical Physics, 67(3–4):795–811, May 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049727.

Kurten:1993:HDA

[KK93a] Karl E. Kürten and Norbert Klingen. Hadamard design and artificial neural nets. Journal of Statistical Physics, 71(1–2):327–339, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048103.

Kuzovkov:1993:ENC

[KK93b] V. Kuzovkov and E. Kotomin. Effect of nonequilibrium charge screening in A + B → 0 bimolecular reactions in condensed matter. Journal of Statistical Physics, 72(1–2):127–144, July 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01048043.

Ketoja:1994:BTA

[KK94a] Jukka A. Ketoja and Juhani Kurkijärvi. Binary tree approach to scaling in unimodal maps. Journal of Statistical Physics, 75 (3–4):643–668, May 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186875.

Kumar:1994:MCS

[KK94b] Susmit Kumar and Stewart K. Kurtz. A Monte Carlo study of size and angular properties of a three-dimensional Poisson– Delaunay cell. Journal of Statistical Physics, 75(3–4):735–748, May 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186878.

Kratky:1998:IFU

[KK98] Karl W. Kratky and Karl E. Kürten. Irregular fluctuations in uncoupled map lattices. Journal of Statistical Physics, 90(3–4): 749–765, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023224903505.

Kumar:1992:PTD

[KKBS92] Susmit Kumar, Stewart K. Kurtz, Jayanth R. Banavar, and M. G. Sharma. Properties of a three-dimensional Poisson– Voronoi tesselation: A Monte Carlo study. Journal of Statistical Physics, 67(3–4):523–551, May 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01049719.

Kumar:1993:PTD

[KKBS93] S. Kumar, S. K. Kurtz, J. R. Banavar, and M. G. Sharma. Properties of a three-dimensional Poisson–Voronoi tesselation: A Monte Carlo study. Journal of Statistical Physics, 71(1–2):349, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048105.

Kotrla:1991:KSV

[KL91] M. Kotrla and A. C. Levi. Kinetic six-vertex model as model of bcc crystal growth. Journal of Statistical Physics, 64(3–4): 579–604, August 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048307.

Klik:1991:MSD

[Kli91] Ivo Klik. Metastable systems driven by colored noise: The stationary state. Journal of Statistical Physics, 63(1–2):389–397, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01026611.

Klik:1992:RMU

[Kli92] Ivo Klik. Rotation of magnetization in unison and Langevin equations for a large spin. Journal of Statistical Physics, 66(1– 2):635–645, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01060085.

Kotecky:1990:SPM

[KLMR90] Roman Kotecký, Lahoussine Laanait, Alain Messager, and Jean Ruiz. The q-state Potts model in the standard Pirogov–Sinai theory: Surface tensions and Wilson loops. Journal of Statistical Physics, 58(1–2):199–248, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020292.

Klosek:1995:HRE

[Kło95] M. M. Kłosek. Half-range expansion analysis for Langevin dynamics in the high-friction limit with a singular absorbing boundary condition: Noncharacteristic case. Journal of Statistical Physics, 79(1–2):313–345, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179392.

Klopp:1998:BEB

[Klo98] Frédéric Klopp. Band edge behavior of the integrated density of states of random Jacobi matrices in dimension 1. Journal of Statistical Physics, 90(3–4):927–947, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023293423978.

Kantz:1994:ETC

[KLR94] Holger Kantz, Roberto Livi, and Stefano Ruffo. Equipartition thresholds in chains of anharmonic oscillators. Journal of Statistical Physics, 76(1–2):627–643, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02188678.

Kessler:1997:ESL

[KLRT97] David A. Kessler, Herbert Levine, Douglas Ridgway, and Lev Tsimring. Evolution on a smooth landscape. Journal of Statistical Physics, 87(3–4):519–544, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02181235.

Karevski:1997:CET

[KLT97] Dragi Karevski, Peter Lajkó, and Loïc Turban. Corner exponents in the two-dimensional Potts model. Journal of Statistical Physics, 86(5–6):1153–1162, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183618.

Keller:1991:FRL

[KM91a] Joseph B. Keller and Gregory J. Merchant. Flexural rigidity of a liquid surface. Journal of Statistical Physics, 63(5–6):1039–1051, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01029998.

Kim:1991:EFN

[KM91b] Bongsoo Kim and Gene F. Mazenko. Equations of fluctuating nonlinear hydrodynamics for normal fluids. Journal of Statistical Physics, 64(3–4):631–652, August 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01048309.

Kedem:1993:CMB

[KM93] Rinat Kedem and Barry M. McCoy. Construction of modular branching functions from Bethe’s equations in the 3-state Potts chain. Journal of Statistical Physics, 71(5–6):865–901, June 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049953.

Koma:1996:FSS

[KM96] Tohru Koma and Nobuyuki Mizukoshi. Finite-size scaling for correlations of quantum spin chains at criticality. Journal of Statistical Physics, 83(3–4):661–726, May 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183744.

Kondratiev:1997:OPS

[KM97] Yu. G. Kondratiev and R. A. Minlos. One-particle subspaces in the stochastic XY model. Journal of Statistical Physics, 87 (3–4):613–642, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181240.

Kehrein:1998:DSP

[KM98] Stefan K. Kehrein and Andreas Mielke. Diagonalization of system plus environment Hamiltonians. Journal of Statistical Physics, 90(3–4):889–898, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023289323069.

Koiller:1996:STD

[KMKdC96] Jair Koiller, Roberto Markarian, Sylvie Oliffson Kamphorst, and Sônia Pinto de Carvalho. Static and time-dependent perturbations of the classical elliptical billiard. Journal of Statistical Physics, 83(1–2):127–143, April 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02183642.

Kamien:1993:DPM

[KN93] Randall D. Kamien and David R. Nelson. Directed polymer melts and quantum critical phenomena. Journal of Statistical Physics, 71(1–2):23–50, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048086.

Knauf:1993:PNT

[Kna93] Andreas Knauf. Phases of the number-theoretic spin chain. Journal of Statistical Physics, 73(1–2):423–431, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01052771.

Kira:1998:PVC

[KNS98] Elisabeti Kira, E. Jordão Neves, and Roberto H. Schonmann. Percolation in a Voronoi competition–growth model. Journal of Statistical Physics, 92(5–6):755–764, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023028207056.

Kiskis:1993:HDR

[KNV93] Joe Kiskis, Rajamani Narayanan, and Pavlos Vranas. The Hausdorff dimension of random walks and the correlation length critical exponent in Euclidean field theory. Journal of Statistical Physics, 73(3–4):765–774, November 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01054349.

Kotecky:1993:DDA

[KO93] R. Kotecký and E. Olivieri. Droplet dynamics for asymmetric Ising model. Journal of Statistical Physics, 70(5–6):1121–1148, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049425.

Kotecky:1994:SGD

[KO94] R. Kotecký and E. Olivieri. Shapes of growing droplets — a model of escape from a metastable phase. Journal of Statistical Physics, 75(3–4):409–506, May 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02186869.

Kobe:1997:EIP

[Kob97] S. Kobe. Ernst Ising — physicist and teacher. Journal of Statistical Physics, 88(3–4):991–995, August 1997. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000015184.19421.03.

Kohring:1990:HPS

[Koh90] G. A. Kohring. A high-precision study of the Hopfield model in the phase of broken replica symmetry. Journal of Statistical Physics, 59(3–4):1077–1086, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01025863.

Kohring:1991:CPP

[Koh91a] G. A. Kohring. Calculation of the permeability of porous media using hydrodynamic cellular automata. Journal of Statistical Physics, 63(1–2):411–418, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01026614.

Kohring:1991:FSN

[Koh91b] G. A. Kohring. Finite-state neural networks. A step toward the simulation of very large systems. Journal of Statistical Physics, 62(3–4):563–576, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01017973.

Kohyama:1991:CGP

[Koh91c] Tamotsu Kohyama. Cluster growth in particle-conserving cellular automata. Journal of Statistical Physics, 63(3–4):637–651, May 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01029203.

Kohmoto:1992:DSR

[Koh92a] Mahito Kohmoto. Dynamical system related to quasiperiodic Schrödinger equations in one dimension. Journal of Statistical Physics, 66(3–4):791–796, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01055701.

Kohring:1992:EHC

[Koh92b] G. A. Kohring. An efficient hydrodynamic cellular automata for simulating fluids with large viscosities. Journal of Statistical Physics, 66(3–4):1177–1184, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01055725.

Kojima:1997:GSC

[Koj97] Takeo Kojima. Ground-state correlation functions for an impenetrable Bose gas with Neumann or Dirichlet boundary conditions. Journal of Statistical Physics, 88(3–4):713–743, August 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000015169.89162.d9.

Kockelkoren:1998:CBS

[KOJ98] Julien Kockelkoren, Fridolin Okkels, and Mogens H. Jensen. Chaotic behavior in Shell models and Shell maps. Journal of Statistical Physics, 93(3–4):833–842, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033165.51531.8c.

Koma:1993:NMC

[Kom93] Tohru Koma. A new Monte Carlo power method for the eigenvalue problem of transfer matrices. Journal of Statistical Physics, 71(1–2):269–297, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048100.

Koo:1995:OPT

[Koo95] W. M. Koo. Oriented polymers: A transfer matrix calculation. Journal of Statistical Physics, 81(3–4):561–578, November 1995.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179248.

Kotulski:1995:ADC

[Kot95] Marcin Kotulski. Asymptotic distributions of continuous-time random walks: A probabilistic approach. Journal of Statistical Physics, 81(3–4):777–792, November 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179257.

Kobryn:1998:MGE

[KOT98] A. E. Kobryn, I. P. Omelyan, and M. V. Tokarchuk. The modified group expansions for construction of solutions to the BBGKY hierarchy. Journal of Statistical Physics, 92(5–6):973–994, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023044610690.

Koukiou:1990:RCI

[Kou90] F. Koukiou. A random covering interpretation for the phase transition of the random energy model. Journal of Statistical Physics, 60(5–6):669–674, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01025988.

Koza:1994:DLR

[Koz94] Zbigniew Koza. The A + B → 0 diffusion-limited reaction with correlated initial condition. Journal of Statistical Physics, 76(3– 4):857–866, August 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02188689.

Koza:1996:LTB

[Koz96] Zbigniew Koza. The long-time behavior of initially separated A + B → 0 reaction-diffusion systems with arbitrary diffusion constants. Journal of Statistical Physics, 85(1–2):179–191, October 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02175561.

Kozitsky:1997:HFV

[Koz97] Yuri V. Kozitsky. Hierarchical ferromagnetic vector spin model possessing the Lee–Yang property. Thermodynamic limit at the critical point and above. Journal of Statistical Physics, 87(3– 4):799–820, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181245.

Klumper:1991:ACS

[KP91a] Andreas Klümper and Paul A. Pearce. Analytic calculation of scaling dimensions: Tricritical hard squares and critical hard hexagons. Journal of Statistical Physics, 64(1–2):13–76, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01057867.

Koukiou:1991:PPP

[KP91b] F. Koukiou and P. Picco. Poisson point processes, cascades, and random coverings of R^{n}. Journal of Statistical Physics, 62(1– 2):481–489, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020884.

Kloeden:1992:HOI

[KP92] P. E. Kloeden and E. Platen. Higher-order implicit strong numerical schemes for stochastic differential equations. Journal of Statistical Physics, 66(1–2):283–314, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01060070.

Kotecky:1994:ESC

[KP94a] R. Kotecký and C.-E. Pfister. Equilibrium shapes of crystals attached to walls. Journal of Statistical Physics, 76(1–2):419– 445, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02188669.

Kreer:1994:PDS

[KP94b] Markus Kreer and Oliver Penrose. Proof of dynamical scaling in Smoluchowski’s coagulation equation with constant kernel. Journal of Statistical Physics, 75(3–4):389–407, May 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02186868.

Kiessling:1995:NVW

[KP95] M. K.-H. Kiessling and J. K. Percus. Nonuniform van der Waals theory. Journal of Statistical Physics, 78(5–6):1337–1376, March 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02180135.

Karttunen:1998:NGS

[KPANG98] Mikko Karttunen, Nikolas Provatas, Tapio Ala-Nissila, and Martin Grant. Nucleation, growth, and scaling in slow combustion. Journal of Statistical Physics, 90(5–6):1401–1411, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023243831128.

Kirsch:1998:AIL

[KPS98] W. Kirsch, L. A. Pastur, and H. Stork. Asymptotics of the interband light absorption coefficient near the band edge for an alloy-type model. Journal of Statistical Physics, 92(5–6):1173– 1191, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023061114325.

Krebs:1995:FSSb

[KPSW95] Klaus Krebs, Markus P. Pfannmüller, Horatiu Simon, and Birgit Wehefritz. Finite-size scaling studies of one-dimensional reactiondiffusion systems. Part II. Numerical methods. Journal of Statistical Physics, 78(5–6):1471–1491, March 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02180139.

Knapp:1991:TWN

[KPW91] R. Knapp, G. Papanicolaou, and B. White. Transmission of waves by a nonlinear random medium. Journal of Statistical Physics, 63(3–4):567–583, May 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01029200.

Krebs:1995:FSSa

[KPWH95] Klaus Krebs, Markus P. Pfannmüller, Birgit Wehefritz, and Haye Hinrichsen. Finite-size scaling studies of one-dimensional reaction-diffusion systems. Part I. Analytical results. Journal of Statistical Physics, 78(5–6):1429–1470, March 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02180138.

Kierlik:1992:CFA

[KR92] E. Kierlik and M. L. Rosinberg. The classical fluid of associating hard rods in an external field. Journal of Statistical Physics, 68(5–6):1037–1063, September 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01048884.

Krapivsky:1992:KRS

[Kra92] P. L. Krapivsky. Kinetics of random sequential parking on a line. Journal of Statistical Physics, 69(1–2):135–150, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01053786.

Krapivsky:1994:KMB

[Kra94] P. L. Krapivsky. Kinetic models of a binary alloy at zero temperature. Journal of Statistical Physics, 74(5–6):1211–1225, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02188224.

Kramli:1996:DCO

[Krá96] András Krámli. Decay of correlations in one and two dimensions. Journal of Statistical Physics, 83(1–2):167–191, April 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02183644.

Kral:1997:GGE

[Krá97] Petr Král. Generalized gradient expansions in quantum transport equations. Journal of Statistical Physics, 86(5–6):1337–1358, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183626.

Kenyon:1996:ANM

[KRS96] Claire Kenyon, Dana Randall, and Alistair Sinclair. Approximating the number of monomer-dimer coverings of a lattice. Journal of Statistical Physics, 83(3–4):637–659, May 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183743.

Kierlik:1997:SCO

[KRT97] E. Kierlik, M. L. Rosinberg, and G. Tarjus. A self-consistent Ornstein–Zernike approximation for the site-diluted Ising model. Journal of Statistical Physics, 89(1–2):215–232, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02770762.

Kierlik:1999:SCO

[KRT99] E. Kierlik, M. L. Rosinberg, and G. Tarjus. A self-consistent Ornstein–Zernike approximation for the random field Ising model. Journal of Statistical Physics, 94(5–6):805–836, March 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1004526931714.

Krug:1992:LSS

[Kru92] Joachim Krug. Landslides on sandpiles: Some moment relations in one dimension. Journal of Statistical Physics, 66(5–6): 1635–1641, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01054439.

Krug:1997:SWC

[Kru97] Joachim Krug. On the shape of wedding cakes. Journal of Statistical Physics, 87(3–4):505–518, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02181234.

Kougias:1990:SIR

[KS90] Ch. F. Kougias and J. Schulte. Simulating the immune response to the HIV-1 virus with cellular automata. Journal of Statistical Physics, 60(1–2):263–273, July 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01013677.

Knackstedt:1992:UGT

[KS92] Mark A. Knackstedt and Muhammad Sahimi. On the universality of geometrical and transport exponents of rigidity percolation. Journal of Statistical Physics, 69(3–4):887–895, November 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01050440.

Karasova:1993:ISL

[KS93a] Iva Karasova and Anton Surda. Incommensurate structure in the lattice-gas ANNNI model. Journal of Statistical Physics, 70(3– 4):675–689, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053589.

Kaufman:1993:DBP

[KS93b] Marcelle Kaufman and Dietrich Stauffer. A dilute bootstrap percolation: Lattice model for unresponsiveness in T-cell immunology. Journal of Statistical Physics, 73(5–6):843–851, December 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01052812.

Kolesik:1993:SEC

[KS93c] M. Kolesík and L. Samaj. Series expansion and CAM study of the nonuniversal behavior of the symmetric 16-vertex model. Journal of Statistical Physics, 72(5–6):1203–1226, September 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01048186.

Kalos:1997:MFM

[KS97a] M. H. Kalos and K. E. Schmidt. Model fermion Monte Carlo with correlated pairs. II. Journal of Statistical Physics, 89(1– 2):425–443, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02770774.

Katsoulakis:1997:SIM

[KS97b] M. A. Katsoulakis and P. E. Souganidis. Stochastic Ising models and anisotropic front propagation. Journal of Statistical Physics, 87(1–2):63–89, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181480.

Kumar:1997:CBT

[KS97c] Sanjay Kumar and Yashwant Singh. Critical behavior of two interacting linear polymer chains in a good solvent. Journal of Statistical Physics, 89(5–6):981–995, December 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02764217.

Korniss:1997:NPT

[KSZ97] G. Korniss, B. Schmittmann, and R. K. P. Zia. Nonequilibrium phase transitions in a simple three-state lattice gas. Journal of Statistical Physics, 86(3–4):721–748, February 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02199117.

Korutcheva:1991:FSE

[KT91] Elka R. Korutcheva and Nicholai S. Tonchev. Finite-size effects in a field-theoretic model with long-range exchange interaction. Journal of Statistical Physics, 62(3–4):553–562, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017972.

Koma:1994:SBF

[KT94] Tohru Koma and Hal Tasaki. Symmetry breaking and finite-size effects in quantum many-body systems. Journal of Statistical Physics, 76(3–4):745–803, August 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02188685.

Katsoulakis:1999:MAI

[KT99] Markos A. Katsoulakis and Athanasios E. Tzavaras. Multiscale analysis for interacting particles: Relaxation systems and scalar conservation laws. Journal of Statistical Physics, 96(3– 4):715–763, August 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004670308361.

Kulske:1997:MDM

[Kül97] Christof Külske. Metastates in disordered mean-field models: Random field and Hopfield models. Journal of Statistical Physics, 88(5–6):1257–1293, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02732434.

Kulske:1998:MDM

[Kül98] Christof Külske. Metastates in disordered mean-field models II: The superstates. Journal of Statistical Physics, 91(1–2):155–176, April 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023040121034.

Kunz:1994:SOM

[Kun94] Hervé Kunz. Surface orbital magnetism. Journal of Statistical Physics, 76(1–2):183–207, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02188660.

Kuske:1999:PDN

[Kus99] R. Kuske. Probability densities for noisy delay bifurcations. Journal of Statistical Physics, 96(3–4):797–816, August 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004658609270.

Kooiman:1992:FFA

[KvL92] A. Kooiman and J. M. J. van Leeuwen. Free fermion approximation for the Ising model with further-neighbor interactions on a triangular lattice. Journal of Statistical Physics, 69(1–2): 247–275, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053793.

Kim:1993:KTG

[KW93] Sang Rak Kim and Leslie V. Woodcock. Kinetic theory of granular shear flow: Constitutive relations for the hard-disk model. Journal of Statistical Physics, 71(1–2):143–162, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01048092.

Klar:1997:ELK

[KW97] A. Klar and R. Wegener. Enskog-like kinetic models for vehicular traffic. Journal of Statistical Physics, 87(1–2):91–114, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02181481.

Kolomeisky:1998:SRM

[KW98] Anatoly B. Kolomeisky and B. Widom. A simplified ‘ratchet’ model of molecular motors. Journal of Statistical Physics, 93(3– 4):633–645, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000033246.14231.e1.

Klyatskin:1996:DPT

[KWG96] V. I. Klyatskin, W. A. Woyczynski, and D. Gurarie. Diffusing passive tracers in random incompressible flows: Statistical topography aspects. Journal of Statistical Physics, 84(3–4):797–836, August 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179658.

Klein:1993:AFO

[KY93a] David Klein and Wei-Shih Yang. Absence of first-order phase transitions for antiferromagnetic systems. Journal of Statistical Physics, 70(5–6):1391–1400, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01049441.

Klein:1993:CFO

[KY93b] David Klein and Wei-Shih Yang. A characterization of first-order phase transitions for superstable interactions in classical statistical mechanics. Journal of Statistical Physics, 71(5–6):1043–1062, June 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049960.

Kesten:1993:TOC

[KZ93] Harry Kesten and Yu Zhang. The tortuosity of occupied crossings of a box in critical percolation. Journal of Statistical Physics, 70(3–4):599–611, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053586.

Landau:1994:OQP

[Lan94] L. J. Landau. Observation of quantum particles on a large space– time scale. Journal of Statistical Physics, 77(1–2):259–309, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186843.

Lang:1995:ECS

[Lan95] Reinhard Lang. Effective conductivity and skew Brownian motion. Journal of Statistical Physics, 80(1–2):125–146, July 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02178356.

Lemke:1995:CDH

[LAT95] Ney Lemke, Jeferson J. Arenzon, and Francisco A. Tamarit. Chaotic dynamics of high-order neural networks. Journal of Statistical Physics, 79(1–2):415–427, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179396.

Lakshminarayan:1994:RLI

[LB94] A. Lakshminarayan and N. L. Balazs. Relaxation and localization in interacting quantum maps. Journal of Statistical Physics, 77(1–2):311–344, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02186844.

Levy:1996:WNC

[LB96] Ohad Levy and David J. Bergman. Weakly nonlinear conductivity of random composites: A series expansion approach. Journal of Statistical Physics, 82(5–6):1327–1344, March 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183385.

Lustfeld:1996:CFN

[LBK96] H. Lustfeld, J. Bene, and Z. Kaufmann. The correlation functions near intermittency in a one-dimensional piecewise parabolic map. Journal of Statistical Physics, 83(5–6):1199–1210, June 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179558.

Lehmani:1997:TIS

[LBT97] Albert Lehmani, Olivier Bernard, and Pierre Turq. Transport of ions and solvent in confined media. Journal of Statistical Physics, 89(1–2):379–402, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02770771.

Leung:1991:TPF

[LC91] Kevin Leung and David Chandler. Theory of percolation in fluids of long molecules. Journal of Statistical Physics, 63(5–6):837– 856, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01029986.

Laughton:1995:MLF

[LC95a] S. N. Laughton and A. C. C. Coolen. Macroscopic Lyapunov functions for separable stochastic neural networks with detailed balance. Journal of Statistical Physics, 80(1–2):375–387, July 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02178364.

Lee:1995:RGS

[LC95b] Benjamin P. Lee and John Cardy. Renormalization group study of the A+B → ⊘ diffusion-limited reaction. Journal of Statistical Physics, 80(5–6):971–1007, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179861. See erra-

tum [LC97].

Lee:1997:ERG

[LC97] Benjamin P. Lee and John Cardy. Erratum: Renormalization group study of the A + B → ⊘ diffusion-limited reaction. Journal of Statistical Physics, 87(3–4):951–954, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF02181256. See [LC95b].

Lemaitre:1999:RGS

[LC99] Anaël Lemaître and Hugues Chaté. Renormalization group for strongly coupled maps. Journal of Statistical Physics, 96(5–6): 915–962, September 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004603332295.

Levitan:1998:ORT

[LD98] Boris Levitan and Eytan Domany. Ostwald ripening in two dimensions: Treatment with pairwise interactions. Journal of Statistical Physics, 93(3–4):501–510, November 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033161.03386.0b.

Lam:1991:F

[LDBA91] Lui Lam, Flonnie Dowell, Helmut Brand, and Guenter Ahlers. Foreword. Journal of Statistical Physics, 64(5–6):899–901, September 1991. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/BF01048802.

Lebowitz:1990:PSM

[Leb90] Joel L. Lebowitz. Program of the 62nd Statistical Mechanics Meeting. Journal of Statistical Physics, 58(5–6):1275–1282, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01026578.

Lebowitz:1991:PSM

[Leb91] Joel L. Lebowitz. Program of the 65th Statistical Mechanics Meeting. Journal of Statistical Physics, 65(3–4):827–834, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01053760.

Lebowitz:1992:BRP

[Leb92a] Joel L. Lebowitz. Book review: Program of the 67th statistical mechanics meeting. Journal of Statistical Physics, 69(3–4): 907–913, November 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01050443.

Lebowitz:1992:PSM

[Leb92b] Joel L. Lebowitz. Program of the 66th Statistical Mechanics Meeting. Journal of Statistical Physics, 67(1–2):421–429, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049044.

Lebowitz:1993:PSMa

[Leb93a] Joel L. Lebowitz. Program of the 68th Statistical Mechanics Meeting. Journal of Statistical Physics, 71(3–4):827–835, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01058451.

Lebowitz:1993:PSMb

[Leb93b] Joel L. Lebowitz. Program of the 69th Statistical Mechanics Meeting. Journal of Statistical Physics, 73(1–2):453–458, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01052776.

Lebowitz:1994:P

[Leb94a] Joel L. Lebowitz. Preface. Journal of Statistical Physics, 77 (1–2):1, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186828.

Lebowitz:1994:PSMa

[Leb94b] Joel L. Lebowitz. Program of the 70th Statistical Mechanics Meeting. Journal of Statistical Physics, 75(1–2):347–354, April 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186293.

Lebowitz:1994:PSMb

[Leb94c] Joel L. Lebowitz. Program of the 71st Statistical Mechanics Meeting. Journal of Statistical Physics, 77(1–2):509–513, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186856.

Lebowitz:1995:LO

[Leb95a] Joel L. Lebowitz. Lars Onsager. Journal of Statistical Physics, 78(1–2):1–3, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183335.

Lebowitz:1995:PSMa

[Leb95b] Joel L. Lebowitz. Program of the 72nd Statistical Mechanics Meeting. Journal of Statistical Physics, 80(5–6):1455– 1462, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179882.

Lebowitz:1995:PSMb

[Leb95c] Joel L. Lebowitz. Program of the 73rd Statistical Mechanics Meeting. Journal of Statistical Physics, 80(5–6):1463– 1467, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179883.

Lebowitz:1996:PSM

[Leb96] Joel L. Lebowitz. Program of the 74th Statistical Mechanics Meeting. Journal of Statistical Physics, 84(5–6):1393– 1400, September 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02174140.

Lebowitz:1997:P

[Leb97a] Joel L. Lebowitz. Preface. Journal of Statistical Physics, 87 (5–6):973–974, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181264.

Lebowitz:1997:PSMa

[Leb97b] Joel L. Lebowitz. Program of the 75th Statistical Mechanics Meeting. Journal of Statistical Physics, 86(1–2):445–452, January 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02180217.

Lebowitz:1997:PSMb

[Leb97c] Joel L. Lebowitz. Program of the 76th Statistical Mechanics Meeting. Journal of Statistical Physics, 88(1–2):525–534, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02508485.

Lebowitz:1997:PSMc

[Leb97d] Joel L. Lebowitz. Program of the 77th statistical mechanics meeting. Journal of Statistical Physics, 89(3–4):885–891, November 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02765550.

Lebowitz:1997:TAA

[Leb97e] Joel L. Lebowitz. Time’s arrow and Archimedes’ point. Journal of Statistical Physics, 87(1–2):463–468, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02181499.

Lebowitz:1998:PSM

[Leb98] Joel L. Lebowitz. Program of the 79th Statistical Mechanics Meeting. Journal of Statistical Physics, 92(5–6):1217– 1223, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023021400212.

Leboeuf:1999:RAC

[Leb99] P. Leboeuf. Random analytic chaotic eigenstates. Journal of Statistical Physics, 95(3–4):651–664, May 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004595310043.

Lefevere:1999:VPS

[Lef99a] R. Lefevere. Variational principle for some renormalized measures. Journal of Statistical Physics, 96(1–2):109–133, July 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004568315437.

Lefevere:1999:WGM

[Lef99b] R. Lefevere. Weakly Gibbsian measures and quasilocality: A long-range pair-interaction counterexample. Journal of Statistical Physics, 95(3–4):785–789, May 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004511913677.

Leggett:1998:SFA

[Leg98] A. J. Leggett. On the superfluid fraction of an arbitrary manybody system at T = 0. Journal of Statistical Physics, 93(3–4): 927–941, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000033170.38619.6c.

Lemberger:1995:LFV

[Lem95] Pirmin Lemberger. Large-field versus small-field expansions and Sobolev inequalities. Journal of Statistical Physics, 79(3–4):525– 568, May 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02184870.

Levermore:1996:MCH

[Lev96] C. David Levermore. Moment closure hierarchies for kinetic theories. Journal of Statistical Physics, 83(5–6):1021–1065, June 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02179552.

Liu:1990:CST

[LF90] Andrea J. Liu and Michael E. Fisher. On the corrections to scaling in three-dimensional Ising models. Journal of Statistical Physics, 58(3–4):431–442, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01112754.

Lauritsen:1993:CEP

[LF93] Kent Bækgaard Lauritsen and Hans C. Fogedby. Critical exponents from power spectra. Journal of Statistical Physics, 72 (1–2):189–205, July 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048046.

Lipscombe:1991:CPS

[LFtH91] T. C. Lipscombe, A. L. Frenkel, and D. ter Haar. On the convection of a passive scalar by a turbulent Gaussian velocity field. Journal of Statistical Physics, 63(1–2):305–313, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01026606.

Lowe:1997:DFL

[LFvdH97] C. P. Lowe, D. Frenkel, and M. A. van der Hoef. Deviations from Fick’s law in Lorentz gases. Journal of Statistical Physics, 87(5– 6):1229–1244, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181281.

Lendi:1998:RQR

[LFvW98] K. Lendi, F. Farhadmotamed, and A. J. van Wonderen. Regularization of quantum relative entropy in finite dimensions and application to entropy production. Journal of Statistical Physics, 92(5–6):1115–1135, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023004929346.

Lin:1992:FKM

[LH92] Bin Lin and Bambi Hu. Frenkel–Kontorova model with Toda interactions. Journal of Statistical Physics, 69(5–6):1047–1068, December 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01058761.

Lee:1994:IME

[LH94] F. T. Lee and M. C. Huang. Ising model in an external field on a hierarchical lattice. Journal of Statistical Physics, 75(5– 6):1119–1135, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186759.

Lach-hab:1997:ISP

[LhBBS97] Mohammed Lach-hab, Estela Blaisten-Barojas, and T. Sauer. Irregular scattering of particles confined to ring-bounded cavities. Journal of Statistical Physics, 87(1–2):137–146, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02181483.

Longuet-Higgins:1995:LON

[LHF95] H. Christopher Longuet-Higgins and Michael E. Fisher. Lars Onsager: November 27, 1903–October 5, 1976. Journal of Statistical Physics, 78(1–2):605–640, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183371.

Li:1990:MIF

[Li90] Wentian Li. Mutual information functions versus correlation functions. Journal of Statistical Physics, 60(5–6):823– 837, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01025996.

Li:1992:PNC

[Li92] Wentian Li. Phenomenology of nonlocal cellular automata. Journal of Statistical Physics, 68(5–6):829–882, September 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01048877.

Liao:1991:KEN

[Lia91] Wuwell Liao. Kolmogorov exponents for near-incompressible turbulence from perturbative quantum field theory. Journal of Statistical Physics, 65(1–2):1–32, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01329848.

Liebovitch:1993:IPS

[Lie93] Larry S. Liebovitch. Interpretation of protein structure and dynamics from the statistics of the open and closed times measured in a single ion-channel protein. Journal of Statistical Physics, 70 (1–2):329–337, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053971.

Lisyansky:1992:RGE

[LIF92] A. A. Lisyansky, Yu. M. Ivanchenko, and A. A. Filippov. Renormalization group equations in local approximation. Journal of Statistical Physics, 66(5–6):1667–1673, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01054443.

Lindblad:1993:IRL

[Lin93] Göran Lindblad. Irreversibility and randomness in linear response theory. Journal of Statistical Physics, 72(3–4):539–554, August 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048023.

Littlejohn:1992:VVF

[Lit92] Robert G. Littlejohn. The Van Vleck formula, Maslov theory, and phase space geometry. Journal of Statistical Physics, 68(1–2):7–50, July 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048836.

Liu:1998:RPA

[Liu98] Pei-Dong Liu. Random perturbations of axiom a basic sets. Journal of Statistical Physics, 90(1–2):467–490, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023280407906.

Liverani:1995:DCP

[Liv95] Carlangelo Liverani. Decay of correlations for piecewise expanding maps. Journal of Statistical Physics, 78(3–4):1111–1129, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183704.

Lee:1995:FPT

[LK95] Jysoo Lee and Joel Koplik. First passage time in a two-layer system. Journal of Statistical Physics, 79(5–6):895–922, June 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02181208.

Levy:1998:DRN

[LK98] Ohad Levy and Robert V. Kohn. Duality relations for non– Ohmic composites, with applications to behavior near percolation. Journal of Statistical Physics, 90(1–2):159–189, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023251701546.

Litz:1992:MDC

[LLH92] Peter Litz, Stefan Langenbach, and Alfred Hüller. Molecular dynamics calculation for the modified xy model. Journal of Statistical Physics, 66(5–6):1659–1665, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01054442.

Lomba:1995:SHC

[LLM95] Enrique Lomba and José Luis López-Martín. On the solutions of the hypernetted chain equation inside the gas-liquid coexistence region. Journal of Statistical Physics, 80(3–4):825–839, August 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02178557.

Laloe:1990:SE

[LM90] F. Laloë and W. J. Mullin. On the Snider equation. Journal of Statistical Physics, 59(3–4):725–744, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01025848.

LeDoussal:1991:SAW

[LM91] P. Le Doussal and J. Machta. Self-avoiding walks in quenched random environments. Journal of Statistical Physics, 64(3–4): 541–578, August 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048306.

Losson:1992:HLE

[LM92] Jérôme Losson and Michael C. Mackey. A Hopf-like equation and perturbation theory for differential delay equations. Journal of Statistical Physics, 69(5–6):1025–1046, December 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01058760.

Lebowitz:1994:LTP

[LM94a] Joel L. Lebowitz and Nicolas Macris. Low-temperature phases of itinerant fermions interacting with classical phonons: The static Holstein model. Journal of Statistical Physics, 76(1–2):91–123, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02188657.

Lerner:1994:FPR

[LM94b] E. Yu. Lerner and M. D. Missarov. Fixed points of renormalization group for the hierarchical fermionic model. Journal of Statistical Physics, 76(3–4):805–817, August 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02188686.

Lebowitz:1996:RLT

[LM96] J. L. Lebowitz and A. E. Mazel. A remark on the lowtemperature behavior of the SOS interface in half-space. Journal of Statistical Physics, 84(3–4):379–397, August 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179648.

Lorinczi:1997:WGM

[LM97] Jôzsef Lörinczi and Christian Maes. Weakly Gibbsian measures for lattice spin systems. Journal of Statistical Physics, 89(3–4): 561–579, November 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02765536.

Lebowitz:1998:IPA

[LM98] J. L. Lebowitz and A. E. Mazel. Improved Peierls argument for high-dimensional Ising models. Journal of Statistical Physics, 90(3–4):1051–1059, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023205826704.

Lebowitz:1992:AVI

[LMM92] Joel L. Lebowitz, Nicolas Macris, and Philippe A. Martin. Atomic versus ionized states in many-particle systems and the spectra of reduced density matrices: A model study. Journal of Statistical Physics, 67(5–6):909–956, June 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01049005.

Lindgren:1998:CTD

[LMN98] Kristian Lindgren, Cristopher Moore, and Mats Nordahl. Complexity of two-dimensional patterns. Journal of Statistical Physics, 91(5–6):909–951, June 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023027932419.

Lebowitz:1999:LVP

[LMP99] J. L. Lebowitz, A. Mazel, and E. Presutti. Liquid-vapor phase transitions for systems with finite–range interactions. Journal of Statistical Physics, 94(5–6):955–1025, March 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004591218510.

Laanait:1993:PCP

[LMR93] Lahoussine Laanait, Noureddine Masaif, and Jean Ruiz. Phase coexistence in partially symmetric q-state models. Journal of Statistical Physics, 72(3–4):721–736, August 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01048030.

Lebowitz:1990:SMP

[LMS90] Joel L. Lebowitz, Christian Maes, and Eugene R. Speer. Statistical mechanics of probabilistic cellular automata. Journal of Statistical Physics, 59(1–2):117–170, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01015566.

Li:1995:CEH

[LMS95] Bin Li, Neal Madras, and Alan D. Sokal. Critical exponents, hyperscaling, and universal amplitude ratios for two- and threedimensional self-avoiding walks. Journal of Statistical Physics, 80 (3–4):661–754, August 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02178552.

Liu:1997:SCB

[LO97] Cliff Z.-W. Liu and Irwin Oppenheim. Spatial correlations in bounded nonequilibrium fluid systems. Journal of Statistical Physics, 86(1–2):179–190, January 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02180203.

Longtin:1993:SRN

[Lon93] André Longtin. Stochastic resonance in neuron models. Journal of Statistical Physics, 70(1–2):309–327, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01053970.

Lopes:1990:FOL

[Lop90] A. O. Lopes. A first-order level-2 phase transition in thermodynamic formalism. Journal of Statistical Physics, 60(3– 4):395–411, August 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01314928.

Lebowitz:1991:PMS

[LOP91] Joel L. Lebowitz, Enza Orlandi, and Errico Presutti. A particle model for spinodal decomposition. Journal of Statistical Physics, 63(5–6):933–974, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01029992.

Leonenko:1996:RCNb

[LOP96a] N. N. Leonenko, E. Orsingher, and V. N. Parkhomenko. On the rate of convergence to the normal law for solutions of the Burgers equation with singular initial data. Journal of Statistical Physics, 84(5–6):1389–1391, September 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02174139.

Leonenko:1996:RCNa

[LOP96b] Nikolai N. Leonenko, Enzo Orsingher, and Victoria N. Parkhomenko. On the rate of convergence to the normal law for solutions of the Burgers equation with singular initial data. Journal of Statistical Physics, 82(3–4):915–930, February 1996.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179795.

Lebowitz:1995:F

[LOQ95] J. L. Lebowitz, S. A. Orszag, and Y. H. Qian. Foreword. Journal of Statistical Physics, 81(1–2):1–4, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179963.

Loss:1990:LQEa

[Los90a] D. Loss. Linear quantum Enskog equation. I. Homogeneous quantum fluids. Journal of Statistical Physics, 59(3–4):691–723, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01025847.

Loss:1990:LQEb

[Los90b] D. Loss. Linear quantum Enskog equation II. Inhomogeneous quantum fluids. Journal of Statistical Physics, 61(1–2):467–493, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01013976.

Loviscach:1994:PMM

[Lov94] Jörn Loviscach. Probabilistic models of multidimensional piecewise expanding mappings. Journal of Statistical Physics, 75 (1–2):189–213, April 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186286.

Li:1990:FTD

[LP90a] Shiwei Li and J. K. Percus. Finite-temperature density functional theory of atoms in strong magnetic fields. Journal of Statistical Physics, 59(1–2):323–332, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01015572.

Litvin:1990:BAS

[LP90b] A. A. Litvin and V. B. Priezzhev. The Bethe ansatz for the six-vertex model with rotated boundary conditions. Journal of Statistical Physics, 60(3–4):307–321, August 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01314922.

Lacker:1991:HDO

[LP91] H. Michael Lacker and Allon Percus. How do ovarian follicles interact? A many-body problem with unusual symmetry and symmetry-breaking properties. Journal of Statistical Physics, 63 (5–6):1133–1161, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01030003.

Ledrappier:1994:DFB

[LP94] François Ledrappier and Anna Porzio. A dimension formula for Bernoulli convolutions. Journal of Statistical Physics, 76(5–6): 1307–1327, September 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02187064.

Ledrappier:1996:MABa

[LP96a] François Ledrappier and Anna Porzio. On the multifractal analysis of Bernoulli convolutions. I. Large-deviation results. Journal of Statistical Physics, 82(1–2):367–395, January 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02189235.

Ledrappier:1996:MABb

[LP96b] François Ledrappier and Anna Porzio. On the multifractal analysis of Bernoulli convolutions. II. Dimensions. Journal of Statistical Physics, 82(1–2):397–420, January 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02189236.

Langlands:1992:UCP

[LPPSA92] R. P. Langlands, C. Pichet, Ph. Pouliot, and Y. Saint-Aubin. On the universality of crossing probabilities in two-dimensional percolation. Journal of Statistical Physics, 67(3–4):553–574, May 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01049720.

Livi:1991:RSHa

[LPR91a] R. Livi, A. Politi, and S. Ruffo. Repeller structure in a hierarchical model. I. Topological properties. Journal of Statistical Physics, 65(1–2):53–72, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01329850.

Livi:1991:RSHb

[LPR91b] R. Livi, A. Politi, and S. Ruffo. Repeller structure in a hierarchical model. II. Metric properties. Journal of Statistical Physics, 65 (1–2):73–95, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01329851.

Lewis:1994:EEL

[LPS94] J. T. Lewis, C.-E. Pfister, and W. G. Sullivan. The equivalence of ensembles for lattice systems: Some examples and a counterexample. Journal of Statistical Physics, 77(1–2):397–419, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186849.

Lepri:1996:CLA

[LPT96] Stefano Lepri, Antonio Politi, and Alessandro Torcini. Chronotopic Lyapunov analysis. I. A detailed characterization of 1D systems. Journal of Statistical Physics, 82(5–6):1429–1452, March 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02183390.

Lepri:1997:CLA

[LPT97] Stefano Lepri, Antonio Politi, and Alessandro Torcini. Chronotopic Lyapunov analysis: II. Toward a unified approach. Journal of Statistical Physics, 88(1–2):31–45, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02508463.

Li:1992:PDR

[LPW92] Weixiong Li, Hyunggyu Park, and Michael Widom. Phase diagram of a random tiling quasicrystal. Journal of Statistical Physics, 66(1–2):1–69, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01060059.

Lim:1998:DFD

[LPY98] Hye Young Lim, Yong Moon Park, and Hyun Jae Yoo. Dirichlet forms and Dirichlet operators for Gibbs measures of quantum unbounded spin systems: Essential self-adjointness and log–Sobolev inequality. Journal of Statistical Physics, 90(3–4):

949–1002, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023245508048.

Lebowitz:1992:PWQ

[LR92] John L. Lebowitz and Peter J. Reynolds. Preface: What is quantum chaos? Journal of Statistical Physics, 68(1–2):1–4, July 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048834.

Luczka:1996:RFD

[LR96] Jerzy Luczka and Ryszard Rudnicki. Randomly flashing diffusion: Asymptotic properties. Journal of Statistical Physics, 83 (5–6):1149–1164, June 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179555.

Li:1990:HPM

[LS90a] Bin Li and Alan D. Sokal. High-precision Monte Carlo test of the conformai-invariance predictions for two-dimensional mutually avoiding walks. Journal of Statistical Physics, 61(3–4): 723–748, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01027299.

Lubachevsky:1990:GPR

[LS90b] Boris D. Lubachevsky and Frank H. Stillinger. Geometric properties of random disk packings. Journal of Statistical Physics, 60(5– 6):561–583, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01025983.

Lebedev:1991:SHC

[LS91a] N. I. Lebedev and A. S. Sigov. The scaling of higher cumulants in a diffusion problem. Journal of Statistical Physics, 63(3–4):701– 705, May 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01029206.

Lieb:1991:CCE

[LS91b] Elliott H. Lieb and Heinz Siedentop. Convexity and concavity of eigenvalue sums. Journal of Statistical Physics, 63(5–6):811–816, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01029984.

Lipowski:1992:CMF

[LS92] Adam Lipowski and Masuo Suzuki. Convergence of mean-field approximations in site percolation and application of CAM to d = 1 further-neighbors percolation problem. Journal of Statistical Physics, 69(1–2):1–16, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053779.

Lidsky:1997:GST

[LS97] David Lidsky and Michael J. Stephen. Ground-state and thermal properties of a long-range Josephson array. Journal of Statistical Physics, 87(3–4):897–908, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02181250.

Lebowitz:1999:GCT

[LS99] Joel L. Lebowitz and Herbert Spohn. A Gallavotti–Cohen-type symmetry in the large deviation functional for stochastic dynamics. Journal of Statistical Physics, 95(1–2):333–365, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004589714161.

Lindenberg:1991:SPD

[LSK91] Katja Lindenberg, Wen-Shyan Sheu, and Raoul Kopelman. Scaling properties of diffusion-limited reactions on fractal and Euclidean geometries. Journal of Statistical Physics, 65(5–6):1269– 1283, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049612.

Luding:1991:BAR

[LSKB91] S. Luding, H. Schnörer, V. Kuzovkov, and A. Blutnen. Bimolecular annihilation reactions: Immobile reactants and multipolar interactions. Journal of Statistical Physics, 65(5–6):1261–1267, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049611.

Lubachevsky:1991:DVS

[LSP91] Boris D. Lubachevsky, Frank H. Stillinger, and Elliot N. Pinson. Disks vs. spheres: Contrasting properties of random packings. Journal of Statistical Physics, 64(3–4):501–524, August 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01048304.

Lieb:1997:SIR

[LSS97] Elliott H. Lieb, Heinz Siedentop, and Jan Philip Solovej. Stability and instability of relativistic electrons in classical electromagnetic fields. Journal of Statistical Physics, 89(1–2):37–59, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02770753.

Lebowitz:1998:P

[LST98] Joel L. Lebowitz, Jack Swift, and Chao Tang. Preface. Journal of Statistical Physics, 93(3–4):407–409, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033228.20564.cd.

Lambert:1993:SPN

[LSV93] A. Lambert, S. Siboni, and S. Vaienti. Statistical properties of a nonuniformly hyperbolic map of the interval. Journal of Statistical Physics, 72(5–6):1305–1330, September 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01048188.

Lu:1999:CEE

[Lu99] Xuguang Lu. Conservation of energy, entropy identity, and local stability for the spatially homogeneous Boltzmann equation. Journal of Statistical Physics, 96(3–4):765–796, August 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004606525200.

Luck:1993:CBA

[Luc93] J. M. Luck. Critical behavior of the aperiodic quantum Ising chain in a transverse magnetic field. Journal of Statistical Physics, 72(3–4):417–458, August 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01048019.

Luo:1997:ASL

[Luo97] Li-Shi Luo. Analytic solutions of linearized lattice Boltzmann equation for simple flows. Journal of Statistical Physics, 88(3– 4):913–926, August 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000015178.19008.78.

Lutken:1995:ATC

[Lüt95] C. A. Lütken. Anisotropic transport of charge and complexified duality. Journal of Statistical Physics, 78(1–2):441–451, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183358.

Levesque:1993:MDT

[LV93] D. Levesque and L. Verlet. Molecular dynamics and time reversibility. Journal of Statistical Physics, 72(3–4):519–537, August 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048022.

Lorinczi:1994:NPG

[LV94] József Lörinczi and Koen Vande Velde. A note on the projecton of Gibbs measures. Journal of Statistical Physics, 77(3–4): 881–887, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179465.

Lu:1992:CSC

[LVY92] Zi-Min Lu, Michel Vallières, and Jian-Min Yuan. Capture by stabilized continuum: Classical and quantum aspects. Journal of Statistical Physics, 68(1–2):189–206, July 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01048842.

Landsberg:1993:CFN

[LW93] P. T. Landsberg and R. P. Woodard. Classical fluids of negative heat capacity. Journal of Statistical Physics, 73(1–2):361–378, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01052765.

Lamb:1998:MSS

[LW98a] Jeroen S. W. Lamb and Frank Wijnands. From multi-site to on-site transfer matrix models for self-similar chains. Journal of Statistical Physics, 90(1–2):261–284, January 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023212020202.

Leonenko:1998:SLS

[LW98b] Nikolai N. Leonenko and Wojbor A. Woyczynski. Scaling limits of solutions of the heat equation for singular non–Gaussian data. Journal of Statistical Physics, 91(1–2):423–438, April 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023060625577.

Lo:1997:DIH

[LY97] ChiCheung Lo and Yeung Yam. Derivation of an improved Hodgkin–Huxley model for potassium channel by means of the Fokker–Planck equation. Journal of Statistical Physics, 89(5–6): 997–1016, December 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02764218.

Lyons:1995:LO

[Lyo95] Philip Lyons. Lars Onsager (1903-1976). Journal of Statistical Physics, 78(1–2):595–603, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183370.

Lloyd:1991:ATS

[LZ91] Seth Lloyd and Wojciech H. Zurek. Algorithmic treatment of the spin-echo effect. Journal of Statistical Physics, 62(3–4):819–839, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017985.

Liebovitch:1998:SUS

[LZ98] Larry S. Liebovitch and Michal Zochowski. Significance of updating schemes in computational models: Dynamics of neutral networks. Journal of Statistical Physics, 90(1–2):253–260, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023259903363.

Marro:1998:ALG

[MA98] J. Marro and A. Achahbar. Anisotropic lattice gases. Journal of Statistical Physics, 90(3–4):817–826, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023281121252.

Machta:1993:CCP

[Mac93] Jonathan Machta. The computational complexity of pattern formation. Journal of Statistical Physics, 70(3–4):949–966, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01053602.

Machlup:1995:LOW

[Mac95a] Stefan Machlup. Lars Onsager was my thesis director. Journal of Statistical Physics, 78(1–2):589–592, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183368.

MacKay:1995:CSM

[Mac95b] R. S. MacKay. The classical statistical mechanics of Frenkel– Kontorova models. Journal of Statistical Physics, 80(1–2):45–67, July 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02178353.

Machado:1997:BEP

[Mac97] Fábio P. Machado. Branching exclusion process on a strip. Journal of Statistical Physics, 86(3–4):765–777, February 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02199119.

Madras:1995:RBC

[Mad95] Neal Madras. A rigorous bound on the critical exponent for the number of lattice trees, animals, and polygons. Journal of Statistical Physics, 78(3–4):681–699, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183684.

Maes:1990:KLC

[Mae90] Christian Maes. Kinetic limit of a conservative lattice gas dynamics showing long-range correlations. Journal of Statistical Physics, 61(3–4):667–681, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01027296.

Maes:1999:FTG

[Mae99] Christian Maes. The fluctuation theorem as a Gibbs property. Journal of Statistical Physics, 95(1–2):367–392, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004541830999.

Majda:1993:EIR

[Maj93] Andrew J. Majda. Explicit inertial range renormalization theory in a model for turbulent diffusion. Journal of Statistical Physics, 73(3–4):515–542, November 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01054338.

Majda:1994:RSD

[Maj94] Andrew J. Majda. Random shearing direction models for isotropic turbulent diffusion. Journal of Statistical Physics, 75 (5–6):1153–1165, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186761.

Manna:1990:LSS

[Man90] S. S. Manna. Large-scale simulation of avalanche cluster distribution in sand pile model. Journal of Statistical Physics, 59 (1–2):509–521, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01015580.

Mantegna:1993:TEP

[Man93] Rosario Nunzio Mantegna. Time evolution of the probability distribution in stochastic and chaotic systems with enhanced diffusion. Journal of Statistical Physics, 70(3–4):721–736, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01053592.

Marchetti:1990:CPL

[Mar90a] Domingos H. U. Marchetti. Comments on ‘Power-law falloff in the Kosterlitz–Thouless phase of a two-dimensional lattice Coulomb gas’. Journal of Statistical Physics, 61(3–4):909– 911, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01027310. See [MKP90].

Martin:1990:ILS

[Mar90b] J. L. Martin. The impact of large-scale computing on lattice statistics. Journal of Statistical Physics, 58(3–4):749–774, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01112773.

March:1992:RSM

[Mar92a] Peter March. Remarks on scaling a model of Witten–Sander type. Journal of Statistical Physics, 67(5–6):1117–1149, June 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01049012.

Marchetti:1992:SPO

[Mar92b] Domingos H. U. Marchetti. Smooth phase in the one-dimensional discrete Gaussian model with 1/(−j)^{2} interaction at inverse temperature β > 1. Journal of Statistical Physics, 66(5–6): 1319–1342, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01054424.

Martens:1992:QQD

[Mar92c] Craig C. Martens. Quantum qualitative dynamics. Journal of Statistical Physics, 68(1–2):207–237, July 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01048843.

Martinelli:1992:DAL

[Mar92d] Fabio Martinelli. Dynamical analysis of low-temperature Monte Carlo cluster algorithms. Journal of Statistical Physics, 66(5– 6):1245–1276, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01054422.

Marchesoni:1993:PCM

[Mar93a] F. Marchesoni. Persistent currents in mesoscopic rings: A stochastic model. Journal of Statistical Physics, 70(1–2):247– 256, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053966.

Markos:1993:WDE

[Mar93b] P. Markos. Weak disorder expansion of Lyapunov exponents of products of random matrices: A degenerate theory. Journal of Statistical Physics, 70(3–4):899–919, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01053599.

Marsili:1994:RTS

[Mar94a] Matteo Marsili. Run time statistics in models of growth in disrrdered media. Journal of Statistical Physics, 77(3–4):733– 754, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179459.

Martin:1994:EIA

[Mar94b] André Martin. Is energy increasing with angular momentum? Journal of Statistical Physics, 76(1–2):497–504, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02188672.

Martinelli:1994:TDD

[Mar94c] F. Martinelli. On the two-dimensional dynamical Ising model in the phase coexistence region. Journal of Statistical Physics, 76(5–6):1179–1246, September 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02187060.

Markarian:1995:SPD

[Mar95] Roberto Markarian. Statistical properties of dynamical systems with singularities. Journal of Statistical Physics, 80(5–6):1207– 1239, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179869.

Martin:1997:OBG

[Mar97] André Martin. ‘Optimal’ bounds on the ground-state energy of n-body systems of bosons and fermions interacting by attractive forces. Journal of Statistical Physics, 89(1–2):445–451, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02770775.

Marder:1998:EKC

[Mar98] M. Marder. Energies of a kinked crack line. Journal of Statistical Physics, 93(3–4):511–525, November 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033239.22129.c1.

Masoliver:1992:BRC

[Mas92] Jaume Masoliver. Book review: Chaotic behavior of deterministic dissipative systems. Journal of Statistical Physics, 66(3– 4):1185, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01055726.

Matsui:1990:LBQ

[Mat90] Taku Matsui. A link between quantum and classical Potts models. Journal of Statistical Physics, 59(3–4):781–798, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01025850.

Mattis:1994:ERD

[Mat94] Daniel Mattis. Electrical resistivity of dilute, interacting fermions. Journal of Statistical Physics, 77(1–2):383–396, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186848.

Mazza:1998:MPC

[Maz98] Christian Mazza. On the mean pair correlation function of ±J Ising spin glasses. Journal of Statistical Physics, 92(3–4): 559–569, August 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023036520918.

Menon:1997:CLI

[MBD97] Gautam I. Menon, Mustansir Barma, and Deepak Dhar. Conservation laws and integrability of a one-dimensional model of diffusing dimers. Journal of Statistical Physics, 86(5–6):1237–1263, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183622.

Mezincescu:1997:DRR

[MBF^{+}97] G. Andrei Mezincescu, Daniel Bessis, Jean-Daniel Fournier, Giorgio Mantica, and Francisc D. Aaron. Distribution of roots of random real generalized polynomials. Journal of Statistical Physics, 86(3–4):675–705, February 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02199115.

McDowell:1992:ETC

[MC92a] H. Keith McDowell and A. M. Clogston. Exact time correlation function for a nonlinearly coupled vibrational system. Journal of Statistical Physics, 67(1–2):331–346, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01049038.

Monge:1992:STB

[MC92b] Alessandro Monge and E. G. D. Cohen. Short-time behavior of the vibrational-energy-time correlation function for a onedimensional model of diatomic molecules. Journal of Statistical Physics, 66(1–2):595–605, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01060083.

Murphy:1993:MNC

[MC93] T. J. Murphy and E. G. D. Cohen. Maximum number of collisions among identical hard spheres. Journal of Statistical Physics, 71 (5–6):1063–1080, June 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049961.

Manificat:1994:MCS

[MC94] G. Manificat and J. M. Caillol. Monte Carlo simulations of the two-dimensional two-component plasma on a line. Journal of Statistical Physics, 74(5–6):1309–1319, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02188234.

Morgado:1990:MMH

[MCC90] W. A. M. Morgado, S. Coutinho, and E. M. F. Curado. Multifractal magnetization on hierarchical lattices. Journal of Statistical Physics, 61(3–4):913–921, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01027311.

McCauley:1995:WEB

[McC95] Joseph L. McCauley. Weak electrolytes, Brownian motion, vortices in superfluid films, and Odins Aker. Journal of Statistical Physics, 78(1–2):531–548, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183363.

McKean:1994:LLG

[McK94] H. P. McKean. A limit law for the ground state of Hill’s equation. Journal of Statistical Physics, 74(5–6):1227–1232, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02188225.

McKean:1995:SMN

[McK95] H. P. McKean. Statistical mechanics of nonlinear wave equations. 3. Metric transitivity for hyperbolic sine-Gordon. Journal of Statistical Physics, 79(3–4):731–737, May 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02184878. See cor-

rection [McK99].

McKean:1999:CSM

[McK99] H. P. McKean. Correction. Statistical Mechanics of Nonlinear Wave Equations (3). Metric Transitivity for Hyperbolic Sine-Gordon. Journal of Statistical Physics, 95(1–2):517, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004554218704. See [McK95].

McQueen:1997:PFS

[McQ97] Philip G. McQueen. Physics and fractal structures. Journal of Statistical Physics, 86(5–6):1397–1398, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183631.

Miller:1994:WDE

[MD94] Jeffrey D. Miller and Bernard Derridda. Weak-disorder expansion for the Anderson model on a tree. Journal of Statistical Physics, 75(3–4):357–388, May 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02186867.

Myshlyavtsev:1997:SAT

[MD97] A. V. Myshlyavtsev and M. D. Dongak. Statistics of adsorption on top and bridge sites of a square lattice: Transfer-matrix approach. Journal of Statistical Physics, 87(3–4):593–606, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02181238.

Morris:1991:POE

[MdBM91] Stephen W. Morris, John R. de Bruyn, and A. D. May. Patterns at the onset of electroconvection in freely suspended smectic films. Journal of Statistical Physics, 64(5–6):1025–1043, September 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048812.

Melnikov:1993:EAD

[Mel93] V. I. Mel’nikov. Enhancement of activated decay of metastable states by resonant pumping. Journal of Statistical Physics, 70 (1–2):77–91, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053955.

Menon:1992:FPT

[Men92] S. V. G. Menon. First passage time distribution in an oscillating field. Journal of Statistical Physics, 66(5–6):1675–1682, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01054444.

Mertens:1990:LAF

[Mer90] S. Mertens. Lattice animals: A fast enumeration algorithm and new perimeter polynomials. Journal of Statistical Physics, 58(5– 6):1095–1108, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01026565.

Meredith:1992:SWN

[Mer92] D. C. Meredith. Semiclassical wavefunctions of nonintegrable systems and localization on periodic orbits. Journal of Statistical Physics, 68(1–2):97–130, July 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01048838.

Merola:1999:AEP

[Mer99] I. Merola. Asymptotic expansion of the pressure in the inverse interaction range. Journal of Statistical Physics, 95(3–4):745– 758, May 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004503611860.

Meyer:1996:QCA

[Mey96] David A. Meyer. From quantum cellular automata to quantum lattice gases. Journal of Statistical Physics, 85(5–6):551–574, December 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02199356.

Martins:1991:BAR

[MF91] M. J. Martins and R. M. Fye. Bethe ansatz results for Hubbard chains with toroidal boundary conditions. Journal of Statistical Physics, 64(1–2):271–276, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01057877.

Mikheev:1992:MDF

[MF92] Lev V. Mikheev and Michael E. Fisher. Microcanonical density functionals for critical systems: An exact one-dimensional example. Journal of Statistical Physics, 66(5–6):1225–1244, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01054421.

Morris:1992:FSSa

[MG92] J. R. Morris and R. J. Gooding. Finite-size scaling study of a first-order temperature-driven symmetry-breaking structural phase transition. Journal of Statistical Physics, 67(3–4):471–506, May 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049717.

Machta:1994:PCG

[MG94] Jonathan Machta and Raymond Greenlaw. The parallel complexity of growth models. Journal of Statistical Physics, 77(3– 4):755–781, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179460.

Morbidelli:1995:SSK

[MG95] Alessandro Morbidelli and Antonio Giorgilli. Superexponential stability of KAM tori. Journal of Statistical Physics, 78(5–6): 1607–1617, March 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02180145.

Machta:1996:CCG

[MG96] Jonathan Machta and Raymond Greenlaw. The computational complexity of generating random fractals. Journal of Statistical Physics, 82(5–6):1299–1326, March 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183384.

McNamara:1995:STL

[MGA95] Guy R. McNamara, Alejandro L. Garcia, and Berni J. Alder. Stabilization of thermal lattice Boltzmann models. Journal of Statistical Physics, 81(1–2):395–408, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179986.

McNamara:1997:HCT

[MGA97] Guy R. McNamara, Alejandro L. Garcia, and Berni J. Alder. A hydrodynamically correct thermal lattice Boltzmann model. Journal of Statistical Physics, 87(5–6):1111–1121, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02181274.

Munoz:1998:PSS

[MGD98] M. A. Muñoz, G. Grinstein, and R. Dickman. Phase structure of systems with infinite numbers of absorbing states. Journal of Statistical Physics, 91(3–4):541–569, May 1998. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023021409588.

Moseley:1992:KDI

[MGJ92] L. L. Moseley, P. W. Gibbs, and Naeem Jan. Kawasaki dynamics with infinite-range spin exchange. Journal of Statistical Physics, 67(3–4):813–818, May 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049728.

Marin:1994:NTD

[MGS94] C. Marín, V. Garzó, and A. Santos. Nonlinear transport in a dilute binary mixture of mechanically different particles. Journal of Statistical Physics, 75(5–6):797–816, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02186744.

Moukarzel:1992:VRL

[MH92] C. Moukarzel and H. J. Herrmann. A vectorizable random lattice. Journal of Statistical Physics, 68(5–6):911–923, September 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048880.

Meo:1990:MCS

[MHB90] Marco D’Onorio De Meo, Dieter W. Heermann, and Kurt Binder. Monte Carlo study of the Ising model phase transition in terms of the percolation transition of ‘physical clusters’. Journal of Statistical Physics, 60(5–6):585–618, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01025984.

Mariz:1990:CSD

[MHdA90] A. M. Mariz, H. J. Herrmann, and L. de Arcangelis. Comparative study of damage spreading in the Ising model using heatbath, Glauber, and Metropolis dynamics. Journal of Statistical Physics, 59(3–4):1043–1050, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01025861.

Matz:1994:DCE

[MHJ94] R. Matz, D. L. Hunter, and Naeem Jan. The dynamic critical exponent of the three-dimensional Ising model. Journal of Statistical Physics, 74(3–4):903–908, February 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02188583.

Manna:1992:SMD

[MHL92] S. S. Manna, H. J. Herrmann, and D. P. Landau. A stochastic method to determine the shape of a drop on a wall. Journal of Statistical Physics, 66(3–4):1155–1163, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01055723.

Mozos:1994:IGD

[MHM94] J. L. Mozos and A. Hernández-Machado. Interfacial growth in driven Ginzburg–Landau models: Short and long-time dynamics. Journal of Statistical Physics, 74(1–2):131–146, January 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02186810.

Miekisz:1990:MMQ

[Mie90] Jacek Miekisz. A microscopic model with quasicrystalline properties. Journal of Statistical Physics, 58(5–6):1137–1149, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01026568.

Mielke:1991:ODH

[Mie91] A. Mielke. The one-dimensional Hubbard model for large or infinite U. Journal of Statistical Physics, 62(3–4):509–528, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017970.

Mieisz:1993:GME

[Mie93] Jacek Mieisz. The global minimum of energy is not always a sum of local minima — a note on frustration. Journal of Statistical Physics, 71(3–4):425–434, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01058430.

Miekisz:1997:SQG

[Mie97] Jacek Miekisz. Stable quasicrystalline ground states. Journal of Statistical Physics, 88(3–4):691–711, August 1997. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000015168.25151.22.

Miekisz:1998:UFC

[Mie98] Jacek Miekisz. An ultimate frustration in classical lattice-gas models. Journal of Statistical Physics, 90(1–2):285–300, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023264004272.

Miekisz:1999:CLG

[Mie99] Jacek Miekisz. Classical lattice–gas models of quasicrystals. Journal of Statistical Physics, 95(5–6):835–850, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004542115011.

Mikheev:1995:RDC

[Mik95] Lev V. Mikheev. Reentrant dimensional crossover in planar Ising superlattices. Journal of Statistical Physics, 78(1–2):79–101, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183339.

Miller:1991:GOD

[Mil91a] Bruce N. Miller. Gravity in one dimension: Diffusion in acceleration. Journal of Statistical Physics, 63(1–2):291–303, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01026605.

Miller:1991:DDC

[Mil91b] Jeffrey D. Miller. Direction-direction correlations of oriented polymers. Journal of Statistical Physics, 63(1–2):89–113, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01026594.

Milotti:1992:SPR

[Mil92] Edoardo Milotti. Survival probabilities for random walks on lattices with randomly distributed traps. Journal of Statistical Physics, 68(5–6):883–893, September 1992. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048878.

Miller:1995:OOK

[Mil95] Donald G. Miller. The origins of Onsager’s key role in the development of linear irreversible thermodynamics. Journal of Statistical Physics, 78(1–2):563–573, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183365.

Monti:1990:MES

[MJ90] F. Monti and H. R. Jauslin. Master equations for subordinated processes. Journal of Statistical Physics, 60(3–4):413–444, August 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01314929.

Marchetti:1991:PLF

[MK91] Domingos H. U. Marchetti and Abel Klein. Power-law falloff in two-dimensional Coulomb gases at inverse temperature β > 8π. Journal of Statistical Physics, 64(1–2):135–162, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01057871.

Medina:1993:NAC

[MK93] Ernesto Medina and Mehran Kardar. Nonuniversality and analytical continuation in moments of directed polymers on hierarchical lattices. Journal of Statistical Physics, 71(5–6):967–980, June 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049956.

Manna:1990:CSO

[MKK90] S. S. Manna, László B. Kiss, and János Kertész. Cascades and self-organized criticality. Journal of Statistical Physics, 61(3–4): 923–932, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01027312.

Morita:1997:QBS

[MKK97] Yoshifumi Morita, Mahito Kohmoto, and Tohru Koma. Quasibound states of two magnons in the spin-1/2 XXZ chain.

Journal of Statistical Physics, 88(3–4):745–780, August 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000015170.66153.4d.

Marchetti:1990:PLF

[MKP90] Domingos H. U. Marchetti, Abel Klein, and J. Fernando Perez. Power-law falloff in the Kosterlitz–Thouless phase of a twodimensional lattice Coulomb gas. Journal of Statistical Physics, 60(1–2):137–166, July 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01013671. See comments [Mar90a].

Mulder:1991:TMW

[MKP91] B. M. Mulder, C. Krikos, and C. Papatriantafillou. Thermodynamics of a model wih interacting annealed bond impurities on the Bethe lattice. Journal of Statistical Physics, 65(3–4): 423–444, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053737.

Monette:1992:MCS

[MKZ92] L. Monette, W. Klein, and M. Zuckermann. Monte Carlo study of the effect of perturbations on critical droplets. Journal of Statistical Physics, 66(1–2):117–132, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01060062.

Maier:1991:ICP

[ML91a] Rolf Maier and W. G. Laidlaw. Invariants for the critical points in network models of flow in porous media. Journal of Statistical Physics, 62(1–2):269–282, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020870.

Mendes:1991:DIR

[ML91b] J. F. F. Mendes and E. J. S. Lage. Dynamics of the infiniteranged Potts model. Journal of Statistical Physics, 64(3–4): 653–672, August 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048310.

Mertens:1992:CLA

[ML92] Stephan Mertens and Markus E. Lautenbacher. Counting lattice animals: A parallel attack. Journal of Statistical Physics, 66(1– 2):669–678, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01060088.

Mackey:1990:NIG

[MLL90] Michael C. Mackey, André Longtin, and Andrzej Lasota. Noiseinduced global asymptotic stability. Journal of Statistical Physics, 60(5–6):735–751, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01025992.

Meijer:1993:ACL

[MLM93] Paul H. E. Meijer, A. H. M. Levelt, and B. R. Miller. Analysis of the critical line network for the van der Waals equation at the van Laar point. Journal of Statistical Physics, 71(1–2):299–312, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048101.

Macris:1990:IEE

[MM90a] N. Macris and Ph. A. Martin. Ionization equilibrium in the electron-proton gas. Journal of Statistical Physics, 60(5–6): 619–637, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01025985.

Maradudin:1990:TCL

[MM90b] A. A. Maradudin and T. Michel. The transverse correlation length for randomly rough surfaces. Journal of Statistical Physics, 58(3–4):485–501, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01112758.

Mezhlumian:1993:ISP

[MM93] A. Mezhlumian and S. A. Molchanov. Infinite-scale percolation in a new type of branching diffusion process. Journal of Statistical Physics, 71(3–4):799–816, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01058448.

Marcelli:1996:SNR

[MM96] Elisabetta Marcelli and Fabio Martinelli. Some new results on the two-dimensional kinetic Ising model in the phase coexistence region. Journal of Statistical Physics, 84(3–4):655–696, August 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179653.

Machta:1997:CCL

[MM97a] J. Machta and K. Moriarty. The computational complexity of the Lorentz lattice gas. Journal of Statistical Physics, 87(5– 6):1245–1252, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181282.

Matsuoka:1997:LTT

[MM97b] H. Matsuoka and R. F. Martin, Jr. Long-time tails of the velocity autocorrelation functions for the triangular periodic Lorentz gas. Journal of Statistical Physics, 88(1–2):81–103, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02508465.

Maass:1998:CLD

[MM98] Alejandro Maass and Servet Martinez. On Cesàro limit distribution of a class of permutative cellular automata. Journal of Statistical Physics, 90(1–2):435–452, January 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023276306998.

Messager:1992:CPS

[MMSR92] Alain Messager, Salvador Miracle-Sole, and Jean Ruiz. Convexity properties of the surface tension and equilibrium crystals. Journal of Statistical Physics, 67(3–4):449–470, May 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01049716.

Macris:1996:FPC

[MN96] Nicolas Macris and Bruno Nachtergaele. On the flux phase conjecture at half-filling: An improved proof. Journal of Statistical Physics, 85(5–6):745–761, December 1996. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02199361.

Moore:1999:CCS

[MN99] Cristopher Moore and Martin Nilsson. The computational complexity of sandpiles. Journal of Statistical Physics, 96(1–2):205– 224, July 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004524500416.

Mehta:1992:LDV

[MND92] Anita Mehta, R. J. Needs, and Sushanta Dattagupta. The Langevin dynamics of vibrated powders. Journal of Statistical Physics, 68(5–6):1131–1141, September 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048889.

Meurice:1997:OBH

[MNO97] Y. Meurice, S. Niermann, and G. Ordaz. The oscillatory behavior of the high-temperature expansion of Dyson’s hierarchical model: A renormalization group analysis. Journal of Statistical Physics, 87(1–2):363–383, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181492.

Martinelli:1993:SRP

[MO93] F. Martinelli and E. Olivieri. Some remarks on pathologies of renormalization-group transformations for the Ising model. Journal of Statistical Physics, 72(5–6):1169–1177, September 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01048184.

Mironov:1994:DTB

[MO94] A. L. Mironov and V. L. Oleinik. The discrete tight binding approximation. Journal of Statistical Physics, 75(1–2):317–335, April 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186291.

Martinelli:1995:IRG

[MO95] F. Martinelli and E. Olivieri. Instability of renormalization-group pathologies under decimation. Journal of Statistical Physics, 79 (1–2):25–42, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179382.

McCoy:1996:AIC

[MO96a] Barry M. McCoy and William P. Orrick. Analyticity and integrability in the chiral Potts model. Journal of Statistical Physics, 83(5–6):839–865, June 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179547.

Meurice:1996:EAB

[MO96b] Y. Meurice and G. Ordaz. The elusive asymptotic behavior of the high-temperature expansion of the hierarchical Ising model. Journal of Statistical Physics, 82(1–2):343–365, January 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02189234.

Molchan:1995:MAB

[Mol95] G. M. Molchan. Multifractal analysis of Brownian zero set. Journal of Statistical Physics, 79(3–4):701–730, May 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02184877.

Molchan:1997:BES

[Mol97] G. M. Molchan. Burgers equation with self-similar Gaussian initial data: Tail probabilities. Journal of Statistical Physics, 88(5–6):1139–1150, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02732428.

Molchan:1998:AMF

[Mol98] G. M. Molchan. Anomalies in multifractal formalism for local time of Brownian motion. Journal of Statistical Physics, 91 (1–2):199–220, April 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023044221942.

Momoi:1994:LLE

[Mom94] Tsutomu Momoi. Low-lying excited states of quantum antiferromagnets on a triangular lattice. Journal of Statistical Physics, 75(3–4):707–734, May 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186877.

Momoi:1996:QFQ

[Mom96] Tsutomu Momoi. Quantum fluctuations in quantum lattice systems with continuous symmetry. Journal of Statistical Physics, 85(1–2):193–210, October 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02175562.

Monroe:1991:PDI

[Mon91a] James L. Monroe. Phase diagrams of Ising models on Husimi trees. I. Pure multisite interaction systems. Journal of Statistical Physics, 65(1–2):255–268, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01329860.

Monroe:1991:RPD

[Mon91b] James L. Monroe. Restrictions on the phase diagrams for a large class of multisite interaction spin systems. Journal of Statistical Physics, 65(3–4):445–452, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053738.

Monroe:1992:PDI

[Mon92] James L. Monroe. Phase diagrams of Ising models on Husimi trees II. Pair wand multisite interaction systems. Journal of Statistical Physics, 67(5–6):1185–1200, June 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01049014.

Monroe:1994:UBO

[Mon94] James L. Monroe. Upper bounds on T_{c} for one-dimensional Ising systems. Journal of Statistical Physics, 76(5–6):1505– 1510, September 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02187074.

Monroe:1997:CIM

[Mon97] James L. Monroe. Comment on: Ising models on hyperbolic graphs. Journal of Statistical Physics, 88(1–2):513–518, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02508482.

Monteil:2004:CET

[Mon04] Thierry Monteil. A counter-example to the theorem of Hiemer and Snurnikov. Journal of Statistical Physics, 114(5–6):1619– 1623, March 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000013974.81162.20. See

[HS98].

Moore:1997:MVC

[Moo97] Cristopher Moore. Majority-vote cellular automata, Ising dynamics, and p-completeness. Journal of Statistical Physics, 88 (3–4):795–805, August 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000015172.31951.7b.

Morita:1990:CVM

[Mor90] T. Morita. Cluster variation method and Möbius inversion formula. Journal of Statistical Physics, 59(3–4):819–825, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01025852.

Mora:1992:LBP

[Mor92a] Peter Mora. The lattice Boltzmann phononic lattice solid. Journal of Statistical Physics, 68(3–4):591–609, August 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01341765.

Morris:1992:FSSb

[Mor92b] Jacob J. Morris. Finite-size scaling of the interfacial tension. Journal of Statistical Physics, 69(3–4):539–571, November 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01050425.

Meurice:1994:NSH

[MOR94] Y. Meurice, G. Ordaz, and V. G. J. Rodgers. A numerical study of the hierarchical Ising model: High-temperature versus epsilon expansion. Journal of Statistical Physics, 77(3–4): 607–626, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179452.

Meyer-Ortmanns:1997:CPC

[MOR97] Hildegard Meyer-Ortmanns and Thomas Reisz. Critical phenomena with convergent series expansions in a finite volume. Journal of Statistical Physics, 87(3–4):755–798, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02181244.

Madras:1990:MCG

[MOS90a] N. Madras, A. Orlitsky, and L. A. Shepp. Monte Carlo generation of self-avoiding walks with fixed endpoints and fixed length. Journal of Statistical Physics, 58(1–2):159–183, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020290.

Martinelli:1990:MEA

[MOS90b] Fabio Martinelli, Enzo Olivieri, and Elisabetta Scoppola. Metastability and exponential approach to equilibrium for lowtemperature stochastic Ising models. Journal of Statistical Physics, 61(5–6):1105–1119, December 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01014367.

Martinelli:1991:SWDa

[MOS91a] Fabio Martinelli, Enzo Olivieri, and Elisabetta Scoppola. On the Swendsen–Wang dynamics. I. Exponential convergence to equilibrium. Journal of Statistical Physics, 62(1–2):117–133, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020862.

Martinelli:1991:SWDb

[MOS91b] Fabio Martinelli, Enzo Olivieri, and Elisabetta Scoppola. On the Swendsen–Wang dynamics. II. Critical droplets and homogeneous nucleation at low temperature for the two-dimensional Ising model. Journal of Statistical Physics, 62(1–2):135–159, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020863.

Meyer-Ortmanns:1990:STF

[MOT90] Hildegard Meyer-Ortmanns and Thomas Trappenberg. Surface tension from finite-volume vacuum tunneling in the 3D Ising model. Journal of Statistical Physics, 58(1–2):185–198, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020291.

Mountford:1992:GVM

[Mou92] T. S. Mountford. Generalized voter models. Journal of Statistical Physics, 67(1–2):303–311, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01049036.

Mountford:1999:ECF

[Mou99] T. S. Mountford. Existence of a constant for finite system extinction. Journal of Statistical Physics, 96(5–6):1331–1341, September 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004652719999.

Muller:1993:OKQ

[MP93] M. Müller and W. Paul. Ordering kinetics in quasi-one-dimensional Ising-like systems. Journal of Statistical Physics, 73 (1–2):209–233, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01052758.

Martin:1994:ODB

[MP94] Philippe A. Martin and Jaroslaw Piasecki. One-dimensional ballistic aggregation: Rigorous long-time estimates. Journal of Statistical Physics, 76(1–2):447–476, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02188670.

Martin:1996:ADS

[MP96] Philippe A. Martin and Jaroslaw Piasecki. Aggregation dynamics in a self-gravitating one-dimensional gas. Journal of Statistical Physics, 84(3–4):837–857, August 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179659.

Mathieu:1998:MCE

[MP98a] P. Mathieu and P. Picco. Metastability and convergence to equilibrium for the random field Curie–Weiss model. Journal of Statistical Physics, 91(3–4):679–732, May 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023085829152.

Mozyrsky:1998:AD

[MP98b] Dima Mozyrsky and Vladimir Privman. Adiabatic decoherence. Journal of Statistical Physics, 91(3–4):787–799, May 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023042014131.

Manas:1993:SRI

[MPdlR93] M. Mañas, J. M. R. Parrondo, and F. J. de la Rubia. Systemreservoir interaction with stochastic coupling parameters. Journal of Statistical Physics, 71(5–6):1157–1169, June 1993. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01049964.

Magnen:1998:WTI

[MPR98] Jacques Magnen, Gilles Poirot, and Vincent Rivasseau. Wardtype identities for the two-dimensional Anderson model at weak disorder. Journal of Statistical Physics, 93(1–2):331–358, October 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1023/B%3AJOSS.0000026737.08422.fd.pdf.

Mei-Qing:1991:EUS

[MQ91] Zhang Mei-Qing. Explicit unitary schemes to solve quantum operator equations of motion. Journal of Statistical Physics, 65(3– 4):793–799, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053755.

Macris:1994:RDS

[MR94a] Nicolas Macris and Jean Ruiz. A remark on the decay of superconducting correlations in one- and two-dimensional Hubbard models. Journal of Statistical Physics, 75(5–6):1179–1184, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186763.

Michel:1994:SMT

[MR94b] Julien Michel and Raoul Robert. Statistical mechanical theory of the Great Red Spot of jupiter. Journal of Statistical Physics, 77(3–4):645–666, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179454.

Morris:1994:POE

[MR94c] Gary P. Morris and Lamberto Rondoni. Periodic orbit expansions for the Lorentz gas. Journal of Statistical Physics, 75 (3–4):553–584, May 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186872.

Michel:1996:SES

[MR96] Julien Michel and Raoul Robert. Statistical equilibrium states and long-time dynamics for a transport equation. Journal of Statistical Physics, 83(3–4):779–789, May 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183750.

Morriss:1995:DPF

[MRC95] Gary P. Morriss, Lamberto Rondoni, and E. G. D. Cohen. A dynamical partition function for the Lorentz gas. Journal of Statistical Physics, 80(1–2):35–43, July 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02178352.

Meester:1994:NCC

[MRS94] Ronald Meester, Rahul Roy, and Anish Sarkar. Nonuniversality and continuity of the critical covered volume fraction in continuum percolation. Journal of Statistical Physics, 75(1–2):123–134, April 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186282.

Maes:1999:RIM

[MRV99] C. Maes, F. Redig, and A. Van Moffaert. The restriction of the Ising model to a layer. Journal of Statistical Physics, 96 (1–2):69–107, July 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004516331366.

Mirollo:1990:ADA

[MS90a] Renato E. Mirollo and Steven H. Strogatz. Amplitude death in an array of limit-cycle oscillators. Journal of Statistical Physics, 60(1–2):245–262, July 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01013676.

Morrow:1990:SCD

[MS90b] T. J. Morrow and E. R. Smith. Simulation calculation of dielectric constants: Comparison of methods on an exactly solvable model. Journal of Statistical Physics, 61(1–2):187–201, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01013960.

Mazel:1991:RST

[MS91] A. E. Mazel and Yu. M. Suhov. Random surfaces with twosided constraints: An application of the theory of dominant ground states. Journal of Statistical Physics, 64(1–2):111–134, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01057870.

Mountford:1992:ERG

[MS92] Thomas S. Mountford and Aidan Sudbury. An extension of a result of Grannan and Swindle on the poisoning of catalytic surfaces. Journal of Statistical Physics, 67(5–6):1219–1222, June 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01049017. See [GS90c].

Mackey:1993:ETS

[MS93a] Michael C. Mackey and Helmut Schwegler. Ensemble and trajectory statistics in a nonlinear partial differential equation. Journal of Statistical Physics, 70(1–2):281–295, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01053968.

Mahato:1993:LDS

[MS93b] Mangal C. Mahato and Subodh R. Shenoy. Langevin dynamic simulation of hysteresis in a field-swept Landau potential. Journal of Statistical Physics, 73(1–2):123–145, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01052753.

Mandell:1993:BSN

[MS93c] Arnold J. Mandell and Karen A. Selz. Brain stem neuronal noise and neocortical ‘resonance’. Journal of Statistical Physics, 70(1– 2):355–373, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053973.

Mukhopadhyay:1994:SBP

[MS94] Sumit Mukhopadhyay and Muhammad Sahimi. Scaling behavior of permeability and conductivity anisotropy near the percolation threshold. Journal of Statistical Physics, 74(5–6):1301–1308, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02188233.

Miracle-Sole:1995:STS

[MS95a] Salvador Miracle-Sole. Surface tension, step free energy, and facets in the equilibrium crystal. Journal of Statistical Physics, 79(1–2):183–214, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179386.

Mittag:1995:MDP

[MS95b] Laurence Mittag and Michael J. Stephen. Motion and diffusion of a passive scalar in a two-dimensional fluid. Journal of Statistical Physics, 78(1–2):377–387, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183354.

Maier:1996:STB

[MS96a] Robert S. Maier and D. L. Stein. A scaling theory of bifurcations in the symmetric weak-noise escape problem. Journal of Statistical Physics, 83(3–4):291–357, May 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183736.

Majda:1996:BET

[MS96b] Andrew J. Majda and Panagiotis E. Souganidis. Bounds on enhanced turbulent flame speeds for combustion with fractal velocity fields. Journal of Statistical Physics, 83(5–6):933–954, June 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179550.

Mantegna:1997:ESB

[MS97] Rosario N. Mantegna and H. Eugene Stanley. Econophysics: Scaling and its breakdown in finance. Journal of Statistical Physics, 89(1–2):469–479, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02770777.

Magnasco:1998:FRP

[MS98a] Marcelo O. Magnasco and Gustavo Stolovitzky. Feynman’s ratchet and pawl. Journal of Statistical Physics, 93(3–4):615– 632, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000033245.43421.14.

Manning:1998:SEP

[MS98b] Anthony Manning and Károly Simon. A short existence proof for correlation dimension. Journal of Statistical Physics, 90(3–4): 1047–1049, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023253709865.

Maes:1999:HQF

[MS99] Christian Maes and Wolfgang Spitzer. Hydrodynamics for quasi– free quantum systems. Journal of Statistical Physics, 94(5– 6):893–912, March 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004535100763.

Mao:1992:COS

[MSD92] J.-M. Mao, J. Shaw, and J. B. Delos. Closed orbits and semiclassical wavefunctions in two-dimensional Hamiltonian systems. Journal of Statistical Physics, 68(1–2):51–96, July 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01048837.

Montanero:1995:DGT

[MSG95] J. M. Montanero, A. Santos, and V. Garzó. Does the Gaussian thermostat maximize the phase-space compression factor? Journal of Statistical Physics, 81(5–6):989–1005, December 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179300.

Montanero:1997:DFL

[MSG97] J. M. Montanero, A. Santos, and V. Garzó. Distribution function for large velocities of a two-dimensional gas under shear flow. Journal of Statistical Physics, 88(5–6):1165–1181, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02732430.

Mandell:1998:THA

[MSS98a] Arnold J. Mandell, Karen A. Selz, and Michael F. Shlesinger. Transformational homologies in amino acid sequences suggest memberships in protein families. Journal of Statistical Physics, 93(3–4):673–697, November 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033248.97129.3f.

Martiouchev:1998:CSN

[MSS98b] L. M. Martiouchev, V. D. Seleznev, and S. A. Skopinov. Computer simulation of nonequilibrium growth of crystals in a two-dimensional medium with a phase-separating impurity. Journal of Statistical Physics, 90(5–6):1413–1427, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023295815199.

Myshlyavtsev:1990:ETB

[MSZZ90] A. V. Myshlyavtsev, J. L. Sales, G. Zgrablich, and V. P. Zhdanov. The effect of three-body interactions on thermal desorption spectra. Journal of Statistical Physics, 58(5–6):1029–1039, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01026561.

Milshtein:1994:NSD

[MT94] G. N. Milshtein and M. V. Tret’yakov. Numerical solution of differential equations with colored noise. Journal of Statistical Physics, 77(3–4):691–715, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179457.

Marro:1999:NNW

[MTG99] J. Marro, J. J. Torres, and P. L. Garrido. Neural networks in which synaptic patterns fluctuate with time. Journal of Statistical Physics, 94(5–6):837–858, March 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004578915784.

Muche:1996:DPT

[Muc96] Lutz Muche. Distributional properties of the three-dimensional Poisson Delaunay cell. Journal of Statistical Physics, 84(1– 2):147–167, July 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179580.

Mukamel:1991:BRN

[Muk91] Shaul Mukamel. Book review: The nonequilibrium statistical mechanics of open and closed systems. Journal of Statistical Physics, 62(1–2):497–498, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020887.

Muller:1993:ODC

[Mül93] V. F. Müller. One-dimensional chiral models with first-order phase transitions. Journal of Statistical Physics, 70(5–6): 1349–1363, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049437.

Muller:1999:DTC

[Mül99] Stefan Müller. Diffusion of a test chain in a quenched background of semidilute polymers. Journal of Statistical Physics, 96 (1–2):169–203, July 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004572416345.

Munton:1992:QDH

[Mun92] David Munton. Quenched disorder in a hierarchical Coulomb gas model. Journal of Statistical Physics, 68(5–6):1105–1125, September 1992. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/BF01048887.

Murmann:1990:NNG

[Mür90] Michael G. Mürmann. The nearest neighbor gradient system. A rigorous model for a version of the minimal entropy production principle. Journal of Statistical Physics, 59(3–4):827–843, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01025853.

Murphy:1994:DHR

[Mur94] T. J. Murphy. Dynamics of hard rods in one dimension. Journal of Statistical Physics, 74(3–4):889–901, February 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02188582.

Manna:1991:MSF

[MV91] S. S. Manna and T. Vicsek. Multifractality of space-filling bearings and Apollonian packings. Journal of Statistical Physics, 64 (3–4):525–539, August 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048305.

Morbidelli:1997:ERH

[MV97] A. Morbidelli and M. Vergassola. Escape rates in Hamiltonian systems. Journal of Statistical Physics, 89(3–4):549–560, November 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02765535.

Michoel:1999:GBN

[MV99] T. Michoel and A. Verbeure. Goldstone Boson normal coordinates in interacting Bose gases. Journal of Statistical Physics, 96(5–6):1125–1161, September 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004648519091.

Madras:1997:MCS

[MvR97] N. Madras and E. J. Janse van Rensburg. Monte Carlo study of the because -point for collapsing trees. Journal of Statistical Physics, 86(1–2):1–36, January 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02180197.

Momont:1997:ASQ

[MVZ97] B. Momont, A. Verbeure, and V. A. Zagrebnov. Algebraic structure of quantum fluctuations. Journal of Statistical Physics, 89(3–4):633–653, November 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02765539.

Margolina:1990:GRS

[MW90] A. Margolina and H. E. Warriner. Growth in a restricted solid on solid model with correlated noise. Journal of Statistical Physics, 60(5–6):809–821, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01025995.

Mecke:1991:ECR

[MW91] K. R. Mecke and H. Wagner. Euler characteristic and related measures for random geometric sets. Journal of Statistical Physics, 64(3–4):843–850, August 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048319.

Merkl:1994:RRW

[MW94] Franz Merkl and Herbert Wagner. Recurrent random walks and the absence of continuous symmetry breaking on graphs. Journal of Statistical Physics, 75(1–2):153–165, April 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02186284.

Mikosch:1995:MCM

[MW95] T. Mikosch and Qiang Wang. A Monte Carlo method for estimating the correlation exponent. Journal of Statistical Physics, 78(3– 4):799–813, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183688.

Mineev-Weinstein:1995:CMN

[MWA95] M. B. Mineev-Weinstein and F. J. Alexander. Conserved moments in nonequilibrium field dynamics. Journal of Statistical Physics, 79(5–6):1013–1022, June 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02181214.

Mazenko:1994:DNF

[MY94] Gene F. Mazenko and Joonhyun Yeo. Density nonlinearities and a field theory for the dynamics of simple fluids. Journal of Statistical Physics, 74(5–6):1017–1032, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02188215.

Marconi:1990:NSB

[MZ90] Umberto Marini Bettolo Marconi and Yi-Cheng Zhang. Novel scaling behavior of directed polymers: Disorder distribution with long tails. Journal of Statistical Physics, 61(3–4):885– 889, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01027307.

Minlos:1996:ADC

[MZ96] R. A. Minlos and E. A. Zhizhina. Asymptotics of decay of correlations for lattice spin fields at high temperatures. I. The Ising model. Journal of Statistical Physics, 84(1–2):85–118, July 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179578.

Meng-Zhao:1990:ERK

[MZMQ90] Qin Meng-Zhao and Zhang Mei-Qing. Explicit Runge–Kuttalike schemes to solve certain quantum operator equations of motion. Journal of Statistical Physics, 60(5–6):839–844, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01025997.

Ning:1991:NST

[NAC91] Li Ning, Guenter Ahlers, and David S. Cannell. Novel states in Taylor–Couette flow subjected to a Coriolis force. Journal of Statistical Physics, 64(5–6):927–944, September 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01048805.

Nadiga:1995:ESB

[Nad95] B. T. Nadiga. An Euler solver based on locally adaptive discrete velocities. Journal of Statistical Physics, 81(1–2):129–146, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179972.

Nagle:1995:OIB

[Nag95] J. F. Nagle. Onsager, ice, biomembranes, dimer models and the F -model. Journal of Statistical Physics, 78(1–2):549–561, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183364.

Nagahata:1998:GCO

[Nag98] Yukio Nagahata. The gradient condition for one-dimensional symmetric exclusion processes. Journal of Statistical Physics, 91 (3–4):587–602, May 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023025510497.

Nasilowski:1991:CAF

[Nas91] Ralf Nasilowski. A cellular-automaton fluid model with simple rules in arbitrarily many dimensions. Journal of Statistical Physics, 65(1–2):97–138, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01329852.

Nawa:1998:TPB

[Naw98] Hayato Nawa. Two points blow-up in solutions of the nonlinear Schrödinger equation with quartic potential on R. Journal of Statistical Physics, 91(1–2):439–458, April 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023012709647.

Nayak:1993:RGA

[Nay93] Chetan Nayak. A renormalization group analysis of turbulent transport. Journal of Statistical Physics, 71(1–2):129–141, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01048091.

Nieuwenhuizen:1990:DSM

[NB90] Th. M. Nieuwenhuizen and H. Brand. Diffusion and survival in a medium with imperfect traps. Journal of Statistical Physics, 59(1–2):53–72, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01015563.

Newman:1990:IEE

[NBM90] T. J. Newman, A. J. Bray, and A. J. McKane. Inertial effects on the escape rate of a particle driven by colored noise: An instanton approach. Journal of Statistical Physics, 59(1–2):357–369, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01015574.

Nouri:1997:IRP

[Nd97] Anne Nouri and Armel de La Bourdonnaye. Ionization and recombination in plasmas. Journal of Statistical Physics, 87 (3–4):643–660, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181241.

Nemirovsky:1992:TDS

[NDF92] A. M. Nemirovsky, Jacek Dudowicz, and Karl F. Freed. Thermodynamics of a dense self-avoiding walk with contact interactions. Journal of Statistical Physics, 67(1–2):395–412, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01049041.

Nikolova:1995:PTP

[NE95] Albena Nikolova and Dotchi Exerowa. Phase transitions in phosphatidylcholine foam bilayers. Journal of Statistical Physics, 78 (1–2):147–160, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183343.

Nemeth:1991:DBC

[Ném91] R. Németh. Difference between canonical and grand canonical ensembles in discrete lattice gas models. Journal of Statistical Physics, 63(1–2):419–424, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01026615.

Neves:1995:DVP

[Nev95] E. Jordão Neves. A discrete variational problem related to Ising droplets at low temperatures. Journal of Statistical Physics, 80 (1–2):103–123, July 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02178355.

Nemirovsky:1992:MEE

[NFID92] A. M. Nemirovsky, Karl F. Freed, Takao Ishinabe, and Jack F. Douglas. Marriage of exact enumeration and 1/d expansion methods: Lattice model of dilute polymers. Journal of Statistical Physics, 67(5–6):1083–1108, June 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01049010.

Nielaba:1999:GOD

[NFL99] P. Nielaba, P. Fratzl, and J. L. Lebowitz. Growth of ordered domains in a computer model alloy with lattice misfit. Journal of Statistical Physics, 95(1–2):23–43, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004569209618.

Noble:1995:DAL

[NGB95] David R. Noble, John G. Georgiadis, and Richard O. Buckius. Direct assessment of lattice Boltzmann hydrodynamics and boundary conditions for recirculating flows. Journal of Statistical Physics, 81(1–2):17–33, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179965.

Narkounskaia:1992:CSO

[NHT92] Galina Narkounskaia, Jie Huang, and Donald L. Turcotte. Chaotic and self-organized critical behavior of a generalized slider-block model. Journal of Statistical Physics, 67(5–6): 1151–1183, June 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049013.

Nicolis:1993:LTC

[Nic93] C. Nicolis. Long-term climatic transitions and stochastic resonance. Journal of Statistical Physics, 70(1–2):3–14, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01053950.

Niu:1991:QCN

[Niu91] Qian Niu. Quantum coherence of a narrow-band particle interacting with phonons and static disorder. Journal of Statistical Physics, 65(1–2):317–361, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01329863.

Niwa:1997:LEI

[Niw97] Toshio Niwa. On the law of entropy increase of some cellular automata on Z^{d}. Journal of Statistical Physics, 89(3–4): 801–816, November 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02765545.

Nakano:1999:ATU

[NK99] Fumihiko Nakano and Masahiro Kaminaga. Absence of transport under a slowly varying potential in disordered systems. Journal of Statistical Physics, 97(5–6):917–940, December 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004657913118.

Nylund:1993:PDH

[NLT93] Eric Nylund, Katja Lindenberg, and George Tsironis. Proton dynamics in hydrogen-bonded systems. Journal of Statistical Physics, 70(1–2):163–181, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053961.

Neumayr:1999:RFF

[NM99] Arne Neumayr and Walter Metzner. Reduction formula for fermion loops and density correlations of the 1D Fermi gas. Journal of Statistical Physics, 96(3–4):613–626, August 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004546206544.

Noirez:1991:NSS

[NMC^{+}91] L. Noirez, F. Moussa, J. P. Cotton, P. Keller, and G. Pépy. Neutron scattering studies of molecular conformations in liquid crystal polymers. Journal of Statistical Physics, 62(5–6):997–1013, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01128173.

Nakamura:1999:SLR

[NMHS99] K.-I. Nakamura, H. Matano, D. Hilhorst, and R. Schätzle. Singular limit of a reaction–diffusion equation with a spatially inhomogeneous reaction term. Journal of Statistical Physics, 95 (5–6):1165–1185, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004518904533.

Nicolis:1993:SRC

[NNM93] G. Nicolis, C. Nicolis, and D. McKernan. Stochastic resonance in chaotic dynamics. Journal of Statistical Physics, 70(1–2): 125–139, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053958.

Nobbe:1995:CMT

[Nob95] Burkhard Nobbe. Classical motion in two-dimensional crystals. Journal of Statistical Physics, 78(5–6):1591–1605, March 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02180144.

Nolden:1992:ASV

[Nol92] I. M. Nolden. The asymmetric six-vertex model. Journal of Statistical Physics, 67(1–2):155–201, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01049030.

Noonan:1998:NUB

[Noo98] John Noonan. New upper bounds for the connective constants of self-avoiding walks. Journal of Statistical Physics, 91(5–6):871– 888, June 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023023831510.

Nossal:1990:BRD

[Nos90] Ralph Nossal. Book reviews: Dealing with real-world nonlinearities. Journal of Statistical Physics, 58(1–2):403–404, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020304.

Nossal:1993:BRS

[Nos93] Ralph Nossal. Book review: Scattering and localization of classical waves in random media. Journal of Statistical Physics, 71 (5–6):1231–1232, June 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049971.

Nouri:1999:BCS

[NOV99] A. Nouri, A. Omrane, and J. P. Vila. Boundary conditions for scalar conservation laws from a kinetic point of view. Journal of Statistical Physics, 94(5–6):779–804, March 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004574814876.

Nardi:1999:IMS

[NOZ99] F. R. Nardi, E. Olivieri, and M. Zahradník. On the Ising model with strongly anisotropic external field. Journal of Statistical Physics, 97(1–2):87–144, October 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004662917583.

Newhouse:1993:ETE

[NP93] Sheldon Newhouse and Thea Pignataro. On the estimation of topological entropy. Journal of Statistical Physics, 72(5–6): 1331–1351, September 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048189.

Nyberg:1994:DFA

[NP94] A. M. Nyberg and J. K. Percus. Density functional approximations for classical fluids with long-range interactions. Journal of Statistical Physics, 76(1–2):397–418, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02188668.

Newman:1999:EEH

[NP99a] M. E. J. Newman and R. G. Palmer. Error estimation in the histogram Monte Carlo method. Journal of Statistical Physics, 97(5–6):1011–1026, December 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004614130865.

Niethammer:1999:NSS

[NP99b] Barbara Niethammer and Robert L. Pego. Non-self-similar behavior in the LSW theory of Ostwald Ripening. Journal of Statistical Physics, 95(5–6):867–902, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004546215920.

Nickel:1990:HTS

[NR90] Bernie G. Nickel and J. J. Rehr. High-temperature series for scalar-field lattice models: Generation and analysis. Journal of Statistical Physics, 61(1–2):1–50, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01013953.

Na:1998:ECL

[NR98] Kyungsun Na and L. E. Reichl. Electron conductance and lifetimes in a ballistic electron waveguide. Journal of Statistical Physics, 92(3–4):519–542, August 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023032420009.

Nannelli:1992:LBE

[NS92] Francesca Nannelli and Sauro Succi. The lattice Boltzmann equation on irregular lattices. Journal of Statistical Physics, 68(3– 4):401–407, August 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01341755.

Neu:1995:RMF

[NS95] Peter Neu and Roland Speicher. Rigorous mean-field model for coherent-potential approximation: Anderson model with free random variables. Journal of Statistical Physics, 80(5–6):1279– 1308, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179871.

Newman:1996:GSS

[NS96] C. M. Newman and D. L. Stein. Ground-state structure in a highly disordered spin-glass model. Journal of Statistical Physics, 82(3–4):1113–1132, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179805.

Newman:1999:EPS

[NS99] C. M. Newman and D. L. Stein. Equilibrium pure states and nonequilibrium chaos. Journal of Statistical Physics, 94(3–4): 709–722, February 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004516908767.

Nummelin:1994:KMP

[Num94] Esa Nummelin. Kink movements and percolation in the binary additive cellular automaton. Journal of Statistical Physics, 75 (5–6):879–889, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186748.

Namachchivaya:1993:MLE

[NV93] N. Sri Namachchivaya and H. J. Van Roessel. Maximal Lyapunov exponent and rotation numbers for two coupled oscillators driven by real noise. Journal of Statistical Physics, 71(3–4):549–567, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01058437.

Nonnenmacher:1998:CEP

[NV98] S. Nonnenmacher and A. Voros. Chaotic eigenfunctions in phase space. Journal of Statistical Physics, 92(3–4):431–518, August 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1023080303171.

Nguyen:1995:GLC

[NY95] Bao Gia Nguyen and Wei-Shih Yang. Gaussian limit for critical oriented percolation in high dimensions. Journal of Statistical Physics, 78(3–4):841–876, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183691.

Oger:1996:DTD

[OAB^{+}96] Luc Oger, Chrystèle Annic, Daniel Bideau, Rongqing Dai, and Stuart B. Savage. Diffusion of two-dimensional particles on an air table. Journal of Statistical Physics, 82(3–4):1047–1061, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179801.

OBrien:1990:MNS

[O’B90] George L. O’Brien. Monotonicity of the number of self-avoiding walks. Journal of Statistical Physics, 59(3–4):969–979, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01025858.

Ohno:1991:MCS

[OB91a] Kaoru Ohno and Kurt Binder. Monte Carlo simulation of manyarm star polymers in two-dimensional good solvents in the bulk and at a surface. Journal of Statistical Physics, 64(3–4):781–806, August 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048315.

Oshanin:1991:RKP

[OB91b] G. S. Oshanin and S. F. Burlatsky. Reaction kinetics in polymer systems. Journal of Statistical Physics, 65(5–6):1109–1122, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049601.

ORourke:1996:ITC

[OB96] M. J. O’Rourke and R. J. Baxter. Interfacial tension of the chiral Potts model. Journal of Statistical Physics, 82(1–2):1–29, January 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02189223.

ORourke:1995:NRT

[OBB95] M. J. O’Rourke, R. J. Baxter, and V. V. Bazhanov. Numerical results for the three-state critical Potts model on finite rectangular lattices. Journal of Statistical Physics, 78(3–4):665–680, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183683.

OCarroll:1993:LCW

[O’C93a] Michael O’Carroll. Lattice and continuum wavelets and the block renormalization group. Journal of Statistical Physics, 71 (3–4):415–423, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01058429.

OCarroll:1993:MRG

[O’C93b] Michael O’Carroll. Multiscale representation of generating and correlation functions for some models of statistical mechanics and quantum field theory. Journal of Statistical Physics, 73(5– 6):945–958, December 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01052817.

Ordonez:1996:SDT

[OD96] Gonzalo E. Ordóñez and Dean J. Driebe. Spectral decomposition of tent maps using symmetry considerations. Journal of Statistical Physics, 84(1–2):269–276, July 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179586.

Ortiz:1996:SPO

[OdAdA96] J. S. Espinoza Ortiz, M. A. M. de Aguiar, and A. M. Ozorio de Almeida. Scars of periodic orbits in the stadium action billiard. Journal of Statistical Physics, 83(1–2):275–287, April 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02183650.

Oerding:1995:RTF

[Oer95] K. Oerding. Relaxation times in a finite Ising system with random impurities. Journal of Statistical Physics, 78(3–4):893–916, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183693.

Oppenheim:1991:BR

[OG91] Irwin Oppenheim and M. Gitterman. Book review. Journal of Statistical Physics, 64(1–2):465–469, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01057887.

Okun:1990:ECP

[Oku90] B. L. Okun. Euler characteristic in percolation theory. Journal of Statistical Physics, 59(1–2):523–527, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01015581.

Ould-Lemrabott:1997:EBS

[OL97] Mohamed Ould-Lemrabott. Effect of the block-spin configuration on the location of β_{c} in two-dimensional Ising models in twodimensional Ising models. Journal of Statistical Physics, 86(5– 6):1109–1115, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183616.

Oleinik:1990:ABE

[Ole90] V. L. Oleinik. Asymptotic behavior of energy band associated with a negative energy level. Journal of Statistical Physics, 59 (3–4):665–678, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01025845.

Oshanin:1993:SFC

[OMM93] G. Oshanin, A. Mogutov, and M. Moreau. Steady flux in a continuous-space Sinai chain. Journal of Statistical Physics, 73 (1–2):379–388, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01052766.

Omnes:1991:ANT

[Omn91] Roland Omnès. About the notion of truth in quantum mechanics. Journal of Statistical Physics, 62(3–4):841–861, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017986.

Onody:1996:CTT

[ON96] Roberto N. Onody and Ubiraci P. C. Neves. Collapse transition of a two-dimensional lattice animal. Journal of Statistical Physics, 84(3–4):399–413, August 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179649.

Oppenheim:1991:UAL

[OO91] Irwin Oppenheim and Alex Orsky. Uses and abuses of the Langevin equation for chemical reactions in condensed phases. Journal of Statistical Physics, 65(5–6):859–872, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01049586.

Olivieri:1990:CED

[OP90] Enzo Olivieri and Pierre Picco. Cluster expansion for ddimensional lattice systems and finite-volume factorization properties. Journal of Statistical Physics, 59(1–2):221–256, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01015569.

Owczarek:1993:ESD

[OP93] Aleksander L. Owczarek and Thomas Prellberg. Exact solution of the discrete (1+1)-dimensional SOS model with field and surface interactions. Journal of Statistical Physics, 70(5– 6):1175–1194, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049427.

Owczarek:1995:CPI

[OP95] A. L. Owczarek and T. Prellberg. The collapse point of interacting trails in two dimensions from kinetic growth simulations. Journal of Statistical Physics, 79(5–6):951–967, June 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02181210.

Owczarek:1993:TBS

[OPB93] A. L. Owczarek, T. Prellberg, and R. Brak. The tricritical behavior of self-interacting partially directed walks. Journal of Statistical Physics, 72(3–4):737–772, August 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048031.

Olarrea:1995:ESSa

[OPdlR95a] J. Olarrea, J. M. R. Parrondo, and F. J. de la Rubia. Escape statistics for systems driven by dichotomous noise. I. General theory. Journal of Statistical Physics, 79(3–4):669–682, May 1995.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02184875.

Olarrea:1995:ESSb

[OPdlR95b] J. Olarrea, J. M. R. Parrondo, and F. J. de la Rubia. Escape statistics for systems driven by dichotomous noise. II. The imperfect pitchfork bifurcation as a case study. Journal of Statistical Physics, 79(3–4):683–699, May 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02184876.

Oppenheim:1991:BRS

[Opp91] Irwin Oppenheim. Book review: Statistical theory of heat. Journal of Statistical Physics, 62(1–2):499–500, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020888.

Oppenheim:1994:NNTa

[Opp94a] Irwin Oppenheim. Nonlinear nonequilibrium thermodynamics I. Linear and nonlinear fluctuation–dissipation theorems. Journal of Statistical Physics, 77(3–4):949–950, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179473.

Oppenheim:1994:NNTb

[Opp94b] Irwin Oppenheim. Nonlinear nonequilibrium thermodynamics I. Linear and nonlinear fluctuation-dissipation theorems. Journal of Statistical Physics, 77(5–6):1109–1110, December 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02183157.

Oppenheim:1995:CPS

[Opp95] Irwin Oppenheim. Contemporary problems in statistical physics. Journal of Statistical Physics, 78(3–4):1181, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02183714.

Oppenheim:1996:SMT

[Opp96a] Irwin Oppenheim. Statistical mechanics and thermodynamics. Journal of Statistical Physics, 82(1–2):455–456, January 1996.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02189242.

Oppenheim:1996:ST

[Opp96b] Irwin Oppenheim. A survey of thermodynamics. Journal of Statistical Physics, 83(3–4):791–792, May 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183751.

Oppenheim:1997:LEA

[Opp97] Irwin Oppenheim. The Langevin equation with applications in physics, chemistry and electrical engineering. Journal of Statistical Physics, 88(1–2):519–520, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02508483.

Oppenheim:1998:BRBa

[Opp98a] Irwin Oppenheim. Book review: Statistical Mechanics For Chemists. J. Goodisman, Wiley, New York, 1997. Journal of Statistical Physics, 90(3–4):1069, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023209927612.

Oppenheim:1998:BRBb

[Opp98b] Irwin Oppenheim. Book review: Time’s Arrows and Quantum Measurement. L. S. Schulman, Cambridge University Press, Cambridge, 1997. Journal of Statistical Physics, 91(1–2):475, April 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023016910556.

Oppenheim:1998:BRS

[Opp98c] Irwin Oppenheim. Book review: Statistical mechanics: Fundamentals and modem applications. Journal of Statistical Physics, 91(5–6):1061–1062, June 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023092219215.

Olivieri:1993:GSO

[OPS93] E. Olivieri, P. Picco, and Yu. M. Suhov. On the Gibbs states for one-dimensional lattice Boson systems with a long-range interaction. Journal of Statistical Physics, 70(3–4):985–1028, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01053604.

ONeil:1991:LDP

[OR91] Kevin A. O’Neil and Richard A. Redner. On the limiting distribution of pair-summable potential functions in many-particle systems. Journal of Statistical Physics, 62(1–2):399–410, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020875.

Olson:1995:TDI

[OR95a] John F. Olson and Daniel H. Rothman. Three-dimensional immiscible lattice gas: Application to sheared phase separation. Journal of Statistical Physics, 81(1–2):199–222, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179976.

Otnes:1995:EMF

[OR95b] K. Otnes and T. Riste. The effect of magnetic field and convective flow on nematic director fluctuations. Journal of Statistical Physics, 78(1–2):453–462, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183359.

Ord:1992:RFC

[Ord92] G. N. Ord. A reformulation of the Feynman chessboard model. Journal of Statistical Physics, 66(1–2):647–659, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01060086.

Ortuno:1991:NUC

[ORG91] M. Ortuño, J. Ruiz, and J. M. F. Gunn. New universality classes in ‘Percolative’ dynamics. Journal of Statistical Physics, 65(3– 4):453–467, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053739.

Orzechowski:1996:PTS

[Orz96] M. E. Orzechowski. On the phase transition to sheet percolation in random Cantor sets. Journal of Statistical Physics, 82(3–4):

1081–1098, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179803.

Olaussen:1991:NMA

[OS91] Kåre Olaussen and George Stell. New microscopic approach to the statistical mechanics of chemical association. Journal of Statistical Physics, 62(1–2):221–237, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020867.

Olivieri:1995:MCE

[OS95] E. Olivieri and E. Scoppola. Markov chains with exponentially small transition probabilities: First exit problem from a general domain. I. The reversible case. Journal of Statistical Physics, 79 (3–4):613–647, May 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02184873.

Olivieri:1996:MCE

[OS96a] E. Olivieri and E. Scoppola. Markov chains with exponentially small transition probabilities: First exit problem from a general domain. II. The general case. Journal of Statistical Physics, 84(5–6):987–1041, September 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02174126.

Olsen:1996:DDP

[OS96b] Peder Olsen and Renming Song. Diffusion of directed polymers in a strong random environment. Journal of Statistical Physics, 83(3–4):727–738, May 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183745.

OConnor:1997:DCL

[OSB97] Denjoe O’Connor, C. R. Stephens, and A. J. Bray. Dimensional crossover in the large-n limit. Journal of Statistical Physics, 87 (1–2):273–291, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181488.

Ossendrijver:1993:LGS

[OSE93] A. J. H. Ossendrijver, A. Santos, and M. H. Ernst. Lattice gases with static disorder: Renormalization of mean field theory.

Journal of Statistical Physics, 71(5–6):1015–1042, June 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01049959.

OConnor:1992:ASD

[OTH92] Patrick W. O’Connor, Steven Tomsovic, and Eric J. Heller. Accuracy of semiclassical dynamics in the presence of chaos. Journal of Statistical Physics, 68(1–2):131–152, July 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01048839.

Orlandini:1992:MSM

[OTT92] E. Orlandini, M. C. Tesi, and G. Turchetti. Meromorphic structure of the Mellin transforms and short-distance behavior of correlation integrals. Journal of Statistical Physics, 66(1–2): 515–533, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01060078.

Orlandini:1995:TES

[OvR95] E. Orlandini and E. J. Janse van Rensburg. Twist in an exactly solvable directed lattice ribbon. Journal of Statistical Physics, 80 (3–4):781–791, August 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02178554.

Orlandini:1996:MCA

[OvRW96] E. Orlandini, E. J. Janse van Rensburg, and S. G. Whittington. A Monte Carlo algorithm for lattice ribbons. Journal of Statistical Physics, 82(3–4):1159–1198, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179807.

Ohlsen:1991:PLT

[OYSK91] Daniel R. Ohlsen, S. Y. Yamamoto, C. M. Surko, and Paul Kolodner. Pinning and long-time-scale behavior in travelingwave convection. Journal of Statistical Physics, 64(5–6):903– 912, September 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048803.

Ozeki:1993:PDC

[Oze93] Yukiyasu Ozeki. Phase diagram and critical properties of the asymmetric mattis model. Journal of Statistical Physics, 71 (3–4):759–773, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01058446.

Petracchi:1993:PFI

[PAB^{+}93] D. Petracchi, C. Ascoli, M. Barbi, S. Chillemi, M. Pellegrini, and M. Pellegrino. Periodic forcing of ion channel gating: An experimental approach. Journal of Statistical Physics, 70(1–2): 393–401, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053975.

Paetzold:1990:DLG

[Pae90] O. Paetzold. Diffusion of lattice gases without double occupancy on three-dimensional percolation lattices. Journal of Statistical Physics, 61(1–2):495–500, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01013977.

Pajevic:1995:NDC

[Paj95] Sinisa Pajevic. Nonlinear dynamics and chaos. Journal of Statistical Physics, 78(5–6):1635–1636, March 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02180148.

Pajevic:1997:WAT

[Paj97] Sinisa Pajevic. Wavelets: An analysis tool. Journal of Statistical Physics, 86(5–6):1399–1400, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183632.

Palmeri:1990:EST

[Pal90] John Palmeri. Exact solutions to the time-dependent Lorentz gas Boltzmann equation: The approach to hydrodynamics. Journal of Statistical Physics, 58(5–6):885–921, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01026556.

Provatas:1995:SPK

[PANG^{+}95] Nikolas Provatas, Tapio Ala-Nissila, Martin Grant, K. R. Elder, and Luc Piché. Scaling, propagation, and kinetic roughening of flame fronts in random media. Journal of Statistical Physics, 81(3–4):737–759, November 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179255.

Parris:1991:QSA

[Par91] P. E. Parris. Quantum and stochastic aspects of low-temperature trapping and reaction dynamics. Journal of Statistical Physics, 65(5–6):1161–1172, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01049605.

Parisi:1993:BST

[Par93] Giorgio Parisi. On the branching structure of the tree of states in spin glasses. Journal of Statistical Physics, 72(5–6): 857–878, September 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048182.

Patrick:1993:PSS

[Pat93] A. E. Patrick. On phase separation in the spherical model of a ferromagnet: Quasiaverage approach. Journal of Statistical Physics, 72(3–4):665–701, August 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01048028.

Patrick:1994:IEB

[Pat94] A. E. Patrick. The influence of external boundary conditions on the spherical model of a ferromagnet. I. Magnetization profiles. Journal of Statistical Physics, 75(1–2):253–295, April 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02186289.

Patrick:1996:DSK

[Pat96] A. E. Patrick. Dynamics in the Sherrington–Kirkpatrick model. I. The first step. Journal of Statistical Physics, 84(5–6):973– 986, September 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02174125.

Patrick:1998:IBC

[Pat98] A. E. Patrick. The influence of boundary conditions on solidon-solid models. Journal of Statistical Physics, 90(1–2):389–433, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023224322927.

Pearce:1990:RTM

[PB90] Paul A. Pearce and Murray T. Batchelor. Row transfer matrix spectra of cyclic solid-on-solid lattice models. Journal of Statistical Physics, 60(1–2):77–135, July 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01013670.

Puri:1994:SDS

[PB94] Sanjay Puri and Kurt Binder. Surface-directed spinodal decomposition in a thin-film geometry: A computer simulation. Journal of Statistical Physics, 77(1–2):145–172, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02186836.

Prellberg:1995:CEN

[PB95] T. Prellberg and R. Brak. Critical exponents from nonlinear functional equations for partially directed cluster models. Journal of Statistical Physics, 78(3–4):701–730, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02183685.

Parreira:1997:NOG

[PBP97] J. Rodrigo Parreira, O. Bolina, and J. Fernando Perez. Néel order in the ground state of spin-1/2 Heisenberg antiferromagnetic multilayer systems. Journal of Statistical Physics, 86(5–6): 1367–1371, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183628.

Prados:1997:DMC

[PBSR97] A. Prados, J. J. Brey, and B. Sánchez-Rey. A dynamical Monte Carlo algorithm for master equations with timedependent transition rates. Journal of Statistical Physics, 89(3– 4):709–734, November 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02765541.

Privman:1995:ESA

[PCG95] Vladimir Privman, António M. R. Cadilhe, and M. Lawrence Glasser. Exact solutions of anisotropic diffusion-limited reactions with coagulation and annihilation. Journal of Statistical Physics, 81(5–6):881–899, December 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179297.

Penna:1990:FPC

[PdO90] T. J. P. Penna and P. M. C. de Oliveira. Fully parallel code for Monte Carlo simulation. Journal of Statistical Physics, 61(3–4): 933–941, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01027313.

Penna:1992:CEP

[PdO92] T. J. P. Penna and P. M. C. de Oliveira. Culture and evolution on populations of neural networks. Journal of Statistical Physics, 69(5–6):1137–1149, December 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01058767.

Pearce:1995:SMM

[Pea95] Paul A. Pearce. 1994 statistical mechanics/mathematical physics meeting. Journal of Statistical Physics, 79(3–4):793–795, May 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02184887.

Peixoto:1995:MBL

[Pei95] Cláudia Peixoto. Metastable behavior of low-temperature Glauber dynamics with stirring. Journal of Statistical Physics, 80(5–6):1165–1184, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179867.

Penrose:1991:BEC

[Pen91a] O. Penrose. Bose–Einstein condensation in an exactly soluble system of interacting particles. Journal of Statistical Physics, 63 (3–4):761–781, May 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01029210.

Penrose:1991:MFE

[Pen91b] O. Penrose. A mean-field equation of motion for the dynamic Ising model. Journal of Statistical Physics, 63(5–6):975–986, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01029993.

Penrose:1994:SAW

[Pen94a] Mathew D. Penrose. Self-avoiding walks and trees in spreadout lattices. Journal of Statistical Physics, 77(1–2):3–15, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186829.

Penrose:1994:SLT

[Pen94b] Roger Penrose. On the second law of thermodynamics. Journal of Statistical Physics, 77(1–2):217–221, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02186840.

Penna:1995:BSM

[Pen95a] T. J. P. Penna. A bit-string model for biological aging. Journal of Statistical Physics, 78(5–6):1629–1633, March 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02180147.

Penrose:1995:LOR

[Pen95b] Joan Penrose. Lars Onsager remembered. Journal of Statistical Physics, 78(1–2):593–594, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183369.

Penrose:1995:MDR

[Pen95c] O. Penrose. Metastable decay rates, asymptotic expansions, and analytic continuation of thermodynamic functions. Journal of Statistical Physics, 78(1–2):267–283, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183348.

Penrose:1997:BDE

[Pen97] O. Penrose. The Becker-Döring equations at large times and their connection with the LSW theory of coarsening. Journal of Statistical Physics, 89(1–2):305–320, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02770767.

Percus:1990:EOD

[Per90] J. K. Percus. Entropy of a one-dimensional mixed lattice gas. Journal of Statistical Physics, 60(1–2):221–243, July 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013675.

Pereyra:1991:RMM

[Per91] Pedro Pereyra. Random-matrix model for dissipative two-level systems. Journal of Statistical Physics, 65(3–4):773–792, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01053754.

Percus:1993:IRS

[Per93] J. K. Percus. Inhomogeneous random sequential adsorption on a lattice. Journal of Statistical Physics, 71(5–6):1201–1211, June 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01049968.

Percus:1995:CLG

[Per95a] Jerome K. Percus. Classical liquid-gas interface in thermal equilibrium. Journal of Statistical Physics, 78(3–4):1165–1169, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183709.

Pereira:1995:OBS

[Per95b] Emmanuel Pereira. Orthogonality between scales in a renormalization group for fermions. Journal of Statistical Physics, 78(3– 4):1067–1082, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183702.

Poschel:1995:GPD

[PER95c] Thorsten Pöschel, Werner Ebeling, and Helge Rosé. Guessing probability distributions from small samples. Journal of Statistical Physics, 80(5–6):1443–1452, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179880.

Percus:1997:NCF

[Per97] J. K. Percus. Nonuniform classical fluid mixture in onedimensional space with next neighbor interactions. Journal of Statistical Physics, 89(1–2):249–272, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02770764.

Pesin:1993:RMD

[Pes93] Ya. B. Pesin. On rigorous mathematical definitions of correlation dimension and generalized spectrum for dimensions. Journal of Statistical Physics, 71(3–4):529–547, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01058436.

Pettersson:1990:SLB

[Pet90] Rolf Pettersson. On solutions to the linear Boltzmann equation with general boundary conditions and infinite-range forces. Journal of Statistical Physics, 59(1–2):403–440, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01015576.

Pettersson:1993:WSC

[Pet93] Rolf Pettersson. On weak and strong convergence to equilibrium for solutions to the linear Boltzmann equation. Journal of Statistical Physics, 72(1–2):355–380, July 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048054.

Peters:1999:FPE

[Pet99] Michael H. Peters. Fokker–Planck equation, molecular friction, and molecular dynamics for Brownian particle transport near external solid surfaces. Journal of Statistical Physics, 94(3–4): 557–586, February 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004552421971.

Peyriere:1991:TMP

[Pey91] Jacques Peyrière. On the trace map for products of matrices associated with substitutive sequences. Journal of Statistical Physics, 62(1–2):411–414, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020876.

Prestipino:1999:SEL

[PG99] Santi Prestipino and Paolo V. Giaquinta. Statistical entropy of a lattice-gas model: Multiparticle correlation expansion. Journal of Statistical Physics, 96(1–2):135–167, July 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004520432275. See errata [PG00].

Prestipino:2000:ESE

[PG00] Santi Prestipino and Paolo V. Giaquinta. Errata: Statistical Entropy of a Lattice-Gas Model: Multiparticle Correlation Expansion. Journal of Statistical Physics, 98(1–2):507–509, January 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018603728546. See [PG99].

Phillies:1991:CSH

[Phi91] George D. J. Phillies. Convergence of spherical harmonic expansions for the evaluation of hard-sphere cluster integrals. Journal of Statistical Physics, 62(3–4):577–585, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01017974.

Phillips:1994:KES

[Phi94] J. C. Phillips. Kohlrausch explained: The solution to a problem that is 150 years old. Journal of Statistical Physics, 77(3–4): 945–947, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179472.

Provatas:1997:GPC

[PHS^{+}97] N. Provatas, M. Haataja, E. Seppälä, S. Majaniemi, J. Åström, M. Alava, and T. Ala-Nissila. Growth, percolation, and correlations in disordered fiber networks. Journal of Statistical Physics, 87(1–2):385–413, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181493.

Petravic:1994:CDS

[PIM94] Janka Petravic, Dennis J. Isbister, and Gary P. Morriss. Correlation dimension of the sheared hard-disk Lorentz gas. Journal of Statistical Physics, 76(3–4):1045–1063, August 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02188697.

Pinson:1994:CPT

[Pin94] Haru T. Pinson. Critical percolation on the torus. Journal of Statistical Physics, 75(5–6):1167–1177, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02186762.

Piza:1997:DPR

[Piz97] M. S. T. Piza. Directed polymers in a random environment: Some results on fluctuations. Journal of Statistical Physics, 89(3– 4):581–603, November 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02765537.

Priezzhev:1997:MSH

[PK97] V. B. Priezzhev and D. V. Ktitarev. Minimal sandpiles on hexagonal lattice. Journal of Statistical Physics, 88(3–4):781–793, August 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000015171.16291.ae.

Platen:1990:SAH

[Pla90] Eckhard Platen. A stochastic approach to hopping transport in semiconductors. Journal of Statistical Physics, 59(5– 6):1329–1353, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01334754.

Provatas:1991:NIA

[PM91] Nicholas Provatas and Michael C. Mackey. Noise-induced asymptotic periodicity in a piecewise linear map. Journal of Statistical Physics, 63(3–4):585–612, May 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01029201.

Percus:1999:STS

[PM99] Allon G. Percus and Olivier C. Martin. The stochastic traveling salesman problem: Finite size scaling and the cavity prediction. Journal of Statistical Physics, 94(5–6):739–758, March 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1004570713967.

Provata:1994:MAM

[PN94] A. Provata and C. Nicolis. A microscopic aggregation model of droplet dynamics in warm clouds. Journal of Statistical Physics, 74(1–2):75–89, January 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186807.

Prakash:1996:DFR

[PN96] S. Prakash and G. Nicolis. Dynamics of fluctuations in a reactive system of low spatial dimension. Journal of Statistical Physics, 82 (1–2):297–322, January 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02189232.

Prakash:1997:DSM

[PN97] S. Prakash and G. Nicolis. Dynamics of the Schlögl models on lattices of low spatial dimension. Journal of Statistical Physics, 86(5–6):1289–1311, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183624.

Posch:1991:EPU

[PNT91] H. A. Posch, Heide Narnhofer, and W. Thirring. Externally perturbed unstable systems. Journal of Statistical Physics, 65(3– 4):555–578, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053744.

Pereira:1993:OBS

[PO93] Emmanuel Pereira and Michael O’Carroll. Orthogonality between scales and wavelets in a representation for correlation functions. The lattice dipole gas and (Δϕ)^{4} models. Journal of Statistical Physics, 73(3–4):695–721, November 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01054346.

Prellberg:1995:SMV

[PO95] Thomas Prellberg and Aleksander L. Owczarek. Stacking models of vesicles and compact clusters. Journal of Statistical Physics, 80 (3–4):755–779, August 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02178553.

Podgornik:1995:SPM

[Pod95a] Rudolf Podgornik. Statistical physics of macromolecules. Journal of Statistical Physics, 78(3–4):1179–1180, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02183713.

Podgornik:1995:STS

[Pod95b] Rudolf Podgornik. Statistical thermodynamics of surfaces, interfaces, and membranes. Journal of Statistical Physics, 78(3–4): 1175–1177, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183712.

Podgornik:1996:PCM

[Pod96] Rudolf Podgornik. Principles of condensed matter physics. Journal of Statistical Physics, 83(5–6):1263–1265, June 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179565.

Podgornik:1997:SMS

[Pod97] Rudi Podgornik. Statistical mechanics and stability of macromolecules. Journal of Statistical Physics, 87(1–2):459–460, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02181497.

Podgornik:1998:BRBb

[Pod98a] R. Podgornik. Book review: The Casimir Effect and its Applications. V. M. Mostepanko and N. N. Trunov, Clarendon Press, Oxford, 1997. Journal of Statistical Physics, 91(1–2):481–483, April 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023021011464.

Podgornik:1998:BRBa

[Pod98b] Rudolf Podgornik. Book review: The Art of Molecular Dynamics Simulation. D. C. Rapaport, Cambridge University Press, Cambridge, 1995. Journal of Statistical Physics, 90(5–6):1493–1495, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023212318833.

Pokorny:1993:CSG

[Pok93a] Martin Pokorny. Continuous spectrum in the ground state of two spin-1/2 models in the infinite-volume limit. Journal of Statistical Physics, 72(1–2):381–403, July 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048055.

Pokorny:1993:CGS

[Pok93b] Martin Pokorny. Convergence to the ground-state energy in the thermodynamic limit of the Ising model in a strong transverse field. Journal of Statistical Physics, 73(1–2):345–360, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01052764.

Poland:1990:ODK

[Pol90a] Douglas Poland. The one-dimensional kinetic Ising model: A series expansion study. Journal of Statistical Physics, 59(3–4):935– 967, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01025857.

Poland:1990:RCS

[Pol90b] Douglas Poland. Relaxation in cooperative systems: Use of mixture virial coefficients. Journal of Statistical Physics, 61(3–4): 765–802, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01027301. See erratum [Pol91a].

Polewczak:1990:GEG

[Pol90c] Jacek Polewczak. Global existence in L^{1} for the generalized Enskog equation. Journal of Statistical Physics, 59(1–2):461–500, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01015578.

Poland:1991:ERC

[Pol91a] Douglas Poland. Erratum: Relaxation in cooperative systems: Use of mixture virial coefficients. Journal of Statistical Physics, 64(1–2):475, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01057889. See [Pol90b].

Pollicott:1991:NAA

[Pol91b] Mark Pollicott. A note on the Artuso–Aurell–Cvitanovic approach to the Feigenbaum tangent operator. Journal of Statistical Physics, 62(1–2):257–267, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020869.

Pollicott:1992:EMG

[Pol92] Mark Pollicott. Exponential mixing for the geodesic flow on hyperbolic three-manifolds. Journal of Statistical Physics, 67 (3–4):667–673, May 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049724.

Poland:1994:LGA

[Pol94] Douglas Poland. Lattice gas activity series from secular equations. Journal of Statistical Physics, 77(3–4):783–806, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179461.

Pomeau:1993:PBC

[Pom93a] Y. Pomeau. Periodic behavior of cellular automata. Journal of Statistical Physics, 70(5–6):1379–1382, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01049439.

Pompe:1993:MSD

[Pom93b] Bernd Pompe. Measuring statistical dependences in a time series. Journal of Statistical Physics, 73(3–4):587–610, November 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01054341.

Porzio:1990:DSA

[Por90] A. Porzio. The dimension spectrum of axiom a attractors. Journal of Statistical Physics, 58(5–6):923–937, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01026557.

Porra:1996:ISP

[Por96] Josep M. Porrà. An introduction to stochastic processes and nonequilibrium statistical physics. Journal of Statistical Physics, 82 (1–2):457–458, January 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02189243.

Porra:1997:IST

[Por97] Josep M. Porrà. An introduction to statistical thermodynamics. Journal of Statistical Physics, 87(1–2):461–462, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02181498.

Porzio:1998:RMS

[Por98] Anna Porzio. On the regularity of the multifractal spectrum of Bernoulli convolutions. Journal of Statistical Physics, 91 (1–2):17–29, April 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023027718308.

Powell:1991:RSR

[Pow91] Robert L. Powell. Rheology of suspensions of rodlike particles. Journal of Statistical Physics, 62(5–6):1073–1094, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01128178.

Paraskevaidis:1991:FIP

[PP91] C. E. Paraskevaidis and C. Papatriantafillou. Field-induced particle pairing in an Ising system. Journal of Statistical Physics, 64(1–2):385–393, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01057883.

Paiva:1993:HMR

[PP93] Cláudio Paiva and J. Fernando Perez. A hierarchical model for random walks in random media. Journal of Statistical Physics, 71(3–4):435–452, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01058431.

Pomraning:1995:TDC

[PP95] G. C. Pomraning and Anil K. Prinja. Transverse diffusion of a collimated particle beam. Journal of Statistical Physics, 80(3– 4):625–660, August 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02178551.

Puri:1994:POD

[PPD94] Sanjay Puri, Nita Parekh, and Sushanta Dattagupta. Phase ordering dynamics in a gravitational field. Journal of Statistical Physics, 75(5–6):839–857, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02186746.

Procacci:1997:CIS

[PPNM97] Aldo Procacci, Emmanuel Pereira, Armando G. M. Neves, and Domingos H. U. Marchetti. Coulomb interaction symmetries and the Mayer series in the two-dimensional dipole gas. Journal of Statistical Physics, 87(3–4):877–889, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02181248.

Pereira:1999:MFC

[PPO99] Emmanuel Pereira, Aldo Procacci, and Michael O’Carroll. Multiscale formalism for correlation functions of fermions. Infrared analysis of the tridimensional Gross–Neveu model. Journal of Statistical Physics, 95(3–4):665–692, May 1999. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004547426881.

Powell:1990:CET

[PPQ90] Jamie R. Powell, David A. Pink, and Bonnie Quinn. Critical exponents for two-dimensional tracer diffusion through a changing background at concentration c=c_{p}. Journal of Statistical Physics, 60(5–6):729–733, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01025991.

Pini:1993:HRT

[PPR93] D. Pini, A. Parola, and L. Reatto. Hierarchical reference theory of fluids: Application to three-dimensional Ising model. Journal of Statistical Physics, 72(5–6):1179–1201, September 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01048185.

Pasquini:1995:RBL

[PPS95] Michele Pasquini, Giovanni Paladin, and Maurizio Serva. Rigorous bounds of the Lyapunov exponents of the one-dimensional random Ising model. Journal of Statistical Physics, 80(1–2):357– 373, July 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02178363.

Pinn:1994:CHR

[PPW94] K. Pinn, A. Pordt, and C. Wieczerkowski. Computation of hierarchical renormalization-group fixed points and their ϵexpansions. Journal of Statistical Physics, 77(5–6):977–1005, December 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183147.

Privman:1990:NFO

[PR90] V. Privman and J. Rudnick. Nonsymmetric first-order transitions: Finite-size scaling and tests for infinite-range models. Journal of Statistical Physics, 60(5–6):551–560, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01025982.

Payandeh:1994:RSR

[PR94a] B. Payandeh and M. Robert. Real space renormalization group theory of the percolation model. Journal of Statistical Physics, 76(1–2):477–495, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02188671.

Pilyavsky:1994:LDP

[PR94b] Anatoly I. Pilyavsky and Alexei L. Rebenko. The large-deviation principle and the BCS model. Journal of Statistical Physics, 74(5–6):1321–1322, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02188235.

Pra:1994:DNC

[Pra94] Paolo Dai Pra. Detecting nonergodicity in continuous-time spin systems. Journal of Statistical Physics, 76(5–6):1247– 1265, September 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02187061.

Prange:1998:QW

[Pra98] R. E. Prange. Quasiclassical wavefunctions. Journal of Statistical Physics, 93(3–4):965–980, November 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033172.86421.91.

Privman:1992:MCG

[Pri92] Vladimir Privman. Model of cluster growth and phase separation: Exact results in one dimension. Journal of Statistical Physics, 69(3–4):629–642, November 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01050428.

Privman:1993:DCT

[Pri93] Vladimir Privman. Discrete to continuous-time crossover due to anisotropy in diffusion-limited two-particle annihilation reactions. Journal of Statistical Physics, 72(3–4):845–854, August 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048035.

Priezzhev:1994:STD

[Pri94] V. B. Priezzhev. Structure of two-dimensional sandpile. I. Height probabilities. Journal of Statistical Physics, 74(5–6):955–979, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02188212.

Pinn:1999:SIC

[PRW99] K. Pinn, M. Rehwald, and Chr. Wieczerkowski. On the stability of the O(N)-invariant and the cubic-invariant three-dimensional n-component renormalization-group fixed points in the hierarchical approximation. Journal of Statistical Physics, 95(1–2):1–22, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004568925547.

Pandey:1990:MPC

[PS90] R. B. Pandey and D. Stauffer. Metastability with probabilistic cellular automata in an HIV infection. Journal of Statistical Physics, 61(1–2):235–240, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01013962.

Pastur:1991:ASA

[PS91a] L. A. Pastur and M. V. Shcherbina. Absence of self-averaging of the order parameter in the Sherrington–Kirkpatrick model. Journal of Statistical Physics, 62(1–2):1–19, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020856.

Polewczak:1991:NPC

[PS91b] Jacek Polewczak and George Stell. New properties of a class of generalized kinetic equations. Journal of Statistical Physics, 64 (1–2):437–464, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01057886.

Privman:1991:IRM

[PS91c] V. Privman and L. S. Schulman. Infinite-range mean-field percolation: Transfer matrix study of longitudinal correlation length. Journal of Statistical Physics, 64(1–2):207–226, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01057874.

Patrascioiu:1992:CBM

[PS92a] A. Patrascioiu and E. Seiler. Critical behavior in a model of correlated percolation. Journal of Statistical Physics, 69(1–2):55–65, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01053782.

Patrascioiu:1992:PST

[PS92b] A. Patrascioiu and E. Seiler. Phase structure of two-dimensional spin models and percolation. Journal of Statistical Physics, 69(3– 4):573–595, November 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01050426.

Prellberg:1992:MII

[PS92c] Thomas Prellberg and Joseph Slawny. Maps of intervals with indifferent fixed points: Thermodynamic formalism and phase transitions. Journal of Statistical Physics, 66(1–2):503–514, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01060077.

Patrascioiu:1993:PST

[PS93a] A. Patrascioiu and E. Seiler. Phase structure of two-dimensional spin models and percolation. Journal of Statistical Physics, 72 (5–6):1407, September 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048194.

Pisani:1993:LYZ

[PS93b] C. Pisani and E. R. Smith. Lee–Yang zeros and Stokes phenomenon in a model with a wetting transition. Journal of Statistical Physics, 72(1–2):51–78, July 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048040.

Poland:1993:CDO

[PS93c] Douglas Poland and Sandra Song. Cooperative diffusion in onedimensional lattice gases. Journal of Statistical Physics, 71(5– 6):1133–1155, June 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049963.

Pajersky:1994:ICP

[PS94a] Pavol Pajerský and Anton Surda. Incommensurate-commensurate phase transitions in an anisotropic antiferromagnetic model on triangular lattice. Journal of Statistical Physics, 76(5–6):1467– 1477, September 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02187071.

Percus:1994:EFE

[PS94b] J. K. Percus and L. Samaj. Exact free energy functionals for non-simply-connected lattices. Journal of Statistical Physics, 77 (1–2):421–440, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186850.

Pastur:1997:ULE

[PS97a] L. Pastur and M. Shcherbina. Universality of the local eigenvalue statistics for a class of unitary invariant random matrix ensembles. Journal of Statistical Physics, 86(1–2):109–147, January 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02180200.

Patrascioiu:1997:SIP

[PS97b] Adrian Patrascioiu and Erhard Sciler. Super-instantons, perfect actions, finite–size scaling, and the continuum limit. Journal of Statistical Physics, 89(5–6):947–961, December 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02764215.

Prahofer:1997:ESM

[PS97c] M. Prähofer and H. Spohn. An exactly solved model of threedimensional surface growth in the anisotropic KPZ regime. Journal of Statistical Physics, 88(5–6):999–1012, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02732423.

Palczewski:1998:ESC

[PS98] Andrzej Palczewski and Jacques Schneider. Existence, stability, and convergence of solutions of discrete velocity models to the Boltzmann equation. Journal of Statistical Physics, 91(1– 2):307–326, April 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023000406921.

Procacci:1999:SMA

[PS99a] A. Procacci and B. Scoppola. Statistical mechanics approach to coding theory. Journal of Statistical Physics, 96(3–4):907–912, August 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004666811087.

Procacci:1999:PGA

[PS99b] Aldo Procacci and Benedetto Scoppola. Polymer gas approach to n-body lattice systems. Journal of Statistical Physics, 96 (1–2):49–68, July 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004564214528.

Pontin:1994:DAI

[PSP94] L. F. Pontin, J. A. Baêta Segundo, and J. Fernando Perez. Dilute antiferromagnets: Imry–Ma argument, hierarchical model, and equivalence to random field Ising models. Journal of Statistical Physics, 75(1–2):51–65, April 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02186280.

Pastor-Satorras:1998:SSE

[PSR98] Romualdo Pastor-Satorras and Daniel H. Rothman. Scaling of a slope: The erosion of tilted landscapes. Journal of Statistical Physics, 93(3–4):477–500, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033160.59155.c6.

Pastur:1994:RSS

[PST94] L. Pastur, M. Shcherbina, and B. Tirozzi. The replica-symmetric solution without replica trick for the Hopfield model. Journal of Statistical Physics, 74(5–6):1161–1183, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02188221.

Pesheva:1993:MEM

[PSZ93] N. C. Pesheva, Yitzhak Shnidman, and R. K. P. Zia. A maximum entropy mean field method for driven diffusive systems. Journal of Statistical Physics, 70(3–4):737–771, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01053593.

Provata:1993:NCD

[PTN93] A. Provata, J. W. Turner, and G. Nicolis. Nonlinear chemical dynamics in low dimensions: An exactly soluble model. Journal of Statistical Physics, 70(5–6):1195–1213, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01049428.

Phan-Thien:1991:FMS

[PTZG91] N. Phan-Thien, R. Zheng, and A. L. Graham. The flow of a model suspension fluid past a sphere. Journal of Statistical Physics, 62(5–6):1173–1195, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01128182.

Pfister:1995:ASQ

[PV95] Charles-Edouard Pfister and Koen Vande Velde. Almost sure quasilocality in the random cluster model. Journal of Statistical Physics, 79(3–4):765–774, May 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02184883.

Pfister:1997:RCR

[PV97] C.-E. Pfister and Y. Velenik. Random-cluster representation of the Ashkin–Teller model. Journal of Statistical Physics, 88(5–6): 1295–1331, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02732435.

Pule:1994:PFI

[PVZ94] J. V. Pulé, A. Verbeure, and V. A. Zagrebnov. Peierls-Fröhlich instability and Kohn anomaly. Journal of Statistical Physics, 76 (1–2):159–182, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02188659.

Park:1990:CII

[PW90] Hyunggyu Park and Mike Widom. Conformal invariance in incommensurate phases. Journal of Statistical Physics, 61(1– 2):51–78, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01013954.

Pollicott:1994:DSS

[PW94] Mark Pollicott and Howard Weiss. The dimensions of some selfaffine limit sets in the plane and hyperbolic sets. Journal of Statistical Physics, 77(3–4):841–866, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179463.

Pesin:1997:MAE

[PW97] Yakov Pesin and Howard Weiss. A multifractal analysis of equilibrium measures for conformal expanding maps and Moran-like geometric constructions. Journal of Statistical Physics, 86(1– 2):233–275, January 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02180206.

Piguet:1997:DLR

[PWG97] C.-A. Piguet, D. F. Wang, and C. Gruber. Off-diagonal longrange order and Meissner effect for lattice systems. Journal of Statistical Physics, 88(5–6):1363–1369, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02732438.

Papanicolaou:1991:RDF

[PX91] George Papanicolaou and Xue Xin. Reaction–diffusion fronts in periodically layered media. Journal of Statistical Physics, 63 (5–6):915–931, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01029991.

Park:1994:CGS

[PY94] Yong Moon Park and Hyun Jae Yoo. A characterization of Gibbs states of lattice boson systems. Journal of Statistical Physics, 75 (1–2):215–239, April 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186287.

Park:1995:UCP

[PY95] Yong Moon Park and Hyun Jae Yoo. Uniqueness and clustering properties of Gibbs states for classical and quantum unbounded spin systems. Journal of Statistical Physics, 80(1–2):223–271, July 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02178359.

Patrick:1991:PDL

[PZ91] A. E. Patrick and V. A. Zagrebnov. On the parallel dynamics for the Little–Hopfield model. Journal of Statistical Physics, 63 (1–2):59–71, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01026592.

Pechersky:1999:UGS

[PZ99] E. Pechersky and Yu. Zhukov. Uniqueness of Gibbs state for nonideal gas in R^{d}: The case of pair potentials. Journal of Statistical Physics, 97(1–2):145–172, October 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004615001653.

Qian:1992:DSD

[QdL92] Y. H. Qian, D. d’Humières, and P. Lallemand. Diffusion simulation with a deterministic one-dimensional lattice-gas model. Journal of Statistical Physics, 68(3–4):563–573, August 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01341763.

Qian:1995:SDD

[QO95] Y. H. Qian and S. A. Orszag. Scalings in diffusion-driven reaction A + B → C: Numerical simulations by lattice BGK models. Journal of Statistical Physics, 81(1–2):237–253, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179978.

Raab:1998:CBS

[Raa98] Andreas Raab. Comment on ‘A Bit-String Model for Biological Aging’. Journal of Statistical Physics, 91(5–6):1055–1060, June 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023040202377.

Radons:1993:NTP

[Rad93] Günter Radons. A new transition for projections of multifractal measures and random maps. Journal of Statistical Physics, 72 (1–2):227–239, July 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048048.

Radin:1999:SQ

[Rad99] Charles Radin. Symmetries of quasicrystals. Journal of Statistical Physics, 95(5–6):827–833, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004516030941.

Ramsden:1993:RNE

[Ram93] J. J. Ramsden. Review of new experimental techniques for investigating random sequential adsorption. Journal of Statistical Physics, 73(5–6):853–877, December 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01052813.

Ramsden:1995:RSI

[Ram95] J. J. Ramsden. Report on the Second International Workshop on Random Sequential Adsorption. Theory and Experiment. Journal of Statistical Physics, 79(1–2):491–496, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179401.

Rapaport:1990:DCP

[Rap90] D. C. Rapaport. On demonstrating cooperative phenomena. Journal of Statistical Physics, 58(3–4):775–791, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01112774.

Rapaport:1992:CSD

[Rap92a] D. C. Rapaport. Cluster size distribution at criticality. Journal of Statistical Physics, 66(1–2):679–682, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01060089.

Rapaport:1992:BRF

[Rap92b] Dennis C. Rapaport. Book review: Finite size scaling and numerical simulation of statistical systems. Journal of Statistical Physics, 69(5–6):1151, December 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01058768.

Rapaport:1998:BRB

[Rap98] D. C. Rapaport. Book review: Lattice–Gas Cellular Automata. D. H. Rothman and S. Zaleski, Cambridge University Press, Cambridge, 1997. Journal of Statistical Physics, 92(5–6):1209– 1211, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023017332071.

Rasmussen:1993:DBD

[Ras93] H. O. Rasmussen. On the determination of box dimensions by means of wavelet transforms. Journal of Statistical Physics, 71 (3–4):817–823, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01058449.

Ray:1991:ESS

[Ray91] T. S. Ray. Evidence for spinodal singularities in high-dimensional nearest-neighbor Ising models. Journal of Statistical Physics, 62 (1–2):463–472, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020882.

Ray:1994:SOA

[Ray94] T. S. Ray. Self-organization of aging in a population approaching the steady state. Journal of Statistical Physics, 74(3–4):929–939, February 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02188586.

Roy:1990:TSA

[RB90] A. K. Roy and A. Blumen. Theory of self-avoiding walks on percolation fractals. Journal of Statistical Physics, 59(5–6): 1581–1588, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01334765.

Rundle:1991:ORD

[RB91] John B. Rundle and Stephen R. Brown. Origin of rate dependence in frictional sliding. Journal of Statistical Physics, 65(1– 2):403–412, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01329869.

Roberts:1994:TMR

[RB94] John A. G. Roberts and Michael Baake. Trace maps as 3D reversible dynamical systems with an invariant. Journal of Statistical Physics, 74(3–4):829–888, February 1994. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02188581.

Rouault:1995:PSS

[RBB95] Y. Rouault, J. Baschnagel, and K. Binder. Phase separation of symmetrical polymer mixtures in thin-film geometry. Journal of Statistical Physics, 80(5–6):1009–1031, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179862.

Rouet:1993:ODB

[RBF93] J. L. Rouet, F. Blasco, and M. R. Feix. The one-dimensional Boltzmann gas: The ergodic hypothesis and the phase portrait of small systems. Journal of Statistical Physics, 71(1–2):209–224, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048095.

Roncaglia:1992:CQI

[RBGW92] Roberto Roncaglia, Luca Bonci, Paolo Grigolini, and Bruce J. West. Chaos and quantum irreversibility. Journal of Statistical Physics, 68(1–2):321–343, July 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01048848.

Risso:1996:TDG

[RC96] Dino Risso and Patricio Cordero. Two-dimensional gas of disks: Thermal conductivity. Journal of Statistical Physics, 82(5–6): 1453–1466, March 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183391.

Rosenfeld:1997:OMY

[RC97] Yaakov Rosenfeld and Gilles Chabrier. Onsager molecules for the Yukawa potentials: Screening potentials and the Jancovici coefficient in strongly coupled electron-screened plasmas. Journal of Statistical Physics, 89(1–2):283–303, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02770766.

Roy:1990:TPE

[RCB90] A. K. Roy, B. K. Chakrabarti, and A. Blumen. Theta-point exponent for polymer chains on percolation fractals. Journal of Statistical Physics, 61(3–4):903–908, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01027309.

Rodrigues:1994:SGE

[RdO94] Eduardo Soares Rodrigues and Paulo Murilo Castro de Oliveira. Spin-glass energy landscape. Journal of Statistical Physics, 74(5– 6):1265–1272, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02188229.

Ruge:1993:STD

[RDWW93] C. Ruge, S. Dunkelmann, F. Wagner, and J. Wulf. Study of the three-dimensional Ising model on film geometry with the cluster Monte Carlo method. Journal of Statistical Physics, 73(1– 2):293–317, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01052762.

Rebenko:1998:NPR

[Reb98] A. L. Rebenko. A new proof of Ruelle’s superstability bounds. Journal of Statistical Physics, 91(3–4):815–826, May 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023098131878.

Redig:1994:EUB

[Red94] F. Redig. An exponential upper bound for the survival probability in a dynamic random trap model. Journal of Statistical Physics, 74(3–4):815–827, February 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02188580.

Reichl:1993:DKB

[Rei93] L. E. Reichl. A delta-kicked Brownian rotor. Journal of Statistical Physics, 70(1–2):213–228, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053964.

Reimann:1996:NODa

[Rei96a] Peter Reimann. Noisy one-dimensional maps near a crisis. I. Weak Gaussian white and colored noise. Journal of Statistical Physics, 82(5–6):1467–1501, March 1996. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183392.

Reimann:1996:NODb

[Rei96b] Peter Reimann. Noisy one-dimensional maps near a crisis. II. General uncorrelated weak noise. Journal of Statistical Physics, 85(3–4):403–425, November 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02174212.

Reis:1998:AIP

[Rei98] F. D. A. Aarão Reis. Adsorption of ideal polymers on an infinitely ramified fractal. Journal of Statistical Physics, 92(3–4):659–674, August 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023096806805.

Rellick:1991:CHD

[REK91] Lorraine M. Rellick, Gerald A. Edgar, and Michael H. Klapper. Calculating the Hausdorff dimension of tree structures. Journal of Statistical Physics, 64(1–2):77–85, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01057868.

Raphael:1997:PCO

[RGdG97] E. Raphaël, C. Gay, and P. G. de Gennes. Progressive construction of an ‘Olympic’ gel. Journal of Statistical Physics, 89(1– 2):111–118, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02770756.

Raghavan:1997:NTC

[RHA97] R. Raghavan, Christopher L. Henley, and Scott L. Arouh. New two-color dimer models with critical ground states. Journal of Statistical Physics, 86(3–4):517–550, February 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02199112.

Roux:1991:MPH

[RHdS^{+}91] Stéphane Roux, Alex Hansen, Luciano R. da Silva, Liacir S. Lucena, and Ras B. Pandey. Minimal path on the hierarchical diamond lattice. Journal of Statistical Physics, 65(1–2):183–204, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01329855.

Rehberg:1991:MSE

[RHH91] I. Rehberg, F. Horner, and G. Hartung. The measurement of subcritical electroconvection. Journal of Statistical Physics, 64(5–6): 1017–1023, September 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048811.

Richardson:1997:EST

[Ric97] M. J. E. Richardson. Exact solution of two-species ballistic annihilation with general pair-reaction probability. Journal of Statistical Physics, 89(3–4):777–799, November 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02765544.

Rieger:1993:FVA

[Rie93] H. Rieger. Fast vectorized algorithm for the Monte Carlo simulation of the random field Ising model. Journal of Statistical Physics, 70(3–4):1063–1073, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053609.

Ray:1990:NNS

[RK90] T. S. Ray and W. Klein. Nucleation near the spinodal in longrange Ising models. Journal of Statistical Physics, 61(3–4): 891–902, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01027308.

Rundle:1993:SCP

[RK93] John B. Rundle and W. Klein. Scaling and critical phenomena in a cellular automaton slider-block model for earthquakes. Journal of Statistical Physics, 72(1–2):405–412, July 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01048056.

Redner:1996:DEN

[RK96] S. Redner and P. L. Krapivsky. Diffusive escape in a nonlinear shear flow: Life and death at the edge of a windy cliff. Journal of Statistical Physics, 82(3–4):999–1014, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179799.

Redner:1991:SOT

[RL91] S. Redner and F. Leyvraz. Spatial organization in two-species annihilation. Journal of Statistical Physics, 65(5–6):1043–1056, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049597.

Ridgway:1998:ESL

[RLK98] Douglas Ridgway, Herbert Levine, and David A. Kessler. Evolution on a smooth landscape: The role of bias. Journal of Statistical Physics, 90(1–2):191–210, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023203818384.

Ray:1993:NDD

[RM93] T. S. Ray and L. L. Moseley. Nonequilibrium dynamics of a diffusion-limited reaction driven by a cluster-memory mechanism. Journal of Statistical Physics, 73(5–6):879–892, December 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01052814.

Rondoni:1997:APO

[RM97] Lamberto Rondoni and Gary P. Morriss. Applications of periodic orbit theory to N-particle systems. Journal of Statistical Physics, 86(5–6):991–1009, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183611.

Rotstein:1998:GCG

[RNCT98] Horacio G. Rotstein, Amy Novick-Cohen, and Rina Tannenbaum. Gelation and cluster growth with cluster–wall interactions. Journal of Statistical Physics, 90(1–2):119–143, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1023247500637.

Robert:1991:MEP

[Rob91]  Raoul Robert.  A maximum-entropy principle for two-

dimensional perfect fluid dynamics. Journal of Statistical Physics, 65(3–4):531–553, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053743.

Roman:1990:DTD

[Rom90] H. Eduardo Roman. Diffusion in three-dimensional random systems at their percolation thresholds. Journal of Statistical Physics, 58(1–2):375–382, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020299.

Rost:1993:BRL

[Ros93] Hermann Rost. Book review: Large scale dynamics of interacting particles. Journal of Statistical Physics, 73(3–4):803– 806, November 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01054353.

Rothman:1993:OBR

[Rot93] Daniel H. Rothman. From ordered bubbles to random stripes: Pattern formation in a hydrodynamic lattice gas. Journal of Statistical Physics, 71(3–4):641–652, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01058440.

Rowlinson:1997:CPH

[Row97] J. S. Rowlinson. The critical point: A historical introduction to the modern theory of critical phenomena. Journal of Statistical Physics, 87(3–4):957–958, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02181258.

Renals:1990:SND

[RR90] Steve Renals and Richard Rohwer. A study of network dynamics. Journal of Statistical Physics, 58(5–6):825–848, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01026553.

Reis:1993:CBS

[RR93] Fábio D. A. Aarão Reis and R. Riera. Critical behavior of selfavoiding walks on fractals. Journal of Statistical Physics, 71 (3–4):453–470, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01058432.

Robert:1997:MSS

[RR97a] Raoul Robert and Carole Rosier. The modeling of small scales in two-dimensional turbulent flows: A statistical mechanics approach. Journal of Statistical Physics, 86(3–4):481–515, February 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02199111.

Ruijgrok:1997:RDE

[RR97b] Th. Ruijgrok and M. Ruijgrok. A reaction-diffusion equation for a cyclic system with three components. Journal of Statistical Physics, 87(5–6):1145–1164, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02181277.

Romero-Rochin:1997:CSA

[RRGT97] Víctor Romero-Rochin and Enrique González-Tovar. Comments on some aspects of Boltzmann H theorem using reversible molecular dynamics. Journal of Statistical Physics, 89(3–4): 735–749, November 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02765542.

Rodriguez-Romo:1998:CTS

[RRT98] Suemi Rodríguez-Romo and Vladimir Tchijov. On continuoustime self-avoiding random walk in dimension four. Journal of Statistical Physics, 90(3–4):767–781, February 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023276920343.

Richards:1991:RTC

[RS91a] Peter M. Richards and Attila Szabo. Reversible trapping on a cubic lattice: Comparison of theory and simulations. Journal of Statistical Physics, 65(5–6):1085–1093, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01049599.

Rieger:1991:DRM

[RS91b] H. Rieger and M. Schreckenberg. Decay of the remanent magnetization in the asymmetric spin chain. Journal of Statistical Physics, 64(1–2):329–361, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01057880.

Ricci:1992:SMD

[RS92a] T. F. Ricci and C. Scherer. A stochastic model for the dynamics of a classical spin. Journal of Statistical Physics, 67(5– 6):1201–1208, June 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049015.

Rondoni:1992:CRD

[RS92b] L. Rondoni and R. F. Streater. Chemical reactions as dynamical systems on the interval. Journal of Statistical Physics, 66(5– 6):1557–1574, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01054433.

Roumeliotis:1992:PFZ

[RS92c] J. Roumeliotis and E. R. Smith. Partition function zeros for the one-dimensional ordered plasma in Dirichlet boundary conditions. Journal of Statistical Physics, 66(1–2):233–247, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01060066.

Raine:1994:MPH

[RS94] D. J. Raine and D. W. Sciama. The membrane picture of Hawking radiation. Journal of Statistical Physics, 77(1–2):223–228, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186841.

Radons:1996:SMG

[RS96] Günter Radons and Ruedi Stoop. Superpositions of multifractals: Generators of phase transitions in the generalized thermodynamic formalism. Journal of Statistical Physics, 82(3–4): 1063–1080, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179802.

Rebenko:1997:CCE

[RS97a] A. L. Rebenko and G. V. Shchepan’uk. The convergence of cluster expansion for continuous systems with many-body interaction. Journal of Statistical Physics, 88(3–4):665–689, August 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000015167.07226.2e.

Ricci:1997:LRS

[RS97b] T. F. Ricci and C. Scherer. Linear response and stochastic resonance of superparamagnets. Journal of Statistical Physics, 86(3– 4):803–819, February 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02199121.

Riethmuller:1997:LEA

[RSGRP97] Tino Riethmüller, Lutz Schimansky-Geier, Dirk Rosenkranz, and Thorsten Pöschel. Langevin equation approach to granular flow in a narrow pipe. Journal of Statistical Physics, 86(1–2):421–430, January 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02180213.

Roerdink:1990:DLA

[RSL90] J. B. T. M. Roerdink, K. E. Shuler, and G. F. Lawler. Diffusion in lattices with anisotropic scatterers. Journal of Statistical Physics, 59(1–2):23–52, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01015562.

Rajewsky:1998:AEP

[RSSS98] N. Rajewsky, L. Santen, A. Schadschneider, and M. Schreckenberg. The asymmetric exclusion process: Comparison of update procedures. Journal of Statistical Physics, 92(1–2):151–194, July 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023047703307.

Reatto:1991:IDD

[RST91] L. Reatto, G. Stell, and M. Tau. On the isothermal density derivative of G(r) and a new theory of the pair correlation function of hard spheres. Journal of Statistical Physics, 64(3–4): 481–500, August 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048303.

Roy:1998:BDP

[RSW98] Rahul Roy, Anish Sarkar, and Damien G. White. Backbends in directed percolation. Journal of Statistical Physics, 91(5–6):889– 908, June 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023075815581.

Ray:1990:PMI

[RT90] T. S. Ray and P. Tamayo. Properties of metastable Ising models evolving under the Swendsen–Wang dynamics. Journal of Statistical Physics, 60(5–6):851–861, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01025999.

Rubin:1991:ODN

[Rub91] Robert J. Rubin. One-dimensional non-nearest-neighbor random walks in the presence of traps. Journal of Statistical Physics, 65(5–6):1207–1216, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01049607.

Ruelle:1990:TST

[Rue90] David Ruelle. Is there screening in turbulence? Journal of Statistical Physics, 61(3–4):865–868, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01027304.

Ruelle:1996:PEP

[Rue96] David Ruelle. Positivity of entropy production in nonequilibrium statistical mechanics. Journal of Statistical Physics, 85 (1–2):1–23, October 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02175553.

Ruelle:1997:PEP

[Rue97] David Ruelle. Positivity of entropy production in the presence of a random thermostat. Journal of Statistical Physics, 86(5– 6):935–951, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183609.

Ruelle:1999:SDN

[Rue99] David Ruelle. Smooth dynamics and new theoretical ideas in nonequilibrium statistical mechanics. Journal of Statistical Physics, 95(1–2):393–468, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004593915069.

Ruskin:1993:CBS

[Rus93] H. J. Ruskin. Critical behavior of sandpile vortices under variable charge. Journal of Statistical Physics, 73(1–2):389–397, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01052767.

Rust:1994:NIR

[Rus94] Heinrich Rust. A note on inherent replication properties of local cellular automata transition functions. Journal of Statistical Physics, 77(3–4):889–897, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179466.

Rutkevich:1992:ISM

[Rut92] S. B. Rutkevich. Instantons in spherical model thermodynamics. Journal of Statistical Physics, 66(3–4):827–847, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01055704.

Robledo:1991:HSO

[RV91] A. Robledo and C. Varea. The hard-sphere order–disorder transition in the Bethe continuum. Journal of Statistical Physics, 63 (5–6):1163–1176, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01030004.

Robledo:1997:IWS

[RV97] A. Robledo and C. Varea. Interfacial width and shape fluctuations and extensions of the Gaussian model of capillary waves. Journal of Statistical Physics, 89(1–2):273–282, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02770765.

Rosenqvist:1996:ADT

[RVW96] Per E. Rosenqvist, Gábor Vattay, and Andreas Wirzba. Application of the diffraction trace formula to the three-disk scattering system. Journal of Statistical Physics, 83(1–2):243–257, April 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183648.

Requardt:1990:WCR

[RW90] M. Requardt and H. J. Wagner. Wigner crystallization and its relation to the poor decay of pair correlations in one-component plasmas of arbitrary dimension. Journal of Statistical Physics, 58(5–6):1165–1180, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01026570.

Requardt:1991:DTD

[RW91] M. Requardt and H. J. Wagner. Does the three-dimensional capillary wave model lead to a universally valid and pathologyfree description of the liquid-vapor interface near g = 0 ? A controversial point of view. Journal of Statistical Physics, 64 (3–4):807–821, August 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048316.

Ruge:1992:SPC

[RW92] C. Ruge and F. Wagner. On scaling properties of cluster distributions in Ising models. Journal of Statistical Physics, 66 (1–2):99–116, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01060061.

Richter:1996:EQC

[RW96a] S. Richter and R. F. Werner. Ergodicity of quantum cellular automata. Journal of Statistical Physics, 82(3–4):963–998, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179798.

Rolf:1996:HTP

[RW96b] J. Rolf and C. Wieczerkowski. The hierarchical ϕ^{4}-trajectory by perturbation theory in a running coupling and its logarithm by perturbation theory in a running coupling and its logarithm. Journal of Statistical Physics, 84(1–2):119–145, July 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179579.

Rondoni:1993:SSI

[RZ93] Lamberto Rondoni and Paul F. Zweifel. Solutions of singular integral equations from gas dynamics and plasma physics. Journal of Statistical Physics, 70(5–6):1297–1312, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01049433.

Sahimi:1991:CST

[SA91] Muhammad Sahimi and Sepehr Arbabi. On correction to scaling for two- and three-dimensional scalar and vector percolation. Journal of Statistical Physics, 62(1–2):453–461, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020881.

Shussman:1994:DTS

[SA94] Y. Shussman and A. Aharony. Different types of self-avoiding walks on deterministic fractals. Journal of Statistical Physics, 77(3–4):545–563, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179449.

Shussman:1995:SAW

[SA95] Yossi Shussman and Amnon Aharony. Self-avoiding walks on random fractal environments. Journal of Statistical Physics, 80 (1–2):147–167, July 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02178357.

Succi:1995:CLB

[SAB95] Sauro Succi, Giorgio Amati, and Roberto Benzi. Challenges in lattice Boltzmann computing. Journal of Statistical Physics, 81 (1–2):5–16, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179964.

Sachdev:1998:BRB

[Sac98] Subir Sachdev. Book review: Quantum Inverse Scattering Method and Correlation Functions. V E. Korepin, N. M. Bogoliubov, and A. G. Izergin, Cambridge University Press, Cambridge, 1996. Journal of Statistical Physics, 90(5–6):1497–1499, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023264302903.

Strick:1998:PAS

[SACB98] T. R. Strick, J.-F. Allemand, V. Croquette, and D. Bensimon. Physical approaches to the study of DNA. Journal of Statistical Physics, 93(3–4):647–672, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000033247.51868.be.

Saied:1995:CSS

[Sai95] Effat A. Saied. Classification of the similarity solutions of free Kramers equation. Journal of Statistical Physics, 78(3–4): 1139–1146, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183706.

Saied:1996:NES

[Sai96] Effat A. Saied. New exact solutions of heat conduction in metals. Journal of Statistical Physics, 82(3–4):951–962, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179797.

Salas:1995:LTS

[Sal95] J. Salas. Low-temperature series for renormalized operators: The ferromagnetic square-lattice Ising model. Journal of Statistical Physics, 80(5–6):1309–1326, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179872.

Salzano:1999:IMC

[Sal99] Marcia Salzano. Infinitely many contact process transitions on a tree. Journal of Statistical Physics, 97(3–4):817–826, November 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004631712692.

Samols:1995:SMQ

[Sam95] T. M. Samols. A stochastic model of a quantum field theory. Journal of Statistical Physics, 80(3–4):793–809, August 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02178555.

Samoletov:1999:RKP

[Sam99] Alex A. Samoletov. A remark on the Kramers problem. Journal of Statistical Physics, 96(5–6):1351–1357, September 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004656820908.

Sasaki:1992:MPU

[Sas92] Kazuo Sasaki. Modulated phases and upsilon points in a spin model with helical ordering. Journal of Statistical Physics, 68 (5–6):1013–1035, September 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01048883.

Sastry:1995:EC

[Sas95a] Srikanth Sastry. An exploration of chaos. Journal of Statistical Physics, 81(3–4):861–863, November 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179264.

Sastry:1995:PCP

[Sas95b] Srikanth Sastry. Physics and chance. Philosophical issues in the foundations of statistical mechanics. Journal of Statistical Physics, 78(3–4):1185–1186, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183716.

Suchanecki:1994:NMP

[SAT94] Z. Suchanecki, I. Antoniou, and S. Tasaki. Nonlocality of the Misra–Prigogine–Courbage semigroup. Journal of Statistical Physics, 75(5–6):919–928, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02186750.

Sali:1997:EMAa

[SB97a] Leora Sali and David J. Bergman. Effective medium approximation for strongly nonlinear media. Journal of Statistical Physics, 86(3–4):455–479, February 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02199110.

Sali:1997:EMAb

[SB97b] Leora Sali and David J. Bergman. Effective medium approximation for trongly nonlinear media. Journal of Statistical Physics, 89(5–6):1105–1106, December 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02764226.

Suarez:1997:NLG

[SB97c] Alberto Suárez and Jean Pierre Boon. Nonlinear lattice gas hydrodynamics. Journal of Statistical Physics, 87(5–6):1123–1130, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02181275.

Schulz-Baldes:1998:KTQ

[SBB98] H. Schulz-Baldes and J. Bellissard. A kinetic theory for quantum transport in aperiodic media. Journal of Statistical Physics, 91 (5–6):991–1026, June 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023084017398.

Schleier:1992:OAS

[SBH92] W. Schleier, G. Besold, and K. Heinz. Overcoming artificial spatial correlations in simulations of superstructure domain growth with parallel Monte Carlo algorithms. Journal of Statistical Physics, 66(3–4):1101–1122, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01055719.

Schmera:1993:LFP

[SBP^{+}93] Gabor Schmera, Adi Bulsara, David Pierson, Frank Moss, and Enrico Di Cera. Looking at Fokker–Planck dynamics with a noisy instrument. Journal of Statistical Physics, 71(5–6):1179–1190, June 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049966.

Schulz-Baldes:1998:RSA

[SBZ98] H. Schulz-Baldes and M. Zarrouati. Rigorous spectral analysis of the metal–insulator transition in a limit-periodic potential. Journal of Statistical Physics, 91(3–4):801–806, May 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023094030969.

Sobkowicz:1996:IMF

[SC96] Mark Sobkowicz and Bulbul Chakraborty. Ising model with frustration, elasticity, and competing interactions. Journal of Statistical Physics, 83(3–4):739–749, May 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183746.

Stauffer:1999:ICK

[SC99] Dietrich Stauffer and Iksoo Chang. Ising cluster kinetics at the critical point. Journal of Statistical Physics, 95(1–2):503–506, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004502216886.

Soto-Campos:1999:SGL

[SCBIR99] Gerardo Soto-Campos, Richard Bowles, Andrey Itkin, and Howard Reiss. Statistical geometry and lattices. Journal of Statistical Physics, 96(5–6):1111–1123, September 1999. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004696402253.

Schonmann:1990:CPT

[Sch90a] Roberto H. Schonmann. Critical points of two-dimensional bootstrap percolation-like cellular automata. Journal of Statistical Physics, 58(5–6):1239–1244, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01026574.

Schroder:1990:BGL

[Sch90b] Manfred Schröder. On the Bose gas with local mean-field interaction. Journal of Statistical Physics, 58(5–6):1151–1163, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01026569.

Schulz:1992:RGA

[Sch92] Michael Schulz. Renormalization group approach for the sitebond percolation in structured stochastic environments. Journal of Statistical Physics, 67(5–6):1109–1116, June 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01049011.

Schinazi:1993:MPT

[Sch93a] Rinaldo Schinazi. On multiple phase transitions for branching Markov chains. Journal of Statistical Physics, 71(3–4):507–511, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01058434.

Schulte:1993:THC

[Sch93b] J. Schulte. Thermal helium clusters at 3.2 Kelvin in classical and semiclassical simulations. Journal of Statistical Physics, 70(5– 6):1343–1347, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049436.

Schutz:1993:GBA

[Sch93c] G. Schütz. Generalized Bethe ansatz solution of a onedimensional asymmetric exclusion process on a ring with blockage. Journal of Statistical Physics, 71(3–4):471–505, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01058433.

Scherer:1994:EIT

[Sch94a] Leopoldo Garcia-Colin Scherer. Extended irreversible thermodynamcis. Journal of Statistical Physics, 75(3–4):773–774, May 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02186882.

Schinazi:1994:ACP

[Sch94b] Rinaldo Schinazi. The asymmetric contact process on a finite set. Journal of Statistical Physics, 74(5–6):1005–1016, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02188214.

Schulman:1994:SSS

[Sch94c] L. S. Schulman. Special states in the spin-boson model. Journal of Statistical Physics, 77(3–4):931–944, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179471.

Schaertl:1995:BDC

[Sch95a] W. Schaertl. Brownian dynamics of colloidal hard spheres. 3. Extended investigations at the phase transition regime. Journal of Statistical Physics, 79(1–2):299–312, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179391.

Schutz:1995:RDP

[Sch95b] Gunter M. Schütz. Reaction-diffusion processes of hard-core particles. Journal of Statistical Physics, 79(1–2):243–264, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02179389.

Schinazi:1996:CPS

[Sch96a] Rinaldo B. Schinazi. A contact process with a single inhomogeneous site. Journal of Statistical Physics, 83(3–4):767–777, May 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183749.

Schuss:1996:NTR

[Sch96b] Zeev Schuss. New trends in reaction rate theory. Journal of Statistical Physics, 84(5–6):1387, September 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02174138.

Schutz:1997:DRA

[Sch97a] Gunter M. Schütz. Duality relations for asymmetric exclusion processes. Journal of Statistical Physics, 86(5–6):1265–1287, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183623.

Schutz:1997:ESM

[Sch97b] Gunter M. Schütz. Exact solution of the master equation for the asymmetric exclusion process. Journal of Statistical Physics, 88 (1–2):427–445, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02508478.

Schonmann:1998:TCC

[Sch98] Roberto H. Schonmann. The triangle condition for contact processes on homogeneous trees. Journal of Statistical Physics, 90(5–6):1429–1440, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023247932037.

Schinazi:1999:SDR

[Sch99] Rinaldo B. Schinazi. On the spread of drug-resistant diseases. Journal of Statistical Physics, 97(1–2):409–417, October 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004635606196.

Soto-Campos:1996:ARP

[SCM96] Gerardo Soto-Campos and Robert M. Mazo. Asymptotic results for a persistent diffusion model of Taylor dispersion of particles. Journal of Statistical Physics, 85(1–2):165–177, October 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02175560.

Scoppola:1993:RGM

[Sco93] Elisabetta Scoppola. Renormalization group for Markov chains and application to metastability. Journal of Statistical Physics, 73(1–2):83–121, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01052752.

Schutz:1993:PTE

[SD93] G. Schütz and E. Domany. Phase transitions in an exactly soluble one-dimensional exclusion process. Journal of Statistical Physics, 72(1–2):277–296, July 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048050.

Shan:1995:MLB

[SD95] Xiaowen Shan and Gary Doolen. Multicomponent lattice-Boltzmann model with interparticle interaction. Journal of Statistical Physics, 81(1–2):379–393, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179985.

Schoenborn:1999:KPO

[SD99] Oliver Schoenborn and Rashmi C. Desai. Kinetics of phase ordering on curved surfaces. Journal of Statistical Physics, 95 (5–6):949–979, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004502500899.

Simone:1996:EGS

[SDJ^{+}96] C. De Simone, M. Diehl, M. Jünger, P. Mutzel, G. Reinelt, and G. Rinaldi. Exact ground states of two-dimensional ±J Ising spin glasses. Journal of Statistical Physics, 84(5–6):1363– 1371, September 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02174135.

Selke:1997:MCM

[Sel97] Walter Selke. Monte Carlo and molecular dynamics of condensed matter systems. Journal of Statistical Physics, 87(3–4):959–960, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02181259.

Sen:1992:GCH

[Sen92] P. Sen. Growth of correlation in the Hopfield model. Journal of Statistical Physics, 67(1–2):413–417, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01049042.

Serinko:1996:ETA

[Ser96] Regis J. Serinko. Ergodic theorems arising in correlation dimension estimation. Journal of Statistical Physics, 85(1–2):25–40, October 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02175554.

Serva:1998:IML

[Ser98] Maurizio Serva. 2D Ising model with layers of quenched spins. Journal of Statistical Physics, 91(1–2):31–45, April 1998.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023079702378.

Saied:1999:ASA

[SER99] Effat A. Saied and Reda G. Abd El-Rahman. Analytic solutions for asymmetric model of a rod in a lattice fluid. Journal of Statistical Physics, 94(3–4):639–652, February 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004560723788.

Sewell:1990:DLR

[Sew90] Geoffrey L. Sewell. Off-diagonal long-range order and the Meissner effect. Journal of Statistical Physics, 61(1–2):415–422, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01013973.

Saied:1998:DLG

[SEW98] Effat A. Saied and S. A. El-Wakil. On the diffusion in a lattice gas model: Group-theoretic approach. Journal of Statistical Physics, 90(1–2):301–310, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023216121110.

Sykes:1991:LAS

[SF91] M. F. Sykes and Sylvia Flesia. Lattice animals: Supplementation of perimeter polynomial data by graph-theoretic methods. Journal of Statistical Physics, 63(3–4):487–489, May 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01029196.

Schimansky-Geier:1993:PLE

[SGH93] L. Schimansky-Geier and H. Herzel. Positive Lyapunov exponents in the Kramers oscillator. Journal of Statistical Physics, 70 (1–2):141–147, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053959.

Sese:1990:DAM

[SGP90] G. Sesé, E. Guàrdia, and J. A. Padró. On the description of atomic motions in dense fluids by the generalized Langevin equation: statistical properties of random forces. Journal of Statistical Physics, 60(3–4):501–518, August 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01314933.

Stovneng:1989:BLM

[SH89] J. A. Støvneng and E. H. Hauge. The Büttiker–Landauer model generalized. Journal of Statistical Physics, 57(3–4):841– 863, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01022836. See response [BL90].

Shneidman:1995:MRN

[SH95] V. A. Shneidman and P. Hänggi. Microscopic reversibility and the nonlinear Einstein–Onsager relation in macroscopic description of nucleation. Journal of Statistical Physics, 78(1–2): 431–439, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183357.

Shapirov:1995:OCR

[Sha95] Felix Shapirov. Onsager–Casimir reciprocity relations for a mixture of rarefied gases interacting with a laser radiation. Journal of Statistical Physics, 78(1–2):413–430, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183356.

Straty:1991:SAN

[SHG91] G. C. Straty, H. J. M. Hanley, and C. J. Glinka. Shearing apparatus for neutron scattering studies on fluids: Preliminary results for colloidal suspensions. Journal of Statistical Physics, 62(5– 6):1015–1023, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01128174.

Shigematsu:1990:STA

[Shi90a] H. Shigematsu. Statistical-thermodynamic approach to a chaotic dynamical system: Exactly solvable examples. Journal of Statistical Physics, 59(1–2):257–297, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01015570.

Shiino:1990:SAD

[Shi90b] Masatoshi Shiino. Stochastic analyses of the dynamics of generalized Little–Hopfield–Hemmen type neural networks. Journal of Statistical Physics, 59(3–4):1051–1075, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01025862.

Shigematsu:1992:PTT

[Shi92] H. Shigematsu. Phase transitions in thermodynamics of a local Lyapunov exponent for fully-developed chaotic systems. Journal of Statistical Physics, 66(3–4):727–754, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01055698.

Shigematsu:1993:ABF

[Shi93a] H. Shigematsu. Asymptotic behavior of fluctuations for the 1D Ising model in zero-temperature limit. Journal of Statistical Physics, 71(5–6):981–1002, June 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01049957.

Shizume:1993:DOA

[Shi93b] Kousuke Shizume. The decrease in the overall algorithmic complexity of the spin-echo effect. Journal of Statistical Physics, 70(3–4):1057–1062, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01053608.

Shirai:1998:FDR

[Shi98] Tomoyuki Shirai. A factorization of determinant related to some random matrices. Journal of Statistical Physics, 90(5–6): 1449–1459, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023252000178.

Shlesinger:1990:BRD

[Shl90] Michael F. Shlesinger. Book review: Dynamical processes in condensed molecular systems. Journal of Statistical Physics, 59 (3–4):1089–1090, May 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01025865.

Shlesinger:1991:BRB

[Shl91] Michael F. Shlesinger. Book review: Schrödinger: Life and thought. Journal of Statistical Physics, 62(3–4):877–878, February 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017988.

Shukla:1993:TDH

[Shu93] Prabodh Shukla. Theory of the dynamics of the Hopfield model of associative memory. Journal of Statistical Physics, 71(3– 4):705–717, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01058443.

Sridhar:1992:MEC

[SHW92] S. Sridhar, D. O. Hogenboom, and Balam A. Willemsen. Microwave experiments on chaotic billiards. Journal of Statistical Physics, 68(1–2):239–258, July 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01048844.

Segundo:1994:QSS

[SHW94] J. A. Baeta Segundo, Heyder Hey, and Walter F. Wreszinski. On quantum stability for systems under quasiperiodic perturbations. Journal of Statistical Physics, 76(5–6):1479–1493, September 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02187072.

Szalma:1999:TDD

[SI99] Ferenc Szalma and Ferenc Iglói. Two-dimensional dilute Ising models: Critical behavior near defect lines. Journal of Statistical Physics, 95(3–4):759–766, May 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004555728698.

Sidharth:1999:AF

[Sid99] B. G. Sidharth. Anomalous fermions. Journal of Statistical Physics, 95(3–4):775–784, May 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004559829607.

Simpelaere:1994:DSAa

[Sim94a] Dominique Simpelaere. Dimension spectrum of axiom a diffeomorphisms. I. The Bowen–Margulis measure. Journal of Statistical Physics, 76(5–6):1329–1358, September 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02187065.

Simpelaere:1994:DSAb

[Sim94b] Dominique Simpelaere. Dimension spectrum of axiom a diffeomorphisms. II. Gibbs measures. Journal of Statistical Physics, 76(5–6):1359–1375, September 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02187066.

Simpelaere:1994:RRT

[Sim94c] Dominique Simpelaere. Recurrence and return times of the Sierpinski carpet. Journal of Statistical Physics, 77(5–6):1099– 1103, December 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183155.

Simonis:1996:MDC

[Sim96] Adilson Simonis. Metastability of the d-dimensional contact process. Journal of Statistical Physics, 83(5–6):1225–1239, June 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02179561.

Simpelaere:1998:CD

[Sim98] Dominique Simpelaere. Correlation dimension. Journal of Statistical Physics, 90(1–2):491–509, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023232624745.

Sinai:1991:TRC

[Sin91] Ya. G. Sinai. Two results concerning asymptotic behavior of solutions of the Burgers equation with force. Journal of Statistical Physics, 64(1–2):1–12, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01057866.

Sinai:1999:SRW

[Sin99] Ya. G. Sinai. Simple random walks on tori. Journal of Statistical Physics, 94(3–4):695–708, February 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004564824697.

Schlijper:1990:VAD

[SK90] Antoine G. Schlijper and Ryoichi Kikuchi. A variational approach to distribution function theory. Journal of Statistical Physics, 61(1–2):143–160, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01013957.

Suematsu:1998:ECP

[SK98] Kazumi Suematsu and Minoru Kohno. Estimation of critical point in branching reactions: Further examination of gel point formula. Journal of Statistical Physics, 93(1–2):293–305, October 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1023/B%3AJOSS.0000026735.03546.a8.pdf.

Seppalainen:1999:HPF

[SK99] Timo Seppäläinen and Joachim Krug. Hydrodynamics and platoon formation for a totally asymmetric exclusion model with particlewise disorder. Journal of Statistical Physics, 95(3–4):525– 567, May 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1007535124155.

Schmidt:1991:IFM

[SL91] K. E. Schmidt and Michael A. Lee. Implementing the Fast Multipole Method in three dimensions. Journal of Statistical Physics, 63(5–6):1223–1235, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01030008. See erra-

tum [SL97a].

Simon:1993:ESB

[SL93a] Adam Simon and Albert Libchaber. Escape and synchronization of a Brownian particle. Journal of Statistical Physics, 70 (1–2):423, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053978.

Stillinger:1993:CAI

[SL93b] Frank H. Stillinger and Boris D. Lubachevsky. Crystallineamorphous interface packings for disks and spheres. Journal of Statistical Physics, 73(3–4):497–514, November 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01054337.

Stillinger:1995:PBS

[SL95] Frank H. Stillinger and Boris D. Lubachevsky. Patterns of broken symmetry in the impurity-perturbed rigid-disk crystal. Journal of Statistical Physics, 78(3–4):1011–1026, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02183698.

Schmidt:1997:EIF

[SL97a] K. E. Schmidt and Michael A. Lee. Erratum: Implementing the fast multipole method in three dimensions. Journal of Statistical Physics, 87(3–4):955, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02181257. See [SL91].

Schmidt:1997:MES

[SL97b] K. E. Schmidt and Michael A. Lee. Multipole Ewald sums for the fast multipole method. Journal of Statistical Physics, 89(1– 2):411–424, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02770773.

Slemrod:1996:MFF

[Sle96] M. Slemrod. Metastable fluid flow described via a discretevelocity coagulation-fragmentation model. Journal of Statistical Physics, 83(5–6):1067–1108, June 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179553.

Slemrod:1998:RCE

[Sle98] M. Slemrod. Renormalization of the Chapman–Enskog expansion: Isothermal fluid flow and Rosenau saturation. Journal of Statistical Physics, 91(1–2):285–305, April 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023048322851.

Shih:1991:ACP

[SLSA91] Wan Y. Shih, Jun Liu, Wei-Heng Shih, and Ilhan A. Aksay. Aggregation of colloidal particles with a finite interparticle attraction energy. Journal of Statistical Physics, 62(5–6):961–984, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01128171.

Strogatz:1991:SIP

[SM91] Steven H. Strogatz and Renato E. Mirollo. Stability of incoherence in a population of coupled oscillators. Journal of Statistical Physics, 63(3–4):613–635, May 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01029202.

Schoenmaker:1999:NAR

[SM99] Wim Schoenmaker and Wim Magnus. Non-Abelian random polygons: A new model in statistical physics. Journal of Statistical Physics, 94(3–4):389–413, February 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004540119245.

Soler:1992:IKM

[SMD92] Mario Soler, Froilán C. Martínez, and José M. Donoso. Integral kinetic method for one dimension: The spherical case. Journal of Statistical Physics, 69(3–4):813–835, November 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01050435.

Smith:1990:ZFS

[Smi90] E. R. Smith. Zeros of the finite-size scaling region partition function for a model with a wetting transition. Journal of Statistical Physics, 60(5–6):529–549, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01025981.

Smith:1994:CPS

[Smi94] E. R. Smith. Calculating the pressure in simulations using periodic boundary conditions. Journal of Statistical Physics, 77(1– 2):449–472, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186852.

Smith:1995:FP

[Smi95] T. G. Smith, Jr. Fractal physiology. Journal of Statistical Physics, 80(5–6):1453–1454, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179881.

Sergeev:1996:VFB

[SMS96] S. M. Sergeev, V. V. Mangazeev, and Yu. G. Stroganov. The vertex formulation of the Bazhanov–Baxter model. Journal of Statistical Physics, 82(1–2):31–49, January 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02189224.

Snider:1990:DCQ

[Sni90] R. F. Snider. A density-corrected quantum Boltzmann equation. Journal of Statistical Physics, 61(1–2):443–465, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013975.

Snider:1991:RDG

[Sni91] R. F. Snider. A reinterpretation of dense gas kinetic theory. Journal of Statistical Physics, 63(3–4):707–718, May 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01029207.

Snider:1995:CBK

[Sni95] R. F. Snider. Conversion between kinetic energy and potential energy in the classical nonlocal Boltzmann equation. Journal of Statistical Physics, 80(5–6):1085–1117, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179865.

Schor:1991:CFG

[SO91] Ricardo Schor and Michael O’Carroll. Correlation functions and the Goldstone picture for the hierarchical classical vector model at low temperatures in three or more dimensions. Journal of Statistical Physics, 64(1–2):163–191, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01057872.

Suematsu:1992:DCSa

[SO92a] Kazumi Suematsu and Tosihiko Okamoto. Distribution of cyclic species in network formation: Microscopic theory of branching process in _{A}g-R-B_{f−g} model. Journal of Statistical Physics, 66 (1–2):661–668, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01060087.

Suematsu:1992:DCSb

[SO92b] Kazumi Suematsu and Tosihiko Okamoto. Distribution of cyclic species in network formation: Microscopic theory of branching processes. Journal of Statistical Physics, 66(3–4):797–802, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01055702.

Soskin:1999:LFM

[Sos99] S. M. Soskin. Large fluctuations in multiattractor systems and the generalized Kramers problem. Journal of Statistical Physics, 97(3–4):609–676, November 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004663224988.

Segel:1991:EDT

[SP91] Lee A. Segel and Alan S. Perelson. Exploiting the diversity of time scales in the immune system: A B-cell antibody model. Journal of Statistical Physics, 63(5–6):1113–1131, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01030002.

Samaj:1993:NCM

[SP93] L. Samaj and J. K. Percus. New collective modes of interaction nature in inhomogeneous Ising networks. Journal of Statistical Physics, 73(1–2):235–251, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01052759.

Sinkovits:1994:CSR

[SP94] Robert S. Sinkovits and Ras B. Pandey. Computer simulation of random sequential adsorption of two interacting species on a lattice. Journal of Statistical Physics, 74(1–2):457–463, January 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186821.

Samaj:1995:FRA

[SP95a] L. Samaj and J. K. Percus. A functional relation among the pair correlations of the two-dimensional one-component plasma. Journal of Statistical Physics, 80(3–4):811–824, August 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02178556.

Samaj:1995:IGD

[SP95b] L. Samaj and J. K. Percus. Inhomogeneous Glauber dynamics and the process of crystallization of a lattice gas. Journal of Statistical Physics, 78(1–2):495–512, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183361.

Solvason:1997:EEF

[SP97] Dorth Sølvason and Henrik G. Petersen. Error estimates for the fast multipole method. Journal of Statistical Physics, 86(1– 2):391–420, January 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02180212.

Speis:1991:IAO

[Spe91] Athanasios Speis. Instability of the anomalies in the onedimensional Anderson model at weak disorder. Journal of Statistical Physics, 63(3–4):541–565, May 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01029199.

Speer:1993:AAS

[Spe93] Eugene R. Speer. Asymmetric abelian sandpile models. Journal of Statistical Physics, 71(1–2):61–74, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01048088.

Speer:1997:FDR

[Spe97] Eugene R. Speer. Finite-dimensional representations of a shock algebra. Journal of Statistical Physics, 89(1–2):169–175, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02770759.

Spohn:1990:TDL

[Spo90] Herbert Spohn. Tracer diffusion in lattice gases. Journal of Statistical Physics, 59(5–6):1227–1239, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01334748.

Spohn:1992:BRT

[Spo92] Herbert Spohn. Book review: Time’s arrow; the origin of thermodynamic behavior. Journal of Statistical Physics, 69(3–4): 905–906, November 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01050442.

Spohn:1993:IMM

[Spo93] Herbert Spohn. Interface motion in models with stochastic dynamics. Journal of Statistical Physics, 71(5–6):1081–1132, June 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01049962.

Spohn:1995:DCS

[Spo95] Herbert Spohn. Disorder and competition in soluble lattice models. Journal of Statistical Physics, 78(3–4):1173, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02183711.

Spohn:1996:SDS

[Spo96] Herbert Spohn. Statistical dynamics, a stochastic approach to nonequilibrium thermodynamics. Journal of Statistical Physics, 85(3–4):521–522, November 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02174220.

Shukla:1991:TIP

[SPR91] K. P. Shukla, B. Payandeh, and M. Robert. Theory of interfacial phase transitions in surfactant systems. Journal of Statistical Physics, 63(5–6):1053–1075, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01029999.

Sen:1990:LPP

[SR90] P. Sen and P. Ray. Longest path in percolating hierarchical lattice. Journal of Statistical Physics, 59(5–6):1573–1580, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01334764.

Siegel:1993:MTD

[SR93] Ralph M. Siegel and Heather L. Read. Models of the temporal dynamics of visual processing. Journal of Statistical Physics, 70 (1–2):297–308, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053969.

Sahimi:1995:PSR

[SR95] Muhammad Sahimi and Hossein Rassamdana. On position-space renormalization group approach to percolation. Journal of Statistical Physics, 78(3–4):1157–1164, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183708.

Stevens:1993:TCB

[SRC93] John G. Stevens, Ronald E. Rosensweig, and A. E. Cerkanowicz. Transient and cyclic behavior of cellular automata with null boundary conditions. Journal of Statistical Physics, 73(1– 2):159–174, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01052755.

Sreenivasan:1998:BRB

[Sre98] Katepalli R. Sreenivasan. Book review: Turbulence, Coherent Structures, Dynamical Systems, and Symmetry. P. Holmes, J. L. Lumley, and G. Berkooz, Cambridge University Press, New York, 1996, pp. 420. Journal of Statistical Physics, 92(1–2):333– 334, July 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023012122871.

Schoolderman:1990:VMI

[SS90] A. J. Schoolderman and L. G. Suttorp. On the validity of magnetohydrodynamics for ionic mixtures. Journal of Statistical Physics, 58(5–6):997–1027, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01026560.

Schlosser:1992:SSF

[SS92a] Till Schlösser and Herbert Spohn. Sample-to-sample fluctuations in the conductivity of a disordered medium. Journal of Statistical Physics, 69(5–6):955–967, December 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01058757.

Schurrer:1992:ESB

[SS92b] F. Schürrer and M. Schaler. Exact solutions of the Boltzmann equation in the VHP model with removal interaction. Journal of Statistical Physics, 66(3–4):1045–1058, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01055715.

Schaertl:1994:BDP

[SS94a] W. Schaertl and H. Sillescu. Brownian dynamics of polydisperse colloidal hard spheres: Equilibrium structures and random close packings. Journal of Statistical Physics, 77(5–6):1007–1025, December 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183148.

Schaertl:1994:BDS

[SS94b] W. Schaertl and H. Sillescu. Brownian dynamics simulations of colloidal hard spheres. Effects of sample dimensionality on self-diffusion. Journal of Statistical Physics, 74(3–4):687–703, February 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02188576.

Simanyi:1994:PBN

[SS94c] Nándor Simányi and Domokos Szász. The K-property of 4D billiards with nonorthogonal cylindric scatterers. Journal of Statistical Physics, 76(1–2):587–604, July 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02188676.

Sreenivasan:1995:TC

[SS95] K. R. Sreenivasan and G. Stolovitzky. Turbulent cascades. Journal of Statistical Physics, 78(1–2):311–333, January 1995.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02183351.

Salas:1996:DCB

[SS96a] Jesús Salas and Alan D. Sokal. Dynamic critical behavior of a Swendsen–Wang-type algorithm for the Ashkin–Teller model. Journal of Statistical Physics, 85(3–4):297–361, November 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02174209.

Schonmann:1996:CVP

[SS96b] Roberto H. Schonmann and Senya B. Shlosman. Constrained variational problem with applications to the Ising model. Journal of Statistical Physics, 83(5–6):867–905, June 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179548.

Salas:1997:APT

[SS97a] Jesús Salas and Alan D. Sokal. Absence of phase transition for antiferromagnetic Potts models via the Dobrushin uniqueness theorem. Journal of Statistical Physics, 86(3–4):551–579, February 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02199113.

Salas:1997:DCB

[SS97b] Jesús Salas and Alan D. Sokal. Dynamic critical behavior of the Swendsen–Wang algorithm: The two-dimensional three-state Potts model revisited. Journal of Statistical Physics, 87(1–2):1– 36, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02181478.

Salas:1997:LCF

[SS97c] Jesús Salas and Alan D. Sokal. Logarithmic corrections and finite-size scaling in the two-dimensional 4-state Potts model. Journal of Statistical Physics, 88(3–4):567–615, August 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000015164.98296.85.

Salas:1998:TSS

[SS98] Jesús Salas and Alan D. Sokal. The three-state square-lattice Potts antiferromagnet at zero temperature. Journal of Statistical Physics, 92(5–6):729–753, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023002322985.

Stephen:1999:PRD

[SS99] Michael J. Stephen and Robin B. Stinchcombe. Persistence in reaction–diffusion problems: I. Bosons. Journal of Statistical Physics, 95(1–2):171–180, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004577411435.

Selke:1997:SCB

[SSLI97] W. Selke, F. Szalma, P. Lajkó, and F. Iglói. Surface critical behavior of two-dimensional dilute Ising models. Journal of Statistical Physics, 89(5–6):1079–1085, December 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02764223.

Scarlatti:1995:LDI

[SSP95] Sergio Scarlatti, Maurizio Serva, and Michele Pasquini. Large deviations for Ising spin glasses with constrained disorder. Journal of Statistical Physics, 80(1–2):337–356, July 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02178362.

Stella:1993:BCB

[SSV93] Attilio L. Stella, Flavio Seno, and Carlo Vanderzande. Boundary critical behavior of d = 2 self-avoiding walks on correlated and uncorrelated vacancies. Journal of Statistical Physics, 73 (1–2):21–46, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01052749.

Shaw:1999:PTD

[SSZ99] L. B. Shaw, B. Schmittmann, and R. K. P. Zia. Phase transitions in a driven lattice gas with anisotropic interactions. Journal of Statistical Physics, 95(5–6):981–996, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004554517737.

Schonmann:1990:ODC

[ST90] Roberto H. Schonmann and Nelson I. Tanaka. One-dimensional caricature of phase transition. Journal of Statistical Physics, 61 (1–2):241–252, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01013963. See erratum [ST91].

Schonmann:1991:EOD

[ST91] Roberto H. Schonmann and Nelson I. Tanaka. Errata: Onedimensional caricature of phase transition. Journal of Statistical Physics, 64(1–2):477, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01057890. See [ST90].

Scacciatelli:1992:FFE

[ST92] E. Scacciatelli and B. Tirozzi. Fluctuation of the free energy in the Hopfield model. Journal of Statistical Physics, 67(5– 6):981–1008, June 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049007.

Shcherbina:1993:FEC

[ST93] M. Shcherbina and B. Tirozzi. The free energy of a class of Hopfield models. Journal of Statistical Physics, 72(1–2):113–125, July 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048042.

Santos:1995:AVM

[ST95] M. A. Santos and S. Teixeira. Anisotropic voter model. Journal of Statistical Physics, 78(3–4):963–970, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183696.

Salazar:1997:SAU

[ST97] R. Salazar and R. Toral. Simulated annealing using hybrid Monte Carlo. Journal of Statistical Physics, 89(5–6):1047–1060, December 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02764221.

Schulz:1999:ANS

[ST99] Michael Schulz and Steffen Trimper. Analytical and numerical studies of the one-dimensional spin facilitated kinetic Ising model. Journal of Statistical Physics, 94(1–2):173–201, January 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004563329723.

Stangroom:1991:BCF

[Sta91] James E. Stangroom. Basic considerations in flowing electrorheological fluids. Journal of Statistical Physics, 64(5–6):1059– 1072, September 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048814.

Standish:1992:SLB

[Sta92] Russell K. Standish. On the spectrum of the linear Boltzmann operator. Journal of Statistical Physics, 66(3–4):1003–1010, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01055713.

Stauffer:1994:IDF

[Sta94a] D. Stauffer. Ising droplets in five dimensions. Journal of Statistical Physics, 74(5–6):1323–1325, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02188236.

Stauffer:1994:EDS

[Sta94b] Dietrich Stauffer. Evolution by damage spreading in Kauffman model. Journal of Statistical Physics, 74(5–6):1293–1299, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02188232.

Stanley:1997:BLK

[Sta97a] H. Eugene Stanley. Turbulence: The legacy of A. N. Kolmogorov. Journal of Statistical Physics, 88(1–2):521–523, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02508484.

Stanley:1997:CAM

[Sta97b] H. Eugene Stanley. Coherent anomaly method. Journal of Statistical Physics, 86(1–2):441, January 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02180215.

Stanley:1997:LMS

[Sta97c] H. Eugene Stanley. The lure of modern science. Journal of Statistical Physics, 86(1–2):443–444, January 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02180216.

Stavans:1998:ASP

[Sta98] Joel Stavans. Axial segregation of powders in a horizontal rotating tube. Journal of Statistical Physics, 93(3–4):467–475, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000033159.91915.f7.

Samuelsen:1995:SOD

[STAJ95] Emil J. Samuelsen, Ellen D. Tuset, Dag Ausen, and Susan Jagner. Stacking order and disorder in layered K_{3}Me(CN)_{6} compounds studied by diffuse X-ray scattering: A realization of the ANNNI model? Journal of Statistical Physics, 78(1–2):135–145, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183342.

Stephenson:1990:HOC

[Ste90] John Stephenson. High-order cycles in the logistic map or centers of cardioids in the Mandelbrot set. Journal of Statistical Physics, 58(3–4):579–597, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01112764.

Stell:1991:SSR

[Ste91] George Stell. Sticky spheres and related systems. Journal of Statistical Physics, 63(5–6):1203–1221, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01030007.

Steif:1995:TAP

[Ste95a] Jeffrey E. Steif. Two applications of percolation to cellular automata. Journal of Statistical Physics, 78(5–6):1325–1335, March 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02180134.

Stell:1995:CPT

[Ste95b] George Stell. Criticality and phase transitions in ionic fluids. Journal of Statistical Physics, 78(1–2):197–238, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02183346.

Stein:1997:FES

[Ste97] Jürgen Stein. Flow equations and the strong-coupling expansion for the Hubbard model. Journal of Statistical Physics, 88 (1–2):487–511, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02508481.

Steinberger:1999:CTE

[Ste99a] Thomas Steinberger. Computing the topological entropy for piecewise monotonic maps on the interval. Journal of Statistical Physics, 95(1–2):287–303, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004585613252.

Stephen:1999:SRS

[Ste99b] Michael J. Stephen. Some results on Sinai diffusion. Journal of Statistical Physics, 96(1–2):403–414, July 1999. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004588819980.

Stoof:1997:MQT

[Sto97a] H. T. C. Stoof. Macroscopic quantum tunneling of a Bose condensate. Journal of Statistical Physics, 87(5–6):1353–1366, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02181289.

Stoop:1997:CHV

[Sto97b] R. Stoop. On convex hull violation by superpositions. Journal of Statistical Physics, 88(5–6):1393–1398, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02732441.

Streater:1994:CGF

[Str94] R. F. Streater. Convection in a gravitational field. Journal of Statistical Physics, 77(1–2):441–448, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02186851.

Strobel:1995:IMP

[Str95] Thomas Strobel. Interface motion in a planar spin-flip model derived from exclusion on the line. Journal of Statistical Physics, 79(5–6):923–950, June 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181209.

Streater:1997:GBP

[Str97] R. F. Streater. A gas of Brownian particles in statistical dynamics. Journal of Statistical Physics, 88(1–2):447–469, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02508479.

Siboni:1994:DTH

[STV94] S. Siboni, G. Turchetti, and S. Vaienti. Diffusion on the torus for Hamiltonian maps. Journal of Statistical Physics, 75(1– 2):167–187, April 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186285.

Suttorp:1992:ABC

[Sut92] L. G. Suttorp. Asymptotic behavior of correlation functions for electric potential and field fluctuations in a classical one- component plasma. Journal of Statistical Physics, 66(5–6):1343–1357, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01054425.

Suto:1996:LDE

[Süt96] András Sütö. Low-density expansion for unstable interactions and a model of crystallization. Journal of Statistical Physics, 82(5–6):1541–1573, March 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02183395.

Shen:1991:STC

[SW91] Hubert H. Shen and Alan A. Wray. Stationary turbulent closure via the Hopf functional equation. Journal of Statistical Physics, 65(1–2):33–52, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01329849.

Soffer:1998:NH

[SW98] A. Soffer and M. I. Weinstein. Nonautonomous Hamiltonians. Journal of Statistical Physics, 93(1–2):359–391, October 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1023/B%3AJOSS.0000026738.52652.6e.pdf.

Sander:1999:MCS

[SW99a] Evelyn Sander and Thomas Wanner. Monte Carlo simulations for spinodal decomposition. Journal of Statistical Physics, 95 (5–6):925–948, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004550416829.

Stuart:1999:AEC

[SW99b] A. M. Stuart and J. O. Warren. Analysis and experiments for a computational model of a heat bath. Journal of Statistical Physics, 97(3–4):687–723, November 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004667325896.

Sweet:1997:ACP

[Swe97] Ted Sweet. The asymmetric contact process at its second critical value. Journal of Statistical Physics, 86(3–4):749–764, February 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02199118.

Sodin:1990:LPF

[SY90] M. Sodin and P. Yuditski. The limit-periodic finite-difference operator on l^{2} (z) associated with iterations of quadratic polynomials. Journal of Statistical Physics, 60(5–6):863–873, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01026000.

Spohn:1995:BDL

[SY95] Herbert Spohn and Horng-Tzer Yau. Bulk diffusivity of lattice gases close to criticality. Journal of Statistical Physics, 79 (1–2):231–241, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179388.

Sauer:1991:E

[SYC91] Tim Sauer, James A. Yorke, and Martin Casdagli. Embedology. Journal of Statistical Physics, 65(3–4):579–616, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01053745.

Szabo:1991:RDI

[SZ91] Attila Szabo and Robert Zwanzig. Reversible diffusion-influenced reactions: comparison of theory and simulation for a simple model. Journal of Statistical Physics, 65(5–6):1057–1083, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049598.

Stroock:1995:EPG

[SZ95] D. Stroock and B. Zegarliński. On the ergodic properties of Glauber dynamics. Journal of Statistical Physics, 81(5–6):1007– 1019, December 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179301.

Song:1996:RDD

[SZ96] Renming Song and Xian Yin Zhou. A remark on diffusion of directed polymers in random environments. Journal of Statistical Physics, 85(1–2):277–289, October 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02175566.

Schmittmann:1998:TPC

[SZ98] B. Schmittmann and R. K. P. Zia. Two-point correlations and critical line of the driven Ising lattice gas in a high-temperature expansion. Journal of Statistical Physics, 91(3–4):525–539, May 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1023069325518.

Spohn:1999:DTP

[SZ99] Herbert Spohn and Wilhelm Zwerger. Decay of the two-point function in one-dimensional O(N) spin models with long-range interactions. Journal of Statistical Physics, 94(5–6):1037–1043, March 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004595419419.

Szatzschneider:1993:MTP

[Sza93] Wojciech Szatzschneider. The motion of a tagged particle and nonhomogeneous media in R^{1}. Journal of Statistical Physics, 70(5–6):1281–1296, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01049432.

Szamel:1997:SMD

[Sza97] Grzegorz Szamel. Statistical mechanics of dissipative transport in crystals. Journal of Statistical Physics, 87(5–6):1067–1082, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02181271.

Szeredi:1996:HCA

[Sze96] T. Szeredi. Hard chaos and adiabatic quantization: The wedge billiard. Journal of Statistical Physics, 83(1–2):259–274, April 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02183649.

Tabachnikov:1996:ADD

[Tab96] Serge Tabachnikov. Asymptotic dynamics of the dual billiard transformation. Journal of Statistical Physics, 83(1–2):27–37, April 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183638.

Taitelbaum:1995:ST

[Tai95] Haim Taitelbaum. Statistical thermophysics. Journal of Statistical Physics, 78(3–4):1183–1184, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183715.

Tang:1992:SSS

[Tan92] Lei-Han Tang. Steady-state scaling function of the (1 + 1)- dimensional single-step model. Journal of Statistical Physics, 67(3–4):819–826, May 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049729.

Tang:1994:TRL

[Tan94] Lei-Han Tang. Two repulsive lines on disordered lattices. Journal of Statistical Physics, 77(3–4):581–606, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179451.

Tang:1998:FDJ

[Tan98] Chao Tang. Fractal dimension of Julia set for nonanalytic maps. Journal of Statistical Physics, 93(3–4):1001–1008, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000033174.46243.d4.

Tasaki:1996:SFH

[Tas96] Hal Tasaki. Stability of ferromagnetism in Hubbard models with nearly flat bands. Journal of Statistical Physics, 84(3– 4):535–653, August 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179652.

Tateno:1998:CSB

[Tat98] Takashi Tateno. Characterization of stochastic bifurcations in a simple biological oscillator. Journal of Statistical Physics, 92 (3–4):675–705, August 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023048923644.

Taylor:1999:VAC

[Tay99] Jean E. Taylor. A variational approach to crystalline triplejunction motion. Journal of Statistical Physics, 95(5–6):1221– 1244, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1004523005442.

Twining:1992:ELC

[TB92] C. J. Twining and P.-M. Binder. Enumeration of limit cycles in noncylindrical cellular automata. Journal of Statistical Physics, 66(1–2):385–401, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01060073.

Tribel:1995:LGI

[TB95] Olivier Tribel and Jean Pierre Boon. Lattice gas with ‘interaction potential’. Journal of Statistical Physics, 81(1–2):361–377, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179984.

Tamayo:1990:SCMa

[TBK90a] P. Tamayo, R. C. Brower, and W. Klein. Single-cluster Monte Carlo dynamics for the Ising model. Journal of Statistical Physics, 58(5–6):1083–1094, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01026564.

Tamayo:1990:SCMb

[TBK90b] P. Tamayo, R. C. Brower, and W. Klein. Single cluster Monte Carlo dynamics for the Ising model. Journal of Statistical Physics, 60(5–6):889, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01026002.

Tamarit:1991:PCP

[TC91] Francisco A. Tamarit and Evaldo M. F. Curado. Pair-correlated patterns in Hopfield model of neural networks. Journal of Statistical Physics, 62(1–2):473–480, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020883.

Tsallis:1993:CMM

[TC93] Constantino Tsallis and Uriel M. S. Costa. Correlated majority model. Journal of Statistical Physics, 70(5–6):1383–1389, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049440.

Taylor:1994:LAS

[TC94] Jean E. Taylor and John W. Cahn. Linking anisotropic sharp and diffuse surface motion laws via gradient flows. Journal of Statistical Physics, 77(1–2):183–197, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02186838.

Tateno:1995:SPL

[TDSR95] Takashi Tateno, Shinji Doi, Shunsuke Sato, and Luigi M. Ricciardi. Stochastic phase lockings in a relaxation oscillator forced by a periodic input with additive noise: A first-passage-time approach. Journal of Statistical Physics, 78(3–4):917–935, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183694.

Tsonis:1995:TSN

[TE95] A. A. Tsonis and J. B. Elsner. Testing for scaling in natural forms and observables. Journal of Statistical Physics, 81(5–6): 869–880, December 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179296.

Telcs:1990:NRR

[Tel90] András Telcs. A note on recurrent random walks on graphs. Journal of Statistical Physics, 60(5–6):801–807, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01025994.

Taucher:1992:AVS

[TF92] T. Taucher and N. E. Frankel. Annealed n-vector p-spin model. Journal of Statistical Physics, 68(5–6):925–976, September 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01048881.

Taucher:1993:PDO

[TF93a] T. Taucher and N. E. Frankel. Probability distributions for the overlaps and self-correlations of the pure states of an nvector model. Journal of Statistical Physics, 71(3–4):361–378, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01058427.

Taucher:1993:QVS

[TF93b] T. Taucher and N. E. Frankel. Quenched n-vector p-spin model. Journal of Statistical Physics, 71(3–4):379–413, May 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01058428.

Tellez:1999:EFS

[TF99] G. Téllez and P. J. Forrester. Exact finite-size study of the 2D OCP at Γ = 4 and Γ = 6. Journal of Statistical Physics, 97(3– 4):489–521, November 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004654923170.

Tetot:1990:SPF

[TFD90] R. Tétot, A. Finel, and F. Ducastelle. Superdegenerate point in FCC phase diagram: CVM and Monte Carlo investigations. Journal of Statistical Physics, 61(1–2):121–141, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013956.

Tasaki:1995:FLF

[TG95] Shuichi Tasaki and Pierre Gaspard. Fick’s law and fractality of nonequilibrium stationary states in a reversible multibaker map. Journal of Statistical Physics, 81(5–6):935–987, December 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179299.

Trizac:1996:DGP

[TH96] Emmanuel Trizac and Jean-Pierre Hansen. Dynamics and growth of particles undergoing ballistic coalescence. Journal of Statistical Physics, 82(5–6):1345–1370, March 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183386.

Thaler:1995:IDM

[Tha95] Maximilian Thaler. The invariant densities for maps modeling intermittency. Journal of Statistical Physics, 79(3–4):739–741, May 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02184879.

Taitelbaum:1991:SPC

[THK^{+}91] Haim Taitelbaum, Shlomo Havlin, James E. Kiefer, Benes Trus, and George H. Weiss. Some properties of the a+b-> c reactiondiffusion system with initially separated components. Journal of Statistical Physics, 65(5–6):873–891, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01049587.

Timoshenko:1997:NAR

[TKD97] E. G. Timoshenko, Yu. A. Kuznetsov, and K. A. Dawson. A nonequilibrium approach for random amphiphilic copolymer model. Journal of Statistical Physics, 89(1–2):347–367, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02770769.

Teich:1993:SCS

[TKG93] Malvin C. Teich, Shyam M. Khanna, and Patrick C. Guiney. Spectral characteristics and synchrony in primary auditory-nerve fibers in response to pure-tone acoustic stimuli. Journal of Statistical Physics, 70(1–2):257–279, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053967.

Leung:1990:TMI

[tL90] Kwan tai Leung. Theory on morphological instability in driven systems. Journal of Statistical Physics, 61(1–2):345–364, October 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01013969.

Ting:1991:EAL

[TLW91] Julian J.-L. Ting, Simon C. Lin, and F. Y. Wu. Exact analysis of a lattice gas on the 3–12 lattice with two- and three-site interactions. Journal of Statistical Physics, 62(1–2):35–43, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020858.

Leung:1996:SDA

[tLZ96] K. t. Leung and R. K. P. Zia. Subtleties in data analysis related to the size of critical region. Journal of Statistical Physics, 83 (5–6):1219–1223, June 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179560.

Theunissen:1999:RTQ

[TNN99] M. Theunissen, C. Nicolis, and G. Nicolis. Recurrence times in quasi-periodic motion: Statistical properties, role of cell size, parameter dependence. Journal of Statistical Physics, 94(3–4): 437–467, February 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004544220154.

Toom:1994:CPIa

[Too94a] Andrei Toom. On critical phenomena in interacting growth systems. Part I: General. Journal of Statistical Physics, 74(1– 2):91–109, January 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186808.

Toom:1994:CPIb

[Too94b] Andrei Toom. On critical phenomena in interacting growth systems.Part II: Bounded growth. Journal of Statistical Physics, 74 (1–2):111–130, January 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186809.

Toom:1995:SOD

[Too95] Andrei Toom. Simple one-dimensional interaction systems with superexponential relaxation times. Journal of Statistical Physics, 80(3–4):545–563, August 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02178547.

Toom:1997:TH

[Too97] Andre Toom. Tails in harnesses. Journal of Statistical Physics, 88(1–2):347–364, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02508475.

Toom:1999:PFL

[Too99] André Toom. On percolation with fibers or layers. Journal of Statistical Physics, 96(1–2):429–437, July 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004593020888.

Torquato:1991:DRA

[Tor91] S. Torquato. Diffusion and reaction among traps: some theoretical and simulation results. Journal of Statistical Physics, 65(5–6): 1173–1206, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049606.

Toth:1990:PTI

[Tót90] Bálint Tóth. Phase transition in an interacting Bose system. An application of the theory of Ventsel’ and Freidlin. Journal of Statistical Physics, 61(3–4):749–764, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01027300.

Toth:1991:UBC

[Tót91] Bálint Tóth. Upper bound on the condensate in the hard-core Bose lattice gas. Journal of Statistical Physics, 65(1–2):373–378, October 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01329865.

Toth:1994:TSA

[Tót94] Bálint Tóth. ‘True’ self-avoiding walks with generalized bond repulsion on Z. Journal of Statistical Physics, 77(1–2):17–33, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186830.

Tucker:1992:MVT

[TP92] Susan C. Tucker and Eli Pollak. Microcanonical variational transition-state theory for reaction rates in dissipative systems. Journal of Statistical Physics, 66(3–4):975–990, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01055711.

Tan:1995:LBA

[TQGO95] M.-L. Tan, Y. H. Qian, I. Goldhirsch, and S. A. Orszag. Lattice– BGK approach to simulating granular flows. Journal of Statistical Physics, 81(1–2):87–103, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179970.

Tij:1994:PAS

[TS94] Mohamed Tij and Andrés Santos. Perturbation analysis of a stationary nonequilibrium flow generated by an external force. Journal of Statistical Physics, 76(5–6):1399–1414, September 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02187068.

Tarjus:1991:RSA

[TST91] Gilles Tarjus, Pierre Schaaf, and Julian Talbot. Random sequential addition: A distribution function approach. Journal of Statistical Physics, 63(1–2):167–202, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01026598.

Temperley:1994:MLV

[TT94] H. N. V. Temperley and D. H. Trevena. Metastability of the liquid-vapor transition and related effects. Journal of Statistical Physics, 77(1–2):501–508, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02186855.

Takayasu:1991:SPA

[TTPH91] Hideki Takayasu, Misako Takayasu, Astero Provata, and Greg Huber. Statistical properties of aggregation with injection. Journal of Statistical Physics, 65(3–4):725–745, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01053751.

Tangerman:1990:AGH

[TV90] F. M. Tangerman and J. J. P. Veerman. Asymptotic geometry of hyperbolic well-ordered Cantor sets. Journal of Statistical Physics, 59(1–2):299–321, April 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01015571.

Toscani:1999:PMU

[TV99] G. Toscani and C. Villani. Probability metrics and uniqueness of the solution to the Boltzmann equation for a Maxwell gas. Journal of Statistical Physics, 94(3–4):619–637, February 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004508706950.

Tesi:1996:MCS

[TvROW96] M. C. Tesi, E. J. Janse van Rensburg, E. Orlandini, and S. G. Whittington. Monte Carlo study of the interacting self-avoiding walk model in three dimensions. Journal of Statistical Physics, 82 (1–2):155–181, January 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02189229.

Tracy:1998:CFC

[TW98] Craig A. Tracy and Harold Widom. Correlation functions, cluster functions, and spacing distributions for random matrices. Journal of Statistical Physics, 92(5–6):809–835, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023084324803.

Tomiya:1996:SNR

[TY96] M. Tomiya and N. Yoshinaga. Scars in nonintegrable and rational billiards. Journal of Statistical Physics, 83(1–2):215–242, April 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183647.

Toroczkai:1997:POD

[TZ97] Z. Toroczkai and R. K. P. Zia. Periodic one-dimensional hopping model with one mobile directional impurity. Journal of Statistical Physics, 87(3–4):545–575, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02181236.

Ueltschi:1999:AHM

[Uel99] Daniel Ueltschi. Analyticity in Hubbard models. Journal of Statistical Physics, 95(3–4):693–717, May 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004599410952.

Ueno:1995:DOP

[Uen95] Yohtaro Ueno. Description of ordering and phase transitions in terms of local connectivity: Proof of a novel type of percolated state in the general clock model. Journal of Statistical Physics, 80 (3–4):841–873, August 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02178558.

Uglov:1996:SOA

[UI96] D. B. Uglov and I. T. Ivanov. sl(n) Onsager’s algebra and integrability. Journal of Statistical Physics, 82(1–2):87–113, January 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02189226.

Ueno:1991:IAD

[UO91] Y. Ueno and Y. Ozeki. Interfacial approach to d-dimensional ±J Ising models in the neighborhood of the ferromagnetic phase boundary. Journal of Statistical Physics, 64(1–2):227–249, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01057875.

Uchida:1991:KMG

[UW91] Takashi Uchida and Koh Wada. Kinetics of monolayer growth. Journal of Statistical Physics, 64(3–4):605–630, August 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01048308.

Vaienti:1992:EPD

[Vai92] S. Vaienti. Ergodic properties of the discontinuous sawtooth map. Journal of Statistical Physics, 67(1–2):251–269, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01049033.

vanBeijeren:1990:LRS

[vB90a] H. van Beijeren. Long-range spatial correlations in a simple diffusion model. Journal of Statistical Physics, 60(5–6):845– 849, September 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01025998.

Vericat:1990:MBF

[VB90b] Fernando Vericat and Lesser Blum. Many-body functions of nonprimitive electrolytes in one dimension. Journal of Statistical Physics, 61(5–6):1161–1185, December 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01014370.

vanBeijeren:1991:FMM

[vB91] H. van Beijeren. Fluctuations in the motions of mass and of patterns in one-dimensional driven diffusive systems. Journal of Statistical Physics, 63(1–2):47–58, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01026591.

Vohra:1993:OSR

[VB93] S. T. Vohra and F. Bucholtz. Observation of stochastic resonance near a subcritical bifurcation. Journal of Statistical Physics, 70 (1–2):413–421, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053977.

Vanderheyden:1998:SCA

[VB98] Benoît Vanderheyden and Gordon Baym. Self-consistent approximations in relativistic plasmas: Quasiparticle analysis of the thermodynamic properties. Journal of Statistical Physics, 93(3– 4):843–861, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000033166.37520.ae.

vanBeijeren:1993:DLL

[vBE93] H. van Beijeren and M. H. Ernst. Diffusion in Lorentz lattice gas automata with backscattering. Journal of Statistical Physics, 70(3–4):793–809, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053595.

Velazquez:1997:VEM

[VBF97] Esov S. Velázquez, Lesser Blum, and Harry L. Frisch. Variational extension of the mean spherical approximation to arbitrary dimensions. Journal of Statistical Physics, 89(1–2):203–214, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02770761.

vanCoevorden:1994:RTF

[vCEBS94] D. V. van Coevorden, M. H. Ernst, R. Brito, and J. A. Somers. Relaxation and transport in FCHC lattice gases. Journal of Statistical Physics, 74(5–6):1085–1115, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02188218.

vanDongen:1990:SFR

[vD90] P. G. J. van Dongen. Spatial fluctuations in reaction-diffusion systems: A model for exponential growth. Journal of Statistical Physics, 58(1–2):87–114, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020286.

vanDongen:1997:ABG

[vD97] P. G. J. van Dongen. Aggregation beyond the gel point: A new class of exactly solvable models. Journal of Statistical Physics, 87 (5–6):1273–1286, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181284.

vandenBerg:1992:BGC

[vdBDP92] M. van den Berg, T. C. Dorlas, and V. B. Priezzhev. The boson gas on a Cayley tree. Journal of Statistical Physics, 69(1– 2):307–328, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053795.

vandenBrule:1991:MCS

[vdBJ91] B. H. A. A. van den Brule and R. J. J. Jongschaap. Modeling of concentrated suspensions. Journal of Statistical Physics, 62(5– 6):1225–1237, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01128184.

vandenBroeck:1993:BRF

[vdBV93] C. van den Broeck and C. Vanderzande. Book review: Fundamental problems in statistical mechanics VII. Journal of Statistical Physics, 70(5–6):1403–1404, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01049443.

Vakarin:1997:CPF

[VDH97] Eduard Vakarin, Yurko Duda, and Myroslav Holovko. Continuum percolation of the four-bonding-site associating fluids. Journal of Statistical Physics, 88(5–6):1333–1352, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02732436.

vanderHofstad:1998:CCL

[vdH98] Remco van der Hofstad. The constants in the Central Limit Theorem for the one-dimensional Edwards model. Journal of Statistical Physics, 90(5–6):1295–1310, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023235529311.

vanDuijneveldt:1995:PBH

[vDL95] Jeroen S. van Duijneveldt and Henk N. W. Lekkerkerker. Phase behavior of hard particles. Journal of Statistical Physics, 78(1– 2):103–116, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183340.

vanderSman:1999:DLB

[vdSE99] R. G. M. van der Sman and M. H. Ernst. Diffusion lattice Boltzmann scheme on a orthorhombic lattice. Journal of Statistical Physics, 94(1–2):203–217, January 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004515413793.

Verberg:1997:SRS

[VdSFC97] R. Verberg, I. M. de Schepper, M. J. Feigenbaum, and E. G. D. Cohen. Square root singularity in the viscosity of neutral colloidal suspensions at large frequencies. Journal of Statistical Physics, 87(5–6):1037–1049, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02181269.

vanEnter:1990:SEN

[vE90] A. C. D. van Enter. Stiffness exponent, number of pure states, and Almeida–Thouless line in spin-glasses. Journal of Statistical Physics, 60(1–2):275–279, July 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01013678.

vanEnter:1996:IDB

[vE96] Aernout C. D. van Enter. Ill-defined block-spin transformations at arbitrarily high temperatures. Journal of Statistical Physics, 83(3–4):761–765, May 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183748.

vanEnter:1990:FSE

[vEAD90] A. C. D. van Enter, Joan Adler, and J. A. M. S. Duarte. Finitesize effects for some bootstrap percolation models. Journal of Statistical Physics, 60(3–4):323–332, August 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01314923.

vanEnter:1991:FSE

[vEAD91] A. C. D. van Enter, Joan Adler, and J. A. M. S. Duarte. Finite size effects for some bootstrap percolation models. Journal of Statistical Physics, 62(1–2):505–506, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020891.

vanEnter:1991:BRG

[vEdH91] Aernout van Enter and Frank den Hollander. Book review: Gibbs measures and phase transitions. Journal of Statistical Physics, 64(1–2):471–473, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01057888.

vanEnter:1995:PBR

[vEFK95] Aernout C. D. van Enter, Roberto Fernández, and Roman Kotecký. Pathological behavior of renormalization-group maps at high fields and above the transition temperature. Journal of Statistical Physics, 79(5–6):969–992, June 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02181211.

vanEnter:1993:RPP

[vEFS93] Aernout C. D. van Enter, Roberto Fernández, and Alan D. Sokal. Regularity properties and pathologies of position-space renormalization-group transformations: Scope and limitations of Gibbsian theory. Journal of Statistical Physics, 72(5–6):879– 1167, September 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048183.

Velazquez:1998:BDE

[Vel98] J. J. L. Velázquez. The Becker-Döring equations and the Lifshitz–Slyozov theory of coarsening. Journal of Statistical Physics, 92(1–2):195–236, July 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023099720145.

vanEnter:1992:HSO

[vEM92] A. C. D. van Enter and Jacek Miekisz. How should one define a (weak) crystal? Journal of Statistical Physics, 66(3–4): 1147–1153, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01055722.

vanEnter:1998:NLR

[vEMZ98] Aernout C. D. van Enter, Jacek Miekisz, and Milos Zahradník. Nonperiodic long-range order for fast-decaying interactions at positive temperatures. Journal of Statistical Physics, 90(5–6): 1441–1447, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023299916107.

vanEnter:1998:AGD

[vES98] Aernout C. D. van Enter and Senya B. Shlosman. (almost) Gibbsian description of the sign fields of SOS fields. Journal of Statistical Physics, 92(3–4):353–368, August 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023024218192.

Velasco:1992:KFE

[VGC92] R. M. Velasco and L. S. García-Colín. The kinetic foundation of extended irreversible thermodynamics revisited. Journal of Statistical Physics, 69(1–2):217–229, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01053791.

Vugmeister:1998:FTD

[VHR98] B. E. Vugmeister, D. L. Huber, and H. Rabitz. Freezing temperature in dilute Ising spin glasses with long-range interactions. Journal of Statistical Physics, 90(3–4):873–888, February 1998.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023237306231.

vanHemmen:1993:LFK

[vHW93] J. L. van Hemmen and W. F. Wreszinski. Lyapunov function for the Kuramoto model of nonlinearly coupled oscillators. Journal of Statistical Physics, 72(1–2):145–166, July 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01048044.

Vieillefosse:1994:CPD

[Vie94] P. Vieillefosse. Coulomb pair density matrix I. Journal of Statistical Physics, 74(5–6):1195–1209, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02188223.

Vieillefosse:1995:CPD

[Vie95] P. Vieillefosse. Coulomb pair density matrix. II. Journal of Statistical Physics, 80(1–2):461–479, July 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02178368.

vanKampen:1991:ORT

[vK91] N. G. van Kampen. Onsager relations for transport in inhomogeneous media. Journal of Statistical Physics, 63(5–6):1019–1033, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01029996.

vanKampen:1993:SFP

[vK93] N. G. van Kampen. Short first-passage times. Journal of Statistical Physics, 70(1–2):15–23, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053951.

vanKampen:1994:NTS

[vK94] N. G. van Kampen. Note on the two-slit experiment. Journal of Statistical Physics, 77(1–2):345–350, October 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02186845.

vanKampen:1995:SMQ

[vK95a] N. G. van Kampen. A soluble model for quantum mechanical dissipation. Journal of Statistical Physics, 78(1–2):299–310, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02183350.

vanKampen:1995:TMB

[vK95b] N. G. van Kampen. The turning of magnetotactic bacteria. Journal of Statistical Physics, 80(1–2):23–33, July 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02178351.

vanKampen:1997:LME

[vKO97] N. G. van Kampen and I. Oppenheim. Langevin and master equation in quantum mechanics. Journal of Statistical Physics, 87(5–6):1325–1334, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02181287.

vanLeeuwen:1997:EIP

[vLH97] J. M. J. van Leeuwen and E. H. Hauge. The effective interface potential for a superconducting layer. Journal of Statistical Physics, 87(5–6):1335–1351, June 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02181288.

Vollmayr:1994:CHA

[Vol94] H. Vollmayr. Cluster hull algorithms for large systems with small memory requirement. Journal of Statistical Physics, 74(3–4): 919–927, February 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02188585.

Voorhees:1992:DFP

[Voo92] Burton Voorhees. Determination of fixed points and shift cycles for nearest neighbor cellular automata. Journal of Statistical Physics, 66(5–6):1397–1414, March 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01054428.

vanRensburg:1997:CSA

[vR97a] E. J. Janse van Rensburg. Crumpling self-avoiding surfaces. Journal of Statistical Physics, 88(1–2):177–200, July 1997. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02508469.

Vugmeister:1997:ELF

[VR97b] B. E. Vugmeister and H. Rabitz. Effect of local field fluctuations on orientational ordering in random-site dipole systems. Journal of Statistical Physics, 88(1–2):471–486, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02508480.

vanRoij:1993:CDL

[vRE93] René H. van Roij and Matthieu H. Ernst. Collective diffusion in a lattice gas automaton. Journal of Statistical Physics, 73 (1–2):47–68, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01052750.

vanRensburg:1996:ECL

[vROS^{+}96] E. J. Janse van Rensburg, E. Orlandini, D. W. Sumners, M. C. Tesi, and S. G. Whittington. Entanglement complexity of lattice ribbons. Journal of Statistical Physics, 85(1–2):103–130, October 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02175557.

Vujic:1999:BPT

[Vuj99] Dragan Vujić. Branched polymers on the two-dimensional square lattice with attractive surfaces. Journal of Statistical Physics, 95 (3–4):767–774, May 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004507812769.

vanVelzen:1993:SSC

[vVBE93] G. A. van Velzen, R. Brito, and M. H. Ernst. Stress-stress correlation functions in lattice gases beyond Boltzmann’s approximation. Journal of Statistical Physics, 70(3–4):811–832, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01053596.

Vogeler:1993:PIL

[VWG93] Armin Vogeler and Dieter A. Wolf-Gladrow. Pair interaction lattice gas simulations: Flow past obstacles in two and three dimensions. Journal of Statistical Physics, 71(1–2):163–190, April 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01048093.

vanWijland:1997:UFS

[vWH97] F. van Wijland and H. J. Hilhorst. Universal fluctuations in the support of the random walk. Journal of Statistical Physics, 89(1– 2):119–134, October 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02770757.

vanWonderen:1995:QTD

[vWL95] A. J. van Wonderen and K. Lendi. Quantum theory of dissipative processes: The Markov approximation revisited. Journal of Statistical Physics, 80(1–2):273–305, July 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02178360.

Verbeure:1992:PTA

[VZ92] A. Verbeure and V. A. Zagrebnov. Phase transitions and algebra of fluctuation operators in an exactly soluble model of a quantum anharmonic crystal. Journal of Statistical Physics, 69(1– 2):329–359, October 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053796.

Verbeure:1994:GNG

[VZ94] A. Verbeure and V. A. Zagrebnov. Gaussian, non-Gaussian critical fluctuations in the Curie–Weiss model. Journal of Statistical Physics, 75(5–6):1137–1152, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02186760.

Verbeure:1995:DQF

[VZ95] A. Verbeure and V. A. Zagrebnov. Dynamics of quantum fluctuations in an anharmonic crystal model. Journal of Statistical Physics, 79(1–2):377–393, April 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179394.

Vespignani:1997:DDR

[VZL97] Alessandro Vespignani, Stefano Zapperi, and Vittorio Loreto. Dynamically driven renormalization group. Journal of Statistical Physics, 88(1–2):47–79, July 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02508464.

Wehr:1990:FEF

[WA90] Jan Wehr and Michael Aizenman. Fluctuations of extensive functions of quenched random couplings. Journal of Statistical Physics, 60(3–4):287–306, August 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01314921.

Wagner:1992:CPB

[Wag92] Wolfgang Wagner. A convergence proof for Bird’s direct simulation Monte Carlo method for the Boltzmann equation. Journal of Statistical Physics, 66(3–4):1011–1044, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01055714.

Wagner:1995:ABE

[Wag95] Wolfgang Wagner. Approximation of the Boltzmann equation by discrete velocity models. Journal of Statistical Physics, 78(5– 6):1555–1570, March 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02180142.

Wagner:1998:TFM

[Wag98] Christoph Wagner. Traffic flow models considering an internal degree of freedom. Journal of Statistical Physics, 90(5–6): 1251–1275, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023231428403.

Walgraef:1991:TFW

[Wal91] D. Walgraef. Temporal forcing of wave patterns. Journal of Statistical Physics, 64(5–6):969–980, September 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01048808.

Wang:1996:ELS

[Wan96a] Feng-Yu Wang. Estimates of logarithmic Sobolev constant for finite-volume continuous spin systems. Journal of Statistical Physics, 84(1–2):277–293, July 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179587.

Wang:1996:AFS

[Wan96b] Jian-Sheng Wang. Anisotropic finite-size scaling analysis of a two-dimensional driven diffusive system. Journal of Statistical Physics, 82(5–6):1409–1427, March 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183389.

Warnaar:1996:FSAa

[War96a] S. Ole Warnaar. Fermionic solution of the Andrews–Baxter– Forrester model. I. Unification of TBA and CTM methods. Journal of Statistical Physics, 82(3–4):657–685, February 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179790.

Warnaar:1996:FSAb

[War96b] S. Ole Warnaar. Fermionic solution of the Andrews–Baxter– Forrester model. II. Proof of Melzer’s polynomial identities. Journal of Statistical Physics, 84(1–2):49–83, July 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179577.

Watson:1999:SRS

[Wat99] Greg I. Watson. Symmetry relations for the six-vertex model. Journal of Statistical Physics, 94(5–6):1045–1054, March 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004547503489.

Wood:1990:PO

[WB90] D. W. Wood and J. K. Ball. On a point of order. Journal of Statistical Physics, 58(3–4):599–615, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01112765.

Wang:1992:MCC

[WB92] Xidi Wang and George A. Baker, Jr. Monte Carlo calculations of the conformal charge. Journal of Statistical Physics, 69(5–6): 1069–1095, December 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01058762.

Weinstein:1994:VKEa

[WB94a] Edward M. Weinstein and H. Benaroya. The van Kampen expansion for the Fokker–Planck equation of a Duffing oscillator. Journal of Statistical Physics, 77(3–4):667–679, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179455.

Weinstein:1994:VKEb

[WB94b] Edward M. Weinstein and H. Benaroya. The van Kampen expansion for the Fokker–Planck equation of a Duffing oscillator excited by colored noise. Journal of Statistical Physics, 77(3–4): 681–690, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179456.

Wright:1999:EME

[WB99] J. D. Wright and David Brydges. Erratum: Mayer Expansions and the Hamiltonian–Jacobi Equation. II. Fermions, Dimensional Reduction Formulas. Journal of Statistical Physics, 97 (5–6):1027, December 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004666114935. See [BW88].

Wagner:1998:IQC

[WBG98] Holger Wagner, Ellen Baake, and Thomas Gerisch. Ising quantum chain and sequence evolution. Journal of Statistical Physics, 92(5–6):1017–1052, September 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023048711599.

Wang:1995:DLL

[WC95] F. Wang and E. G. D. Cohen. Diffusion in Lorentz lattice gas cellular automata: The honeycomb and quasi-lattices compared with the square and triangular lattices. Journal of Statistical Physics, 81(1–2):467–495, October 1995. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179989.

Wang:1996:DRL

[WC96] F. Wang and E. G. D. Cohen. Diffusion on random lattices. Journal of Statistical Physics, 84(1–2):233–261, July 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02179584.

Wreszinski:1998:MTL

[WC98] Walter F. Wreszinski and S. Casmeridis. Models of two-level atoms in quasiperiodic external fields. Journal of Statistical Physics, 90(3–4):1061–1068, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023257810774.

Willaime:1991:ROL

[WCT91] H. Willaime, O. Cardoso, and P. Tabeling. Regimes of oscillation in a linear array of vortices. Journal of Statistical Physics, 64(5– 6):961–968, September 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048807.

Weeks:1991:CCW

[Wee91] John D. Weeks. Comment on the capillary wave model in three dimensions. Journal of Statistical Physics, 64(3–4):823–827, August 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01048317.

Wehr:1997:LBV

[Weh97a] Jan Wehr. A lower bound on the variance of conductance in random resistor networks. Journal of Statistical Physics, 86(5– 6):1359–1365, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183627.

Wehr:1997:NIG

[Weh97b] Jan Wehr. On the number of infinite geodesics and ground states in disordered systems. Journal of Statistical Physics, 87 (1–2):439–447, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181495.

Wehr:1997:SLL

[Weh97c] Jan Wehr. A strong law of large numbers for iterated functions of independent random variables. Journal of Statistical Physics, 86(5–6):1373–1384, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02183629. See erra-

tum [Ano01].

Weiss:1991:BRF

[Wei91a] George H. Weiss. Book review: Finite-size scaling, vol. 2. Journal of Statistical Physics, 62(1–2):501, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020889.

Weiss:1991:BRK

[Wei91b] George H. Weiss. Book review: Kinetics of diffusion controlled chemical processes. Journal of Statistical Physics, 65(3– 4):823–824, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053758.

Weiss:1992:SVF

[Wei92] Howard Weiss. Some variational formulas for Hausdorff dimension, topological entropy, and SRB entropy for hyperbolic dynamical systems. Journal of Statistical Physics, 69(3–4): 879–886, November 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01050439.

Weiss:1993:BRM

[Wei93a] George H. Weiss. Book review: Maximum entropy in action. Journal of Statistical Physics, 70(3–4):1081–1082, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01053611.

Weiss:1993:BRS

[Wei93b] George H. Weiss. Book review: Stochastic processes in physics and chemistry. Journal of Statistical Physics, 73(1–2):447, October 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01052774.

Weiss:1995:RPP

[Wei95] Howard Weiss. A remark on papers by Pixton and Oliveira: Genericity of symplectic diffeomorphisms of S^{2} with positive topological entropy. Journal of Statistical Physics, 80(1–2):481– 485, July 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02178369.

Weiss:1996:RWR

[Wei96] George H. Weiss. Random walks and random environments, volume 1: Random walks. Journal of Statistical Physics, 82(5–6): 1675–1677, March 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183400.

Weiss:1999:LSC

[Wei99] Howard Weiss. The Lyapunov spectrum for conformal expanding maps and axiom — a surface diffeomorphisms. Journal of Statistical Physics, 95(3–4):615–632, May 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004591209134.

Wennberg:1997:EDM

[Wen97] Bernt Wennberg. Entropy dissipation and moment production for the Boltzmann equation. Journal of Statistical Physics, 86(5– 6):1053–1066, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183613.

Wennberg:1999:ENS

[Wen99] Bernt Wennberg. An example of nonuniqueness for solutions to the homogeneous Boltzmann equation. Journal of Statistical Physics, 95(1–2):469–477, April 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004546031908.

West:1991:BRS

[Wes91] Bruce J. West. Book review: Synergetics and dynamical instabilities. Journal of Statistical Physics, 62(1–2):493–495, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020886.

Weiss:1990:D

[WF90] George H. Weiss and Michael E. Fisher. Dedication. Journal of Statistical Physics, 58(3–4):407–409, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01112752.

Williams:1991:DSD

[WF91] G. O. Williams and H. L. Frisch. Diffusion from sessile droplets through membranes. Journal of Statistical Physics, 63(5–6): 1035–1037, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01029997.

Weiss:1993:MPP

[WG93] George H. Weiss and Moshe Gitterman. Motion in a periodic potential driven by rectangular pulses. Journal of Statistical Physics, 70(1–2):93–105, January 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01053956.

Wolf-Gladrow:1995:LBE

[WG95] Dieter Wolf-Gladrow. A lattice Boltzmann equation for diffusion. Journal of Statistical Physics, 79(5–6):1023–1032, June 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02181215.

Wang:1996:CHM

[WG96] D. F. Wang and C. Gruber. On the chiral Hubbard model and the chiral Kondo lattice model. Journal of Statistical Physics, 82(1– 2):421–430, January 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02189237.

Weiss:1991:SPF

[WH91] George H. Weiss and Shlomo Havlin. Some properties of a fractaltime continuous-time random walk in the presence of traps. Journal of Statistical Physics, 63(5–6):1005–1018, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01029995.

Wagner:1995:LBS

[WH95] Lukas Wagner and Fernand Hayot. Lattice Boltzmann simulations of flow past a cylindrical obstacle. Journal of Statistical Physics, 81(1–2):63–70, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179968.

Wheeler:1999:CHV

[Whe99] A. A. Wheeler. Cahn–Hoffman ξ-vector and its relation to diffuse interface models of phase transitions. Journal of Statistical Physics, 95(5–6):1245–1280, June 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004575022280.

Weyersberg:1992:CBS

[WHF92] A. Weyersberg, T. Holey, and M. Fähnle. Critical behavior of the specific heat for pure and site-diluted simple cubic Ising systems. Journal of Statistical Physics, 66(1–2):133–138, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01060063.

Whittle:1994:PMG

[Whi94] P. Whittle. Polymer models and generalized Potts–Kasteleyn models. Journal of Statistical Physics, 75(5–6):1063–1092, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF02186757.

Wiener:1991:ECF

[WHS^{+}91] Richard J. Wiener, Philip W. Hammer, Charles E. Swanson, David C. Samuels, and Russell J. Donnelly. The effect of a Coriolis force on Taylor–Couette flow. Journal of Statistical Physics, 64(5–6):913–926, September 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01048804.

Widom:1998:BRB

[Wid98] Benjamin Widom. Book review: The Critical Point. A Historical Introduction to the Modern Theory of Critical Phenomena. Cyril Domb, Taylor and Francis, London, 1996. Journal of Statistical Physics, 92(5–6):1213–1216, September 1998. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023069316142.

Widom:1999:RBO

[Wid99] Harold Widom. On the relation between orthogonal, symplectic and unitary matrix ensembles. Journal of Statistical Physics, 94(3–4):347–363, February 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004536018336.

Wieczerkowski:1997:RCE

[Wie97] C. Wieczerkowski. Running coupling expansion for the renormalized Φ_{4}^{4} trajectory from renormalization invariance. Journal of Statistical Physics, 89(5–6):929–945, December 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02764214.

Wieczerkowski:1998:CHT

[Wie98a] C. Wieczerkowski. Construction of the hierarchical ϕ^{4}-trajectory. Journal of Statistical Physics, 92(3–4):377–430, August 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023028319101.

Wiese:1998:PEK

[Wie98b] Kay Jörg Wiese. On the perturbation expansion of the KPZ equation. Journal of Statistical Physics, 93(1–2):143–154, October 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1023/B%3AJOSS.0000026730.76868.c4.pdf.

Wilemski:1991:NBD

[Wil91] Gerald Wilemski. Nonequilibrium Brownian dynamics simulations of shear thinning in concentrated colloidal suspensions. Journal of Statistical Physics, 62(5–6):1239–1253, March 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01128185.

Weiss:1990:SRM

[WK90] Jeffrey B. Weiss and Edgar Knobloch. A stochastic return map for stochastic differential equations. Journal of Statistical Physics, 58(5–6):863–883, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01026555.

Weron:1997:EPC

[WK97] Karina Weron and Marcin Kotulski. On the equivalence of the parallel channel and the correlated cluster relaxation models. Journal of Statistical Physics, 88(5–6):1241–1256, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02732433.

Weingartner:1995:CNC

[WKWS95] Hermann Weingärtner, M. Kleemeier, S. Wiegand, and W. Schröer. Coulombic and non-Coulombic contributions to the criticality of ionic fluids. An experimental approach. Journal of Statistical Physics, 78(1–2):169–196, January 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02183345.

Wilson:1992:MBF

[WL92] W. G. Wilson and W. G. Laidlaw. Microscopic-based fluid flow invasion simulations. Journal of Statistical Physics, 66(3–4): 1165–1176, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01055724.

Wu:1995:EPV

[WL95] Jiang Wu and Xiufang Liu. The existence of phase V in the Mandelbrot percolation process. Journal of Statistical Physics, 80(5–6):1405–1414, September 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179876.

Wilson:1994:MBF

[WLC94] W. G. Wilson, W. G. Laidlaw, and D. A. Coombe. Microscopicbased fluid flow simulation of invasion on a two-dimensional lattice. II. Mobilization and cohesion. Journal of Statistical Physics, 75(5–6):1185–1195, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02186764.

Widom:1998:VFP

[WM98] Michael Widom and José A. Miranda. Viscous fingering patterns in ferrofluids. Journal of Statistical Physics, 93(3–4): 411–426, November 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000033156.44251.15.

Weiss:1990:ARW

[WMS90] George H. Weiss, Jaume Masoliver, and Kurt E. Shuler. On the asymmetry of a random walk in the presence of a field. Journal of Statistical Physics, 58(3–4):643–652, February 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01112768.

Wright:1990:IPA

[WMTR90] H. Wright, R. Muralidhar, T. Tobin, and D. Ramkrishna. Inverse problems of aggregation processes. Journal of Statistical Physics, 61(3–4):843–863, November 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01027303.

Wolansky:1992:RTD

[Wol92] Gershon Wolansky. Resonance trapping in dissipative and antidissipative systems: An ergodic approach. Journal of Statistical Physics, 67(1–2):33–64, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01049026.

Wu:1995:CPM

[WPK95] F. Y. Wu, P. Pant, and C. King. The chiral Potts model and its associated link invariant. Journal of Statistical Physics, 78(5– 6):1253–1276, March 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02180131.

Warnaar:1994:OPD

[WPSN94] S. Ole Warnaar, Paul A. Pearce, Katherine A. Seaton, and Bernard Nienhuis. Order parameters of the dilute a models. Journal of Statistical Physics, 74(3–4):469–531, February 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02188569.

Wieczerkowski:1997:IPE

[WR97] C. Wieczerkowski and J. Rolf. Interpolation parameter and expansion for a three-dimensional nontrivial scalar infrared fixed point. Journal of Statistical Physics, 89(3–4):817–845, November 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02765546.

Walsh:1995:ABO

[WRJ95] C. Walsh, T. S. Ray, and Naeem Jan. Anomalous biennial oscillations in a Fisher equation with a discretized Verhulst term. Journal of Statistical Physics, 81(3–4):761–775, November 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179256.

Widder:1992:ITG

[WT92] M. E. Widder and U. M. Titulaer. The influence of temperature gradients on the kinetic boundary layer problem for a condensing droplet. Journal of Statistical Physics, 67(1–2):347–367, April 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049039.

Widder:1993:KBL

[WT93] M. E. Widder and U. M. Titulaer. Kinetic boundary layers in gas mixtures: Systems described by nonlinearly coupled kinetic and hydrodynamic equations and applications to droplet condensation and evaporation. Journal of Statistical Physics, 70(5–6): 1255–1279, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049431.

Wu:1995:NMF

[Wu95] C. Chris Wu. A note on mean-field behavior for self-avoiding walk on branching planes. Journal of Statistical Physics, 81(3– 4):673–680, November 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179252.

Wu:1996:IMH

[Wu96] C. Chris Wu. Ising models on hyperbolic graphs. Journal of Statistical Physics, 85(1–2):251–259, October 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02175564.

Wu:1997:CPP

[Wu97a] C. Chris Wu. Continuity of percolation probability on hyperbolic graphs. Journal of Statistical Physics, 87(3–4):909–913, May 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02181251.

Wu:1997:ICP

[Wu97b] C. Chris Wu. Inhomogeneous contact processes on trees. Journal of Statistical Physics, 88(5–6):1399–1408, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02732442.

Weron:1995:RFD

[WWW95] A. Weron, K. Weron, and W. A. Woyczynski. Relaxation functions in dipolar materials. Journal of Statistical Physics, 78(3–4): 1027–1038, February 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02183699.

Wehr:1997:FSB

[WX97] J. Wehr and J. Xin. Front speed in the Burgers equation with a random flux. Journal of Statistical Physics, 88(3–4):843–871, August 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000015175.70862.77.

Xin:1993:ENT

[Xin93] Jack X. Xin. Existence and nonexistence of traveling waves and reaction-diffusion front propagation in periodic media. Journal of Statistical Physics, 73(5–6):893–926, December 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01052815.

Xu:1995:NCG

[Xu95] Kun Xu. A new class of gas-kinetic relaxation schemes for the compressible Euler equations. Journal of Statistical Physics, 81 (1–2):147–164, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02179973.

Yamada:1996:BAE

[Yam96] Yuji Yamada. Bethe ansatz equations for the broken Z_{N}- symmetric model. Journal of Statistical Physics, 82(1–2):51–86, January 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02189225.

Yang:1994:GME

[Yan94] Dah-Yen Yang. Generalized moment expansion for nonadiabatic transition reaction. Journal of Statistical Physics, 74(3– 4):631–661, February 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02188574.

Yau:1994:MGL

[Yau94] Horng-Tzer Yau. Metastability of Ginzburg–Landau model with a conservation law. Journal of Statistical Physics, 74(3–4): 705–742, February 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02188577.

Yaldram:1991:MCS

[YB91] Khwaja Yaldram and Kurt Binder. Monte Carlo simulation of phase separation and clustering in the ABV model. Journal of Statistical Physics, 62(1–2):161–175, January 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020864.

Yepez:1995:LGC

[Yep95] J. Yepez. Lattice-gas crystallization. Journal of Statistical Physics, 81(1–2):255–294, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179979.

Yamanaka:1996:EGS

[YHHK96] Masanori Yamanaka, Shinsuke Honjo, Yasuhiro Hatsugai, and Mahito Kohmoto. Exact ground-state correlation functions of one-dimensional strongly correlated electron models with resonating-valence-bond ground state. Journal of Statistical Physics, 84(5–6):1133–1208, September 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02174131.

Yukhnovskii:1995:IHM

[YIK95] I. R. Yukhnovskii, I. M. Idzyk, and V. O. Kolomiets. Investigation of a homogeneous many-particle system in the vicinity of the critical point. Journal of Statistical Physics, 80(1–2):405– 443, July 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02178366.

Yuan:1991:CBN

[YK91] Jian-Yang Yuan and T. R. Kirkpatrick. Crossover behavior for a noninteracting disordered electronic system in the presence of a weak magnetic field. Journal of Statistical Physics, 64(1–2):309– 327, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01057879.

Yuste:1996:OSF

[YL96] S. Bravo Yuste and Katja Lindenberg. Order statistics for first passage times in one-dimensional diffusion processes. Journal of Statistical Physics, 85(3–4):501–512, November 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02174217.

Yeung:1993:SDI

[YMHMJ93] Chuck Yeung, J. L. Mozos, A. Hernánez-Machado, and David Jasnow. Surface-driven instability and enhanced relaxation in the dynamics of a nonequilibrium interface. Journal of Statistical Physics, 70(5–6):1149–1174, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01049426.

Yoshida:1997:RCD

[Yos97] Nobuo Yoshida. Relaxed criteria of the Dobrushin–Shlosman mixing condition. Journal of Statistical Physics, 87(1–2): 293–309, April 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02181489.

Yoshida:1998:FVG

[Yos98] Nobuo Yoshida. Finite-volume Glauber dynamics in a small magnetic field. Journal of Statistical Physics, 90(3–4):1015–1035, February 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023249608957.

Yukhnovskii:1995:GCD

[YP95] I. R. Yukhnovskii and O. V. Patsahan. Grand canonical distribution for multicomponent system in the collective variables method. Journal of Statistical Physics, 81(3–4):647–672, November 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179251.

Yeung:1992:PSD

[YRHMJ92] C. Yeung, T. Rogers, A. Hernandez-Machado, and David Jasnow. Phase separation dynamics in driven diffusive systems. Journal of Statistical Physics, 66(3–4):1071–1088, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01055717.

Yuste:1993:RDF

[YS93] S. Bravo Yuste and A. Santos. Radial distribution function for sticky hard-core fluids. Journal of Statistical Physics, 72(3– 4):703–720, August 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01048029.

Yang:1997:ABM

[YS97] Z. R. Yang and Li Song. Application of bond-moving renormalization-group approach to fractal lattices. Journal of Statistical Physics, 88(5–6):1409–1418, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02732443.

Yamaguchi:1990:TFH

[YT90] Yoshihiro Yamaguchi and Kiyotaka Tanikawa. A theorem on the first heteroclinic tangency in two-dimensional maps. Orientation-preserving cases. Journal of Statistical Physics, 59 (5–6):1297–1310, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01334752.

Yamaguchi:1991:NTH

[YT91] Yoshihiro Yamaguchi and Kiyotaka Tanikawa. New type of heteroclinic tangency in two-dimensional maps. Journal of Statistical Physics, 64(3–4):741–754, August 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01048313.

Yang:1992:NDC

[YZ92] Wei-Shih Yang and Yu Zhang. A note on differentiability of the cluster density for independent percolation in high dimensions. Journal of Statistical Physics, 66(3–4):1123–1138, February 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01055720.

Zegarlinski:1990:ESF

[Zeg90] Bogusław Zegarliński. On equivalence of spin and field pictures of lattice systems. Journal of Statistical Physics, 59(5– 6):1511–1530, June 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01334761.

Zegarlinski:1994:SDE

[Zeg94] Bogusław Zegarliński. Strong decay to equilibrium in onedimensional random spin systems. Journal of Statistical Physics, 77(3–4):717–732, November 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02179458.

Zeng:1998:EIH

[ZFB98] Chen Zeng, D. J. J. Farnell, and R. F. Bishop. An efficient implementation of high-order coupled-cluster techniques applied to quantum magnets. Journal of Statistical Physics, 90(1–2): 327–361, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023220222019.

Zhang:1991:NHR

[Zha91] M. Q. Zhang. The nonuniform hard-rod fluid revisited. Journal of Statistical Physics, 63(5–6):1191–1202, June 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01030006.

Zhang:1992:MLS

[Zha92] Qiang Zhang. A multi-length-scale theory of the anomalous mixing-length growth for tracer flow in heterogeneous porous media. Journal of Statistical Physics, 66(1–2):485–501, January 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01060076.

Zhang:1996:REN

[Zha96a] Yingjie Zhang. Rate of escape from nonattracting chaotic sets. Journal of Statistical Physics, 82(5–6):1371–1384, March 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02183387.

Zhang:1996:DBR

[Zha96b] Yu Zhang. Divergence of the bulk resistance at criticality in disordered media. Journal of Statistical Physics, 84(1–2):263– 267, July 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179585.

Zou:1995:IIL

[ZHCD95] Qisu Zou, Shuling Hou, Shiyi Chen, and Gary D. Doolen. A improved incompressible lattice Boltzmann model for timeindependent flows. Journal of Statistical Physics, 81(1–2):35–48, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02179966.

Zou:1995:ASL

[ZHD95] Qisu Zou, Shuling Hou, and Gary D. Doolen. Analytical solutions of the lattice Boltzmann BGK model. Journal of Statistical Physics, 81(1–2):319–334, October 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02179981.

Zhizhina:1998:TPS

[Zhi98] E. A. Zhizhina. Two-particle spectrum of the generator for stochastic model of planar rotators at high temperatures. Journal of Statistical Physics, 91(1–2):343–368, April 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023004507830.

Ziegler:1991:SCF

[Zie91] K. Ziegler. Statistics of colored flux lines. Journal of Statistical Physics, 64(1–2):277–308, July 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01057878.

Ziegler:1993:BCL

[Zie93] Donald P. Ziegler. Boundary conditions for lattice Boltzmann simulations. Journal of Statistical Physics, 71(5–6):1171–1177, June 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01049965.

Ziff:1991:FT

[Zif91] Robert M. Ziff. Flux to a trap. Journal of Statistical Physics, 65(5–6):1217–1233, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01049608.

Zimmer:1993:FNS

[Zim93] Michael F. Zimmer. Fluctuations in nonequilibrium systems and broken supersymmetry. Journal of Statistical Physics, 73(3–4): 751–764, November 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01054348.

Zimmer:1994:EFN

[Zim94] M. F. Zimmer. Errata: Fluctuations in nonequilibrium systems and broken supersymmetry. Journal of Statistical Physics, 75(5–6):1205, June 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02186766.

Zhang:1993:BQM

[ZK93] Shiwei Zhang and M. H. Kalos. Bilinear quantum Monte Carlo: Expectations and energy differences. Journal of Statistical Physics, 70(3–4):515–533, February 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01053583.

Zhdanov:1998:SRT

[ZK98] V. P. Zhdanov and B. Kasemo. Surface restructuring, thermal desorption, kinetic bistability, and chemical waves. Journal of Statistical Physics, 90(1–2):79–101, January 1998. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023243432497.

Zumofen:1991:TRF

[ZKB91a] G. Zumofen, J. Klafter, and A. Blumen. Transient a+b->0 reaction on fractals: stochastic and deterministic aspects. Journal of Statistical Physics, 65(5–6):1015–1023, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01049595.

Zumofen:1991:TAE

[ZKB91b] G. Zumofen, J. Klafter, and A. Blumen. Trapping aspects in enhanced diffusion. Journal of Statistical Physics, 65(5–6): 991–1013, December 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01049594.

Zhang:1993:GMR

[ZM93] M. Q. Zhang and T. G. Marr. Genome mapping by random anchoring: A discrete theoretical analysis. Journal of Statistical Physics, 73(3–4):611–623, November 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01054342.

Zhang:1993:RFS

[ZP93] G. Zhang and J. K. Percus. Response factorization of simply connected Ising lattices with application to Bethe lattice spin glasses. Journal of Statistical Physics, 70(5–6):1365–1377, March 1993. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01049438.

Zebende:1994:DKC

[ZP94] G. F. Zebende and T. J. P. Penna. The Domany–Kinzel cellular automaton phase diagram. Journal of Statistical Physics, 74(5– 6):1273–1279, March 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02188230.

Zobov:1999:EVE

[ZP99] V. E. Zobov and M. A. Popov. Excluded volume effects for frequency moments of the spin autocorrelation function of the Heisenberg model on a square lattice at high temperatures. Journal of Statistical Physics, 97(3–4):793–802, November 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004675527713.

Zaks:1997:CDS

[ZPK97] Michael A. Zaks, Arkady S. Pikovsky, and Jürgen Kurths. On the correlation dimension of the spectral measure for the Thue– Morse sequence. Journal of Statistical Physics, 88(5–6):1387– 1392, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02732440.

Zverev:1991:CON

[ZR91] V. V. Zverev and B. Ya. Rubinstein. Chaotic oscillations and noise transformations in a simple dissipative system with delayed feedback. Journal of Statistical Physics, 63(1–2):221–239, April 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01026600.

Zhang:1990:DRP

[ZSP90] Yi-Cheng Zhang, Maurizio Serva, and Mikhail Polikarpov. Diffusion reproduction processes. Journal of Statistical Physics, 58 (5–6):849–861, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01026554.

Zhi-Xiong:1994:LTD

[ZXZY94] Wen Zhi-Xiong and Wen Zhi-Ying. On the leading term and the degree of the polynomial trace mapping associated with a substitution. Journal of Statistical Physics, 75(3–4):627–641, May 1994. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02186874.

Zylka:1990:IPF

[Zyl90] Ch. Zylka. An inequality for partition functions with disturbed Hamiltonians. Journal of Statistical Physics, 60(1–2):281–284, July 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01013679.

zic:1996:CEN

[zzMZ96] Suncica Elezović-Had zić, Milan Kne zević, Sava Milosević, and Ivan Zivić. Critical exponents for numbers of differently anchored polymer chains on fractal lattices with adsorbing boundaries. Journal of Statistical Physics, 83(5–6):1241–1253, June 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02179562.